A Bibliography of Publications about PVM (Parallel Virtual Machine) and MPI (Message Passing Interface)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

05 March 2025 Version 3.257

Title word cross-reference

+ [BDV03, Cha02, HDB^{+}13, Lee12]. 0 [ICC02]. 1 [ICC02, LRQ01, VDL^{+}15]. $19.95 [Ano95b]. 2 [Bha98, BAS13, CGU12, ES11, KRKS11, KO14, WMRR17, WRMR19]. $24.95 [Ano95c]. $27.50 [Ano96a]. 3 [And98, BCL00, BAS13, CP15, DYN^{+}06, EFR^{+}05, GCN^{+}13, HF14a, HF14b, JR10, KO14, KD13, KHS01, KLR16, MSZG17, NSM12, SSS99, SC19, SH14, TPD15, WR01, YSL^{+}12]. $35 [Ano00a, Ano00b]. $35.00 [Ano99a, Ano99c, Ano99b, Ano99d]. 3D [KA13]. $60 [Ano00a, Ano00b]. ^{3} [PBC^{+}01]. A [ARYT17]. α [JMdVG^{+}17]. Ax = b [BG95]. D [UZC^{+}12]. H^{2}/H^{∞} [GWC95]. hp [BCM^{+}16]. k [She95, TK16]. ↔ [GRW^{+}19]. M^{3} [JSH^{+}05]. PVM^{+} [Wil94]. N [IHM05, Per99, Rol08b, SP99, SRK^{+}12]. P_{N} [OGM^{+}19]. P_{N−2} [OGM^{+}19]. SU(3) [BW12]. t [MPZ21]. τ [RGDM15, RGDML16]. XY [KO14].

* [MMAH20].

-based [Rót19]. -body

[IHM05, Per99, SP99, SRK^{+}12]. -D [DYN^{+}06, SSS99, SH14, Bha98, ES11, KHS01, NSM12]. -Dimensional [LRQ01]. -Lop [RGDM15, RGDML16]. -Means [TK16]. -Queens [Rol08b]. -set [She95]. -SNE [MPZ21]. -stable [JMdVG^{+}17].

. [Wil94].

/Fortran [TBG^{+}02]. /many [KSG13]. /MPI [BKK20]. /OpenMP [VDL^{+}15].

’00 [RV00].

1 [HMKV94, SOHL^{+}98]. 1/Pascal [GDS^{+}20]. 10 [LLVM21a]. 10-Gigabit [HcF05]. 100 [Str94]. 1007 [AEW^{+}20]. 100k [SC19]. 1012 [CWL^{+}20]. 1016 [HFB21]. 10th [DLO03, IEE96e]. ’11 [ACM11, QM21]. 11th [IEE97b, KKD04]. ’12 [Hol12]. 128-processor [LL01]. 12th [DKD05, Bil95]. 13th [Ano95d, MTWD06, PSB^{+}94]. 14th [CHD07, RV00, CHD09]. 15-18 [SL94a]. 15th [IEE95i, LKD08]. 16th [RWD09]. 17th [KGRD10, MC94]. 18-21 [DKD07]. 18th [CDND11]. 1990 [ACM90]. 1991 [DE91, EJL92, IEE91]. 1992 [KG93, R^{+}92, VW92]. 1993 [Ano94c, GGK^{+}93, IEE93a, IEE93e, JPTE94, MMH93]. 1994 [Ano94a, Ano94e, DSZ94, DT94, GN95, GT94, HK95, IEE94h, PSB^{+}94, SPE95, SPH95, VV95]. 1995 [ACM95a, ACM96a, AGH^{+}95, BH95, Gat95, Ham95a, IEE95b, IEE95a, IEE95d, IEE95h, IEE95i, JB96, NM95, Nar95, Ten95, UCW95, ZL96]. 1996 [ACM96b, Abr96, Boi97, ERS96, IEE96f, IEE96e, IEE96i, Ree96]. 1998 [ACM98b]. 1999 [ACM99]. 19th [TBD12, IEE05]. 1st [Abr96, BR95a, CGB^{+}10, Kum94, Van95, Fer92].

2 [AKL99, BCAD06, BHS^{+}02, BMPZ94a, CwCW^{+}11, CD96, DPSD08, FST98a, FST98b, GFD03, GGHL^{+}96, GT01, GHLL^{+}98, GLT99, GLT00b, GLT00a, HGMW12, Jon96, LC97b, LSK04, MS02a, MK04, PS00a, SS99, SSL97, TRH00, VAT95, bT01a]. 2-D [BMPZ94a]. 2.0 [Bae20, BO01, LPD^{+}11, LW97, Mat00b, NSM12, RS22, WYH^{+}21]. 2.2 [HRR^{+}11]. 2.X [KS96]. 2000 [ACM00, CLBS17, LL01, LSK04, NU05, RV00, ZSnH01]. 2001 [ACM01, Old02]. 2003 [ACM03, AS14, Don06, OL05]. 2004 [ACM04]. 2005 [ACM05, DKD07]. 2006 [ACM06a, MTW07]. 2007 [SM07]. 2008 [SMCH15]. 2010 [CGB^{+}10]. 2011 [LCK11]. 2012 [Hol12, TB14]. 2015 [IS16]. 2017 [GT19]. 2019 [TH20]. 202X [Taf21]. 21st [IEE95a]. 25nm [Ano03]. 26th [Ano93a, SL94a]. 27th [Ano94h]. 28th [ZL96]. 2D [TPV20, ZZZ^{+}15]. 2D-DWT [ZZZ^{+}15]. 2nd [FK95, IEE93c, Nag05, YM97].

3 [Bri95, Che10, FCS^{+}19, GBH14, GBH18, GPL^{+}96, GLT12, Gro12, HDT^{+}15, QM21]. 3-D [Bri95]. 3.0 [Ano97, Bra97, BMR02, BRM03, DBB^{+}16, KaM10, OP10]. 3.06 [Ano03]. 3.1 [WCC12]. 3.4 [Gei97, GKPS97]. 3.X [KS96]. 3000 [HWM02]. 33rd [ACM95a]. 37th [ACM06a]. 3D [GAP97, Gra97, KB21, LO96]. 3D-Fall [Gra97]. 3rd [ACM06b, CZG^{+}08, Ano95a, IEE96a].

4 [Ano03, HRZ97, KSHS01, NU05, SD13, SBT04]. 4.0 [BPJ22, DSGS17, DWS^{+}21, JCP15, dOSMM^{+}16]. 4.5 [CBYG18, DFP^{+}19, TMT^{+}20]. 43 [UZC^{+}12]. 45-degree [CT13]. 48th [IEE94e]. 4th [BDW97, EdS08, FF95, USE00].

5 [TRH00]. 512 [RBB97c]. 5th [AD98, Cha05, IEE94a, MdSC09].

600 [LSK04]. 6000 [AL93, NMW93]. 64 [dCZG06]. 64-bit [Wil93]. 6th [ACDR94, DLM99, GT94, PW95, SHM^{+}10, Sin93].

7th [ACM95b, CGKM11, DKP00, GN95, PBG^{+}95].

857 [SMSW06]. 897 [HWS09]. 8th

[CMMR12, CD01].

90 [Ben95, SM03]. 9076 [Bri95]. ’91

[BG91, EJL92, IEE91]. ’92 [Sie92a, Sie92b, VW92]. ’93 [Ano93g, GGK^{+}93, GHH^{+}93, IEE93a, IEE93e]. 93SC038 [FS93]. 93SC041 [Gle93]. ’94 [BS94, DW94, GT94, IEE94b, IEE94h, PSB^{+}94, SPE95, WPH94, dGJM94]. 947 [LTDD14]. ’95 [ACM95b, AH95, BH95, CLM^{+}95, CJNW95, DMW96, FF95, HAM95b, IEE95l, Lev95, NM95, Van95, Ano98, FD97, KaM10]. 95/NT [FD97]. ’96 [ACM96b, ACM96c, BDLS96, BFMR96, CH96, IEE96g, IEE96e, IEE96d, LHHM96, Li96, Sil96, Was96, YH96]. ’97 [ACM97a]. 978 [Che10, SD13]. 978-0-12-415933-4 [SD13]. 978-0-13-138768-3 [Che10]. 981 [Riz17]. 997 [Spe19]. 9th [IEE95f, Kra02, YH96].

Aachen [Ano93a, GHH^{+}93]. Abortable [CAWL17]. Abortable-locking [CAWL17]. Abstract [MKW11, Wel94, BG94b, HTA08]. Abstraction [DSU20, SW12, YWTC15]. Abstractions [RHM^{+}17]. Abstracts [IS16]. ACC [APJ^{+}16]. accelerate [SdM10, TBB12, VGP^{+}19]. Accelerated [AB13, EADT19, KF16, KA13, LHZ^{+}20, NRdA^{+}20, SCSL12, VZT^{+}19, BBB^{+}22, BMS19, CGK^{+}16, CP15, DCD^{+}14, HTJ^{+}16, JCP^{+}20, KM10, PGdCJ^{+}18, PTMF18, Sai10, iSYS12, SKM15, ZWL^{+}17, ARYT17]. Accelerating [BBC^{+}19, Dab19, GM18, HF14a, HF14b, HKOO11, JK10, JLS^{+}14, JNL^{+}15, LSSZ15, LSVMW08, LSMW11, LAFA15, PSV19, RJ21, SCJH19, TMP16, TS12b, UZC^{+}12, YEG^{+}13, vdLJR11, HWX^{+}13]. Acceleration [CGBS^{+}15, GDEBC20, RVKP19, STH22, TK16, WTS19, CBYG18, CLBS17, CBS18, HE13, JPL22, MGS^{+}15, MPS20, OGM^{+}19, PRS16, RVKP18, SWS^{+}12]. Accelerator [APJ^{+}16, CLA^{+}19, SSAS12, SXMX^{+}18, YCA18, KLV15, MKP22, WHMO19]. Accelerator-Aware [APJ^{+}16].

Accelerator-bound [CLA^{+}19]. Accelerators [AKL16, AC17, NTR16, SHM^{+}10, TCM18, TL19, KHBS19, MSZG17, UGT09, vdP17]. Access [Bri10, HDT^{+}15, IFA^{+}16, JJPL17, LB98, SGH12, WTR03, CLA^{+}19, CG99b, GBH14, GBH18, HGMW12, LOHA01, MN91, SFL^{+}94]. Accesses [CVPS19, TGL02]. accessible [BHW^{+}12]. Accident [Smi93a, SBR95]. According [LGM00]. ACCT [FVD00]. Accumulated [KS15b]. Accumulative [IH04]. accuracy [SSH^{+}19]. accuracy-aware [SSH^{+}19]. Accurate [HD00b, MLA^{+}14, RSPM98, HD00a, LZC^{+}20]. Accurately [BGdS09]. achievable [HMS^{+}19]. Achieving [CBPP02, Gro01a, HLK^{+}20, KKLL11, RH01]. ACM [ACM90, ACM95a, ACM95b, ACM97b, ACM98b, ACM04, ACM05, IEE02]. ACM/IEEE [ACM97b, ACM98b, ACM05]. ACO [Tsu12]. acoustic [ZWC21]. ACPC [Bos96, Vol93]. Across [NE98, AL96, CZ95b, KW20]. ACSCI [Van95]. action [Hol95]. Active [CSAGR98, Pla02, SKH96]. Activities [MSS97, CMV^{+}94]. activity [Vet02]. Ad [IBC^{+}10, ITT02]. Ad-Hoc [IBC^{+}10]. Ada [Tou96, KP96, PRQ21, Taf21, Tou96]. Adam [Ano95b, NMC95]. Adaptable [SPH^{+}18, BCM^{+}16]. Adaptation [WST95, RS21]. Adapted [Uhl95a]. Adapting [VFD02]. Adaptive [Ano94b, BCMR00, BKdSH01, Bir94, CKO^{+}94, FLS20, FSC^{+}11, HWX^{+}13, KK98, KT02, LFL11, LYGG20, MKC^{+}12, MBES94, MRB17, MAGR01, OKW95, Ran05, RA09, SHM^{+}12, SGZ00, SS09, STY99, Sta95a, TMW17, ZSG12, ADK22, BDP^{+}10, CLSP07, DLR94, EZBA16, EASS95, IDS16, KTXP21, LCL^{+}12, SWCB20, SLGZ99, TCBV10, Was95a, WK20, Wil94, FSC^{+}11]. Adaptive-CoMPI [FSC^{+}11]. Adaptive-Length [FLS20]. Adas [HHC^{+}18]. Adding [CB00, GRV01, PSM^{+}14]. Address [SS01, DO96]. addresses [CGL^{+}93]. ADDT [SR96]. ADI [Sch01]. adjacent [Kan12]. adjoint [RMNM^{+}12]. Adjusting [GSHL02]. Adjustment [DSCL05]. ADOL [BGK08]. ADOL-C [BGK08]. adoption [CMV^{+}94]. Adsmith [LKL96]. Advanced [Ano98, Ano00a, D^{+}95, Gei96, Gei97, GLT99, GLT00b, GLT00a, GLT12, KG93, SSAS12, TG94, Ben95, DMK19]. Advances [Bha93, BBH^{+}08, CHD07, CDND11, KGRD10, KKDV03, KKD04, KKD05, LKD08, LK10, MTWD06, RWD09, TBD12, AD98, BC14, BDW97, CD01, DKD05, DLM99, DKP00, DLO03, HPS^{+}12, Kra02, HPS^{+}13, IEE97a]. Advection [AKK^{+}94, CT94a, TC94, CT94b]. Advection-Chemistry [AKK^{+}94]. Advisor [GVF^{+}18]. Aerospace [MAB05]. AES [HMKG19]. Affine [DMB16]. Affinity [ETWaM12, AGG^{+}95, NAAL01, vdP17]. Affordable [Rol94]. After [DF21]. again [Har94]. against [GHD12]. Age [MdSC09, Ano94f, GJLT11, HK95]. AGEB [SAS01]. Agent [Mat01b, MCB05, ZWZ^{+}95]. agent-based [MCB05]. agents [KBA02]. Aggregation [KLH^{+}20]. Aging [LRBG15]. Aging-Aware [LRBG15]. AIMS [Yan94]. Air [AKK^{+}94, BZ97, MPD04, MSML10, BTC^{+}17, SH94, Syd94]. airspace [TCP15]. Aix [GA96, Ano01a]. Aix-les-Bains [GA96]. Al [Ano95b, NMC95]. Alamos [Old02]. Albuquerque [IEE91, IEE95d]. Alchemist [GRW^{+}19]. ALDY [GS96]. ALE [HAA^{+}11]. Algebra [BDT08, CDD^{+}13, Coo95b, DGH^{+}19, IS16, MGMH97, Neu94, van97, BKvH^{+}14, Cal94, Coo95a, LRLG19, PMZM16, VLCM^{+}20, dCH93]. Algebraic [CGPR98, Lev95]. Algorithm [AEW^{+}20, AiIS^{+}21, ACMR14, BST^{+}13, BP99, BT01b, DYN^{+}06, FJBB^{+}00, HA10, HD02b, ITT02, MW98, MPZ21, Per21, PKD95, PB12, RDMB99, Rót19, SAS01, Sch96a, SSLMW10, SWH15, Sta95b, TK16, WHDB05, ZJHS20, ZWLZ21, ART17, AAAA16, ARL^{+}94, AD95, BBC^{+}19, BB95a, BAV08, BY12, BCM^{+}16, CCU95, CT13, CSW99, GM94, GCN^{+}13, GGL^{+}08, GKK09, GP95, HWS09, IM95, JPL22, JR13, KDSO12, KY10, KWEF18, Kan12, KBP16, KN17, KO14, Kom15, KRC17, LYIP19, LYZ13, MM92, MLVS16, MK00, NB96, NAJ99, OKW95, OGM^{+}19, OMK09, PGBF^{+}07, PSLT99, Ram07, RJC95, RAGJ95, Sch96b, SOA11, SOYHDD19, Sur95a, TNIB17, TSCS14, TGKL19, Was95a, YULMTS^{+}17, ZSK15, ZWL^{+}17, dH94, van93, AEW^{+}20, CWL^{+}20, HFB21, HWS09, LTDD14, Riz17, Spe19, SMSW06]. Algorithm-based [PKD95]. Algorithm-Dependant [BP99]. Algorithmic [Stp20, HHSM19, RJDH14]. Algorithms [ACM95b, ATC94, ADRCT98, ABG20, ASA97, CDT05, CCSM97, DK20, DALD18, DAK98, DK06, FB94, GAMR00, GK10, HO14, HHK94, IEE96d, KTAB^{+}19, KK02a, LHHM96, Li96, LAD16, lLLmH^{+}21, MTSS94, MGMH97, MBS15, Nar95, Pet97, PBK00, Pro21, SG15, SGS^{+}21, VRS00, AK99, AL92, BHJ96, BMS^{+}17, BID95, DDLM95, FR95, FP92, GWC95, HL17, HPLT99, HKOO11, HS95b, JKN22, Jou94, JRM^{+}94, KL95, KRG13, LFL11, LNW^{+}12, LRLG19, MTK16, MJG^{+}12, NP12, Ols95, PP16, Pan95b, PBK99, PD11, PCS94, RHG^{+}96, SPE95, Sur95b, TSZC94, WCVR96, YLZ13]. alias [SOA11]. alias-free [SOA11]. aligned [AGIS94]. Aligners [SMM^{+}16]. Alignment [dOSMM^{+}16, AMHC11]. All-gather [Pro21]. all-port [RJMC93]. All-to-All [LZH17, LZH18, Trä02b]. Allgather [KTAB^{+}19]. Allgatherv [KTAB^{+}19]. Allocation [AGS97, BS01, DGG^{+}12, RFRH96, SPNB14]. alloy [TG94]. AllReduce [NWT21]. ALM [PZ12]. almost [LFW20]. alpha [WLYL20]. Altera [RGB^{+}18, TK16]. Alternative [EM94, SWHP05, Trä12a, EKTB99]. ALWAN [HB96a, HB96b, MSB97]. Amazon [ZLZ^{+}11]. AMBER [SL95]. AMBER4 [VM95]. American [Ara95]. AMIP [Gat95]. Among [CB16]. AMPI [ZHK06]. AMPIC [CCHW03]. amplified [EZBA16]. AMR [NLRH07, TK19]. AMRVAC [KTXP21, TK19]. AN2 [HBT95]. analogue [WWZ^{+}96]. analyses [ANS95]. Analysis [BHW^{+}17, BR02, BGG^{+}02, BBC^{+}00, BDL98, CGLD01, CLA^{+}19, DFP^{+}19, EML00, FK01, FJK^{+}17, Hol12, JF95, KL94, KNT02, KRG13, LCK11, MK17, MCLD01, NAW^{+}96, NMS^{+}14, Ost94, PZ12, PGAB^{+}05, SPL^{+}12, SBR95, SGL^{+}20, SN01, TFGM02, WYH^{+}21, Whi04, WM01, BB93, BBDH14, BBH^{+}15, Che99, DSGS17, EPP^{+}17, GR95, GFB^{+}14, GSM^{+}00, GKS^{+}11, GE95, GE96, GT07, JB96, JLG05, LC07, LLG12, LRLG19, LL16, LBH12, MMB^{+}94, MMW96, MLA^{+}14, MJPB16, Pat93, PHJM11, PSV19, PGAB^{+}07, RTN21, SdSCP13, iSYS12, SS94, SDJ17, SPH95, Shi94, Sil96, SWL^{+}01, SSG95, TMC09, TW12, TFZZ12, Uhl95a, Uhl95c, VM94, YCL14]. analytic [THDS19]. analytical [BAE22, BHW^{+}12, HK09, JS13, KN17]. analytics [MMAH20]. Analyzer [JJPL17, KKM15]. Analyzers [Ano01a]. Analyzing [BRU05, DF17, FM09, HG12, HcF05, PFG97, RPS19, MH21]. anasslich [Ano94c]. Anatomy [KWEF18]. Andrew [Ano99c, Ano99d]. animal [LM99]. anisotropic [LBB^{+}16, SSB^{+}16, YSVM^{+}16]. ’Annai [CEF^{+}95]. Annapolis [IEE96c]. annealing [WHMO19, FH97]. Annecy [VW92]. Anniversary [Ano92, Ano93f]. annotated [GGH99]. Annotation [MGA^{+}17]. announcement [WRMR19]. Announcements [Ano98]. Annual [ACM95b, Ano93b, Ano94h, IEE95b, USE00, Van95, Y^{+}93, ACM95a, Eng00, IEE94e, IEE95l]. Ant [ITT02]. ante [Ano03]. antenna [DSOF11]. Anthony [Ano95c, Ano00b]. Antonio [Ano95d, IEE95g, IEE97c]. Any [Gro02a, Mar07]. AP [PBC^{+}01, SMTW96]. AP/Linux [SMTW96]. AP1000 [SH96, IM94, SWJ95]. AP3000 [TD99]. Apache [GRW^{+}19]. API [DM98, KQT^{+}21, LPD^{+}11]. APIs [WCS^{+}13]. APOLLO [Sta95b]. APOLLO-II [Sta95b]. Appendix [Ano01a]. Appendixes [Ano01a]. APPL [AB93b, AB93a]. Application [AKE00, BSN95, BGdS09, BS07, BFM97, BBH^{+}15, Cha02, CRGM14, DFMD94, FDG97a, FDG97b, FSC^{+}11, GB98, HT08, IADB19, JFY00, JCH^{+}08, KNT02, LD01, LMRG14, Mal01, MTSS94, MBB^{+}12, NSLV16, NS16, PSSS01, Riz17, SBF^{+}04, ST02a, SCL97, UTY02, WYZ^{+}19, YNJS21, ZZ04, ABC^{+}00, ADMV05, ADR^{+}05, BBB^{+}22, BvdB94, BFLL99, BL97, BBC^{+}99, BMPS03, CBYG18, CRM14, CRGM16, DS22, EPML99, FMFM15, GVF^{+}18, GWVP^{+}14, HTJ^{+}16, HZ96, KME09, LSG12, LFS^{+}19, LCMG17, LBB^{+}19, MMW96, MM03, MLA^{+}14, MvWL^{+}10, NMW93, RBAI17, Rol08b, SM12, SCJH19, SSS99, SFSV13, SL00, TCP15, Wor96, ZZZ^{+}15, CG99a, PGPCK21]. application-centric [SFSV13]. Application-Level [CRGM14, LMRG14, SBF^{+}04, SCL97, BMPS03, CRM14, CRGM16, LCMG17, LBB^{+}19]. Applications [APJ^{+}16, AGS97, Ano89, Ano96c, AZG17, BCLN97, Ben18, BHV12, BBH^{+}06, BRU05, BFMT96b, BFBW01, CGS15, CBL10, CGLD01, CBB^{+}20, CBB^{+}21, Cha05, CTBT21, CJNW95, CRGM14, Cot98, CTK00, Cot04, Cza02, Cza03, DW02, DLM^{+}17, DERC01, DHK97, DGF97, DDN^{+}22, DGMJ93, EV01, EML00, FLD98, FD00, FGRD01, Fer92, FK95, Fin00, FC05, FM09, GKP97, GK10, HMK09, HDW21, Hus98, IEE95l, ITT02, Jes93b, JJPL17, KB98, KBS04, KGK^{+}03, KSB^{+}20, KKP01, KK02b, Kuh98, Laf01, LAdS^{+}15, LWSB19, LRG14, kLCCW07, LBB^{+}21, LdSB19, LMRG14, dLR04, MSOGR01, MS02a, Mar02, Mat01b, MAB05, MC98, MG15, MANR09, NFK98, PSM^{+}14, Rei01, RPM^{+}08, RBB15, RRBL01, SPL^{+}12, SdR^{+}21, SG12, SPH^{+}18, SC04, SPB^{+}17, SSB^{+}17, TTSY00, TFGM02, VdS00, VY02, Vos03, Wal96a, WC09, WZM17, WJA^{+}19]. Applications [Wis96a, WSN99, WBH97, WM01, dGJM94, AC07, ACH^{+}11, ACC^{+}21, ACJ12, Ano93a, Ano94f, Ano03, ABB20, Ara95, Arn95, ASB18, AGMJ06, BAE22, BKH^{+}13, BR04, BDV03, BAG17, BFM96, BFMT96a, CGK^{+}16, CGBS^{+}15, CDMS15, CLE^{+}20, CLSP07, CBM^{+}08, CZP21, CIJ^{+}10, CFPS95, CCHW03, CCM^{+}06, DZ98a, DSZ94, DPFT19, D^{+}95, DCH02, EKTB99, EGH99, EDSV09, FE17a, FE17b, FNSW99, FCS^{+}12, Fin94, Fin95, FF95, GBR15, GS02, GHD12, GJMM18, GS96, GSM^{+}00, GHH^{+}93, HD00a, HZ99, HAJK01, JC17, JPTE94, KC19, KSC^{+}19, LRG^{+}16, LBG^{+}20, LZZ^{+}20, LMG17, LCMG17, LBB^{+}19, LGM^{+}20, LZHY19, LS08, MA09, MBKM12, MLC04, MSMC15, MS96b, NSBR07, NCB^{+}12, NFG^{+}10, PK05, PTL^{+}16, Rab99, RS95, RGGP^{+}18, RGP22, SJLM14, SPE95, SBG^{+}12, SDJ17, SGH12, SG05, SPBR20, SIC^{+}19, SLG95, SB01]. applications [SD16, SRS^{+}19, TMC09, TBB12, TPLY18, Vet02, Wis96b, Wol92, WT13, WMP14, XLW^{+}09, XJR21, YZ14, ZLZ^{+}11, BP93, TDBEE11, ATC94]. Applied [FGRD01, HC06, KaM10, GFIS^{+}18, HMKV94, MM92, MPS20, NF94, PGK^{+}10, DMW96, Was96]. Applying [GSM^{+}00]. Approach [AZG17, BHM94, BJ93, BHNW01, CRGM14, CD98, DLM^{+}17, FFP03, GCBL12, HMKG19, HD00b, KBA02, KK02a, KmWH10, LGM00, Mar06, PPR01, Pet00a, Pet00b, RGD13, Ros13, SdR^{+}21, TJPF12, BK11, Bis04, BTC^{+}17, CLYC16, CDP99, CRGM16, DiN96, EO15, FMS15, HDB^{+}13, JS13, KPL^{+}12, KSSS07, KJEM12, LSG12, MGG05, MS99b, NEM17, OHG19, OW92, QM21, SVC^{+}11, SEC15, TWFO09, VGP^{+}19, WO09, YW21]. Approaches [JCH^{+}08, Ney00, SWHP05, SM02, AKB^{+}19, BFLL99, CB11, PS00b]. ApproxHPVM [SSH^{+}19]. Approximate [FLS20, Huc96, MYL21, MM02, GGC^{+}07, GG09, MM03]. approximately [LFW20]. Approximation [SLJ^{+}14, SJLM14]. April [ANS95, AH95, Ano93h, Ano94h, CH96, DR94, GH94, Ham95a, IEE92, IEE93b, IEE95f, IEE96e, IEE97b, IEE05, LCHS96, MC94, Nar95, Sie94, SW91, Ten95]. APS [GT94]. AQsort [LTS16]. AQUAgpusph [CP15]. arbitrary [HP11]. ARCH [Ada97, Ada98]. architectural [GGC^{+}07]. Architecture [BG94a, CGC^{+}11, CLOL18, EBKG01, EM02, FDG19, FD97, Fuj08, HRZ97, IEE97c, ITKT00, LSZL02, PT01, PS01b, SMM^{+}16, SC04, SYL19, WKP11, YTH^{+}12, ZGNZ22, BBCR99, BG94c, CSPM^{+}96, CS96, CBIGL19, DiN96, FHC^{+}95, HIIG16, HK09, MMDA19, MRH^{+}96, PWD^{+}12, SWYC94, SBK21, SSGF00, Squ03, SP11, WCC^{+}07, YÁJG^{+}15, YEG^{+}13, ZWZ^{+}95]. architecture-independent [DiN96]. Architectures [ACM95b, BDT08, BBD^{+}20, BFG^{+}10, CHPP01, HD02a, HD02b, HHK94, IEE96d, KDT^{+}12, LHHM96, Li96, LZH17, LAD16, MS02b, MTSS94, MPZ21, MCS00, NO02b, Nar95, PZ12, SXMX^{+}18, TSCaM12, WYZ^{+}19, YKW^{+}18, ZTD19, BDP^{+}10, BN00, BKML95, CLM^{+}95, CDZ^{+}98, DM93, DZZY94, GDC15, GP95, HHS18, HCC^{+}20, Hos12, LCL^{+}12, LDJK13, MLC04, NO02a, PY95, RFH^{+}95, RMNM^{+}12, SPL99, TDG13, TSZC94, Uhl95a, VDL^{+}15, WST95, dlAMC11]. Area [CDHL95, Fis01, BHW^{+}12, FGT96, FGG^{+}98, KHB^{+}99, Qu95]. area-based [Qu95]. arising [ARvW03].

Aristotle [FSV14]. Arithmetic [Ano98, JPT14, Sur95a]. Arithmetics [HD00b, HD00a]. Arizona [IEE95b, JB96]. ARM [AFGR18, MGL^{+}17]. ARM-based [AFGR18]. Array [DDPR97, HD02b, LTS16, MYK19, WG17, CCM12, DK13, HSE^{+}17, JKN^{+}13, Ott93, TOC18, Wal02]. arrays [HCL05, RBS94]. Arria [LLVM21b, LLVM21a]. Arrival [FPY08, Pro21, MLVS16]. art [LF93b]. artifact [ZZZ^{+}15]. Artificial [BPG94]. ARTUR [FJBB^{+}00]. ARVO [BHW^{+}12]. ARVO-CL [BHW^{+}12]. ary [Pan95a]. Ascona [DR94]. Ashes [Thr99]. ASL [FGRT00]. ASME [LF^{+}93a]. aspects [CG99a]. Assembly [PGF18, TPD15]. Assessing [LMG17, dLR04, MABG96, TSCaM12, CMV^{+}94]. Assessment [Mat01b, TAH^{+}01, Boi97, LH98, LSB^{+}20]. Assignment [Cza13, CK99]. assist [Kik93]. Assisted [GTH96, ADK22, GM13, MBBD13]. Astro [CC17]. Astronomical [JB96, SPH95]. asymmetric [GCN^{+}10]. asynchronization [FSG19a, FSG19b]. Asynchronous [Ada97, Cav93, CZ95a, CDP99, HE02, SPH^{+}18, BBDH14, BCK^{+}09, CZ95b, DDYM99, PHM^{+}22, RSC^{+}19, Sch99]. Athapascan [CP98]. Atlanta [AGH^{+}95, Ara95, USE00, UCW95]. ATM [GFV99, HBT95, Jon96, LHD^{+}94, LHD^{+}95]. Atmosphere [BS93]. Atmospheric [HK93, KHBS19, RSBT95]. atom [MGG05, SPBR20]. atom-based [SPBR20]. Atomic [LRT07, LAFA15, SYF96, DS13, Hin11, SY95, XF95]. atomics [BDW16]. atoms [JLS^{+}14]. Attacks [PV97, GHD12]. attempt [GM18]. Attraction [GB96]. attribution [GADM20]. audio [BJ13]. Augmented [GFJT19]. Augmenting [TL19]. August [ATC94, Agr95a, BFMR96, DMW96, GT94, HAM95b, IEE94g, IEE95k, IEE95l, IEE96f, LF^{+}93a, Ost94, PSB^{+}94, PBG^{+}95, Ree96, VV95, Was96]. Aurora [TSEE21, LdSB19]. Austin [IEE94b]. Australasian [Bil95]. Australia [GN95, Nar95, ACDR94, Bil95]. Australian [ACDR94, GN95]. Austria [Bos96, BH95, Kra02, TBD12, Vol93]. Austrian [Fer92, FK95]. Austrian-Hungarian [Fer92, FK95]. Auto [CC17, DWM12, DBLG11, PSB^{+}19, RDLQ12, WG17, FE17a, FE17b, SH14, TWFO09, VLCM^{+}20]. Auto-Generation [CC17, DWM12]. auto-parallelization [TWFO09]. Auto-scoping [RDLQ12]. Auto-tuned [PSB^{+}19, VLCM^{+}20]. Auto-Tuning [WG17, DBLG11, FE17a, FE17b, SH14]. AutoLink [GMPD98]. AutoMap [GMPD98]. Automata [Car07, BBK^{+}94, SC19]. Automated [BMPS03, MVY95, RVKP18, LLG12, LCH^{+}22, RFRH96, Yan94]. Automatic [AAB^{+}16, BVML12, BBH^{+}08, BGK08, BHK^{+}06, CBL10, Cza03, DW02, EML98, EML00, FAFD15, FFM11, GKCF13, HZ99, HH22, JFY00, JJY^{+}03, JJPL17, KOI01, KHS12, MB18, MGA^{+}17, NCB^{+}17, OWSA95, Rab99, RGD13, SZ11, SR96, SSB^{+}17, TJPF12, WC15, WZW21, WM01, APBcF16, AMuHK15, ABB20, AGG^{+}95, BR04, BHRS08, CHKK15, CdGM96, CPR^{+}95, HZ96, LME09, LF93b, PHM^{+}22, VLCM^{+}20, WMP14, ZHK06, FVD00]. Automatically [VZT^{+}19, WBSC17]. Automating [RHM^{+}17]. automation [Ano93a]. automotive [Ano93a, Ano93a]. autoregressive [CBS18]. Autotuning [BAG17, PSH^{+}20]. Auxiliary [STMK97]. Available [Bak98, BF98]. Averaging [RJ21]. Avoidance [CRGM14]. avoiding [GKD^{+}18]. AVTP [FHC^{+}95]. award [Str94]. Awards [Str94]. Aware [APJ^{+}16, AAB^{+}16, BHP^{+}03, Ben18, DS22, EGR15, GFIS^{+}18, HVA^{+}16, LRBG15, MJB15, Pan14, ZLP17, BLVB18, CLA^{+}19, CGH^{+}14, FA18, GMA20, GHZ12, HJYC10, HG12, JKN^{+}13, KBG16, MBBD13, MSMC15, MMAH20, SHM^{+}12, SSH^{+}19, SPK^{+}12, WRSY16]. awareness [HK09, VGS14]. AXAF [NH95]. AXC [CBIGL19]. azTotMD [RS22].

B [Ano01a]. Back [BIC^{+}10]. Backend [IOK00]. backtracking [PGdCJ^{+}18]. Backup [Gua16]. Bains [GA96]. Balance [HE02]. balanced [EZBA16]. Balancing [BKdSH01, DBA97, DI02, DK06, FSG19a, GCBL12, KSB^{+}20, MM02, PT01, Pus95, ST97, Wal01a, Bir94, BS05, DZ96, DLR94, DvdLVS94, DR95, FMBM96, FH97, Hum95, JH97, MM03, NP94, SGS95, SY95]. Balatonfured [DKP00]. balls [BBH^{+}15]. Baltimore [IEE02, SPH95]. Bamboo [NCB^{+}12]. banded [DG95]. Bandwidth [NE01, RK01]. Bangalore [Kum94, PBPT95]. Barbara [ACM95b, AH95, IEE95f]. Barcelona [DLM99]. Barnes [MPZ21]. BARRACUDA [EPP^{+}17]. Barrier [CLdJ^{+}15, SDB^{+}16, YLZ13]. Based [Ada97, AHD12, AAB^{+}17, ABG20, AP96, BHW^{+}17, BDG^{+}91b, BBD^{+}20, BoFBW00, CAM12, CGC^{+}02, CLOL18, CLP^{+}99, CDPM03, DW02, DLLZ19, DLLZ20, DBK^{+}09, FSC^{+}11, FC05, For95, FSLS98, GSxx, GFJT19, HF14a, HF14b, HM01, Hus00, KLR16, LSZL02, LZH18, kL11, LWP04, LAFA15, MDM17, MGL^{+}17, MMH98, MZLS20, NSLV16, NE01, NHT02, NPS12, PPT96a, PCY14, PFG97, PSSS01, RDMB99, SPL^{+}12, SM03, Smi93a, ST02b, ST97, SJK^{+}17a, SJK^{+}17b, THS^{+}15, TD98, WTTH17, WC09, WZHZ16, WYH^{+}21, WJG^{+}21, Wis96a, WM01, WJB14, YG96, YTH^{+}12, ZJHS20, ZWJK05, AKB^{+}19, Ada98, AASB08, AAAA16, AVA^{+}16, Ano03, AFG21, ABB20, AFGR18, BLPP13, BDG^{+}92a, BLVB18, BCH^{+}03, Bri95, BFMT96a, CwCW^{+}11, CC10, CPM^{+}18, CKmWH16, CRM14, CXB^{+}12, DXB96, FE17a, FE17b, FFB99, FJZ^{+}14, FNSW99, FSTG99, FLPG18, FFFC99, FWS^{+}17]. based [GS91a, GS92, GKS^{+}11, Gra97, Gra09, GFPG12, HDZ^{+}20, HZ94, HWX^{+}13, IM95, ITT99, JCP^{+}20, JL18, JKM^{+}17, KLV15, KB21, KPL^{+}12, KSC^{+}19, KPNM16, LV12, LRW01, LKL96, LNW^{+}12, LZC^{+}20, LGG16, LMM^{+}15, MYB16, MMO^{+}16, MB21, MKP^{+}96, MCB05, MT96, MS99a, MS99b, MMAH20, MFPP03, NRdA^{+}20, Neu94, NHT06, OLG^{+}16, OP98, PARB14, PES99, PPT96b, PK05, PS19a, PAdS^{+}17, PGK^{+}10, PSHL11, PKD95, PSK^{+}10, PSLT99, Qu95, Rag96, RBP^{+}21, RJH^{+}20, Rót19, STP^{+}19, SJLM14, SSN^{+}21, SS09, SG05, SSS99, SZ11, SPBR20, SVC^{+}11, SXMX^{+}18, SLS96, SKB^{+}14, Sto98, Stp18, Stp20, Str96, SLN^{+}12, SPNB14, TBB12, TSCS14, TGKL19, TY14, TBD96, TWFO09, TMPJ01, VLCM^{+}20, WHMO19, WO09, WTFO14, WTS19, WGG^{+}19, Wis96b, WCS99, YC98, YL09, YWC11, YSL^{+}12, ZAFAM16, ZLP17, ZHK06, ZZG^{+}14]. based [ZWZ^{+}95, ZWC21, vHKS94, BFMT96b, FH97, KSJ95, WAS95b, FO94, GK97, KSJ96, PY95, Sut96, TSZC94, ZPLS96]. Basel [Ano94i]. Basic [PGC02, BKvH^{+}14, BR94]. basierte [Gra97]. Basis [OMK09, RB01]. batch [VLMPS^{+}18]. batched [GNP19]. Batching [LML^{+}19]. Bath [BP93]. Bayesian [CBS18, Fer10]. BC [IEE95i]. BCS [FFP03]. BCS-MPI [FFP03]. be [CB00]. Beach [IEE93b]. beam [OIH10, RCFS96]. bearings [NF94]. Beguelin [Ano95b, NMC95]. Behavior [BFM97, DeP03, Ros13, FGL^{+}20, LLG12, PPF89, YMYI11]. behaviour [EPML99]. Beijing [CZG^{+}08, LHHM96, Li96]. Beitrage [Ano94c]. Belgium [LCHS96]. belts [NS20]. Benard [TVV96]. Benchmark [BWV^{+}12, DS16, HC10, Luo99, Mül02, MBB^{+}12, RSPM98, RTH00, SGJ^{+}03, Trä12b, UTY02, Ano03, BKML95, DWM12, DH95, DHS96, MKP22, Mül03, MvWL^{+}10, PHJM11, PSH^{+}20, Reu01, RST02, Wor96, YSWY14]. Benchmarking [GC05, HCA16, LCY96, MMU99, MCS00, WRA02, RST02]. Benchmarks [CRE99, KS96, KAC02, MM07, NA01, RK01, TSB02, TSB03, WAS95b, ZSnH01, CDD^{+}96, MKP22, MMH99, Ste94, WT11, CE00, WT12]. Beneficial [CB00]. benefit [SBG^{+}12]. Benefits [LB16, PSM^{+}14, SIRP17]. Benutzerprofile [Wil94]. Benutzertreffens [Ano94c]. Beowulf [CMM03, Ste00, UP01]. Beowulf-Class [Ste00]. Berlin [PW95]. Bessel [KT10]. best [GT19]. Betriebssystemkern [Sei99]. Better [Str94, RS21]. Between [AAB^{+}17, BS07, ASS^{+}17, AKE00, BID95, GFV99, JAT97, LDCZ97, MSP93, WK20]. Beverly [IEE93f]. Beyond [Gei93a, GKPS97, Gei98, Gro12, LBB^{+}21, Olu14, Gei93b, LSG12, Sch93, SC19, SHM^{+}10]. Biconjugate [GFPG12]. bidirectional [HE15]. Big [CLOL18, GTS^{+}15, LK14, VPS17, ASS^{+}17, Str94]. Biharmonic [RB01]. Bill [Ano99c, Ano99d]. billion [KTJT03]. Billions [MRB17]. binary [CG93, EPP^{+}17, SGS95, TCBV10]. binary-level [EPP^{+}17]. binary-splitting [TCBV10]. Binding [CLL03, Coo95b, MG97, Coo95a]. Bindings [Ano98, VGRS16]. Bioinformatics [BBH12]. Biological [CNM11, VBB18, BA06]. Biology [SYL19]. Biomolecular [BCGL97, PZKK02]. BIP [CDP99, Tou00]. BIP-Myrinet [Tou00]. BIP/Myrinet [CDP99]. bit [HLO^{+}16, Wil93]. bit-parallel [HLO^{+}16]. bitonic [PSHL11]. Bitsliced [HMKG19]. Black [FSXZ14, Kha13, van93]. BLACS [DSW96, DS96a, Wal95]. blame [DSGS17]. BLAS [Add01, ARvW03, FMFM15]. BLASTP [HWW21, LSMW11]. Blaze [PWPD19]. Blaze-Tasks [PWPD19]. Block [ABG20, DDPR97, MYL21, SMM^{+}16, SBB20, SSB21, WO95, ZB97, ADDR95, DR18, GP95, HKMCS94, HC08, LYIP19, WO96]. Block-Based [ABG20]. Block-Cyclic [DDPR97, WO95, HKMCS94, HC08, WO96]. block-tridiagonal [DR18]. Blocking [FH98, BCH^{+}08, HKT^{+}12, Nak03, HTA08, STP^{+}19, TGKL19]. Blood [Pat93]. Blue [KMH^{+}14, AAC^{+}05, BGH^{+}05, EFR^{+}05, LM13, MV17, MSW^{+}05]. blurred [Wil94]. BMMC [CC99]. Boca [Edd18]. bodies [AGIS94, LHLK10]. Body [RB01, RTRG^{+}07, IHM05, NS16, Per99, SP99, SRK^{+}12, ADB94]. BOF [Mat00a]. Boltzmann [OTK15, CGK^{+}16, JCP^{+}20, MS95, Pri14, STA20, SJK^{+}17a, SJK^{+}17b]. bond [THDS19]. bond-order [THDS19]. Bonn [MTWD06]. Book [Ano95b, Ano95c, Ano96a, Ano99a, Ano99c, Ano99b, Ano99d, Ano00a, Ano00b, Che10, Edd18, Mar06, Nag05, NMC95, Per97, SD13, Vog13, Vre04, YM97]. books [YM97, Nov95]. Boost [CVPS19]. Boosting [LRG14, SFO95]. Bose [KLM^{+}19, MBA21]. Boston [IEE94e]. Both [BGD12, KP96, LSM^{+}18]. Bottleneck [MWG97]. bottlenecks [DSG17, JKHK08]. Boulevard [ACM99]. Bound [ASA97, SGS^{+}21, CLA^{+}19, MBKM12, ADMV05]. boundaries [KGB^{+}09]. Boundary [BS21, PTT94, STA20, SBQZ14, SP11, SD99]. boundary-value [SP11]. Bounded [CPKG17, MdSAS^{+}18, PAdS^{+}17]. BowMapCL [NTR16]. Box [JR13, JPP95]. Box-counting [JR13]. brackets [GSMK17]. Braga [IEE96g]. brain [VLSPL19]. Branch [ASA97, LW20, ADMV05]. Branches [SGS^{+}21]. Breaking [OS97]. breast [Str94]. Brest [IEE94c]. Bridge [VDL^{+}15]. Bridges [DSS00]. Bridging [ACM04, AAB^{+}17, ASS^{+}17]. Bringing [FKKC96]. Brisbane [ACDR94, Nar95]. Bristol [MC94]. British [IEE95a, IEE95e]. Broadband [OIS^{+}06, CLLASPDP99]. Broadcast [PSM^{+}14, YSP^{+}05, AMC^{+}19, MTK16]. Broadcasts [SE02]. Brownian [SKM15]. Bruijn [PGF18]. Brussels [LCHS96]. Brute [JRG21]. Brute-forcing [JRG21]. BSGP [HZG08]. BSP [Mar06, Bis04, GRRM99, Mar09, Röh00]. BSP2OMP [Mar09]. BT [WT11, WT12]. Budapest [FK95, KKD04]. Buffer [SEF^{+}16, Tsu07]. buffers [MR96]. Build [HRSA97]. Building [FD04, Gei01, Gro02a, LBD^{+}96, LVP04, WADC99, Arn95, HS95b, MSL12, PW95, Sur95b, Kos95b]. Bulk [Cer99, CLE^{+}20, DLRR99, HZG08, SRS^{+}19, TNIB17]. bulk-synchronous [CLE^{+}20, HZG08]. burden [AV18]. Burrows [NTR16]. Burst [SEF^{+}16]. BUS [ITT99]. BUSTER [XWZS96]. Butterfly [ST17]. Butterfly-Patterned [ST17].

C [Edd18, Gal97, Pri14, SM12, SSL97, TBG^{+}02, VDL^{+}15, Vre04, ABB20, BKK20, BGK08, BB00, CNC10, CCHW03, DARG13, Don06, FLMR17, FHK01, GTH96, GSI97, Gör01, KK02a, KPO00, KLM^{+}19, LYSS^{+}16, Mat16, MHSK16, QM21, Qui03, RBC20, Rót19, SSB^{+}17, SC95, TNIB17, UZC^{+}12, YULMTS^{+}17, YSVM^{+}16, ZT17]. C# [WLR05]. C-to-CUDA [UZC^{+}12]. C/C [SM12, KPO00]. C/OpenCL [RBC20]. C11 [BDW16, QM21]. C11/C [QM21]. C2CU [TNIB17]. CA [ACM95b, Ano89, BBG^{+}95]. Cache [CVPS19, LWKA15, LZH17, LZH18, MC18, MM07, NIO^{+}02, NIO^{+}03, SS01, SVC^{+}11]. Cache-Coherent [SS01]. cache-friendly [SVC^{+}11]. Cache-Oblivious [LZH17, LZH18]. Caches [LB16]. Caching [kLCCW07, DO96, WMRR17, WRMR19]. CAE [KDL^{+}95a, KDL^{+}95b]. CAF [GBR15, Mar05]. Caffe [AHHP17]. calculating [EZBA16, KD12]. Calculation [GDM18, QRMG96, GSMK17, KN17, MM95, NS16, SR11]. Calculations [RB01, Sta95b, ART17, AiIS^{+}21, Hol95, WH96]. calculus [PQ07]. Calif [IEE93f]. California [ACM97b, Gat95, IEE93a, NM95, USE94, AH95, GE95, GE96, Has95, IEE93b, IEE93f, IEE94g, IEE95c, IEE95f, LF^{+}93a]. Call [DW02, MCP17]. Call-Graph [DW02]. Callback [SSN^{+}21]. Callback-based [SSN^{+}21]. Calls [FHK01, AGLv96]. CALPHAD [TKP15]. Cambridge [Ano95b, Ano95c, Ano96a, Ano99a, Ano99c, Ano99b, Ano99d, Ano00a, Ano00b]. CAMeL [KDL^{+}95a, KDL^{+}95b]. CAMeL/PVM [KDL^{+}95a, KDL^{+}95b]. CAMP [CLM^{+}95]. Can [Gro02a, SBG^{+}12]. Canada [BG91, GGK^{+}93, IEE95a, IEE95i, Vos03, IEE95e, Lev95]. Cancellation [TBS12]. cancer [Str94]. Cancun [Sie94]. CAP [GTH96, MGMH97]. CAP-Specified [MGMH97]. Capabilities [Gei97, CG99a, IMS16]. capability [BBH^{+}13b]. Capable [KF16, KYL03]. Capacity [LYGG20, RCG95]. Capture [DW02]. Capturing [FM09]. card [SR11]. Cardiac [ORA12]. cards [KY10, KME09]. Carlo [ADRCT98, AK99, DAK98, HJBB14, NSLV16, RR00, RP95, SK00, SKM15, WH96, ZZ04]. Carnegie [IEE94d]. Carolina [ACM95a]. cars [Str94]. Cartesian [Gro19]. CASCON [GGK^{+}93]. Case [AIM97, BF01, BWW^{+}12, BfDA94, BHLS^{+}95, CML04, DARG13, DHP97, GL97a, GMdMBD^{+}07, HHC^{+}18, KCR^{+}17, LSB15, PS19b, RRBL01, SCL01, Tha98, AML^{+}99, BJ13, BJS99, Bri00, FO94, GADM20, HIIG16, JLG05, MS96b, PGK^{+}10, Pri14, SIRP17, TPD15, THMH21, Wal01b, ZSK15, LPD^{+}11]. CASTEP [STH22]. casting [KGB^{+}09]. CATCH [DW02]. Causal [ZJHS20]. Cavanaugh [IEE93c]. CAVE [BBH^{+}15]. CAVE-CL [BBH^{+}15]. cavities [BBH^{+}15]. Cavity [PKYW95, RM99]. CBFEM [OMK09]. CC [GB96, KYL03]. CC-COMA [GB96]. ccNUMA [CHPP01, CBPP02, MCS00, SSGF00]. CCp [BB00]. cCUDA [SNN^{+}20]. CE2014 [MBS15]. CEBAF [DZDR95]. Celebrating [EO15]. Cell [DBK^{+}09, SYL19, JMS14, VDL^{+}15, OOS^{+}08, OIS^{+}06]. Cell-Centered [SYL19]. Cells [MRB17]. Cellular [Car07, SC19]. Cenju [GPL^{+}96, KSHS01]. Cenju-3 [GPL^{+}96]. Cenju-4 [KSHS01]. Center [ACM98b, ACM99, ACM00, Hol12, IEE94b]. Centered [SYL19, JPOJ12]. Centers [EGR15]. Centre [IEE95e]. centric [SFSV13]. century [IEE95a]. CERN [VV95]. Cesena [CH96]. Cetraro [D^{+}95, KG93]. cf4ocl [FLMR17]. CFD [SPE95, AMS94, ADT14, CP97, HDZ^{+}20, HAJK01, HT01, JR10, DK02, PBK00, XJR21, YPAE09]. CFD-DEM [ADT14]. CG [ABF^{+}17]. CGM [CDT05]. CGMGRAPH [CDT05]. CGMGRAPH/CGMLIB [CDT05]. CGMLIB [CDT05]. CGPredict [WZM17]. Ch [CNC10]. Chain [FK01]. Challenge [DGMJ93, MKP22, LB96]. Challenges [Agr95a, Gro01a, Gro12, Ree96, Ten95, Wit16, BDG^{+}92c, GScFM13, WLK^{+}18]. CHAMELEON [KSB^{+}20]. Chamfer [YPZC95]. Chandra [Stp02]. Channel [GK97, LBD^{+}96, SG05]. CHAOS [BLW98, JL18]. Characteristic [OMK09]. Characteristics [WR01, WT12, BN00, GL99, WT11]. Characterization [AJC^{+}20, KB98, LCY19, LPJ98, MM07, Wor96]. Characterizing [BCM11, BGdS09, FLPG18, GScFM13, OdSSP12]. Charge [BL95]. Charm [ZHK06]. Charts [DSS00]. Chebyshev [Rót19]. Check [MC17, LCC^{+}03]. checkerboard [BW12]. Checking [CGZQ13, Gro00, HMK09, LCC^{+}03, MdSAS^{+}18, PAdS^{+}17, RAS16, SMAC08, YYW^{+}12]. Checkpoint [AKB^{+}19, SSB^{+}05, SBF^{+}04, CZP21, CRM14, ZWZ05, ZHK06, BDB^{+}13]. checkpoint-based [CRM14, ZHK06]. Checkpoint-on-Failure [BDB^{+}13].

Checkpoint-Recovery [SBF^{+}04].

Checkpoint/Restart

[SSB^{+}05, AKB^{+}19, CZP21].

Checkpointing [DCH02, LMRG14,

SSB^{+}05, TSS00b, BMPS03, BCH^{+}08, CG96, LCMG17, LBB^{+}19, PKD95, SSCC95, Ste96]. checkpoints [LFW20]. chemical [NMW93]. ChemIO [NFK98]. Chemistry [AKK^{+}94, NFK98, BR95a, DMW96, SSGF00]. Chemkin [Ano97, Bra97]. CHEMPI [RR01]. Chicago [CGKM11]. China [CZG^{+}08, IEE97a, LHHM96, Li96]. Chip [Jes93b, URKG12, WYZ^{+}19, TDG13, dCZG06, MYK19]. Cholesky [DG95, LC97b]. Chromosome [BM97, dOSMM^{+}16]. Chromosome-Wide [dOSMM^{+}16]. CICADA [MK94]. Cilk [Stp18]. Circuit [WPC07, BJ95]. Circuits [GJN97]. Circular [Tsu07]. Circulation [GAM^{+}02, Nes10, RSBT95]. CIS [AH00]. citation [Squ03]. City [Hol12]. civil [PW95]. CL [BHW^{+}12, BBH^{+}15, LW95]. CL-PVM [LW95]. CL ARRAY [ZT17]. clarified [WBBD15]. CLAS [DZDR95]. Class [AFGR18, DFN12, Rót19, Ste00, Dem96, MSL96, RFH^{+}95]. Classes [DeP03, GG09, Ott93]. classic [HL17]. Classical [BCGL97]. Classification [SNN^{+}19, TPLY18]. clauses [WC15]. Clemson [ACM95a]. Client [Ano93f, FSLS98, KS97, kLCCW07, Mat01b, HIIG16, Sch93, Sto98, Vis95]. Client-Agent-Server [Mat01b]. Client-Server [FSLS98, Sto98, Vis95]. Client-Side [kLCCW07]. Client/Server [Ano93f, Sch93]. climate [Str94]. CLIPS [Ano95a, Ano95e]. clMAGMA [CDD^{+}13]. clock [NB96]. clocks [TPLY18]. CLOMP [BGdS09]. clone [ZWL^{+}17]. Closer [HCZ16]. Closure [CGPR98, KH15, PPR01]. Cloud [HC17, LSB^{+}18, SIS17, URKG12, ZLZ^{+}11, ZLP17, GFIS^{+}18, GHZ12, GWVP^{+}14, KSC^{+}19]. cloud-based [KSC^{+}19]. Cluster [AUR01, BKGS02, BL95, BM97, CRE99, CMM03, HD02a, ES11, GGGC99, Gei94, Gei00, GSN^{+}01, GT01, GC05, HD02b, ITKT00, IDD94, KKH03, KS96, KS01, KHS01, LR01, MFTB95, MM01, NO02b, OF00, PFG97, RB01, RšT06, RLL01, SCR92, SHHI01, SHTS01, ST02a, TOTH99, TSN21, Trä02b, YCA18, bT01a, AL93, BLP93, BALU95, BTC^{+}17, BID95, CCF^{+}94, Cou93, ED94, GK97, GMU95, Heb93, KEGM10, KO14, Kom15, LC07, LZZ^{+}20, Liu95, MW93, MM03, NO02a, PDY14, RJDH14, SS94, SR95, ST02b, SLS96, SY95, SSN94, Tho94, THM^{+}94, Tsu95, UH96, YWO95, ZLZ^{+}11, MS04]. cluster-based [SLS96]. Cluster-enabled [SHHI01]. clustered [KHB^{+}99]. Clustering [BBH12, HA10, RJC95, GGL^{+}08, YCL14]. Clustern [MS04]. Clusters [AH00, AHHP17, AJC^{+}20, BDH^{+}95, BDH^{+}97, BWV^{+}12, CDT05, CLOL18, CSC96, DK06, GDM18, GMdMBD^{+}07, GSY^{+}13, HPP02, HSMW94, HVA^{+}16, HC17, Hus00, JNL^{+}15, LC97a, LH95, LVP04, LHCW05, MS98, MFPP03, Pan14, PKB01, PT01, PS00a, Pus95, Rei01, dOSMM^{+}16, SFG98, SvL99, Ste00, Tou00, UP01, WLNL03, WT12, YWCF15, YKI^{+}96, AB95, ALR94, ADB94, ABG^{+}96, ADMV05, BWT96, BDV03, Bru95, CRE01, EKTB99, GBF95, GDMME22, HCL05, Hus99, JKHK08, Jon96, JR10, JRM^{+}94, KYL03, KYL05, KSL^{+}12, KJEM12, LBD^{+}96, Lee12, LLC13, LL95, LKYS04, NMW93, NN95, PS07, PRS^{+}14, PM95, PR94c, PRS16, PL96, RCFS96, RGDML16, SPBR20, Slo05, SC96a, SBK21, SL95, TFZZ12, WLNL06, WLYC12, YST08, YL09, YHL11, YWC11, ZHS99, dCH93, NWT21]. CM [SBG^{+}02]. CMMD [Har94, Har95]. CMPI [GHZ12]. CMS [FMS15]. CNF [IKM^{+}01, IKM^{+}02]. CNN [MZLS20]. CO [ACM01, AHHP17, GDM18, HJ98, SNN^{+}20, PSB^{+}19, TOC18, Wal02]. co-array [TOC18, Wal02]. Co-designing [AHHP17]. co-execution [PSB^{+}19]. Co-Expression [GDM18]. Co-processed [HJ98]. Co-Scheduling [SNN^{+}20]. Coarray [GBR15, YBMCB14]. coarrays [SMCH15, SC19]. Coarse [ADRCT98, IOK00, KOI01, LGM00, NIO^{+}02, NIO^{+}03, SSK^{+}18, HDZ^{+}20, Heb93, RJC95]. Coarse-Grain [IOK00]. coarse-grained [HDZ^{+}20, Heb93, RJC95]. Coarse-Grid [SSK^{+}18]. coarsening [PSLT99]. Coast [IS16]. Coastal [GAM^{+}02]. CoCheck [MS96b, Ste96]. Code [AHP01, And98, BCGL97, CB00, CP97, CCK12, CCBPGA15, Cre16, DDL00, DZDR95, HE02, KaM10, KAMAMA17, KHS01, LD01, MMD98, MS02b, MM07, PBC^{+}01, RGD13, SM03, SZBS95a, Sta95b, TGBS05, AMS94, ADB94, AFST95, BCAD06, BADC07, BW12, Bha98, Bri95, Cou93, DLR94, EZBA16, FMFM15, GSMK17, Heb93, IJM^{+}05, JL18, KPL^{+}12, KH10, MGS^{+}15, MRH^{+}96, MWO95, PKE^{+}10, PSK^{+}10, RP95, RVKP18, SZBS95b, SK00, SFLD15, SMSW06, TBD96, VBLvdG08, VDL^{+}15, WLYL20, Wor96, XR21, YL09, ZGZS20, ZT20]. codebooks [PMM95]. Codes [FAFD15, JFY00, SWH15, HTJ^{+}16, HWS09, HASnP00, JPP95, KBG^{+}09, LRW01, Mal01, OLG^{+}16, WB96]. Coding [FLS20, Uhl94, Uhl95b, SCC96]. Coefficients [MW98, ARYT17]. COFFEE [DFSW19]. cognitive [PWD^{+}12]. Coherence [MM07]. Coherent [SS01]. Collaborative [DCPJ12, MZLS20, DCPJ14]. Collapse [PKYW95]. Collecting [BMR01]. Collection [LTRA02, DH95, MGC^{+}15]. collection-oriented [MGC^{+}15]. Collections [JFGRF12]. Collective [BIL99, BIC05, CCA00, FVD00, FCLG07, FPY08, GLB00, GMdMBD^{+}07, Hus99, KH96, KLH^{+}20, MJG^{+}12, PGAB^{+}05, SG15, TRG05, VFD02, WRA02, BPJ22, FA18, HS12, HMS^{+}19, HG12, HWW97, KHB^{+}99, KBHA94, KMH^{+}14, LFW20, MBBD13, MB21, Pan95b, PGBF^{+}07, PGAB^{+}07, RJMC93, SCB14, SCB15, SS99, TD99, Trä12a, THMH21, TFZZ12]. Collectives [CSW12, SvL99, ZGZS20, DJJ^{+}19, HGX^{+}22, Zah12]. Collector [GTS^{+}15, WK08a, WK08c, WK08b]. College [AGH^{+}95, Ano94h]. Collision [QRMG96, Sta95b, ART17, FFFC99, LHLK10]. Collocative [MKW11]. Colony [ITT02]. Colorado [R^{+}92, IEE05]. Colt [WN10]. Columbia [IEE95a, IEE95e, MAB05]. column [HSP^{+}13]. column-stores [HSP^{+}13]. COMA [GB96]. Combined [CBHH94, TJPF12, AiIS^{+}21]. Combining [DP94, LSM^{+}18, PQR18, Rab98, SCB14, Sch96a, SMAC08, YPAE09, Bor99, QM21, Sch96b]. comes [Ano94f]. Coming [HK95]. Commands [OLG01]. comments [Str94]. commerce [Ano94f]. commercial [Ano93h]. commodity [GGL^{+}08]. Common [HEH98, Per21, DK13, WLR05]. Communicating [FKK^{+}96b, GMPD98, FKK96a]. Communication [ABF^{+}17, AJC^{+}20, BCG^{+}10, BIL99, BIC05, DCPJ12, DZZY94, EM02, FST98a, FJK^{+}17, FGKT97, FBSN01, GFD03, GFB^{+}03, GGS99, GMA20, GKD^{+}18, GFV99, GLB00, GC05, HB96b, HC10, HDB^{+}12, HC06, HIP02, KB98, KV98, KBG16, LRT07, LC93, LBB^{+}21, LCVD94a, MH01, MMH98, MR96, Nit00, PLK^{+}04, RK01, RRAGM97, RšT06, SWHP05, SCP97, SGH12, SBG^{+}02, SJ02, ST02b, SGL^{+}00, SKH96, Sum12, TRG05, TGT05, TRH00, Trä02b, UMK97, WBH97, XH96, YC98, ZSG12, AC07, FH98, BHJ96, BVML12, BBH^{+}13b, BS94, BMG07, CAHT17, CGL^{+}93, Dem96, DWM12, DCPJ14, DGB^{+}14, DBB^{+}16, DS96b, DWS^{+}21, GK97, GM13, Gra97, GL94, GB94, HB96a, HWX^{+}13, Hus99, HWW97, KH96, KB01, KYL03, KYL05, KHB^{+}99, LR06b, LFL11, MLAV10, MMU99, MABG96, OGM^{+}16, Pan95b, Par93, PGK^{+}10]. communication [PM95, PKE^{+}10, PSK^{+}10, PS00b, RS21, SH14, SC95, TG09, TGKL19, Trä12a, THMH21, Vet02, WK20, Wu99, WMP14]. Communication-avoiding [GKD^{+}18]. Communication-aware [GMA20]. communication-based [PGK^{+}10]. Communication-buffers [MR96]. Communication/Computation [HIP02]. Communications [BPS01, CP98, CDHL95, CDH^{+}95, FVD00, FST98b, GT01, GBS^{+}07, GMdMBD^{+}07, IEE95b, IEE95e, LZH17, LZH18, MB00, VFD02, YTH^{+}12, bT01a, ADLL03a, ADLL03b, AiIS^{+}21, BBW19, CDP99, FA18, HS12, KBHA94, MBBD13, McR92, MN91, MS99c, RGDML16, SCB14, SCB15, TD99, WLYC12]. Communicators [DFKS01, GFD03, GFD05, FKS96, GJMM18, KH96, MJG^{+}12]. communities [ACM04]. Community [BHW^{+}17, FCP^{+}01]. Como [CLM^{+}95]. COMOPS [Luo99]. Compact [Uhl94, Uhl95b, Wor96]. compaction [VSW^{+}13, WK08a, WK08b, WK08c]. Compactly [KLR16]. Comparative [KB98, MYL21, PSK08, SN01, AGR^{+}95b, ED94, YCL14]. Comparing [BF01, DSU20, Fin97, GBR15, HVSH95, ICC02, LKJ03, ORA12, SSG95, JLG05, WBSC17]. Comparison [BvdB94, BS07, HC10, KBM97, LCW^{+}03, Mat94, Mat95, Ney00, OP10, OF00, PPJ01, Pok96, RS93, RBB97a, SS01, SR98, SHH94b, VS00, Wal02, ZBd12, Ahm97, AB93b, BLP93, BID95, EVMP20, dFdOSR^{+}19, GMU95, Har94, Har95, JS13, KDSO12, KNH^{+}18, KC06, MSP93, Ols95, PS07, PSHL11, Pri14, SdM10, SYR^{+}09, SWS^{+}12, SHH94a, TOC18, TSZC94]. comparison-based [PSHL11]. Comparisons [GGS99, PGC02, CLYC16]. Compass [PWD^{+}12]. Compatible [MM14, LBH12, OIH10]. Compcon [IEE93a]. compete [Ano96a]. CoMPI [FSC^{+}11, FCS^{+}12]. Compilation [FSSD17, HKMCS94, LRBG15, RBC20, RVKP19, SBW91, ABB20, Coe94, FM90, PGS^{+}13, PG18, SHM^{+}12]. Compile [GB94, TSY99, JE95]. Compile-time [GB94]. Compile/run [TSY99]. Compile/run-time [TSY99]. compiled [KYL03, KYL05]. Compiler [ADK22, Ano98, Dan12, IOK00, KSS00, KSHS01, MB12, Mar09, MKW11, SSE12, SKS01, TJPF12, TBG^{+}02, TGBS05, BAG17, HEHC09, LME09, LHC^{+}07, LLCD15, MA09, Mül03, PP16, RKBA^{+}13, SHHI01, SSH^{+}19, THH^{+}05, TMT^{+}20]. Compiler-assisted [ADK22]. Compilers [Ano01a, CFF^{+}94, LZ97, MKV^{+}01, SBT04, SS96, Hos12, PBG^{+}95, ZT17]. Compiling [DMB16, Hos12, CGK11]. Complete [BdS07, GHLL^{+}98, Nag05, Per97, SOHL^{+}98, YM97, Ano99a, Ano99c, Ano99b, Ano99d, PRS^{+}14, SOHL^{+}96]. Completed [PTT94]. completion [PHM^{+}22, SSN^{+}21]. Complex [BCGL97, GMPD98, MBS15, SOYHDD19, ZT20]. Complexity [NPS12, LCH^{+}22]. component [HLP10, KRKS11, Squ03]. Components [ABG20, BT01b, CT02, Fin00, Gro02a, Lus00, Wis01, GKD^{+}18, LRW01]. Composable [MLGW18]. Composed [Wel94]. Composing [PHA10, RHM^{+}17]. composite [MALM95, YPA94]. Compositing [GPC^{+}17]. Composition [CTK00, Cot04, DLB07, FC05, KH15, CFP96, SOYHDD19]. compound [LLC13, SAP16]. Comprehensive [MZLS20, RST02]. compressible [HHSM19]. Compression [BKK20, FSC^{+}11, KBS04, VPS17, AAAA16, HE15, UH96, Wu99]. compression-based [AAAA16]. COMPSAC [IEE95l]. Compton [BCD96]. Computation [BKGS02, B^{+}05, Cer99, DSM94, DSS00, EMO^{+}93, ESM^{+}94, Fer10, FF95, GS91b, HIP02, IEE94a, IEE96c, KF16, KS15b, Mar06, MR12, MSCW95, Nag05, PPR01, Sie92a, Sie92b, SMOE93, SSB21, VZT^{+}19, WTTH17, ACM97a, AC07, ABDP15, Bis04, BALU95, Bos96, BHKR95, CL93, CMH99, CKP^{+}93, Dab19, DZZY94, HLM^{+}17, HK94, KB01, KHBS19, KJJ^{+}16, KG93, Lev95, MLAV10, Neu94, NZZ94, NCKB12, PF05, PKE^{+}10, Röh00, Shi94, SH14, TBB12, TPD15, TW12, Vol93, Wan97, Was96, SM07]. computation-communication [SH14]. Computational [ALR94, CMM03, DFMD94, JFY00, KH15, Liv00, MBS15, NFK98, R^{+}92, SZBS95a, SM07, SYL19, SN01, TDBEE11, TGEM09, WPH94, Whi04, AGMJ06, BvdB94, BDG^{+}92c, BR95a, HVSC11, KBG^{+}09, PBK99, RBB15, SPE95, SZBS95b, STT96, Str94, VDL^{+}15, XR21, BR95a, CCHW03, R^{+}92, SL94a, WPH94]. Computationally [DFN12]. Computations [AGH^{+}95, ACGR97, CGU12, CGPR98, IH04, PBK00, PMvdG^{+}13, WJ12, ANS95, AASB08, BL99, CG93, DMW96, EGDK92, HJYC10, KD13, MRRP11, MR96, RBC20, Smi93b, SAP16, TS12b]. Compute [DBK^{+}09, LSM^{+}18, KKLL11, OHG19, VLMPS^{+}18, ZLZ^{+}11]. Compute-intensive [LSM^{+}18]. computed [FWS^{+}17, SSS99]. Computer [ACM06a, Ano94a, GTH96, IEE95l, IEE96h, IEE97c, IS16, KCR^{+}17, Neu94, Old02, PSB^{+}94, ST02a, Sum12, Ten95, URKG12, YTH^{+}12, BN00, BS94, BKML95, BFM96, Cal94, CLM^{+}95, GRTZ10, JWB96, Str94]. Computer-Assisted [GTH96]. Computers [Ano89, BP99, BCL00, DDP^{+}19, DGMJ93, FFP03, GC05, IEE95b, IEE95e, ITKT00, LF^{+}93a, MFTB95, PSZÉ00, SPM^{+}10, SS96, BvdB94, BB93, BBK^{+}94, DLR94, Duv92, ESB13, GBF95, KOS^{+}95a, LR06a, MMB^{+}94, NF94, POL99, PBK99, Wal94a, Wal94b]. Computing [ACM97b, ACM98b, ACM00, ACM01, ACM04, ACM06b, AJYH18, ACDR94, AIM97, BJ93, BBG^{+}95, BDG^{+}93a, BGR97a, BL95, BCP^{+}97, BRST94, BDH^{+}95, BDH^{+}97, BHNW01, BBH12, CZ95a, CGB^{+}10, CLL03, CLOL18, CNC10, Cze16, DDS^{+}94, DERC01, DPP01, DKM^{+}92, DGMS93, DT94, Edd18, FTVB00, Fer98b, FGKT97, Fos98, FS93, GLN^{+}08, GS92, Gei93a, GBD^{+}94, GSxx, Gei00, GN95, GL97a, GT94, Gua16, Hol12, HT01, IEE92, IEE93d, IEE93c, IEE94g, IEE95c, IEE95k, IEE95i, IEE96a, IEE96f, IFI95, KK02a, KS97, LCK11, LRG14, LC93, LR01, Lus00, dlFMBdlFM02, ME17, Mat16, Mat94, Mat95, MS04, Nov95, PKYW95, PR94b, PWPD19, SHTS01, SCSL12, Sin93, SSSS97, Ste00, SGS10, SW91, SGS^{+}21, Sun90a, Sun90b, Sun92, Sun93, Sun94a, Ten95, VV95, VW92, WN10, YH96, YG96, ZL17, ZL18, ACGdT02, AMKM20]. computing [ARYT17, AL92, AH95, ASCS95, Ano93h, Ano94e, Ano94h, Ano03, ADDR95, AMV94, BPG94, BDG^{+}92a, BDG^{+}94, BBB^{+}20, BKML95, Bru95, BHW^{+}12, CZ95b, CZ96, CHKK15, DLRR99, DKD08, DKB20, DW94, D^{+}95, DMW96, DE91, EKTB99, EJL92, FBD01a, FGRD01, FO94, FS95, Fer98a, FS98, FME^{+}12, FHC^{+}95, GGGC99, GS02, GS91a, GS93, Gei93b, Gei94, GH94, GkLyCY97, HP05, HW11, HH14, HPY^{+}93, HS95a, HH95, mH12, IEE97a, IM95, JPOJ12, JY95, JJM^{+}11, JPTE94, KO14, Kos95b, KSSS07, LV12, LH98, LCHS96, LHD^{+}94, LHD^{+}95, LM13, Maf94, MZK93, Mal95, Mar07, NRdA^{+}20, PGS^{+}13, PKB06, Pen95, PGK^{+}10, PTT94, PBG^{+}95, PNV01, PWD^{+}12, RBS94, RJDH14, RRJ^{+}20, Sch93, SGS95, SMS00, STT96, SBK21, Sti94, SP11, Sun94b, SGDM94, Sun95, Swa01, SD99, TJD09, TKP15, TDB00]. computing [Tho94, TSS98, VM94, Vis95, Was96, XJR21, YULMTS^{+}17, YLC16, YSL^{+}12, Zem94, ZWL13, ZGC94, ZHS99, ZKRA14, ACM98a, Kon00, PW95, Per96, SCR92, TGEM09, NMC95, Ano95b]. Concept [KaM10, LTR00, SB95]. concern [Ano94i]. Concurrency [ME17, NPS12, DGB^{+}14, EBB^{+}20, PTG13]. Concurrent [Ano89, BDG^{+}91b, BRS92, BHV12, BKH^{+}13, DG95, GS91b, GS92, GSxx, Gre94, HS93, SNN^{+}20, SPB^{+}17, Sun92, Sun93, ZDR01, BDG^{+}92a, FS95, GS91a, GS93, LPD^{+}11, NP12, RGDML16, RCG95, Sun94b, SGDM94, Wal94a, Wal94b, WK08a, WK08b, WK08c, ZWZ^{+}95]. condensates [KLM^{+}19, MBA21]. condensed [MC99]. Condition [GK10]. Conditional [JCP^{+}20, SGS^{+}21, CBS18]. conditions [STA20]. Condor [CF01, PL96]. conduction [iSYS12]. Cone [RCFS96, OIH10]. Conference [ACM90, ACM94, ACM96b, ACM96c, ACM97b, ACM98b, ACM04, Abr96, ATC94, AGH^{+}95, Ano89, Ano93g, Ano94a, Ano94e, Ano94i, ACDR94, BBG^{+}95, B^{+}05, Boi97, Bos96, BFMR96, BH95, CGB^{+}10, CH96, DSM94, DSZ94, DKD07, DKM^{+}92, ERS95, ERS96, EJL92, FF95, Gat95, GN95, GT94, Ham95a, HAM95b, HS95a, HS94, Hol12, IEE92, IEE94f, IEE95b, IEE95a, IEE95e, IEE95i, IEE95l, IEE95j, IEE96a, IEE96d, IEE96h, IEE96i, IEE02, LCK11, LF^{+}93a, MMH93, Nar95, OL05, PR94b, Ree96, R^{+}92, SPE95, Sil96, SM07, Sin93, SW91, USE95, USE00, VW92, Vol93, WPH94, Y^{+}93, YH96, ACM95a, ACM05, ACM06b, ANS95, Ano93b, Ano93c, Ano95a, BR95a, Bil95, BDLS96, DR94, Eng00, GH94, JPTE94, LCHS96, Mal95, PW95, RV00, Van95, ZL96, ACM94, Ano94g, IEE95b, KKDV03]. Configurable [IEE94d, MYK19, PKB^{+}16, BB94]. configurations [PTL^{+}16]. conflict [TCP15]. conformational [MK94]. Congress [CJNW95, GHH^{+}93, PSB^{+}94, BH95, dGJM94]. Congressi [GT94]. Conjugate [BG95, GFPG12, SSK^{+}18, BAC20, MM92, Ols95]. Connected [ABG20, BT01b, KRKS11, OF00, Pet01, GKD^{+}18].

Connectivity [Whi94]. Conquer [CTK01, Cza02, Cza03]. conscious [ZA14]. Considerations [CJPC19, FA18]. Considers [WYZ^{+}19]. consistency [DPFT19, KSTM20, WBSC17, YYW^{+}12]. Consistent [TGT10, CG96, CG99a]. Console [PES99]. Consortium [BRST94]. Constrained [BSH15, EGR15, TSCS14]. constraints [BMS19]. Construct [DP94, EM94]. Constructing [DM93]. construction [ART17]. Constructor [MYK19]. Constructs [KDT^{+}12, PGC02, BKH^{+}13, BN00]. consumer [ACJ12]. Contact [Nak03]. CONTAIN [SBR95]. containers [Str12, ZT17]. content [GFB^{+}14]. Contention [ALB^{+}18, ALW^{+}15, DSG17, SSD^{+}20, Zah12]. Context [DGG^{+}12, ZL18, DR18, EVMP20, MdSAS^{+}18, OLG^{+}16, PAdS^{+}17, SCB15]. context-bounded [MdSAS^{+}18, PAdS^{+}17]. Contexts [CS14]. Contiguous [KLH^{+}20, WTR03]. continual [NS16]. continuation [VY15]. Continuations [SSN^{+}21]. Continuous [TA14]. Contour [GFJT19]. Contract [KPNM16]. Contract-based [KPNM16]. contrarian [KSSS07]. Contrasts [GGS99]. Control [FLD98, FM09, IEE94e, MSS97, CMZ99, MBKM12, MH18, OHG19, RRJ^{+}20, SFL^{+}94, SHPT00]. control-flow [MH18]. Controlled [DSU20]. controller [GWC95]. convection [BB95b, CEGS07, TVV96]. Convention [ACM98b, ACM99, ACM00, Hol12, IEE94b]. Converse [BK96]. Conversion [ZG95b]. convex [GCN^{+}13]. Convolution [ADGA20, WTS19]. convolutions [DZZY94]. Cook [SD13]. Cooperation [Wis01, Str94]. Cooperative [DGF97, DiN96, HRSA97, kLCCW07, Pet00a, Pet00b, JKN^{+}13, SHLM14]. Coordinate [OP98, LFW20]. coordinated [BCH^{+}08]. COORDINATION [CH96, KAHS96, FKK96a, CH96]. copies [RS19]. Copley [IEE94e]. Copperhead [CGK11]. Coprocessor [BB18]. Copy [SWHP05, HCC^{+}20]. copying [SH96]. CORBA [DPP01, Fin97, LRW01]. Core [ABB^{+}10, Bri10, CZG^{+}08, LZH17, SOHL^{+}98, TCM18, YGH^{+}14, YTH^{+}12, ACMZR11, AV18, BBC^{+}19, BBG^{+}14, BL99, FHB^{+}13, HTA08, JR13, JJM^{+}11, JR10, KSG13, LLCD15, LLH^{+}14, MBBD13, PZ12, SFSV13, SVC^{+}11, TFZZ12, VDL^{+}15, WCC^{+}07, WYLC12, dCZG06, MMH98, Nag05, Ano99a, Ano99b]. Cores [BBG^{+}11, DT17, BMS^{+}17, DJJ^{+}19, SC19, WO09]. Corfu [SM07]. correct [DM93]. Correction [SSLMW10, BCD96, FME^{+}12]. Corrections [BL95, DLLZ20, Spe19]. Correctness [DFP^{+}19, HMK09]. Correlated [MM07]. corruption [FME^{+}12]. Coscheduling [GRV01, SGHL01]. Cosenza [KG93]. cosmological [BADC07, Sai10]. Cost [FLS20, KS15b, LSB^{+}18, RLL01, GK97, GWVP^{+}14, MH21, Wu99]. Cost-Efficient [LSB^{+}18]. costs [GB94, LFS^{+}19]. Cots [HHC^{+}18]. count [KVGH11]. Countdown [CBB^{+}20, CBB^{+}21]. counters [Rab99]. counting [JR13]. County [ACM98b]. Coupled [MBS15, SS01, SBR95, Gra97, MBA21, TK19]. Coupling [BS93, KR09, SB95, WB96]. course [STT96]. Coverage [GSYT21]. Covering [MYK19]. CoW [KMG99]. CPPvm [Gör01]. CPS [Mat94]. CPU [BB18, CLOL18, DF17, EBB^{+}20, HSO^{+}21, HCC^{+}20, JR13, KSL^{+}12, Lee12, LRG14, LLC13, LFL11, OFA^{+}15, PDY14, PHO^{+}15, Pri14, RBC20, SdR^{+}21, SPB^{+}17, SSB^{+}17]. CPU-GPU [HSO^{+}21]. CPU-MIC [BB18]. CPU/GPU [EBB^{+}20, KSL^{+}12, Lee12, LLC13, OFA^{+}15, RBC20, SSB^{+}17]. CPU/multi [SAP16]. CPUs [ASB18, KH12, LNK^{+}15, ÓN12, SFSV13, YSWY14]. CPVM [CG96]. Cracow [BDW97]. cranial [NAJ99]. CRANIUM [MBES94]. Crash [LCVD94b]. Crash-simulation [LCVD94b]. crashworthiness [LCVD94a]. Crawler [Wal01a]. Cray [BL94, GRRM99, MP95, Sch96a, Sch96b, ABG^{+}96, AZ95, AFST95, BBW19, CCSM97, LKJ03, LSK04, MWO95, Oed93, RBB97c, SWS^{+}12, SCC95]. CRAY-T3D [Sch96a, Sch96b]. CRAY-T3E [Che99]. CRC [Edd18]. Creation [Hat98, MFC98, PS00a]. Crew [GHL97]. CRI [MSCW95]. CRI-MAP [MSCW95]. Critical [DSGS17, SLN^{+}12, KSC^{+}19, SDJ17]. Critical-blame [DSGS17]. critical-path [SDJ17]. cross [JR13]. cross-platform [JR13]. Crossbar [ZL17]. CRState [CZP21]. cryptanalysis [BSN95]. Cryptographic [PV97, ABDP15]. cryptosystem [WLC07]. Crystals [LHZ^{+}20, lLLmH^{+}21]. CS [FST98a, FST98b, Jon96]. CS-2 [FST98a, FST98b]. CS/2 [Jon96]. CT [DYN^{+}06, NAJ99]. CT-scans [NAJ99]. cube [Pan95a]. Cubes [DERC01]. CUDA [DLLZ20, Pri14, AMuHK15, AMKM20, AAAA16, ACMZR11, AC17, Ano12, AFG21, ASB18, BHS18, BY12, BTC^{+}17, BAG17, BSH15, BBH12, CAM12, CGU12, CNM11, CLYC16, CBM^{+}08, CSV12, CFF19, CB11, Cza13, DCD^{+}14, DSU20, DS13, DR18, DS22, DARG13, DLLZ19, DLV16, DWL^{+}10, DWL^{+}12, DM12, Edd18, EADT19, EPP^{+}17, ER12, FJZ^{+}14, Fer10, FMFM15, FFM11, FWS^{+}17, Fuj08, GDC15, GScFM13, GLN^{+}08, GÓ19, GML^{+}16, GDEBC20, GFPG12, GWVP^{+}14, GRTZ10, HE13, HJBB14, HVA^{+}16, HLM^{+}17, HD11, HLP10, HP11, HLP11, Hog13, HF14a, HF14b, HKOO11, HT08, HWW21, HLO^{+}16, JRG21, JL18, JPL22, JK10, JC17, JLS^{+}14, JFGRF12, KRKS11, KHBS19, KD12, KAMAMA17, Kha13, KS13, KC19, KSC^{+}19, KKB^{+}21, KF16, KVGH11, KME09, KO14, KH15, KD13, KA13, Lan09, LRG14, LGKQ10, LLG12]. CUDA [LSSZ15, LBH12, LSVMW08, LSMW11, LAD16, LBB^{+}16, LYSS^{+}16, LYIP19, LYZ13, MMO^{+}16, MV20, MNYN21, MR12, Mat16, MSML10, MdSAS^{+}18, MGL^{+}17, MM14, MH21, NSLV16, NS20, NS16, NBGS08, OIH10, ORA12, OHG19, PGS^{+}13, PRS^{+}14, PGD18, PHJM11, PAdS^{+}17, PGdCJ^{+}18, PSHL11, PSH^{+}20, PTMF18, PSV19, PRS16, RBW^{+}20, RS22, RBAI17, Ros13, RTN21, SSE12, STA20, SK10, iSYS12, SDJ17, STK08, SS09, Seg10, SSLMW10, SKM15, SP11, Stp20, SR11, SJK^{+}17a, SJK^{+}17b, TNIB17, TVCB18, TS12b, TA14, TCP15, Tsu12, UZC^{+}12, VLMPS^{+}18, WGG^{+}19, WG17, WJ12, WMRR17, WRMR19, WWFT11, WJB14, XXL13, YULMTS^{+}17, YHL11, YZ14, YW21, YMYI11, ZJHS20, ZSK15, ZAFAM16, ZWLZ21, ZZG^{+}14, ZBd12, ZLS^{+}15, ZZZ^{+}15, dlAMC11, dlAMCFN12, vdLJR11, Che10, SD13, Vog13]. CUDA-Aware [DS22, HVA^{+}16]. CUDA-Based [DLLZ20, DLLZ19, ZJHS20, AAAA16, AFG21, WGG^{+}19]. CUDA-BLASTP [LSMW11]. CUDA-C [YULMTS^{+}17]. CUDA-compatible [LBH12]. CUDA-Enabled [LSMW11, SSLMW10, DS13, KHBS19, PSV19, SR11, ZLS^{+}15]. CUDA-JMI [GDEBC20]. CUDA-NP [YZ14]. CUDA-powered [RTN21]. CUDA-quicksort [MMO^{+}16]. CUDA-sharing [PRS^{+}14]. CUDA-streams [TVCB18]. CUDA-to-OpenCL [GScFM13]. CUDA/MPI [LYSS^{+}16]. cudaBayesreg [Fer10]. CUDAEASY [Sai10]. CUDAlign [SdM10, dOSMM^{+}16]. CUDAs [KMM15]. CUDA^{T M} [SM12]. culling [LHLK10]. CUMODP [HLM^{+}17]. CUMULVS [GKP97]. cuPC [ZJHS20]. cuPentBatch [GNP19]. CURAND [Ano12]. CURD [PGD18]. Current [Bak98, GFD05, IFI95, BDG^{+}93b, FK94, FHP^{+}95]. Curse [OS97].

Curve [Rót19]. cuSten [GÓ19]. Customization [GSY^{+}13]. cut [CG99a, CXB^{+}12]. cut-through [CXB^{+}12]. cuTensor [ZLWW20]. cuTensor-Tubal [ZLWW20]. cuThomasBatch [VLMPS^{+}18]. cuThomasVBatch [VLMPS^{+}18]. cuts [GKD^{+}18]. CVFuzz [LCH^{+}22]. CVL [Har94]. Cybernetics [IEE95a]. cycles [PL96]. Cyclic [DDPR97, WO95, HKMCS94, HC08, WO96]. Cyclops [dCZG06]. Cyclops-64 [dCZG06].

D [And98, DYN^{+}06, SSS99, SH14, VDL^{+}15, Bha98, BCL00, Bri95, BMPZ94a, BAS13, CGU12, CP15, EFR^{+}05, ES11, GCN^{+}13, HF14a, HF14b, JR10, KRKS11, KO14, KD13, KHS01, KLR16, MK94, MSZG17, NSM12, SC19, TPD15, WMRR17, WRMR19, WR01, YSL^{+}12, vHKS94]. D-CICADA [MK94]. DAC [Cza02, Cza03]. Daemon [LB98]. DAG [SGL^{+}20]. Dagum [Stp02]. d’Aix [GA96]. d’Aix-Marlioz [GA96]. Dallas [ACM00, IEE95l]. Dame [IEE96i, PG18]. damping [YPA94]. DAMPVM [Cza02, Cza03]. DAMPVM/DAC [Cza02, Cza03]. DAMS [CD98]. Dangers [BCP^{+}97]. DaReL [KN95]. Data [AJF16, BMR01, BCG^{+}10, BKK20, BGD12, CKmWH16, CLOL18, DK20, DERC01, DiN96, EGR15, Edd18, EASS95, FLS20, GTS^{+}15, GSYT21, GB98, GMPD98, Gua16, HA10, HB96b, HC06, IADB19, JDB^{+}14, KA13, LK14, LSM^{+}18, LHCW05, LDJK13, LBB^{+}21, MV17, Man01, MK17, ME17, Mat16, MGA^{+}17, MJB15, NJ01, NPP^{+}00b, NPP^{+}00c, NA01, NLRH07, PCY14, RJ21, Rei01, SGH12, SPK96, SSLMW10, SR96, Str12, THS^{+}15, TPK^{+}19, WO95, Wel94, ZDR01, ZG95b, Zho21, AB95, ASS^{+}17, AGG^{+}95, BK11, Ben95, BR12, BID95, CFKL00, CGK11, CGL^{+}93, DRUE12, EP96, FB97, Fan98, FVLS15, FME^{+}12, FKK^{+}96b, FWS^{+}17, GE95, GE96, HB96a, HC08, JB96, JCP15, JE95, JPOJ12, KN95, KJJ^{+}16, KRG13, LOHA01, LF^{+}93a, LL16, LW20, MA09, MMB^{+}94, MMM13, MR96, NCB^{+}12, NCB^{+}17, NPP^{+}00a, OPP00]. data [PDY14, PG18, RJMC93, SJLM14, SSS99, SPH95, SK92, TW12, TGKL19, WO96, WZW21, WLK^{+}18, YCL14, YWO95, ZJDW18, ZRQA11]. Data- [LSM^{+}18]. data-centered [JPOJ12]. Data-Driven [ME17, NCB^{+}12, NCB^{+}17]. Data-Intensive [LBB^{+}21, Rei01]. Data-Parallel [AJF16, GB98, RJ21, CKmWH16, SPK96, CGL^{+}93, FKK^{+}96b, MMB^{+}94, MR96, SK92]. data-parallelism [BR12]. data-privatization [KRG13]. Data-Structures [GMPD98]. Databank [FCP^{+}01]. Database [AR01, BFZ97, EK97, LBB^{+}21, MWG97, MM14, PPT96a, MN91, PPT96b, PPT96c, PMZM16]. Databases [RGB^{+}18, BA06, Bos96, ZWL13]. Dataflow [DT17, CSPM^{+}96]. Datasets [DLLZ19, DLLZ20, VPS17, KGB^{+}09]. Datatype [Gro00, SWHP05, HCC^{+}20, KHS12]. Datatypes [JDB^{+}14, RTH00, SGH12, Tha98, CAHT17, THRZ99]. Dave [Stp02]. David [Ano96a, Ano99a, Ano99b, Nag05]. DawnCC [MGA^{+}17]. DAWNING [HWM02]. DAWNING-3000 [HWM02]. Day [IS16]. dbx [NE98, NE01]. DC [B^{+}05, IEE94h, IEE95k]. DCE [Sch93, FLD96, RS93, Sch93]. DDL [FB97]. Deadlock [LZC^{+}02, SG12, HPS^{+}12, HPS^{+}13]. Deadlocks [FJK^{+}17]. Debbuger [WCS99]. Debugger [HM01, NE01, CH94, CG99b, MT96, XWZS96]. Debuggers [Ano01a]. Debugging [BDGS93, GKP96, KKV01, KV98, Mor95, NE98, Wis97, ZLL^{+}12, BL97, BS96a, DKF93, DH22, HLOC96, KCD^{+}97, MLA^{+}14]. December [Bil95, Eng00, HHK94, IEE96a, Kum94, NM95, PBPT95, Y^{+}93].

Decimation [PCY14]. Declarative [EADT19]. decoder [MC17]. Decomposition [BKK20, BJS97, CP97, EGH^{+}14, KDHZ18, DBVF01, ETV94, OMK09, SHHC18, TGS^{+}20]. decompositions [NZZ94]. deconfliction [TCP15]. Dedicated [WLNL03, DJJ^{+}19, Hus99, RSC^{+}19, WLNL06]. Deep [AHHP17, AJC^{+}20, GDS^{+}20, SM19, TWLL19, AMC^{+}19, NWT21, SEC15]. Deep-Learning [AJC^{+}20]. Deferred [Spe19]. Defined [Gua16]. Defining [GAML01]. Deformable [STK08]. Deforming [GAP97]. Degree [CTBT21, CT13]. degrees [KTJT03]. Delaunay [CWL^{+}20]. DELAUNAYSPARSE [CWL^{+}20]. Delegation [YTH^{+}12]. Delegation-Based [YTH^{+}12]. Delft [DSZ94]. Delivering [Hus98]. Delphi [ACGdT02]. Demand [CTK00, LSB^{+}18]. Denmark [DW94, DMW96, Was96]. Dense [AKL16, BDT08, CDD^{+}13, Fuj08, Hog13, PMvdG^{+}13, ZBd12, BRR99, LRLG19]. Densities [MW98]. Density [BL95, MC17, CBHH94, ZWHS95]. Denver [ACM01, IEE05, R^{+}92]. Dependable [GM95]. Dependant [BP99]. Dependence [LAdS^{+}15, BLVB18]. dependence-aware [BLVB18]. Dependency [PPR01]. Dependent [DFA^{+}09, HO14, MFTB95, DM12, LBB^{+}16, LYSS^{+}16, ÓN12, RS22, SSB^{+}16, TVV96, YPA94, YSVM^{+}16, YSMA^{+}17]. DEPICT [HM01]. Deploying [PKB01, CLLASPDP99]. depth [MKP22, SSS99]. Derivation [GB98]. Derived [JDB^{+}14, RTH00, SWHP05, Tha98, CAHT17, HCC^{+}20, Jou94, THRZ99]. Descent [Sch01]. description [TKP15]. descriptors [LNW^{+}12]. Design [AS92, AAC^{+}05, Ano01b, ACD^{+}09, BCD^{+}15, BBH^{+}13b, BS96b, BMR02, BRM03, CLP^{+}99, ETWaM12, FD02a, FA18, FFP03, GG09, HWM02, JSH^{+}05, KVGH11, LML^{+}19, kLCC^{+}06, kL11, LVP04, Man94, MMSW02, MZLS20, NPS12, OFA^{+}15, Pan14, PLK^{+}04, PCS94, SBG^{+}02, SWYC94, SSL97, SPK^{+}12, Sum12, THM^{+}94, TPV20, USE94, VGRS16, ZGZS20, BR91, BMS19, CARB10, CSS95, DS96b, FD02b, FGL^{+}20, GL94, GkLyCY97, HDZ^{+}20, KA95, LC07, MAS06, OA17, PGK^{+}10, PTW99, RSC^{+}19, SL94b, Sep93, Sil96, SSD^{+}94, SWL^{+}01, WHMO19, Wal94a, Wal94b]. design-pattern [MAS06]. designed [BSH15]. Designing [BS21, GKZ12, LAD16, SWHP05, SH14, WYLC12, ZLP17, AHHP17, DSOF11, Pan95b]. Designs [HVA^{+}16, SM19, AAAA16, MC17, Shi94]. desktop [Mar07]. detach [PHM^{+}22]. Detailed [DLV16, RSPM98, BTC^{+}17, LR06b]. detect [DPFT19, Str94]. Detecting [AGG^{+}95, LCH^{+}22, PPJ01, ZRQA11]. Detection [BHW^{+}17, CSW12, CBL10, CFMR95, DMMV97, EML98, FME^{+}12, HHC^{+}18, KSJ14, SG12, ZDD97, BBH^{+}15, DKF94a, HDDG09, HGMW12, HPS^{+}12, HPS^{+}13, LZC^{+}02, RAGJ95, TCP15, TDG13, TWFO09, WTFO14, YULMTS^{+}17]. Detector [DZDR95, PGD18]. Determination [LAFA15]. Determine [BP99]. Determinism [CTBT21]. Deterministic [CFMR95, DK02, ZLL^{+}12, MV20]. Develop [PD98]. Developer [IEE96i]. developers [Str94]. Developing [BFZ97, CCSM97, Cot98, DDLM95, Reu03]. Development [AC17, Ano01a, BDG^{+}91b, BR95c, CHPP01, Cha02, Cot97, Cza02, DF21, DeP03, PS01a, SK00, SB01, TBD96, TDBEE11, ARvW03, ABC^{+}00, BL97, BDG^{+}92a, DSZ94, DHP97, KCD^{+}97, LLC13, MMW96, PES99, SM12, TBB12, ZL96, Sei99]. Developments [Mat00a]. device [KKLL11, LS10, SBQZ14, YWTC15].

Devices [GJN97, RVKP18, ZJDW18]. DFB [WWZ^{+}96]. DFN [RS93]. DFN-RPC [RS93]. DG [MV20]. DG-MOSFETs [MV20]. DGX [GDS^{+}20]. DGX-1 [GDS^{+}20]. DGX-1/Pascal [GDS^{+}20]. Diagnosis [AP96, LAdS^{+}15]. diagnostic [RSBT95]. dictionary [LSSZ15]. Diego [Has95, LF^{+}93a, NM95]. Difference [UZC^{+}12, CdOO^{+}20, GÓ19, GFPG12, HE13, NZZ94, NB96, Pri14, Ram07, Str94, VM94]. Differences [AKE00, LDCZ97]. Different [AIM97, DSU20, GL97b, JCH^{+}08, Ney00, Rab98, RBB97a, BN00, PY95]. Differential [MFTB95, MKK21, Riz17, DFSW19, JK10, MPS20, NF94, RBB15, SP11]. Differentiating [Cer99]. Differentiation [BBH^{+}08, BGK08, HH22, CdGM96, HHSM19]. Diffusion [HF14a, HF14b, MW98, CEGS07, DM93, MM92]. Digest [IEE93a, IEE95c]. Digit [DALD18, LAD16]. Digital [KLR16, CIJ^{+}10]. Dijon [YH96]. Dimemas [GLB00]. Dimensional [Car07, GA96, HD02b, KD12, LRQ01, MW98, SJK^{+}17a, SJK^{+}17b, ZWLZ21, AL93, KT02, LSSZ15, lLLmH^{+}21, MKK21, Ols95, PR94c, Ram07, RG18, ZWC21]. Dimensions [CWL^{+}20, SAS01, Ano93h, HP11, LZC^{+}20]. Diophantine [ZTD19]. dipolar [LBB^{+}16, LYSS^{+}16]. DIPORSI [GGCGO01]. DipSystem [SPL99]. Direct [Bri10, GPC^{+}17, LB98, WJB14, BCM^{+}16, Gra09, HWS09, MM11, SWH15]. direction [BDG^{+}93b]. Directions [IFI95, FK94, FHP^{+}95, Sun96]. directive [CPM^{+}18, LV12, NO02a, YL09]. directive-based [CPM^{+}18, LV12, YL09]. directive/MPI [NO02a]. Directives [AAB^{+}16, BBG^{+}99, BBG^{+}01, BKO00, CCBPGA15, JFY00, BC19b, LOHA01, VGS14]. directory [JCP15]. discharges [LZC^{+}20]. Disciplined [LWKA15]. Discontinuous [CF19, KK19]. Discovering [FJK^{+}17]. discovery [ASAK19, BK11, GWVP^{+}14]. Discrete [SSB21, ST17, WMC^{+}18, YW21]. Discrete-Event [WMC^{+}18]. diskless [PKD95]. Disks [dlFMBdlFM02]. Dispersion [RSV^{+}05]. Displacement [BJS97, PSSS01]. Dissemination [GL97a]. Distance [MR12]. Distances [LAFA15]. Distributed [AGS97, Ano95e, BMS^{+}17, BME02, BGR97a, BL95, Bha93, BJ95, BRST94, BT01b, BHKR95, CGB^{+}10, CLL03, CSW97, CC99, DMB16, DBA97, DFMD94, DGF97, DHHW92, DHHW93a, EMO^{+}93, ESM^{+}94, FH95, Fan98, FTVB00, FK01, Fos98, FS93, FFFC99, GGCM99, GGCGO01, GCGS98, GCBM97, GWC95, GM95, HJ98, HC10, HRSA97, IEE93d, IEE93c, IEE94d, IEE94g, IEE95h, IEE95k, IEE95i, IEE95g, IEE96b, IEE96g, IEE96f, IEE05, JML01, KBA02, KP96, KDL^{+}95b, KL95, KK02b, KSHS01, LC93, LHD^{+}94, LHD^{+}95, MC18, MZK93, MB12, MFTB95, MSCW95, Mat95, MBE03, NSBR07, NZZ94, NH95, Pen95, PKYW95, Pet00a, Pet00b, PTT94, PMM95, PBK00, PD98, PMvdG^{+}13, RGD97, Sch94, SA93, SMOE93, SW91, Sun90a, Sun90b, SPNB14, TSS00b, THN00, Wil93, WO97, WCS99, YH96, ZDD97, ZDR01, AMBG93]. distributed [AGR^{+}95b, AB95, Ano94e, Arn95, ADMV05, BSC99, BB95a, Bir94, BMPZ94a, CBPP02, CH94, CEF^{+}95, CBHH94, CLLASPDP99, CPR^{+}95, CK99, DLR94, DR94, DHHW93b, DR95, EGH99, FB97, FS95, FS98, FHC^{+}95, FHB^{+}13, GBR97, GCN^{+}10, GKK09, GkLyCY97, GP95, HPY^{+}93, HHA95, IEE97a, JWB96, KN95, KSG13, KJJ^{+}16, KDL^{+}95a, LR06b, LFS93a, LFS93b, LH98, LKL96, Liu95, LYIP19, LGMdRA^{+}19, Maf94, MVTP96, Man98, MLC04, NAJ99, OLG^{+}16, PK05, POL99, Par93, PR94c, RBW^{+}20, RAGJ95, RFH^{+}95, SSH08, SHHI01, SL94b, Sch93, SFL^{+}94, SSC96, SPL99, Smi93b, SD99, THDS19, TSP95, THM^{+}94, Uhl95a, VM94, VB99, Vet02, Vis95, Wal94a, Wal94b, WPL95, Wan97, YLC16, YWO95, YX95, YPZC95, YZPC95, ZL96, ZGC94, ZHS99, Pet01]. distributed-data [FB97]. Distributed-Memory [CSW97, CC99, KN95, SSH08]. distributed-shared [ADMV05]. Distributing [AL92]. Distribution [HB96b, LHCW05, MJB15, NPP^{+}00b, NPP^{+}00c, NA01, SR96, AGG^{+}95, CSW99, GS96, HB96a, JMdVG^{+}17, KRC17, NPP^{+}00a, RJMC93, Wil94]. Distributions [ST17, WO95, HKMCS94, WO96, vHKS94]. Divergence [SdSCP13, WYH^{+}21, LW20, VSW^{+}13]. Divergent [WJA^{+}19]. diversity [EO15]. Divide [CTK01, Cza02, Cza03]. Divide-and-Conquer [CTK01, Cza02, Cza03]. DMMP [BB93]. DMPI [HWM02, ZLL^{+}12]. DNA [dFdOSR^{+}19, GDMME22, PGF18]. DNAml [CDZ^{+}98]. DNMR [SR11]. do [JLG05]. dOCAL [RBW^{+}20]. docking [ESB13, IPG^{+}18, RJH^{+}20, VGP^{+}19, ZWL13]. Document [MHSK16, AD95]. Documentation [BDG^{+}xx]. Documents [Ano98]. does [KC94]. dog [LK14]. Domain [BMR01, CP97, EGH^{+}14, KDHZ18, kL11, ETV94, HE13, Nel93, NZZ94, Olu14, OMK09, Ram07, SHHC18, VM94]. Domaine [GA96]. Domains [KR09]. Dongarra [Ano95b, Ano96a, Ano99a, Ano99b, NMC95, Nag05]. dOpenCL [KSG13]. dot2dot [GDMME22]. Double [FKKC96, PTT94]. down [Str94]. Downloadable [Ano98]. DP [Arn95, KLR^{+}15]. DPVM [IHvA^{+}00]. DQN [PS19a]. DQN-based [PS19a]. draft [DHHW93b, GL92]. Draw [ST17]. Dresden [MdSC09]. Driven [AIM97, LWSB19, ME17, PCY14, FSG19a, FSG19b, Hin11, NCB^{+}12, NCB^{+}17, Qu95, SIS17, TWFO09, WTFO14].

Dror [Stp02]. drug [GWVP^{+}14]. drugs [Str94]. DSIR [LTR00, RTL99]. DSM [KBVP07]. DSMC [JL18]. DSMPI [SSC96, SSC97]. DTM [PS07]. DTS [BHKR95]. Dual [BBC^{+}00, GAM^{+}02, DK02, CT13, LSSZ15]. dual-dictionary [LSSZ15]. Dual-Level [BBC^{+}00, GAM^{+}02, DK02]. dual-scanline [CT13]. Dublin [LKD08]. During [DeP03, DH22]. Dust [dlFMBdlFM02]. DVFS [PTL^{+}16]. DWT [ZZZ^{+}15]. Dyn [WLNL03, WLNL06]. Dyn-MPI [WLNL03, WLNL06]. Dynamic [ACGR97, AGS97, AUR01, BBD^{+}20, CGLD01, CKmWH16, CML04, CK99, CTK01, DMB16, DBA97, DFMD94, FMBM96, FD00, GFD03, GFD05, GRV01, GCBL12, GMPD98, GL95a, KFL05, LZZ^{+}20, MK17, NPP^{+}00c, NLRH07, PK98, PLK^{+}04, PT01, PGdCJ^{+}18, Ran05, SPH^{+}18, Smi93b, SY95, TS12a, TPK^{+}19, VdS00, Vet02, Wal01a, Wil94, YST08, Zel95, DDLM95, EO15, FH97, FCS^{+}12, FKLB08, JC17, MSMC15, NSBR07, NF94, OKW95, PGD18, PSH^{+}20, RBAI17, RCG95, SCB14, SCB15, SKK^{+}12, SKB^{+}14, WRSY16, YPA94, DvdLVS94, FCS^{+}12]. Dynamically [HDW21, SSS99]. DynamicPVM [DvdLVS94]. Dynamics [BST^{+}13, BCGL97, DR97, JFY00, KBM97, dlFMBdlFM02, MH01, OS97, SZBS95a, SA93, TDBEE11, TGEM09, YWCF15, ZB94, ALR94, AiIS^{+}21, ABG^{+}96, AGMJ06, BvdB94, BHS18, BvdSvD95, BBK^{+}94, BMPZ94b, BMPZ94a, CC00b, FHSO99, HHS18, HVSC11, JAT97, JMS14, KFA96, KPK13, KRG13, LHZ^{+}20, LSVMW08, NS20, OKM12, PARB14, PBK99, PIR^{+}20, RS22, RBB15, SPE95, SZBS95b, SKM15, TG94, WPH94, XR21]. Dynamische [Wil94]. dynamite [IvdLH^{+}00, IHvA^{+}00]. Dynamite/DPVM [IHvA^{+}00]. dynamo [Hol95]. DySel [CKmWH16].

E-scale [Gua16]. EA [Ben18]. each [Ano00a, Ano00b]. Early [CD96, LV12, SLG95, EFR^{+}05, HHK^{+}19, KJA^{+}93]. Earth [KTJT03, Nak03, Nak05a, Nak05b, UTY02]. Earthquake [UZC^{+}12, KTJT03, KME09]. Easily [PKB01]. East [IS16]. Easy [HCA16, TDG13, MJPB16, SBF94]. EasyGrid [BR04]. EASYPVM [Saa94]. ECMWF [HK93, HK95]. ed [Nag05]. EDEM [Tsu95]. Edge [ZDD97, Gra97, RAGJ95]. edition [Ano99a, Ano99b, Ano00b]. Editor [GT19]. Editors [AM07, GSA08]. EDP [SdR^{+}21]. education [ACM06a]. EDV [Ano94c]. EDV-Benutzertreffens [Ano94c]. Edward [Che10]. Effect [DK06, LFS^{+}19]. Effective [MLAV10, RK01, SNN^{+}20, TMC09, Tsu95, BC19b, Cza13, JH97, KS15a]. Effects [SSE12]. efficacy [GScFM13]. Efficiency [KS96, MTU^{+}15, CZ96, MMU99, RS95]. Efficient [ADT14, Att96, BHW^{+}17, BGBP01, BCK^{+}09, BHLS^{+}95, BFG^{+}10, BGD12, Bru95, BDH^{+}95, BDH^{+}97, BMPZ94b, CVPS19, CAWL17, CFP96, DZ98a, DDN^{+}22, DGG^{+}12, FHPS94a, FHPS94b, FCS^{+}19, GGZ^{+}20, HBT95, HKT^{+}12, HT08, HC06, HLO^{+}16, KGK^{+}03, KD13, LSB^{+}18, LHCW05, LAD16, MDM17, MB12, MRB17, MKK21, NBK99, NWT21, OWO98, PGS^{+}13, RJMC93, RRBL01, RSC^{+}19, SPB^{+}17, SOYHDD19, TGBS05, WQKH20, WSN99, WWFT11, YPZC95, YT20, ZWHS95, ZLWW20, ZT20, BfDA94, BHW^{+}12, CLE^{+}20, CGH^{+}14, FM90, FNSW99, FHB^{+}13, HCL05, KKB^{+}21, KVGH11, LML^{+}19, LKL96, LZC^{+}20, LA06, MNYN21, MMDA19, Pan95b, PRS^{+}14, PSH^{+}20, PGPCK21, RR01, STA20, SOA11, TPD15, TDG13, YLC16, dCZG06, CRD99, THRZ99]. Efficiently [CC99, CCM^{+}06, PHA10]. effortless [ITT99]. eigenproblem [BV99, GG99]. eigensolvers [DR18]. Eigenvalue [DAK98, BSC99, THM^{+}94].

Eighth [ERS95, Sie94, IEE96b]. Eilean [CSS95]. einem [BL94]. Einflüß [Gra97]. Einführung [MS04]. einit [CLE^{+}20]. Einstein [ARYT17, KLM^{+}19, MBA21]. Einstein- [ARYT17]. Ejector [CCBPGA15]. elastic [PTG13]. elasticity [PTT94]. Elastodynamic [MAIVAH14]. electric [BALU95, Ano03]. electrical [Sil96]. electroabsorption [WWZ^{+}96]. electromagnetic [DSOF11, NZZ94, OMK09, WGG^{+}19]. electromagnetics [OGM^{+}16]. electron [ART17, JL18]. electron-molecule [ART17]. Electronic [GJN97]. Electronics [IEE95d]. Electrosoft [Sil96]. electrostatic [VDL^{+}15]. Element [BS21, DK20, KK19, MMD98, MS02b, OD01, OMK09, RHM^{+}17, SM02, VRS00, BB93, BCM^{+}16, Gra09, HMKV94, KME09, KEGM10, MGS^{+}15, MKK21, Nak05a, Nak05b, PTT94, PSV19, TOC18]. Elemental [PMvdG^{+}13]. elements [KB13]. Eliminating [DSG17]. elimination [ACMZR11]. elision [CLdJ^{+}15]. elliptic [AGIS94, PR94c]. ELLPACK [BBH12, MKP^{+}96]. ELLPACK-R [BBH12]. Else [Gei00]. elucidation [MK94]. embarrassingly [RGP22]. Embedded [TCM18, WZM17, YGH^{+}14, ADK22, ACJ12, CGK11, NEM17, TMW17, WCS^{+}13]. Embedding [FS97, SML17, SML19, MS96a]. Embodying [Ser97]. Emerging [WJA^{+}19, RMNM^{+}12]. Emission [Pat93, EZBA16]. emphasis [Bos96]. eMPI [MS96a]. eMPI/eMPICH [MS96a]. eMPICH [MS96a]. Empirical [SS94, VY02, KSTM20]. Employing [AGMJ06, GVF^{+}18, LB16]. emulation [Bae20, MS99b]. emulator [LTLC94]. enable [SPK^{+}12]. Enabled [Fos98, GSY^{+}13, LSMW11, Pan14, SSLMW10, ZL17, ZLP17, DS13, GLM^{+}08, HJBB14, KHBS19, KTF03, PSV19, RA09, SHHI01, SBK21, SR11, ZLS^{+}15]. Enabling [APBcF16, BGG^{+}15, CLSP07, DGB^{+}14, GBH14, GBH18, HJYC10, NPS12, TY14, ZPI06, BR04, MA09, SHHC18, WDR^{+}19]. encapsulation [DRUE12]. encoding [AAAA16, PGBF^{+}07, SM12]. endpoint [LLH^{+}14]. endpoints [DGB^{+}14, WK20]. energies [TKP15]. Energy [BPG94, CBB^{+}20, CBB^{+}21, EGR15, KFL05, LML^{+}19, RBAI17, SPB^{+}17, VW92, FKLB08, KN17, LRLG19, MNYN21, PTL^{+}16, TDG13]. Energy-Aware [EGR15]. Energy-Efficient [SPB^{+}17, LML^{+}19, MNYN21, TDG13]. Engine [Wal01a, NPP^{+}00a, Wal01b, WGG^{+}19]. Engineering [Ano98, BPG94, BP93, EGH^{+}14, IEE96h, KaM10, LSB15, LF^{+}93a, MS02a, MBS15, Nag05, SM07, Str94, DMW96, IEE94c, PW95, RMS^{+}18, Sil96, LF^{+}93a]. engineers [HW11]. Engines [SLJ^{+}14, HSW^{+}12, SHM^{+}12]. Engine^{T M} [OIS^{+}06]. English [Wil94]. Enhance [AR01]. Enhanced [Ano98, CDHL95, CDH^{+}95, FMSG17, KY10, PLR02, Saa94, BR95b, FE17a, FE17b, TSCS14]. enhancement [ARL^{+}94, Boi97]. Enhancements [BDG^{+}95, BCKP00, DM95b, DM95a]. Enhancing [BFIM99, CMZ99, FSC^{+}11, HMS^{+}19, IPG^{+}18, MVTP96, MSMC15, OFA^{+}15]. Ensemble [Cot97, Cot98, BY12, FH97]. Ensemble-Based [FH97]. ENSOLV [AMS94]. Entwicklung [Sei99]. Environment [BDGS93, BFG^{+}10, BFM97, BGL00, CHPP01, CTK01, DLB07, DI02, DHHW92, DHHW93a, DDL00, FTVB00, FWR^{+}95, GJN97, GL97a, HRSA97, KBA02, KKH03, KDL^{+}95b, KVH97, LC93, Lus00, MSOGR01, MM02, MFG^{+}08, MSS97, NJ01, Ong02, Rol94, SDN99, SGL^{+}00, SGHL01, TTP97, WL96a, ASAK19, ABG^{+}96, BDG^{+}92b, BDG^{+}94, BK96, BT96, CEF^{+}95, CLLASPDP99, DZ96, DL10, DHHW93b, EASS95, FMBM96, FB95, Fan98, Fra95, GBR97, GGH99, GPL^{+}96, GkLyCY97, HZ94, IJM^{+}05, IvdLH^{+}00, KCD^{+}97, Kat93, KDL^{+}95a, Kos95b, KFSS94, wL94, MSL12, MK97, NP94, PES99, PVKE01, PQ07, RNPM13, SSKF95, Sch93, SPK96, SBF94, SWYC94, Skj93, SSG95, TJD09, TSCS14, Tho94, WCC^{+}07, WL96b, WLC07, ZPLS96]. environmental [ANS95]. Environments [Ano95e, Ano01a, Bak98, BF98, DT94, GFB^{+}03, Laf01, Mat94, Mat95, MFC98, PS01a, RB01, SHH94b, SSSS97, SCL00, TAH^{+}01, ACGdT02, ARL^{+}94, ALR94, ADDR95, AMV94, Bon96, BFIM99, CDH^{+}94, CK99, DR94, DR95, EO15, HS93, HVSH95, LC07, LGMdRA^{+}19, MSP93, SS94, SHH94a, SAP16, TSS98, VB99, YS93, ZL96]. environments-the [CDH^{+}94]. EPS [GT94]. EPS-APS [GT94]. Epstein [BL95]. Epstein-Nesbet [BL95]. Equation [ES11, LZ97, SAS01, VRS00, DM12, LBB^{+}16, LYSS^{+}16, MS95, NP94, ÓN12, Ols95, Pri14, iSYS12, SSB^{+}16, YSVM^{+}16, YSMA^{+}17]. Equations [And98, BG95, GK10, Huc96, LLY93, lLLmH^{+}21, MFTB95, MKK21, ORA12, ZB97, BHW^{+}12, Che99, DFSW19, IM95, JK10, Jou94, MPS20, MM11, NF94, RBB15, SP11, SMSW06, ZZG^{+}14, dH94]. Equi [LTRA02]. Equi-Join [LTRA02]. equivalencing [LLG12]. Era [ABB^{+}10, CZG^{+}08, CGKM11, EdS08]. Erratum [Ano01b, HF14b, Wal94b]. Error [DFC^{+}07, SSLMW10, HPS^{+}12, HPS^{+}13]. Errors [FCLG07, DPFT19, SD16]. Erweiterung [GBR97]. ESA [Whi94]. ESBMC [MdSAS^{+}18]. ESBMC-GPU [MdSAS^{+}18]. Espoo [RWD09]. ESPRIT [CDH^{+}94]. Estimation [GK10, TSN21, WZM17, WQKH20, YNJS21, AMHC11, CCU95, GB94, JMdVG^{+}17, KS13, ZWHS95]. Estuarine [LRQ01]. Ethernet [CC00a, Fin97, HcF05, KYL03, KYL05, OF00, PFG97]. EU [Ano03]. Eugene [MCdS^{+}08]. Euler [DLR94, IDD94]. Euler/Navier [DLR94, IDD94]. EURO [HAM95b, BFMR96, HAM95b, BFMR96]. Euro-Par [BFMR96, HAM95b, BFMR96]. Euromicro [IEE95h, IEE96g]. EuroMPI [CDND11, KGRD10, TBD12, GT19, TB14, TH20]. EuroMPI/USA [GT19]. EUROPE [LCHS96, Ano92, Ano93f, Ano93g, Ano94g, Tou96]. European [AD98, Ano94i, BR95a, BDLS96, BC00, BDW97, CHD07, CHD09, CD01, CDND11, DKD05, DLM99, DKP00, DLO03, KGRD10, Kra02, KKD04, LKD08, MTWD06, RWD09, TBD12, WPH94, DHK97]. EuroPVM [BDLS96, OL05, DKD07, MTW07]. EUROPVM/MPI [OL05, DKD07, MTW07]. EuroPVMMPI [KKDV03]. EUROSIM [BH95, DSZ94, BH95]. Eurospace [Tou96]. Eurospace-Ada-Europe [Tou96]. Evaluate [MW98]. Evaluating [BWV^{+}12, FVLS15, FL21, FST98a, GFD03, GFD05, GGCGO01, GB96, HWW97, LRG^{+}16, LH95, SSSS97, ZSnH01, GScFM13, LTLC94, TG09, ZLZ^{+}11]. Evaluation [ATM01, BF98, BIC^{+}10, BFM97, BEG^{+}10, BB18, CLP^{+}99, DI02, FST98b, FSSD17, Han98, JCH^{+}08, KS96, KK19, KK02b, KSS00, LGCH99, LNK^{+}15, LZ97, kL11, LVP04, MH01, MGC12, NNON00, OTK15, OM96, Pan14, Par93, RB01, SWHP05, SCP97, SEF^{+}16, SBF^{+}04, SM02, Sou01, SJK^{+}17a, SJK^{+}17b, TOTH99, TSB02, TSB03, TTSY00, UMK97, VY02, AB13, BBG^{+}14, BBH… 13a, BMG07, CdOO^{+}20, CB11, DBB^{+}16, DWS^{+}21, HPR^{+}95, HHK^{+}19, HASnP00, HPS95, IM94, JC17, JMdVG^{+}17, KC19, KKB^{+}21, LV12, LNW^{+}12, MKP^{+}96, MKP22, MM03, MT96, MMH99, MH21, NN95, PSK08, RLFdS13, SL94b, SWS^{+}12, SWYC94, SFSV13, TSP95, THM^{+}94, TMPJ01, Wor96, YWO95, YS93, ZHK06]. Evaluations [KNH^{+}18, MM14]. Event [KKV01, NSLV16, THS^{+}15, WM01, WMC^{+}18, FSG19a, FSG19b]. Event-Based [NSLV16]. event-driven [FSG19a, FSG19b]. events [HHK^{+}19]. everything [CCM^{+}06]. everything-shared [CCM^{+}06]. Evolution [Mat01a, PS01a, RBB17, SSL97, SGDM94, DFSW19, GS93, SSD^{+}94]. Evolutionary [B^{+}05, DSM94, Rag96]. Evolving [Bad16, ER12, MdSC09]. Ewing [Ano95c, Ano99c, Ano99d, Ano00a, Ano00b]. EWOMP’99 [BC00]. Exact [dOSMM^{+}16]. examine [LFS^{+}19]. ExaMPI [SBG20]. Example [Che10, SK10, NB96, Pat93]. Examples [Cre16, Edd18, Mat16]. Exascale [Bad16, SBG20, BBB^{+}20, LV12, LSG12, LGM^{+}20, RPS19]. Exception [FMSG17]. exchange [MMM13, Pan95a]. excluded [BHW^{+}12]. executable [WMP14]. Execution [AHD12, BME02, DT17, FC05, FM09, GR07, KGK^{+}03, MK17, Mar05, MFG^{+}08, MAGR01, Ney00, STY99, SAP16, BLVB18, EPML99, Mor95, PSB^{+}19, SMAC08, TNIB17, TSY99, TSY00, UGT09]. Executions [GAML01]. exhaustive [CPKG17]. Exhibition [HS95a, GH94, LCHS96]. Existing [CB00]. EXOCHI [WCC^{+}07]. Expand [CGC^{+}02]. Expanding [LA02]. expected [CAHT17]. Experience [BCP^{+}97, BT96, CP98, PS01a, Tou00, AMS94, BC19b, CARB10, KJA^{+}93, RSC^{+}15]. Experiences [AHP01, BFZ97, CMV^{+}94, CLLASPDP99, GLN^{+}08, GS91a, GSI97, GB96, GL95d, ITT02, JR10, KS97, LBG^{+}20, Mar02, TGEM09, ZPLS96, ZKRA14, AL92, BBB^{+}22, CCF^{+}94, Sch94, SGDM94, BDG^{+}93b]. Experiment [Luo99]. Experimental [BIL99, BIC05, BB18, EGC02, Ser97, UMK97]. Experiments [BPMN97, Coe94, LGM00, OS97, RR00, ZB97, RHG^{+}96, HAJK01]. Expert [BPG94]. experts [EO15]. ExpEther [NMS^{+}14]. Explicit [BHV12, GFPG12, SGHL01, CdOO^{+}20, DS22, KW20, LC97b]. Explicitly [Mai12, SYR^{+}09]. exploit [ZPI06]. Exploitation [GGL^{+}08, GAM^{+}02, BK11, GAM^{+}00]. Exploiting [Add01, AML^{+}99, Bri10, FKLB08, HEHC09, KFL05, LWKA15, LFW20, NAAL01, VGP^{+}19, Nob08, SWCB20, THH^{+}05]. Exploration [AMuHK15, HSO^{+}21, MZLS20, OFA^{+}15, ABDP15, GE95, GE96, PDY14]. Explorations [BGG^{+}15]. Exploring [CPM^{+}18, IFA^{+}16, IMS16, LGMdRA^{+}19, MBKM12, MTU^{+}15]. Expose [SAL^{+}17]. Exposing [SD16]. Exposition [IEE95d, LF^{+}93a]. EXPRESS [KS96, Ahm97, FK94, LH95, SHH94a, SHH94b]. expressing [HLK^{+}20]. Expression [BN12, GDM18, KH15, Per21, Sur95a]. Expressions [VZT^{+}19, SFLD15]. expressive [Trä12a, YLC16]. Extend [DFA^{+}09]. Extended [BR02, Rót19, HTA08, SS99]. Extending [ABB^{+}10, BCC^{+}00a, BCC^{+}00b, BDB^{+}13, CS96, CG99a, KDT^{+}12, LMRG14, Mar03, OFA^{+}15, RGDML16, SDV^{+}95, TMTP96, CG96, GGHL^{+}96, KSC^{+}19, LRG^{+}16]. Extensible [BL97, GS94]. Extension [AELGE16, BGR97a, CSAGR98, VAT95, Hum95, JH97, SG14, SC95, ZT17, GBR97]. Extensions [Fis01, GOM^{+}01, GHLL^{+}98, HVA^{+}16, HE15, DPSD08, HP05, Kat93, VLCM^{+}20, ZCBD22, Ano99c, Ano99d]. Extent [kL11]. Extent-Based [kL11]. exterior [HMKV94]. external [BBB^{+}94]. Extraction [CBL10, HLO^{+}16, RTN21, dAT17]. Extreme [MdSC09, ZKRA14]. Extreme-scale [ZKRA14]. eyes [Str94].

F [FHPS94b, FHP^{+}94]. F90 [DP94]. Fabric [ZL17, ZL18]. face [HDDG09]. Faces [Gro12]. facilitate [PKB06]. Facilitating [MC99, ZLL^{+}12, ESB13]. Facilities [MMH98, MN91]. Facility [KG96, SHTS01, KZCS96, LHCT96].

Factorisation [BB18]. Factorization [KF16, OPJ^{+}19, AZ95, BSvdG91, BRS92, DG95, KBP16, WLC07]. Factorizations [TD98, LC97b]. Fail [LFS92, LFS93a, LFS93b]. Fail-safe [LFS92, LFS93a, LFS93b]. Failure [BBH… 13a, CRGM14, SRS^{+}19, BBH^{+}13b, CGH^{+}14, BDB^{+}13]. failure-aware [CGH^{+}14]. failures [JS13]. Faithful [KLR16]. FALCON [HCC^{+}20]. FALCON-X [HCC^{+}20]. Fall [Gra97]. false [JE95]. FAME [lLLmH^{+}21]. family [AVA^{+}16]. far [KKB^{+}21]. farming [Str94]. Fast [ADGA20, Ben01, BHS^{+}02, BDA^{+}18, BBH12, CS14, DMK19, DFN12, EM02, HMKG19, Hog13, Hol95, JFGRF12, JMdVG^{+}17, KK19, LYIP19, lLLmH^{+}21, PSHL11, PR94c, PBC^{+}01, RB01, SE02, SS09, STY99, SR11, TPLY18, UP01, WTR03, AiIS^{+}21, KKB^{+}21, Lan09, LCL^{+}12, NYNT12, STA20, TDG13, YULMTS^{+}17, YLZ13, YBZL03, ZA14, AAB^{+}17, DBLG11, PFG97]. Faster [Tsu12, ZG95a, ZG96]. Fat [Zah12]. Fat-tree [Zah12]. FATCOP [CF01]. Fault [BBC^{+}02, BCH^{+}03, BHK^{+}06, CF01, CFDL01, FBD01a, FBVD02, FD02a, FD04, GFB^{+}03, GKP97, GJR09, GL04, Gua16, IEE95c, JSH^{+}05, LMRG14, LGM^{+}20, LNLE00, dLR04, MSF00, RPM^{+}08, TS12a, WC09, Wil93, BCH^{+}08, CLE^{+}20, FBD01b, FD02b, HG12, LRG^{+}16, LMG17, LS08, MB21, PKD95, RGP22, SG05, WDR^{+}19, ZHK06, FD00]. Fault-Management [GJR09]. fault-tolerance [CLE^{+}20, WDR^{+}19]. Fault-Tolerant [BHK^{+}06, FD04, GFB^{+}03, IEE95c, JSH^{+}05, LMG17, LS08, MB21]. Faults [LAdS^{+}15]. FCRC [ACM96b]. FD [And98]. FD-TD [And98]. FDDI [LC93]. FDTD [DSOF11, VM94, WGG^{+}19]. Fe [Old02, RV00, BJS99]. feasibility [KBG16]. Feature [Qu95, GDEBC20, RTN21, ZWL^{+}17]. Feature-driven [Qu95]. Features [GLT99, GLT00b, GLT00a, GLT12, KAHS96, Ano00a, BPJ22, CMZ99, CRD99, IMS16, LSB^{+}20, WKS96, ZKRA14, dAT17]. February [Ano95d, GE95, GE96, IEE93a, IEE94a, IEE97c]. FEM [EVMP20, GEW98]. FEM-Systeme [GEW98]. Fermi [SP11, WKP11]. fermions [GM18]. FETI [KLR^{+}15]. few [NS16]. few-body [NS16]. Feynman [NS16]. FFT [DMK19, DALD18, GB98, JKM^{+}17, NSM12, SH14, WJB14]. FFT-Based [WJB14]. FFTs [EFR^{+}05]. FFTW [KT10]. FHP [BMS94a]. Fibonacci [GFJT19]. Field [KNT02, Goe02, KKB^{+}21, LK20, RS22, TKP15]. fields [BALU95, RSBT95]. Fifth [DKM^{+}92, HK93, IEE96f, SM07, IEE95c]. filamentary [YPA94]. File [BIC^{+}10, CGC^{+}02, LRT07, kLCCW07, kL11, PLR02, RK01, TSS00b, Tsu07, WTR03, DL10, LL95, SBQZ14, iSYS12]. File-I [PLR02, RK01]. File-I/O [PLR02, RK01]. film [SL00]. Filter [FDG19, BY12, CCU95]. find [GDMME22]. Finding [FCLG07, GÁVRRL17, PCS94]. Fine [AZG17, BBG^{+}10, JCP15, SFL^{+}94, TCM18, YSS^{+}17, BK11, KW14, LZHY19]. Fine-Grain [AZG17, JCP15, SFL^{+}94, BK11, KW14]. Fine-Grained [BBG^{+}10, TCM18, YSS^{+}17, LZHY19]. Finite [DK20, DFN12, KK19, MMD98, MS02b, MAIVAH14, MKK21, OD01, OMK09, Pri14, RHM^{+}17, SM02, UZC^{+}12, VM94, VRS00, BB93, CdOO^{+}20, DS22, GÓ19, Gra09, GFPG12, HE13, HMKV94, KME09, KEGM10, KB13, Nak05a, Nak05b, NZZ94, NB96, PSV19, Ram07, TOC18]. Finite-Difference [UZC^{+}12, VM94, CdOO^{+}20, HE13, NZZ94, Ram07]. Finite-Element [MS02b, MKK21, BB93, KME09, KEGM10, Nak05a, Nak05b]. Finland [RWD09]. Fire [JML01, SJ02]. Firedrake [RHM^{+}17]. First [AGH^{+}95, BCD96, BC00, CH96, Dem96, DFN12, DW94, Gat95, HAM95b, Kum94, Nar95, PBPT95, SSP^{+}94, USE94, AH95, BS94, GM18, MMDA19, PTMF18, PBPT95]. Fix [DLV16]. fixed [PSV19]. fixed-grid [PSV19]. FLAME [VBLvdG08]. flat [Nak05b]. Flattening [THRZ99]. flavors [GM18]. FlexCL [LWZ18]. Flexibility [KK02b]. Flexible [CS14, GR95, GBS^{+}07, SHPT00, CARB10, DGB^{+}14, GAM^{+}00, HGX^{+}22, HC08]. Flink [KWEF18]. FlinkCL [CLOL18]. flip [KO14, Kom15]. Floating [LWSB19]. Floating-Point [LWSB19]. Florida [ACM98b]. Flow [BHW^{+}17, BGD12, CGZQ13, CCBPGA15, FM09, MK17, Pat93, AMS94, AFST95, EP96, ED94, HK94, HTHD99, HHSM19, JAT97, LL16, MBKM12, MH18, Ols95, PTT94, RM99, SCC95, SU96, TS12b, TOC18, TGS^{+}20]. Flow-Based [BHW^{+}17]. Flows [GAP97, BCM^{+}16, BTC^{+}17, DS22, Heb93, LLG12]. flowshop [CB11]. Fluid [DFMD94, GAP97, JFY00, SZBS95a, TDBEE11, TGEM09, ALR94, ATL^{+}12, AGMJ06, BvdB94, BHS18, Bil95, HVSC11, MRRP11, PBK99, SPE95, SZBS95b, WPH94, XR21]. fluid-particulate [ATL^{+}12]. fluids [HK94, WB96]. Flux [QRMG96, QRG95]. Fly [WMC^{+}18, KSJ14, THRZ99, BCAD06, BADC07]. FM [LC97a]. FMA [LO96]. Fock [MMDA19, CBHH94]. Focus [Cla98, CFF19]. foolish [Rol08a]. Footprint [CBB^{+}20, TS12b]. force [Goe02, RS22]. forcing [JRG21]. Forecast [AHP01]. forecasting [Bjo95, KOS^{+}95a]. Forest [JML01, MPZ21, NCKB12]. ForestGOMP [BFG^{+}10]. Foreword [CHD09, SBG20]. FORGE [WCVR96]. Fork [BGD12, SML17, SML19]. Fork-Join [BGD12, SML17, SML19]. form [NCB^{+}12, NCB^{+}17]. Formal [BG94a, BdS07, GKS^{+}11, GB98, LPD^{+}11, PGK^{+}10, VVD^{+}09, BG94c, SZ11]. Formalizing [FGRT00]. Format [BBH12, MDM17, CBIGL19]. Forschung [Ano94c]. Fortran [Ano97, Ben95, Bra97, GBR15, TOC18, AC17, Ano98, AS14, BW12, BC19b, DZ98b, Don06, GML^{+}16, HE13, HH14, HZ99, KaM10, Kuh98, KLM^{+}19, LC97b, LCC^{+}03, MWO95, iSYS12, SM03, SMCH15, SC19, TBG^{+}02, Wal02, YBMCB14, YSVM^{+}16, YSMA^{+}17, vHKS94]. Fortran/PVM [MWO95]. Forum [Str94]. Forward [RMNM^{+}12, BDB^{+}13]. forwarding [CXB^{+}12]. foster [SM12]. Foundation [Gei01]. Foundations [KSTM20]. four [GSMK17, MGG05]. four-atom [MGG05]. four-particle [GSMK17]. Fourier [DBLG11, BCM^{+}16, YW21]. Fourteenth [IEE95b]. Fourth [Ano89, IEE93d, IEE95k, Sie92a, Sie92b, Ano94i, IEE96g]. FPGA [KNH^{+}18, LLVM21a, LLVM21b, MKP22, MTU^{+}15, MZLS20, PWP^{+}16, PGF18, RGB^{+}18, WTTH17, WHMO19, WTS19]. FPGA-based [WTS19]. FPGA-Platform [WTTH17]. FPGAs [AJYH18, CJPC19, JCP^{+}20, LLVM21a, LLVM21b, LWZ18, MC17, MKP22, OFA^{+}15, PGS^{+}13, WZHZ16, Röh00]. fractal [Wu99]. fragment [KS15a]. fragments [OA17]. frame [MNYN21]. Framework [Ben18, DGMS93, FC05, GGCGO01, GR07, GDDM17, HDW21, MGL^{+}17, NSZS13, PWPD19, PMvdG^{+}13, RBP^{+}21, SSB^{+}05, SSAS12, Sun90a, Sun90b, VT19, WZHZ16, Ano93c, BA06, BLVB18, BR04, BAG17, EFR^{+}05, FLMR17, GM13, HDZ^{+}20, JCP^{+}20, KKM15, KJJ^{+}16, KKJ^{+}08, KH10, LLVM21a, LLVM21b, LME09, LGG16, LCMG17, LS08, MW21, PTL^{+}16, RSC^{+}15, SL00, TDB00, XJR21, YLC16, YWTC15, ZT17, dAT17]. Frameworks [OP10, ASS^{+}17, KDSO12]. France [ACM90, BR95a, BFMR96, CHD07, DE91, FR95, JPTE94, MCdS^{+}08, VW92, YH96, GA96, IEE94c]. Francisco [BBG^{+}95, IEE93a, IEE94g]. Frankfurt [Tou96]. Frankfurt/Main [Tou96]. Fredericton [BG91]. Free [DK20, KK19, PKYW95, CP15, MKK21, SOA11, Zah12]. freedom [KTJT03]. Frequency [IEE94e, SdR^{+}21]. friendly [SVC^{+}11]. Frontiers [ACM06b, IEE94a, IEE96c, Sie92a, Sie92b, Sie92a]. Frontiers’95 [IEE94a]. Frontiers’96 [IEE96c]. FSI [HAA^{+}11]. FT [FD00, LNLE00, WTS19]. FT-MPI [FD00]. Fujitsu [Ano98, AKL99, BHS^{+}02, SWJ95, SH96]. full [CFF19]. full-orbit [CFF19]. Fully [GA96, ZL17, SSB^{+}16, VLCM^{+}20]. Function [AGS97, Bri02, HHS18, MCP17, Rót19, RB01, SW12, HE15, JMdVG^{+}17, KRC17]. Functional [ACM90, AJF16, CNM11, NW98, Ser97, CBHH94, EP96, HLK^{+}20, HSE^{+}17, SFLD15, WZWS08]. functionality [BFIM99]. functionally [PSV19]. Functions [BKGS02, Brü12, Hat98, MDM17, CdGM96, HWX^{+}13, PNV01]. Fundamentals [Wal96a]. fused [TW12]. fusing [BAC20]. Fusion [FHK01, FMFM15, LK20, PKE^{+}10]. fusions [FFM11]. Futhark [HSE^{+}17]. Future [Dar01, IEE93d, Mat00a, BDG^{+}93b, FK94, FHP^{+}95, Gei94, RPS19, Sni18]. Futures [Kuh98]. fuzzing [LLCD15]. Fuzzy [MDM17, TVCB18].

G [OPM06]. G2 [Cot04, KTF03, OPM06]. GA [Ara95]. GAIN [ARYT17]. GAIN-MPI [ARYT17]. Gains [CMM03]. Galerkin [CF19, KK19]. Gallipoli [Ano93b]. GAMMA [CC00a]. Gap [AAB^{+}17, ASS^{+}17]. Garbage [GTS^{+}15]. Gas [BMS94b, BBK^{+}94, BMS94a]. GASPI [SIC^{+}19]. gather [MTK16, Pro21]. gauge [BW12]. Gauss [BG95, LM99, Ols95]. GCel [SHH94a, SHH94b]. GECCO [B^{+}05]. Geist [Ano95b, NMC95]. gem5 [PHO^{+}15]. gem5-gpu [PHO^{+}15]. Gemini [SWS^{+}12]. gems [Fer04, mH12, Ngu08, PF05]. Gene [GDM18, PCS94, AAC^{+}05, BGH^{+}05, EFR^{+}05, KMH^{+}14, LM13, MV17, MSW^{+}05]. gene-finding [PCS94]. Gene/L [AAC^{+}05, BGH^{+}05, EFR^{+}05, MSW^{+}05]. Gene/Q [KMH^{+}14, LM13, MV17]. General [AJYH18, Che10, IH04, MW98, SK10, SZBS95a, Sun94a, TPV20, ABDP15, ADLL03a, ADLL03b, CBM^{+}08, FLD96, KPNM16, PF05, RSBT95, SSD^{+}20, SZBS95b, SMSW06, YPA94]. General-Purpose [AJYH18, Che10, SK10, ABDP15, CBM^{+}08, KPNM16, PF05]. generalised [TGS^{+}20]. Generalized [DFKS01, FKS96, BSC99, SD99, van93]. Generating [AZG17, CGL^{+}93, ER12, IJM^{+}05, PKB^{+}16, SFLD15]. Generation [AB93a, CC17, FAFD15, Gei98, GTH96, GSYT21, HT08, JFY00, LTDD14, RGD13, SSK^{+}18, SSB^{+}17, TGBS05, VPS17, AB93b, CPKG17, CPR^{+}95, DCD^{+}14, DWM12, EYP^{+}20, KHS12, KPL^{+}12, KH10, LCH^{+}22, MMDA19, RBC20, SP11, TGKL19, WKS96, WMP14, ZKRA14]. generational [WK08a, WK08b, WK08c]. generative [MAS06]. generator [Lan09, Stp20, TNIB17, YL09]. generators [CCS19]. Generic [ARS89, AKL99, GB98, BAS13, GM13, ZT17]. Genetic [FTVB00, MTSS94, MSCW95, PB12, TGKL19, WKS96, Wal01a, WHDB05, AB13, BB95a, FSTG99, HPLT99, JPL22, RJC95, Wal01b, B^{+}05]. genetics [LM99]. Geneva [IEE97b]. genomic [SdM10]. genomics [CJPC19]. GeoComputation [Abr96, Abr96]. GeoFEM [NO02b, NO02a, Nak03]. geomechanics [BJS99]. Geometric [DDP^{+}19, TK19, VGP^{+}19]. geometrical [FMS15]. Geometry [STK08, Hol95, STT96]. geophysical [Has95]. Georeferring [GCGS98]. Georgia [USE00, UCW95]. German [EGH99, GBR97, Gra97, GEW98, Sei99, Wer95]. Germany [BDLS96, GH94, KGRD10, MTWD06, MdSC09, PSB^{+}94, Sch93, Tou96, Ano93a, BPG94, Cal94, GHH^{+}93, WPH94]. Gesellschaft [Ano94c]. get [Str94]. Getting [Nob08]. GF100 [WKP11]. gHull [GCN^{+}13]. GHz [Ano03]. Gibbs [TKP15]. Gigabit [CC00a, HcF05, EGH99, OF00]. Giganet [GT01, Trä02b, bT01a]. GIS [CFPS95, CCSM97]. Give [DZ98b]. glass [JRG21]. Glenda [SBF94, Bic95]. Global [BSG00, DSS00, Pan95a, Ros13, SHTS01, STK08, SWH15, TTP97, HWS09, HCL05, HEHC09, LF^{+}93a, Str94, Wan02, YLZ13, Zah12, ZWHS95]. Globally [BHS^{+}02]. GLUE [Rab98]. GMRES [dH94]. Gmunden [Vol93]. GNU [YSMA^{+}17]. go [KC94]. good [Mat03]. Göttingen [Ano94c]. GP [LRBG15]. GP-GPUs [LRBG15]. GPFS [AHP01, BIC^{+}10, PTH^{+}01a, PTH^{+}01b]. GPGPU [AAB^{+}16, ASB18, BGG^{+}15, CVPS19, CPM^{+}18, HA11, HCZ16, JKN^{+}13, LME09, LDJK13, LCY19, LYZ13, MBKM12, PTG13, TWLL19, TY14, YZ14, YNJS21, YEG^{+}13]. GPGPUs [CS19, JMdVG^{+}17, LSB15]. gprMax [WGG^{+}19]. gprof [GJLT11]. GPU [Che10, KA13, SPB^{+}17, AKL16, ADGA20, AFG21, AHHP17, BDP^{+}10, BR12, BCD^{+}12, BCD^{+}15, BBD^{+}20, BTC^{+}17, BMS19, BWV^{+}12, BBH12, CLOL18, CBYG18, CCBPGA15, DSU20, DF17, DS22, DS16, DK13, DALD18, DSOF11, DWL^{+}10, DWL^{+}12, EBB^{+}20, ER12, FA18, Fer04, FFM11, FSSD17, GCN^{+}13, HSO^{+}21, HVA^{+}16, HCC^{+}20, HSE^{+}17, HDW21, HK09, HK10, HZG08, mH12, JDB^{+}14, JLS^{+}14, JR13, JNL^{+}15, JJPL17, JPT14, KDSO12, Kha13, KSL^{+}12, KPL^{+}12, KF16, KI17, KPNM16, KEGM10, KO14, KNH^{+}18, KMM15, LWSB19, LV12, Lee12, LRG14, LLC13, LML^{+}19, LW20, LAD16, LYGG20, MMO^{+}16, MPS20, MPZ21, MdSAS^{+}18, MGL^{+}17, NRdA^{+}20, Ngu08, NWT21, NMS^{+}14, NSM12, OFA^{+}15, Pan14, PDY14, PGdCJ^{+}18, PF05, PS19b, Pri14, RSC^{+}15, RS19, RBC20, RMNM^{+}12, Sai10, SK10, SdM10, dOSMM^{+}16, iSYS12, SS09, SNN^{+}19, SSD^{+}20, SCSL12, SIRP17]. GPU [SBK21, SAP16, SYL19, SD16, SSB^{+}17, SKM15, SKB^{+}14, SG14, TBB12, TS12b, TMT^{+}20, TPV20, VZT^{+}19, VT19, WZM17, WJA^{+}19, WGG^{+}19, WKP11, WYZ^{+}19, XJR21, XR21, YULMTS^{+}17, YHL11, YCL14, YSS^{+}17, YSS^{+}19, ZJHS20, ZGNZ22, ZRQA11, ZZG^{+}14, ARYT17, PHO^{+}15]. GPU-Accelerated [KA13, KF16, SCSL12, PGdCJ^{+}18]. GPU-Aware [Pan14, FA18]. GPU-based [MMO^{+}16, SS09]. GPU-clusters [NWT21]. GPU-code [EZBA16]. GPU-enabled [SBK21]. GPU-Job [PS19b]. GPU-programming [HSE^{+}17]. GPU-Resident [JDB^{+}14]. GPUDirect [OGM^{+}16, YWCF15]. GPUMixer [LWSB19]. GPUMP [ZC10]. GPUrpc [IFA^{+}16]. GPUs [AJYH18, ABG20, BLVB18, BY12, BC19b, BDA^{+}18, CJPC19, CPKG17, DS13, DS16, GNP19, GML^{+}16, GFPG12, GPC^{+}17, GM18, HTJ^{+}16, HLP10, HP11, HLP11, Hos12, HWW21, IFA^{+}16, JKM^{+}17, JAK17, KGB^{+}09, KKM15, KKLL11, KC19, KVGH11, KW20, LWKA15, LBH12, LRBG15, MA09, MYL21, NS20, ÓN12, OIH10, PP16, PSV19, PB12, SHLM14, SNN^{+}20, STH22, SDB^{+}16, SKK^{+}12, TPK^{+}19, Tsu12, VLMPS^{+}18, VY15, WRSY16, WQKH20, WJ12, WJB14, YLZ13, YSWY14, ZLWW20, ZC10, Zho21, ZZZ^{+}15]. gpuSPHASE [WMRR17, WRMR19]. GPUVerify [BCD^{+}12]. GQ [RFG^{+}00]. gQoS [LYGG20]. GRACE [YKI^{+}96, ZRQA11]. GRADE [DDL00]. graded [PSV19]. Gradient [BG95, GFPG12, SSK^{+}18, BAC20, KN17, MM92, Ols95]. Grain [AZG17, IOK00, KOI01, MJPB16, NIO^{+}02, NIO^{+}03, BK11, JCP15, KW14, SFL^{+}94].

Grained

[ADRCT98, BBG^{+}10, LGM00, TCM18, YSS^{+}17, HDZ^{+}20, Heb93, LZHY19, RJC95]. GRAM [HDW21]. Grammatical [RBB17]. Grand [DGMJ93, Ten95, BDG^{+}92c]. Graph [BHW^{+}17, CDT05, CTBT21, DW02, MM14, NPS12, PPR01, STV97, Zho21, HLP10, HKOO11, MMAH20, PP16, PD11, RJH^{+}20]. Graph-Based [NPS12]. Graph-Partitioning [STV97]. Graphic [HJBB14]. Graphical [BDG^{+}91b, DDL00, BDG^{+}92a, KCD^{+}97, KFSS94, SSKF95, VDL^{+}15]. Graphics [JPL22, KS15b, LSVMW08, LSMW11, SLJ^{+}14, SSLMW10, vdLJR11, ABDP15, BHS18, CBM^{+}08, DBLG11, Fer04, GKL95, HTA08, HSW^{+}12, KFA96, KY10, KB21, KME09, LHLK10, MSZG17, PF05, SHM^{+}12, SR11, WWFT11, ZLS^{+}15, MSML10]. graphics-scalable [GKL95]. Graphite [MMAH20]. Graphs [LGM00, OP10, PGF18, VZT^{+}19, EP96, MC99, MJPB16]. Gravitational [ZSK15, KM10]. Greece [CD01, CDND11, SM07, TG94]. green [PTL^{+}16, LWKA15]. Grenoble [JPTE94]. Grid [AB93a, CGB^{+}10, CLL03, DPP01, Fos98, KT02, Laf01, Liv00, MRB17, PLK^{+}04, Rei01, SSK^{+}18, TGEM09, AMKM20, AB93b, Eng00, GLM^{+}08, KRKS11, KTXP21, PSV19, WYLC12, AASB08, BR04, CCHW03, DKD08, FC05, GFB^{+}03, GL02, KTF03, KGK^{+}03, KSSS07, LC07, LS08, NSBR07, RPM^{+}08, RTRG^{+}07, SHTS01]. Grid-Adaptive [KT02, KTXP21]. Grid-Enabled [Fos98, GLM^{+}08, KTF03]. Grids [NO02b, ACH^{+}11, CC10, KBG^{+}09, NO02a, NB96, TK19, XJR21, BBH^{+}06, GR07, Ram07, SN01]. GROMACS [BvdSvD95]. Gropp [Ano95c, Ano99c, Ano99d, Ano00a, Ano00b]. Gross [LBB^{+}16, LYSS^{+}16, SSB^{+}16, YSVM^{+}16, YSMA^{+}17]. Ground [HTHD99, NS16]. Groundwater [MMD98, AFST95, EGDK92]. Group [AD98, Ano98, Ara95, ACDR94, CHD07, CHD09, CD01, CDND11, DKD05, DLM99, DKP00, GN95, KGRD10, Kra02, KKD04, LKD08, MC94, MTWD06, RWD09, TBD12, UMK97, WQKH20, BDW97, DLO03, MMU99]. grouping [WPL95]. Groups [GOM^{+}01]. Grover [LYZ13]. Growth [PKYW95, BB95a]. GTS [PKE^{+}10]. Guest [AM07, GSA08, GT19]. GUI [VGS14]. GUI-awareness [VGS14]. guidance [SDJ17]. Guide [Ano12, D^{+}91, GBD^{+}94, Lad04, Nov95, NMC95, Per96, Ano95b, BDG^{+}91a, McK94]. Guided [FDG19]. Guideline [Trä12b]. Guidelines [TGT10]. GVirtuS [MGL^{+}17].

Hack [DLV16]. Hadoop [LSM^{+}18]. Hague [Ano93f]. Halide [RKBA^{+}13]. halo [BBW19]. halo-swapping [BBW19]. Hamburg [PSB^{+}94]. Hamiltonian [ART17]. Handling [DFC^{+}07, FMSG17, LSB15, LGM00, RC97, FFFC99, LNW^{+}12, THRZ99]. Hands [KmWH10]. Hands-on [KmWH10]. Harbor [BBC^{+}00]. Hardware [BGG^{+}15, BWW^{+}12, Brü12, BCKP00, CDPM03, DW02, EADT19, FGL^{+}20, GJMM18, HSP^{+}13, KF16, LSMW11, MFC98, PSM^{+}14, PKB^{+}16, SSK^{+}18, SSLMW10, ZGNZ22, vdLJR11, ER12, GGL^{+}08, PMZM16, Rab99, RS21, SBG^{+}12, SH94, SWS^{+}12, YÁJG^{+}15, ZLS^{+}15]. Hardware-Based [CDPM03]. Hardware-oblivious [HSP^{+}13]. harmonic [GSMK17]. Harness [EBKG01, MS99b, PL96, FBD01a, FBD01b, FBVD02, FD02a, FD02b, MSF00, Gei98]. HARP [FDG19]. Harrogate [CJNW95]. Hartree [CBHH94, MMDA19]. HASEonGPU [EZBA16]. Haskell [WO97]. Hate [Dan12]. Hawaii [ERS95, ERS96, HS94, MMH93, ZL96].

HCA [KBG16]. HDL [Kat93, KMK16]. HDMR [KD12]. Heading [Sch99]. Heaps [GFJT19]. Heat [SAS01, NP94, iSYS12]. Hector [RFRH96, RRG^{+}99]. Heijen [Van95]. held [AGH^{+}95, GA96, JB96, KG93, MMH93, Old02, R^{+}92, SPH95, TG94]. Helios [SPK96]. Helmholtz [HMKV94]. Helps [Stp02]. HeNCE [BDG^{+}92a, BDG^{+}92b, BDG^{+}93a, BDG^{+}94]. Hénon [JPT14]. Herzliya [IEE96h]. HeSSE [MRV00]. Heterogeneous [ABB^{+}10, BDG^{+}93a, BDGS93, BL95, BCP^{+}97, BGR97b, BCKP00, CMMR12, CLOL18, CLBS17, DKB20, DGMS93, DGMJ93, FDG97a, FDG97b, FLD98, Fos98, GS91b, GDDM17, HSO^{+}21, IEE93f, KR09, KCR^{+}17, LC93, LSB^{+}18, MRV00, MM01, MM02, NTR16, OPJ^{+}19, PD98, PHO^{+}15, RVKP19, SM19, SMS00, SGS10, TQDL01, VLO^{+}08, ACGdT02, ADB94, ADDR95, AMV94, BDG^{+}92c, BDG^{+}94, BALU95, BRR99, BAG17, CCM12, CFPS95, FMBM96, GKZ12, GCN^{+}10, GDEBC20, GKCF13, HHS18, HK94, IPG^{+}18, KSG13, KSL^{+}12, Kos95b, KSS^{+}18, LBG^{+}20, LCL^{+}12, LR06a, Lee12, Mai12, MSL12, MM03, NP94, NEM17, Pen95, PSB^{+}19, RCFS96, RVKP18, SCJH19, Skj93, Smi93b, Sun94b, Sun95, TBB12, TMW17, TKP15, TDG13, VB99, VGP^{+}19, WCC^{+}07, WZW21, YST08, YSL^{+}12, ZJDW18]. HeteroMPI [LR06a, VLO^{+}08]. Heuristic [BHM96, STV97, WH94]. HI [ERS96, HS94, IEE96e, ACM97a]. HICSS [ERS96, MMH93]. HICSS-26 [MMH93]. HICSS-29 [ERS96]. hiCUDA [HA11]. Hierarchical [BMR01, FBSN01, HA10, HL17, MB18, MALM95, RR02, ADMV05, BDV03, GJMM18, LZZ^{+}20, OKM12, YPZC95]. hierarchies [SYR^{+}09]. High [ACM97b, ACM98a, ACM98b, ACM00, ACM01, ACM04, AJC^{+}20, BPG94, BS21, BRST94, BS07, BDA^{+}18, CDD^{+}13, CNM11, CDHL95, CWL^{+}20, CS14, DPP01, DDL00, DE91, FGKT97, GSHL02, GBH99, GBS^{+}07, GLDS96, HMKG19, HVA^{+}16, HA11, Hol12, IEE92, IEE93c, IEE94g, IEE95k, IEE96a, IEE96f, IEE97c, IFI95, JJM^{+}11, KLH^{+}20, Kha13, KQT^{+}21, KMK16, KEGM10, KH15, Laf01, LCK11, LC97a, LkLC^{+}03, LML^{+}19, LBH12, LWP04, MW98, MPD04, ME17, MAB05, MKK21, NFK98, NU05, OPJ^{+}19, OIH10, OLG01, PKB01, PR94b, PTH^{+}01b, Rab98, RH01, SPM^{+}10, SSLMW10, SCSL12, SJ02, Slo05, SVC^{+}11, SSSS97, Tou00, Tsu07, VW92, WN10, YCL14, YWCF15, YSP^{+}05, Zho21, AH95, Ano03, BADC07, Ber96, BWT96, BID95, CHKK15, CBYG18, DL10, Duv92, EZBA16, EVMP20, ESB13, FME^{+}12, GS02, GGC^{+}07, GL96, GL97c, HDDG09, HLK^{+}20]. high [HW11, HGX^{+}22, Hos12, KBP16, KME09, Lan09, LBD^{+}96, MNYN21, MSL12, MSZG17, NS91, NFG^{+}10, Old02, OGM^{+}16, PGS^{+}13, PGK^{+}10, PF05, PTW99, RBW^{+}20, Reu03, RJDH14, SG14, SFLD15, ZSK15, ZWL13, dAT17, CDH^{+}95, DZ98b, D^{+}95, DE91, GH94, HS95a, KD12, LCHS96, LC97b, SSH08, Ten95]. High-Dimensional [MW98, MKK21]. high-frame-rate [MNYN21]. High-Level [CS14, DDL00, HA11, Hos12, RBW^{+}20, SG14, SFLD15]. High-order [KEGM10, EVMP20, KME09, OGM^{+}16]. High-Performance [ACM98a, AJC^{+}20, BS21, FGKT97, IEE97c, LkLC^{+}03, OPJ^{+}19, OLG01, PKB01, PR94b, PTH^{+}01b, Rab98, RH01, SPM^{+}10, SCSL12, WN10, GLDS96, LML^{+}19, OIH10, SVC^{+}11, Ano03, ESB13, FME^{+}12, GL96, GL97c, HDDG09, HLK^{+}20, KBP16, LBD^{+}96, Old02, PGS^{+}13, PGK^{+}10, PF05, Reu03, RJDH14, SFLD15, ZSK15, HS95a, GH94, LCHS96, SSH08]. High-Precision [Kha13]. High-Quality [BDA^{+}18]. High-Scalability [BS07]. High-Speed [CDHL95, KMK16, AH95, BWT96, CDH^{+}95]. High-Throughput [HMKG19, SSLMW10, ESB13]. Higher [MYB16, KB13, wL94]. higher-level [wL94]. Higher-order [MYB16]. Highly [MM95, PV97, TMP16, CARB10, GBH14, GBH18, JCP^{+}20, KKB^{+}21, PSH^{+}20, VM95]. highly-efficient [PSH^{+}20]. highly-scalable [GBH14]. Hills [IEE93f]. HiNet [AH95]. HIRLAM [Bjo95, HE02, KOS^{+}95a]. histogramming [KRC17]. History [OWSA95]. Hitachi [Ano03, NNON00, TSB02, TSB03]. HLA [RTRG^{+}07]. Hoare [KI17]. Hoc [IBC^{+}10, ITT02]. Högskolan [Eng00]. Hole [Kha13]. holistic [TWFO09]. Homomorphisms [RG18]. homotopy [GWC95, SMSW06, VY15]. Honolulu [IEE96e]. honor [Str94]. Host [Ano95e, LLRS02]. Host-Parasite [LLRS02]. HOTB [GSMK17]. Hotel [IEE94e]. Hotel-Copley [IEE94e]. Hough [YULMTS^{+}17]. house [ZLZ^{+}11]. Houston [ACM06a, Ano95a, Cha05, DKM^{+}92, Y^{+}93]. HP [CGB^{+}10]. HPC [ASS^{+}17, CGBS^{+}15, DH22, EYP^{+}20, GDC15, GKK09, LZZ^{+}20, LCVD94b, MKP22, MMAH20, OLG^{+}16, PRS^{+}14, RGGP^{+}18, VGP^{+}19, WDR^{+}19, ZLP17]. HPC2002 [Ano03]. Hpcfolder [JKN22]. HPCN [LCHS96]. HPF [BP98, BF01, BID95, Bri00, BDV03, CM98, CDD^{+}96, Coe94, FKK^{+}96b, FKKC96, FKK96a, LZ97, OP98, OPP00, SM02, Str94]. HPF-MPI [BP98]. HPL [Lee12]. HPVM [BCKP00, CLP^{+}99, KSS^{+}18]. HPVM-Based [CLP^{+}99]. hull [GCN^{+}13]. human [VLSPL19]. Hungarian [Fer92, FK95, LYIP19]. Hungary [DKP00, KKD04, VV95, FK95]. hunting [JPP95]. Husky [YLC16]. Huss [Ano96a, Ano99a, Ano99c, Ano99b, Ano99d, Nag05]. Huss-Lederman [Ano96a, Ano99a, Ano99c, Ano99b, Ano99d]. Hut [MPZ21]. Hybrid [BBG^{+}10, BBH^{+}06, BB18, CGC^{+}11, CNM11, Cha02, DR97, EBB^{+}20, GPC^{+}17, HVSC11, IDS16, KS15a, KLR^{+}15, KSB^{+}20, LLRS02, LRG14, MS02b, MV20, MYK19, NO02b, PZ12, SSB^{+}16, VPS17, WT12, YHL11, YPAE09, YTH^{+}12, AC07, ADR^{+}05, BBG^{+}14, CSPM^{+}96, FMS15, GÁVRRL17, GKK09, HDZ^{+}20, HDB^{+}13, JR10, JMS14, KN17, KRG13, KJEM12, LLC13, LLH^{+}14, MLAV10, MRRP11, NO02a, Nak05a, Nak05b, PARB14, PHJM11, QM21, SDJ17, SVC^{+}11, THDS19, WT11, WYLC12, WLYC12, WT13, YWC11, YW21, ZWL13, ZGZS20, ZWC21]. hybrid-core [BBG^{+}14]. Hybridizing [LSG12]. Hydra [DKB20]. HYDRA MPI [PBC^{+}01]. Hydrodynamics [RBP^{+}21]. hydrographic [LSB^{+}20]. Hyper [CSW99, SBT04, TBG^{+}02, ZAT^{+}07]. Hyper-Rectangle [CSW99]. Hyper-Threading [SBT04, TBG^{+}02, ZAT^{+}07]. hyper.deal [MKK21]. hyperbolic [PGPCK21]. hypercube [HS95b, Sur95b]. Hypercubes [Ano89, RJMC93, She95]. Hypercubic [HP11]. hyperelastic [OKW95]. hypersonic [BTC^{+}17]. Hyperspectral [VLO^{+}08].

I-SPAN [LHHM96, Li96]. I-WAY [FGT96]. I/O [Bos96, CFF^{+}96, DRUE12, IRU01, IBC^{+}10, KLH^{+}20, LkLC^{+}03, kLCC^{+}06, LPJ98, MMD98, MV17, MC18, MGC12, MG15, NFK98, OWO98, PSK08, PLR02, RK01, SBQZ14, SR98, Tha98, Tsu07, WSN99, ZJDW18]. IASTED [Ham95a]. IBM [AL93, Ano03, BBB^{+}94, BGBP01, BR95c, BR95b, Bri95, CE00, CDM93, FHPS94b, FHP^{+}94, FHP^{+}95, Fra95, FWR^{+}95, GL95d, HSMW94, HMKV94, Heb93, JF95, KB98, KAC02, KHS01, KMH^{+}14, LC97b, MP95, MW93, MABG96, NMW93, WZWS08, XH96]. IBM-SP1 [FHPS94b]. ICA [IEE96d]. ICAPP [Nar95]. ICCMSE [SM07]. ICIP [IEE94b]. ICPP [Agr95a]. ICS [RV00]. ID [DGG^{+}12]. Idaho [Str94]. Ideas [IEE95d]. identification [HPLT99]. Identifying [CTBT21]. identity [KN17]. IEEE [ACM97b, ACM98b, ACM04, ACM05, Bha93, IEE94e, IEE94g, IEE95b, IEE95a, IEE95k, IEE95g, IEE96b, IEE96f, IEE96d, IEE02, Nar95]. IEEE/ACM [ACM04]. IFIP [Boi97, DR94, PSB^{+}94]. IFS [AHP01]. Igniteg [SBK21]. Igniteg-GPU [SBK21]. Igniting [ACM03]. II [DE91, GE95, HS94, BPS01, BWW^{+}12, EM00b, GÁVRRL17, Sta95b]. III [BPG94, BP93, DSM94, GE96, Has95, OKW95, SSGF00]. ILDJIT [CARB10]. I’ll [Har94]. Illumination [STK08, ZWHS95]. ILU [ABF^{+}17]. ILU-preconditioned [ABF^{+}17]. im [Gra97]. Image [DYN^{+}06, FDG19, FLS20, FJBB^{+}00, GA96, GPC^{+}17, KBA02, KS01, LSZL02, MC18, NJ01, PLR02, RRBL01, WN10, WYZ^{+}19, ARL^{+}94, ASB18, DZZY94, GDC15, JC96, KKLL11, LK20, RKBA^{+}13, SLS96, UH96, Wu99, YULMTS^{+}17, YPZC95, YZPC95, dAT17, SBB20]. Imagery [GGCM99, GGCGO01, GCGS98, GGGC99]. Images [SSB21, Uhl94, Uhl95b, VLO^{+}08, NAJ99, RTN21]. Imaging [NH95, Has95, LM13, Pat93]. Imbalanced [Pro21]. imbalances [MLVS16]. IMEC [ZL17]. immunodominance [ZWL^{+}17]. Impact [ADLL03a, ADLL03b, BRU05, Brü12, TSS00a, WHDB05, DO96, FSV14, SHHC18]. impacts [Str94]. Implement [GM95, Gro19, PPT96c]. Implementation [AB93a, AKL99, BGG^{+}15, BGBP01, BPS01, BG95, BHP^{+}03, BBS99, Ben01, BP98, BCD^{+}15, Bjo95, BJS97, BIC^{+}10, BMR02, BRM03, BMS94b, BMG07, BDA^{+}18, CGC^{+}02, CFMR95, DYN^{+}06, DAK98, DWS^{+}21, EFR^{+}05, ES11, FH97, FD04, FHSO99, FSXZ14, FJBB^{+}00, FHPS94a, FHPS94b, FHP^{+}94, FSLS98, GBH99, GB98, GBS^{+}07, Gro02a, HPP02, HMKG19, HRZ97, HKT^{+}12, Huc96, HHA95, HAA^{+}11, IBC^{+}10, ITT02, IM94, JSS^{+}15, JSH^{+}05, LSZL02, LTRA02, LZ97, LWP04, LHCW05, MS02b, MW98, MN91, MT96, MRH^{+}96, NSS12, NNON00, OTK15, OLG01, Pan14, PLK^{+}04, PS00a, Per21, Pet97, PBK99, PTH^{+}01a, PTH^{+}01b, PB12, RDMB99, RG18, RSV^{+}05, SH94, SBF^{+}04, SBG^{+}02, Ser97, SCC96, SSC97, SZBS95a, SWJ95, SYF96, Sum12, Sur95a, TOTH99, TBG^{+}02, TRH00, TMPJ01, USE94, VT97, WH94, WPC07, YGH^{+}14]. Implementation [YWO95, ZZG^{+}14, ACGdT02, AS92, AAAA16, AAC^{+}05, ADLL03a, ADLL03b, AFG21, AB93b, BR91, BvdSvD95, BR95b, Ber96, BBCR99, BK96, BCK^{+}09, BS01, BS05, Bor99, BRR99, BS96b, BDV03, Bri95, BB00, BAS13, CDZ^{+}98, CEGS07, CG99a, CdGM96, CBHH94, CD96, DS22, DSW96, DS96a, DL10, DBB^{+}16, DSOF11, DM12, FFB99, FWNK96, FGT96, FGG^{+}98, FCS^{+}19, GCC99, GG99, GG09, GÁVRRL17, GL92, GL94, GL96, GLDS96, GL97c, GT07, GkLyCY97, HBT95, HCL05, HWW21, HS95b, ITT99, IvdLH^{+}00, JRM^{+}94, JC96, KY10, KTF03, KB21, KBVP07, KL95, KVGH11, KNH^{+}18, KB13, Lee12, LC07, LYIP19, LO96, MMO^{+}16, Man94, MV20, MAIVAH14, MS95, MSZG17, ÓN12, OKW95, OA17, OGM^{+}16, PHJM11, PR94a, PGPCK21, PTW99, PCS94, Ram07, RRFH96, Sep93, SZBS95b, SCL97, SBB20]. implementation [Sto98, SNMP10, Sur95b, Swa01, SL95, TSCS14, TKP15, TPD15, TS12b, TA14, TCP15, Tsu95, TVV96, VDL^{+}15, VGRS16, VM95, Was95a, WMRR17, WRMR19, YPA94, ZLS^{+}15, dH94, dlAMCFN12, van93]. Implementations [AKK^{+}94, Ano01a, ACMR14, AJF16, BM00, BS07, BEG^{+}10, DFP^{+}19, FB94, Gro02b, kLCC^{+}06, LCW^{+}03, Mar02, ORA12, Sap97, TSCaM12, TGEM09, VS00, WT12, ZDD97, CLSP07, ER12, ED94, GML^{+}16, ICC02, KWEF18, MKP^{+}96, NN95, Pri14, RLFdS13, WLK^{+}18, WT11, YCL14]. implemented [BBDH14, EP96, VLCM^{+}20]. Implementing [CDT05, DPZ97, Fin94, Fin95, GL95b, HB96a, HB96b, LRT07, MMH98, MS99c, MSB97, SSC96, SS99, SMTW96, SGHL01, SCC95, Tra02a, Wil93, BWT96, LHZ97, YX95]. Implementor [GL95b]. Implicit [LHCW05, MS02b, NA01, SGHL01, Bjo95, EVMP20, TSP95, WADC99]. Importance [BCG^{+}10, PCY14]. Importance-Driven [PCY14]. Improve [KBS04, SKH96, Tha98, ZWLZ21, GK97, HD00a, RHG^{+}96]. Improved [Trä02b, AFG21, MMO^{+}16, XJR21, dlAMCFN12]. improvements [DPSD08]. Improving [CGZQ13, DZ96, DCPJ12, DCPJ14, GSY^{+}13, HE02, IRU01, KLH^{+}20, KH12, KW20, KK02b, LB98, MK97, MPZ21, PTG13, RSC^{+}15, SM12, SPBR20, SCL00, XF95, CZ96, JKN^{+}13]. Imputation [Zho21]. In-depth [MKP22]. in-house [ZLZ^{+}11]. in-kernel [CZP21]. In-Memory [CLOL18, ZL17, CRM14, HSP^{+}13, SBK21]. In-Place [LTS16, HSE^{+}17, PSHL11]. Including [BWW^{+}12, GLT12]. Incomplete [MYL21]. incompressible [BCM^{+}16, Lou95, RM99, TS12b, TGS^{+}20]. Incorporating [LM94, LYZ13, TKP15]. Incremental [dOSMM^{+}16]. Indefinite [YKW^{+}18]. Independent [BCL00, BRU05, BDA^{+}18, CSW12, CBS18, CDMS15, DiN96, MV17, YBZL03]. Index [DALD18, LAD16]. Index-Digit [DALD18, LAD16]. Indexers [Wal01a]. Indexers/Crawler [Wal01a]. Indexing [LTR00]. India [CGB^{+}10, IEE96a, Kum94, PBPT95]. indicator [FSV14]. Industrial [BPMN97, DHK97, ALR94, ABCI95a, ABCI95b, BT96, EKTB99, Was96, Kon00]. industries [Ano93a]. Industry [DM98, Ano94f]. Industry-Standard [DM98]. inefficiency [HGMW12]. Inertial [Str97]. Infer [VBB18]. Inference [BBD^{+}20, LAdS^{+}15, TVCB18]. Inference-Based [BBD^{+}20]. InfiniBand [LCW^{+}03, LVP04, LWP04, PK05, PRS16, SPK^{+}12, ZLP17, SWHP05]. InfiniBand-based [PK05]. inflation [OdSSP12]. influence [Gra97]. influencing [KSC^{+}19]. inform [FGL^{+}20]. Information [Ano98, CGB^{+}10, Ano93c, CG99b, Gro19, IMS16, MMR99, WADC99, PSB^{+}94]. infrastructure [GFIS^{+}18, WLR05]. infrastructures [GWVP^{+}14]. Initial [LLH^{+}14, VDL^{+}15, AL96, LSR95]. Initiated [SSB^{+}05]. initiatives [Sun95]. initio [SSGF00, SEC15]. Injection [RRAGM97, SAL^{+}17]. Inn [IEE93c]. Innovation [ACM03]. Input [CFF^{+}94, YNJS21, CPKG17, LCH^{+}22, SHM^{+}12, JWB96]. input-aware [SHM^{+}12]. Input-Output [CFF^{+}94]. Input/output [JWB96]. Insight [IEE02]. Insights [FGL^{+}20]. Inspection [BPMN97, DLLZ19, DLLZ20]. inspired [NEM17, TDB00]. instances [RBAI17, ZLZ^{+}11]. Institute [Old02, TG94]. Instrumentation [MVY95, Yan94]. Insurance [PZ12]. Integer [ASA97, CF01, Ger18, WLC07, ZC10, BHJ96, KVGH11]. InteGrade [CC10]. integral [HK94]. Integrals [FBSN01, NS16]. Integrate [GLRS01]. Integrated [CFDL01, DGMS93, HKN^{+}01, KSV01, WL96a, DF17, HK10, KW14, VDL^{+}15, WWZ^{+}96, WL96b, XWZS96]. Integrating [BCLN97, CM98, Fin00, GJP01, KJA^{+}93, KAHS96, wL94, STP^{+}19, WTFO14, TWFO09]. Integration [CGC^{+}11, CSW97, FD96, FB94, MAIVAH14, Sei99, AL96, CSW99, KB13, RMS^{+}18, RBB15, STA20]. Integrator [Per99, SP99]. Integrity [KQT^{+}21]. Intel [Ano96c, Ano03, CBIGL19, DSGS17, GDS^{+}20, MP95, MKP22, OTK15, URKG12, VDL^{+}15, YSMA^{+}17]. Intelligence [BPG94]. intelligent [IEE95a, ZWZ^{+}95].

Intel(R) [TBG^{+}02, MMDA19, SBT04]. INtensities [ARYT17]. Intensive [LBB^{+}21, Rei01, BFLL99, BKML95, LSM^{+}18, SL94a]. Inter [KFL05, LAFA15, FKLB08, LFL11, RS19, SDB^{+}16]. Inter-Atomic [LAFA15]. Inter-Node [KFL05, FKLB08, LFL11, RS19]. inter-workgroup [SDB^{+}16]. Interaction [DMMV97, GFV99, NSLV16, Sou01]. interactions [PARB14]. Interactive [Coo95b, KPK13, KA13, NE98, RTRG^{+}07, STK08, Coo95a, IJM^{+}05]. Intercommunication [TMP16]. Interconnect [Brü12, SJ02, BWT96, SWS^{+}12, TBD96]. Interconnected [Hus00]. Interconnecting [MC98]. Interconnection [MANR09, SB95, AVA^{+}16]. Interconnects [AJC^{+}20, RA09]. Interface [Ano93d, Ano01b, BCFK99, BC19a, BDH^{+}97, CHD07, Cer99, CGH94, CDND11, DFKS01, DHHW92, DHHW93a, DBK^{+}09, FKKC96, FSLS98, Gle93, GLS94, GL95c, GLDS96, GLT00b, HDB^{+}12, HRSA97, KSJ95, KGRD10, KKDV03, KKD04, LKD08, LkLC^{+}03, LW97, MPI98a, MPI98b, MS98, MSS98, MBES94, MMSW02, MTWD06, PS01b, RWD09, SSL97, TDB00, TW01, TBD12, WD96, Wer95, YHGL01, Ada98, AD98, Ano93e, Ano94d, BBB^{+}94, BBCR99, Bru95, BDW97, BK00, BR94, CFKL00, CFF^{+}96, CD01, CG99b, DKD05, DBB^{+}16, DS96b, DLM99, DKP00, DLO03, EYP^{+}20, GRW^{+}19, HPY^{+}93, HHK^{+}19, HRR^{+}11, IMS16, JKN22, KOB01, KSJ96, KBHA94, Kra02, NS91, Pie94, PR94a, RMS^{+}18, SL94a, SWJ95, SDV^{+}95, VM95, Wal94a, Wal94b, ZWL13, ZKRA14, AMHC11, BC14, BBH^{+}06, BRU05, BDH^{+}95, Cot04, DKD08, DiN96]. Interface [FKS96, FGT96, FGG^{+}98, GGHL^{+}96, GLT99, GLS99, GLT00a, GL04, Han98, IBC^{+}10, KTF03, KKD05, LK10, MSL96, RRFH96, SWHP05, SLG95, SWL^{+}01, TGT05, YGH^{+}14, Ano95c, Ano00a, Ano00b]. InterfaceArchitecture [Sei99]. Interfaces [LBB^{+}21, MGC12, Wit16, FCS^{+}19, RJDH14, Trä12a, THMH21]. Interfacing [Lus00, PL96]. interference [ZJDW18]. Intergroup [KTAB^{+}19]. Intermediate [SML17, SML19]. internal [BBH^{+}15]. International [ACM94, ACM96b, ANS95, Abr96, ATC94, AGH^{+}95, Ano93a, Ano94a, Ano94e, BPG94, Bos96, BFMR96, Cha05, CZG^{+}08, CGKM11, CMMR12, CGB^{+}10, CH96, DSM94, DW94, EV01, EdS08, ERS95, ERS96, EJL92, Gat95, GA96, GT94, Ham95a, HAM95b, HS95a, HS94, Hol12, IEE93c, IEE93b, IEE94d, IEE94g, IEE95b, IEE95c, IEE95a, IEE95k, IEE95i, IEE95f, IEE95l, IEE96a, IEE96f, IEE96e, IEE96d, IEE97b, IEE97c, IEE05, Kum94, LCK11, LF^{+}93a, Lev95, LHHM96, Li96, MMH93, MCdS^{+}08, MdSC09, Nar95, Ost94, PW95, PBG^{+}95, PBPT95, Ree96, R^{+}92, SHM^{+}10, Sie94, Sil96, SM07, Tou96, VW92, Vol93, Vos03, Was96, YH96, ACM97a, AH95, BS94, DMW96, FR95, GH94, HJB^{+}21, JPTE94, LCHS96, Mal95, RV00, ZL96, Ano93b, HHK94, Sch93]. Internet [NE98]. Interoperabilität [GBR97]. Interoperability [BoFBW00, Don06, PLR02, SIC^{+}19, CPM^{+}18, GBR97]. Interoperable [Rab98, MSL12, YBMCB14]. Interoperation [FDG97a, FDG97b, FLD98]. Interpolants [RB01]. Interpolation [CWL^{+}20, BAS13]. interposition [GSM^{+}00]. Interpretative [MKW11]. Interpreted [FSSD17]. Interpretive [CNC10]. interprocess [SC95]. interprocessor [DS96b]. interrupts [CXB^{+}12, SH96]. Intervals [MDM17]. Intra [KLH^{+}20, GM13, VSW^{+}13]. Intra-Node [KLH^{+}20, GM13]. intra-warp [VSW^{+}13]. intrinsics [Stp18]. Introducción [VP00]. Introducing [JKM^{+}17, TBS12]. Introduction [Ano96b, AM07, Che10, Cze16, DOSW95, GSA08, HW11, Mar02, Mat00b, SK10, GT19, VP00]. Intrusive [SdR^{+}21]. Invariant [BBD^{+}20]. Invasive [URKG12]. inventory [OHG19]. Inverse [Huc96, BV99, GGC^{+}07, GG09, Wan02]. Inverses [MYL21]. Inversion [ACMR14, Kan12]. Investigating [GMdMBD^{+}07, Ros13]. investigation [PHW^{+}13]. Invisible [Wis97]. Invited [Gei93a]. IO [AHP01, BIC^{+}10, CGC^{+}02, CFF^{+}96, DL10, FGRD01, FWNK96, FSLS98, LRT07, LGG16, PSK08, PTH^{+}01a, PTH^{+}01b, SW12, Sto98, TGL02, ZZ04]. IO/GPFS [PTH^{+}01a]. IOMMU [YWCF15]. IOV [YWCF15, ZLP17]. IP [CCA00]. IPCC [SC95]. IPPS [IEE96e]. IR [SSH^{+}19, ZJDW18]. Ireland [LKD08]. IRREGULAR [FR95, BMR01, Cza02, Cza03, BL99, HASnP00, HY20, LOHA01, MR96, NP12]. irregularly [FR95, Smi93b]. ISA [Wit16]. ISBN [Che10, SD13]. ISBN-13 [Che10]. ISCA [Ano94e, YH96]. Ischia [ACM06b]. iScore [RJH^{+}20]. Iserver [SHH94a, SHH94b]. Iserver-Occam [SHH94a, SHH94b]. Ising [AL93, KO14]. island [JPL22]. Isolating [Lus00]. Isosurface [PCY14]. ISPAN [HHK94]. Israel [DSM94, IEE96h]. Israeli [IEE96h]. ISSAC [Lev95]. ISSTA [Ost94]. Issue [AM07, BDB^{+}13, BC00, GSA08, MPI98a, MPI98b, SBG20, BC19a, CHD09, DKD07, GT19, Mar02, Old02, TH20]. Issues [BDT08, FD02a, KGK^{+}03, MW98, Pan95b, PRQ21, PS01b, ZDD97, ARvW03, EGH99, FD02b, HHA95, PBK99]. Italy [CMMR12, CH96, DKD05, DKD07, D^{+}95, DLO03, HS95a, IEE95h, KG93, OL05, ACM06b, Ano93b, CLM^{+}95, DR94, Sil96]. Iteration [BAC20, HF14a, HF14b, OHG19]. Iteration-fusing [BAC20]. iterations [Lou95, YST08]. Iterative [CCSM97, DK06, NO02b, Nak03, SC04, ADDR95, EDSV09, LSR95, MGG05, NO02a, Nak05a, Nak05b, OMK09, dH94]. Ithaca [PBG^{+}95, Ree96]. IV [SPH95]. IWOMP [CZG^{+}08, CGKM11, CMMR12, EdS08, MCdS^{+}08, MdSC09, SHM^{+}10]. IWPP [Kum94, PBPT95]. IWPP-94 [Kum94, PBPT95]. IWWP [Kum94]. IX [R^{+}92].

Jack [Ano95b, Ano96a, Ano99a, Ano99b, Nag05, NMC95]. Jacobi [BBDH14, CGU12, LM99]. JaMP [KBVP07]. January [ERS96, GE96, HS94, IEE95h, IEE96g, MMH93, USE95]. Janus [GJP01]. Japan [SHM^{+}10, SPE95, HHK94, IFI95]. Jason [Che10]. Java [ACM98a, Ano97, BCFK99, BDY99, Bra97, BK00, BKO00, CGJ^{+}00, CFKL00, CLL03, DeP03, Fer98b, Fer98a, GGS99, KOB01, KBVP07, LRW01, MSS98, MG97, NE98, RAS16, SMS00, SZ99, TDB00, VGRS16, VGS14, WN10, WCS99, YC98, YHGL01]. Java-based [WCS99]. Java-MPI [GGS99]. Java/CORBA [LRW01]. JavaNOW [TDB00]. Jaypee [CGB^{+}10]. Jeff [Stp02]. Jersey [Bha93]. Jerusalem [DSM94]. Jiang [Ano95b, NMC95]. JMI [GDEBC20]. Job [KSC^{+}19, NSS12, PS19b]. Jobs [GSHL02, OPM06, WDR^{+}19, ZA14]. Join [BGD12, LTRA02, SML17, BMS^{+}17, SML19, She95]. Joint [GT94, Ano03, YHGL01, Ano93c]. JOMP [BK00]. Jose [ACM97b, GE95, GE96]. JPEG [CLBS17, NU05]. JPT [BDY99]. JPVM [Fer98b, Fer98a, LGCH99]. Jr [ACM99]. Juggler [BLVB18]. Julia [Cre16]. July [ACM95b, ACM97a, Boi97, EV01, GA96, Has95, IEE93c, IEE96i, Lev95, PW95, TG94]. Jumpshot [ZLGS99]. June [ACM90, Ano94f, B^{+}05, BG91, CZG^{+}08, CGKM11, CMMR12, DSZ94, DW94, D^{+}95, IEE94e, IEE95c, IEE95i, IEE96d, IEE96h, KG93, LHHM96, Li96, MCdS^{+}08, MdSC09, R^{+}92, SL94a, SHM^{+}10, TG94, Vos03]. Jupiter [Str94]. Just [FKLB08, FSSD17, KFL05, FK94]. Just-In-Time [FSSD17, FKLB08]. JVMPI [DeP03].

k-ary [Pan95a]. K20 [MYL21]. Kalman [BY12]. Kanazawa [HHK94]. Kandrot [Che10]. Karlsruhe [Cal94, Sch93]. Karlsruher [Reu01]. Katsevich [DYN^{+}06]. Kaufmann [SD13]. KBLAS [AKL16]. Keele [Ano93c]. KENO [RP95]. KENO-Va [RP95]. Kernel [CKmWH16, CFDL01, EBKG01, HKT^{+}12, MBBD13, PWP^{+}16, STA20, SNN^{+}19, TY14, CZP21, FMFM15, GM13, MMW96, PSB^{+}19, RJH^{+}20, SAP16, YBZL03, AKL99, PSH^{+}20]. Kernel-assisted [MBBD13, GM13]. Kernel-based [CKmWH16, TY14]. kernel-independent [YBZL03]. Kernel-Level [HKT^{+}12]. Kernels [BCD^{+}15, CTBT21, KI17, KAC02, LCY19, Pet01, Ros13, SNN^{+}20, SSB^{+}17, VZT^{+}19, WQKH20, ARS89, BCD^{+}12, FSV14, FVLS15, FFM11, KKM15, LCH^{+}22, MH21, PTG13, PGS^{+}13, PSH^{+}20, TBB12]. Kerr [Kha13]. key [LF^{+}93a]. kind [SP11]. Kinect [KPK13]. kinetic [JL18]. Kinetics [LD01, BTC^{+}17]. King [ACM99]. Kingdom [Boi97]. Kirchhoff [SSS99]. Klagenfurt [Bos96]. Knapsack [ICC02]. KNEM [GM13]. Knights [GDS^{+}20]. KNN [MPZ21]. knowledge [FNSW99]. knowledge-based [FNSW99]. Knoxville [PR94b]. Kohr [Stp02]. Kokkos [EVMP20]. Kolmogorov [Str97]. KOP3D [KR09]. Koppelrandkommunikation [Gra97]. Kpi [EML00]. KPN2GPU [BK11]. KPP [AC17]. Kremlin [GJLT11]. Kronecker [LNW^{+}12]. KSIX [AUR01]. KSR1 [BL94]. KU [IM94]. Kungl [Eng00]. Kyoto [IFI95, SPE95, IFI95].

L [AAC^{+}05, BGH^{+}05, EFR^{+}05, MSW^{+}05].

LA-MPI [YSP^{+}05]. Lab [Str94]. Label [ABG20]. Labeling [PPJ01, KRKS11]. labelling [HLP10]. laboratory [JY95]. LABS [RRJ^{+}20]. Lafayette [EV01, EdS08]. Lagrangian [CT94a, CT94b, RSV^{+}05, TC94]. Lahey [Ano98]. Lake [Hol12]. LAM [OF00, RšT06, SSB^{+}05, Squ03, Swa01, ZWZ05]. LAM/MPI [OF00, RšT06, SSB^{+}05, Squ03, ZWZ05]. lambda [PQ07]. lambda-calculus [PQ07]. LAMGAC [MSOGR01, MS02a]. Lamport [TPLY18]. LAN [CCU95, CDH^{+}95, MSOGR01, MTSS94, TSZC94, ZGC94]. LAN-based [TSZC94]. LAN-Message [MTSS94]. Lanczos [GP95, Sch96a, Sch96b]. Landing [dCZG06, GDS^{+}20]. Landsat [GGCM99, GCGS98]. Landsat-TM [GGCM99, GCGS98]. Lane [HHC^{+}18]. Language [ACM96a, NM95, PD98, Stp18, TA14, WLR05, Ben95, CGK11, Hos12, Nob08, RKBA^{+}13, Röh00, Stp20]. Language-based [Stp18, Stp20]. Languages [CFF^{+}94, FMSG17, FSSD17, CH96, Mar05, Olu14, SWS^{+}12, PBG^{+}95, SS96]. LANs [Fin97]. LAPACK [Add01, ARvW03]. LaPerm [WRSY16]. LAPI [BGBP01]. Laplace [ACMR14]. Large [AKE00, BHW^{+}17, BKK20, BZ97, BJS99, BHNW01, CGC^{+}11, CBB^{+}20, DALD18, FFP03, HC17, Huc96, JFGRF12, LLY93, MKC^{+}12, MFPP03, PCY14, Rót19, RGB^{+}18, SGJ^{+}03, SM03, SvL99, TGEM09, WMC^{+}18, WT12, ZWJK05, AASB08, AMS94, AMC^{+}19, BCA^{+}06, BA06, BCH^{+}08, Che99, CCHW03, DZZY94, FME^{+}12, GG99, IPG^{+}18, IM95, JLS^{+}14, KEGM10, Kos95b, KA95, LS10, MLA^{+}14, NWT21, NFG^{+}10, PTL^{+}16, PGPCK21, PD11, RMNM^{+}12, SIC^{+}19, SC96a, TBB12, TOC18, WT11, WT13, ZWL13, ZA14]. large-message [AMC^{+}19]. Large-Scale [AKE00, BHW^{+}17, BZ97, CBB^{+}20, FFP03, HC17, MFPP03, SM03, WMC^{+}18, WT12, BKK20, BJS99, SvL99, AASB08, BCH^{+}08, Che99, FME^{+}12, IPG^{+}18, LS10, MLA^{+}14, NWT21, PD11, RMNM^{+}12, SIC^{+}19, WT11, WT13, ZA14]. large-sized [JLS^{+}14]. Larger [NB96]. LargeScale [LAdS^{+}15]. laser [EZBA16, WWZ^{+}96]. LASs [VLCM^{+}20]. Lastverteilung [Wil94]. Latency [Jes93a, Jon96, KBHA94, NCB^{+}12, NCB^{+}17, TBD96]. latency-tolerant [NCB^{+}12, NCB^{+}17]. Lattice [BBK^{+}94, BMS94b, HLP11, SJK^{+}17a, SJK^{+}17b, BW12, BMS94a, CGK^{+}16, GM18, Sai10, STA20, SVC^{+}11, BLPP13, OTK15]. launches [Ano03]. Layer [CSAGR98, HEH98, FKK96a, PTT94, RS21, dlAMC11, dlAMCFN12]. Layered [Taf21, DiN96]. Layering [Hus01]. Layers [VZT^{+}19, KC94]. Layout [WG17, BGH^{+}05, HP11, LDJK13, Str12]. Lazy [TCBV10]. Leaks [DLV16]. Learned [GKPS97, MWO95]. Learning [AHHP17, AJC^{+}20, GDS^{+}20, Gro01b, TWLL19, ZJHS20, ZLWW20, AMC^{+}19, FE17a, FE17b, KWEF18, LSSZ15, NWT21, SEC15, TWFO09, WO09, WTFO14]. learning-based [FE17a, FE17b]. Least [PWP^{+}16, VRS00, DK13]. Least-Squares [VRS00]. Lecture [Gei93a]. Lederman [Ano96a, Ano99a, Ano99c, Ano99b, Ano99d, Nag05]. Leeds [Abr96]. legacy [BR04, LP00, LRW01]. Legio [RGP22]. Lemon [DRUE12]. Length [FLS20]. Lengths [GSHL02]. LEO [CCBPGA15]. Leonardo [Stp02]. Lessons [MWO95]. Level [AELGE16, BGG^{+}15, BBC^{+}00, CS14, CRGM14, DHHW92, DHHW93a, DDL00, GS91b, GAM^{+}02, HA11, HKT^{+}12, DK02, KCP^{+}94b, KOW97, LVP04, LMRG14, NPP^{+}00c, SHM^{+}10, SBF^{+}04, TS12a, TW01, XF95, BMPS03, CAWL17, CRM14, CRGM16, EPP^{+}17, GGS99, HE15, HK09, Hos12, KCP^{+}94a, LRG^{+}16, wL94, LCY19, LCMG17, LBB^{+}19, LM13, MALM95, NS91, Nak05b, RBW^{+}20, STY99, SCL97, SG14, SFLD15, WDR^{+}19, YZ14, ZWZ05, ZZZ^{+}15, BBH… 13a]. levels [AML^{+}99]. Leveraging [BBW19, HDB^{+}12, NPP^{+}00c, SHLM14, BPJ22, LFL11]. LFIB4 [Stp20]. LIB [NPP^{+}00d]. libefp [KS15a]. libOMP [BGD12]. Libraries [BHLS^{+}95, BWV^{+}12, CGZQ13, DARG13, GFD05, IEE94f, IEE95j, MLGW18, MM14, ARvW03, BCM11, BfDA94, CRD99, DWS^{+}21, GS94, PS07, Skj93, SDB94, SSG95, DHK97]. Library [AKL16, Ada97, BS21, Boo01, BLW98, CBB^{+}20, CBB^{+}21, Coo95b, DHP97, EM02, FHK01, For95, GFB^{+}03, GSI97, Gro02a, HB96b, ITKT00, JPT14, KBG16, MKK21, OD01, PLK^{+}04, PS01a, RR02, Rót19, Saa94, SBG^{+}02, Sta95b, SKH96, TD98, UTY02, WN10, YKLD17, ZC10, Ada98, AMHC11, Arn95, CSS95, CGG10, CCS19, Coo95a, DRUE12, DXB96, FB97, Fan98, FKK^{+}96b, GDC15, GÓ19, GLM^{+}08, GL94, HB96a, HLM^{+}17, Har94, Har95, JKM^{+}17, JC96, KS15a, KN95, LR06a, MSL96, PKB06, PS00b, RFH^{+}95, SSC96, SH96, TK19, VLCM^{+}20, ZT17, CC95, McD96, Sum12]. Life [PZ12, Str94]. Lifting [vdLJR11]. light [LK20]. Lightweight [CKmWH16, DT17, FLB^{+}05, KMK16, SWCB20, TCM18, FS95, HWW21, Ott93]. Like [BST^{+}13, BK00, BKO00, CGJ^{+}00, HY20, KOB01, TSEE21, VGS14, CSS95]. Likelihood [TSN21]. Likelihoods [MSCW95]. LIME [DRUE12]. Limits [GB96, MBKM12]. LINDA [KS96, MSP93, BLP93, CSS95, Gal97, Mat94, Mat95, TDB00]. Linda-like [CSS95]. Line [BoFBW00, CGS15, Wis98, Bor99]. Linear [ASA97, BDT08, BG95, CDD^{+}13, DGH^{+}19, Gao03, Huc96, LLY93, LZ97, MB18, MGMH97, MSB97, YKW^{+}18, ZTD19, van97, BSN95, BAC20, BKvH^{+}14, BAV08, BRR99, CEGS07, DR18, Gra09, GFPG12, Jou94, LSB^{+}20, LRLG19, MW98, MM11, OKW95, SCC96, SMSW06, VLCM^{+}20, dCH93, dH94]. Linear-scaling [Gao03]. linearization [MH18]. Lines [NE01, YULMTS^{+}17]. Link [BGR97b, SJ02]. Linked [WJ12]. Linköping [FF95]. LINPACK [JNL^{+}15]. Linux [Sei99, USE00, SSSS97, Ano01a, GSN^{+}01, MK04, OF00, PS07, PKB01, RšT06, Sei99, SMTW96, Slo05, SGL^{+}00, YL09]. Linz [Kra02]. lipid [FHSO99]. Liquid [DSS00, JLS^{+}14, ZL18]. Lisbon [IEE93d]. LISP [ACM90]. List [Tra98, WJ12]. Lithe [PHA10]. Lithography [RDMB99]. Liverpool [AD98]. LLVM [SML17, SML19]. Load [Ano94b, BKdSH01, BS05, DI02, DR95, DK06, GCBL12, HE02, KSB^{+}20, MM02, NP94, PT01, Pus95, SGS95, ST97, Wal01a, Bir94, CKO^{+}94, DZ96, DLR94, DvdLVS94, EZBA16, FMBM96, FH97, GS96, Hum95, JH97, MM03, SCL97, SY95, Wil94]. load-balanced [EZBA16]. Local [BSG00, CDHL95, CCSM97, IKM^{+}01, LBB^{+}19, AMHC11, BY12, CGL^{+}93, FSV14, IKM^{+}02, LHD^{+}94, LHD^{+}95, PHM^{+}22, RRJ^{+}20]. Locality [AAB^{+}16, MJB15, TPK^{+}19, ZLP17, BHRS08, CMZ99, HJYC10, KW20, RKBA^{+}13, WRSY16]. Locality-Aware [AAB^{+}16, MJB15, HJYC10]. localization [HC08]. Localized [DDN^{+}22]. Locally [BHS^{+}02]. Locating [PNV01]. Lock [ALB^{+}18]. Lockheed [Str94]. Locking [kL11, CAWL17, PGK^{+}10]. Log [DDN^{+}22]. Logging [BCH^{+}03, DDN^{+}22, LBB^{+}19]. Logic [KI17, BJ95, KMC96, KMC97, POL99]. Logical [SR98, TPLY18]. LogP [CKP^{+}93]. London [EJL92, Ano93h, Ano94f]. long [dFdOSR^{+}19, ZCBD22]. Longest [Per21]. Look [HCZ16]. lookup [BJ13]. Loop [DMB16, HC17, SHM^{+}10, TJPF12, AV18, SHLM14, WYLC12, WLYC12, YST08, YWC11]. Loops [AHD12, CLA^{+}19, COE20, DSCL05, HH22, LOHA01, RRJ^{+}20]. Loosely [Ada97]. Lop [RGDML16, RGDM15]. Louisiana [USE95, IEE96b]. Love [Dan12]. Love-Hate [Dan12]. Low [BGG^{+}15, FLS20, GGS99, Jon96, MC17, NE01, RLL01, SM19, Str94, GK97, KBHA94, LZHY19, TBD96, ZRQA11]. Low-Bandwidth [NE01]. Low-Cost [FLS20, RLL01, GK97]. Low-Density [MC17]. Low-Level [BGG^{+}15, GGS99]. Low-life [Str94]. low-overhead [ZRQA11]. Low-power [SM19]. LPN [TSCS14]. LPVM [ZG98]. LSS [BCAD06, BADC07]. LU [AZ95, BRS92, BB18, LC97b]. Lugano [GT94]. Luminous [KNT02]. Lumsdaine [Ano99c, Ano99d]. Lusk [Ano95c, Ano99c, Ano99d, Ano00a, Ano00b]. Lustre [DL10]. Luther [ACM99]. Lyngby [DW94, DMW96, Was96]. Lyon [BFMR96, FR95].

M [PBC^{+}01]. M-SPH [PBC^{+}01]. M2L [KKB^{+}21]. M6A [EM00a]. M6B [EM00b]. MA [Ano95b, Ano95c, Ano96a, Ano99a, Ano99c, Ano99b, Ano99d, Ano00a, Ano00b]. Machine [AS92, AGIS94, BJ93, BS93, CHD07, D^{+}91, FE17a, FE17b, Fis01, GBD^{+}94, Gre94, JCP^{+}20, KNT02, KKDV03, KKD04, LKD08, MTWD06, Nov95, NMC95, Pat93, Per96, RWD09, TY14, VS00, Wel94, AD98, AL92, Ano95b, BR91, BDG^{+}91a, BPC94, Bir94, BDLS96, BDW97, CARB10, CLM^{+}95, Cav93, Cha96, Che99, CD01, CC00b, DM93, DKD05, DLM99, DKP00, DLO03, FM90, KWEF18, KMC97, KSS^{+}18, Kra02, LG93, MN91, MRH^{+}96, NB96, Sch94, SK92, SCC96, SL00, TVCB18, TW12, TWFO09, WO09, WTFO14, ARL^{+}94, BG94b, JPP95, KKD05, LK10, QRG95, SSSS96]. machine-learning [TWFO09]. machine-learning-based [WTFO14]. Machines [BP99, BZ97, BCC^{+}00a, BT01b, CDT05, DR97, EGR15, GB96, GTS^{+}15, HC10, MGL^{+}17, STY99, SCSL12, ZWJK05, BCA^{+}06, BSC99, BCC^{+}00b, BBW19, BB95b, DDS^{+}94, DCH02, GKZ12, Hol95, KN95, PRS16, RJH^{+}20, SL94b, TSY99, TSY00, WPL95, ZWL13, Gei01, YC98]. made [MJPB16]. MAFFT [ZLS^{+}15]. Magnetic [Y^{+}93, PKE^{+}10]. Magnetism [Y^{+}93]. magnetized [CFF19]. Magnetohydrodynamic [KT02, WWFT11]. magnetohydrodynamics [ZT20]. Magnetostatic [BB93]. MagPIe [KHB^{+}99]. Main [Tou96]. Maintaining [PKB01]. maintenance [ZDR04, ZDR01]. major [WLK^{+}18]. Makes [ZG95b, Str94]. Malleable [EDSV09, MSMC15]. Mambo [WZWS08]. Man [IEE95a]. Manageable [PKB01]. Managed [KCR^{+}17, LB16, SYR^{+}09]. Management [AJ97, ALB^{+}18, AUR01, BGR97b, BGL00, CVPS19, EK97, FDG97a, FDG97b, GJR09, PPT96a, PS00a, SIS17, STY99, THS^{+}15, ARS89, DZ96, DF17, FLD96, GJMM18, GL95a, JCP15, LF^{+}93a, PPT96b, PPT96c, YWTC15]. manager [Sep93, SSD^{+}20]. managers [FLD96]. Managing [FLD98, FGKT97, Liv00, NPS12, Obe96]. Manchek [Ano95b, NMC95]. Mangrove [BBD^{+}20]. Manipulation [KKV01]. Mantle [BB95b]. Manual [CSW12, NSLV16, Reu01]. Many [DT17, LZH17, LLCD15, RB01, SXMX^{+}18, TCM18, YTH^{+}12, ACMZR11, AV18, BBC^{+}19, VDL^{+}15, dCZG06]. Many-Accelerator [SXMX^{+}18]. Many-Core [LZH17, TCM18, YTH^{+}12, LLCD15, ACMZR11, AV18, BBC^{+}19, KSG13, MBBD13, dCZG06]. Many-Cores [DT17]. Manycore [MJB15, CdOO^{+}20, DJJ^{+}19, KGB^{+}09]. Map [JPT14, FFM11, FJBB^{+}00, MSCW95]. MAPA [JJPL17]. Maple [Pet00a, Pet00b, Pet01]. Mapping [BB18, DDP^{+}19, FDG19, GAMR00, HC06, NTR16, RRBL01, SPB^{+}17, Taf21, TSZC94, WO09, ASAK19, DDLM95, EO15, GFIS^{+}18, HC08, TWFO09, WCS^{+}13, WTFO14, WK08a, WK08c, dCZG06, WK08b]. MapReduce [EADT19, GGZ^{+}20, JS13, MMM13, PD11, WZHZ16]. Maps [BM97, KRC17]. Marc [Ano96a, Ano99a, Ano99c, Ano99b, Ano99d, Nag05]. March [ACM95a, ACM06a, Ano89, Ano93c, Cal94, DKM^{+}92, IEE93f, IEE94d, IEE95b, IEE97a]. Marine [LLRS02]. market [LF^{+}93a]. Markov [BBH12, FK01]. Marlioz [GA96]. Marsa [Stp20]. Marsa-LFIB4 [Stp20]. marshaling [CFKL00]. MARTE [RGD13]. Martin [ACM99]. Maryland [IEE96c, SPH95]. MASA [dFdOSR^{+}19, SMM^{+}16]. MASA-OpenCL [dFdOSR^{+}19]. MasPar [ARL^{+}94]. Massachusetts [IEE94e]. masses [Cla98]. Massive [Sie92a, MALM95, OLG^{+}16]. Massively [BJ93, BHS18, BBH12, DSZ94, IEE94a, IEE96c, KHBS19, KmWH10, LPJ98, Oed93, Sie92b, Sta95b, CS96, DR94, HVSC11, KN17, KB21, LCL^{+}12, MYB16, RBB17, SRK^{+}12, DSZ94]. massively-parallel [MYB16]. Master [FH98, EML00, LTR00, HP05]. master-slave [HP05]. Master-Workerproblem [FH98]. Master/Slave [LTR00]. Master/Worker [EML00]. Matching [GGC^{+}07, KMM15, KS01, MM02, OWSA95, WH94, FLPG18, FGL^{+}20, GMA20, LFS^{+}19, MM03, Qu95, YPZC95, YZPC95]. Materials [STH22, Y^{+}93, PSV19, SSP^{+}94]. Mathematical [Per21, VZT^{+}19, Wan97, Has95]. Mathematics [Whi04, ANS95]. MATLAB [BKGS02, RBC20, Whi04, Ano97, Bra97, ZZG^{+}14]. MATLAB-MPI [BKGS02]. MatlabMPI [KA04, Kep05]. Matloff [Edd18]. MATOG [WG17]. matrices [DR18, GG99, GSMK17, Kan12]. Matrix [AKL16, BSvdG91, Cha96, DS13, DK20, Fuj08, GK10, KF16, KK19, MKK21, PMvdG^{+}13, TQDL01, TD98, ART17, CMH99, ER12, FAF16, FJZ^{+}14, KBP16, MPS20, PKD95, TPD15, XXL13]. Matrix-Free [DK20, KK19, MKK21]. Matrix-Vector [AKL16, DS13, Fuj08, XXL13]. matting [WLYL20]. Maui [ACM97a]. Max [Ano94c]. Max-Planck-Gesellschaft [Ano94c]. Maximal [BDA^{+}18]. maximisation [CCU95]. Maximizing [PIR^{+}20]. Maximum [TSN21, HKOO11]. Maxwell [And98, lLLmH^{+}21]. May [ACM96b, ACM06b, AGH^{+}95, BR95a, BS94, Cha05, DT94, EdS08, Gat95, HS95a, IEE95e, IEE95d, IEE95i, PR94b, RV00, SPE95, SW91, SS96, Van95]. Maydan [Stp02]. MBCF [MMH99]. MCA [WCS^{+}13]. McDonald [Stp02]. MCHF [SYF96]. McLean [IEE94a, Sie92a, Sie92b]. MCNP [MW93, McK94, WH96]. MD [IEE02, TMPJ01]. mdb [DKF94a]. MDE [RGD13]. Means [TK16]. Measurement [BFBW01, BFIM99, KRS99, Shi94, TMC09]. Measurements [IHvA^{+}00, EFR^{+}05, GL99]. MECCA [AC17]. mechanics [Bil95, MGG05, SL95]. Mechanism [CGLD01, KSV01, MH01, THS^{+}15, TSS00b, Tra02a, HWX^{+}13, SIRP17, ZRQA11, ZA14]. Mechanisms [Wal01a, CGBS^{+}15, Ott93, TMTP96]. Mechatronic [KDL^{+}95b, KDL^{+}95a]. mEDA [VAT95]. mEDA-2 [VAT95]. media [EZBA16, MAIVAH14]. Medical [WYZ^{+}19, RTN21]. Medicine [GA96]. MEDINA [AC17]. Medium [CWL^{+}20, WLNL06]. medium-scale [WLNL06]. Meeting [AD98, Ano93f, CHD07, CD01, CDND11, DKD05, DLM99, DKP00, DLO03, GA96, KGRD10, Kra02, KKD04, LKD08, MC94, MTWD06, RWD09, TBD12, BDW97, JB96, SPH95, Ano92, CHD09]. megabase [SdM10]. Meiko [FST98a, FST98b, Jon96]. Melia [WZHZ16]. Mellon [IEE94d].

Membership [BMS19, MDM17]. membrane [FHSO99]. Memory [ADGA20, Att96, BME02, BWW^{+}12, Bri10, BdS07, BT01b, CVPS19, CDT05, CLOL18, CLA^{+}19, CSW97, CC99, DM98, DMB16, DR97, DHHW92, DHHW93a, EADT19, FB94, GGZ^{+}20, GCBM97, GB96, GSN^{+}01, GSHL02, GLRS01, HC10, HDB^{+}12, HDT^{+}15, HT01, JJPL17, KB98, KS13, KC19, KSHS01, LSB15, LWKA15, LML^{+}19, Luo99, MB12, MRB17, MBE03, MMH98, MCdS^{+}08, Mül02, NPP^{+}00d, PBK00, Pok96, PMvdG^{+}13, Ros13, STY99, ST02b, SW91, Thr99, VS00, VT97, WJA^{+}19, ZL17, ZL18, ARS89, ABCI95a, ABCI95b, ADMV05, BCA^{+}06, BVML12, BSC99, BMG07, CdOO^{+}20, CBPP02, Cha05, CJvdP08, Cha96, CBHH94, CRM14, CC00b, DF17, DLR94, DBVF01, DPFT19, DS96b, DHHW93b, DPZ97, EVMP20, EV01, FSV14, FHB^{+}13, GCN^{+}10, GBH14, GBH18, GKK09, GL96, GL97c, GP95, GADM20, HSP^{+}13, HGMW12, HDB^{+}13, HK09]. memory [JC17, JE95, KN95, KSTM20, KJA^{+}93, KC06, LKL96, MLC04, NAJ99, NAAL01, OLG^{+}16, PK05, PS00b, QM21, RS19, RGDM15, SSH08, SHHI01, SL94b, SBG^{+}12, SYR^{+}09, SFL^{+}94, SSC96, SPL99, SBK21, SD16, SPNB14, TSY99, TSY00, THDS19, TSCS14, Uhl95a, Vos03, Wal94a, Wal94b, WPL95, WK08a, WK08b, WK08c, WK20, WBSC17, WMRR17, WRMR19, YX95, LBD^{+}96, GK97, SG05]. Memory-access-aware [CLA^{+}19]. Memory-Based [MMH98]. memory-constrained [TSCS14]. Memory-Divergent [WJA^{+}19]. Memory-Efficient [GGZ^{+}20, MRB17]. memory-level [HK09]. Memory-Oriented [ZL18]. Memory/Message [ST02b]. MemTo [GSN^{+}01]. Menon [Stp02]. Mesh [DDP^{+}19, HAA^{+}11, MRB17, Ran05, BAS13, CLSP07, Cou93, GBR15, HDZ^{+}20, IDS16, SWCB20]. mesh-oriented [HDZ^{+}20]. mesh-particle [BAS13]. Meshes [MRB17, TPD15]. Mesoscopic [VT19]. Message [Ano93d, AKL99, Att96, BC19a, BZ97, BCH^{+}03, BBG^{+}99, BBG^{+}01, BDH^{+}97, BGR97b, BFM97, CHD07, Cer99, CGZQ13, CGH94, Cot97, Cot98, CTK00, CDND11, DFKS01, DDN^{+}22, DHHW92, DHHW93a, DDL00, FKKC96, Fos98, FB94, GR07, GB96, Gle93, GLRS01, GLS94, GL95c, GLT00b, Hem94, KGRD10, KS97, KSV01, KKDV03, KKD04, LKD08, Luo99, MPI98a, MPI98b, MP95, MS98, MBES94, MG97, MTWD06, MSS97, NW98, PBK00, Pok96, RC97, RRBL01, RWD09, RFG^{+}00, SAL^{+}17, ST02b, TBD12, WD96, Wer95, Wis97, YHGL01, ZWL13, ZG95a, ZG96, ZLL^{+}12, Ada98, AD98, AAC^{+}05, Ano93e, Ano94d, Ano95c, Ano00a, Ano00b, AMC^{+}19, BBG^{+}14, BL97, BvdSvD95, Bjo95, Bru95, BDW97, BFIM99, CGJ^{+}00, CDZ^{+}98, CRD99, CD01, CG99b, DKF93, DM93, DKD05, DS96b, DHHW93b, DOSW96, DLM99]. message [DKP00, DLO03, FGL^{+}20, FK94, GMA20, GL92, HP05, HPY^{+}93, Hem96, JKN22, KJA^{+}93, Kra02, LR06a, LBD^{+}96, wL94, LFS^{+}19, LCY96, LMM^{+}15, LBB^{+}19, LC97b, NS91, PS07, PKB06, Pie94, PR94a, PS00b, Sei99, SWJ95, SDV^{+}95, SZ99, SSG95, Sti94, TSZC94, VM95, Wal94a, Wal94b, ZKRA14, ZA14, AMHC11, BC14, BBH^{+}06, BRU05, BDH^{+}95, Cot04, DKD08, DiN96, FKS96, FGT96, FGG^{+}98, GGHL^{+}96, GLDS96, GLT99, GLS99, GLT00a, GL04, Han98, IBC^{+}10, KTF03, KKD05, LK10, MTSS94, MSL96, PS01b, RRFH96, SWHP05, SLG95, SWL^{+}01, TGT05, TDB00, Wer95, YGH^{+}14]. Message-Passing [Ano93d, Att96, Cot97, Cot98, DHHW92, DDL00, GLS94, GL95c, GLT00b, MPI98a, MPI98b, PBK00, Pok96, RRBL01, AAC^{+}05, Ano94d, Ano95c, Ano00a, Ano00b, BvdSvD95, CDZ^{+}98, GL92, Hem96, KJA^{+}93, LR06a, LBD^{+}96, wL94, LMM^{+}15, PS00b, SSG95, Sti94, DiN96, GGHL^{+}96, Han98, RRFH96, SLG95, Wer95, YGH^{+}14]. Message-Passing-Interface [Wer95]. MessagePassing [Sei99]. Messages [KBS04, SKH96]. Messaging [HEH98, KC94]. Meta [BCLN97, FBD01a, FGRD01]. Meta-Applications [BCLN97]. Meta-computing [FBD01a, FGRD01]. Metacomputer [OS97]. Metacomputing [Fin00, MSF00, MS99b, FBVD02]. Metagenomics [LSM^{+}18]. MetaHaskell [Mai12]. metaheuristics [ZSK15]. metal [JLS^{+}14]. MetaMP [OW92]. metaprogramming [Mai12, TSEE21]. meteorological [RSBT95]. Meteorology [HK93, HK95]. Method [ADGA20, ACMR14, BP99, BJS97, CGU12, DAD19, FCLG07, GSI97, HFB21, HC06, KMK16, OMK09, RHM^{+}17, Riz17, STA20, TSS00a, ARYT17, AiIS^{+}21, AFG21, BBDH14, BCM^{+}16, DSOF11, ETV94, GFIS^{+}18, HE13, HMKV94, HJBB14, HPLT99, JMS14, KS15a, KD12, KKB^{+}21, LCL^{+}12, MMDA19, Nak05b, NS16, PTT94, PGPCK21, Pri14, Qu95, RTN21, SHHC18, TKP15, YBZL03, dlAMCFN12, AAB^{+}17, OTK15]. Methodologies [Sun94b]. Methodology [MOL05, WTTH17, HPR^{+}95, LM94, WMP14]. Methods [BCMR00, CMK00, DFN12, EGH^{+}14, FGKT97, GFPG12, KLR^{+}15, kL11, NA01, Sch01, SM07, TDBEE11, Whi04, ZGNZ22, ZB97, CdOO^{+}20, CEGS07, DF17, D^{+}95, Gra09, Has95, KW20, LSR95, MM11, Nak05a, PGK^{+}10, PGPCK21, R^{+}92, SL94a, SGS95]. Metric [SNN^{+}19]. Metrics [DW02, PARB14]. Metropolis [HJBB14]. Mexico [IEE91, RV00, Sie94]. MGCG [TSS00a]. MGF [GLM^{+}08]. MGRIT [HFB21]. MIAOW [BGG^{+}15]. MIC [BB18, CCBPGA15, LCY19]. MICE [BK96]. Micro [Ano03, BWV^{+}12, SGH12, YSWY14].

Micro-applications [SGH12].

Micro-Benchmark [BWV^{+}12, YSWY14]. microbenchmark [BO01]. Microcoded [PWP^{+}16]. microtask [OIS^{+}06]. MIDAS [BFZ97]. Middleware [AUR01, CLL03, CC10, RPS19]. Middlewares [DPP01]. Midpoint [JMS14]. Migol [LS08]. Migratable [KOW97]. Migrating [VSRC94, VSRC95, IvdLH^{+}00, KBG^{+}09]. Migration [Ano94b, CCK^{+}95, CLL03, CML04, CCBPGA15, CTK01, NPP^{+}00c, NLRH07, Ott94, OS97, PS19b, ST97, AMBG93, BBGL96, CKO^{+}94, CRM14, CRGM16, CK99, DDYM99, HZ99, LCVD94b, LM13, QHCC17, RRFH96, SSS99, SCL97, Ste96]. Milan [HS95a]. million [LHLK10]. Millions [BBG^{+}11]. MIMD [BvdB94, BB93, BCL00, Uhl95a, WST95]. MIMD/DMMP [BB93]. MiMPI [GCC99]. mini [LBG^{+}20, SCJH19]. mini-application [SCJH19]. mini-applications [LBG^{+}20]. MINIME [DS16]. MINIME-GPU [DS16]. minimization [POL99]. minimize [AiIS^{+}21]. Minimum [KA95, Wu99, GKD^{+}18, NCKB12]. Mining [BBD^{+}20, MA09]. minisweep [SCJH19]. Mississippi [IEE94f, IEE95j, IEE94f, IEE95j]. mitigating [OdSSP12]. Mitigation [BBH… 13a]. Mitsubishi [Ano03]. mittels [Wil94]. Mixed [ASA97, BEG^{+}10, CF01, OPP00, ST02a, MRH^{+}96, SK00, SB01]. Mixed-Mode [BEG^{+}10]. Mixing [CP98, GAP97, HDW21, CBYG18]. mixture [EO15]. MK [NS91]. MLP [JLG05]. mm par2.0 [OKM12]. MN [Ano94h]. Mob [STV97]. Mobile [ITT02, TWLL19]. Mode [BGK08, Bri02, BEG^{+}10, LRT07, HHSM19, SB01, YX95]. Model [AP96, BGG^{+}02, BdS07, CKmWH16, Cha02, CZG^{+}08, Dar01, DFA^{+}09, FSXZ14, FBSN01, GLB00, GLRS01, HLP11, KD12, LWKA15, LWZ18, LGG16, LPJ98, LA02, LRQ01, MKW11, NSLV16, NO02b, PRQ21, Ran05, RSV^{+}05, RRBL01, SPM^{+}10, SB95, SPH^{+}18, THN00, VT97, Wal01a, WYZ^{+}19, YCA18, AL93, BSC99, Bir94, BG94b, BDV03, CMV^{+}94, CL93, CKP^{+}93, ED94, GKZ12, GCN^{+}10, GkLyCY97, GWVP^{+}14, GRTZ10, HPLT99, HK09, HK10, HY20, JPL22, KOS^{+}95a, KSL^{+}12, KLV15, LR06b, LA06, LLH^{+}14, Mar05, MMAH20, MdSAS^{+}18, MSZG17, MGC^{+}15, NO02a, Nak05a, PAdS^{+}17, PQR18, QM21, RAS16, RGDML16, RCG95, Sch93, SH94, Sch99, SMAC08, Str94, VBLvdG08, Vis95, Wan02, WC15, WLK^{+}18, WYLC12, YX95, ZWC21, TA14]. Model-Based [AP96, LGG16]. Modeling [ACM96a, ATM01, BS07, COE20, CSC96, CDM93, FST98a, GAM^{+}02, HSO^{+}21, MOL05, MZLS20, MH21, NM95, RGDM15, Rót19, SEF^{+}16, STH22, TD99, VFD02, WJA^{+}19, WMC^{+}18, XH96, AC07, BDP^{+}10, BAE22, Bic95, BB95b, JL18, KM10, KME09, KEGM10, LZHY19, MS99a, WT13, XXL13, YMYI11]. Modelling [FST98b, GC05, Ham95a, KDL^{+}95b, BJS99, HTHD99, KDL^{+}95a, MSML10, QHCC17]. Models [AKK^{+}94, BS93, BZ97, CMK00, Cer99, CNM11, DK06, EMO^{+}93, ESM^{+}94, GJN97, PPF89, SS01, SMOE93, SYL19, TSN21, Whi04, BB95a, CPM^{+}18, CH96, CBS18, Duv92, EVMP20, KO14, LV12, MCB05, Nes10, RSBT95, RBAI17, RJH^{+}20, STP^{+}19, SYR^{+}09, Wal00, WBSC17]. moderate [Uhl95a]. Modern [AHHP17, DARG13, KDT^{+}12, LNK^{+}15, MPZ21, SM07, EYP^{+}20, HH14, HCC^{+}20, PMZM16]. modernization [WLYL20]. modes [WZWS08]. Modified [Riz17, GP95, KD12]. Modular [CT02, HPP02, FWS^{+}17, HLM^{+}17]. modulator [WWZ^{+}96]. modulator/DFB [WWZ^{+}96]. Module [Ano98]. Modules [AKK^{+}94, DS96b]. modules-design [DS96b]. Molecular [ABG^{+}96, BST^{+}13, BCGL97, BL95, BS07, DR97, DI02, KBM97, LAFA15, MH01, SA93, YWCF15, ZB94, AiIS^{+}21, BvdSvD95, BBK^{+}94, BMPZ94b, BMPZ94a, CC00b, DCD^{+}14, Dab19, FHSO99, HHS18, JAT97, JMS14, KFA96, KRG13, LHZ^{+}20, LSVMW08, OKM12, PARB14, PIR^{+}20, SL95, VGP^{+}19, ZWL13, RS22]. molecule [ART17]. Møller [BL95, KN17]. Moment [SSB21]. MONC [BBW19]. Monito [SGL^{+}00]. Monitor [KRS99, Whi94]. Monitoring [AH00, BCLN97, Beg93b, BFM96, BFMT96b, CD98, DBK^{+}09, GSN^{+}01, IADB19, LY93, LW97, MWG97, MVY95, SGL^{+}00, UP01, Wis98, Wis01, Yan94, Beg92, Beg93c, Beg93a, BB94, BS96a, BFMT96a, FLB^{+}05, LC07]. Monodomain [ORA12]. Monona [ZL18]. Monte [HJBB14, RP95, WH96, ADRCT98, AK99, DAK98, NSLV16, RR00, SK00, SKM15, ZZ04]. Monterey [Ano89, Gat95, USE94]. Montpellier [DE91]. Montréal [Lev95]. MOPS [GJN97]. Morehouse [AGH^{+}95]. Morgan [SD13]. Morphable [ZL17]. morphology [VLSPL19]. Morton [LZH18]. MOSFETs [MV20]. MOSIX [BBGL96]. motif [FMS15]. motors [SKM15]. movement [MV17, PG18]. Moving [HAA^{+}11, KQT^{+}21, LSG12]. MPC [BPJ22]. MPE [GKL95, KFA96]. MPEG [NU05]. MPEG-4 [NU05]. MPI [ARYT17, AD98, Ano95c, Ano99a, Ano99c, Ano99b, Ano99d, Ano00a, Ano00b, BDW97, CHD07, CHD09, CD01, CDND11, DKD05, DLM99, DKP00, DLO03, GBR97, GEW98, IEE96i, JKN22, JMS14, KGRD10, Kra02, KKD04, LKD08, MTWD06, Nag05, Per97, PS01b, RWD09, RLVRGP12, SBG20, ST02a, TDB00, TBD12, Vre04, WSN99, YM97, ST02b, ACGdT02, AKB^{+}19, Ada97, Ada98, AC07, ACH^{+}11, APJ^{+}16, AASB08, ART17, ATM01, ACC^{+}21, ACGR97, AK99, ABF^{+}17, AHP01, ACMZR11, ALW^{+}15, ALB^{+}18, ADLL03a, ADLL03b, And98, AiIS^{+}21, FH98, AVA^{+}16, Ano93e, Ano94d, Ano98, Ano01a, Ano03, AKE00, AKL99, AJF16, AIM97, ADR^{+}05, AHHP17, AMC^{+}19, Bad16, BV99, BCMR00, Bak98, BF98, BCFK99, BBG^{+}10, BCG^{+}10, BBG^{+}11, BKK20, BGBP01, BBS99, BAC20, BBG^{+}14, BA06, BCAD06, BADC07]. MPI [BGR97a, BKGS02, Ben01, BW12, BHV12, BKH^{+}13, BIL99, BIC05, BBB^{+}20, BP98, BF01, BBCR99, BBDH14, BK96, BKdSH01, Bha98, BfDA94, BHLS^{+}95, BHS^{+}02, Bis04, BBH… 13a, BBH^{+}13b, BDB^{+}13, BIC^{+}10, BR04, BCM^{+}16, BTC^{+}17, BM00, Boo01, BBC^{+}02, BPJ22, BCH^{+}03, BHK^{+}06, BBC^{+}99, BBC^{+}00, BS96b, BMR02, Bri02, BRM03, Bri10, BMPS03, BS07, BBW19, BDL98, Bru95, BDH^{+}95, BDH^{+}97, Brü12, BLW98, BFBW01, BEG^{+}10, BCH^{+}08, BWV^{+}12, CdOO^{+}20, CGC^{+}02, CSW12, CGC^{+}11, CwCW^{+}11, CRE99, CE00, CRE01, CC10, CP98, CAHT17, CGJ^{+}00, CFKL00, CSS95, CGBS^{+}15, CGG10, CB00, CDMS15, CGS15, CBL10, CBB^{+}20, CBB^{+}21, CLE^{+}20, Cha02, CTBT21, CEGS07, CDP99, CCA00, CFDL01, CLL03, CGZQ13, CC17, CSAGR98, CNC10, CC00a, CGH94, CCSM97, CFMR95, CDD^{+}96, Coo95a, Coo95b, CFF^{+}96, CRGM14]. MPI [CRM14, CRGM16, CC99, CT02, CD96, CG99b, Cre16, DPS05, DPSD08, DMK19, Dan12, DSG17, DZ96, DZ98a, DR18, DK20, DW02, DLM^{+}17, DZ98b, Dem96, DPP01, DJJ^{+}19, DLB07, DSW96, DS96a, DRUE12, DKD07, DI02, DDN^{+}22, DL10, DCPJ12, DCPJ14, DPFT19, DAK98, DGG^{+}12, DGB^{+}14, DBB^{+}16, HD02a, DXB96, DOSW95, DWS^{+}21, DFSW19, DCH02, DH22, DBK^{+}09, EZBA16, EGH99, EDSV09, EYP^{+}20, ES11, FH97, FD96, FDG97a, FDG97b, FLD98, FD00, FBD01a, FBD01b, FGRD01, FBVD02, FD02a, FD02b, FD04, FCLG07, FB95, FB96, FB97, Fan98, FPY08, FA18, FFB99, FNSW99, FTVB00, FFP03, FLPG18, FGL^{+}20, FL21, FMS15, FHK01, FKH02, FSC^{+}11, FCS^{+}12, Fin97, Fin94, Fin95, FWNK96, Fin00, FLB^{+}05, FC05, FST98a, FST98b, FJK^{+}17, FKK^{+}96b, FKK96a, FGT96, Fos98, FHPS94a, FHPS94b, FHP^{+}94, FHP^{+}95, Fra95]. MPI [FWR^{+}95, FKLB08, FBSN01, FSLS98, FCS^{+}19, GBR97, GFD03, GFD05, GDC15, GVF^{+}18, GGGC99, GGCM99, Gao03, GGZ^{+}20, GBR15, GCGS98, GCC99, GCBL12, GGHL^{+}96, Gei00, GR07, GGL^{+}08, GJR09, GSI97, GBH14, GBH18, GGS99, GMA20, GR95, GLB00, GRW^{+}19, Gle93, GM13, GJMM18, GT01, GBH99, GFIS^{+}18, GHZ12, GSYT21, GÁVRRL17, GDMME22, GRRM99, GAMR00, GKS^{+}11, GB98, GMPD98, GPL^{+}96, Gra97, GEW98, GBS^{+}07, GLM^{+}08, GL92, GL94, GLS94, GL95a, GL95b, GKL95, GL95c, GL96, GLDS96, GL97c, GL97b, GHLL^{+}98, GL99, GLT99, GLS99, Gro00, GLT00b, GLT00a, Gro01a, Gro01b, Gro02a, GL02, Gro02b, GT07, GLT12, Gro12, Gro19, GPC^{+}17, GC05, GSY^{+}13, Gua16, GADM20, HGX^{+}22, HJ98, HC10, Har94, Har95, HL17, HCC^{+}20, Hat98, HO14, HD02b, HDZ^{+}20, HE02, Hem94, HZ96]. MPI [Hem96, HRZ97, HZ99, HEH98, HGMW12, HMK09, HPS^{+}12, HPS^{+}13, Hin11, HRR^{+}11, HDB^{+}12, HDB^{+}13, HDT^{+}15, HKN^{+}01, HMS^{+}19, HLOC96, HKT^{+}12, HJB^{+}21, HVSC11, HWX^{+}13, HM01, HCA16, HG12, HcF05, Hus98, Hus00, Hus01, HWW97, IDS16, IRU01, ITKT00, IPG^{+}18, ICC02, IMS16, JL18, JF95, JDB^{+}14, Jes93b, JJM^{+}11, JS13, JNL^{+}15, Jon96, JLG05, JR10, JSH^{+}05, KB01, KFA96, KS15a, KPW05, KW14, KWEF18, KD12, Kan12, KTAB^{+}19, KLH^{+}20, KFL05, KB98, KK02a, KL94, KYL03, KYL05, KSJ95, KSJ96, KN17, KBS04, KGK^{+}03, KTXP21, KHB^{+}99, KBM97, KLR^{+}15, KR09, KSB^{+}20, KMG99, KEGM10, KRC17, KV98, KAC02, KC06, KBG16, KMH^{+}14, KRG13, LK14, LAdS^{+}15, LRG^{+}16, LLRS02, LTDD14, LGM00, LRT07, LC97a, LR06b, LTRA02, Lee12, LFS^{+}19, LFW20, LZ97, LRW01, LPD^{+}11, LLC13, LZH17, LZH18]. MPI [LK20, kLCC^{+}06, kLCCW07, kL11, LZZ^{+}20, LFL11, LS10, LSM^{+}18, LZC^{+}20, LCY96, LCW^{+}03, LVP04, LWP04, LGG16, LBB^{+}21, LYSS^{+}16, LB96, LGMdRA^{+}19, LMG17, LCMG17, LBB^{+}19, LGM^{+}20, LNLE00, LO96, dLR04, LZHY19, LS08, LL01, LZC^{+}02, LKJ03, LCC^{+}03, LKYS04, LSK04, LLH^{+}14, MBBD13, MMR99, MS02a, MS02b, MV17, MC18, MTK16, Man01, Man98, MK17, MLVS16, MB21, MLAV10, MKP^{+}96, MSMC15, MSL12, MH01, MSL96, MS96a, MC98, MGG05, MAS06, MM02, MM03, MOL05, MCS00, MANR09, MRRP11, MG97, MMDA19, MMAH20, MMM13, MTW07, MK04, MCLD01, MMH98, MMH99, MS99c, MB00, MvWL^{+}10, NAW^{+}96, NO02b, NO02a, Nak05a, Nak05b, NSBR07, NE98, NE01, Nes10, NSS12, NH95, NCB^{+}12, NCB^{+}17, NWT21, NAJ99, NW98, Nit00, NHT02, NHT06, NFG^{+}10, NN95, OM96, OLG^{+}16, OKM12, OIS^{+}06]. MPI [OD01, OF00, Ong02, OP98, OL05, OGM^{+}16, OMK09, Pac97, PARB14, Pan14, PK98, PES99, PLK^{+}04, PSK08, PDY14, PS00a, PS01a, PHJM11, PTL^{+}16, Per99, PZ12, PGK^{+}10, PFG97, PLR02, PGAB^{+}05, PGBF^{+}07, PGAB^{+}07, Pla02, PD11, PSSS01, PSK^{+}10, PTH^{+}01a, PTH^{+}01b, PS00b, PHM^{+}22, PTW99, QB12, QM21, Qui03, Rab98, Rab99, RDMB99, RR01, Ram07, RSBT95, RMS^{+}18, Ran05, RA09, RAS16, RCFS96, RJH^{+}20, RBB97a, RBB97b, RBB97c, RSPM98, RTH00, RH01, Reu01, RST02, Reu03, RGDM15, RGDML16, RGGP^{+}18, RGP22, RNPM13, RPM^{+}08, Röh00, Rol08b, RšT06, RSC^{+}19, RFRH96, RRG^{+}99, RTRG^{+}07, SE02, SCB14, SCB15, STP^{+}19, SPM^{+}10, SWCB20, SSB^{+}05, Sap97, SSB^{+}16, SDJ17, SGH12, SSN^{+}21, SBF^{+}04, SCJH19, SW12, SBG^{+}02, SG05, Ser97, SS01, SWS^{+}12, SG12, STY99]. MPI [SM02, SM03, SC19, SPH^{+}18, SP99, SZ11, SC04, SSC96, SS99, SIC^{+}19, SZBS95a, SZBS95b, SDN99, SvL99, SJ02, SWJ95, SMTW96, SH96, SDB94, SLG95, SDV^{+}95, SPH96, Slo05, SVC^{+}11, SK00, SB01, SOHL^{+}96, SOHL^{+}98, Sni18, SHHC18, SSL97, Squ03, Ste96, ST97, Sto98, SU96, Str96, SRS^{+}19, Sum12, SN01, Swa01, TOTH99, TAH^{+}01, TSY99, TSY00, THDS19, TSCS14, TKP15, TK19, Tha98, TGL02, TG09, TGKL19, TPLY18, TW01, TD99, TOC18, Tra98, THRZ99, TRH00, Trä02b, Tra02a, TGT10, Trä12a, Trä12b, THMH21, TMPJ01, TFGM02, Tsu07, TFZZ12, TPV20, UTY02, URKG12, VFD02, VLSPL19, VS00, VPS17, VSRC94, VSRC95, VGRS16, VdS00, VP00, VVD^{+}09, WH96, Wal95, WO95, Wal96a, WD96, WO96, Wal01a, Wal01b, Wal00, WC09, WLNL03, WLNL06, Wer95, WST95]. MPI [Whi04, WK20, WLR05, WWZ^{+}96, Wis98, WB96, WM01, WADC99, Wor96, WRA02, WDR^{+}19, WCS99, WT11, WYLC12, WT12, WLYC12, WT13, WMP14, XH96, XLW^{+}09, YM97, YL09, YHL11, YWC11, YCL14, YBMCB14, YW21, YPAE09, YTH^{+}12, YSP^{+}05, Zah12, ZZ04, ZLZ^{+}11, ZWZ05, ZLP17, ZJDW18, ZWLZ21, ZCBD22, ZLL^{+}12, ZGZS20, ZT20, ZWC21, ZZ95, ZSnH01, ZKRA14, ZA14, bT01a, dlAMCFN12, KH96, Mar06, YM97, Ano96a, Ano99a, Ano99c, Ano99b, Ano99d]. MPI-1 [SOHL^{+}98]. MPI-2 [Ano99c, Ano99d, Ano00a, AKL99, BCAD06, BHS^{+}02, CwCW^{+}11, CD96, DPSD08, GFD03, GGHL^{+}96, GT01, GHLL^{+}98, GLT99, GLT00b, GLT00a, HGMW12, LSK04, MS02a, MK04, PS00a, SS99, SSL97, TRH00, bT01a, BADC07]. MPI-3 [FCS^{+}19, GBH14, GBH18, GLT12, HDT^{+}15, QM21]. MPI-ACC [APJ^{+}16]. MPI-AllReduce [NWT21]. MPI-AMRVAC [KTXP21, TK19]. MPI-Based [Ada97, FSC^{+}11, RDMB99, SM03, Ada98, AVA^{+}16, GKS^{+}11, Gra97, LRW01, LZC^{+}20, OLG^{+}16, OP98, SZ11, TSCS14, TMPJ01]. MPI-basierte [Gra97]. MPI-benchmark [Reu01]. MPI-CHECK [LCC^{+}03]. MPI-CUDA [DR18, YW21, dlAMCFN12]. MPI-DDL [FB97]. MPI-Delphi [ACGdT02]. MPI-dot2dot [GDMME22]. MPI-driven [Hin11]. MPI-F [FHPS94b, FHP^{+}94]. MPI-FM [LC97a]. MPI-FT [LNLE00]. MPI-GLUE [Rab98]. MPI-GPU [TPV20]. MPI-Hybrid [CGC^{+}11]. MPI-I [IRU01, Tsu07]. MPI-I/O [IRU01, Tsu07]. MPI-interoperable [YBMCB14]. MPI-IO [BIC^{+}10, CGC^{+}02, CFF^{+}96, DL10, FWNK96, FSLS98, LRT07, LGG16, PSK08, PTH^{+}01a, SW12, Sto98, TGL02, ZZ04]. MPI-IO/GPFS [PTH^{+}01a]. MPI-LAPI [BGBP01]. MPI-Level [LVP04]. MPI-like [CGJ^{+}00]. MPI-only [LS10]. MPI-OpenCL [JNL^{+}15]. MPI-OpenMP [MS02b]. MPI-Parallel [DK20]. MPI-parallelized [DFSW19, KMG99]. MPI-Performance-Aware-Reallocation [GFIS^{+}18]. MPI-StarT [Hus98]. MPI-The [Ano99c, Ano99d]. MPI-thread [IDS16]. MPI-Umgebung [GBR97]. MPI/CUDA [PHJM11]. MPI/GAMMA [CC00a]. MPI/GPU [EZBA16]. MPI/GPU-code [EZBA16]. MPI/MBCF [MMH99]. MPI/OpenACC [OGM^{+}16]. MPI/OpenMP [ADR^{+}05, GÁVRRL17, HDZ^{+}20, HKN^{+}01, JLG05, JR10, KS15a, KN17, KLR^{+}15, KRG13, LLRS02, MMDA19, PZ12, SB01, WT11, WT12, WT13]. MPI/PVM [ES11]. MPI/RT [SKD^{+}04]. MPI/RT-1.1 [SKD^{+}04]. MPI/SMPSs [MLAV10]. MPI1 [Sti94]. MPI2 [MPI98a, MPI98b, Wal96b]. MPI2007 [MvWL^{+}10]. mpi4py [DF21]. MPI Allgather [GMdMBD^{+}07]. MPI Connect [FGRD01]. MPI T [GVF^{+}18, HHK^{+}19]. MPICH [BBC^{+}02, BCH^{+}03, BHK^{+}06, Cot98, Cot04, GL97a, KTF03, LKJ03, OPM06, OF00, RFG^{+}00, RšT06, SBG^{+}02, TRG05]. MPICH-CM [SBG^{+}02]. MPICH-G2 [Cot04, KTF03, OPM06]. MPICH-GQ [RFG^{+}00]. MPICH-V [BBC^{+}02, BHK^{+}06]. MPICH-V2 [BCH^{+}03]. MPICH2 [BMG07, Gro02b, ZSG12]. MPIConnect [FLD98]. mpicroscope [Trä12b]. MPIGeneNet [GDM18]. mpiJava [BCFK99]. MPINE [Sou01]. MPIPOV [FFB99]. MPIT [HIP02]. MPIWiz [XLW^{+}09]. MPJ [CGJ^{+}00]. MPL [XH96]. MPL0* [CRD99]. MPP [CDJ95, DOSW96, GBR97]. MPP-Systeme [GBR97]. MPPs [BGR97a, RBB97a]. MPSoC [KKJ^{+}08, KH10, PSM^{+}14]. MPSoCs [MB12, NEM17, SPB^{+}17]. MPVM [CCK^{+}95]. MRI [LSSZ15]. MRO [MMM13]. MRO-MPI [MMM13]. Multi [Ada98, ABB^{+}10, Bri10, BCKP00, CAWL17, CZG^{+}08, COE20, DK20, DS22, DWL^{+}10, EBKG01, FSXZ14, HD02b, HRZ97, JCH^{+}08, JNL^{+}15, KBA02, KT02, LTS16, LCY19, LM13, MLGW18, MG15, MB00, NMS^{+}14, PZ12, RG18, RR02, Smi93a, ST02a, ST02b, SSB^{+}17, TPV20, WBH97, XR21, YGH^{+}14, ZL18, ACMZR11, AGMJ06, BBC^{+}19, BCK^{+}09, CdOO^{+}20, DCH02, DWL^{+}12, Fin94, Fin95, FHB^{+}13, HTA08, HE15, JR13, JJM^{+}11, JR10, KSG13, KLV15, KO14, Kom15, LSG12, LS10, LLH^{+}14, MALM95, NSM12, SCB15, SFSV13, SVC^{+}11, SAP16, Str12, TS12b, TFZZ12, VLSPL19, WCC^{+}07, WO09, WADC99, WYLC12, ZAFAM16, ZWZ^{+}95, ZZZ^{+}15, SAP16, SG14]. multi- [ACMZR11, BBC^{+}19, CdOO^{+}20, KSG13]. multi-/many-core [KSG13]. multi-accelerator [KLV15]. multi-agent [ZWZ^{+}95]. Multi-agents [KBA02]. Multi-Array [LTS16]. Multi-cluster [ST02b, KO14, Kom15]. Multi-Context [ZL18]. Multi-Core [ABB^{+}10, Bri10, CZG^{+}08, YGH^{+}14, PZ12, FHB^{+}13, HTA08, JR13, JJM^{+}11, JR10, LLH^{+}14, SFSV13, SVC^{+}11, TFZZ12, WCC^{+}07, WYLC12]. multi-cores [WO09]. multi-CPU [SAP16]. multi-CPU/multi-GPU [SAP16]. Multi-Dimensional [HD02b, KT02, RG18]. multi-endpoint [LLH^{+}14]. Multi-GPU [DS22, JNL^{+}15, NMS^{+}14, XR21, NSM12, TS12b, SAP16, SG14]. multi-kernel [SAP16]. Multi-level [CAWL17, LCY19, LM13, HE15, MALM95, ZZZ^{+}15]. multi-morphology [VLSPL19]. Multi-Network [BCKP00]. Multi-Node [HRZ97]. multi-petaflops [LSG12]. multi-phase [ZAFAM16]. Multi-Physics [WBH97]. multi-place [BCK^{+}09]. Multi-platform [DWL^{+}10, DWL^{+}12]. Multi-Processing [MLGW18]. Multi-Processor [RR02, Smi93a, DCH02]. multi-programming [WADC99]. Multi-protocol [MB00]. Multi-Resolution [TPV20]. Multi-Socket [COE20, LS10]. Multi-Stage [FSXZ14]. Multi-Threaded [MG15, Ada98, EBKG01, SCB15]. Multi-Threading [MLGW18]. multi-valued [Str12]. Multi-Vectors [DK20]. Multi-versioned [SSB^{+}17]. multi-zonal [Fin94, Fin95]. Multi-Zone [JCH^{+}08, AGMJ06]. Multiblock [IDD94, DLR94]. Multicast [CCA00, CDPM03, ZGN94]. Multicasting [SE02]. multicenter [CwCW^{+}11]. MultiCL [APBcF16]. multicomputer [SWJ95, TD99]. multicomputers [HWW97, Yan94, YX95]. Multiconference [Ten95]. Multicore [BDT08, CGC^{+}11, CB16, DS16, DGH^{+}19, GDM18, KDT^{+}12, LNK^{+}15, WT12, YKW^{+}18, ASB18, CLYC16, GJLT11, GDMME22, HWX^{+}13, JPOJ12, KN17, LS10, MBBD13, MM11, Nob08, OPW^{+}12, PDY14, QB12, RGDML16, WCS^{+}13, WT11, WLYC12, WT13, YHL11, YWC11, dlAMC11, BAE22]. multicore/many [MBBD13]. multicore/many-core [MBBD13].

Multicores

[Ger18, GDDM17, ADK22, UGT09].

multidestination [Pan95a].

multidimensional

[CSW99, DMK19, PDY14, ZT17]. multidisciplinary [Fin94, Fin95]. multifold [PIR^{+}20]. multifrontal [IM95]. Multigrain [AZG17, IOK00]. Multigrid [BCMR00, SSK^{+}18, AGIS94, IHM05, Lou95, Mic93, Mic95, PSLT99, RM99, Sta95a, TK19, ZZG^{+}14]. Multigroup [QRG95, QRMG96]. Multilevel [JLG05, PSSS01, BAV08, ETV94, GAM^{+}00, JJY^{+}03]. multimedia [GFB^{+}14]. multimethod [FGT96]. Multiobjective [RLVRGP12]. Multiparadigm [FS98]. Multiphase [SPH^{+}18]. Multiphysics [NPS12]. Multiplatform [SMM^{+}16]. Multiple [BSG00, CB16, FGKT97, FBSN01, JPT14, JSH^{+}05, KMM15, LTR00, NTR16, Pet01, Tsu12, ZC10, Zho21, AML^{+}99, ESB13, GM18, KGB^{+}09, KKLL11, SHHC18]. Multiple-Precision [ZC10, JPT14]. Multiplication [AKL16, DS13, Fuj08, TQDL01, FAF16, FJZ^{+}14, XXL13]. Multipole [AAB^{+}17, AiIS^{+}21, KKB^{+}21, LCL^{+}12, YBZL03]. Multiported [SG15]. Multiprocessing [MW93, VGS14]. Multiprocessor [Pet97, ABCI95a, ABCI95b, ADMV05]. MultiProcessors [BDV03, CC99, HPP02, NPP^{+}00d, SBW91, SS01, Tra98, JE95, KC06, SYR^{+}09, AGIS94]. multiprogrammed [TSY99]. Multiprogramming [BHP^{+}03]. Multiprotocol [BHK^{+}06]. Multirail [LVP04]. multiscale [CwCW^{+}11]. multiservice [CLLASPDP99]. multisource [ZDR04]. multistage [ZGN94]. Multistart [Cza13]. Multitasking [ZGNZ22, FH95]. multithread [GCC99, SWYC94, ZG98]. multithread-safe [GCC99]. Multithreaded [ALB^{+}18, AZG17, DGG^{+}12, PS01b, RBAA05, TGBS05, WJ12, DSG17, TMC09, TG09, WCC^{+}07]. Multithreading [BBG^{+}10, ZWL13]. Munich [BDLS96, GH94]. Mushy [Wit16]. MUST [HPS^{+}12, HPS^{+}13]. mutual [She95]. MV [TWLL19]. MV-Net [TWLL19]. MVAPICH [RMS^{+}18]. MVICH [OF00]. Myocardial [Pat93]. Myrinet [CDP99, GBH99, JSH^{+}05, LCW^{+}03, PTW99, Tou00].

n [DDN^{+}22, Pan95a, ADB94, RTRG^{+}07]. N-body [ADB94, RTRG^{+}07]. n-cube [Pan95a]. NAG [DHP97, For95, McD96]. NAMD [PZKK02]. Naming [MSF00]. Nancy [BR95a]. NanosCompiler [GAM^{+}00]. Narrow [YSS^{+}17, YSS^{+}19]. NAS [CRE99, CE00, CCF^{+}94, CDD^{+}96, KS96, KAC02, MMH99, WAS95b, WT11, WT12]. NASA [MAB05]. NASLU [PHJM11]. National [Str94, BRST94]. Native [SZ99]. NATO [KG93, TG94]. NATUG [Ara95]. NATUG-7 [Ara95]. nature [DSM94]. Navier [Che99, DLR94, HSMW94, IDD94, Lou95, SCC95]. NB [BG91]. NC [Agr95a, SL94a]. NCCL [AMC^{+}19]. NCCL2 [AMC^{+}19]. NCS [AL92]. nCUBE2 [BL94]. Near [PKYW95]. Nearest [DI02]. Nearest-Neighbor [DI02]. Nebelung [MFG^{+}08]. NEC [GPL^{+}96, HRZ97, TRH00]. Necessary [NPP^{+}00b]. Needed [Gei00]. Negative [KF16]. Neighbor [DI02]. neighborhood [HS12]. Nek5000 [MGS^{+}15, OGM^{+}19]. Nekbone [GML^{+}16]. Nemesis [BMG07]. Nesbet [BL95]. Nested [AHD12, BR12, BS01, DLRR99, DSCL05, GLP^{+}00, HA10, MMS07, SGL^{+}20, TTSY00, ZLP17, aMST07, AGMJ06, BS05, HSE^{+}17, HY20, LW20, THH^{+}05, YZ14, JLG05]. Nesting [BBC^{+}99]. Nests [DMB16]. Net [CNM11, NE98, NE01, PES99, TWLL19]. Net-Console [PES99]. Net-dbx [NE98, NE01]. netCDF [LkLC^{+}03].

Netherlands [DSZ94, Ano93f, Van95]. Nets [Sou01, Str94]. Network [ACM98a, AR01, BDG^{+}91b, BDG^{+}93a, BCKP00, CZ95a, CDHL95, CSC96, DM95b, DM95a, DBA97, DFMD94, DGMS93, DGMJ93, EK97, Fer98b, Fis01, GS91b, GS92, Gei93a, GSxx, Hus98, ITT02, LB98, LH95, MSCW95, MANR09, OF00, OWSA95, RJ21, TW01, VZT^{+}19, AL92, AH95, AVA^{+}16, BDG^{+}92a, BDG^{+}92c, BDG^{+}94, BSvdG91, BJ95, Bon96, BBK^{+}94, BID95, BFM96, Coe94, CLLASPDP99, Fer98a, GS91a, Gei93b, GK97, GHZ12, HBT95, HK94, HH95, IM95, KMC96, KMC97, KA95, LH98, LK20, LHD^{+}94, LHD^{+}95, MK94, MRH^{+}96, POL99, PR94c, PTW99, Rag96, SEC15, SPK^{+}12, TSS98, YS93, ZPLS96, GK97]. Network-Balancing [DBA97]. Network-Based [BDG^{+}91b, GS92, BDG^{+}92a, IM95]. Network-Specific [DM95b, DM95a]. network-topology-aware [SPK^{+}12]. Networked [FGKT97, GBD^{+}94, Nov95, NMC95, Per96, Ano95b, BMPZ94b, BMS94a, BMPZ94a, GM94, HS93, RRG^{+}99]. Networking [ACM97b, ACM98b, ACM00, ACM01, ACM04, Hol12, LCK11, CXB^{+}12, GH94, HS95a, ITT99, LCHS96, MZK93]. Networks [CSV12, CDM93, DDP^{+}19, DDPR97, GFV99, GDM18, GHL97, HHK94, HLCZ00, HIP02, LHHM96, Li96, LHZ98, MBES94, QMGR00, SG15, SM19, TQDL01, Tou00, VLO^{+}08, VBB18, WAS95b, WMC^{+}18, BK11, BRS92, CZ95b, CFPS95, DG95, DZ98a, Jou94, LR06a, LTLC94, LHD^{+}94, LHD^{+}95, NFG^{+}10, Pan95a, SOYHDD19, TDB00, ZGN94]. Neural [AGH^{+}95, CAM12, CSV12, QMGR00, RJ21, SM19, Str94, GkLyCY97, Rag96]. Neurocomputing [PSZÉ00]. Neutral [CBB^{+}21]. neutrino [KHBS19]. Neutron [LD01, RS97, VRS00, WR01, MM92]. Nevada [Ano94e]. never [Har94]. Neville [ACMZR11]. Newport [IEE93b]. News [Ano97, Ano03, Bra97, ESB13, KS15a, Str94]. Newton [AEW^{+}20, ZB97]. Next [GKPS97, Gei98, Gei01, VPS17, VZT^{+}19, EYP^{+}20, SP11, ZKRA14, vdP17]. Next-Generation [VPS17, ZKRA14]. NFS [CGC^{+}02]. NHPDCC [BRST94]. NIC [MFPP03]. NIC-based [MFPP03]. Nice [ACM90]. nineteenth [IEE95l]. Ninth [ERS96, R^{+}92]. NIST [SNMP10]. Nitzberg [Ano99c, Ano99d]. NLP [VB99]. NM [IEE95d, Old02]. NMF [KF16]. nmfgpu4R [KF16]. NoC [HWX^{+}13]. NoC-based [HWX^{+}13]. Node [HRZ97, KLH^{+}20, KFL05, FKLB08, GM13, Gro19, JR10, LFL11, MKP22, RS19, Zah12]. Nodes [BBC^{+}02, BCH^{+}03, DBK^{+}09, JNL^{+}15, MKC^{+}12, BBB^{+}22, VGP^{+}19]. Noise [SAL^{+}17]. Non [BCG^{+}10, CTBT21, CCSM97, Gua16, HTA08, KLH^{+}20, KF16, MW98, Man01, SdR^{+}21, WLNL03, WTR03, FH98, BCH^{+}08, OKW95, OMK09, STP^{+}19, TVCB18, WLNL06]. Non-blocking [HTA08, FH98, BCH^{+}08, STP^{+}19]. Non-Contiguous [KLH^{+}20, WTR03]. Non-Data-Communication [BCG^{+}10]. non-dedicated [WLNL06]. Non-Determinism [CTBT21]. Non-Intrusive [SdR^{+}21]. non-iterative [OMK09]. Non-linear [MW98, OKW95]. Non-Local [CCSM97]. Non-Negative [KF16]. Non-persistent [Man01]. non-singleton [TVCB18]. Non-stop [Gua16]. nonaligned [AGIS94]. nonblocking [DJJ^{+}19]. Noncontiguous [JDB^{+}14, TGL02]. Nondeterminacy [DKF93]. nondeterminism [Obe96]. Nondeterministic [KSV01, CRD99]. nonequispaced [YW21]. Nonintrusive [TGS^{+}20]. Nonlinear [Nak03, Was95a, ZB97, CEGS07, Jou94, NS20]. nonnegative [KBP16]. nonsymmetric [dH94]. Nordic [FF95]. Norfolk [Sin93]. normal [CBS18]. normalized [Gra09]. Norman [Edd18]. North [CJNW95]. Note [BR02, Cre16, SGHL01]. notification [SSN^{+}21]. Notre [IEE96i]. novel [DDYM99, GKK09, MLVS16, MSL12, QM21]. November [ACM96c, ACM97b, ACM98b, ACM99, ACM00, ACM01, ACM03, ACM04, ACM05, Ano94c, ACDR94, BDW97, GN95, HK95, Hol12, IEE91, IEE93e, IEE94b, IEE94h, IEE02, LCK11, USE94]. novice [CGG10]. Novices [Stp02]. NOWs [SLGZ99]. NP [YZ14]. NPACI [PKB01]. NPB [EGC02]. NR [Gua16]. NR-MPI [Gua16]. NRC [LD01]. NScluster [TSN21]. NSGA [GÁVRRL17]. NSW [GN95]. NT [Ano01a, Bak98, BF98, CLP^{+}99, FD97, GGGC99, PS00a, SFG98, TAH^{+}01]. NTRUEncrypt [KY10]. NTUG [FF95]. Nuclear [BPG94, GA96]. nuclei [NS16]. NUMA [BCC^{+}00a, BCC^{+}00b, BFG^{+}10, CAWL17, GTS^{+}15, MKC^{+}12, MMAH20, MJB15, OPW^{+}12, SLN^{+}12, TSCaM12, ZLP17]. NUMA-aware [MMAH20]. NumaGiC [GTS^{+}15]. Numba [BS21]. Number [BP99, HT08, WHDB05, CCS19, CBYG18, Lan09, Stp20]. Numeric [MLGW18]. Numerical [ACMR14, BS93, BCP^{+}97, CSW97, DHK97, DHP97, FK01, For95, FB94, HH14, Hol95, Hus98, IFI95, KM10, Kha13, McD96, NS20, NHT02, PKYW95, TDBEE11, TPV20, YKLD17, AL92, Boi97, BCM^{+}16, CSW99, DFSW19, FP92, GS94, HD00a, JK10, KB13, Nob08, NHT06, Pri14, SMAC08, SU96]. Numerically [BKML95, BFLL99]. nur [BL94]. Nutzung [GEW98]. NVIDIA [GDS^{+}20, GNP19, KC19, KME09, Seg10, VLMPS^{+}18, XXL13, KKM15, Lan09]. NVRAM [MC18]. NX [Pie94, PR94a]. NY [IEE96f, PBG^{+}95, Ree96, SS96].

O [Bos96, CFF^{+}96, DRUE12, IRU01, IBC^{+}10, KLH^{+}20, LkLC^{+}03, kLCC^{+}06, LPJ98, MMD98, MV17, MC18, MGC12, MG15, NFK98, OWO98, PSK08, PLR02, RK01, SBQZ14, SR98, Tha98, Tsu07, WSN99, ZJDW18]. O2000 [CML04]. O2WebCL [CHKK15]. Oberammergau [BPG94]. Object [Ada97, BCFK99, CFKL00, FMSG17, MSL96, PD98, SWL^{+}01, YHGL01, YX95, Ada98, BR91, DM12, LKL96, OKM12, RFH^{+}95, SL94b, TDG13]. object-based [LKL96]. Object-Oriented [BCFK99, PD98, SWL^{+}01, Ada98, DM12, OKM12, RFH^{+}95]. Objects [KH15, Man01, MFC98, HS93, SOA11, SC95, YWO95, ZPLS96]. Oblivious [LZH17, LZH18, UALK17, UALK19, HSP^{+}13]. observations [ZKRA14]. observed [CAHT17]. Occam [ACDR94, GN95, MC94, EM94, SHH94a, SHH94b]. Ocean [BS93, GAM^{+}02, Bic95, Mal01, Nes10, Sch99, Wal00]. Oceans [IEE94c, IEE94c]. OCLoptimizer [FAFD15]. OCM [BoFBW00]. OCM-Based [BoFBW00]. October [Ano93f, Ano94e, Ano94i, Ara95, BPG94, Bha93, BDLS96, CHD07, CGB^{+}10, DSM94, DLO03, DE91, FK95, GGK^{+}93, IEE94f, IEE95a, IEE95g, IEE95j, IEE96b, IEE96c, IFI95, JB96, Kra02, Old02, OL05, Sch93, Sie92a, Sie92b, Tou96, USE00, UCW95, Vol93]. octree [JL18, TK19]. octree-based [JL18]. ODE [Ano97, Bra97]. ODEs [Pet97]. OdinMP [BB00]. OdinMP/CCp [BB00]. Off [CGS15]. Off-Line [CGS15]. Offering [EK97]. Official [Ano98]. Offload [BRU05]. Offloading [DFP^{+}19, MGA^{+}17, DSGS17, KBG16, MNYN21, SWCB20, TSEE21, TMT^{+}20, WZW21]. oft [Rol08a]. Oil [FSXZ14, ZAFAM16]. OKs [Ano03]. old [LK14]. OMB [BWV^{+}12]. OMB-GPU [BWV^{+}12]. OMIS [LW97]. Omni [KSS00, KSHS01]. OmniRPC [SHTS01]. OMP [SGJ^{+}03]. OMP2001 [TSB03]. OMP2012 [MBB^{+}12]. OMPI [ACH^{+}11, OM96]. OmpSs [ABF^{+}17, BAC20, PSB^{+}19, VLCM^{+}20, YÁJG^{+}15]. On-Chip [WYZ^{+}19, TDG13]. On-Demand [CTK00, LSB^{+}18]. On-GPU [LW20]. On-Line [BoFBW00, Wis98]. On-the-fly [KSJ14]. ONC [RS93]. One [BPS01, GFD03, GFD05, GBH14, GT01, HDB^{+}12, LRT07, MH01, TGT05, TRH00, ZSG12, bT01a, DPFT19, DBB^{+}16, GBH18, KW20, LSK04, MS99c, Ols95, PGK^{+}10, dlAMC11]. one-dimensional [Ols95]. one-layer [dlAMC11]. One-Sided [BPS01, GFD03, GFD05, GT01, HDB^{+}12, LRT07, MH01, TGT05, TRH00, ZSG12, bT01a, DPFT19, DBB^{+}16, LSK04, MS99c, PGK^{+}10]. one-step [KW20]. only [LS10, Squ03]. Ontario [GGK^{+}93]. onto [OFA^{+}15]. OOMPI [MSL96]. OOPS [RFH^{+}95]. OPAL [CwCW^{+}11, NW98]. OPAL-MPI [NW98]. opaque [SOA11]. Open [BGG^{+}15, KDL^{+}95b, WGG^{+}19, AVA^{+}16, KDL^{+}95a, LSB^{+}20, Nob08, GBS^{+}07, VGRS16]. Open-Source [BGG^{+}15, AVA^{+}16, LSB^{+}20, Nob08]. OpenACC [ACC^{+}21, CGK^{+}16, CCBPGA15, GML^{+}16, GM18, HTJ^{+}16, HY20, JCP15, KDHZ18, KLV15, Kom15, LLVM21a, LLVM21b, LBG^{+}20, LB16, LSG12, LHZ^{+}20, MGS^{+}15, OGM^{+}19, OGM^{+}16, QHCC17, RLFdS13, SCJH19, STH22, Stp20, VGP^{+}19, WLK^{+}18, XJR21, XR21, EVMP20]. OpenACC-based [KLV15]. OpenACC-like [HY20]. OpenACC-to-FPGA [LLVM21a, LLVM21b]. OpenCL [ABDP15, APBcF16, ASAK19, AB13, BLPP13, BBC^{+}19, BDW16, BN12, BS21, BHW^{+}12, BBH^{+}15, BAS13, CJPC19, CDD^{+}13, CP15, CLOL18, CZP21, CIJ^{+}10, CHKK15, CCS19, CCK12, CS14, CLBS17, CBIGL19, CBS18, DARG13, Di 14, DWL^{+}10, DWL^{+}12, FAFD15, FLMR17, FDG19, FE17a, FE17b, FSV14, FVLS15, dFdOSR^{+}19, GScFM13, GDDM17, HSO^{+}21, HHS18, HD11, HE15, HHC^{+}18, JSS^{+}15, JCP^{+}20, JKM^{+}17, JR13, JNL^{+}15, JMdVG^{+}17, KKM15, KH12, KM10, KKLL11, KSL^{+}12, KJJ^{+}16, KNH^{+}18, KB13, KPK13, Lee12, LWKA15, LNK^{+}15, LCH^{+}22, LWZ18, LL16, LAFA15, MC17, MKP22, MAIVAH14, MTU^{+}15, MSZG17, MZLS20, MHSK16, ÓN12, OTK15, ORA12, PS19a, PCY14, PHW^{+}13, PSB^{+}19, PSH^{+}20, PB12, RG18, RBW^{+}20, RBC20, RVKP18, RVKP19, RGD13, RBB15, RGB^{+}18, RRJ^{+}20, RBB17, SFSV13, SPB^{+}17, SAP16, SXMX^{+}18, SSB^{+}17, SG14, SFLD15]. OpenCL [SGS10, Str12, THS^{+}15, TSEE21, TK16, TMW17, TKP15, TY14, TL19, WTTH17, WHMO19, WZHZ16, WTS19, WQKH20, WYH^{+}21, YSWY14, YWTC15, YSL^{+}12, ZWL^{+}17, ZT17, dAT17, KB21]. OpenCL-accelerated [ZWL^{+}17]. OpenCL-Based [CLOL18, MZLS20, WTTH17, WZHZ16, JKM^{+}17, SXMX^{+}18, WHMO19]. OpenCL-like [TSEE21]. OpenCL-to-WebCL [CHKK15]. OpenCL-written [KNH^{+}18]. OpenCLC [LSB^{+}20]. openFabrics [FCS^{+}19]. OpenFOAM [TGS^{+}20]. OpenGL [Ano98, Bae20, LHZ97, ORA12, Rót19]. OpenGL- [Rót19]. OpenHMPP [AAB^{+}16]. openMosix [Slo05]. OpenMP [Cha05, CZG^{+}08, CGKM11, CMMR12, EV01, JMS14, MdSC09, SHM^{+}10, Vos03, OKM12, ST02a, ST02b, Add01, ARvW03, ABC^{+}00, AC07, AHD12, ADK22, AAB^{+}17, AELGE16, ACC^{+}21, ACMZR11, ATL^{+}12, ADT14, ACJ12, Ano97, Ano01b, Ano03, ABB20, AKE00, ADMV05, ADR^{+}05, ASB18, AML^{+}99, AGMJ06, AM07, ACD^{+}09, ABB^{+}10, BST^{+}13, BBB^{+}22, BR02, BAE22, BHP^{+}03, BME02, Ben18, BN00, BF01, BBDH14, BWW^{+}12, BCC^{+}00a, BCC^{+}00b, BGK08, BGG^{+}02, BS01, BS05, BBC^{+}99, BBC^{+}00, Bra97, Bri00, BDV03, BdS07, BGdS09, BFG^{+}10, BGD12, BC00, BS07, BB00, BC19b, BK00, BKO00, BO01, BEG^{+}10, BB18, CdOO^{+}20, CRE99, CE00, Car07, CB00, CGLD01, CDK^{+}01, CLYC16, CM98, CMZ99, CHPP01, CBPP02, Cha02, CM05, CJvdP08, CGKM11, CMMR12, CLA^{+}19, Cla98, CBYG18, CCM^{+}06, CCBPGA15, CC00b]. OpenMP [CF19, Dab19, DM98, DW02, DBVF01, DFP^{+}19, DKB20, DSGS17, HD02a, DGH^{+}19, DFC^{+}07, DFA^{+}09, ETWaM12, EBB^{+}20, EM00a, EM00b, EV01, EdS08, FGRT00, FMSG17, FSG19a, FSG19b, FSXZ14, FM09, GSA08, GJP01, GSMK17, GG09, Goe02, GÁVRRL17, GSM^{+}00, GAM^{+}00, GAML01, GOM^{+}01, GAM^{+}02, Gra09, HPP02, HP05, HDDG09, HA10, HO14, HD02b, HDZ^{+}20, HMK09, HASnP00, HKN^{+}01, HAJK01, HVSC11, HLCZ00, HT01, HCL05, HEHC09, HJYC10, HHSM19, HH22, HAA^{+}11, IJM^{+}05, ICC02, IOK00, ITT02, JCP15, JKHK08, JPOJ12, JFY00, JJY^{+}03, JCH^{+}08, JJM^{+}11, JLG05, JR10, KB01, KS15a, KOB01, KaM10, KOI01, KN17, KKH03, KT02, KSJ14, KLR^{+}15, KBVP07, KBG^{+}09, KQT^{+}21, KSB^{+}20, KKV01, KT10, KH15, KAC02, KC06, Kuh98, KPO00, KLM^{+}19, KRG13, KSS00, KSHS01, KJEM12, LOHA01, LP00]. OpenMP [LLRS02, LTS16, LBG^{+}20, LD01, LME09, LLC13, LHC^{+}07, LNW^{+}12, LRLG19, LHCW05, LYSS^{+}16, LA02, LA06, LdSB19, LMRG14, LHZ98, LL01, LLH^{+}14, MKC^{+}12, MS02b, Mal01, MV20, MM07, MB12, Mar02, Mar03, MLC04, Mar05, Mar09, MPD04, MCB05, Mat00a, Mat00b, Mat01a, Mat03, MGG05, MGC12, MG15, MM11, MFG^{+}08, MKV^{+}01, MBE03, MRRP11, MMDA19, MMSW02, MKW11, MM14, MMS07, MJB15, MJPB16, MCdS^{+}08, Mül01, Mül02, Mül03, MBB^{+}12, MBA21, NO02b, Nak05a, NIO^{+}02, NIO^{+}03, NEM17, NPP^{+}00b, NPP^{+}00c, NPP^{+}00a, NPP^{+}00d, NAAL01, NA01, NNON00, Nob08, NU05, NHT02, NHT06, OOS^{+}08, OP10, OPW^{+}12, PARB14, PPJ01, PVKE01, PK05, Per21, PZ12, PQR18, PRQ21, PGC02, PKE^{+}10, Qui03, Ran05, RDLQ12, RLVRGP12, RBAA05, SSE12, SSB^{+}16, SHHI01, SHTS01, SKS01, SLGZ99, SGZ00]. OpenMP [SPL^{+}12, SdR^{+}21, SHPT00, SSAS12, SK00, SB01, SBB20, SSB21, Stp02, Stp18, Stp20, SGL^{+}20, SGS^{+}21, Taf21, TCM18, TBS12, TS12a, TSB02, TTSY00, TSN21, TSS00a, THDS19, TSCaM12, TJPF12, Thr99, TBG^{+}02, THH^{+}05, TGBS05, TMT^{+}20, VLSPL19, VLCM^{+}20, VDL^{+}15, VPS17, VGS14, VGP^{+}19, Vos03, Vre04, Wal00, Wal02, Wan02, WCC12, WC15, WZW21, WJG^{+}21, WMK^{+}19, WPC07, WLYL20, WT11, WYLC12, WT12, WLYC12, WT13, YKW^{+}18, YHL11, YWC11, YCL14, YKLD17, YPAE09, YSVM^{+}16, YSMA^{+}17, YYW^{+}12, YCA18, ZAT^{+}07, ZT20, ZWC21, ZSnH01, aMST07, dCZG06, vdP17, RM99, SSGF00, WCS^{+}13, EVMP20]. OpenMP* [KDT^{+}12]. OpenMP-based [ABB20, LNW^{+}12]. OpenMP-like [BK00, BKO00, KOB01, VGS14]. OpenMP-oriented [MLC04]. OpenMP-parallel [HHSM19]. OpenMP-style [JPOJ12]. OpenMP/MPI [BEG^{+}10, HMK09, LLC13, LYSS^{+}16, MGG05, NO02b, Nak05a, SSB^{+}16, SK00]. OpenMPI [DS22]. OpenSHMEM [HVA^{+}16]. OpenTuner [BAG17]. OpenUH [HEHC09, LHC^{+}07]. Operating [MMH98, RGD97, TL19, USE94, Wil93, ARS89, Sei99]. operational [KOS^{+}95a]. Operations [BIL99, BIC05, CCA00, FCLG07, FPY08, GFD05, GLB00, PSM^{+}14, PGAB^{+}05, TRG05, TGT05, WRA02, ZLWW20, BMG07, DS13, HMS^{+}19, IDS16, KHB^{+}99, KMH^{+}14, LFW20, MB21, PGAB^{+}07, PKD95, SS99, TFZZ12]. Operators [DK20, KK19, NHT02, NHT06]. opportunistic [CC10]. Opportunities [LB16]. optical [MRH^{+}96]. Optimal [BP99, GAMR00, ZGN94, BB95a, ER12, PQ07, PTL^{+}16, Sur95a]. optimiertes [Sei99]. optimisation [AMuHK15]. Optimising [Boo01, FKH02]. Optimistic [SCL00, CXB^{+}12, PY95]. Optimization [AEW^{+}20, BSG00, BHNW01, DBA97, Goe02, HS12, Hus00, ITT02, KGK^{+}03, KMH^{+}14, LLVM21a, LLVM21b, LCY19, LdSB19, MC17, MBS15, Mül01, NIO^{+}02, NIO^{+}03, PSSS01, SM03, SvL99, SWH15, TRG05, WTTH17, WJ12, AMKM20, BMS19, Cou93, DSOF11, DH22, FCS^{+}12, HWS09, HDZ^{+}20, KHS12, LME09, LDJK13, MALM95, PP16, PS19a, PMM95, SKS01, SDJ17, Stp20, Str12, TMW17, TMT^{+}20, TFZZ12, VSW^{+}13, Was96, XXL13, XR21, ZWC21]. Optimizations [NSLV16, SSE12, iSYS12, TSS00a, BVML12, HLK^{+}20, HEHC09, LL16, MV17, SSH^{+}19]. Optimize [SdR^{+}21, BBW19, GVF^{+}18, GFIS^{+}18, WLYC12]. Optimized [AKL16, ABG20, AMC^{+}19, Bri02, FAFD15, MAIVAH14, PM95, PTH^{+}01a, THS^{+}15, THDS19, WJB14, BKvH^{+}14, EBB^{+}20, MMM13, Sei99]. optimizer [BHRS08, Rag96]. Optimizing [BGH^{+}05, CXB^{+}12, FMFM15, KKP01, MBE03, MZLS20, NSZS13, OM96, SSAS12, TGL02, TGT05, WK20, GS02, LHC^{+}07, RKBA^{+}13]. Options [RR00]. Orange [ACM98b]. orbit [CFF19, MBA21, SSN94]. Order [BL95, DFN12, LZH18, EVMP20, KN17, KME09, KEGM10, KB13, MYB16, OGM^{+}16, THDS19]. ordering [Zah12]. ordinary [NF94, RBB15, SP11]. Oregon [ACM99, IEE93e, SW91]. Organization [BPC94, JFGRF12]. Oriented [Ada97, BCFK99, FMSG17, LYGG20, MSL96, PD98, YHGL01, ZL18, Ada98, BR91, CJPC19, CBIGL19, DM12, HDZ^{+}20, MGC^{+}15, OKM12, RFH^{+}95, SWL^{+}01, MLC04]. Origin [LL01, LSK04, ZSnH01]. Origin2000 [Bri00, MH01]. original [RNPM13]. Orlando [ACM98b]. Orleans [IEE96b, USE95]. ORNL [Bor99]. Orthogonal [SSB21]. orthogonality [THMH21]. OSCAR [IOK00, Slo05]. oscillations [KHBS19]. oscillator [BJ13, GSMK17]. OSDI [USE94]. OSF [Sch93]. OSWALD [RGB^{+}18]. Other [OP10]. OtOt [DKF94b]. Otto [Ano96a, Ano99a, Ano99b, Nag05]. out-of-core [BL99]. Output [CFF^{+}94, HE02, JWB96]. Outstanding [LSB15]. Overcoming [JKHK08]. Overhauling [BDW16]. Overhead [BR02, DFP^{+}19, FST98a, XH96, CRGM16, KC94, KRS99, LZHY19, ZRQA11]. Overheads [BCG^{+}10, BGdS09, BCM11, SS94]. Overlap [ADGA20, BRU05, DCPJ12, DCPJ14, MLAV10, PSK08, SH14]. Overlap-and-Save [ADGA20]. Overlapped [GPC^{+}17]. Overlapping [KB01, kLCC^{+}06, PKE^{+}10, BBH^{+}15, DJJ^{+}19, MMM13]. overlay [BMS19, CXB^{+}12]. overlay-based [CXB^{+}12]. oversubscription [KC19]. Overview [CFF^{+}96, Gre95, GL95c, Zol93, GHZ12, GPL^{+}96, HHK^{+}19, Wer95]. OWL [JKN^{+}13]. Ownership [FHB^{+}13]. Oxford [Boi97].

P [CAM12, WHDB05]. P-RnaPredict [WHDB05]. P03M [BJ93]. P2P [GR07, GGL^{+}08, GJR09, RS19, SBG^{+}02]. P2P-MPI [GGL^{+}08, GJR09]. P4 [KS96, Mat94, Mat95]. PA [ACM04, Ham95a, ACM96c]. Pablo [BFMT96a, BFMT96b]. Pablo-based [BFMT96a, BFMT96b]. Pacific [IEE95e]. Package [BKK20, BS93, HFB21, KCP^{+}94b, KOW97, LW95, OD01, SYF96, TSN21, van97, BHW^{+}12, BBH^{+}15, CwCW^{+}11, DFSW19, Gao03, KCP^{+}94a, LFS93a, LFS93b, SL95]. Packet [MBES94]. Packets [Uhl94, Uhl95b]. PaCT [Mal95]. PaCT-95 [Mal95]. PACX [FGRD01, KR09, RBB97b]. PACX-MPI [KR09, RBB97b]. Page [CML04, NPP^{+}00c]. pages [Ano95b, Ano95c, Ano96a, Ano99a, Ano99c, Ano99b, Ano99d, Ano00a, Ano00b].

Pagoda [YSS^{+}17, YSS^{+}19]. pairwise [AMHC11]. Palazzo [GT94]. PALLAS [KVH97]. Palm [TSN21]. Papers [BDB^{+}13, OL05, TB14, ACM90, CHD09, DKD07, GT19, IEE93a, IEE95c, KKDV03, MTW07, Old02, TH20, Ano93g, Cha05]. PARA [DW94, DMW96, Was96, CD96]. parabolized [SCC95]. ParaCells [SYL19]. ParADE [KKH03]. Paradigm [HIP02]. Paradigms [BGD12, CM98, DSU20, HD02a, HD02b, CdOO^{+}20]. Paradyn [MHC94a, MHC94b]. Paragon [Ano96c, HWW97, MP95, PR94a]. Parallel [ACM95b, Ada97, ATC94, Agr95a, AMHC11, AGH^{+}95, AS92, ADRCT98, AK99, AMBG93, ASA97, AL96, AP96, Ano95b, ACMR14, AB93a, AJF16, BHM94, BJ93, BBG^{+}95, BCGL97, BKK20, BFLL99, BP99, BG95, BS93, BDG^{+}91a, BKGS02, Ben01, BP98, Bha93, Bic95, BGK08, Bis04, BALU95, BCL00, BSG00, BBG^{+}99, BBC^{+}00, BBG^{+}01, BFZ97, BDL98, BDH^{+}95, BDH^{+}97, BT01b, BMS94b, BMPZ94a, BFM97, BKO00, BBH12, BGL00, CGC^{+}02, CHD07, Cer99, CDZ^{+}98, CCU95, CDK^{+}01, Cha02, CGB^{+}10, COE20, CNC10, CFF^{+}94, CSW97, CMH99, CFPS95, CCSM97, Coo95b, CT94a, CT94b, CC00b, Cze16, DSM94, DK20, DERC01, DYN^{+}06, DK13, DDP^{+}19, Di 14, DI02, DAD19, DSS00, D^{+}91, DKM^{+}92, DGMJ93, DT94, DGH^{+}19, DZDR95, DK06, DSCL05, EKTB99, EGR15, EM00a, EM00b, EGDK92, EJL92, ES11, FGRD01]. Parallel [FHSO99, FJBB^{+}00, FFP03, Fer98b, FHK01, dFdOSR^{+}19, Fis01, For95, FP92, FB94, FS93, FF95, GCBM97, GLN^{+}08, GBD^{+}94, GKP97, GR07, GSI97, GSMK17, GDM18, GB98, GHL97, GK10, GFPG12, GJN97, Gre94, GLS94, GL97a, GLS99, GkLyCY97, HFB21, HJ98, HLP10, HO14, HK94, HK93, HK95, HHK94, HT01, HH22, HAA^{+}11, IEE93b, IEE94a, IEE94f, IEE95h, IEE95f, IEE95g, IEE95j, IEE96b, IEE96c, IEE96g, IEE96e, IEE96d, IEE97b, IEE05, ITKT00, IBC^{+}10, IOK00, IDD94, IH04, IHM05, JAT97, JML01, JLG05, Jou94, JRM^{+}94, KFA96, Kan12, KDHZ18, KK02a, KOI01, KNT02, Kat93, KBS04, Kep05, KmWH10, KR09, KSB^{+}20, Kon00, KKP01, KMC96, KMC97, KS96, KKDV03, KKD04, KS01, KVH97, KHS01, Kuh98, KBG16, Kum94, Lad04, LTDD14, LTR00, LKD08, LSZL02]. Parallel [LTRA02, LHHM96, Li96, LZ97, LHZ97, kLCC^{+}06, LPJ98, LO96, Lus00, MSOGR01, MMD98, MS02b, MM92, MC18, MWG97, dlFMBdlFM02, Mar06, Mar07, MFTB95, MSCW95, Mat16, Mat94, Mat95, MSM05, MBS15, MGC12, MG15, MRB17, MYK19, MM11, Mic93, Mic95, MTWD06, MCLD01, MS95, MCdS^{+}08, MBB^{+}12, MSB97, NO02b, NO02a, Nak03, Nak05a, Nak05b, NSZS13, Nar95, NSS12, NAJ99, NJ01, NFK98, Nov95, NMC95, Oed93, OWO98, OP10, OLG01, Ong02, Ott93, OWSA95, Pac97, PPT96a, PVKE01, Pat93, PSZÉ00, PV97, Per99, Per21, Per96, PRQ21, PLR02, PWPD19, PKB^{+}16, PBC^{+}01, Qui03, RR00, RDMB99, RJ21, RBS94, Ree96, RS95, RC97, RSV^{+}05, Röh00, Rol94, RWD09, RTL99, RLL01, SCP97, SPE95, SGZ00, Sch01, Sch96a, Sch96b, Seg10, Ser97, Sev98, She95, SSLMW10]. Parallel [SM03, SP99, Sie94, Sie92a, Sie92b, SR98, Sin93, STV97, SWH15, Sou01, SBB20, SSB21, Sta95b, Ste94, SSN94, SGS10, Str96, Str97, Str94, SNMP10, Sun90a, Sun90b, Sun94a, Syd94, TMP16, TSS00b, TTP97, TC94, TCP15, TQDL01, THN00, TDBEE11, Tsu07, TVV96, Uhl94, Uhl95b, UH96, UCW95, VLO^{+}08, VRS00, VB99, WH96, Wal01a, Wel94, WAS95b, WHDB05, WO97, WSN99, WMC^{+}18, WTR03, WT12, YM97, YHGL01, YH96, YPA94, YG96, YTH^{+}12, YZPC95, YSL^{+}12, ZTD19, ZJHS20, ZB94, ZZ04, ZDR04, ZWLZ21, ZWJK05, ZAT^{+}07, ZLS^{+}15, ZZZ^{+}15, ZWC21, ZGC94, ZB97, van97, ACM97a, ARvW03, APBcF16, ART17, AAAA16, AD98, AL92, ABF^{+}17, ASCS95, ADT14, AD95, ACJ12, Ano93h, Ano95c, Ano00b, ADB94, AV18, ADDR95, AB93b, AFST95, AB13, AGIS94, ADMV05, ASB18]. parallel [BHJ96, BBB^{+}94, BR91, BA06, BHS18, BB95a, BCAD06, BB93, BDG^{+}92b, BB94, BPC94, Ben95, BvdSvD95, BKH^{+}13, BAV08, BN00, Bir94, BCM^{+}16, BKML95, Bos96, BFMR96, BID95, Bri95, Bru95, BDW97, BSH15, BB95b, CARB10, CL93, CGK11, Cav93, CLdJ^{+}15, CLSP07, CT13, CLYC16, CKmWH16, Cha05, CJvdP08, Cha96, CGL^{+}93, CEGS07, CH94, CZ96, Che99, CIJ^{+}10, CS96, CSW99, CCS19, Cla98, CEF^{+}95, CDD^{+}96, CdGM96, CBHH94, Coo95a, CCHW03, CLLASPDP99, CFF^{+}96, CPR^{+}95, CD01, CDH^{+}94, CKP^{+}93, CB11, DMK19, DKF93, DKF94b, DR18, DLR94, DLRR99, DDS^{+}94, DR94, DSZ94, DM93, DRUE12, DBVF01, DKD05, DvdLVS94, DXB96, DMW96, DLM99, DKP00, DLO03, Duv92, DZZY94, EASS95, EVMP20, EV01, FB96, FFB99, FM90, FO94, FSTG99, Fer98a, FMS15, FCS^{+}12, FKK^{+}96b]. parallel [FFM11, FHC^{+}95, GG99, GCN^{+}10, GGL^{+}08, GBF95, GKD^{+}18, GG09, GFB^{+}14, GÁVRRL17, GDMME22, GSM^{+}00, GKS^{+}11, GEW98, GKK09, GKCF13, Gra09, GP95, HHS18, HAM95b, HPY^{+}93, HD00a, HWS09, HDZ^{+}20, Heb93, HPS^{+}96, HZ94, HZ99, HPLT99, HDB^{+}13, HVSH95, Hol95, HH95, HLOC96, HVSC11, HHSM19, HLO^{+}16, IEE97a, IM95, JWB96, JC17, JY95, JJM^{+}11, JC96, JMdVG^{+}17, KCD^{+}97, KHBS19, KOB01, KBP16, KN17, KOS^{+}95a, KTXP21, KB21, KC19, KL95, Kos95b, KSS^{+}18, KRC17, KG93, KFSS94, Kra02, KKJ^{+}08, KH10, LM99, LCL^{+}12, LH98, LS10, LZC^{+}20, LCVD94a, LGMdRA^{+}19, LMM^{+}15, Lou95, LG93, LM13, LL95, LC97b, LSR95, MMR99, MYB16, MMB^{+}94, MZK93, MV20, MM95, Mar05, MSP93, MW21, MK00, MN91, MHC94a, MRRP11, MALM95, MLA^{+}14, MRH^{+}96, MMH99, Mor95, MC99]. parallel [MR96, MvWL^{+}10, NSBR07, Neu94, NB96, NBGS08, NCKB12, NF94, OdSSP12, Ols95, Olu14, OW92, PHA10, PPT96b, PPT96c, PKB06, PBG^{+}95, PNV01, PBK99, PPF89, PY95, PBPT95, PSLT99, PCS94, Ram07, RJC95, RGP22, RBB15, Rol08b, RBB17, SJLM14, SWCB20, SM12, SSKF95, SH94, Sch94, Sch99, SPK96, SBF94, SWYC94, SK92, SCC96, SL00, SMAC08, SZ11, SPL99, SMS00, SVC^{+}11, Smi93b, STT96, SH14, SRK^{+}12, SLS96, Sta95a, Sti94, SMSW06, Sun95, Sur95a, Sut96, Swa01, SL95, TJD09, THDS19, TDB00, TGKL19, TMPJ01, Uhl95a, Uhl95c, VM95, Vis95, Vos03, Wan97, WZW21, Was96, Was95a, WK08a, WK08b, WK08c, Wol92, WT11, WYLC12, WLYC12, WMP14, YULMTS^{+}17, YHL11, YWC11, YBZL03, YYW^{+}12, ZL96, ZWHS95, ZAFAM16, ZWL13, ZJDW18, ZT20, ZWL^{+}17, dH94, ARL^{+}94, Ano94e, Ano94f, ACDR94]. Parallel [BDLS96, BS94, BG94b, Bos96, CC95, Cza13, DSM94, DHK97, DW94, Edd18, EJL92, FR95, FF95, GN95, JPTE94, JPP95, KKD05, Kum94, LK10, LkLC^{+}03, Mal95, MKP^{+}96, OKW95, PQ07, QRG95, SSSS96, SPE95, Stp02, TDBEE11, TGEM09, Vol93, Vre04, WN10, YC98, ZPLS96, ZDR01, ZHS99]. Parallel-in-time [HFB21]. parallel-programming [KKJ^{+}08]. parallel/distributed [FHC^{+}95, Wan97]. parallele [GEW98]. paralleles [BL94]. Parallelisation [SJK^{+}17a, SJK^{+}17b, WCVR96, LF93b]. Parallelism [CGC^{+}11, EdS08, EK97, FKKC96, GLP^{+}00, GAM^{+}02, GPC^{+}17, DK02, KT02, Mar03, MGA^{+}17, MMS07, MdSC09, RBAA05, SHM^{+}10, SML17, SML19, SGZ00, SGL^{+}20, TCM18, TTSY00, TPK^{+}19, Thr99, YPAE09, ATL^{+}12, AML^{+}99, BK11, BR12, BS01, BS05, CCM12, GAM^{+}00, HSP^{+}13, HSE^{+}17, HK09, HY20, JC17, JPOJ12, Kos95b, MMAH20, OPP00, RKBA^{+}13, SLGZ99, SHPT00, THH^{+}05, TWFO09, WO09, WTFO14, WRSY16, WZW21, YZ14, PGdCJ^{+}18]. Parallelization [AL93, And98, AAB^{+}16, AIM97, BCM11, BS07, CRE99, CP97, Cou93, CF19, Cza03, ETV94, HA10, JR10, Kik93, KLR^{+}15, LP00, MB18, OD01, Pok96, QMGR00, Rag96, RP95, RM99, RS97, SAS01, WPL95, WZWS08, WR01, aMST07, ACC^{+}21, ABB20, AGMJ06, BW12, BDY99, BJS99, CDD^{+}96, FSG19a, Gao03, Goe02, IDS16, IJM^{+}05, JL18, JJY^{+}03, JMS14, KS15a, KD12, KRG13, MCB05, MGG05, MMDA19, Nes10, NEM17, OLG^{+}16, Stp18, TWFO09, VBLvdG08, ZT20]. parallelize [JKN22]. Parallelized [FBSN01, OMK09, AiIS^{+}21, DFSW19, KMG99, OKM12]. parallelizer [BHRS08]. Parallelizing [BST^{+}13, Car07, GGH99, IOK00, IKM^{+}01, IKM^{+}02, SR95, ZZ95, AMS94, BY12]. Parallelldatorcentrum [Eng00]. Parallizing [LRQ01]. parameter [DH22, HPLT99, JMdVG^{+}17]. parameterizable [JCP^{+}20]. parameterized [CT13]. Parameters [GFV99, BAG17, KSC^{+}19]. Parametric [LLG12, Pat93]. parametrised [TGS^{+}20]. Paramid [Ste94]. Paraperm [LTDD14]. Paraprox [SJLM14]. Parasite [LLRS02]. paravirtualization [SBQZ14]. ParCo93 [JPTE94]. PARCOACH [SCB14]. PARCS [LD01]. Paris [CHD07, Har94, Har95]. Parity [MC17]. Parix [HVSH95, RS95, SHH94a, SHH94b]. Park [SL94a, IEE93c]. PARKBENCH [DHS96, DH95]. PARMACS [GR95, HZ96, HZ99]. PARMACS-to-MPI [HZ96]. ParNSS [HSMW94]. PARRAY [CCM12]. parsing [Sur95a]. Parsytec [SHH94a, SHH94b]. part [VSRC95, EM00a, EM00b, GK10]. Partial [DERC01, DLV16, FSSD17, KK02b, MK17, MFTB95, MH18, MKK21, OM96, ST17]. partially [CdGM96]. Particle [GSI97, KHS01, NSLV16, RBP^{+}21, ZZ04, BAS13, CFF19, FFFC99, GSMK17, KPK13, RFH^{+}95, VDL^{+}15]. particle-based [FFFC99]. particle-in-cell [VDL^{+}15]. particle-mesh [BAS13]. particulate [ATL^{+}12]. Partition [DAD19, PS19a]. partitioned [DWS^{+}21]. Partitionierung [Gra97]. Partitioning [CTK01, DAD19, kL11, SPB^{+}17, STV97, WJG^{+}21, CT13, Cha96, Gra97, GKCF13, YST08]. Partitioning-Based [WJG^{+}21]. partners [Str94]. Pasadena [IEE95c]. Pascal [GDS^{+}20, KC19]. PASCO [ACM97a]. passage [PTMF18]. Passing [AMHC11, Ano93d, AKL99, Att96, BC19a, BZ97, BC14, BBH^{+}06, BBG^{+}99, BBG^{+}01, BRU05, BDH^{+}95, BDH^{+}97, BGR97b, BFM97, CHD07, Cer99, CGH94, Cot97, Cot98, CTK00, Cot04, CDND11, DFKS01, DKD08, DHHW92, DHHW93a, DDL00, FKKC96, FKS96, FGT96, Fos98, FGG^{+}98, FB94, GR07, GB96, Gle93, GLRS01, GLS94, GL95c, GLDS96, GLT99, GLS99, GLT00b, GLT00a, GL04, IBC^{+}10, KTF03, KGRD10, KS97, KSV01, KKDV03, KKD04, KKD05, LKD08, LK10, Luo99, MPI98a, MPI98b, MTSS94, MS98, MSL96, MBES94, MG97, MTWD06, MSS97, NW98, PBK00, Pok96, PS01b, RRBL01, RWD09, RFG^{+}00, SWHP05, SWL^{+}01, ST02b, TGT05, TDB00, TBD12, WD96, Wer95, Wis97, YHGL01, ZG95a, ZG96, ZLL^{+}12, Ada98, AD98, AAC^{+}05, Ano93e, Ano94d, Ano95c, Ano00a, Ano00b, BL97, BvdSvD95]. passing [Bjo95, Bru95, BDW97, BFIM99, CGJ^{+}00, CDZ^{+}98, CRD99, CD01, DKF93, DM93, DKD05, DS96b, DHHW93b, DOSW96, DLM99, DKP00, DLO03, FK94, FHB^{+}13, GL92, HP05, HPY^{+}93, Hem96, JKN22, KJA^{+}93, Kra02, LR06a, LBD^{+}96, wL94, LCY96, LMM^{+}15, LC97b, MP95, NS91, PS07, PKB06, Pie94, PR94a, PS00b, Sei99, SWJ95, SDV^{+}95, SZ99, SSG95, Sti94, TSZC94, VM95, Wal94a, Wal94b, ZWL13, ZKRA14, DiN96, GGHL^{+}96, Han98, Hem94, RRFH96, SLG95, Wer95, YGH^{+}14]. Past [Dar01]. Path [CGPR98, GSYT21, GAMR00, SDJ17, SLN^{+}12, Zel95]. path-based [SLN^{+}12]. pathological [LCH^{+}22]. Pathway [CNM11]. PATOP [BFBW01]. Pattern [CSW12, CC17, JJPL17, RDMB99, MAS06, SJLM14]. pattern-based [SJLM14]. Pattern-Independent [CSW12]. Patterned [ST17]. Patterns [DMMV97, FPY08, KB98, MSM05, PKB^{+}16, Pro21, RRAGM97, SGH12, SR98, DZZY94, GÁVRRL17, HGMW12, LGMdRA^{+}19, PM95, PSK^{+}10]. PC [AH00, CDT05, EKTB99, KS01, LKYS04, RLL01, Ste00, WLYC12, YST08, YL09, ZJHS20, MMB^{+}94]. PC-Cluster [RLL01]. PCAT [ACDR94, GN95]. PCAT-93 [ACDR94]. PCAT-94 [GN95]. PCG [BJS97]. PCI [GK97]. PCI-based [GK97]. PCRCW [BS94]. PCs [CRE99]. PCSC [LM94]. PCTE [HZ94]. PCTRAN [KHS01]. PDCS [YH96]. PDE [GBR15, KTXP21, NHT02, NHT06, NPS12]. PDES [PT01, SCL00, SCL01, CdOO^{+}20, HO14, HHA95]. PDGC [CGB^{+}10]. PDP [IEE96g]. pearl [HLK^{+}20]. Peer [GR07]. Peer-to-Peer [GR07]. PELCR [PQ07]. PEMPI [FB95]. PEMPIs [MOL05]. Pennsylvania [ACM96b, IEE94d]. pentadiagonal [GNP19, Kan12]. Pentium [Ano03]. Pentium(R) [SBT04]. PENTRAN [KHS01]. people [ASCS95, Ano94i]. per-triangle [SOA11]. perception [CLM^{+}95]. perceptual [WPL95]. perform [CBIGL19]. Performance [ACM97b, ACM98a, ACM98b, ACM00, ACM01, ACM04, AC07, ATM01, AR01, Ano01a, Ano01b, ADR^{+}05, AJC^{+}20, Bak98, BBGL96, Ben18, BN00, BS21, BBDH14, BGG^{+}02, BY12, BRM03, BRST94, BS07, BDL98, BCKP00, BHNW01, BFMT96b, BFBW01, BEG^{+}10, CGK^{+}16, CVPS19, CDD^{+}13, CRE99, CDJ95, CGLD01, CBB^{+}21, CNM11, Che99, COE20, CSC96, CCBPGA15, DPSD08, DM95b, DW02, DZ98b, DPP01, DWL^{+}10, DBK^{+}09, EGH99, EGC02, EML98, EML00, FD02a, FGRT00, FCP^{+}01, FSC^{+}11, FST98b, FGKT97, GFD03, GKP96, GGS99, GBH99, GFIS^{+}18, GRRM99, GBS^{+}07, GC05, GMdMBD^{+}07, GSY^{+}13, HVA^{+}16, HKN^{+}01, Hol12, HF14a, HF14b, HPS95, Hus98, IEE92, IEE93c, IEE94g, IEE95k, IEE96a, IEE96f, IEE97c, IFI95, IRU01, IHvA^{+}00, IADB19, JSS^{+}15, JC17, JCH^{+}08, JS13, JLG05, KDSO12, KaM10, KL94, KH12, KBS04, KBM97]. Performance [KC19, KKP01, KH15, KC06, KK02b, KHS01, KSS00, Laf01, LAdS^{+}15, LWSB19, LCK11, LC97a, LB98, LGCH99, LNK^{+}15, LH98, LC93, LkLC^{+}03, LWZ18, LNW^{+}12, LRLG19, LS10, LCW^{+}03, LVP04, LWP04, LDCZ97, LZHY19, LC97b, LKYS04, MMB^{+}94, MKP^{+}96, MPD04, ME17, MGMH97, MGC12, MM02, MM03, MOL05, MS99a, MHC94b, MMSW02, MK04, MCLD01, MMH99, MM14, MMS07, MZLS20, NSLV16, NMW93, NFK98, NPP^{+}00d, NMS^{+}14, NN95, OTK15, OPJ^{+}19, OF00, OLG01, PARB14, PKB01, PHJM11, PZ12, PR94b, PFG97, PGAB^{+}05, PGAB^{+}07, PGC02, PY95, PTH^{+}01b, PS01b, QHCC17, QB12, Rab98, RBB97a, RBB97c, RH01, RRAGM97, Ros13, RšT06, SGJ^{+}03, SPM^{+}10, SLJ^{+}14, SWHP05, SCP97, SEF^{+}16, SPL^{+}12, SCSL12, SM02, SM03, SSC97, SJ02, SSSS97, SC96b, SKH96, SJK^{+}17a, SJK^{+}17b, TSB02, TSB03, TTSY00]. Performance [Ten95, Tha98, TBG^{+}02, TGT10, Trä12b, TFGM02, TFZZ12, VFD02, VY02, WZM17, WQKH20, WN10, WAS95b, WM01, WT11, WT12, WT13, WYZ^{+}19, XF95, XH96, XXL13, YC98, Yan94, YWC11, YS93, YWCF15, YSP^{+}05, ZLGS99, ZWLZ21, ZWJK05, ZHK06, Zho21, ZSnH01, ABDP15, Ahm97, ADLL03a, ADLL03b, Ano03, AFST95, BDP^{+}10, BAE22, Ber96, BPJ22, BDV03, BFM96, BFMT96a, BFIM99, CRE01, CAHT17, CLYC16, CBPP02, CBM^{+}08, CHKK15, DM95a, DL10, DO96, D^{+}95, DWL^{+}12, DE91, Duv92, EFR^{+}05, ESB13, FAF16, FD02b, FE17a, FE17b, FSV14, FME^{+}12, Fin97, GVF^{+}18, GS02, GGC^{+}07, GK97, GR95, GHZ12, GML^{+}16, GSM^{+}00, GL96, GLDS96, GL97c, GL99, GWVP^{+}14, HDDG09, HLK^{+}20, HW11, HGX^{+}22, HASnP00, HAJK01, HMS^{+}19, HK10, HVSC11, HHA95, HG12, HcF05, JKHK08, JJM^{+}11, JKN^{+}13, KBP16, KKM15, KS13, KSC^{+}19, LBD^{+}96]. performance [LTLC94, LFS^{+}19, LC07, LML^{+}19, LBH12, LCY96, LB96, LL01, LKJ03, LSK04, MC17, MP95, MSMC15, MSW^{+}05, MSL12, MKP22, MABG96, MHC94a, MSZG17, MJPB16, MGC^{+}15, NU05, NFG^{+}10, OIH10, Old02, PGS^{+}13, PS19a, PHW^{+}13, PGK^{+}10, PF05, PMZM16, PTW99, Rab99, RMS^{+}18, RPS19, Reu03, RGDM15, RJDH14, Sep93, SFO95, SPBR20, SWJ95, Slo05, SVC^{+}11, SK00, SFLD15, TMC09, TSP95, TG09, THM^{+}94, VDL^{+}15, Wor96, XR21, YCL14, ZSK15, ZWL13, ZGZS20, dAT17, HS95a, GH94, LCHS96, SSH08]. performance-aware [MSMC15]. Performance-based [YWC11]. Performance-Driven [LWSB19]. Performance-Neutral [CBB^{+}21]. Performance-Portable [JSS^{+}15, DWL^{+}10, DWL^{+}12, FAF16]. performance-prediction [BDV03]. performance/cost [GWVP^{+}14]. performance/power [RPS19]. Performances [GFV99, DS96b, IM94]. Performing [CC99]. Peridynamic [MSZG17]. Periscope [LGG16]. perishable [OHG19]. Permutations [CC99, LTDD14]. Persistent [Man01, SG12, HMS^{+}19]. Persistent-Sets [SG12]. Personal [SSSS97]. personalized [BHJ96]. perspective [Sni18]. perturbation [KN17]. Perverse [Rol08a]. PES [MK94]. Pessimistic [BCH^{+}03]. petaflops [LSG12]. Petascale [CGKM11, CBYG18, ZWL13, Gei01]. Petersburg [Mal95]. Petri [CNM11]. PFACC [HY20]. PFSLib [LL95]. PGAS [SWS^{+}12, SJK^{+}17a, SJK^{+}17b]. Phase [CBL10, DH22, ED94, TKP15, TG94, ZAFAM16]. phase-field [TKP15]. PHAT [BBC^{+}19]. Phi [BB18, CBIGL19, DSGS17, MTK16, OTK15]. Philadelphia [ACM96b]. Phi^{T M} [MMDA19]. PHOENICS [SZBS95b, SZBS95a]. Phoenix [ACM03, IEE95b, Ten95]. Photo [JFGRF12]. Photonic [lLLmH^{+}21]. Phylogenetic [MR12, LBH12]. Physical [BM97, GJN97, SR98, GWVP^{+}14]. Physics [GT94, KH15, VW92, WBH97, ANS95, BPG94, DMW96, SPBR20]. PIC [BDV03, HTJ^{+}16, JL18]. Picos [YÁJG^{+}15]. Pilot [OS97, CGG10]. PINEAPL [DHK97]. Pinhole [NH95]. Pipe [MTU^{+}15]. Pipeline [GAMR00, KB21]. Pipelined [GAML01]. Pipelines [MAGR01, FWS^{+}17, RKBA^{+}13]. pipelining [MM11]. Pisa [Sil96]. Pitaevskii [LBB^{+}16, LYSS^{+}16, SSB^{+}16, YSVM^{+}16, YSMA^{+}17]. Pittsburgh [ACM96c, ACM04, Ham95a, IEE94d]. Place [IEE94e, LTS16, BCK^{+}09, HSE^{+}17, PSHL11]. placement [DJJ^{+}19, SLN^{+}12, SPK^{+}12]. Planck [Ano94c]. Planing [GAMR00]. Planning [HMS^{+}19, Zel95]. plant [FO94]. PLAPACK [van97]. plasma [JL18, DGH^{+}19, YKLD17]. Plasmafusionsforschung [BL94]. plasmas [CFF19]. Platform [BKGS02, BB18, NO02b, PGF18, WTTH17, BSH15, CB11, Cza13, DWL^{+}10, DWL^{+}12, HTJ^{+}16, HHA95, JPL22, JR13, KSC^{+}19, NO02a, XXL13, YSL^{+}12]. Platforms [AIM97, COE20, HSO^{+}21, HD00b, JML01, OPJ^{+}19, RVKP19, ZB97, BBC^{+}19, EYP^{+}20, GGC^{+}07, GFB^{+}14, MBBD13, TKP15, TS12b]. Plesset [BL95, KN17]. PLIERS [MMR99]. plug [MS99b]. plug-in [MS99b]. plume [JL18]. plus [HDB^{+}13, Stp18]. PMaC [PTL^{+}16]. PMD [Che99]. PML [Ram07]. PMPIO [FWNK96]. PMPIO-a [FWNK96]. pocl [JSS^{+}15]. Point [GBS^{+}07, HC10, KV98, LWSB19, TSN21, ADLL03a, ADLL03b, WK20]. Point-to-Point [GBS^{+}07, HC10, KV98, ADLL03a, ADLL03b, WK20]. Pointer [WYH^{+}21]. Pointer-Based [WYH^{+}21]. Pointers [LRT07]. Poisson [BP98, WJB14]. Poland [BDW97]. Polder [OS97]. Policies [CML04, PZ12, OHG19]. policy [MMM13]. Polling [DCPJ12, Pla02, DCPJ14, SH96]. Pollutant [RSV^{+}05]. Pollution [AKK^{+}94, BZ97, MPD04, MSML10, SH94, Syd94]. POLSYS GLP [SMSW06]. polygonization [TSP95]. polygons [CT13]. polyhedral [BHRS08, KGB^{+}09]. polymers [JAT97]. Polynomial [VY15, HLM^{+}17, SMSW06]. port [CCHW03, Har94, RJMC93]. Portability [KaM10, RS95, RH01, ABDP15, CGK^{+}16, FE17a, FE17b, HHS18, MGC^{+}15, PHW^{+}13, QHCC17, Reu03]. Portable [Ano95c, Ano00b, BHV12, BHLS^{+}95, CDH^{+}94, DHK97, Di 14, FCLG07, FSLS98, GLS94, GL97a, GLS99, JSS^{+}15, LNLE00, Man98, MKV^{+}01, MG97, PPT96a, PBC^{+}01, SSCC95, STH22, SDB^{+}16, Sti94, Tra98, WCS^{+}13, YBMCB14, YT20, Arn95, BCK^{+}09, BfDA94, BB00, BL99, BAS13, CJvdP08, CH94, CEF^{+}95, DWL^{+}10, DWL^{+}12, FAF16, FWNK96, GR95, GL94, GS94, GLDS96, HTJ^{+}16, HZ94, HSW^{+}12, JC96, KN95, LFS93a, LFS93b, LHC^{+}07, MMB^{+}94, PPT96b, PPT96c, PMZM16, SSH^{+}19, SFLD15, Sto98, VM95]. portal [AASB08]. portals [BS96b, BMR02, BRM03]. Portfolio [SIS17]. Portfolio-driven [SIS17]. Porting [Ano96c, BBB^{+}22, BSC99, BLW98, EM02, HSO^{+}21, Har94, Har95, HASnP00, KGK^{+}03, KME09, SR96, YKLD17, dCH93, BvdB94, HD11, LBG^{+}20, MWO95, ZPLS96]. Portland [ACM99, ANS95, IEE93e, SW91]. Portugal [IEE93d, IEE96g]. Positron [Pat93]. POSIX [LD01]. Post [BBH^{+}13b, Wit16, ABC^{+}00]. Post-failure [BBH^{+}13b]. Post-ISA [Wit16]. Poster [JJPL17, LZH17]. POSYBL [Mat94]. Potential [EGC02, Gro01a, KS15a]. potentials [THDS19]. Potts [KO14]. POV [FFB99]. POV-Ray [FFB99]. Power [DDN^{+}22, LWZ18, LB96, EZBA16, FO94, HK10, Nel93, RPS19, SM19, Bri95, DDN^{+}22]. Power-Efficient [DDN^{+}22]. Powered [NE98, RTN21]. PP [IEE96d]. PPARDB [PPT96b, PPT96a, PPT96c]. PPARDB/PVM [PPT96b, PPT96c]. PPPE [CDH^{+}94]. PPSN [DSM94]. PPT [BAE22]. PPT-Multicore [BAE22]. Practical [ACC^{+}21, BHJ96, BCP^{+}97, CZG^{+}08, RHG^{+}96, TGBS05, AMS94, BHRS08, LPD^{+}11, McK94, Pan95b, VVD^{+}09, WDR^{+}19]. Practice [ACM11, GN95, ZGZS20]. Praktische [MS04]. Pre [AC17]. Pre-processor [AC17]. Precedence [EGR15]. Precedence-Constrained [EGR15]. Precise [FJK^{+}17]. Precision [Ano98, Kha13, ZC10, JPT14]. Precisions [HDW21]. Preconditioned [GFPG12, ABF^{+}17, MM92]. Preconditioner [BBS99, FSXZ14]. Preconditioners [Huc96]. Preconditioning [MYL21, Nak03, GGC^{+}07]. predictability [GRRM99]. Predicting [RRAGM97]. Prediction [MOL05, WHDB05, ZWJK05, ADR^{+}05, BAE22, BDV03, CMV^{+}94, HHA95, RBAI17, SEC15, SC96b, SSN94, Was95a, ZAT^{+}07]. Predictive [FJK^{+}17]. Preemptive [BBH^{+}06, BBGL96]. Preface [DKD07, OL05]. Prefetching [BIC^{+}10, KC19]. Prefix [WJ12, DK13, MYB16]. Preliminary [BF98, Wal01a, WLK^{+}18, RJC95, RLFdS13, SWS^{+}12]. PREMER [VBB18]. Preprocessors [Ano01a]. prescription [MRH^{+}96]. Present [Dar01]. presented [ACM90]. preservation [IEE94c]. Preserving [RNPM13]. Press [Ano95b, Ano95c, Ano96a, Ano99a, Ano99c, Ano99b, Ano99d, Ano00a, Ano00b, Edd18]. Pricing [RR00]. Primitives [DDL00, FST98a, ZLWW20, ABDP15, CIJ^{+}10, STP^{+}19]. Princeton [Bha93]. principles [BSC99, HS12, SSP^{+}94]. printing [YM97]. priority [DR95, Man98]. Prism [SDN99]. private [Str94]. privatization [KRG13]. Probabilistic [LAdS^{+}15]. Probability [QRMG96, Sta95b]. Problem [BSH15, DALD18, DAK98, GAMR00, ICC02, Lee06, MTSS94, RLVRGP12, ZSnH01, AB93b, DSM94, GM94, GKCF13, GADM20, HMKV94, IHM05, MM92, RRJ^{+}20, SL00, SP11, TSCS14, Cza13]. Problems [ASA97, BHM94, BHM96, BMR01, BPMN97, CGPR98, EML98, HAA^{+}11, DK02, LSM^{+}18, MBS15, Nak03, Riz17, AL96, CEGS07, FR95, JRG21, LSR95, NZZ94, OMK09, SC96a, SD99, TGS^{+}20]. procedure [AGLv96]. Proceedings [ACM94, ACM96c, ACM97a, ACM97b, ACM98b, ACM04, ACDR94, CJNW95, GN95, Hol12, IEE93f, IEE95d, IEE02, KG93, LCK11, MC94, RV00, R^{+}92, SM07, Ten95, TG94, dGJM94, ACM96b, Ano94e, Ano94i, BPG94, Boi97, BH95, CLM^{+}95, DSZ94, DE91, EJL92, FF95, GHH^{+}93, HK95, HHK94, IEE94a, IEE94b, IEE94c, IEE95b, IEE95e, IEE96a, IEE97c, IEE05, JPTE94, Kum94, LF^{+}93a, Li96, PSB^{+}94, PBPT95, SPE95, SW91, WPH94, ACM90, ACM95a, ACM05, ACM06b, ACM06a, ATC94, Agr95a, AGH^{+}95, AH95, Ano89, Ano92, Ano94a, BBG^{+}95, Bha93, CHD07, CZG^{+}08, CGKM11, CMMR12, CGB^{+}10, CDND11, DKM^{+}92, DT94, DLO03, EV01, EdS08, ERS95, ERS96, Fer92, FK95, Gat95, GGK^{+}93, GA96, GT94, Ham95a, HS94, HK93, IEE91, IEE92, IEE93d, IEE93c, IEE93b, IEE93e, IEE94e, IEE94d, IEE94f, IEE94h, IEE94g, IEE95h]. Proceedings [IEE95k, IEE95i, IEE95f, IEE95l, IEE95g, IEE95j, IEE96g, IEE96f, IEE96e, IEE96d, IEE96h, KGRD10, LKD08, MTWD06, MMH93, MCdS^{+}08, MdSC09, Ost94, PR94b, Ree96, RWD09, SCR92, SHM^{+}10, Sie94, TBD12, USE94, USE95, USE00, VW92, Vos03, Y^{+}93, YH96, AD98, BG91, BDLS96, BS94, Bos96, BFMR96, BDW97, CH96, CD01, DSM94, DKD05, DW94, DMW96, DLM99, DKP00, Eng00, FR95, GH94, HAM95b, HS95a, IEE96c, IEE97a, Kra02, KKD04, LCHS96, Mal95, PBG^{+}95, Sch93, Tou96, VV95, Vol93, Was96]. Proceedings. [Ano93f, Ano94g, IEE96i, IEE97b, LHHM96]. Process [AUR01, BGL00, CLL03, DeP03, DK06, FDG97a, FDG97b, FLD98, FPY08, KCP^{+}94b, KOW97, PS00a, Pro21, SC04, ST97, TSN21, Tra02a, BK11, BBGL96, CK99, FLD96, GL95a, HRR^{+}11, HG12, JLS^{+}14, KCP^{+}94a, MLVS16, MK00, SHHC18, Ste96]. Process-Management [BGL00]. processed [HJ98]. Processes [CB16, MW98, Pet00a, Pet00b, FS95, GFIS^{+}18, SOYHDD19, SPK^{+}12]. Processing [ATC94, Agr95a, AR01, BBG^{+}95, DKM^{+}92, GGCM99, GGCGO01, HJBB14, IEE93b, IEE93f, IEE95e, IEE95h, IEE95f, IEE95g, IEE96b, IEE96g, IEE96e, IEE96d, IEE97b, IEE05, IOK00, JDB^{+}14, KOI01, KS15b, LSVMW08, MLGW18, MC18, MSML10, Nar95, NH95, NJ01, OWO98, PLR02, PD98, Ree96, RRBL01, Rol94, SCP97, Sev98, Sie94, Sin93, VLO^{+}08, WN10, AB95, Ano94f, ASB18, BJ13, BHS18, BFMR96, CFPS95, CLLASPDP99, DSZ94, FWS^{+}17, GDC15, GGGC99, Gre94, HAM95b, HCC^{+}20, HPS^{+}96, JPL22, JC96, Kat93, KB21, Kum94, LHLK10, LG93, PSB^{+}94, PBPT95, RKBA^{+}13, Röh00, RCG95, SSS99, SLS96, VDL^{+}15, Wol92, WWFT11]. Processor [HC06, Oed93, Ott94, PWP^{+}16, RR02, Smi93a, SBT04, UALK17, UALK19, ABDP15, AC17, DJJ^{+}19, DCH02, HC08, LL01, MMDA19, OIS^{+}06, RNPM13]. Processor-Oblivious [UALK17, UALK19]. Processors [AJ97, Bri10, DDP^{+}19, HK93, HK95, KmWH10, MJB15, OLG01, PZKK02, AV18, BBG^{+}14, CBM^{+}08, DBLG11, HTA08, HWX^{+}13]. Producing [HAJK01]. product [CMH99, ER12, SMSW06]. Production [IADB19, CLdJ^{+}15, SL00]. productive [LV12]. Productivity [BS07, DSU20, KaM10, Wit16]. products [Ano97, Bra97]. profile [TWFO09, WTFO14]. profile-driven [TWFO09, WTFO14]. profiler [AS92]. profiles [BAE22, Wil94]. Profiling [AJC^{+}20, EYP^{+}20, GPL^{+}96, LZHY19, Rab99, Vet02]. Profitability [CLA^{+}19]. Program [Ano96d, AB93a, BMS94b, CHPP01, Cot97, EML98, MM95, MK17, MRV00, Ney00, PS01b, TSY00, THN00, UTY02, CDZ^{+}98, CZP21, JF95, LP00, LLC13, OKM12, PPF89, Sai10, TNIB17, TMPJ01, ZL96]. programación [VP00]. Programmable [OA17]. Programmcode [BL94]. Programmer [Gua16, Wit16]. programmers [CGG10]. Programming [ACM90, Ada97, ACGR97, ASA97, ACJ12, Ano96b, BBG^{+}10, BLP93, BHV12, BF01, BBG^{+}99, BBG^{+}01, BKO00, CMK00, CDK^{+}01, CKmWH16, Cha02, CZG^{+}08, CF01, Cza03, DM98, DSU20, DARG13, DDL00, DK06, DWL^{+}10, EM00a, EM00b, FTVB00, FWR^{+}95, GLRS01, GLS94, GLS99, HSO^{+}21, HA11, HDB^{+}12, HDT^{+}15, KKH03, Kep05, KP96, KmWH10, KVH97, Lad04, Laf01, LLRS02, MSOGR01, Mat94, Mat95, MSM05, MCdS^{+}08, NO02b, SPM^{+}10, SK10, SS01, SDN99, SHH94b, ST02a, ST02b, SGS10, Stp02, TTP97, VT97, Vre04, Wal01a, Wal02, WO97, YM97, YHGL01, YCA18, ACGdT02, AMuHK15, Ano95c, Ano00b, AB13, BJ13, BCA^{+}06, BB94, BS96a, BKH^{+}13, CPM^{+}18, CLYC16, Cha05, CJvdP08, CEF^{+}95, CDH^{+}94, CGH^{+}14, DWL^{+}12, Duv92, EASS95, EVMP20, EV01, FSG19b, FB95, FB96, Fan98, FSTG99, Fer04]. programming [Fra95, FHB^{+}13, FF95, GKZ12, Gei96, GBH14, GBH18, GRTZ10, HTA08, HS93, HDZ^{+}20, HZ94, HDB^{+}13, HVSH95, HSW^{+}12, HZG08, HY20, JPL22, KDSO12, KOB01, KSG13, KSL^{+}12, KLV15, KPNM16, KFSS94, KKJ^{+}08, LV12, LFS93a, LFS93b, LH98, LPD^{+}11, LLH^{+}14, MMB^{+}94, MVTP96, MSP93, MC99, MGC^{+}15, NO02a, Nak05a, NYNT12, NBGS08, OIS^{+}06, Olu14, OW92, Pac97, PVKE01, PF05, Qui03, RBW^{+}20, RJDH14, STP^{+}19, iSYS12, SSKF95, SYR^{+}09, Seg10, SPK96, SBF94, SPL99, SHH94a, SD99, VP00, Vos03, Wal01b, Wan02, WCC^{+}07, WADC99, WYLC12, WLYC12, YHL11, YWC11, YX95, YS93, ZWC21, ZGC94, DR94, HSE^{+}17, Che10, SD13]. Programs [AJF16, Beg93b, BKdSH01, BGK08, BGG^{+}02, BDL98, BGL00, CSW12, CRE99, CHPP01, CD98, DLB07, DMMV97, Di 14, FKH02, FJK^{+}17, GR07, GTH96, GSYT21, GL04, GC05, HC10, HKN^{+}01, HM01, JLG05, KFL05, KL94, KSJ14, KKV01, KSV01, Mar09, MVY95, MOL05, MBE03, MKW11, MCLD01, MJB15, NSZS13, NE98, NE01, NPP^{+}00d, OM96, PPJ01, RH01, RFG^{+}00, SGZ00, SBF^{+}04, SR96, TGBS05, WYH^{+}21, Wel94, Wis97, ZLL^{+}12, Beg92, Beg93c, Beg93a, BCK^{+}09, BMPS03, CRE01, CLdJ^{+}15, CGL^{+}93, CH94, CRM14, CFP96, DKF93, DKF94b, EP96, EPP^{+}17, FSG19a, FLB^{+}05, FKLB08, GGH99, GRRM99, GKS^{+}11, GB94, HD11, HZ96, HLOC96, HEHC09, KCD^{+}97, KS13, KO14, Kom15, KLM^{+}19, LGKQ10, LLG12, LL16, LBB^{+}16, LYSS^{+}16, LMM^{+}15, LZC^{+}02, LCC^{+}03, MT96, MdSAS^{+}18, Mor95, NBK99, Obe96]. programs [OdSSP12, PES99, PAdS^{+}17, RAS16, Reu03, RRG^{+}99, SSB^{+}16, SKS01, SMAC08, SZ11, SR95, SY95, SC96b, TMW17, THH^{+}05, TGKL19, UGT09, VVD^{+}09, WZW21, YSVM^{+}16, YSMA^{+}17, YYW^{+}12, ZJDW18, ZRQA11]. Progress [BRU05, LAdS^{+}15, SPH^{+}18, DJJ^{+}19, MLA^{+}14, RSC^{+}19, MC94]. Progress-Dependence [LAdS^{+}15]. Project [BHK^{+}06, BSH15, DHK97, MRV00, ABC^{+}00, BBB^{+}20, CDH^{+}94]. Promise [Ano93f]. Promotion [OCY^{+}15, WBBD15]. Propagation [EMO^{+}93, ESM^{+}94, JML01, SMOE93, ASAK19, KEGM10, RMNM^{+}12, ZWC21]. proper [TGS^{+}20]. Properties [FGRT00, JL18, MS96b, SSP^{+}94]. Proposal [DHHW92, DHHW93a, DFC^{+}07, DFA^{+}09, ZKRA14]. Proposals [Wal96b]. protected [GHD12]. Protein [RGB^{+}18, GÁVRRL17, RJH^{+}20, SEC15, ZAT^{+}07]. protein-protein [RJH^{+}20]. proteins [BHW^{+}12, BBH^{+}15, FMS15]. Protocol [CAWL17, GSY^{+}13, kL11, LMM^{+}15, RA09, XF95, BDB^{+}13, CwCW^{+}11, DDYM99, MN91, MB00, ZPI06]. Protocol-based [LMM^{+}15]. Protocols [BCH^{+}08, DDN^{+}22, DM93, LH98, LZZ^{+}20]. Protoplanetary [dlFMBdlFM02]. Prototype [Ano01b, FHP^{+}94, MMSW02, BK96, CCF^{+}94, KYL03, KYL05]. Prototyping [SXMX^{+}18, Spe19]. prover [Sut96]. Provide [Add01, LMRG14]. Provides [Ano98, Nel93]. Providing [GKP97, Zah12]. Proving [MS96b]. PRS [UCW95]. pruned [dFdOSR^{+}19]. Pruning [SMM^{+}16, WQKH20]. PS [AMV94]. Pseudo [Wal01a, Lan09]. Pseudo-search [Wal01a]. Pseudorandom [WHDB05, Stp20]. Pseudospectra [BKGS02]. pseudospectral [Bri95, MRRP11]. PSPVM [BWT96]. Pthread [ZAT^{+}07]. Pthreads [AS14, TS12b]. PTX [iSYS12]. Public [Str94, GWVP^{+}14, Nel93, RST02]. Public-private [Str94]. Pulsar [WTS19]. pulse [ASAK19]. Puma [BS96b]. purely [HSE^{+}17]. Purpose [AJYH18, BDT08, Che10, SZBS95a, Sun94a, ABDP15, CBM^{+}08, KPNM16, PF05, SK10, SSD^{+}20, SZBS95b]. PVaniM [BCLN97, TSS98]. PVFS [IRU01]. PVM [AD98, BL94, BDLS96, BDW97, CHD07, CHD09, CD01, DKD05, DLM99, DKP00, DLO03, Kra02, KKD04, LKD08, McD96, MTWD06, RWD09, Wil94, AJ97, Ahm97, AS92, ACGR97, ADRCT98, AL92, AGR^{+}95b, AB95, ASA97, AL96, ARL^{+}94, AKK^{+}94, AP96, Ano94b, Ano95e, Ano96b, Ano96c, ABCI95a, ABCI95b, ABG^{+}96, AGLv96, AB93b, AB93a, ADMV05, BSN95, BLP93, BFLL99, BBGL96, BG95, BS93, BDG^{+}91a, BDG^{+}92b, Beg92, BDG^{+}93b, BDG^{+}93a, Beg93b, Beg93c, Beg93a, BDG^{+}95, BS96a, BDG^{+}xx, BL95, BR95b, Ber96, BJS97, BT96, BWT96, BG94a, Bon96, BG94b, BG94c, Bor99, BCD96, BRR99, BFZ97, BID95, BMS94b, BFM96, BFMT96a, BFMT96b, CMV^{+}94, CP97, CDJ95, CKO^{+}94, CCK^{+}95, CSPM^{+}96, CZ95a, CGPR98, CG93, CDHL95, CDH^{+}95, CF01, CZ96, CS96, CG96, CG99a]. PVM [CSC96, CDM93, CdGM96, CPR^{+}95, CT94a, CT94b, CFP96, CT02, CD98, CTK01, DG95, DKF94a, DDYM99, DM95b, DM95a, DP94, DMMV97, DGF97, DFN12, D^{+}91, DGMS93, DGMJ93, DHP97, DPZ97, EP96, EM94, EGDK92, ED94, EM02, EML98, EML00, ES11, EMO^{+}93, ESM^{+}94, EK97, FMBM96, FD96, FLD96, FH95, FHSO99, FO94, FSTG99, FJBB^{+}00, Fin97, FD97, FS97, For95, FS93, GRV01, Gal97, GCBM97, GS91a, GS91b, GS92, GS93, Gei93a, Gei93b, GDB^{+}93, GBD^{+}94, Gei96, GKP96, Gei97, GKPS97, Gei98, GSxx, Gei00, Gei01, GTH96, GB96, GM95, GSHL02, GFV99, GGH99, GS96, Gör01, GHL97, Gre95, Gre94, GL97b, GMU95, GkLyCY97, HB96a, HB96b, HSMW94, HJ98, Har94, Har95, HBT95, HPS^{+}96, Hem96, HEH98, HTHD99, HVSH95, HH95, HRSA97, Huc96, Hum95, HS95b]. PVM [ITT99, IvdLH^{+}00, IDD94, IKM^{+}01, IKM^{+}02, JAT97, JH97, JML01, JW96, JC96, KBA02, Kat93, KK98, KP96, KBM97, KDL^{+}95a, KDL^{+}95b, KG96, KCP^{+}94a, KCP^{+}94b, KOW97, KMC96, KS96, KZCS96, KS97, KV98, KAHS96, KK02b, LGM00, LB98, LSZL02, LHCT96, wL94, LFS92, LFS93a, LFS93b, LH95, LC93, LY93, LLY93, LW95, LHZ97, LKL96, LDCZ97, MW98, Man94, MVTP96, Man01, MP95, dlFMBdlFM02, MTSS94, MFTB95, MSCW95, MSP93, Mat94, Mat95, MMU99, Mat01b, MRV00, MK97, McK94, MC98, MFC98, MVY95, MS96b, Mic93, Mic95, MT96, MS99a, MS99b, MHC94a, MHC94b, MRH^{+}96, MS95, MC99, MWO95, Nel93, NP94, Neu94, NBK99, Ney00, NB96, NAJ99, Nov95, Obe96, Ols95, OPP00, Ott94, OWSA95, PPR01, PK98, PPT96b, PPT96a, PPT96c, POL99, PT01, PKYW95]. PVM [Per96, Pet97, PTT94, PGPCK21, Pla02, PNV01, PD98, PY95, PL96, Pus95, QRG95, QRMG96, Qu95, QMGR00, RR00, RS93, Rag96, RS95, RHG^{+}96, RRAGM97, Rol94, RGD97, Saa94, SAS01, Sch94, Sch96a, Sch96b, SB95, SFG98, SGS95, SSS99, SPK96, Sep93, Sev98, Shi94, SA93, SR96, SHH94a, SHH94b, Smi93a, SBR95, SC96a, STT96, SMOE93, SGL^{+}00, SGHL01, SCL97, SSSS97, Sta95b, SY95, SYF96, SC96b, Str94, SKH96, Sun90a, Sun90b, Sun92, Sun93, Sun94a, SGDM94, Sun96, STMK97, SN01, SCL00, Sur95b, Sut96, SL95, TMTP96, TC94, TBD96, TD98, Tsu95, Uhl94, Uhl95b, UH96, UMK97, VSRC94, VSRC95, VB99, VAT95, WKS96, WH94, WCVR96, WAS95b, WO97, Wis96a, WL96a, Wis98, Wis96b, WL96b, WCS99, Wu99, WLC07, XWZS96, XF95, YG96, YKI^{+}96]. PVM [ZPLS96, ZPI06, ZB94, Zem94, ZDR01, ZG95a, ZG95b, ZG96, ZG98, Zol93, van93, NMC95, Ano95b]. PVM-AMBER [SL95]. PVM-Based [WAS95b, FO94, PY95, Sut96, ZPLS96, LSZL02, TD98]. PVM-GRACE [YKI^{+}96]. PVM-Implementation [BJS97, Huc96]. PVM-RPC [KS97]. PVM/C [GTH96]. PVM/MPI [AD98, BDW97, CHD07, CHD09, CD01, DKD05, DLM99, DKP00, DLO03, Kra02, KKD04, LKD08, MTWD06, RWD09, ACGR97, SN01]. PVM3 [IM94]. PVM3/AP1000 [IM94]. PVMaple [Pet00a, Pet00b, Pet01]. PVMe [BR95c, BR95b]. PVMGeant [DZDR95]. PVMPI [FD96, FDG97a, FDG97b]. PyCUDA [KPL^{+}12]. PyMGRIT [HFB21]. PyOpenCL [KPL^{+}12]. pySDC [Spe19]. pySDC-Prototyping [Spe19]. PySPH [RBP^{+}21]. Python [BL97, DPS05, DPSD08, Di 14, DFSW19, GFB^{+}14, HFB21, RBP^{+}21, SSH08]. Python-based [RBP^{+}21]. PyTrilinos [SSH08].

Q [KMH^{+}14, LM13, MV17]. QAPs [Tsu12]. QCD [BLPP13, GM18, SVC^{+}11]. QCG [ACH^{+}11]. QCG-OMPI [ACH^{+}11]. QCMPI [TJD09]. QMPI [EYP^{+}20]. QNSTOP [AEW^{+}20]. QoS [LYGG20]. QoS-Oriented [LYGG20]. QR [GKK09, LC97b]. QSATS [Hin11]. QSW MPI [MW21]. Quadratic [Cza13]. Quadrics [YSP^{+}05, LCW^{+}03]. quadtree [HS95b, PGBF^{+}07, SCC96, Sur95b, TK19]. quadtree/octree [TK19]. qualitative [BLP93]. Quality [Boi97, BDA^{+}18, RFG^{+}00, WHDB05, Ano94i, Lan09, Boi97]. Quality-of-Service [RFG^{+}00]. Quantifying [AKE00, LDCZ97, TPK^{+}19]. quantitative [BLP93, BBH^{+}15]. quantization [HE15]. Quantum [BCGL97, BCL00, GRTZ10, Hin11, MW21, MGG05, NMW93, SK00, SSGF00, TJD09, WHMO19].

Quasi [AEW^{+}20, DDYM99, Pla02, ZB97]. Quasi- [Pla02]. Quasi-asynchronous [DDYM99]. Quasi-Newton [AEW^{+}20, ZB97]. Queens [Rol08b]. Queensland [ACDR94]. Query [AR01]. Quest [MWG97]. Queue [NSS12, CG99b, PTL^{+}16, Sep93, ZA14]. Queueing [COE20]. queues [Man98]. quicksort [MMO^{+}16, MMO^{+}16].

R [Edd18, BBH12, JPOJ12, LR01, Mat16, TSN21]. R&D [Str94]. R&D-100 [Str94]. Rabi [MBA21]. Rabi-coupled [MBA21]. Race [CFMR95, KSJ14, DKF94a, PGD18]. Races [PPJ01, SAL^{+}17, DKF94b, LLG12, ZRQA11, EPP^{+}17]. Radial [RB01, KRC17]. Radiance [GCBM97, KMG99, RC97]. radiation [NS20, SCJH19]. radiative [RS22]. Radiology [GA96]. Rajeev [Ano00a]. Raleigh [Agr95a]. Ramesh [Stp02]. Random [HT08, LTDD14, MPZ21, CCS19, Lan09, RJH^{+}20]. Randomized [DSU20, Tra98]. Range [KBM97, MH01, BMPZ94a, PARB14, She95]. range-join [She95]. Rank [Hat98, ZLWW20]. Ranking [Tra98, RJH^{+}20]. Rapid [FWS^{+}17]. RASC [YCL14]. rate [BBG^{+}14, MNYN21, YPA94]. rationale [BBH^{+}13b]. Raton [Edd18]. Ray [CG93, DP94, KGB^{+}09, FWS^{+}17, SGS95, FFB99]. Ray-Tracing [DP94]. Rayleigh [TVV96]. Rayleigh-Benard [TVV96]. rCUDA [CPM^{+}18, IPG^{+}18, PRS16, PS19b, PIR^{+}20, RSC^{+}15, RPS19, RS19, RS21, SIRP17, SPBR20]. RDMA [GSY^{+}13, LWP04, Pan14, RA09]. RDMA-Based [LWP04]. RDMA-Enabled [GSY^{+}13, Pan14, RA09]. Re [MCP17]. Re-Vectorization [MCP17]. Reaching [BHS^{+}02]. Reaction [HF14a, HF14b]. Reactive [BCL00, KSB^{+}20, Heb93]. reactor [ANS95]. Read [SSLMW10]. readability [SM12]. Reading [HK95]. Ready [Bri02, DZ98b].

Ready-Mode [Bri02]. Real [ASB18, LHLK10, NSLV16, PRQ21, SM19, SGL^{+}20, TWLL19, Tho94, UP01, YGH^{+}14, Ano94f, Fer04, FLB^{+}05, JR10, ZWZ^{+}95, SKD^{+}04]. Real-Time [SGL^{+}20, TWLL19, UP01, YGH^{+}14, ASB18, LHLK10, PRQ21, SM19, Fer04, ZWZ^{+}95, SKD^{+}04]. Real-World [NSLV16]. Realistic [YMYI11, ZSnH01, CKP^{+}93]. Reality [ACM96a, Ano93f, NM95, Wit16]. realizing [YZ14]. Reallocation [GFIS^{+}18]. rebooting [GJLT11]. Receive [Bri02]. Receiver [ZG95b]. receptor [ESB13]. Rechnen [Ano94c, BL94, MS04]. Recognition [CC17]. recomputation [RKBA^{+}13]. Reconfigurable [FDG19, HGX^{+}22, MFC98, SPM^{+}10, SSK^{+}18, ZL18, NYNT12, RRJ^{+}20]. Reconfiguration [CS14, MSMC15]. Reconstruction [BM97, DYN^{+}06, GA96, LSSZ15, OIH10, RAGJ95]. Record [UALK17, UALK19, CRD99]. Record&Replay [KSV01]. record/replay [CRD99]. Recovery [SBF^{+}04, BBH^{+}13b, BDB^{+}13, LFS93a, LFS93b, LZZ^{+}20, SSCC95, SRS^{+}19, ZWZ05]. Rectangle [CSW99]. rectified [WBBD15]. Recurrences [ACGR97, MB18]. Recursive [DSS00, PWP^{+}16, SML19, SD99]. Red [van93]. redesign [HL17]. Redesigning [RS21]. Redistribution [DDPR97, HC06, WO95, WO96, HC08, KN95]. Reduce [CBB^{+}20, PSM^{+}14]. Reduced [SW12]. Reducing [AV18, CRGM16, JE95, BCM11]. Reduction [DAD19, FKH02, MFPP03, SG12, HL17, Jes93a, MLVS16, Pan95a, PQ07]. Reductions [PWPD19, ZCBD22]. Redundancy [TS12a]. redundant [KJJ^{+}16]. Reference [GHLL^{+}98, Nag05, SOHL^{+}98, YM97, Ano99a, Ano99c, Ano99b, Ano99d, SOHL^{+}96, Per97, Ano96a]. Refine [ZWLZ21]. Refinement [MRB17, Ran05, CLSP07, DLR94, SWCB20]. region [SPNB14]. region-based [SPNB14]. regions [LFL11]. Registration [WYZ^{+}19]. regression [RBAI17]. Regular [HLP11, NHT02, NHT06]. Reims [MCdS^{+}08]. rejection [SOYHDD19]. RELAP5 [SBR95]. related [SD16]. Relating [EPML99]. relation [DO96, Hem96]. Relationship [Dan12]. relativistic [BHS18]. relaxation [OKW95]. Reliability [CGZQ13]. Reliable [SE02, Arn95]. RELION [ZWLZ21]. remapping [LW20]. Remark [SWH15]. remedies [ALW^{+}15]. Remo [IEE95h]. Remote [BMR01, HDT^{+}15, IFA^{+}16, OCY^{+}15, Tsu07, WBBD15, AGLv96, CPM^{+}18, FHC^{+}95, GBH14, GBH18, HGMW12, RSC^{+}15, SIRP17, SH96]. Remote-Scope [OCY^{+}15, WBBD15]. Remotely [GGCM99, GGCGO01, GCGS98, VLO^{+}08, GGGC99]. Remoting [MGL^{+}17]. removal [ZZZ^{+}15]. Removing [ZJDW18]. Rendering [DLLZ19, DLLZ20, GCBM97, LSZL02, SU96, UCW95, KB21]. Rendezvous [RA09]. Reordering [Hat98]. Reparallelization [KBG^{+}09]. Reparameterisation [RJ21]. Repeated [WH94, Shi94]. repeats [GDMME22]. Replacement [GHD12]. Replay [CFMR95, HLOC96, UALK17, UALK19, CRD99, MT96, NBK99, XLW^{+}09]. replay-based [MT96]. Replication [WC09, KJJ^{+}16, ZJDW18]. Replication-Based [WC09]. Report [DZ98b]. Reports [Ano98, ACM11]. Representation [BMR01, KD12, MDM17, SML17, SML19, CCM12, SBB20]. Represented [SSB21]. reproduce [AVA^{+}16]. reproducibility [HD00a]. Reproducible [GL99, HCA16, XLW^{+}09]. Requests [KLH^{+}20]. Requirements [GSHL02, GT07, LPJ98, Ber96, KBG16, LCVD94a]. Research [Ano96d, BR02, MC94, SL94a, SGHL01, Ara95, BPG94, LP00, Oed93]. Reservoir [KDHZ18, OWSA95, ZAFAM16, ZZ95, Ano95d]. Resident [JDB^{+}14]. Resilience [YNJS21]. resiliency [RGP22]. Resilient [CGH^{+}14, Gua16, LCMG17, Pro21, LMG17, LBB^{+}19, MLVS16]. Resistive [ZL17]. Resolution [MAB05, Str94, TPV20, ZWLZ21, BADC07, KN17]. Resolving [Str97]. Resource [BGR97b, BSH15, KK98, SIS17, YSS^{+}17, BMS19, DZ96, FLD96, FL21, NEM17, PIR^{+}20, ZA14]. resource-conscious [ZA14]. resource-restricted [NEM17]. Resources [LSB15, NAW^{+}96, WYZ^{+}19, Kos95b, RSC^{+}19, R^{+}92]. Response [BBC^{+}00]. Responsibility [KQT^{+}21]. Restart [SSB^{+}05, AKB^{+}19, CZP21, LMG17]. restarted [dH94]. Restoration [FJBB^{+}00]. Restore [Gua16]. Restricted [JCP^{+}20, NEM17]. Restructuring [KAMAMA17]. Results [BIL99, BIC05, HSMW94, Wal01a, BR95c, DHS96, VDL^{+}15]. retargetable [KKJ^{+}08]. rethinking [GJLT11]. Retrieval [Per21, RLL01, MMR99, MRH^{+}96, RTL99]. reusable [LTLC94]. reuse [BAE22, BVML12, LM94, NAAL01]. Reverse [BGK08, HHSM19, LSB15, LM13, QHCC17]. Reverse-mode [HHSM19]. Review [Ano95b, Ano95c, Ano96a, Ano99a, Ano99c, Ano99b, Ano99d, Ano00a, Ano00b, BDL98, Che10, Edd18, Mar06, MCLD01, Nag05, NMC95, Per96, Per97, SD13, Vre04, AMKM20, Stp02, Vog13]. Reviews [Ano97, Bra97, YM97]. Revised [Cha05]. Revision [MHSK16]. rewrite [HLK^{+}20, SFLD15]. REYES [LSZL02]. RFSA [SW12]. Rhine [Cal94]. Rhodes [TG94]. RHODOS [RGD97]. Rich [MKW11]. Riemann [PGPCK21]. Right [ZG95b]. Rim [IEE95e]. ring [ZZZ^{+}15]. RISC [AL93, NMW93, BSvdG91]. RMA [BBW19, FCS^{+}19, SPH^{+}18]. RNA [WHDB05]. RnaPredict [WHDB05].

Robert [Ano95b, NMC95]. robotic

[ZWZ^{+}95]. Robust

[Att96, GR07, LSB^{+}18, PSLT99]. Rocks [PKB01, Slo05]. Roe [PGPCK21, dlAMCFN12]. Rohit [Stp02]. Roll [DDN^{+}22]. Rollback [DDN^{+}22, LZZ^{+}20, LBB^{+}19]. rollback-recovery [LZZ^{+}20]. rolling [NF94]. Rome [CMMR12]. Roothaan [MMDA19]. roots [PNV01]. rotating [KLM^{+}19, MBA21]. routed [Pan95b, RJMC93, ZGN94]. routers [Jes93a]. Routines [Add01, Sch96a, LSK04, Sch96b, VLMPS^{+}18]. Routing [BHM94, BHM96, MTSS94, MBES94, WH94, BS94, Zah12]. RPC [KZCS96, KS97, RS93, SHTS01]. RPVM [CMM03, LR01]. RS [BGBP01, Cou93, Heb93, MW93]. RS/ [Cou93, Heb93, MW93]. RS/6000 [BGBP01]. RS6000 [CDM93]. RSA [WLC07]. RT [KAMAMA17]. RT-1.1 [SKD^{+}04]. RT-CUDA [KAMAMA17]. RTL [BGG^{+}15]. RUBIS [BR94]. Ruby [Ong02]. rules [SFLD15]. Run [CBB^{+}20, CBB^{+}21, DLR94, DGMJ93, FHK01, GOM^{+}01, OP98, SBW91, SPB^{+}17, SS96, KPL^{+}12, RRG^{+}99, Str94, TCBV10]. Run-Time [CBB^{+}20, CBB^{+}21, FHK01, GOM^{+}01, OP98, SPB^{+}17, SS96, DLR94, SBW91, KPL^{+}12, TSY99, TCBV10]. Running [BZ97, CCM^{+}06, YKI^{+}96, CRE01, ZLZ^{+}11]. Runtime [AAB^{+}17, BGD12, CFF^{+}94, DMB16, DT17, DSCL05, Gro00, KBS04, KCR^{+}17, NPP^{+}00d, PG18, SdR^{+}21, TJPF12, YSS^{+}19, ZLP17, AKB^{+}19, ADK22, ALW^{+}15, BL99, BR94, EPP^{+}17, EO15, HPS^{+}12, HPS^{+}13, KW14, LRLG19, LLH^{+}14, MA09, NPP^{+}00a, TSY00, YÁJG^{+}15]. Runtime-compilation [PG18]. Runtimes [AHHP17]. Russia [Mal95]. RWA [RLVRGP12].

S [AHHP17, Röh00]. S-Caffe [AHHP17]. S-language [Röh00]. S1 [GLT00b]. S3D [LSG12]. SAEO [GSYT21]. Safe [Pla02, GCC99, LFS92, LFS93a, LFS93b, NYNT12]. Safety [CLA^{+}19, GT07]. salesman [GM94]. Salt [Hol12]. sampling [CBS18, SOYHDD19, WLYL20]. San [ACM97b, Ano95d, BBG^{+}95, GE95, GE96, Has95, IEE93a, IEE94g, IEE95h, IEE95g, IEE97c, LF^{+}93a, NM95]. Sanders [Che10]. Sandy [VDL^{+}15]. Santa [ACM95b, AH95, IEE95f, Old02, RV00]. Santorini [CD01, CDND11]. Santorini/Thera [CD01]. Saphir [Ano99c, Ano99d]. SAR [AB95]. Satellite [Uhl94, Uhl95b, SSN94]. Satisfiability [IKM^{+}01, IKM^{+}02]. saturated [TOC18]. Saturday [B^{+}05]. Saturday-Wednesday [B^{+}05]. Save [ADGA20, KFL05, FKLB08]. Saving [CBB^{+}21]. SBS [MSB97, WWZ^{+}96]. SBS-Type [MSB97]. SC’11 [LCK11]. SC2000 [ACM00]. SC2001 [ACM01]. SC2002 [IEE02]. SC2003 [ACM03]. SC97 [ACM97b, ACM97b]. SC98 [ACM98b, ACM98b]. SC’99 [ACM99]. Scalability [Ben18, BS07, FSC^{+}11, KBS04, LL01, LKYS04, LSK04, VLSPL19]. Scalable [Add01, AHHP17, BHW^{+}17, BBC^{+}02, BHNW01, BGL00, CGS15, CLE^{+}20, CDPM03, EFR^{+}05, GFB^{+}14, GS94, HC17, HGMW12, IEE92, IEE94f, IEE95j, IBC^{+}10, KTAB^{+}19, KK98, LTS16, kLCC^{+}06, MFPP03, NBGS08, NPP^{+}00d, NCKB12, NSM12, OLG01, PPJ01, PR94b, PBK00, SDJ17, SBF^{+}04, Skj93, SS96, TPD15, TPV20, UP01, VBLvdG08, VY02, ZLGS99, ZL18, BBB^{+}94, Bri95, CLSP07, FWS^{+}17, GBH14, GBH18, GM13, GKL95, HRR^{+}11, HAJK01, KRC17, KRG13, LM99, LTLC94, MMB^{+}94, MRRP11, PWD^{+}12, SPK^{+}12, Trä12a]. ScaLAPACK [BV99, BRR99, DHP97]. Scale [AKE00, AFGR18, BHW^{+}17, BZ97, BHNW01, CBB^{+}20, FFP03, HC17, MFPP03, SM03, TGEM09, WMC^{+}18, WT12, AASB08, BKK20, BCA^{+}06, BJS99, BCH^{+}08, Che99, DZZY94, FME^{+}12, Gua16, IPG^{+}18, Kos95b, LS10, MLA^{+}14, NWT21, PTL^{+}16, PD11, RMNM^{+}12, SIC^{+}19, SvL99, TBB12, WLNL06, WT11, WT13, ZKRA14, ZA14, Ben18]. SCALE-EA [Ben18]. Scale-Out [AFGR18]. Scale-Up [AFGR18]. SCALEA [TFGM02]. Scaling [CC17, GDS^{+}20, KFL05, SLJ^{+}14, FKLB08, Gao03, LFL11, PDY14]. scan [AAAA16, YLZ13]. scanline [CT13]. scans [NAJ99]. SCASH [SHHI01]. SCATCI [ART17]. scatter [BCD96, MTK16]. Scattering [BCL00, NZZ94, OMK09]. SCF [MM95]. schedule [NAAL01]. scheduler [ADDR95, TCBV10, WRSY16]. schedulers [AV18, NP12]. Scheduling [BBH^{+}06, BSH15, CML04, DMB16, EGR15, GDDM17, GSHL02, GHL97, HC06, JW96, MJB15, NIO^{+}02, NIO^{+}03, SM19, SNN^{+}20, SGL^{+}20, TJPF12, WJG^{+}21, APBcF16, DZ98a, HC17, JKN^{+}13, KSC^{+}19, LHCT96, MBKM12, NSBR07, OPW^{+}12, Smi93b, SKK^{+}12, SKB^{+}14, WYLC12, WLYC12, YWC11]. Scheme [CTK01, LNLE00, MW98, SBF^{+}04, Bae20, BBGL96, Bjo95, MRRP11, OKM12, SCC96, YPZC95, FM90]. Schemes [HC17, PPJ01, MPS20, WYLC12, WLYC12, ZAT^{+}07]. Schmidt [CBYG18]. School [VV95]. Schrödinger [DM12, ÓN12]. SCI [FS97, HEH98, Hus00, RR01, ZHS99]. SCIDDLE [ABG^{+}96, AGLv96]. SCIDDLE-PVM [ABG^{+}96]. Science [Edd18, EGH^{+}14, IEE95d, Mat16, MMH93, Old02, SM07, ACM06a, DMW96, HK93]. Sciences [ERS96, HS94, ZL96, ERS95]. Scientific [AGH^{+}95, APJ^{+}16, BBG^{+}95, DKM^{+}92, DT94, Gat95, GL97a, HJ98, KK02a, LWSB19, LkLC^{+}03, Mar06, Nag05, Sin93, SSB^{+}17, VY02, WN10, ACC^{+}21, Bis04, DW94, SBG^{+}12, SIC^{+}19, TBB12, WT13, Ano97, Bra97]. scientists [HW11, Str94]. SciPAL [KH15]. SCIPVM [ZHS99]. Scope [OCY^{+}15, BDB^{+}13, WBBD15]. scoping [RDLQ12, WC15]. Scottsdale [IEE95b]. Scratchpad [JAK17, MB12]. Scripting [Ong02, KPL^{+}12, Nob08]. scripting-based [KPL^{+}12]. SCTP [KPW05, ZPI06]. SDK [TK16]. SDSM [CCM^{+}06]. Sea [LPJ98]. Seamless [KK02a, LdSB19]. Search [BSH15, Cza13, IKM^{+}01, Wal01b, WTS19, FMS15, IKM^{+}02, RRJ^{+}20, Wal01a, ZSK15, CB11]. Searches [BSG00]. Searching [JPT14, MM01, BA06, Wal01b]. Seattle [ACM05, BS94, LCK11, Ost94]. Second [Ano00b, BL95, DT94, DE91, IEE94d, IEE96d, IEE96i, LHHM96, Tou96, Vol93, WPH94, ACM97a, Ano99a, Ano99b, BFMR96, DMW96, FR95, KN17, Li96]. Second-Order [BL95, KN17]. Secondary [WHDB05, SEC15, ZAT^{+}07]. section [Ano93b, DKD08]. segment [FJZ^{+}14]. segment-based [FJZ^{+}14]. Segmentation [KBA02, AD95, CCU95]. Seidel [BG95, LM99, Ols95]. Seismic [OWO98, AMBG93, KL95, KEGM10, LM13, QHCC17, RMNM^{+}12, SSS99, WCVR96]. Seismograms [DP94]. Select [KKDV03]. Selected [DHS96, MTW07, OL05, TB14, TH20, CHD09, Cha05, DKD07, JC17, KC19]. selecting [PTL^{+}16]. Selection [CKmWH16, SNN^{+}19, GDEBC20, PGBF^{+}07, WKS96, ZWL^{+}17]. Selective [Nak03]. Self [HC17, NSS12, SLJ^{+}14, TGT10, VFD02, NSBR07, WYLC12, WLYC12, YWC11]. Self-Consistent [TGT10]. Self-scheduling [HC17, NSBR07, WYLC12, WLYC12, YWC11]. Self-Submitting [NSS12]. Self-Tuning [SLJ^{+}14]. Semantic [EADT19, MTU^{+}15, DKF94a, OA17]. Semantically [MKW11]. semantics [RNPM13]. Semaphores [TTP97]. Semi [CT94a, Bjo95, PSLT99, TC94, CT94b]. semi-coarsening [PSLT99]. semi-implicit [Bjo95]. Semi-Lagrangian [CT94a, TC94, CT94b]. Semiconductor [GJN97, Ano03, LS10]. Seminar [Ano94f, Ano93h]. Send [GPC^{+}17]. Sender [BCH^{+}03]. Sensed [GGCM99, GGCGO01, GCGS98, VLO^{+}08, GGGC99]. sensitive [GKCF13]. Sensitivity [dLR04]. Separable [Ben01, CdGM96]. September [Abr96, AD98, Ano93a, Ano93b, Ano95a, Bos96, BP93, BH95, CLM^{+}95, CHD07, CJNW95, CD01, CDND11, DKD05, DKD07, DLM99, DKP00, DLO03, EJL92, FK95, FR95, GHH^{+}93, IEE93d, IEE94c, JPTE94, KGRD10, Kra02, KKD04, LKD08, Mal95, MTWD06, OL05, PSB^{+}94, RWD09, SPH95, SM07, TBD12, VV95, VW92, WPH94, YH96]. Sequence [GMU95, SMM^{+}16, AMHC11, TSZC94]. sequences [dFdOSR^{+}19, GÁVRRL17, SdM10]. Sequencing [VPS17]. Sequential [EK97, RPM^{+}08, GGH99, SR95, TNIB17, TSZC94]. Serial [SWH15, HPS^{+}96, HWS09]. serialization [CFKL00]. Serialized [KH10]. Serielles [BL94]. Series [Nag05, BR94]. Server [Ano93f, AFGR18, FSLS98, KS97, Mat01b, Sch93, Sto98, Vis95]. Server-Class [AFGR18]. serverless [NRdA^{+}20]. Servers [CGC^{+}02, SIS17, GK97]. Service [RFG^{+}00, LS08, SPK^{+}12]. Services [FC05, LSB^{+}18, AAC^{+}05, ZKRA14]. Session [NYNT12, ZL96]. Set [BDA^{+}18, SW12, WL96a, Ano00a, Ano00b, PSH^{+}20, She95, THMH21, WL96b]. Sets [SG12, CGL^{+}93]. setting [GL95a]. Setup [NSLV16]. Seventh [BBG^{+}95, HS94, IEE93b, IEE95g, IEE96h, Eng00, Y^{+}93]. several [GBR15]. SGI [Che99, CML04, KMG99, LB96, LL01, LKJ03, LSK04, TW12, ZSnH01]. SGI/CRAY [Che99]. SGI/CRAY-T3E [Che99]. shadow [SOA11]. shallow [DS22, STA20, dlAMC11, dlAMCFN12]. Shane [SD13]. Shanghai [IEE97a]. shaped [ZWC21]. SHARE [Ano92, Ano93f, Ano94g]. Shared [ADGA20, BCA^{+}06, BME02, Bri10, CDT05, DM98, DMB16, FKH02, FB94, GB96, GLRS01, HC10, HDB^{+}12, HT01, KB98, KSHS01, LRT07, Luo99, MBE03, MCdS^{+}08, Mül02, NPP^{+}00d, PBK00, Pok96, PS00b, Ros13, SS01, STY99, ST02b, Thr99, VS00, VT97, ABCI95a, ABCI95b, ADMV05, BMG07, CdOO^{+}20, CBPP02, CJvdP08, Cha96, CCM^{+}06, CC00b, DBVF01, DS96b, DPZ97, EVMP20, EV01, GCN^{+}10, GL96, GL97c, HS93, HDB^{+}13, JE95, KJA^{+}93, KC06, LKL96, MLC04, PK05, QM21, RGDM15, SHHI01, SL94b, SFL^{+}94, SSC96, TSY99, TSY00, THDS19, Vos03, WLYL20, WK20, WMRR17, WRMR19, YWO95, YX95, Cha05]. Shared-Memory [DM98, HDB^{+}12, NPP^{+}00d, Pok96, Thr99, PS00b, ABCI95a, ABCI95b, BMG07, CdOO^{+}20, EVMP20, GL96, GL97c, KJA^{+}93, PK05, TSY00]. shared/distributed [THDS19]. Sharing [Att96, CML04, CB16, DiN96, JAK17, KK98, LYGG20, JE95, Ott93, PRS^{+}14]. shear [JAT97]. ShearLab [KLR16]. Shearlet [KLR16]. Shearlets [KLR16]. Shelf [LPJ98]. SHMEM [BBDH14, Hus01, LSK04, Sch96a, Sch96b, SS01]. Short [KBM97, MH01, SSLMW10, BMPZ94a, PARB14]. Short-Range [KBM97, MH01, BMPZ94a, PARB14]. Short-Read [SSLMW10]. shorter [NB96]. Showcase [USE00]. SHPCC [IEE92]. SHPCC-92 [IEE92]. SIAM [BBG^{+}95, DKM^{+}92, Sin93]. Side [kLCCW07]. Sided [BPS01, GFD03, GFD05, GT01, HDB^{+}12, LRT07, MH01, MB00, TGT05, TRH00, ZSG12, bT01a, BM00, DPFT19, DBB^{+}16, GBH18, LSK04, MS99c, PGK^{+}10, GBH14]. SIGCSE [ACM06a]. Signal [IEE95e]. signals [Uhl95c]. Signatures [Gro00]. significance [AMHC11]. silent [FME^{+}12]. Silicon [LHZ^{+}20, Ano03, Goe02, ZL18].

Silicon-Monona [ZL18]. SIMD

[BvdB94, HS95b, KDT^{+}12, LL16, Sur95b, VSW^{+}13, WMK^{+}19, vdP17]. similarity [LSB^{+}20]. Simple [MSF00, Mül01, SC04, BC19b, ITT99, JH97, JKN22, Nes10, PGPCK21, PNV01]. simulate [Heb93]. Simulated [BHM94, BHM96, FH97, MPZ21, RSBT95]. Simulating [DLM^{+}17, KDL^{+}95b, KDL^{+}95a, NFG^{+}10]. Simulation [CDMS15, CCBPGA15, DMMV97, DZDR95, GSI97, GM95, GJN97, Ham95a, JML01, KDHZ18, KBM97, KMK16, LLRS02, MFTB95, MPD04, MANR09, PCY14, PKYW95, PZKK02, RR00, RDMB99, SSAS12, SXMX^{+}18, Str97, Ten95, UZC^{+}12, VT19, WMC^{+}18, ZZ04, ZWJK05, dlAMC11, ASAK19, Ano95d, ADR^{+}05, BJ95, BCM^{+}16, BH95, BMPZ94b, CwCW^{+}11, CSPM^{+}96, DSOF11, FHSO99, FO94, FLPG18, FFFC99, GRTZ10, IPG^{+}18, JAT97, JLS^{+}14, KTJT03, KNH^{+}18, KMC96, KMC97, LFS^{+}19, LHZ^{+}20, LCVD94b, LCVD94a, LYZ13, MMW96, MW21, MALM95, NS20, NB96, NF94, OKM12, PARB14, PY95, RFH^{+}95, SWYC94, SSP^{+}94, SKM15, Str96, Syd94, Tho94, WHMO19, WGG^{+}19, YPA94, YEG^{+}13, YSL^{+}12, Eng00]. Simulation-Based [ZWJK05]. Simulations [CGS15, CNM11, DFMD94, DI02, GAP97, HLP11, HF14a, HF14b, KT02, Kha13, NH95, RTRG^{+}07, SM02, YPAE09, ADT14, ABG^{+}96, BHS18, BADC07, CFF19, GM18, Hin11, JMS14, LS10, LSVMW08, MNYN21, RMNM^{+}12, SU96, THDS19, TOC18, VLSPL19, WWFT11]. Simulator [CAM12, MRV00, PHO^{+}15, UTY02, WPC07, AMV94, LS10, LZC^{+}20, PWD^{+}12, WZWS08, ZAFAM16, ZZ95, KTJT03, Nak03, Nak05a, Nak05b]. Simulators [SB95, AVA^{+}16]. Singapore [IEE96d]. Single [BM00, HF14a, HF14b, MB00, URKG12, WZM17, AGIS94, KKLL11, LK20, MKP22, THMH21]. Single-Chip [URKG12]. Single-sided [BM00]. Single-Threaded [WZM17]. single/multigrid [AGIS94]. singleton [TVCB18]. Sinks [JPT14]. Sites [Ano98]. Sixth [HK95, IEE96c, MMH93, SW91]. Size [WQKH20, YT20, GKCF13]. sized [JLS^{+}14]. Sizes [DALD18, ZSnH01]. Sizing [YNJS21]. SKaMPI [KRS99, RSPM98, RH01, Reu01, RST02, Reu03]. SkelCL [SG14]. Skeleton [GB98, IH04, RJDH14]. Skeletons [Ser97]. Skew [GGZ^{+}20]. Skew-Tolerant [GGZ^{+}20]. Skjellum [Ano95c, Ano00b]. Slack [CBB^{+}20, KFL05, FKLB08]. SLAE [ADRCT98, AK99]. sLASs [VLCM^{+}20]. Slave [LTR00, HP05]. SLEPc [DR18]. SLICC [KBHA94]. Slices [GSHL02]. Slim [WMC^{+}18]. Small [HLP11, TS12b, Ano94h]. small-footprint [TS12b]. Small-World [HLP11]. Smith [KDSO12, RGB^{+}18]. Smithsonian [Str94]. smoking [YSL^{+}12]. Smoothed [RBP^{+}21]. SMP [Add01, CRE99, CRE01, CCBPGA15, HD02a, DK06, GT01, GMdMBD^{+}07, HD02b, Hus00, HIP02, JKHK08, KOI01, KKH03, KMG99, KAC02, NO02b, NO02a, ST02a, TOTH99, Trä02b, YWC11, bT01a]. SMPCkpt [DCH02]. SMPI [DLM^{+}17]. SMPs [HLCZ00, NU05, SvL99]. SMPSs [MLAV10]. SMPSuperscalar [GCBL12]. SMT [PAdS^{+}17]. SMT-based [PAdS^{+}17]. snake [JPP95]. snake-in-the-box [JPP95]. SNE [MPZ21]. Snir [Ano96a, Ano99a, Ano99c, Ano99b, Ano99d, Nag05]. SnuCL [Lee12]. soccer [YMYI11]. Socket [COE20, Gro19, LS10]. SoCs [AFGR18]. Soft [AJYH18]. Softshell [SKK^{+}12]. Software [Ano94i, BKK20, BME02, BPG94, BDG^{+}xx, CZ95b, DGH^{+}19, ESB13, FFP03, GBF95, Gre95, HPR^{+}95, HS94, HHA95, IEE95l, IEE96h, IFI95, KS15a, KC94, KAMAMA17, KG93, LB16, MBE03, NPS12, Ost94, PZ12, Sil96, STH22, Swa01, TDBEE11, VdS00, Wis01, Wol92, Ano97, BSC99, Boi97, Bra97, BR94, CMV^{+}94, CBPP02, DPZ97, Hum95, JH97, JB96, LSB^{+}20, LM94, MK94, Neu94, Old02, PHA10, PK05, PGK^{+}10, RAS16, RJH^{+}20, SHHI01, Sch94, Sei99, SPH95, SSD^{+}20, Str94, WGG^{+}19, ZGN94, Ano94i, KG93, Sil96]. Software-Managed [LB16]. Solan [CGB^{+}10]. Solaris [Ano01a]. Solidification [HSO^{+}21, JLS^{+}14]. solids [Hin11]. Solution [DWL^{+}10, FBSN01, HO14, MC18, RPM^{+}08, SEF^{+}16, SSK^{+}18, Tsu12, VRS00, DWL^{+}12, GADM20, IM95, JK10, LGM^{+}20, LSR95, MALM95, ÓN12, PRS^{+}14, SC96a]. solutions [AGIS94, LMG17]. Solve [Hog13, LSM^{+}18, Riz17, BAV08, Che99, GGGC99, TSCS14]. Solver [Ben01, BP98, CF01, CF19, HSMW94, IDD94, LZ97, SJK^{+}17a, SJK^{+}17b, TPV20, WJB14, YKW^{+}18, AMS94, CP15, CFF19, DS22, DM12, GNP19, HDZ^{+}20, HHSM19, JR10, LM99, Lou95, MV20, MBA21, OGM^{+}16, RM99, STA20, SRK^{+}12, SCC95, THM^{+}94, ZZG^{+}14]. Solvers [DFN12, DALD18, GK10, MSB97, NO02b, Nak03, NHT02, NLRH07, QRMG96, RS97, SSK^{+}18, WR01, ABF^{+}17, ADLL03a, ADLL03b, ADDR95, BRR99, CL93, DR18, EVMP20, MKP^{+}96, MS95, NO02a, Nak05a, Nak05b, NHT06, PGPCK21, PR94c, QRG95, SSH08]. Solving [ADRCT98, BHM94, BHM96, BV99, BG95, BDG^{+}92c, BSH15, DALD18, DAD19, GFPG12, Huc96, LLY93, MS02a, NF94, SAS01, SP11, SD99, ZTD19, BB95a, DSM94, HHA95, LBB^{+}16, LYSS^{+}16, MM11, SSB^{+}16, SMSW06, YSVM^{+}16, YSMA^{+}17]. SOM [GkLyCY97]. Some [BDT08, Mül01, Pet97, AL92, NN95, RSBT95]. Sopron [VV95]. Sorrento [DKD05, DKD07]. sort [KVGH11, PSHL11]. Sorting [Ger18, LTS16, BHJ96, PSHL11]. Sound [SG12]. Source [ABB20, BGG^{+}15, HH22, MM07, AC17, AVA^{+}16, LSB^{+}20, NCB^{+}17, Nob08, PSK^{+}10, WGG^{+}19]. Source-Code-Correlated [MM07]. Source-to-Source [HH22, ABB20, AC17]. Sources [CTBT21, ZDR01, KM10]. South [ACM95a]. southeast [ACM95a]. Sowing [GL97a]. SP [BGBP01, CE00, HMKV94, LC97b, WT11, WT12]. SP-1 [HMKV94]. SP-2 [LC97b]. SP1 [BR95c, FHPS94b, FHP^{+}94, FHP^{+}95, Fra95, FWR^{+}95, GL95d, HSMW94, MP95]. SP1/SP2 [FHP^{+}95, Fra95, FWR^{+}95]. SP2 [BR95b, FHP^{+}95, Fra95, FWR^{+}95, HWW97, JF95, KB98, KHS01, MABG96, XH96]. SPAA [ACM95b]. Space [CML04, CB16, HO14, MSF00, MZLS20, OFA^{+}15, SAS01, SS01, TA14, SRK^{+}12]. Space-Sharing [CML04]. Space-Time [HO14, SRK^{+}12]. Spaces [Rót19]. SPAI [BBS99]. Spain [DLM99]. SPAN [LHHM96, Li96]. Spanish [VP00]. spanning [NCKB12]. Spark [GRW^{+}19, KWEF18]. Sparse [AZ95, BBH12, CWL^{+}20, DS13, DK20, Huc96, MYL21, NHT02, TD98, ZB97, AK99, ADLL03a, ADLL03b, BAC20, ER12, FJZ^{+}14, GG99, Gra09, NHT06, XXL13]. SPEC [Ano03, MvWL^{+}10, MBB^{+}12, NA01, SGJ^{+}03, TSB03]. Special [AM07, BDT08, BC19a, BDB^{+}13, BC00, CHD09, DKD07, DKD08, GSA08, GT19, MPI98a, MPI98b, SBG20, TH20, Bos96, Mar02, PNV01, Reu01, Old02]. Specific [DM95b, DM95a, Olu14]. Specification [BG94a, BdS07, MGC12, MHSK16, BG94c, LPD^{+}11]. Specifications [OFA^{+}15, WMP14]. Specified [MGMH97]. specifying [LPD^{+}11]. specimen [Rol08b]. SPECT [BCD96]. spectator [YMYI11]. Spectra [Str97, SR11]. Spectral [MW98, Spe19, BCM^{+}16, MGS^{+}15]. spectral/ [BCM^{+}16]. spectrum [NS20]. Speculation [AELGE16, SHLM14]. Speculative [RA09, dOSMM^{+}16]. Speed [CDHL95, Tou00, AH95, Ano03, BWT96, BID95, KMK16, CDH^{+}95]. Speeding [CSV12, YNJS21]. Speedup [VPS17]. SPH [AFG21, CP15, OLG^{+}16, PBC^{+}01, WMRR17, WRMR19]. Sphere [CT94a, CT94b]. spherical [Hol95, KT10]. SPICE3 [WPC07]. Spiking [CAM12]. Spin [HLP11, JRG21, KO14, Kom15, MBA21]. spin-1 [MBA21]. spin-glass [JRG21]. spin-orbit- [MBA21]. splitting [MPS20, TCBV10]. SPMD [BST^{+}13, Dar01, KAC02, Wal00, Wal02]. SPMD-Like [BST^{+}13]. SpMV [CBIGL19]. Spokane [IEE93c]. Sponge [HSW^{+}12]. spontaneous [EZBA16]. spreading [SOYHDD19]. Spring [Ano94g, IEE93a]. SPTHEO [Sut96]. SPY [SSG95]. Squares [PWP^{+}16, VRS00]. SR [YWCF15, ZLP17]. SR-IOV [YWCF15]. SR8000 [NNON00, TSB02, TSB03]. SRP [BBC^{+}19]. SS7 [LTLC94]. SSGM [HPS^{+}96]. SSS [MMH98]. SSS-CORE [MMH98]. St [Mal95]. Stability [DSS00, HD00a]. stable [JMdVG^{+}17]. Stage [FSXZ14]. stages [KW20, SRS^{+}19]. staggered [GM18]. Stampi [ITKT00]. stamping [DPFT19]. Standard [DM98, GSI97, GLP^{+}00, GL95c, Hem94, MPI98a, MPI98b, NH95, SKD^{+}04, SGS10, Wer95, YKLD17, Ano94d, BDB^{+}13, Bor99, Cla98, CG99b, DHHW93b, DOSW96, FB95, GK97, GL92, Hem96, Sti94, VM95, Wal94a, Wal94b, WD96, Ano97, Bra97, CGH94, DOSW95, GLDS96]. Standards [FKKC96, Thr99]. Star [CDM93, Coo95a, Coo95b]. STAR/MPI [Coo95a, Coo95b]. Start [Gro02b, Hus98]. Startup [PS07]. State [ACM11, IEE94f, IEE95j, Wis96a, Wis96b, BTC^{+}17, LF93b]. state-to-state [BTC^{+}17]. states [NS16]. Static [NIO^{+}02, NIO^{+}03, RLVRGP12, SCB15, SCB14]. Static/dynamic [SCB15]. Statics [TG94, TG94]. Stationary [MW98]. Statistical [LR01, SNMP10, AMHC11, KKM15, Röh00, SL94a, Vet02]. statistics [FL21]. Status [Bak98, DF21, DZ98b, GL95c, BDG^{+}93b, FHP^{+}95, Hem96, Sun96]. stealing [TCBV10]. Steepest [Sch01]. Steering [GKP97, PK98]. Stencil [CGU12, WTTH17, GÓ19, KD13, TBB12]. stencil-based [TBB12]. step [KW20, Kos95b, ZG98, vdP17]. steps [KW20]. Stereo [ZBd12, Qu95]. Steve [Ano96a, Ano99a, Ano99b, Nag05]. Steven [Ano96a, Ano99a, Ano99c, Ano99b, Ano99d, Nag05]. Still [HCA16]. Stochastic [AEW^{+}20, DK02, LLRS02, MW98, PTMF18, RSV^{+}05, JK10, MW21]. Stockholm [Eng00, HAM95b]. Stokes [Che99, DLR94, HSMW94, IDD94, Lou95, PTT94, SCC95, ZZG^{+}14]. stop [Gua16, LMG17]. stop-and-restart [LMG17]. Storage [ACM04, FLS20, Hol12, LCK11, HP11, NFG^{+}10, RGGP^{+}18, ZJDW18]. stores [HSP^{+}13]. straight [YULMTS^{+}17]. Strategies [CF19, MM02, BVML12, CG99a, DBVF01, HLK^{+}20, MM03, OPW^{+}12, PSK08, SIC^{+}19, TSZC94, VB99]. Strategy [AIM97, DI02, Hat98, VPS17, ZB94, ZSG12, DKF94b, DR95, LZZ^{+}20, MSL12, PSV19]. Stratix [LLVM21b, LLVM21a]. strayed [Rol08a]. stream [HSW^{+}12, LGMdRA^{+}19, UGT09]. streamer [LZC^{+}20]. Streaming [IADB19]. Streamline [CGC^{+}11]. streams [TVCB18]. StreamScan [YLZ13]. Strength [Kon00]. String [KMM15, MM02, MM03]. striped [KDSO12]. Strongly [GAP97, ZZG^{+}14]. Structural [PSSS01]. Structure [CBL10, LAFA15, SYF96, WHDB05, ZJHS20, EPML99, SEC15, SY95, ZAT^{+}07]. Structured [FB96, HDZ^{+}20, Mar06, MRB17, NLRH07, Ran05, AMKM20, Bis04, CLSP07, FR95, GBR15, JAT97, Smi93b, XJR21]. Structures [DK20, GMPD98, JY95, KA95, OKW95, SHPT00, WB96, YPA94]. studies [DHP97]. Study [AIM97, AFGR18, BF01, BHLS^{+}95, DARG13, DJJ^{+}19, EGC02, FPY08, Ger18, GL97a, HHC^{+}18, KCR^{+}17, LSB15, MYL21, MMD98, MM02, NSLV16, NA01, PK05, RRBL01, SCL01, TG94, AGR^{+}95b, AML^{+}99, BJ13, BfDA94, BJS99, BY12, Bri00, CBM^{+}08, DXB96, ED94, FO94, GADM20, HIIG16, IPG^{+}18, JR13, JLG05, KBG16, LPD^{+}11, LLH^{+}14, MS96b, NS20, PSK08, PGK^{+}10, PSHL11, RSBT95, RJC95, RRJ^{+}20, TPD15, Wal01b, WLK^{+}18, ZSK15]. Stuttgart [KGRD10, WPH94]. style [JPOJ12]. sub [MJG^{+}12]. sub-communicators [MJG^{+}12]. subcircuit [HLO^{+}16]. subdomain [CEGS07]. subdomains [SHHC18]. subgroup [XLW^{+}09]. Submitting [NSS12]. Subrange [Str97]. Subroutine [Saa94]. subroutines [dCH93]. Subsequence [Per21]. Subset [CWL^{+}20]. subsurface [ED94]. Subsystem [CVPS19, BMG07, MABG96]. Subsystems [STMK97]. Subtle [SAL^{+}17]. Success [Gro01b, LF^{+}93a]. Successes [Gro01a]. Successful [Gro12]. suffix [DK13]. SUGAR [YNJS21]. Suitability [Mat01b]. suitable [MAS06]. Suite [ACMR14, AKE00, BWV^{+}12, MBB^{+}12, Riz17, Ano03, BO01, MKP22, MvWL^{+}10, TG09, YSWY14, SNMP10]. Suites [MCS00, SGJ^{+}03]. summary [FL21]. summation [IHM05]. Summit [BC19b]. Sums [ST17, MYB16]. SUN [BM00, SJ02, WSN99]. Sunderam [Ano95b, NMC95]. Super [Gua16, YX95]. Super-Object [YX95]. Supercomputer [Ano93a, CLP^{+}99, Str94, AAC^{+}05, BGH^{+}05, EFR^{+}05, GL96, GL97c, KMH^{+}14, NSM12, Ste94, GS91b, MAB05]. Supercomputers [BP93, BDG^{+}92c, EKTB99, KN17, WT11, WT13]. Supercomputing [ACM96b, ACM04, ACM05, BDG^{+}91b, GGZ^{+}20, HK93, IEE91, IEE93e, IEE94h, RV00, Liu95, Sch94, ACM94, ACM96c, Ano93g, BG91]. superlattice [Pri14]. Supernode [CS19]. superscalar [ACJ12]. Supersonic [CCBPGA15]. Support [Ano98, BBG^{+}10, BFBW01, CFF^{+}94, DMMV97, FGRD01, GRV01, GOM^{+}01, HRSA97, LMRG14, MK04, OP98, PSM^{+}14, RR02, SDN99, SBT04, TW01, Wis98, Wis01, YSP^{+}05, ZL18, ADK22, BBH… 13a, BPJ22, BL99, CC10, CZ95b, DLR94, Hos12, Maf94, RS19, RJH^{+}20, TSY99, TSY00, TY14, WK08a, WK08b, WK08c, YÁJG^{+}15]. Supported [KLR16, ZGNZ22, CDD^{+}96, RJH^{+}20]. Supporting [FD00, FMSG17, FSG19b, GAML01, Gua16, MMS07, OOS^{+}08, SGL^{+}20, WLNL03, WLNL06, WCS99, YWCF15, FLD96, GAM^{+}00]. Supports [AELGE16, CLL03, DGMS93]. suppression [WWZ^{+}96]. Surface [KS15b, PKYW95, Rót19, BHW^{+}12, DCD^{+}14, RAGJ95, TSP95]. surfaces [Dab19]. Survey [Sap97, ZGNZ22, BBB^{+}20, HJB^{+}21]. Survive [ABB^{+}10]. sustainable [CGBS^{+}15]. SVD [CMH99]. Swan [HD11]. Swapping [SC04, BBW19]. Sweden [Eng00, HAM95b, FF95]. Swendsen [KO14, Kom15]. Switch [SCL01, TBD96, KSC^{+}19]. Switched [LC93, KYL03, KYL05]. switches [HGX^{+}22, DT17]. Switzerland [GT94, Ano94i, IEE97b]. SX [HRZ97, TSEE21, TRH00]. SX-4 [HRZ97]. SX-5 [TRH00]. SX-Aurora [TSEE21]. Sydney [Bil95]. Sylvester [GK10]. Sylvester-Type [GK10]. Symbolic [CCK12, Coo95b, Ste00, YYW^{+}12, ACM97a, BHKR95, Coo95a, Lev95, LGKQ10, LLG12, SMAC08]. Symmetric [BDV03, MDM17, YKW^{+}18, BAV08, DCH02, GG99]. Symposium [ACM95b, ACM96a, Ano94a, Ano95d, BG91, DE91, HHK94, IEE93c, IEE93b, IEE94a, IEE94e, IEE94g, IEE95c, IEE95d, IEE95k, IEE95f, IEE95g, IEE96b, IEE96c, IEE96f, IEE96e, IEE97b, IEE97c, IEE05, LHHM96, Li96, NM95, Ost94, SL94a, Sie94, Sie92a, Sie92b, Ten95, Tou96, USE94, UCW95, ACM97a, ACM06a, Ano93a, Ano94h, Lev95, Old02]. synchronisation [SDB^{+}16]. Synchronization [LA02, OCY^{+}15, TGT05, BMG07, LA06, SPNB14, TMTP96, YLZ13]. Synchronizing [VT97]. Synchronous [Ada97, BJ13, Cer99, CLE^{+}20, DLRR99, HZG08, SRS^{+}19]. Synergia [SSAS12]. Synergistic [UGT09]. Synthesis [CS14, GWC95]. synthesized [MC17]. Synthesizer [DS16]. Synthesizing [AJF16, LK20, NP12]. Synthetic [CC17, DP94]. Syracuse [IEE96f]. SYSMO [MM95]. System [Ada97, AJ97, AH00, BG95, BDG^{+}xx, BL95, BFZ97, BGD12, CAM12, CGC^{+}02, DBA97, DALD18, ERS95, ERS96, EK97, FBD01a, FBVD02, FFP03, Fis01, Gal97, GCBM97, GS91b, GS92, GSxx, GM95, Gre95, HS94, IADB19, KBA02, LLRS02, LTR00, LLY93, Maf94, MRV00, MM02, MSF00, MMH98, MMS07, MMH93, NPP^{+}00d, NMS^{+}14, Oed93, PPT96a, RGD97, SGJ^{+}03, SSB^{+}05, SCP97, SA93, ST02b, Sun93, TSS00b, Tsu07, UP01, Wil93, YSS^{+}19, ARS89, ADK22, AS92, AL92, BB94, Bri95, BBH^{+}15, DL10, DPFT19, DH22, FNSW99, FK94, GS91a, GS93, GS96, GMU95, GkLyCY97, HDDG09, Hum95, HS95b, IBC^{+}10, ITT99, JH97, JLS^{+}14, KW14, Kik93, LBD^{+}96, LKL96, LL95, MA09, MMR99, MMB^{+}94, MAS06, MM11, MS99b, MALM95, MMAH20, NAJ99, PPT96b, PPT96c, PK05]. system [RJDH14, RTL99, SHHI01, SL94b, Sei99, SPL99, SGDM94, Sun96, Sur95b, VSRC94, VSRC95, WCC^{+}07, WZWS08, YPZC95, YZPC95, ZL96, ZPLS96, ZWZ^{+}95, dCZG06, AL93, NMW93, Yan94]. System-Initiated [SSB^{+}05]. system-on-a-chip [dCZG06]. System/6000 [AL93, NMW93]. Systeme [GBR97, GEW98]. Systems [AAB^{+}17, Ano94b, Att96, BCGL97, BGBP01, BME02, BPG94, Bha93, CDJ95, CAWL17, COE20, CFF^{+}94, CSW97, CJNW95, Coo95b, DAD19, EADT19, FD96, FGKT97, Fos98, GGZ^{+}20, Gua16, HC17, HRSA97, IEE93d, IEE94d, IEE95a, IEE95i, IEE96h, KKH03, KP96, KQT^{+}21, KDL^{+}95b, KCR^{+}17, KS97, LY93, LBB^{+}21, LW97, MWG97, MBE03, MJB15, MBB^{+}12, SM03, SGS10, SGS^{+}21, SS96, TMP16, TWLL19, THN00, TL19, USE94, WJG^{+}21, YGH^{+}14, YH96, ZTD19, ZB97, dGJM94, AGR^{+}95b, ACMZR11, ATL^{+}12, Ano94e, BBB^{+}94, BAC20, BAV08, CdOO^{+}20, CKO^{+}94, CLYC16, CBPP02, Coo95a, CPR^{+}95, DF17, DR94, DBVF01, DvdLVS94, FHB^{+}13, GBR97, GCN^{+}10, GDEBC20, GEW98, GKK09, GKCF13, Gra09, GFPG12, GHH^{+}93, HHA95, IPG^{+}18, IM95, JB96, JJM^{+}11, KSG13, KHB^{+}99, KLV15, KDL^{+}95a, KFSS94, LBG^{+}20, LR06b, LH98, LRLG19, LCVD94b]. systems [LGM^{+}20, LLH^{+}14, MSL12, MvWL^{+}10, Old02, OPW^{+}12, Pan95b, Par93, PSB^{+}19, PGPCK21, QB12, RPS19, SSKF95, SCJH19, SPH95, SVC^{+}11, Smi93b, SG14, SMSW06, SLN^{+}12, Sun94b, TBB12, TMW17, TVCB18, TSP95, VLMPS^{+}18, WCS^{+}13, WWZ^{+}96, WADC99, WYLC12, ZL96, ZGC94, dH94, dlAMC11, dlAMCFN12, JWB96]. Systemsoftware [Sei99]. systolic [BSC99].

T3D [AZ95, AFST95, CCSM97, HWW97, MP95, MWO95, Oed93, Sch96a, Sch96b, SCC95]. T3E [BBS99, Boo01, Che99, GRRM99, LSK04, RBB97c]. T3E-512 [RBB97c]. T3E-600 [LSK04]. T9000 [BR94]. table [BJ13]. tablets [MNYN21]. Tabu [BSH15, Cza13, CB11]. Tags [Wis97]. TaihuLight [LHZ^{+}20]. Tails [Kha13]. takes [GDB^{+}93]. Talbot [ACMR14, Riz17]. tandem [GDMME22]. Tapir [SML17, SML19]. Targeting [BC19b, ABB20, JKM^{+}17, RVKP18]. Task [AHD12, AAB^{+}17, FKKC96, GDDM17, GPC^{+}17, GFJT19, IOK00, KOI01, KSB^{+}20, LHCT96, Mar03, MJB15, NIO^{+}02, NIO^{+}03, NSZS13, NJ01, OP10, OS97, SGZ00, SPL^{+}12, SGS^{+}21, TBS12, TS12a, WJG^{+}21, YKW^{+}18, APBcF16, ABF^{+}17, BLVB18, BGH^{+}05, GKCF13, OdSSP12, OPW^{+}12, OPP00, RRFH96, RFRH96, STP^{+}19, SWCB20, SKB^{+}14, WC15, WDR^{+}19]. Task-Based [AHD12, AAB^{+}17, GFJT19, SPL^{+}12, BLVB18, STP^{+}19, SKB^{+}14]. task-level [WDR^{+}19]. Task-Overlapped [GPC^{+}17]. Task-Parallel [KSB^{+}20, NSZS13, APBcF16, ABF^{+}17]. Taskers [FLD96]. Tasking [DFA^{+}09, KaM10, SHM^{+}10, TCM18, TSCaM12, VLSPL19, WC15, vdP17]. tasklet [PQR18]. Tasks [ACD^{+}09, DDP^{+}19, DT17, DFA^{+}09, JW96, OP98, PWPD19, RR02, RDLQ12, SGL^{+}20, WJG^{+}21, YSS^{+}17, YSS^{+}19, BS01, DDYM99, DR95, EBB^{+}20, FKK^{+}96b, FKK96a, IvdLH^{+}00, PKE^{+}10, PWPD19]. TAU [MMS07, RMS^{+}18]. taxonomy [SPH96]. TBB [Stp18]. TBSCM [BP98]. TC2 [Boi97]. TC2/WG2.5 [Boi97]. TCGMSG [GB96, Mat94, Mat95]. TCP [KPW05]. TD [And98]. Teaching [MK00, JY95, MK97, PKB06]. Technical [Ano93c, Ano98, MC94, USE95, ACM06a, Sni18]. Technique [BCD^{+}15, HC06, HAA^{+}11, MK17, HC08, Nes10, RBB17, MAIVAH14]. Techniques [CP97, GS02, Mül01, SAL^{+}17, SPL^{+}12, TGBS05, Wis01, AMKM20, BPG94, Fer04, FCS^{+}12, GSM^{+}00, HKMCS94, JKN^{+}13, KBG^{+}09, NFG^{+}10, PF05, SKS01, WST95]. technologies [Mal95]. Technology [Ano97, Bra97, CGB^{+}10, CSV12, Dan12, GN95, HS94, PWP^{+}16, SBT04, TBG^{+}02, Ano93a, Ano93c, D^{+}95, DM12, IEE94c, NS16, ZAT^{+}07]. Tekniska [Eng00]. Telegraphic [ES11]. TELMAT [BR94]. temperature [Hin11, RS22]. temperature-dependent [RS22].

Template [GSI97, PKB06]. Templates [BN12, KH15]. Tennessee [PR94b]. Tensor [BKK20, ZLWW20]. terabyte [KTJT03]. Terabytes [IEE02]. teraflops [KTJT03]. Terms [KD12]. Tesla [MYL21]. Tessellation [SS09]. Test [GSYT21, SNMP10, TG09, AAAA16, CPKG17, CPR^{+}95, GL92, TGKL19]. test-input [CPKG17]. Testbed [Mat01b, EGH99, PY95]. Testing [CDT05, CCK12, DKF94b, DLLZ19, DLLZ20, Ost94, VdS00, CMV^{+}94, DKF93, KSTM20]. Testsuite [WCC12]. Texas [ACM06a, IEE94b, IEE95l, IEE95g, IEE97c, Y^{+}93]. Text [LTR00, MM01, RLL01, RTL99]. Textbook [Ano98]. textural [WKS96]. texture [HE15]. TFETI [SHHC18]. TH [CFDL01]. TH-MPI [CFDL01]. Thakur [Ano00a]. Their [Brü12, GOM^{+}01, RG18, GSMK17]. theorem [Sut96]. Theory [GK10, BW12, CBHH94]. Thera [CD01]. thermostat [RS22]. Think [HCA16]. Third [BPG94, Bos96, DSM94, GA96, IEE94g, Sil96, Was96, BDLS96, Mal95, IEE97c]. Thirty [Y^{+}93]. Thirty-seventh [Y^{+}93]. Thousands [PZKK02, BMS^{+}17]. Thread [AELGE16, BB18, ETWaM12, GOM^{+}01, GT07, LML^{+}19, Nit00, Pla02, STY99, SPB^{+}17, AKB^{+}19, HK09, IDS16, JKN^{+}13, LW20, SPH96, SLN^{+}12, YZ14]. thread-based [AKB^{+}19]. thread-data [LW20]. Thread-Level [AELGE16, HK09, YZ14]. Thread-Safe [Pla02]. Thread-safety [GT07]. Threaded [BBG^{+}10, MG15, WZM17, Ada98, EBKG01, SCB15, SVC^{+}11, TSY99, TSY00]. threaded-MPI [SVC^{+}11]. Threading [BHV12, MLGW18, SBT04, TBG^{+}02, WMK^{+}19, KPO00, KRG13, QB12, ZAT^{+}07]. Threads [CP98, LD01, Lee06, SdR^{+}21, BS01, DJJ^{+}19, MVTP96, ALW^{+}15]. Three [Car07, GA96, lLLmH^{+}21, Nak05b, Ram07, SAS01, ZWLZ21, GSMK17, LSSZ15, LZC^{+}20, Mar05, PR94c, ZWC21]. three- [GSMK17]. Three-Dimensional [GA96, ZWLZ21, lLLmH^{+}21, LSSZ15, PR94c, ZWC21]. Three-level [Nak05b]. Throughput [HMKG19, SSLMW10, Tsu07, CJPC19, ESB13, PP16]. throughput-oriented [CJPC19]. Thrust [DSU20]. Tied [WJG^{+}21]. Tightly [SS01]. Tightly-Coupled [SS01]. Tilewise [KS15b]. tiling [KW20]. Time [BCL00, CBB^{+}20, CBB^{+}21, DLLZ19, DLLZ20, FHK01, FSSD17, GSHL02, GOM^{+}01, HO14, KFL05, MFTB95, OP98, SPB^{+}17, SGL^{+}20, SCL01, SS96, TWLL19, TSP95, UP01, YGH^{+}14, AL96, ASB18, CDMS15, DLR94, DS22, DPFT19, DM12, Fer04, FLB^{+}05, FKLB08, GB94, HE13, HFB21, JE95, KC94, KPL^{+}12, KSC^{+}19, KW20, LHLK10, LBB^{+}16, LYSS^{+}16, LM13, MMW96, NZZ94, ÓN12, OdSSP12, PTMF18, PRQ21, QHCC17, Ram07, SBW91, SSB^{+}16, SM19, SK92, SRK^{+}12, TSY99, Tho94, TVV96, TCBV10, Uhl95c, VM94, YSVM^{+}16, YSMA^{+}17, ZWZ^{+}95, SKD^{+}04]. time-critical [KSC^{+}19]. time-dependent [DM12, LBB^{+}16, LYSS^{+}16, ÓN12, SSB^{+}16, YSVM^{+}16, YSMA^{+}17]. time-domain [HE13, NZZ94, Ram07, VM94]. time-explicit [DS22]. time-independent [CDMS15]. time-stamping [DPFT19]. Time-Varying [DLLZ19, DLLZ20, Uhl95c]. times [MLVS16, NB96, SWCB20, SSS99]. timing [Ols95]. tips [Fer04]. TLM [SC96a]. TM [GGCM99, GCGS98, KHS01]. TN [DT94, BR94]. TOD [GPC^{+}17]. TOD-Tree [GPC^{+}17]. today [IEE94c]. Toeplitz [BV99, BAV08]. Tolerance [GKP97, GL04, LMRG14, LNLE00, RPM^{+}08, TS12a, WC09, Wil93, CLE^{+}20, LRG^{+}16, LGM^{+}20, SG05, WDR^{+}19, ZHK06]. Tolerant [BBC^{+}02, BCH^{+}03, BHK^{+}06, CF01, CFDL01, FD00, FBD01a, FBVD02, FD02a, FD04, GFB^{+}03, GGZ^{+}20, IEE95c, JSH^{+}05, MSF00, BCH^{+}08, FBD01b, FD02b, HG12, LMG17, LS08, MB21, NCB^{+}12, NCB^{+}17, PKD95]. Tomographic [Pat93]. tomography [FWS^{+}17, RCFS96]. tomorrow [IEE94c]. Tool [Ano01b, Beg93b, BFMT96b, DW02, GSN^{+}01, KAMAMA17, KSJ14, KKP01, LMRG14, MMSW02, MK04, NE98, SR96, SGL^{+}00, Trä12b, VBB18, WL96a, AGG^{+}95, BDP^{+}10, Beg92, Beg93c, Beg93a, BDY99, BFMT96a, BHW^{+}12, CPR^{+}95, DKF94a, FSTG99, GDMME22, HPR^{+}95, HD11, IMS16, JKN22, LSB^{+}20, LCC^{+}03, MdSAS^{+}18, RMS^{+}18, TSS98, WL96b, WL96b]. Tool-Set [WL96a]. Toolbox [Ano97, Bra97]. Toolkit [Ano12, KTXP21, LC07, LLC13, SLS96, PSH^{+}20]. Tools [ABC^{+}00, BDG^{+}91b, BDG^{+}93a, BS96a, BDL98, BoFBW00, Cha05, CDD^{+}96, DT94, EV01, GMPD98, MHC94b, MCLD01, PKB01, STMK97, Vos03, Wan97, AMKM20, AVA^{+}16, BDG^{+}92a, BFIM99, Fan98, GBF95, LH98, MSW^{+}05, MHC94a, ZL96]. Tools-supported [CDD^{+}96]. Top [AHP01, Gal97, Hus01, Man01, PTH^{+}01b, Ser97, BBCR99, PTH^{+}01a, SSC96, SCL97, CCHW03]. TOP-C [CCHW03]. ToPe [JKM^{+}17]. topologies [BCM^{+}16, Gro19, MK00]. Topology [DK06, Hat98, HM01, Tra02a, GJMM18, HRR^{+}11, MBBD13, SPK^{+}12]. topology-aware [MBBD13]. Topology-Based [HM01]. TOPPER [KKP01]. Toronto [GGK^{+}93, Vos03]. Torus [DDP^{+}19, SG15]. Townsend [DT94]. TPVM [FS95, FS98]. Trace [Ney00, FLPG18]. trace-based [FLPG18]. Traceback [dOSMM^{+}16]. Tracefiles [FCP^{+}01]. Traces [CC17, MANR09, WM01, CDMS15, DWM12]. Tracing [CGLD01, DP94, KG96, CG93, Mor95, SGS95]. Tracking [GAP97, HD02b]. tradeoff [RPS19]. Trading [BHM94, BHM96]. Traffic [VT19, Zah12]. Training [CSV12, RJ21]. Transactional [BWW^{+}12, MFG^{+}08, SBG^{+}12]. Transactions [BWW^{+}12, SSD^{+}20]. Transcoding [LSB^{+}18]. Transfer [BKGS02]. Transfers [THS^{+}15]. Transform [YULMTS^{+}17, KT10, DBLG11]. Transformation [CLA^{+}19, EP96, NSZS13, CS19, GSMK17, HZ96, TSY00, YW21]. transformations [JE95, TG94]. transformed [BY12]. Transforming [PSK^{+}10]. Transforms [ACMR14, KLR16, HP11, Uhl95c, Zem94]. Transient [SIS17]. transistor [Ano03]. transistors [Ano03]. Transition [MRV00]. Transitive [CGPR98, PPR01]. Translating [Mar09, NCB^{+}12]. Translation [DDL00, SSE12, HCL05, LME09, NCB^{+}17, WZW21]. Translator [KMK16, UZC^{+}12, CHKK15, GScFM13]. transmitters [WWZ^{+}96]. Transparent [CCK^{+}95, IFA^{+}16, NPP^{+}00c, RVKP19, SLGZ99, LFS93a, LFS93b, LFL11, NPP^{+}00a, SOA11]. Transparently [CB16]. Transport [KHS01, MMD98, RS97, VRS00, WR01, ZZ04, Pri14, SH94, SCJH19, WH96]. Transporter [Fer92]. transpose [Bha98]. Transposition [HD02b]. Transputer [Ara95, ACDR94, CJNW95, FK95, FF95, GN95, GHH^{+}93, MC94, dGJM94, ZPLS96, Ara95, CJNW95, GHH^{+}93, dGJM94]. Transputers [ACDR94, AGR^{+}95b, dCH93]. Transtech [Ste94]. trap [LBB^{+}16, SSB^{+}16, YSVM^{+}16]. TRAPPER [KFSS94, SSKF95]. travel [SSS99]. travel-times [SSS99]. traveling [GM94]. traversing [BDG^{+}92b]. TreadMarks [LDCZ97]. Tree [DAD19, GPC^{+}17, MB21, ADB94, AB13, BCAD06, CG93, SGS95, Zah12]. Tree-based [MB21]. Trees [CDPM03, GFJT19]. Trends [Duv92, IEE93d, MBS15, JPTE94, SGDM94, Sun96]. Trial [DSU20]. Triangle [SL94a, SOA11]. Triangular [Hog13, MRB17]. triangulated [Dab19]. Triangulation [CWL^{+}20]. tricks [Fer04, LK14]. Tridiagonal [DALD18, DAD19, DR18, VLMPS^{+}18]. Triolet [RJDH14]. Trivandrum [IEE96a]. Troy [SS96]. Truncated [ZB97]. truncating [Ram07]. TSMC [Ano03]. TSUBAME [NSM12]. TSUBASA [TSEE21]. Tsukuba [SHM^{+}10]. tsunami [KNH^{+}18]. TTIG [RRBL01]. Tubal [ZLWW20, ZLWW20]. Tubal-Rank [ZLWW20]. Tucker [BKK20, OPJ^{+}19]. TuckerMPI [BKK20]. Tucson [JB96]. tuned [PSB^{+}19, VLCM^{+}20]. Tuning [Ben18, Cza02, Cza03, LWSB19, NPP^{+}00d, PSH^{+}20, SLJ^{+}14, SdR^{+}21, WG17, YT20, DBLG11, FE17a, FE17b, LGG16, SH14, Yan94, FVD00]. tuple [MYB16]. tuple-based [MYB16]. Turbulence [Str97, MRRP11, Str96]. turbulent [BCM^{+}16, CBYG18, NS20]. Tutorial [EM00a, EM00b, GBD^{+}94, GLT00b, Nov95, NMC95, Per96, Ano95b]. TV [CIJ^{+}10]. Twenty [ERS95, ERS96, HS94, IEE95c, MMH93]. Twenty-Eighth [ERS95]. Twenty-fifth [IEE95c]. Twenty-Ninth [ERS96]. Twenty-Seventh [HS94]. Twenty-Sixth [MMH93]. Two [CM98, STY99, SJK^{+}17a, SJK^{+}17b, YM97, AGR^{+}95b, AL93, ADLL03a, ADLL03b, CB11, ED94, HAJK01, LK20, MSP93, dlAMCFN12]. Two-Dimensional [SJK^{+}17a, SJK^{+}17b, AL93]. two-layer [dlAMCFN12]. Two-level [STY99]. two-phase [ED94]. TX [ACM00, Cha05, DKM^{+}92, Ano95a, Ano95d]. Type [GK10, MSB97, FVLS15, GFPG12]. Types [Wel94, NYNT12]. typy [OA17].

U.S. [LD01]. U.S.A [Ano94e]. Überblick [Wer95]. UK [Abr96, AD98, EJL92, HK95, BP93, CJNW95, MC94]. UKMO [RSBT95]. ULFM [LCMG17, LGM^{+}20]. Ultra [SJ02]. Ultra-High [SJ02]. Ultrafast [KRC17, FWS^{+}17]. ultrashort [MV20].

Ultrasonic [ASAK19, DLLZ19, DLLZ20]. Umgebung [GBR97]. UML [RGD13]. UML/MARTE [RGD13]. Umpire [VdS00]. Unbalanced [OP10]. Uncertainty [MBS15]. underlying [RS21]. Understand [DeP03]. Understanding [CRE01]. underwater [ZWC21]. unexpected [LFW20]. Unibus [KSSS07]. UNICOM [Ano93h]. Unified [KC19, GKZ12, JC17, KSL^{+}12, KLV15, STA20]. unifies [RJDH14]. uniform [KSG13]. uniformly [Trä12a]. Unify [VSRC94, VSRC95]. unifying [CCM12]. Unintended [SAL^{+}17]. unit [JPL22, VDL^{+}15, MSML10]. United [Boi97]. Units [KS15b, LSVMW08, ABDP15, BHS18, LHLK10, WWFT11, HJBB14]. Universal [LW97, DDLM95]. University [CGB^{+}10, IEE94d, IEE95j, R^{+}92]. Unix [OLG01, RBS94]. Unleashing [TCM18]. unscharfer [Wil94]. Unstructured [AB93a, NO02b, SM02, SM03, AB93b, NO02a, TPD15]. unsupervised [RTN21]. unveils [Ano03]. UPC [EGC02, MTK16, Mar05, SJK^{+}17a, SJK^{+}17b]. Update [DF21, KT10, GSMK17]. Updates [ESB13, KS15a, ZDR01, HSE^{+}17]. UPM [NPP^{+}00d]. ups [Ano03]. USA [ACM96b, ACM98b, ACM00, ACM06a, AGH^{+}95, BBG^{+}95, BS94, Cha05, CGKM11, DT94, EV01, EdS08, ERS96, Gat95, Ham95a, Hol12, IEE95b, IEE95d, IEE96f, IEE96e, IEE96i, MCdS^{+}08, Old02, PBG^{+}95, Ree96, RV00, Sin93, Ten95, ACM95b, ACM97b, Agr95a, Ano89, B^{+}05, DKM^{+}92, GT19, HS94, IEE94e, IEE95k, IEE02, Ost94, SL94a, SS96, USE94, USE95, USE00]. Usage [FD02a, FCLG07, BBB^{+}20, FD02b, FVLS15, FL21, PIR^{+}20]. Use [FJBB^{+}00, Gro02a, HK93, HK95, MB12, PSZÉ00, Shi94, AB95, GEW98]. used [JKN22]. USENIX [USE94, USE95]. User [AD98, ACDR94, BDG^{+}91a, CHD07, CD01, CDND11, DKD05, D^{+}91, DHHW92, DHHW93a, DLM99, DKP00, DLO03, FCLG07, GBD^{+}94, GN95, KGRD10, KCP^{+}94b, KOW97, Kra02, KKD04, LKD08, MC94, MTWD06, NPP^{+}00c, Nov95, NMC95, Per96, RWD09, TBD12, XF95, ZWZ05, Ano95b, BBB^{+}94, BDW97, KCP^{+}94a, LRG^{+}16, RSC^{+}15, Reu01, Wil94, BBH… 13a]. User-Level [DHHW92, DHHW93a, KCP^{+}94b, KOW97, NPP^{+}00c, XF95, ZWZ05, KCP^{+}94a, LRG^{+}16, BBH… 13a]. Users [Ara95, CHD09, KQT^{+}21, HJB^{+}21]. uses [SH96]. Using [AR01, ADRCT98, AHP01, And98, AP96, Ano95e, AKE00, AZG17, AB93a, BST^{+}13, BPMN97, BG95, BS93, BKGS02, BM97, Bon96, BBC^{+}00, BBH12, CGC^{+}11, CRE99, CMM03, CP97, CSPM^{+}96, CJvdP08, CC17, Che99, COE20, CCSM97, CDM93, CCHW03, CRGM14, CT94a, CCBPGA15, CD98, DeP03, DARG13, DDN^{+}22, DAK98, DGMJ93, DGH^{+}19, EM02, EMO^{+}93, ESM^{+}94, EK97, FAFD15, FD04, FDG19, FTVB00, FS93, GGCM99, GCGS98, GTH96, GM95, GK97, GS96, GSYT21, GMPD98, GHL97, GJN97, GLS94, GLT99, GLS99, GLT00b, GLT00a, Gro19, HB96b, HSMW94, HJ98, HLP11, HD00a, HT08, HRSA97, HT01, IOK00, IDD94, IKM^{+}01, JFGRF12, JPP95, KB98, KOI01, KKV01, KF16, KS96, KA13, LLRS02, LTR00, LRT07, LTRA02, LFS^{+}19, LY93, LLY93, LZ97, LSB^{+}18, LAFA15, MK17, MTSS94, MPD04, MR12]. Using [MSCW95, MANR09, MBB^{+}12, MSB97, NO02b, NIO^{+}02, NIO^{+}03, Neu94, NH95, NA01, OM96, OCY^{+}15, OWSA95, PWP^{+}16, PK98, PPT96c, POL99, PT01, Per99, Pet97, PBK00, PD98, PGF18, Pus95, QRMG96, QMGR00, RR00, Reu03, RRBL01, RLVRGP12, RLL01, RRG^{+}99, SAS01, Sev98, SSAS12, SP99, SA93, Smi93a, SBR95, STV97, SMOE93, SSB21, Sta95b, ST17, SKH96, SCL01, SJK^{+}17a, SJK^{+}17b, TS12a, TSB02, TSB03, TSN21, TK16, TBB12, Tha98, Tra98, Tsu07, VLO^{+}08, WO95, Wal01a, WTS19, WJ12, WLR05, Wis97, Wis01, WMC^{+}18, WLYC12, YKW^{+}18, ZWLZ21, ZCBD22, ZBd12, van97, vdLJR11, vdP17, AMHC11, ASAK19, AK99, ABF^{+}17, AL96, ADT14, ABG^{+}96, AB93b, AGIS94, AGG^{+}95, BV99, BBC^{+}19, BFLL99, BAE22, BSC99, BDG^{+}92c, Bic95, Bis04, BCM^{+}16, BTC^{+}17, BCD96, BID95, BAG17]. using [BSH15, BMG07, CJPC19, CPM^{+}18, CG93, CBM^{+}08, CBYG18, CdGM96, CS14, CLBS17, CT94b, CC00b, DG95, DMK19, DS13, DS22, DRUE12, DSOF11, DCH02, DM12, EGDK92, FB96, FSV14, FSC^{+}11, Fin94, Fin95, FHC^{+}95, FWS^{+}17, GGGC99, GSMK17, GG09, Goe02, GFB^{+}14, GMU95, GM18, GRTZ10, GADM20, HB96a, HDDG09, HTJ^{+}16, HP11, HPS^{+}96, HPLT99, HASnP00, Hol95, HLO^{+}16, HAA^{+}11, IJM^{+}05, IM95, IKM^{+}02, JL18, JKN22, JF95, JPL22, JKHK08, JLS^{+}14, JJY^{+}03, JJM^{+}11, JPT14, JR10, JMdVG^{+}17, KFA96, KRKS11, KY10, Kat93, KJJ^{+}16, KR09, KMK16, KME09, KMC96, KMC97, KRC17, KMM15, KD13, KPK13, LP00, LSG12, LSSZ15, LHZ^{+}20, LCY96, LSVMW08, LCMG17, LO96, MMR99, MP95, Mar06, MSMC15, MNYN21, MAB05, McK94, MKP22, MM11, Mic93, Mic95, MRH^{+}96, MMM13, MSML10]. using [MS95, MM14, MC99, MvWL^{+}10, NO02a, Nak05a, NZZ94, NB96, NAJ99, NU05, OKM12, OIH10, Ols95, OHG19, Pat93, PDY14, PGdCJ^{+}18, PSV19, PNV01, PKE^{+}10, QRG95, RJC95, RAS16, RCFS96, RBAI17, RM99, RCG95, SHLM14, SdM10, SLGZ99, SSN^{+}21, SGS95, SSS99, SMS00, SOA11, SVC^{+}11, SSGF00, SBB20, SOYHDD19, SFLD15, SSN94, SU96, SP11, Stp18, Stp20, TC94, TPLY18, Tsu95, Uhl94, Uhl95b, UH96, VM94, VB99, VGS14, VM95, WO96, Wal01b, WCS^{+}13, WCVR96, WST95, WMRR17, WRMR19, WADC99, Wor96, WYLC12, XF95, XJR21, XR21, YULMTS^{+}17, YWC11, YWCF15, YCA18, ZWHS95, ZSK15, ZAT^{+}07, ZZ95, Ano95c, Ano00a, Ano00b]. UT [Hol12]. UTE [JF95]. Utilising [SC96a]. Utilities [CC95]. UV2 [TW12]. UVM [NSLV16].

V [JB96, BBC^{+}02, BHK^{+}06]. V100 [MYL21]. V2 [BCH^{+}03]. VA [Sin93, RP95]. Vacancy [HD02b]. Vaidy [Ano95b, NMC95]. Validation [BDV03, GLB00, WCC12, CMV^{+}94, SCB14, SCB15]. Value [vHKS94, AL96, LSR95, OHG19, SP11, SD99]. Value-based [vHKS94]. valued [Str12]. VAMPIR [BHNW01, NAW^{+}96]. Vancouver [IEE95a, IEE95i]. Vapour [PKYW95]. Variable [Ano98, LK20, ZZG^{+}14]. Variables [FKH02]. variably [TOC18]. Various [LH95]. Varying [DLLZ19, DLLZ20, Uhl95c]. VASP [WMK^{+}19]. VCMON [Whi94]. vCUDA [SCSL12]. Vector [AKL16, DS13, Fuj08, KDT^{+}12, LL16, Uhl95c, ER12, FVLS15, FJZ^{+}14, GL96, GL97c, Har94, Har95, HE15, PMZM16, RJH^{+}20, XXL13, ZCBD22]. Vectorization [IKM^{+}01, MCP17, IKM^{+}02, Stp18]. Vectorized [KB13]. Vectors [DK20, AAAA16]. Vegas [Ano94e]. Vehicle [BHM94, BHM96, WH94, BKvH^{+}14]. Vendor [Rab98, Bor99]. Vendors [KQT^{+}21]. Venice [DLO03, OL05]. venture [Ano03]. Verification [BCD^{+}15, RAS16, Trä12b, LMM^{+}15, SZ11, VVD^{+}09]. verified [WBBD15]. verifier [BCD^{+}12, LGKQ10]. verify [MdSAS^{+}18, SMAC08]. Verilog [Kat93, KMK16]. Versatile [KSJ14]. Version [BCGL97, CCK^{+}95, MHSK16, Bjo95, BHW^{+}12, BBH^{+}15, DS22, Man94, RS22, Str94, Wal95, WRMR19]. versioned [SSB^{+}17]. Versions [Ano98]. Versus [RTRG^{+}07, Ahm97, CE00, KPW05, KAC02, KPO00, LMG17, LC97b, MFTB95, NSLV16, NHT02, NHT06, RS95, SZ99, Wal00, ZLZ^{+}11]. verteilter [GBR97]. VGRIDSG [AB93a]. VIA [Sei99, CTBT21, FKKC96, ADGA20, BKK20, BHW^{+}12, CWL^{+}20, CGZQ13, DS96b, FLPG18, GB96, Hos12, HCL05, LAdS^{+}15, LSSZ15, LCH^{+}22, NPP^{+}00c, QHCC17, RBC20, SLJ^{+}14, Sti94, VBLvdG08, YPZC95, ZJDW18, ZLL^{+}12, EM02, RR01]. VIA/SCI [RR01]. viable [Ano03]. Victoria [IEE95e]. Video [KSJ95, KSJ96, LSB^{+}18]. videogames [YMYI11]. Vienna [BH95, TBD12, Ben95]. View [ZDR01, ZDR04]. ViMPIOS [Sto98]. VinaMPI [ESB13]. ViPIOS [Sto98]. Virginia [IEE92, IEE94a, Sie92a, Sie92b]. VirtCL [YWTC15]. Virtual [ACM96a, AS92, ARL^{+}94, BJ93, BP99, BS93, BG94b, CHD07, D^{+}91, EGR15, Fis01, GBD^{+}94, Gei01, Gre94, ITT99, JPP95, KNT02, KKDV03, KKD04, KKD05, LKD08, LK10, MTWD06, NM95, Nov95, NMC95, Pat93, Per96, QRG95, RWD09, SSSS96, Sei99, SCSL12, SXMX^{+}18, TY14, Tsu07, Wel94, YC98, ARS89, AD98, AL92, Ano95b, BR91, BDG^{+}91a, BPC94, BBCR99, Bir94, BDLS96, BCM^{+}16, BFM96, BDW97, BB95b, CARB10, Cav93, Cha96, CD01, CXB^{+}12, DDS^{+}94, DM93, DKD05, DLM99, DKP00, DLO03, DPZ97, ESB13, FM90, Hol95, KMC97, KSS^{+}18, Kra02, LG93, MN91, MRH^{+}96, NB96, PRS16, Sch94, SK92, SCC96, SL00, WK08a, WK08b, WK08c, AGIS94, Sei99]. virtual-time [SK92]. Virtualization [FC05, LYGG20, MGL^{+}17, Ott94, YSS^{+}17, ZLP17, CPM^{+}18, HIIG16, NRdA^{+}20, RSC^{+}15, SIRP17]. Virtualized [EGR15, YWCF15, RNPM13]. virtue [LFW20]. viruses [Str94]. viscoelastic [HK94, MAIVAH14]. viscosity [ZZG^{+}14]. viscous [RM99]. Vision [KCR^{+}17, JRM^{+}94]. VISPAT [HPS95]. Visual [BPMN97, FNSW99, PDY14, Ros13, ACGdT02, LC07, GE95, GE96]. Visualization [BDGS93, GKP96, GKP97, HJ98, KA13, MVY95, NAW^{+}96, PK98, PCY14, Wis96a, ZLGS99, Bor99, Eng00, FHC^{+}95, HPS95, KFA96, TSS98, WST95, Wis96b]. Visualizer [HKN^{+}01]. VLSI [Jes93a]. VM [GHD12, McR92, Whi94]. VM-protected [GHD12]. VM/ESA [Whi94]. VMPP [LG93]. VOBLA [BKvH^{+}14]. Vol [ATC94, HS94, Nag05]. Volatile [BBC^{+}02, BCH^{+}03]. Volta [KC19]. Voltage [KFL05, FKLB08]. Volume [Ano99a, Ano99c, Ano99b, Ano99d, DLLZ19, DLLZ20, DFN12, GHLL^{+}98, KLH^{+}20, SOHL^{+}98, BHW^{+}12, DS22, WST95]. Volumes [GAP97, SOA11]. Volumetric [KA13, CLBS17, KGB^{+}09]. Voodoo [PMZM16]. VOOM [BR91]. VORD [KSJ14]. VR [DBA97]. VRML [ACM96a, NM95, KSJ95, KSJ96]. VRML-Based [KSJ95, KSJ96]. vs [FH98, AFGR18, BCH^{+}08, IPG^{+}18, Luo99, Nak05b, SC19]. VTC [NU05]. VTDIRECT95 [HWS09, SWH15]. vulnerabilities [LCH^{+}22]. VxWorks [YGH^{+}14].

WA [ACM05, LCK11]. Wailea [ERS96, HS94, MMH93]. wait [SWCB20]. Waknaghat [CGB^{+}10]. walk [RJH^{+}20]. Walker [Ano96a, Ano99a, Ano99b, Nag05]. walks [MW21]. wall [NB96]. wall-clock [NB96]. walls [JAT97]. WAMM [BCLN97]. Wang [KO14, Kom15]. Warehousing [DERC01]. Warp [MPZ21, SCL01, HKOO11, MMW96, VSW^{+}13]. WARPED [MMW96]. WARPmemory [SFO95]. Washington [B^{+}05, BS94, IEE93c, IEE94h, IEE95k, Ost94]. watching [JLG05]. water [DS22, HTHD99, R^{+}92, STA20, dlAMC11, dlAMCFN12]. Waterman [KDSO12, RGB^{+}18]. watershed [NAJ99]. Wave [BBC^{+}00, EMO^{+}93, ESM^{+}94, NSLV16, SMOE93, Gei94, KM10, KEGM10, Mal01, NS20, NB96, RMNM^{+}12].

Wave-Particle [NSLV16]. Waveform [LSR95]. Wavelet [Uhl94, Uhl95b, Zem94, vdLJR11, Uhl95a, Uhl95c]. Way [Vog13, HLK^{+}20, WDR^{+}19, FGT96]. ways [CZ96]. WCRT [SGS^{+}21]. weak [SD16]. Weather [AHP01, HE02, Bjo95, KOS^{+}95a, Mal01]. web [CHKK15, AASB08, NE01, PES99, Wal01b]. Web-Based [NE01, PES99]. WebCL [CHKK15]. WebCom [OPM06]. WebCom-G [OPM06]. wedge [ZWC21]. wedge-shaped [ZWC21]. Wednesday [B^{+}05]. Weicheng [Ano95b, NMC95]. weight [KA95]. Weighted [RJ21]. Weighted-Averaging [RJ21]. welcomes [Str94]. West [EV01, EdS08]. Westin [IEE94e]. We’ve [GKPS97]. WG10.3 [DR94]. WG2.5 [Boi97]. Wheeler [NTR16]. where [KC94]. which [SH96]. Whippletree [SKB^{+}14]. whistler [NS20]. Wide [FGG^{+}98, MPZ21, dOSMM^{+}16, FGT96, KHB^{+}99]. Wide-area [FGG^{+}98, FGT96]. Wide-Warp [MPZ21]. WIEN [Gao03]. Will [CB00]. William [Ano95c, Ano99c, Ano99d, Ano00a, Ano00b]. Williamsburg [IEE92]. Win32 [MS98]. windows [QB12, QM21, RGGP^{+}18, Ano01a, CLP^{+}99, FD97, GGGC99, PS01a, SFG98, SSSS97, TAH^{+}01]. Windows95 [SSSS96]. Winona [Ano94h]. wireless [Bon96]. wissenschaftliche [MS04]. wissenschaftliches [Ano94c]. within [WDR^{+}19]. without [BW12, Pla02, RSC^{+}19, YLZ13]. WLAN [MSOGR01]. WMPI [BPS01, MS98, MSS98, MS99c, PS01a, SMS00]. WOMPAT [Cha05, EV01, Vos03]. Woollongong [GN95]. Work [HRSA97, Pet00a, Pet00b, WQKH20, OdSSP12, TCBV10]. Work-Group [WQKH20]. work-stealing [TCBV10]. Worker [EML00, YG96]. Worker-Based [YG96]. Workerproblem [FH98]. Workflow [LYZ13]. workflows [WDR^{+}19]. Workforce [Liv00]. Workgroup [YT20, SDB^{+}16]. Working [Ano98, Boi97, MCS00, Pet01, DR94]. Workload [AGS97, DBVF01, PS19a]. Workloads [AJC^{+}20, AFGR18, CC17, LWZ18, APBcF16, AVA^{+}16, AMC^{+}19, CJPC19, GDS^{+}20, JCP^{+}20, SKB^{+}14]. WorkPlace [Ano97, Bra97]. workqueuing [VBLvdG08]. Workshop [ACM98a, Agr95a, BPG94, Bha93, BC00, Cha05, CZG^{+}08, CGKM11, CMMR12, DW94, DT94, EV01, EdS08, Fer92, FK95, FF95, HK93, HK95, IEE93d, IEE93f, IEE94d, IEE95h, IEE96g, IFI95, KG93, Kuh98, Kum94, MdSC09, PBG^{+}95, PBPT95, SCR92, SHM^{+}10, Sch93, Vos03, Was96, AH95, BS94, Cal94, D^{+}95, DMW96, FR95, GL95b, IEE93f, SBG20]. Workshops [MCdS^{+}08]. Workstation [GHL97, HSMW94, KS96, LC97a, MFTB95, Pus95, YKI^{+}96, AB95, ALR94, BLP93, BSvdG91, BRS92, BALU95, BWT96, CCU95, DG95, ED94, GBF95, Heb93, JRM^{+}94, LL95, NMW93, NN95, PM95, PL96, RBS94, RCFS96, SC96a, SSN94, SL95, THM^{+}94, Tsu95, UH96, YWO95, ZHS99, MS04]. workstation-cluster [Heb93]. Workstation-Clustern [MS04]. Workstations [AR01, BL94, BL95, BM97, BDH^{+}95, BDH^{+}97, BMS94b, DDPR97, EK97, GS91b, HIP02, IDD94, Liu95, LHZ98, MSCW95, MM01, OWSA95, PFG97, TQDL01, VLO^{+}08, AL93, BJ95, BID95, Bru95, BMPZ94b, BMS94a, BMPZ94a, CCF^{+}94, Coe94, DZ98a, DOSW96, GM94, GMU95, HK94, Hus99, KMC96, KMC97, KA95, MK94, MM03, RRG^{+}99, SFO95, SR95, TDB00, dCH93]. World [CMMR12, CJNW95, FD00, GHH^{+}93, HLP11, MC94, NSLV16, PSB^{+}94, Wit16, dGJM94, GDB^{+}93, JR10]. Worlds [Rab98]. wormhole [Pan95a, Pan95b, RJMC93, ZGN94]. wormhole-routed [Pan95b, RJMC93, ZGN94]. worms [Pan95a]. WoTUG [MC94]. WoTUG-17 [MC94]. WPVM [ASCS95, BPMN97]. Wrapper [AS14]. Wrapping [LRW01]. Write [BIC^{+}10]. Write-Back [BIC^{+}10]. Writing [FAF16, SDB94, FNSW99]. Written [KaM10, KNH^{+}18]. WWW [KSJ95, KSJ96].

X [Bad16, FWS^{+}17, HCC^{+}20, MMAH20]. X-ray [FWS^{+}17]. X10 [CGH^{+}14]. X11 [GKL95]. x86 [MGL^{+}17]. Xab [Beg92, Beg93b, Beg93c, Beg93a]. Xen [PRS16]. Xeon [CBIGL19, DSGS17, MMDA19, OTK15, BB18, MTK16]. Xilinx [MKP22]. XPVM [KG96]. XXI [EGH^{+}14].

Years [DF21]. YLC [Gal97]. YMP [BL94]. Yorkshire [CJNW95].

Zero [HCC^{+}20, SWHP05, Hin11]. Zero-Copy [SWHP05, HCC^{+}20]. ZEUS [FF95]. Zipcode [wL94, SSD^{+}94]. zonal [Fin94, Fin95]. Zone [JCH^{+}08, AGMJ06]. zum [Wer95]. zur [GBR97, Sei99].

References

AlQuraishi:2016:CBP

[AAAA16] Eman AlQuraishi, Eman AlDwaisan, Alaa AlSaqaa, and Imtiaz Ahmad. A CUDA-based parallel implementation of a test vectors encoding algorithm in compression-based scan designs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(3):280– 293, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Andion:2016:LAA

[AAB^{+}16] José M. Andión, Manuel Arenaz, François Bodin, Gabriel Rodríguez, and Juan Touriño. Locality-aware automatic parallelization for GPGPU with OpenHMPP directives. International Journal of Parallel Programming, 44(3):620–643, June 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10766-015-0362-9.

Agullo:2017:BGB

[AAB^{+}17] Emmanuel Agullo, Olivier Aumage, Berenger Bramas, Olivier Coulaud, and Samuel Pitoiset. Bridging the gap between OpenMP and task-based runtime systems for the Fast Multipole Method. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2794–2807, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/10/07912335-
abs.html.

Almasi:2005:DIM

[AAC^{+}05] G. Almási, C. Archer, J. G.

Castaños, J. A. Gunnels, C. C. Erway, P. Heidelberger, X. Martorell, J. E. Moreira, K. Pinnow, J. Ratterman, B. D. Steinmacher-Burow, W. Gropp, and B. Toonen. Design and implementation of messagepassing services for the Blue Gene/L supercomputer. IBM Journal of Research and Development, 49 (2/3):393–406, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/almasi.
pdf.

Akzhalova:2008:WPL

[AASB08] Assel Zh. Akzhalova, Daniar Y. Aizhulov, Galymzhan Seralin, and Gulnar Balakayeva. Web portal for large-scale computations based on Grid and MPI. Scalable Computing: Practice and Experience, 9(2): 135–142, June 2008. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.
org/vols/vol09/no2/SCPE_
9_2_06.pdf;   http://www.
scpe.org/vols/vol09/no2/
SCPE_9_2_06.zip.

Arthur:1993:PIU

[AB93a] T. Arthur and M. Bockelie. A parallel implementation of the unstructured grid generation program VGRIDSG using PVM and APPL. In Sincovec [Sin93], pages 899– 902. ISBN 0-89871-315-3. LCCN QA 76.58 S55 1993. Two volumes.

Arthur:1993:CUA

[AB93b] Trey Arthur and Michael J.

Bockelie. A comparison of using APPL and PVM for a parallel implementation of an unstructured grid generation problem. Technical Report NASA CR-191425, National Aeronautics and Space Administration, Langley Research Center; National Technical Information Service, distributor, Hampton, VA, USA, 1993. ?? pp.

Aloisio:1995:UPW

[AB95] G. Aloisio and M. A. Bochicchio. The use of PVM with workstation clusters for distributed SAR data processing. In Hertzberger and Serazzi [HS95a], pages 570–581. ISBN 3-540-59393-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Augusto:2013:APG

[AB13] Douglas A. Augusto and Helio J. C. Barbosa. Accelerated parallel genetic programming tree evaluation with OpenCL. Journal of Parallel and Distributed Computing, 73(1): 86–100, January 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S074373151200024X.

Ayguade:2010:EOS

[ABB^{+}10] Eduard Ayguadé, Rosa M. Badia, Pieter Bellens, Daniel Cabrera, Alejandro Duran Roger Ferrer, Marc González, Francisco Igual, Daniel Jiménez-González, Jesús Labarta, Luis Martinell, Xavier Martorell, Rafael Mayo, Josep M. Pérez, Judit Planas, and Enrique S. Quintana-Ortí. Extending OpenMP to survive the heterogeneous multi-core era. International Journal of Parallel Programming, 38(5– 6):440–459, October 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=5&spage=440.

Arabnejad:2020:SSC

[ABB20] Hamid Arabnejad, João Bispo, and Jorge G. Barbosa. Source-to-source compilation targeting OpenMPbased automatic parallelization of C applications. The Journal of Supercomputing, 76(9):6753–6785, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

https://link.springer.
com/article/10.1007/s11227-
019-03109-9.

Adhianto:2000:TOA

[ABC^{+}00] L. Adhianto, F. Bodin, B. Chapman, L. Hascoet, A. Kneer, D. Lancaster, I. Wolton, and M. Wirtz. Tools for OpenMP application development: the POST project. Concurrency: practice and experience, 12(12):1177–1191, October 2000. CODEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76500357/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500357&PLACEBO=IE.
pdf.

Appiani:1995:PSI

[ABCI95a] E. Appiani, M. Bologna, M. Corvi, and M. Iardella. PVM in a shared-memory industrial multiprocessor. In Hertzberger and Serazzi [HS95a], pages 588–593. ISBN 3-540-59393-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Appiani:1995:PSM

[ABCI95b] E. Appiani, M. Bologna, M. Corvi, and M. Iardella. PVM in a shared-memory industrial multiprocessor. In Hertzberger and Serazzi [HS95a], pages 588–593. ISBN 3-540-59393-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Agosta:2015:OPP

[ABDP15] Giovanni Agosta, Alessandro Barenghi, Alessandro Di Federico, and Gerardo Pelosi. OpenCL performance portability for general-purpose computation on graphics processor units: an exploration on cryptographic primitives. Concurrency and Computation: Practice and Experience, 27(14):3633–3660, September 25, 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aliaga:2017:CTP

[ABF^{+}17] José I. Aliaga, María Barreda, Goran Flegar, Matthias Bollhöfer, and Enrique S. Quintana-Ortí. Communication in taskparallel ILU-preconditioned CG solvers using MPI + OmpSs. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arbenz:1996:MDS

[ABG^{+}96] P. Arbenz, M. Billeter, P. Güntert, P. Luginbühl, M. Taufer, and U. von Matt. Molecular dynamics simulations on Cray clusters using the SCIDDLE-PVM environment. In Bode et al. [BDLS96], pages 142–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Allegretti:2020:OBB

[ABG20] S. Allegretti, F. Bolelli, and C. Grana. Optimized block-based algorithms to label connected components on GPUs. IEEE Transactions on Parallel and Distributed Systems, 31(2):423– 438, February 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Abrahart:1996:GIC

[Abr96] R. J. Abrahart, editor. Geo-Computation 96. 1st International Conference on Geo-Computation: Leeds, UK, 17–19 September 1996. ????, ????, 1996. ISBN ???? LCCN ????

Adhianto:2007:PMC

[AC07] Laksono Adhianto and Barbara Chapman. Performance modeling of communication and computation in hybrid MPI and OpenMP applications. Simulation Modelling Practice and Theory, 15(4): 481–491, April 2007. CO-DEN SMPTCA. ISSN 1569-190X (print), 1878-1462 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S1569190X06001109.

Alvanos:2017:PMM

[AC17] Michail Alvanos and Theodoros Christoudias. MEDINA:

MECCA development in accelerators — KPP Fortran to CUDA source-to-source pre-processor. Journal of Open Research Software, 5(1):13–??, April 28, 2017. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.158/.

Aldinucci:2021:PPS

[ACC^{+}21] Marco Aldinucci, Valentina Cesare, Iacopo Colonnelli, Alberto Riccardo Martinelli, Gianluca Mittone, Barbara Cantalupo, Carlo Cavazzoni, and Maurizio Drocco. Practical parallelization of scientific applications with OpenMP, OpenACC and MPI. Journal of Parallel and Distributed Computing, 157 (??):13–29, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731521001295.

Ayguade:2009:DOT

[ACD^{+}09] Eduard Ayguade, Nawal Copty, Alejandro Duran, Jay Hoeflinger, Yuan Lin, Federico Massaioli, Xavier Teruel, Priya Unnikrishnan, and Guansong Zhang. The design of OpenMP tasks. IEEE Transactions on Parallel and Distributed Systems, 20(3):404–418, March 2009. CODEN ITDSEO.

ISSN 1045-9219 (print), 1558-2183 (electronic).

Arnold:1994:PCT

[ACDR94] D. Arnold, R. Christie, J. Day, and P. Roe, editors. Parallel Computing and Transputers. PCAT-93. Proceedings of the 6th Australian Transputer and Occam User Group Conference, November 3–4, 1993, Brisbane, Queensland, Australia, volume 37 of Transputer and Occam Engineering Series. IOS Press, Postal Drawer 10558, Burke, VA 2209-0558, USA, 1994. ISBN 90-5199-149-5. LCCN ????

Acacio:2002:MDM

[ACGdT02] M. Acacio, O. Cánovas, J. M. García, and P. E. López de Teruel. MPI-Delphi: an MPI implementation for visual programming environments and heterogeneous computing. Future Generation Computer Systems, 18(3):317–333, January 2002. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/32/28/abstract.
html.

Alexandrov:1997:PMC

[ACGR97] V. Alexandrov, K. Chan, A. Gibbons, and W. Rytter. On the PVM/MPI computations of dynamic programming recurrences. Lecture Notes in Computer Science, 1332:305–312, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agullo:2011:QOM

[ACH^{+}11] Emmanuel Agullo, Camille Coti, Thomas Herault, Julien Langou, Sylvain Peyronnet, Ala Rezmerita, Franck Cappello, and Jack Dongarra. QCG-OMPI: MPI applications on grids. Future Generation Computer Systems, 27(4):357– 369, April 2011. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Andersch:2012:PPE

[ACJ12] Michael Andersch, Chi Ching Chi, and Ben Juurlink. Programming parallel embedded and consumer applications in OpenMP superscalar. ACM SIGPLAN Notices, 47(8):281–282, August 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

ACM:1990:PAC

[ACM90] ACM, editor. Proceedings of the 1990 ACM Conference on LISP and Functional Programming: papers presented at the conference, Nice, France, June 27– 29, 1990. ACM Press, New York, NY 10036, USA, 1990. ISBN 0-89791-368-X. LCCN QA 76.73 L23 A24 1990. ACM order no. 552900.

ACM:1994:CPI

[ACM94] ACM, editor. Conference Proceedings. 1994 International Conference on Supercomputing. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-665-4. LCCN

????  URL http://www.
acm.org/pubs/contents/
proceedings/supercomputing/
181181/.

ACM:1995:PAS

[ACM95a] ACM, editor. Proceedings of the 33rd annual southeast conference [ACM]: Clemson, South Carolina, March 17– 18, 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-747-2. LCCN ????

ACM:1995:SAA

[ACM95b] ACM, editor. SPAA ’95, 7th Annual ACM Symposium on Parallel Algorithms and Architectures: July 17– 19, 1995, Santa Barbara, CA, USA, volume 7. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-717-0. LCCN QA76.642 .A25 1995.

ACM:1996:SVR

[ACM96a] ACM, editor. 1995 Symposium on the Virtual Reality Modeling Language (VRML ‘95). ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-818-5. LCCN

????  URL http://www.
acm.org/pubs/contents/
proceedings/graph/217306/

.

ACM:1996:FCP

[ACM96b] ACM, editor. FCRC ’96: Conference proceedings of the 1996 International Conference on Supercomputing: Philadelphia, Pennsylvania, USA, May 25–28, 1996. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

ACM:1996:SCP

[ACM96c] ACM, editor. Supercomputing ’96 Conference Proceedings: November 17– 22, Pittsburgh, PA. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-89791-854-1. LCCN QA 76.88 S8573 1996. URL http://

www.supercomp.org/sc96/
proceedings/. ACM Or-

der Number: 415962, IEEE Computer Society Press Order Number: RS00126.

ACM:1997:PPS

[ACM97a] ACM, editor. PASCO ’97. Proceedings of the second international symposium on parallel symbolic computation, July 20–22, 1997, Maui, HI. ACM Press, New York, NY 10036, USA, 1997. ISBN ???? LCCN ????

ACM:1997:SHP

[ACM97b] ACM, editor. SC’97: High Performance Networking and Computing: Proceedings of the 1997 ACM/IEEE SC97 Conference: November 15–21, 1997, San Jose, California, USA. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-89791-985-8. LCCN QA76.9.A25 A265

1997. URL http://www.
acm.org/pubs/contents/
proceedings/commsec/266741/
;   http://www.supercomp.
org/sc97/proceedings/.

ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

ACM:1998:AWJ

[ACM98a] ACM, editor. ACM 1998 Workshop on Java for High-Performance Network Computing. ACM Press, New York, NY 10036, USA, 1998. ISBN ???? LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/
program.html. Possibly un-

published, except electronically.

ACM:1998:SHP

[ACM98b] ACM, editor. SC’98: High Performance Networking and Computing: Proceedings of the 1998 ACM/IEEE SC98 Conference: Orange County Convention Center, Orlando, Florida, USA, November 7–13, 1998. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN ???? LCCN ???? URL http://

                                        www.supercomp.org/sc98/
                                        papers/.

ACM:1999:SPO

[ACM99] ACM, editor. SC’99: Oregon Convention Center 777 NE Martin Luther King Jr. Boulevard, Portland, Oregon, November 11–18, 1999. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999.

ACM:2000:SHP

[ACM00]  ACM, editor.  SC2000:

High Performance Networking and Computing. Dallas Convention Center, Dallas, TX, USA, November 4–10, 2000. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA,

2000. URL http://www.
sc2000.org/proceedings/
info/fp.pdf.

ACM:2001:SHP

[ACM01]  ACM, editor.  SC2001:

High Performance Networking and Computing. Denver, CO, November 10–16, 2001. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 1-58113-293-X. LCCN ????

ACM:2003:SII

[ACM03] ACM, editor. SC2003: Igniting Innovation. Phoenix, AZ, November 15–21, 2003. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 1-58113-695-1. LCCN ????

ACM:2004:SHP

[ACM04] ACM, editor. SC 2004:

High Performance Computing, Networking and Storage: Bridging communities: Proceedings of the IEEE/ ACM Supercomputing 2004 Conference, Pittsburgh, PA, November 6–12, 2004. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2153-3. LCCN ????

ACM:2005:PAI

[ACM05] ACM, editor. Proceedings of the 2005 ACM/IEEE conference on Supercomputing 2005, Seattle, WA, November 12–18 2005. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 1-59593-061-2. LCCN ????

ACM:2006:PST

[ACM06a]  ACM, editor.  Proceed-

ings of the 37th SIGCSE technical symposium on Computer science education 2006, Houston, Texas, USA, March 03–05, 2006. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-259-3. ACM order number 457060.

ACM:2006:PCC

[ACM06b] ACM, editor. Proceedings of the 3rd conference on Computing Frontiers, May 3–5, 2006, Ischia, Italy. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-302-6. ACM order number 104060.

ACM:2011:SSP

[ACM11] ACM, editor. SC ’11 State of the Practice Reports. ACM Press, New York, NY 10036, USA, 2011. ISBN 1-4503-1139-3. LCCN ????

Antonelli:2014:ATS

[ACMR14] Laura Antonelli, Stefania Corsaro, Zelda Marino, and Mariarosaria Rizzardi. Algorithm 944: Talbot suite: Parallel implementations of Talbot’s method for the numerical inversion of Laplace transforms. ACM Transactions on Mathematical Software, 40(4):29:1–29:18, June 2014. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Alonso:2011:NEM

[ACMZR11] P. Alonso, R. Cortina, F. J. Martínez-Zaldívar, and J. Ranilla. Neville elimination on multi- and many-core systems: OpenMP, MPI and CUDA. The Journal of Supercomputing, 58(2): 215–225, November 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
58&issue=2&spage=215.

Ancona:1995:PAD

[AD95] M. Ancona and M. De Benedetto. A parallel algorithm for ‘document segmentation’. In IEEE [IEE95h], pages 516–521. ISBN 0-8186-7031-2, 0-8186-7032-0. LCCN QA76.58 .E97 1995.

Alexandrov:1998:RAP

[AD98] Vassil Alexandrov and J. J. Dongarra, editors. Recent advances in parallel virtual machine and message passing interface: 5th European PVM/MPI User’s Group Meeting, Liverpool, UK, September 7–9, 1998: proceedings, volume 1497 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-65041-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1497. Jointly sponsored by the Computer Science Dept., University of Liverpool and Oak Ridge National Laboratory.

Adamo:1997:AOO

[Ada97] J.-M. Adamo. ARCH, an object oriented MPI-based library for asynchronous and loosely synchronous parallel system programming. Lecture Notes in Computer Science, 1332:67–74, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adamo:1998:MTO

[Ada98] Jean-Marc Adamo. Multithreaded object-oriented MPIbased message passing interface: the ARCH library, volume SECS 446 of The Kluwer international series in engineering and computer science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998. ISBN 0-7923-8165-3. xiv + 185 pp. LCCN TK5102.5.A293 1998. US$120.00.

Antonuccio-Delogu:1994:PTN

[ADB94] V. Antonuccio-Delogu and U. Becciani. A parallel tree N-body code for heterogeneous clusters. In Dongarra and Wasniewski [DW94], pages 17–32. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1994. DM104.00.

Addison:2001:EOP

[Add01]  Cliff Addison.  Exploit-

ing OpenMP to provide scalable SMP BLAS and LAPACK routines. Lecture Notes in Computer Science, 2073:3–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2073/20730003.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2073/20730003.
pdf.

Arioli:1995:PSB

[ADDR95] M. Arioli, A. Drummond, I. S. Duff, and D. Ruiz. A parallel scheduler for block iterative solvers in heterogeneous computing environments. In Bailey et al. [BBG^{+}95], pages 460–465. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Adamek:2020:GFC

[ADGA20] Karel Adámek, Sofia Dimoudi, Mike Giles, and Wesley Armour. GPU fast convolution via the overlapand-save method in shared memory. ACM Transactions on Architecture and Code Optimization, 17(3): 18:1–18:20, August 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3394116.

Agathos:2022:CAA

[ADK22] Spiros N. Agathos, Vassilios V. Dimakopoulos, and Ilias K. Kasmeridis. Compilerassisted, adaptive runtime system for the support of OpenMP in embedded multicores. Parallel Computing, 110(??):??, May 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819122000035.

Amestoy:2003:IIMa

[ADLL03a] Patrick R. Amestoy, Iain S. Duff, Jean-Yves L’Excellent, and Xiaoye S. Li. Impact of the implementation of MPI point-to-point communications on the performance of two general sparse solvers. Report TR/PA/03/14 and RR-4372 and LBNL-48968 and RT/APO/01/4, CER-FACS, Toulouse, France, 2003. ???? pp.

Amestoy:2003:IIMb

[ADLL03b] Patrick R. Amestoy, Iain S.

Duff, Jean-Yves L’Excellent, and Xiaoye S. Li. Impact of the implementation of MPI point-to-point communications on the performance of two general sparse solvers. Parallel Computing, 29(7):833–849, July 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Aversa:2005:HDS

[ADMV05] Rocco Aversa, Beniamino Di Martino, Nicola Mazzocca, and Salvatore Venticinque. A hierarchical distributed-shared memory parallel Branch & Bound application with PVM and OpenMP for multiprocessor clusters. Parallel Computing, 31(10–12):1034–1047, October/December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Aversa:2005:PPT

[ADR^{+}05] Rocco Aversa, Beniamino Di Martino, Massimiliano Rak, Salvatore Venticinque, and Umberto Villano. Performance prediction through simulation of a hybrid MPI/ OpenMP application. Parallel Computing, 31(10– 12):1013–1033, October/ December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Alexandrov:1998:CGP

[ADRCT98] V. Alexandrov, F. Dehne, A. Rau-Chaplin, and K. Taft. Coarse grained parallel Monte Carlo algorithms for solving SLAE using PVM. Lecture Notes in Computer Science, 1497:323–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amritkar:2014:EPC

[ADT14] Amit Amritkar, Surya Deb, and Danesh Tafti. Efficient parallel CFD-DEM simulations using OpenMP. Journal of Computational Physics, 256(??):501–519, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0021999113006128.

Aldea:2016:OES

[AELGE16] Sergio Aldea, Alvaro Estebanez, Diego R. Llanos, and Arturo Gonzalez-Escribano.

An OpenMP extension that supports thread-level speculation. IEEE Transactions on Parallel and Distributed Systems, 27(1):78– 91, January 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07014262-
abs.html.

Amos:2020:AQQ

[AEW^{+}20] Brandon D. Amos, David R.

Easterling, Layne T. Watson, William I. Thacker, Brent S. Castle, and Michael W. Trosset. Algorithm 1007: QNSTOP — quasi-Newton algorithm for stochastic optimization. ACM Transactions on Mathematical Software, 46(2):17:1–17:20, June 2020. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3374219.

Antonelli:2021:CBI

[AFG21] L. Antonelli, E. Francomano, and F. Gregoretti. A CUDA-based implementation of an improved SPH method on GPU. Applied Mathematics and Computation, 409(??):Article 125482, November 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300320304410.

Azimi:2018:SVS

[AFGR18] Reza Azimi, Tyler Fox, Wendy Gonzalez, and Sherief Reda. Scale-out vs scaleup: A study of ARM-based SoCs on server-class workloads. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3 (4):18:1–18:??, September 2018. CODEN ???? ISSN 2376-3639. URL https:

                                        //dl.acm.org/citation.
                                        cfm?id=3232162.

Ashby:1995:PPG

[AFST95] S. F. Ashby, R. D. Falgout, S. G. Smith, and A. F. B. Tompson. The parallel performance of a groundwater flow code on the Cray T3D. In Bailey et al. [BBG^{+}95], pages 131–136. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Ayguade:1995:DUA

[AGG^{+}95] E. Ayguade, J. Garcia, M. Girones, J. Labarta, J. Torres, and M. Valero. Detecting and using affinity in an automatic data distribution tool. In Pingali et al. [PBG^{+}95], pages 61– 75. ISBN 3-540-58868-X. LCCN QA76.58 .W656 1994.

Aityan:1995:PFI

[AGH^{+}95] S. K. Aityan, L. T. Grujic, R. J. Hathaway, G. S. Ladde, N. Medhin, and M. Sambandham, editors. Proceedings of the First International Conference on Neural, Parallel and Scientific Computations held at Morehouse College, Atlanta, USA, May 28–31, 1995, Proceedings of Neural Parallel and Scientific Computations 1995. Dynamic Publishers, Atlanta, GA, USA, 1995. ISBN 0-9640398-9-3 (hardback) 0-9640398-8-5 (paperback). LCCN QA76.87 .I58 1995.

Averbuch:1994:PES

[AGIS94] A. Averbuch, E. Gabber, S. Itzikowitz, and B. Shoham. On the parallel elliptic single/multigrid solutions about aligned and nonaligned bodies using the Virtual Machine for Multiprocessors. Scientific Programming, 3(1):13– 32, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Arbenz:1996:SRP

[AGLv96] P. Arbenz, W. Gander, H. P.

Lüthi, and U. von Matt. Sciddle 4.0, or, remote procedure calls in PVM. In Liddell et al. [LCHS96], pages 820–?? ISBN 3-540-61142-8 (paperback). LCCN QA76.88 .H52 1996.

Ayguade:2006:ENO

[AGMJ06] Eduard Ayguade, Marc Gonzalez, Xavier Martorell, and Gabriele Jost. Employing nested OpenMP for the parallelization of multizone computational fluid dynamics applications. Journal of Parallel and Distributed Computing, 66(5): 686–697, May 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Agrawal:1995:PIW

[Agr95a] D. P. Agrawal, editor. Proceedings of the 1995 ICPP Workshop on Challenges for Parallel Processing, August 14, 1995, Raleigh, NC, USA. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1995. ISBN 0-8493-2618-4. LCCN QA76.58.I34 1995.

Almeida:1995:CST

[AGR^{+}95b] F. Almeida, F. Garcia, J. Roda, D. Morales, Rodriguez, and C. A comparative study of two distributed systems: PVM and transputers. In Cook et al. [CJNW95], pages 244– 258. ISBN 90-5199-235-1 (IOS Press), 4-274-90062-2 (Ohmsha). LCCN ????

Alfaro:1997:FDW

[AGS97] F. J. Alfaro, J. A. Gallud, and J. L. Sanchez. A function to dynamic workload allocation in distributed applications. Lecture Notes in Computer Science, 1332:

219–225, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alnuweiri:1995:PHF

[AH95] Hussein M. Alnuweiri and Mounir Hamdi, editors. Proceedings of HiNet ’95: first international workshop on high-speed network computing, April 25, 1995, Santa Barbara, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7124-6. LCCN TK5105.5 .H56 1995.

Astalos:2000:CMS

[AH00] Ján Astalos and Ladislav Hluchý. CIS — a monitoring system for PC clusters. Lecture Notes in Computer Science, 1908:225–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080225.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080225.
pdf.

Agathos:2012:TBE

[AHD12] Spiros N. Agathos, Panagiotis E. Hadjidoukas, and Vassilios V. Dimakopoulos. Task-based execution of nested OpenMP loops. Lecture Notes in Computer Science, 7312: 210–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-30961-8_
16/.

Awan:2017:CCD

[AHHP17] Ammar Ahmad Awan, Khaled Hamidouche, Jahanzeb Maqbool Hashmi, and Dhabaleswar K. Panda. S-Caffe: Co-designing MPI runtimes and Caffe for scalable deep learning on modern GPU clusters. ACM SIGPLAN Notices, 52(8):193–205, August 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ahmad:1997:EVP

[Ahm97]  Ishfaq Ahmad.  Express

versus PVM: a performance comparison. Parallel Computing, 23(6):783– 812, June 20, 1997. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=23&
issue=6&aid=1138.

Allsopp:2001:EUM

[AHP01] Nicholas K. Allsopp, John F. Hague, and Jean-Pierre Prost. Experiences in using MPI–IO on top of GPFS for the IFS weather forecast code. Lecture Notes in Computer Science, 2150:380–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2150/21500380.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2150/21500380.
pdf.

Andoh:2021:AMM

[AiIS^{+}21] Yoshimichi Andoh, Shin ichi Ichikawa, Tatsuya Sakashita, Noriyuki Yoshii, and Susumu Okazaki. Algorithm to minimize MPI communications in the parallelized fast multipole method combined with molecular dynamics calculations. Journal of Computational Chemistry, 42(15): 1073–1087, June 5, 2021. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Aversa:1997:MDP

[AIM97] R. Aversa, G. Iannello, and N. Mazzocca. An MPI driven parallelization strategy for different computing platforms: a case study. Lecture Notes in Computer Science, 1332:401–408, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aguilar:1997:PMS

[AJ97] J. Aguilar and T. Jimenez. A processors management system for PVM. Lecture Notes in Computer Science, 1300:158–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Awan:2020:CPC

[AJC^{+}20] A. A. Awan, A. Jain, C. Chu, H. Subramoni, and D. K. Panda. Communication profiling and characterization of deep-learning workloads on clusters with high-performance interconnects. IEEE Micro, 40 (1):35–43, January 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Aubrey-Jones:2016:SMI

[AJF16] Tristan Aubrey-Jones and Bernd Fischer. Synthesizing MPI implementations from functional dataparallel programs. International Journal of Parallel Programming, 44(3):552– 573, June 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10766-015-0359-4.

AlKadi:2018:GPC

[AJYH18] Muhammed Al Kadi, Benedikt Janssen, Jones Yudi, and Michael Huebner. Generalpurpose computing with soft GPUs on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(1):5:1– 5:??, March 2018. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Alexandrov:1999:PMC

[AK99] V. Alexandrov and A. Karaivanova.

Parallel Monte Carlo algorithms for sparse SLAE using MPI. In Dongarra et al. [DLM99], pages 283–

290. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Adam:2019:CRA

[AKB^{+}19] Julien Adam, Maxime Kermarquer, Jean-Baptiste Besnard, Leonardo Bautista-Gomez, Marc Pérache, Patrick Carribault, Julien Jaeger, Allen D. Malony, and Sameer Shende. Checkpoint/restart approaches for a thread-based MPI runtime. Parallel Computing, 85(??):204–219, July 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819118303247.

Armstrong:2000:QDB

[AKE00] Brian Armstrong, Seon Wook Kim, and Rudolf Eigenmann. Quantifying differences between OpenMP and MPI using a largescale application suite. Lecture Notes in Computer Science, 1940:482–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1940/19400482.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1940/19400482.
                                        pdf.

Andersen:1994:PIA

[AKK^{+}94] B. S. Andersen, P. Kaae, C. Keable, W. Owczarz, J. Wasniewski, and Z. Zlatev. PVM implementations of advection-chemistry modules of air pollution models. In Dongarra and Wasniewski [DW94], pages 11– 16. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1994. DM104.00.

Asai:1999:MIF

[AKL99] Noboru Asai, Thomas Kentemich, and Pierre Lagier. MPI-2 implementation on a Fujitsu Generic Message Passing Kernel. In ACM [ACM99], page ??

Abdelfattah:2016:KOL

[AKL16] Ahmad Abdelfattah, David Keyes, and Hatem Ltaief. KBLAS: an optimized library for dense matrixvector multiplication on GPU accelerators. ACM Transactions on Mathematical Software, 42(3):18:1– 18:31, May 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Alfano:1992:DNA

[AL92] M. Alfano and G. Lo Re.

Distributing numerical algorithms: some experiences with network computing system (NCS) and parallel virtual machine (PVM). In SCRI WCC’92 [SCR92], page ?? ISBN ???? LCCN ???? Proceedings available via anonymous

ftp from ftp.scri.fsu.edu
in directory pub/parallel-
workshop.92.

Altevogt:1993:PTD

[AL93] P. Altevogt and A. Linke.

Parallelization of the twodimensional Ising model on a cluster of IBM RISC System/6000 workstations. Parallel Computing, 19(9): 1041–1052, September 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Alt:1996:PIA

[AL96] R. Alt and J. L. Lamotte. Parallel integration across time of initial value problems using PVM. In Bode et al. [BDLS96], pages 323–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Amer:2018:LCM

[ALB^{+}18] Abdelhalim Amer, Huiwei Lu, Pavan Balaji, Milind Chabbi, Yanjie Wei, Jeff Hammond, and Satoshi Matsuoka. Lock contention management in multithreaded MPI. ACM Transactions on Parallel Computing (TOPC), 5(3):12:1–12:??, January 2018. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3275443.

Alund:1994:CFD

[ALR94] A. Alund, P. Lotstedt, and R. Ryden. Computational fluid dynamics on workstation clusters in industrial environments. In Dongarra and Wasniewski [DW94], pages 1–10. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1994. DM104.00.

Amer:2015:MRC

[ALW^{+}15] Abdelhalim Amer, Huiwei Lu, Yanjie Wei, Pavan Balaji, and Satoshi Matsuoka. MPI+Threads: runtime contention and remedies. ACM SIGPLAN Notices, 50(8):239–248, August 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ayguade:2007:SIO

[AM07] Eduard Ayguadé and Matthias S.

Mueller. Special issue on OpenMP — Guest Editors’ introduction. International Journal of Parallel Programming, 35(4):331– 333, August 2007. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=4&spage=331.

Almasi:1993:PDS

[AMBG93] G. S. Almasi, T. McLuckie, J. Bell, and A. Gordon. Parallel distributed seismic migration. Concurrency: practice and experience, 5(2): 105–131, April 1993. CO-DEN CPEXEI. ISSN 1040-3108.

Awan:2019:OLM

[AMC^{+}19] Ammar Ahmad Awan, Karthik Vadambacheri Manian, Ching-Hsiang Chu, Hari Subramoni, and Dhabaleswar K. Panda. Optimized large-message broadcast for deep learning workloads: MPI, MPI + NCCL, or NCCL2? Parallel Computing, 85(??):141–152, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167819118303284.

Agrawal:2011:PPS

[AMHC11] Ankit Agrawal, Sanchit Misra, Daniel Honbo, and Alok Choudhary. Parallel pairwise statistical significance estimation of local sequence alignment using Message Passing Interface library. Concurrency and Computation: Practice and Experience, 23(17):2269– 2279, December 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Al-Mouhamed:2020:RCO

[AMKM20] Mayez A. Al-Mouhamed, Ayaz H. Khan, and Nazeeruddin Mohammad. A review of CUDA optimization techniques and tools for structured grid computing. Computing, 102(4):977– 1003, April 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ayguade:1999:EML

[AML^{+}99] E. Ayguade, X. Martorell, J. Labarta, M. Gonzalez, and N. Navarro. Exploiting multiple levels of parallelism in OpenMP: a case study. In ????, editor, Proceedings of the 1999 International Conference on Parallel Processing, pages 172– 180. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999.

Amato:1994:PEP

[AMS94] M. Amato, A. Matrone, and P. Schiano. A practical experience in parallelizing a large CFD code: the ENSOLV flow solver. In Gentzsch and Harms [GH94], pages 508–513. ISBN 0-387-57981-8 (New York), 3-540-57981-8 (Berlin). LCCN QA76.88.I57 1994. DM96.00. Two volumes.

anMey:2007:NPO

[aMST07] Dieter an Mey, Samuel Sarholz, and Christian Terboven. Nested parallelization with OpenMP. International Journal of Parallel Programming, 35(5): 459–476, October 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=5&spage=459.

Al-Mouhamed:2015:EAO

[AMuHK15] Mayez Al-Mouhamed and Ayaz ul Hassan Khan. Exploration of automatic optimisation for CUDA programming. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 30(4):309–324, 2015. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

                                        //www.tandfonline.com/
                                        doi/abs/10.1080/17445760.
                                        2014.953158.

Aversa:1994:PSH

[AMV94] R. Aversa, N. Mazzocca, and U. Villano. PS: a simulator for heterogeneous computing environments. In Dekker et al. [DSZ94], pages 335– 343. ISBN 0-444-81784-0. LCCN QA76.58.E98 1994.

Andersson:1998:PFT

[And98] U. Andersson. Parallelization of a 3D FD-TD code for the Maxwell equations using MPI. Lecture Notes in Computer Science, 1541:12– 19, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1989:PFC

[Ano89] Anonymous, editor. Proceedings of the Fourth Conference on Hypercubes, Concurrent Computers and Applications, 6–8 March 1989, Monterey, CA, USA. Golden Gate Enterprises, Los Altos, CA, USA, 1989. LCCN QA76.5.C619215 1989. Two volumes.

Anonymous:1992:PSE

[Ano92] Anonymous, editor. Proceedings SHARE Europe Anniversary Meeting. SHARE Eur. Assoc, Geneva, Switzerland, 1992.

Anonymous:1993:ATA

[Ano93a] Anonymous, editor. Automotive technology and automation: Supercomputer applications in the automotive industries: 26th International symposium — September 1993, Aachen, Germany, ISATA — Proceedings — 26th. Automotive Automation Ltd, Croydon, UK, 1993. ISBN 0-947719-62-8. LCCN ????

Anonymous:1993:ISA

[Ano93b] Anonymous, editor. International section: Annual conference — September 1993, Gallipoli, Italy, Atti del Congresso Annuale — Associazione Italiana per l’Informatica ed il Calcolo Automatico 1993. AICA, ????, 1993. ISBN ???? LCCN ????

Anonymous:1993:JFI

[Ano93c] Anonymous, editor. Joint framework for information technology: Technical conference — March 1993, Keele, JFIT Technical Conference Digest. Dept. of Trade and Industry, Information and Manufacturing Division, London, UK, 1993. ISBN ???? LCCN ????

Anonymous:1993:MPI

[Ano93d]  Anonymous.  Message-

passing interface. The International Journal of Supercomputer Applications, 7 (2):179, June 1993. CODEN IJSAE9. ISSN 0890-2720.

                                        URL http://journals.
                                        sagepub.com/doi/pdf/10.
                                        1177/109434209300700208.

Anonymous:1993:MMP

[Ano93e] Anonymous. MPI: a message passing interface. Proceedings of the Supercomputing Conference, pages 878– 883, ???? 1993. CODEN ???? ISBN 0-8186-4340-4. ISSN 1063-9535.

Anonymous:1993:PSE

[Ano93f] Anonymous, editor. Proceedings. SHARE Europe Anniversary Meeting. Client/Server — the Promise and the Reality: October 25–28, 1993, the Hague, the Netherlands. SHARE Europe, Geneva, Switzerland, 1993. ISBN ???? ISSN 0254-6213. LCCN ????

Anonymous:1993:SEC

[Ano93g] Anonymous, editor. Supercomputing Europe ’93. Conference Papers. Royal Dutch Fairs, Utrecht, Netherlands, 1993. ISBN ???? LCCN ????

Anonymous:1993:CDP

[Ano93h] Anonymous, editor. The commercial dimensions of parallel computing: UNI-COM seminar — April 1993, London. Unicom Seminars Ltd, ????, 1993. ISBN ???? LCCN ????

Anonymous:1994:ICS

[Ano94a] Anonymous, editor. 1994 International Computer Symposium Conference Proceedings. Nat. Chiao Tung Univ, Hsinchu, Taiwan, 1994. ISBN ???? LCCN ???? 2 vol.

Anonymous:1994:ALM

[Ano94b] Anonymous. Adaptive load migration systems for PVM. In IEEE [IEE94h], pages 390–399. ISBN 0-8186-6607-2, 0-8186-6605-6, 0-8186-6606-4. ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:FWR

[Ano94c] Anonymous, editor. Forschung und wissenschaftliches Rechnen: Beitrage anasslich des 10. EDV-Benutzertreffens der Max-Planck-Gesellschaft in Göttingen, November 1993, number 1 in Berichte und Mitteilungen — Max Planck Gesellschaft. Max-Planck-Gesellschaft, München, Germany, 1994. ISBN ????

ISSN 0341-7778. LCCN Q180.55.E4 M39 1993.

Anonymous:1994:MMP

[Ano94d]  Anonymous.  MPI: a

message-passing interface standard. International Journal of Supercomputer Applications and High Performance Computing, 8 (3/4):159–416, Fall-Winter 1994. CODEN IJSAE9. ISSN 0890-2720.

Anonymous:1994:PDC

[Ano94e] Anonymous, editor. Parallel and distributed computing systems: proceedings of the ISCA International Conference, Las Vegas, Nevada, U.S.A., October 6–8, 1994. ISCA, Raleigh, NC, USA, 1994. ISBN 1-880843-09-9. LCCN QA76.58.I543 1994.

Anonymous:1994:PPC

[Ano94f] Anonymous, editor. Parallel processing comes of age: real applications from industry and commerce: Seminar — June 1994, London. Unicom Seminars, ????, 1994. ISBN ???? LCCN ????

Anonymous:1994:PSE

[Ano94g] Anonymous, editor. Proceedings. SHARE Europe Spring Conference. SHARE Europe (SEAS), Carouge/Geneva, Switzerland, 1994. ISBN ???? LCCN ????

Anonymous:1994:SCC

[Ano94h] Anonymous, editor. Small college computing: 27th Annual symposium — April 1994, Winona, MN, SCCS — Proceedings — 27th. SCCS, ????, 1994. ISBN ???? LCCN ????

Anonymous:1994:SQC

[Ano94i] Anonymous, editor. Software quality concern for people: proceedings of the fourth European Conference on Software Quality, October 17–20, 1994, Basel, Switzerland. vdf Verlag der Fachvereine, Zurich, Switzerland, 1994. ISBN 3-7281-2153-3. LCCN ????

Anonymous:1995:CCS

[Ano95a] Anonymous, editor. 3rd CLIPS conference — September 1994, Houston, TX, NASA Publications N N95-19625-647, N95-19747-768. National Aeronautics and Space Administration, Washington, DC, USA, 1995. ISBN ???? LCCN ????

Anonymous:1995:BRPb

[Ano95b] Anonymous. Book review:

PVM: Parallel virtual machine: a users’ guide and tutorial for networked parallel computing: By Al Geist, Adam Beguelin, Jack Dongarra, Weicheng Jiang, Robert Manchek and Vaidy Sunderam. MIT Press, Cambridge, MA. (1994). 279 pages. $19.95. Computers and Mathematics with Applications, 30(9):122, November 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0898122195901973.

Anonymous:1995:BRU

[Ano95c] Anonymous. Book review: Using MPI: Portable parallel programming with the message-passing interface: By William Gropp, Ewing Lusk and Anthony Skjellum. MIT Press, Cambridge, MA. (1994). 307 pages. $24.95. Computers and Mathematics with Applications, 30(9): 122, November 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/089812219590199X.

Anonymous:1995:RSS

[Ano95d] Anonymous, editor. Reservoir simulation: 13th Symposium — February 1995, San Antonio, TX, Papers

— Society of Petroleum Engineers of AIME. Society of Petroleum Engineers, Richardson, TX, USA, 1995. ISBN ???? LCCN ????

Anonymous:1995:UPH

[Ano95e] Anonymous. Using PVM to host CLIPS in distributed environments. In 3rd CLIPS conference — September 1994, Houston, TX [Ano95a], pages 203–

211. ISBN ???? LCCN ????

Anonymous:1996:BRMh

[Ano96a]  Anonymous.  Book re-

view: MPI: the compete reference: By Marc Snir, Steve Otto, Steven Huss-Lederman, David Walker, and Jack Dongarra. MIT Press, Cambridge, MA. (1996). 336 pages. $27.50. Computers and Mathematics with Applications, 31 (11):140, June 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0898122196873494.

Anonymous:1996:IPP

[Ano96b]  Anonymous.  An intro-

duction to PVM programming. World-Wide Web,

1996. URL http://www.
epm.ornl.gov/pvm/intro.
html.

Anonymous:1996:PPA

[Ano96c] Anonymous. Porting PVM applications to the Intel Paragon. World-Wide Web, 1996. URL http:/

/www.ccs.ornl.gov/news/
guide/xps_pvm.html.

Anonymous:1996:RP

[Ano96d] Anonymous. Research program. World-Wide Web,

1996. URL http://www.

epm.ornl.gov/networking/

. Anonymous:1997:TNR

[Ano97]  Anonymous.  Technology

news & reviews: Chemkin software; OpenMP Fortran Standard; ODE toolbox for Matlab; Java products; Scientific WorkPlace 3.0. IEEE Computational Science & Engineering, 4(4):75–??, October/December 1997. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1997/pdf/c4075.
pdf.

Anonymous:1998:ANO

[Ano98]  Anonymous.  Announce-

ments: New official Fortran technical reports; working group 5 documents; OpenGL Fortran 95 bindings; MPI module provides enhanced Fortran support; variable precision arithmetic; Fortran information sites; new Fortran compiler versions from Lahey and Fujitsu; downloadable advanced Fortran textbook; Fortran engineering textbook. ACM Fortran Forum, 17(3):1–2, December 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1999:BRMa

[Ano99a]  Anonymous.  Book re-

view: MPI — The complete reference: Volume 1, the MPI core, second edition: By Marc Snir, Steve Otto, Steven Huss-Lederman, David Walker and Jack Dongarra. MIT Press, Cambridge, MA. (1998). 426 pages. $35.00. Computers and Mathematics with Applications, 37(3): 130, February 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122199903590.

Anonymous:1999:BRMf

[Ano99b]  Anonymous.  Book re-

view: MPI — The complete reference: Volume 1, the MPI core, second edition: By Marc Snir, Steve Otto, Steven Huss-Lederman, David Walker and Jack Dongarra. MIT Press, Cambridge, MA (1998). 426 pages. $35.00. Computers and Mathematics with Applications, 37 (6):130, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122199902237.

Anonymous:1999:BRMb

[Ano99c]  Anonymous.  Book re-

view: MPI-The complete reference: Volume 2, the MPI-2 extensions: By William Gropp, Steven Huss-Lederman, Andrew Lumsdaine, Ewing Lusk, Bill Nitzberg, William Saphir and Marc Snir. MIT Press, Cambridge, MA. (1998). 344 pages. $35.00. Computers and Mathematics with Applications, 37(3): 130, February 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122199903619.

Anonymous:1999:BRMg

[Ano99d]  Anonymous.  Book re-

view: MPI-The complete reference: Volume 2, the MPI-2 extensions: By William Gropp, Steven Huss-Lederman, Andrew Lumsdaine, Ewing Lusk, Bill Nitzberg, William Saphir and Marc Snir. MIT Press, Cambridge, MA. (1998). 344 pages. $35.00. Computers and Mathematics with Applications, 37(6): 130, March 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122199902250.

Anonymous:2000:BRUd

[Ano00a] Anonymous. Book review: Using MPI-2: Advanced features of the message-passing interface: By William Gropp, Ewing Lusk and Rajeev Thakur. The MIT Press, Cambridge, MA.

(1999). 382 pages. $35 (each); $60 (set). Computers and Mathematics with Applications, 40(2–3):419, July/August 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122100902098.

Anonymous:2000:BRUe

[Ano00b]  Anonymous.  Book re-

view: Using MPI: Portable parallel programming with the message-passing interface: Second edition. By William Gropp, Ewing Lusk and Anthony Skjellum. The MIT Press, Cambridge, MA. (1999). 371 pages. $35 (each); $60 (set). Computers and Mathematics with Applications, 40(2–3):419, July/August 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122100902074.

Anonymous:2001:AAL

[Ano01a] Anonymous. Appendixes:

Appendix A: Linux, Windows NT, AIX, Solaris; appendix B: Compilers and preprocessors, MPI implementations, development environments, debuggers, performance analyzers. The International Journal of High Performance Computing Applications, 15(2):191– 194, Summer 2001. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

                                        //journals.sagepub.com/
                                        doi/pdf/10.1177/109434200101500213.

Anonymous:2001:EDP

[Ano01b] Anonymous. Erratum: Design and prototype of a performance tool interface for OpenMP. The Journal of Supercomputing, 23 (1):105–128, May 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0920-8542&volume=
                                        23&issue=1&spage=105.

Anonymous:2003:MNIc

[Ano03] Anonymous. Micro news: IBM ups the ante in silicon transistor speed; new benchmark suite based on high-performance computing applications, MPI and OpenMP [SPEC HPC2002]; EU OKs Hitachi, Mitsubishi Electric semiconductor joint venture; Intel launches Pentium 4 at 3.06 GHz; TSMC unveils viable 25nm transistors. IEEE Micro, 23(1):6–6, 87, January/February 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2003/pdf/m1006.
                                        pdf.

Anonymous:2012:CTC

[Ano12] Anonymous. CUDA Toolkit 5.0 CURAND guide. Web document, 2012. URL http:

//docs.nvidia.com/cuda/
pdf/CURAND_Library.pdf.

ANS:1995:MCR

[ANS95] ANS, editor. Mathematics and computations, reactor physics, and environmental analyses: International conference — April 1995, Portland, OR. American Nuclear Society, La Grange Park, IL, USA, 1995. ISBN 0-89448-198-3. LCCN TK9006.M37 1995. Two volumes.

Anglano:1996:PMB

[AP96] C. Anglano and L. Portinale. Parallel model-based diagnosis using PVM. Lecture Notes in Computer Science, 1156:331–334, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aji:2016:MEA

[APBcF16] Ashwin M. Aji, Antonio J.

Peña, Pavan Balaji, and Wu chun Feng. MultiCL: Enabling automatic scheduling for task-parallel workloads in OpenCL. Parallel Computing, 58(??):37– 55, October 2016. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819116300357.

Aji:2016:MAA

[APJ^{+}16] Ashwin M. Aji, Lokendra S. Panwar, Feng Ji, Karthik Murthy, Milind Chabbi, Pavan Balaji, Keith R. Bisset, James Dinan, Wu chun Feng, John Mellor-Crummey, Xiaosong Ma, and Rajeev Thakur. MPI-ACC: Accelerator-aware MPI for scientific applications. IEEE Transactions on Parallel and Distributed Systems, 27 (5):1401–1414, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07127020-
abs.html.

AlHaddad:2001:UNW

[AR01] Mohammed Al Haddad and Jerome Robinson. Using a network of workstations to enhance database query processing performance. Lecture Notes in Computer Science, 2131:352–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310352.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310352.
pdf.

Arabnia:1995:TRA

[Ara95] Hamid Arabnia, editor. Transputer research and applications 7: American Transputer Users Group, October 23–25, 1994, Atlanta, GA (NATUG-7), volume 42 of Transputer and occam engineering series. IOS Press, Postal Drawer 10558, Burke, VA 2209-0558, USA, 1995. ISBN 90-5199-187-8 (IOS Press), 4-274-90017-7 (Ohmsha). ISSN 0925-4986. LCCN ????

Altas:1994:NIE

[ARL^{+}94] I. Altas, M. Rezny, J. Louis, K. Burrage, R. Moore, and J. Belward. A new image enhancement algorithm on MasPar and Parallel Virtual Machine (PVM) environments. In Dekker et al. [DSZ94], pages 819– 826. ISBN 0-444-81784-0. LCCN QA76.58.E98 1994.

Arnow:1995:DLB

[Arn95] D. M. Arnow. DP: a library for building portable, reliable distributed applications. In USENIX [USE95], pages 235–247. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

Abrossimov:1989:GVM

[ARS89] V. Abrossimov, M. Rozier, and M. Shapiro. Generic virtual memory management for operating system kernels. Operating Systems Review, 23(5):123–136, 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Al-Refaie:2017:PAH

[ART17] Ahmed F. Al-Refaie and Jonathan Tennyson. A parallel algorithm for Hamiltonian matrix construction in electron-molecule collision calculations: MPI– SCATCI. Computer Physics Communications, 221(??): 53–62, December 2017. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465517302436.

Addison:2003:OIA

[ARvW03] C. Addison, Y. Ren, and M. van Waveren. OpenMP issues arising in the development of parallel BLAS and LAPACK libraries. Scientific Programming, 11(2): 95–104, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Al-Refaie:2017:PCT

[ARYT17] Ahmed F. Al-Refaie, Sergei N. Yurchenko, and Jonathan Tennyson. GPU Accelerated INtensities MPI (GAIN-MPI): a new method of computing Einstein-A coefficients. Computer Physics Communications, 214(??): 216–224, May 2017. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465517300255.

Al-Salman:1992:DIP

[AS92] Abdulmalik Salman Al-Salman. Design and implementation of a profiler for the parallel virtual machine (PVM) system. M.s. thesis, University of Georgia, Athens, GA, USA, 1992. vi + 51 pp. Directed by Steven C. Cater.

Awile:2014:PWF

[AS14] Omar Awile and Ivo F.

Sbalzarini. A Pthreads wrapper for Fortran 2003. ACM Transactions on Mathematical Software, 40(3): 19:1–19:15, April 2014. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Alonso:1997:PBB

[ASA97] J. L. Alonso, H. Schmidt, and V. N. Alexandrov. Parallel branch and bound algorithms for integer and mixed integer linear programming problems under PVM. Lecture Notes in Computer Science, 1332:313–320, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Al-Shorman:2019:UPP

[ASAK19] Mohammad Y. Al-Shorman and Majd M. Al-Kofahi. Ultrasonic pulse propagation simulation using OpenCL for environment mapping and discovery. The International Journal of High Performance Computing Applications, 33(5):1019–1029, September 1, 2019. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:

//journals.sagepub.com/
doi/full/10.1177/1094342019846290.

Aydin:2018:RTP

[ASB18] Semra Aydin, Refik Samet, and Omer Faruk Bay. Realtime parallel image processing applications on multicore CPUs with OpenMP and GPGPU with CUDA. The Journal of Supercomputing, 74(6):2255–2275, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alves:1995:WPC

[ASCS95] A. Alves, L. Silva, J. Carreira, and J. G. Silva. WPVM: parallel computing for the people. In Hertzberger and Serazzi [HS95a], pages 582–587. ISBN 3-540-59393-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Anderson:2017:BGB

[ASS^{+}17] Michael Anderson, Shaden Smith, Narayanan Sundaram, Mihai Capota, Zheguang Zhao, Subramanya Dulloor, Nadathur Satish, and Theodore L. Willke. Bridging the gap between HPC and big data frameworks. Proceedings of the VLDB Endowment, 10(8):901–912, April 2017. CODEN ???? ISSN 2150-8097.

Agrawal:1994:PIC

[ATC94] Dharma P. Agrawal, K. C.

(Kuo Chung) Tai, and Jagdish Chandra, editors. Proceedings of the 1994 International Conference on Parallel Processing, August 15–19, 1994. Vol 3: Algorithms and applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1994. ISBN 0-8493-2496-3, 0-8493-2495-5. ISSN 0190-3918. LCCN QA 76.58 I55 1994. Three volumes.

Amritkar:2012:OPF

[ATL^{+}12] Amit Amritkar, Danesh Tafti, Rui Liu, Rick Kufrin, and Barbara Chapman. OpenMP parallelism for fluid and fluid-particulate systems. Parallel Computing, 38(9):501–517, September 2012. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167819112000476.

Al-Tawil:2001:PME

[ATM01] Khalid Al-Tawil and Csaba Andras Moritz. Performance modeling and evaluation of MPI. Journal of Parallel and Distributed Computing, 61(2):202–223, February 1, 2001. CODEN JPDCER.

ISSN 0743-7315 (print), 1096-0848 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        jpdc.2000.1677;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/jpdc.
                                        2000.1677/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/jpdc.
                                        2000.1677/ref.

Attiya:1996:ERS

[Att96] H. Attiya. Efficient and robust sharing of memory in message-passing systems. Lecture Notes in Computer Science, 1151:56– ??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Angskun:2001:DPM

[AUR01] Thara Angskun, Putchong Uthayopas, and Arnon Rungsawang. Dynamic process management in KSIX cluster middleware. Lecture Notes in Computer Science, 2131:209–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2131/21310209.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2131/21310209.
                                        pdf.

Arif:2018:RBP

[AV18] Mahwish Arif and Hans Vandierendonck. Reducing the burden of parallel loop schedulers for manycore processors. ACM SIG-PLAN Notices, 53(1):383– 384, January 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andujar:2016:OSF

[AVA^{+}16] Francisco J. Andújar, Juan A.

Villar, Francisco J. Alfaro, José L. Sánchez, and Jesus Escudero-Sahuquillo. An open-source family of tools to reproduce MPI-based workloads in interconnection network simulators. The Journal of Supercomputing, 72(12):4601–4628, December 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Asenjo:1995:SLF

[AZ95] R. Asenjo and E. L. Zapata. Sparse LU factorization of the Cray T3D. In Hertzberger and Serazzi [HS95a], pages 690–696. ISBN 3-540-59393-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Arteaga:2017:GFG

[AZG17] Jaime Arteaga, Stéphane Zuckerman, and Guang R. Gao. Generating fine-grain multithreaded applications using a multigrain approach. ACM Transactions on Architecture and Code Optimization, 14(4):47:1–47:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Beyer:2005:GEC

[B^{+}05] Hans-Georg Beyer et al., editors. Genetic and Evolutionary Computation Conference: GECCO 2005, June 25–29, 2005 (Saturday-Wednesday) Washington, DC, USA. ACM Press, New York, NY 10036, USA, 2005. ISBN 1-59593-010-8 (paperback). LCCN QA76.623 .G44 2005. ACM order number 910050.

Battre:2006:MFP

[BA06] Dominic Battré and David Sigfredo Angulo. MPI framework for parallel searching in large biological databases. Journal of Parallel and Distributed Computing, 66(12): 1503–1511, December 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Barreda:2020:IFC

[BAC20] María Barreda, José I. Aliaga, and Marc Casas. Iteration-fusing conjugate gradient for sparse linear systems with MPI + OmpSs. The Journal of Supercomputing, 76(9):6669– 6689, September 2020. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

https://link.springer.
com/article/10.1007/s11227-
019-03100-4.

Bader:2016:EMT

[Bad16] David A. Bader. Evolving MPI+X toward exascale. Computer, 49(8):10, August 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
co/2016/08/mco2016080010.
html.

Becciani:2007:FMH

[BADC07] U. Becciani, V. Antonuccio-Delogu, and M. Comparato. FLY: MPI-2 high resolution code for LSS cosmological simulations. Computer Physics Communications, 176(3):211–217, February 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465506003687.

Baek:2020:ESO

[Bae20] Nakhoon Baek. An emulation scheme for OpenGL SC 2.0 over OpenGL. The Journal of Supercomputing, 76(10):7951–7960, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s11227-
                                        018-2399-1.

Barai:2022:PMP

[BAE22] Atanu Barai, Yehia Arafa, and Stephan Eidenbenz.

                                        PPT-Multicore:  perfor-

mance prediction of OpenMP applications using reuse profiles and analytical modeling. The Journal of Supercomputing, 78(2):2354– 2385, February 2022. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s11227-
                                        021-03949-4.

Bruel:2017:ACC

[BAG17] Pedro Bruel, Marcos Amarís, and Alfredo Goldman. Autotuning CUDA compiler parameters for heterogeneous applications using the Open-Tuner framework. Concurrency and Computation: Practice and Experience, 29 (22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Baker:1998:MNC

[Bak98] M. Baker. MPI on NT: The current status and performance of the available environments. Lecture Notes in Computer Science, 1497:63– ??, 1998. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Blaszczyk:1995:PCE

[BALU95] A. Blaszczyk, Z. Andjelic, P. Levin, and A. Ustundag. Parallel computation of electric fields in a heterogeneous workstation cluster. In Hertzberger and Serazzi [HS95a], pages 606–611. ISBN 3-540-59393-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Buyukkececi:2013:POI

[BAS13] Ferit Büyükkeçeci, Omar Awile, and Ivo F. Sbalzarini. A portable OpenCL implementation of generic particle-mesh and meshparticle interpolation in 2D and 3D. Parallel Computing, 39(2):94–111, February 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167819112000920.

Bernabeu:2008:MPA

[BAV08] Miguel O. Bernabeu, Pedro Alonso, and Antonio M. Vidal. A multilevel parallel algorithm to solve symmetric Toeplitz linear systems. The Journal of Supercomputing, 44(3):237–256, June 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0920-8542&volume=
                                        44&issue=3&spage=237.

Bedrosian:1993:MFA

[BB93] G. Bedrosian and R. W. Benway. Magnetostatic finite-element analysis on MIMD/DMMP parallel computers. In Yelon et al. [Y^{+}93], pages 6772–6777. CODEN JAPIAU. ISBN 1-56396-212-8. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. LCCN QC753 .C748 1990. Two volumes.

Beguelin:1994:CMS

[BB94] A. Beguelin and B. Bruegge. A configurable monitoring system for parallel programming. In IEEE [IEE94d], page 206. ISBN 0-8186-5390-6. LCCN QA76.9.D5I595 1994. IEEE catalog no. 94TH0651-0.

Beaumont:1995:DPG

[BB95a] P. M. Beaumont and P. T. Bradshaw. A distributed parallel genetic algorithm for solving optimal growth models. Computational Economics, 8(3):159–179, August 1995. CODEN CNOMEL. ISSN 0927-7099.

Bunge:1995:MCM

[BB95b] Hans-Peter Bunge and John R. Baumgardner. Mantle convection modeling on parallel virtual machines. Computers in Physics, 9(2):207– ??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.168525.

Brunschen:2000:OCP

[BB00] Christian Brunschen and Mats Brorsson. OdinMP/ CCp — a portable implementation of OpenMP for C. Concurrency: practice and experience, 12(12):1193– 1203, October 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76500347/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500347&PLACEBO=IE.
pdf.

Bylina:2018:EEO

[BB18] Beata Bylina and Jaroslaw Bylina. An experimental evaluation of the OpenMP thread mapping for LU factorisation on Xeon Phi coprocessor and on hybrid CPU-MIC platform. Scalable Computing: Practice and Experience, 19(3):259– 274, ???? 2018. CO-DEN ???? ISSN 1895-

1767.  URL https://
www.scpe.org/index.php/
scpe/article/view/1373.

Bala:1994:IEU

[BBB^{+}94] V. Bala, J. Bruck, R. Bryant, R. Cypher, and P. De Jong. The IBM external user interface for scalable parallel systems. Parallel Computing, 20(4):445–??, April 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bernholdt:2020:SMU

[BBB^{+}20] David E. Bernholdt, Swen Boehm, George Bosilca, Manjunath Gorentla Venkata, Ryan E. Grant, Thomas Naughton, Howard P. Pritchard, Martin Schulz, and Geoffroy R. Vallee. A survey of MPI usage in the US exascale computing project. Concurrency and Computation: Practice and Experience, 32(3):e4851:1– e4851:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bak:2022:OAE

[BBB^{+}22] Seonmyeong Bak, Colleen Bertoni, Swen Boehm, Reuben Budiardja, Barbara M. Chapman, Johannes Doerfert, Markus Eisenbach, Hal Finkel, Oscar Hernandez, Joseph Huber, Shintaro Iwasaki, Vivek Kale, Paul R. C. Kent, JaeHyuk Kwack, Meifeng Lin, Piotr Luszczek, Ye Luo, Buu Pham, Swaroop Pophale, Kiran Ravikumar, Vivek Sarkar, Thomas Scogland, Shilei Tian, and P. K. Yeung.

OpenMP application experiences: Porting to accelerated nodes. Parallel Computing, 109(??):??, March 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819121001009.

Bova:1999:NOM

[BBC^{+}99] S. W. Bova, C. P. Breshears, C. Cuicchi, Z. Demirbilek, and H. Gabb. Nesting OpenMP in an MPI application. In ????, editor, Proceedings of the ISCA 12th International Conference. Parallel and Distributed Systems, pages 566–571. ISCA, Raleigh, NC, USA, 1999.

Bova:2000:DLP

[BBC^{+}00] Steve W. Bova, Clay P. Breshears, Christine E. Cuicchi, Zeki Demirbilek, and Henry A. Gabb. Dual-level parallel analysis of harbor wave response using MPI and OpenMP. The International Journal of High Performance Computing Applications, 14(1):49–64, Spring 2000. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Bosilca:2002:MVT

[BBC^{+}02] George Bosilca, Aurelien Bouteiller, Franck Cappello, Samir Djilali, Gilles Fedak, Cecile Germain, Thomas Herault, Pierre Lemarinier, Oleg Lodygensky, Frederic Magniette, Vincent Neri, and Anton Selikhov. MPICH-V: Toward a scalable fault tolerant MPI for volatile nodes. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN

                                        ???? URL http://www.sc-
                                        2002.org/paperpdfs/pap.
                                        pap298.pdf.

Badia:2019:ASP

[BBC^{+}19] Jose M. Badía, Jose A. Belloch, Maximo Cobos, Francisco D. Igual, and Enrique S. Quintana-Ortí. Accelerating the SRP–PHAT algorithm on multi- and many-core platforms using OpenCL. The Journal of Supercomputing, 75(3):1284– 1297, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bertozzi:1999:MIT

[BBCR99] M. Bertozzi, F. Boselli, G. Conte, and M. Reggiani. An MPI implementation on the top of the virtual interface architecture. In Dongarra et al. [DLM99], pages 199–206. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Bombieri:2020:MIB

[BBD^{+}20] N. Bombieri, F. Busato, A. Danese, L. Piccolboni, and G. Pravadelli. Mangrove: An inference-based dynamic invariant mining for GPU architectures. IEEE Transactions on Computers, 69(4):606–620, April 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bethune:2014:PAA

[BBDH14] Iain Bethune, J. Mark Bull, Nicholas J. Dingle, and Nicholas J. Higham. Performance analysis of asynchronous Jacobi’s method implemented in MPI, SHMEM and OpenMP. The International Journal of High Performance Computing Applications, 28(1):97–111, February 2014. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/28/
1/97.full.pdf+html.

Bailey:1995:PSS

[BBG^{+}95] D. H. Bailey, P. E. Bjorstad, J. R. Gilbert, M. V. Mascagni, R. S. Schreiber, H. D. Simon, V. J. Torczon, and L. T. Watson, editors. Proceedings of the Seventh SIAM Conference on Parallel Processing for Scientific Computing (San Francisco, CA, USA). Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1995. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Bova:1999:PPM

[BBG^{+}99] Steve W. Bova, Clay P. Breshears, Henry Gabb, Rudolf Eigenmann, Greg Gaertner, Bob Kuhn, Bill Magro, and Stefano Salvini. Parallel programming with message passing and directives. SIAM News, 32(9):??, November 1999. ISSN 0036-1437.

Bova:2001:PPM

[BBG^{+}01] Steve W. Bova, Clay P. Breshears, Henry Gabb, Bob Kuhn, Bill Magro, Rudolf Eigenmann, Greg Gaertner, Stefano Salvini, and Howard Scott. Parallel programming with message passing and directives. Computing in Science and Engineering, 3(5):22–37, September/October 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c5022abs.htm;   http:/
                                        /dlib.computer.org/cs/
                                        books/cs2001/pdf/c5022.
                                        pdf.

Balaji:2010:FGM

[BBG^{+}10] Pavan Balaji, Darius Buntinas, David Goodell, William Gropp, and Rajeev Thakur. Fine-grained multithreading support for hybrid threaded MPI programming. The International Journal of High Performance Computing Applications, 24(1):49– 57, February 2010. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/24/
1/49.full.pdf+html.

Balaji:2011:MMC

[BBG^{+}11] Pavan Balaji, Darius Buntinas, David Goodell, William Gropp, Torsten Hoefler, Sameer Kumar, Ewing Lusk, Rajeev Thakur, and Jesper Larsson Träff. MPI on millions of cores. Parallel Processing Letters, 21(1):45– 60, March 2011. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Barrett:2014:EMM

[BBG^{+}14] Brian W. Barrett, Ron Brightwell, Ryan Grant, Simon D. Hammond, and K. Scott Hemmert. An evaluation of MPI message rate on hybrid-core processors. The International Journal of High Performance Computing Applications, 28(4): 415–424, November 2014. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

http://hpc.sagepub.com/
content/28/4/415.

Barak:1996:PPM

[BBGL96] A. Barak, A. Braverman, I. Gilderman, and O. Laden.

Performance of PVM with the MOSIX preemptive process migration scheme. In IEEE [IEE96h], pages 38–45. ISBN 0-8186-7536-5. LCCN QA75.5 .I75 1996. IEEE Computer Society Press Order Number PR07536.

Bouteiller:2006:HPS

[BBH^{+}06] Aurélien Bouteiller, Hinde-Lilia Bouziane, Thomas Herault, Pierre Lemarinier, and Franck Cappello. Hybrid preemptive scheduling of Message Passing Interface applications on Grids. The International Journal of High Performance Computing Applications, 20(1):77– 90, Spring 2006. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/20/
1/77.full.pdf+html.

Bischof:2008:AAD

[BBH^{+}08] Christian H. Bischof, H. Martin Bücker, Paul Hovland, Uwe Naumann, and Jean Utke, editors. Advances in Automatic Differentiation, volume 64 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CO-DEN LNCSA6. ISBN 3-540-68935-4 (print), 3-540-68942-7 (e-book). ISSN 1439-7358. LCCN QA304 .I58 2008. URL http://

link.springer.com/book/
10.1007/978-3-540-68942-
3;   http://www.springerlink.
com/content/978-3-540-
68942-3.

Bustamam:2012:FPM

[BBH12] Alhadi Bustamam, Kevin Burrage, and Nicholas A. Hamilton. Fast parallel Markov clustering in bioinformatics using massively parallel computing on GPU with CUDA and ELLPACK-R sparse format. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 9(3):679–692, May 2012. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Bland:2013:EUL

[BBH… 13a] Wesley Bland, Aurelien Bouteiller, Thomas Herault, and Joshua Hursey …. An evaluation of User-Level Failure Mitigation support in MPI. Computing, 95(12):1171–1184, December 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00607-013-0331-3.

Bland:2013:PFR

[BBH^{+}13b] Wesley Bland, Aurelien Bouteiller, Thomas Herault, George Bosilca, and Jack Dongarra. Post-failure recovery of MPI communication capability: Design and rationale. The International Journal of High Performance Computing Applications, 27(3):244–254, August 2013. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/27/
3/244.full.pdf+html.

Busa:2015:CCO

[BBH^{+}15] Ján Busa, Jr., Ján Busa, Shura Hayryan, Chin-Kun Hu, and Ming-Chya Wu. CAVE-CL: an OpenCL version of the package for detection and quantitative analysis of internal cavities in a system of overlapping balls: Application to proteins. Computer Physics Communications, 190(??): 224–227, May 2015. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465514004378.

Boryczko:1994:LGA

[BBK^{+}94] K. Boryczko, M. Bubak, J. Kitowski, J. Moscinski, and R. Slota. Lattice gas automata and molecular dynamics on a network of computers. In Gentzsch and Harms [GH94], pages 177–180. ISBN 0-387-57981-8 (New York), 3-540-57981-8 (Berlin). LCCN QA76.88.I57 1994. DM96.00. Two volumes.

Barnard:1999:MIS

[BBS99] Stephen T. Barnard, Luis M.

Bernardo, and Horst D. Simon. An MPI implementation of the SPAI preconditioner on the T3E. The International Journal of High Performance Computing Applications, 13(2):107– 123, Summer 1999. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Brown:2019:LMR

[BBW19] Nick Brown, Michael Bareford, and Michèle Weiland. Leveraging MPI RMA to optimize halo-swapping communications in MONC on Cray machines. Concurrency and Computation: Practice and Experience, 31 (16):e5008:1–e5008:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brorsson:2000:SIE

[BC00] Mats Brorsson and Barbara Chapman. Special issue: EWOMP’99 — First European Workshop on OpenMP. Concurrency: practice and experience, 12(12):1117– 1119, October 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/

                                        cgi-bin/abstract/76500352/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=76500352&PLACEBO=IE.
                                        pdf.

Blas:2014:RAM

[BC14] Javier Garcia Blas and Jesus Carretero. Recent advances in the Message Passing Interface. The International Journal of High Performance Computing Applications, 28(4):387–389, November 2014. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

                                        tronic). URL http://hpc.
                                        sagepub.com/content/28/
                                        4/387.

Balaji:2019:SIM

[BC19a] Pavan Balaji and Marc Casas. Special issue on the message passing interface. Parallel Computing, 86(??):14–15, August 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016781911930095X.

Budiardja:2019:TGO

[BC19b] Reuben D. Budiardja and Christian Y. Cardall. Targeting GPUs with OpenMP directives on Summit: a simple and effective Fortran experience. Parallel Computing, 88(??):Article 102544, ???? 2019. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819119301358.

Barton:2006:SMP

[BCA^{+}06] Christopher Barton, Călin Casçaval, George Almási, Yili Zheng, Montse Farreras, Siddhartha Chatterje, and José Nelson Amaral. Shared memory programming for large scale machines. ACM SIGPLAN Notices, 41(6):108–117, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Becciani:2006:FMP

[BCAD06] U. Becciani, M. Comparato, and V. Antonuccio-Delogu. FLY MPI-2: a parallel tree code for LSS. Computer Physics Communications, 174(7):605–606, April 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465506000713.

Bircsak:2000:EONa

[BCC^{+}00a] John Bircsak, Peter Craig, RaeLyn Crowell, Zarka Cvetanovic, Jonathan Harris, C. Alexander Nelson, and Carl D. Offner. Extending OpenMP for NUMA machines. In ACM [ACM00], pages 68–

                                        69.  URL http://www.
                                        sc2000.org/proceedings/
                                        techpapr/papers/pap226.
                                        pdf.

Bircsak:2000:EONb

[BCC^{+}00b] John Bircsak, Peter Craig, RaeLyn Crowell, et al. Extending OpenMP for NUMA machines. Scientific Programming, 8(3):163–181, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bouchard:1996:FCS

[BCD96] V. Bouchard, P. Cinquin, and L. Desbat. First Compton scatter correction in SPECT using PVM. In Grangeat and Amans [GA96], pages 109–111. ISBN 0-7923-4129-5. LCCN R857.T47 T485 1996.

Betts:2012:GVG

[BCD^{+}12] Adam Betts, Nathan Chong, Alastair Donaldson, Shaz Qadeer, and Paul Thomson. GPUVerify: a verifier for GPU kernels. ACM SIG-PLAN Notices, 47(10):113– 132, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Betts:2015:DIV

[BCD^{+}15] Adam Betts, Nathan Chong, Alastair F. Donaldson, Jeroen Ketema, Shaz Qadeer, Paul Thomson, and John Wickerson. The design and implementation of a verification technique for GPU kernels. ACM Transactions on Programming Languages and Systems, 37 (3):10:1–10:??, June 2015. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Baker:1999:MOO

[BCFK99] M. Baker, B. Carpenter, G. Fox, and Sung Hoon Koo. mpiJava: An object-oriented Java interface to MPI. Lecture Notes in Computer Science, 1586:748–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balaji:2010:IND

[BCG^{+}10] Pavan Balaji, Anthony Chan, William Gropp, Rajeev Thakur, and Ewing Lusk. The importance of non-data-communication overheads in MPI. The International Journal of High Performance Computing Applications, 24(1):5–15, February 2010. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/24/
1/5.full.pdf+html.

Bala:1997:PVQ

[BCGL97] P. Bala, T. Clark, P. Grochowski, and B. Lesyng. Parallel version of a quantum classical molecular dynamics code for complex molecular and biomolecular systems. Lecture Notes in Computer Science, 1332:409–416, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouteiller:2003:MVF

[BCH^{+}03] Aurelien Bouteiller, Franck Cappello, Thomas Herault, Geraud Krawezik, Pierre Lemarinier, and Frederic Magniette. MPICH-V2: a fault tolerant MPI for volatile nodes based on pessimistic sender based message logging. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://
www.sc-conference.org/
sc2003/inter_cal/inter_
cal_detail.php?eventid=
10696#1;   http://www.
sc-conference.org/sc2003/
paperpdfs/pap209.pdf.

Buntinas:2008:BVN

[BCH^{+}08] Darius Buntinas, Camille Coti, Thomas Herault, Pierre Lemarinier, Laurence Pilard, Ala Rezmerita, Eric Rodriguez, and Franck Cappello. Blocking vs. nonblocking coordinated checkpointing for large-scale fault tolerant MPI protocols. Future Generation Computer Systems, 24(1):73–84, January 2008. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bikshandi:2009:EPI

[BCK^{+}09] Ganesh Bikshandi, Jose G.

Castanos, Sreedhar B. Kodali, V. Krishna Nandivada, Igor Peshansky, Vijay A. Saraswat, Sayantan Sur, Pradeep Varma, and Tong Wen. Efficient, portable implementation of asynchronous multiplace programs. ACM SIG-PLAN Notices, 44(4):271– 282, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruno:2000:PEH

[BCKP00] G. Bruno, A. A. Chien, M. J. Katz, and P. M. Papadopoulos. Performance enhancements for HPVM in multi-network and heterogeneous hardware. In Engquist [Eng00], pages 17–32. ISBN 3-540-67264-8. ISSN 1439-7358. LCCN QA76.9.C65 S535 2000.

Bolloni:2000:TIQ

[BCL00] Alessandro Bolloni, Stefano Crocchianti, and Antonio Laganà. Time independent 3D quantum reactive scattering on MIMD parallel computers. Lecture Notes in Computer Science, 1908:338–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1908/19080338.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1908/19080338.
                                        pdf.

Baraglia:1997:IPW

[BCLN97] R. Baraglia, M. Cosso, D. Laforenza, and M. Nicosia. Integrating PVaniM into WAMM for monitoring meta-applications. Lecture Notes in Computer Science, 1332:226–233, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhattacharjee:2011:PLC

[BCM11] Abhishek Bhattacharjee, Gilberto Contreras, and Margaret Martonosi. Parallelization libraries: Characterizing and reducing overheads. ACM Transactions on Architecture and Code Optimization, 8(1):5:1–5:??, April 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bolis:2016:APA

[BCM^{+}16] A. Bolis, C. D. Cantwell, D. Moxey, D. Serson, and S. J. Sherwin. An adaptable parallel algorithm for the direct numerical simulation of incompressible turbulent flows using a Fourier spectral/hp element method and MPI virtual topologies. Computer Physics Communications, 206(??): 17–25, September 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S001046551630100X.

Baiardi:2000:AMM

[BCMR00] Fabrizio Baiardi, Sarah Chiti, Paolo Mori, and Laura Ricci. Adaptive multigrid methods in MPI. Lecture Notes in Computer Science, 1908:80–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080080.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080080.
pdf.

Blackford:1997:PEN

[BCP^{+}97] L. S. Blackford, A. Cleary, A. Petitet, R. C. Whaley, J. Demmel, I. Dhillon, H. Ren, K. Stanley, J. Dongarra, and S. Hammarling. Practical experience in the numerical dangers of heterogeneous computing. ACM Transactions on Mathematical Software, 23(2):133– 147, June 1997. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.

                                        acm.org/pubs/citations/
                                        journals/toms/1997-23-
                                        2/p133-blackford/.

Burtscher:2018:HQF

[BDA^{+}18] Martin Burtscher, Sindhu Devale, Sahar Azimi, Jayadharini Jaiganesh, and Evan Powers. A high-quality and fast maximal independent set implementation for GPUs. ACM Transactions on Parallel Computing (TOPC), 5(2):8:1– 8:??, January 2018. CO-DEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Bland:2013:SIP

[BDB^{+}13] Wesley Bland, Peng Du, Aurelien Bouteiller, Thomas Herault, George Bosilca, and Jack J. Dongarra. Special issue papers: Extending the scope of the Checkpoint-on-Failure protocol for forward recovery in standard MPI. Concurrency and Computation: Practice and Experience, 25(17):2381–2393, December 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Beguelin:1991:UGP

[BDG^{+}91a] A. Beguelin, J. Dongarra, A. Geist, R. Manchek, and V. Sunderam. A user’s guide to PVM: Parallel virtual machine. Technical Report ORNL/TM-11826, Mathematical Sciences Section, Oak Ridge National Laboratory, Knoxville, TN, USA, September 1991.

Beguelin:1991:GDT

[BDG^{+}91b] Adam Beguelin, Jack J.

Dongarra, A. Geist, Robert Manchek, and V. S. Sunderam. Graphical development tools for network-based concurrent supercomputing. In IEEE [IEE91], pages 435– 444. ISBN 0-8186-9158-1 (IEEE: case), 0-8186-2158-3 (IEEE: paper), 0-8186-6158-5 (IEEE: microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. IEEE catalog no. 91CH3058-5.

Beguelin:1992:HGD

[BDG^{+}92a] A. Beguelin, J. Dongarra, A. Geist, R. Manchek, K. Moore, R. Wade, and V. Sunderam. HeNCE: graphical development tools for network-based concurrent computing. In IEEE [IEE92], pages 129– 136. ISBN 0-8186-2775-1. LCCN QA76.76.A65S33 1992. IEEE catalog no. 92TH0432-5.

Beguelin:1992:PHT

[BDG^{+}92b] A. Beguelin, J. Dongarra, A. Geist, R. Manchek, and V. Sunderam. PVM and HeNCE: traversing the parallel environment. CRAY Channels, 14(4):22–25, Fall 1992. CODEN CRCHE8.

Beguelin:1992:SCG

[BDG^{+}92c] A. Beguelin, J. Dongarra, A. Geist, R. Manchek, and V. Sunderam. Solving computational grand challenges using a network of heterogeneous supercomputers. In Dongarra et al. [DKM^{+}92], pages 596–601. ISBN 0-89871-303-X. LCCN QA76.58.P76 1992.

Beguelin:1993:PHT

[BDG^{+}93a] A. Beguelin, J. Dongarra, A. Geist, R. Manchek, K. Moore, and V. Sunderam. PVM and HeNCE: Tools for heterogeneous network computing. In Kowalik and Grandinetti [KG93], page ?? ISBN 3-540-56451-9 (Berlin), 0-387-56451-9 (New York). LCCN QA76.58 .S629 1993.

Beguelin:1993:PEC

[BDG^{+}93b] A. Beguelin, J. Dongarra, A. Geist, R. Manchek, S. Otto, and J. Walpole. PVM: Experiences, current status and future direction. In IEEE [IEE93e], pages 765–766. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Beguelin:1994:HHN

[BDG^{+}94] A. Beguelin, J. J. Dongarra, G. Al Geist, R. Manchek, and K. Moore. HeNCE: a heterogeneous network computing environment. Scientific Programming, 3(1):49– 60, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Beguelin:1995:REP

[BDG^{+}95] Adam Beguelin, Jack Dongarra, Al Geist, Robert Manchek, and Vaidy Sunderam. Recent enhancements to PVM. International Journal of Supercomputer Applications and High Performance Computing, 9 (2):108–127, Summer 1995. CODEN IJSCFG. ISSN 1078-3482.

Beguelin:19xx:PSS

[BDG^{+}xx] A. Beguelin, J. J. Dongarra, G. A. Geist, R. Manchek, and V. S. Sunderam. PVM software system and documentation. Email to

netlib@ornl.gov, ???? 19xx.

Beguelin:1993:VDH

[BDGS93] Adam Beguelin, Jack Dongarra, Al Geist, and V. Sunderam. Visualization and debugging in a heterogeneous environment. Computer, 26 (6):88–95, June 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bruck:1995:EMPb

[BDH^{+}95] Jehoshua Bruck, Danny Dolev, Ching-Tien Ho, Marcel-Catalin Rosu, and Ray Strong. Efficient Message Passing Interface (MPI) for parallel computing on clusters of workstations. In ACM [ACM95b], pages 64–

73. ISBN 0-89791-717-0. LCCN QA76.642 .A25 1995.

Bruck:1997:EMP

[BDH^{+}97] Jehoshua Bruck, Danny Dolev, Ching-Tien Ho, Marcel-Cătălin Roşu, and Ray Strong. Efficient message passing interface (MPI) for parallel computing on clusters of workstations. Journal of Parallel and Distributed Computing, 40(1): 19–34, January 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.idealibrary.com/
links/doi/10.1006/jpdc.
1996.1267/production;
http://www.idealibrary.
com/links/doi/10.1006/
jpdc.1996.1267/production/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/
jpdc.1996.1267/production/
ref.

Browne:1998:RPA

[BDL98] Shirley Browne, Jack Dongarra, and Kevin London. Review of performance analysis tools for MPI parallel programs. NHSE Review, 3, 1998. CODEN ???? ISSN ???? URL http://

www.cs.utk.edu/~browne/

perftools-review/. Ac-

cepted, to appear.

Bode:1996:PVM

[BDLS96] Arndt Bode, Jack Dongarra, T. Ludwig, and V. Sunderam, editors. Parallel virtual machine, EuroPVM ’96: third European PVM conference, Munich, Germany, October 7–9, 1996: proceedings, volume 1156 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Baghsorkhi:2010:APM

[BDP^{+}10] Sara S. Baghsorkhi, Matthieu Delahaye, Sanjay J. Patel, William D. Gropp, and Wen mei W. Hwu. An adaptive performance modeling tool for GPU architectures. ACM SIGPLAN Notices, 45(5): 105–114, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bronevetsky:2007:CFS

[BdS07] Greg Bronevetsky and Bronis R. de Supinski. Complete formal specification of the OpenMP memory model. International Journal of Parallel Programming, 35(4):335–392, August 2007.

CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        35&issue=4&spage=335.

Baboulin:2008:SID

[BDT08] Marc Baboulin, Jack J. Dongarra, and Stanimire Tomov. Some issues in dense linear algebra for multicore and special purpose architectures. LAPACK Working Note 200, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May 2008. URL http:/

                                        /www.netlib.org/lapack/
                                        lawnspdf/lawn200.pdf.

Briguglio:2003:PPM

[BDV03] Sergio Briguglio, Beniamino Di Martino, and Gregorio Vlad. A performanceprediction model for PIC applications on clusters of symmetric multiprocessors: Validation with hierarchical HPF + OpenMP implementation. Scientific Programming, 11(2):159–176, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bubak:1997:RAP

[BDW97] Marian Bubak, J. J. Dongarra, and Jerzy Wasniewski, editors. Recent advances in parallel virtual machine and message passing interface: 4th European PVM/MPI user’s group meeting Cracow, Poland, November 3–5, 1997: proceedings, volume 1332 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CO-DEN LNCSD9. ISBN 3-540-63697-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E973 1997.

Batty:2016:OSA

[BDW16] Mark Batty, Alastair F.

Donaldson, and John Wickerson. Overhauling SC atomics in C11 and OpenCL. ACM SIGPLAN Notices, 51 (1):634–648, January 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beyls:1999:JJP

[BDY99] K. Beyls, E. D’Hollander, and Y. Yu. JPT: a Java parallelization tool. In Dongarra et al. [DLM99], pages 173–180. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Beguelin:1992:XTM

[Beg92] Adam Louis Beguelin. Xab:

a tool for monitoring PVM programs. Technical report, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, June 5, 1992.

Beguelin:1993:XTMb

[Beg93a] A. L. Beguelin. Xab: a tool for monitoring PVM programs. In Mudge et al. [MMH93], pages 102–103 (vol. 2) (or 4–??). ISBN 0-8186-3230-5. LCCN ???? Four volumes. IEEE catalog number 93TH0501-7.

Beguelin:1993:XAT

[Beg93b] Adam Beguelin. Xab: a tool for monitoring PVM programs. In IEEE [IEE93f], pages 92–97. ISBN 0-8186-2702-6. LCCN QA76.58 .W654 1992.

Beguelin:1993:XTMa

[Beg93c] Adam L. Beguelin. Xab: a tool for monitoring PVM programs. Research paper CMU-CS-93-164, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, 1993. 8 pp.

Bull:2010:PEM

[BEG^{+}10] J. Mark Bull, James Enright, Xu Guo, Chris Maynard, and Fiona Reid. Performance evaluation of mixed-mode OpenMP/MPI implementations. International Journal of Parallel Programming, 38(5– 6):396–417, October 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=5&spage=396.

Benkner:1995:VFA

[Ben95] S. Benkner. Vienna Fortran 90 — an advanced data parallel language. In Malyshkin [Mal95], pages 142–156. ISBN 3-540-60222-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I547 1995.

Bencheva:2001:MPI

[Ben01] G. Bencheva. MPI parallel implementation of a fast separable solver. Lecture Notes in Computer Science, 2179:454–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2179/21790454.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2179/21790454.
pdf.

Benedict:2018:SES

[Ben18] Shajulin Benedict. SCALE-EA: A scalability aware performance tuning framework for OpenMP applications. Scalable Computing: Practice and Experience, 19(1):15–30, ???? 2018. CODEN ???? ISSN 1895-1767. URL https://

                                        www.scpe.org/index.php/
                                        scpe/article/view/1390.

Bernaschi:1996:RHP

[Ber96] Massimo Bernaschi. The requirements of a high performance implementation of PVM. Future Generation Computer Systems, 12(1): 3–11, May 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Baker:1998:MNP

[BF98] M. Baker and G. Fox. MPI on NT: a preliminary evaluation of the available environments. Lecture Notes in Computer Science, 1388: 549–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berthou:2001:COH

[BF01] Jean-Yves Berthou and Eric Fayolle. Comparing OpenMP, HPF, and MPI programming: a study case. The International Journal of High Performance Computing Applications, 15(3):297– 309, Fall 2001. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Bubak:2001:PMS

[BFBW01] Marian Bubak, Włodzimierz Funika, Bartosz Bali, and Roland Wismüller. Performance measurement support for MPI applications with PATOP. Lecture Notes in Computer Science, 1947:288–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1947/19470288.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1947/19470288.
pdf.

Bischof:1994:CSM

[BfDA94] Christian Bischof and Institute for Defense Analyses. A case study of MPI: portable and efficient libraries. Technical report SRC-TR-94-130, Supercomputing Research Center: IDA, Lanham, MD, USA, 1994. 6 pp.

Broquedis:2010:FEO

[BFG^{+}10] François Broquedis, Nathalie Furmento, Brice Goglin, Pierre-André Wacrenier, and Raymond Namyst. ForestGOMP: An efficient OpenMP environment for NUMA architectures. International Journal of Parallel Programming, 38(5– 6):418–439, October 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/

                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        38&issue=5&spage=418.

Bubak:1999:EFP

[BFIM99] M. Bubak, W. Funika, K. Iskra, and R. Maruszewski. Enhancing the functionality of performance measurement tools for message passing environments. In Dongarra et al. [DLM99], pages 67–74. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Baraglia:1999:PAN

[BFLL99] R. Baraglia, R. Ferrini, D. Laforenza, and A. Lagana. Parallel approaches to a numerically intensive application using PVM. In Dongarra et al. [DLM99], pages 364–371. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Bubak:1996:MPP

[BFM96] M. Bubak, W. Funika, and J. Moscinski. Monitoring of performance of PVM applications on virtual network computer. In Wasniewski [Was96], pages 147– 156. ISBN 3-540-62095-8. LCCN QA76.58 .P35 1996.

Bubak:1997:EPA

[BFM97] M. Bubak, W. Funika, and J. Moscinski. Evaluation of parallel application’s behavior in message passing environment. Lecture Notes in Computer Science, 1332: 234–241, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouge:1996:EPP

[BFMR96] Luc Bouge, P. Fraigniaud, A. Mignotte, and Y. Robert, editors. Euro-Par ’96 parallel processing: second International Euro-Par Conference, Lyon, France, August 26–29, 1996: proceedings, volume 1123–1124 of Lecture notes in computer science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61626-8 (vol. 1), 3-540-61627-6 (vol. 2). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I554 1996, QA267.A1 L43 no.1123-1124. Two volumes.

Bubak:1996:PBP

[BFMT96a] M. Bubak, W. Funika, J. Moscinski, and D. Tasak. Pablo-based performance monitoring tool for PVM applications. In Dongarra et al. [DMW96], pages 69–78. ISBN 3-540-60902-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.P35 1995.

Bubak:1996:PPM

[BFMT96b] M. Bubak, W. Funika, J. Moscinski, and D. Tasak. Pablo-Based performance monitoring tool for PVM applications. In Dongarra et al. [DMW96], pages 69–78. ISBN 3-540-60902-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.P35 1995.

Bozas:1997:PED

[BFZ97] G. Bozas, M. Fleischhauer, and S. Zimmermann. PVM experiences in developing the MIDAS parallel database system. Lecture Notes in Computer Science, 1332:427–434, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhavsar:1991:SSJ

[BG91] Virendrakumar Chhabulal Bhavsar and Uday Govinddas Gujar, editors. Supercomputing Symposium ’91, June 3–5, 1991, Fredericton, NB, Canada: symposium proceedings. University of New Brunswick Press, Fredericton, NB, Canada, 1991. ISBN 0-920114-14-8. LCCN QA76.88.S87 1991.

Boerger:1994:FSP

[BG94a] E. Boerger and U. Glaesser. A formal specification of the PVM architecture. In Pehrson et al. [PSB^{+}94], pages 402–409. CODEN ITATEC. ISBN 0-444-81990-8, 0-444-81989-4. ISSN 0926-5473. LCCN QA75.5.I3785 1994. Three volumes.

Borger:1994:AMP

[BG94b] E. Borger and U. Glasser.

An abstract model of the Parallel Virtual Machine (PVM). In Anonymous [Ano94e], pages 308–309. ISBN 1-880843-09-9. LCCN QA76.58.I543 1994.

Borger:1994:FSP

[BG94c] E. Borger and U. Glasser.

A formal specification of the PVM architecture. IFIP Transactions. A. Computer Science and Technology, A-51:402–409, ???? 1994. CO-DEN ITATEC. ISSN 0926-5473.

Barbour:1995:PIG

[BG95] A. E. Barbour and M. F.

Gabre. Parallel implementation of Gauss–Seidel and conjugate gradient for solving system of linear equations Ax = b using PVM. In Aityan et al. [AGH^{+}95], pages 33–36. ISBN 0-9640398-9-3 (hardback) 0-9640398-8-5 (paperback). LCCN QA76.87 .I58 1995.

Banikazemi:2001:MLE

[BGBP01] Mohammad Banikazemi, Rama K. Govindaraju, Robert Blackmore, and Dhabaleswar K. Panda. MPI-LAPI: An efficient implementation of MPI for IBM RS/6000 SP systems. IEEE Transactions on Parallel and Distributed Systems, 12(10):1081–1093, October 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://dlib.
computer.org/td/books/
td2001/pdf/l1081.pdf;
http://www.computer.org/
tpds/td2001/l1081abs.htm.

Broquedis:2012:LEO

[BGD12] François Broquedis, Thierry Gautier, and Vincent Danjean. libOMP, an efficient OpenMP runtime system for both fork-join and data flow paradigms. Lecture Notes in Computer Science, 7312: 102–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-30961-8_
8/.

Bronevetsky:2009:CAC

[BGdS09] Greg Bronevetsky, John Gyllenhaal, and Bronis R. de Supinski. CLOMP: Accurately characterizing OpenMP application overheads. International Journal of Parallel Programming, 37(3):250–265, June 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&

issn=0885-7458&volume=
37&issue=3&spage=250.

Blanco:2002:PMA

[BGG^{+}02] V. Blanco, L. García, J. A.

González, C. Rodríguez, and G. Rodríguez. A performance model for the analysis of OpenMP programs. Parallel and Distributed Computing Practices, 5(2):139– 151, June 2002. CODEN ???? ISSN 1097-2803.

Balasubramanian:2015:EGL

[BGG^{+}15] Raghuraman Balasubramanian, Vinay Gangadhar, Ziliang Guo, Chen-Han Ho, Cherin Joseph, Jaikrishnan Menon, Mario Paulo Drumond, Robin Paul, Sharath Prasad, Pradip Valathol, and Karthikeyan Sankaralingam. Enabling GPGPU low-level hardware explorations with MIAOW: an open-source RTL implementation of a GPGPU. ACM Transactions on Architecture and Code Optimization, 12(2):21:1–21:??, July 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bhanot:2005:OTL

[BGH^{+}05] G. Bhanot, A. Gara, P. Heidelberger, E. Lawless, J. C. Sexton, and R. Walkup. Optimizing task layout on the Blue Gene/L supercomputer. IBM Journal of Research and Development, 49 (2/3):489–500, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/bhanot.
pdf.

Bischof:2008:PRM

[BGK08] Christian Bischof, Niels Guertler, and Andreas Kowarz. Parallel reverse mode automatic differentiation for OpenMP programs with ADOL-C. In Bischof et al. [BBH^{+}08], pages 163–173. CO-DEN LNCSA6. ISBN 3-540-68935-4 (print), 3-540-68942-7 (e-book). ISSN 1439-7358. LCCN QA304 .I58 2008. URL http://

link.springer.com/content/
pdf/10.1007/978-3-540-
68942-3_15.

Butler:2000:SPM

[BGL00] Ralph Butler, William Gropp, and Ewing Lusk. A scalable process-management environment for parallel programs. Lecture Notes in Computer Science, 1908: 168–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080168.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080168.
pdf.

Beisel:1997:EMD

[BGR97a] T. Beisel, E. Gabriel, and M. Resch. An extension to MPI for distributed computing on MPPs. Lecture Notes in Computer Science, 1332:75–82, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brune:1997:HMP

[BGR97b] Matthias Brune, Jörn Gehring, and Alexander Reinefeld. Heterogeneous message passing and a link to resource management. The Journal of Supercomputing, 11 (4):355–369, December 1997. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
11&issue=4&spage=355;
http://www.wkap.nl/oasis.
htm/147011.

Breitenecker:1995:ESC

[BH95] Felix Breitenecker and Irmgard Husinsky, editors. EUROSIM ’95: simulation congress: proceedings of the EUROSIM Conference, EUROSIM ’95, Vienna, Austria, 11–15 September 1995. Elsevier, Amsterdam, The Netherlands, 1995. ISBN 0-444-82241-0. LCCN A76.9.C65E966 1995.

Bhargava:1993:PIW

[Bha93] Bharat Bhargava, editor. Proceedings of the IEEE Workshop on Advances in Parallel and Distributed Systems, October 6, 1993, Princeton, New Jersey. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-5250-0, 0-8186-5251-9. LCCN QA76.58.I444 1993.

Bhanot:1998:DTM

[Bha98] Gyan Bhanot. A 2-d transpose MPI code. Research report RC 21217, T. J. Watson Research Center, IBM Corporation, Almaden, CA, USA, 1998.

Bader:1996:PPA

[BHJ96] David A. Bader, David R. Helman, and Joseph JáJá. Practical parallel algorithms for personalized communication and integer sorting. ACM Journal of Experimental Algorithmics, 1:3:1–3:??, ???? 1996. CODEN ???? ISSN 1084-6654.

Bouteiller:2006:MVP

[BHK^{+}06] A. Bouteiller, T. Herault, G. Krawezik, P. Lemarinier, and F. Cappello. MPICH-V project: a multiprotocol automatic fault-tolerant MPI. The International Journal of High Performance Computing Applications, 20(3): 319–333, Fall 2006. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/20/
3/319.full.pdf+html.

Bubeck:1995:DSC

[BHKR95] T. Bubeck, M. Hiller, W. Kuchlin, and W. Rosenstiel. Distributed symbolic computation with DTS. In Ferreira and Rolim [FR95], pages 231– 248. ISBN 3-540-60321-2. LCCN QA76.642.I59 1995.

Bischof:1995:CSM

[BHLS^{+}95] C. Bischof, S. Huss-Lederman, Xiaobai Sun, A. Tsao, and T. Turnbull. A case study of MPI: Portable and efficient libraries. In Bailey et al. [BBG^{+}95], pages 728–733. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Bachem:1994:PCT

[BHM94] A. Bachem, W. Hochstättler, and M. Malich. Simulated trading — a new parallel approach for solving vehicle routing problems. In Joubert et al. [JPTE94], pages 471–475. ISBN 0-444-81841-3. LCCN QA76.58 .P3794 1993.

Bachem:1996:STH

[BHM96] A. Bachem, Hochstättler, and M. Malich. The simulated trading heuristic for solving vehicle routing problems. Discrete Applied Mathematics, 65(1-3):47–72, ????

1996. CODEN DAMADU. ISSN 0166-218X (print), 1872-6771 (electronic).

Brunst:2001:POL

[BHNW01] Holger Brunst, Hans-Christian Hoppe, Wolfgang E. Nagel, and Manuela Winkler. Performance optimization for large scale computing: The scalable VAMPIR approach. Lecture Notes in Computer Science, 2074:751–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2074/20740751.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2074/20740751.
pdf.

Barekas:2003:MAO

[BHP^{+}03] Vasileios K. Barekas, Panagiotis E. Hadjidoukas, Eleftherios D. Polychronopoulos, et al. A multiprogramming aware OpenMP implementation. Scientific Programming, 11(2):133–141, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bondhugula:2008:PAP

[BHRS08] Uday Bondhugula, Albert Hartono, J. Ramanujam, and P. Sadayappan. A practical automatic polyhedral parallelizer and locality optimizer. ACM SIG-PLAN Notices, 43(6):101– 113, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bisseling:2002:FMF

[BHS^{+}02] Georg Bißeling, Hans-Christian Hoppe, Alexander Supalov, Pierre Lagier, and Jean Latour. Fujitsu MPI-2: Fast locally, reaching globally. Lecture Notes in Computer Science, 2474:401–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/
service/series/0558/bibs/
2474/24740401.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740401.pdf.

Bazow:2018:MPS

[BHS18] Dennis Bazow, Ulrich Heinz, and Michael Strickland. Massively parallel simulations of relativistic fluid dynamics on graphics processing units with CUDA. Computer Physics Communications, 225(??):92– 113, April 2018. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465517300279.

Berka:2012:PET

[BHV12] Tobias Berka, Helge Hagenauer, and Marian Vajteršic. Portable explicit threading and concurrent programming for MPI applications. Lecture Notes in Computer Science, 7204: 81–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-31500-8_
9/.

Busa:2012:ACO

[BHW^{+}12] Ján Busa, Jr., Shura Hayryan, Ming-Chya Wu, Ján Busa, and Chin-Kun Hu. ARVO-CL: the OpenCL version of the ARVO package — an efficient tool for computing the accessible surface area and the excluded volume of proteins via analytical equations. Computer Physics Communications, 183(11): 2494–2497, November 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465512001580.

Bae:2017:SEF

[BHW^{+}17] Seung-Hee Bae, Daniel Halperin, Jevin D. West, Martin Rosvall, and Bill Howe. Scalable and efficient flow-based community detection for large-scale graph analysis. ACM Transactions on Knowledge Discovery from Data (TKDD), 11 (3):32:1–32:??, April 2017. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic).

Bickham:1995:POM

[Bic95]  J. L. Bickham.  Paral-

lel ocean modeling using Glenda. In ACM [ACM95a], pages 58–63. ISBN 0-89791-747-2. LCCN ????

Bernaschi:2005:ERA

[BIC05] Massimo Bernaschi, Giulio Iannello, and Saverio Crea. Experimental results about MPI collective communication operations. Parallel Processing Letters, 15(1/2): 223–236, March/June 2005. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Blas:2010:IEF

[BIC^{+}10] Javier Garcia Blas, Florin Isaila, Jesus Carretero, David Singh, and Felix Garcia-Carballeira. Implementation and evaluation of file write-back and prefetching for MPI-IO over GPFS. The International Journal of High Performance Computing Applications, 24(1):78– 92, February 2010. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/24/
1/78.full.pdf+html.

Branca:1995:CBH

[BID95] A. Branca, M. Ianigro, and A. Distante. A comparison between HPF and PVM for data parallel algorithms on a cluster of workstations using a high speed network. In Hertzberger and Serazzi [HS95a], pages 930–931. ISBN 3-540-59393-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Bilger:1995:AFM

[Bil95] R. W. Bilger, editor. 12th Australasian fluid mechanics conference: – December 1995, Sydney, Australia, Australasian Fluid Mechanics Conference 1995; EDIT 12//V2. University of Sydney, ????, 1995. ISBN 0-86934-034-4. LCCN ????

Bernaschi:1999:ERA

[BIL99] M. Bernaschi, G. Iannello, and M. Lauria. Experimental results about MPI collective communication operations. Lecture Notes in Computer Science, 1593: 774–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biradar:1994:ADL

[Bir94] Umesh V. Biradar. Adaptive distributed load balancing model for parallel virtual machine. Master of science in computer science, Department of Computer Science, College of Engineering, Lamar University, Beaumont, TX, USA, 1994. viii + 44 pp.

Bisseling:2004:PSC

[Bis04] Rob H. Bisseling. Parallel scientific computation: a structured approach using BSP and MPI. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-852939-2. xviii + 305 pp. LCCN QA76.58 .B57 2004.

URL http://www.loc.gov/
catdir/enhancements/fy0617/
2004046141-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0617/2004046141-t.html.

Baiardi:1993:PVM

[BJ93] F. Baiardi and M. Jazayeri.

P03M: a virtual machine approach to massively parallel computing. Proceedings of the International Conference on Parallel Processing, pages I–340–??, ???? 1993. CO-DEN PCPADL. ISSN 0190-3918.

Boianov:1995:DLC

[BJ95] L. Boianov and I. Jelly. Distributed logic circuit simulation on a network of workstations. In IEEE [IEE95h], pages 304–310. ISBN 0-8186-7031-2, 0-8186-7032-0. LCCN QA76.58 .E97 1995.

Barkati:2013:SPA

[BJ13] Karim Barkati and Pierre Jouvelot. Synchronous programming in audio processing: a lookup table oscillator case study. ACM Computing Surveys, 46(2): 24:1–24:??, November 2013. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Bjorge:1995:ISS

[Bjo95] D. Bjorge. Implementation of the semi-implicit scheme in a message passing version of HIRLAM (weather forecasting). In Hoffmann and Kreitz [HK95], pages 75–90. ISBN 981-02-2211-4. LCCN QC866.E26 1994.

Blaheta:1997:PIP

[BJS97] R. Blaheta, O. Jakl, and J. Stary. PVM-implementation of the PCG method with displacement decomposition. Lecture Notes in Computer Science, 1332:321–328, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blaheta:1999:LFM

[BJS99] R. Blaheta, O. Jakl, and J. Stary. Large-scale FE modelling in geomechanics: a case study in parallelization. In Dongarra et al. [DLM99], pages 299–306. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Bhandarkar:1996:MPM

[BK96] M. A. Bhandarkar and L. V.

Kale. MICE: a prototype MPI implementation in Converse environment. In IEEE [IEE96i], pages 26–31. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Bull:2000:JOL

[BK00] J. M. Bull and M. E. Kambites. JOMP: an OpenMPlike interface for Java. In ????, editor, Proceedings of the ACM 2000 conference on Java Grande, pages 44–53. ACM Press, New York, NY 10036, USA, 2000.

Balevic:2011:KAD

[BK11] Ana Balevic and Bart Kienhuis. KPN2GPU: an approach for discovery and exploitation of fine-grain data parallelism in process networks. ACM SIGARCH Computer Architecture News, 39(4):66–71, September 2011. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Bhandarkar:2001:ALB

[BKdSH01] Milind Bhandarkar, L. V.

Kalé, Eric de Sturler, and Jay Hoeflinger. Adaptive load balancing for MPI programs. Lecture Notes in Computer Science, 2074:108–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2074/20740108.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2074/20740108.
pdf.

Bekas:2002:PCP

[BKGS02] Constantine Bekas, Efrosini Kokiopoulou, Efstratios Gallopoulos, and Valeria Simoncini. Parallel computation of pseudospectra using transfer functions on a MATLAB-MPI cluster platform. Lecture Notes in Computer Science, 2474: 199–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.de/link/
service/series/0558/bibs/
2474/24740199.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740199.pdf.

Berka:2013:CPC

[BKH^{+}13] Tobias Berka, Giorgos Kollias, Helge Hagenauer, Marian Vajteršic, and Ananth Grama. Concurrent programming constructs for parallel MPI applications. The Journal of Supercomputing, 63(2):385–406, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11227-011-0739-5.

Ballard:2020:TPC

[BKK20] Grey Ballard, Alicia Klinvex, and Tamara G. Kolda. TuckerMPI: a parallel C++/ MPI software package for large-scale data compression via the Tucker tensor decomposition. ACM Transactions on Mathematical Software, 46(2):13:1–13:31, June 2020. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3378445.

Boryczko:1995:NIC

[BKML95] I. Boryczko, J. Kitowski, J. Moscinski, and A. Leszczynski. Numerically intensive computing as a benchmark for parallel computer architectures. In Hertzberger and Serazzi [HS95a], pages 118–123. ISBN 3-540-59393-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Bull:2000:PPJ

[BKO00] J. Mark Bull, Mark E.

Kambites, and Jan Obdrzalek. Parallel programming in Java with OpenMP-like directives. In ACM [ACM00],

                                        page 150. URL http://www.
                                        sc2000.org/proceedings/
                                        info/fp.pdf.

Beaugnon:2014:VVO

[BKvH^{+}14] Ulysse Beaugnon, Alexey Kravets, Sven van Haastregt, Riyadh Baghdadi, David Tweed, Javed Absar, and Anton Lokhmotov. VOBLA: a vehicle for optimized basic linear algebra. ACM SIG-PLAN Notices, 49(5):115– 124, May 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ballico:1994:PSP

[BL94] M. Ballico and H. Lederer. Plasmafusionsforschung: Serielles und paralleles Rechnen mit nur einem Programmcode auf Cray YMP, nCUBE2, Workstations mit PVM und KSR1. In Anonymous [Ano94c], pages 232– 234. ISBN ???? ISSN 0341-7778. LCCN Q180.55.E4 M39 1993.

Bendrider:1995:SME

[BL95] M. Bendrider and J.-M. Leclercq. Second-order Møller–Plesset and Epstein-Nesbet corrections to the molecular charge density: Distributed computing on a cluster of heterogeneous workstations with the PVM system. In Bernardi and Rivail [BR95a], pages 73– ?? ISBN 1-56396-457-0. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. LCCN QD39.3.E46 E15 1995.

Beazley:1997:EMP

[BL97] D. M. Beazley and P. S.

Lomdahl. Extensible message passing application development and debugging with Python. In IEEE [IEE97b], pages 650– 655. ISBN 0-8186-7793-7. LCCN QA76.58 .I56 1997. IEEE catalog number 97TB100107. IEEE Computer Society Press order number PR07792.

Bubak:1999:TPR

[BL99] M. Bubak and P. Luszczek.

Towards portable runtime support for irregular and out-of-core computations. In Dongarra et al. [DLM99], pages 59–66. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Baraglia:1993:PWC

[BLP93] R. Baraglia, D. Laforenza, and R. Perego. Programming a workstation cluster with PVM and Linda: a qualitative and quantitative comparison. In Anonymous [Ano93b], pages 101– 114. ISBN ???? LCCN ????

Bach:2013:LQB

[BLPP13] Matthias Bach, Volker Lindenstruth, Owe Philipsen, and Christopher Pinke. Lattice QCD based on OpenCL. Computer Physics Communications, 184(9):2042– 2052, September 2013. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465513001288.

Belviranli:2018:JDA

[BLVB18] Mehmet E. Belviranli, Seyong Lee, Jeffrey S. Vetter, and Laxmi N. Bhuyan. Juggler: a dependence-aware task-based execution framework for GPUs. ACM SIG-PLAN Notices, 53(1):54–67, January 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bubak:1998:PCL

[BLW98] M. Bubak, P. Luszczek, and A. Wierzbowska. Porting CHAOS library to MPI. Lecture Notes in Computer Science, 1497:131–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhandarkar:1997:CRP

[BM97] Suchendra M. Bhandarkar and Salem Machaka. Chromosome reconstruction from physical maps using a cluster of workstations. The Journal of Supercomputing, 11(1):61–86, March 1997. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
11&issue=1&spage=61;
http://www.wkap.nl/oasis.
htm/141471.

Booth:2000:SSM

[BM00] S. Booth and E. Mourao.

Single-sided MPI implementations for SUN MPI. In ACM [ACM00], page 46.

URL http://www.sc2000.
org/proceedings/techpapr/
papers/pap182.pdf.

Basumallik:2002:TOE

[BME02] Ayon Basumallik, Seung-Jai Min, and Rudolf Eigenmann. Towards OpenMP execution on software distributed shared memory systems. Lecture Notes in Computer Science, 2327: 457–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2327/23270457.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2327/23270457.
pdf.

Buntinas:2007:IES

[BMG07] Darius Buntinas, Guillaume Mercier, and William Gropp. Implementation and evaluation of shared-memory communication and synchronization operations in MPICH2 using the Nemesis communication subsystem. Parallel Computing, 33(9):634–644, September 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bronevetsky:2003:AAL

[BMPS03] Greg Bronevetsky, Daniel Marques, Keshav Pingali, and Paul Stodghill. Automated application-level checkpointing of MPI programs. ACM SIGPLAN Notices, pages 84–94, 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bubak:1994:PDS

[BMPZ94a] M. Bubak, J. Mosciniski, M. Pogoda, and W. Zdechlikiewicz. Parallel distributed 2-D short-range molecular dynamics on networked workstations. In Dongarra and Wasniewski [DW94], pages 127–135. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1994. DM104.00.

Bubak:1994:EMD

[BMPZ94b] M. Bubak, J. Moscinski, M. Pogoda, and W. Zdechlikiewicz. Efficient molecular dynamics simulation on networked workstations. In Gruber and Tomassini [GT94], pages 191–194. ISBN 2-88270-011-3. LCCN QC20.7.E4I58 1994.

Baiardi:2001:CRD

[BMR01] Fabrizio Baiardi, Paolo Mori, and Laura Ricci. Collecting remote data in irregular problems with hierarchical representation of the domain. Lecture Notes in Computer Science, 2131: 304–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310304.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310304.
pdf.

Brightwell:2002:DIM

[BMR02] Ron Brightwell, Arthur B.

Maccabe, and Rolf Riesen. Design and implementation of MPI on Portals 3.0. Lecture Notes in Computer Science, 2474:331–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/
service/series/0558/bibs/
2474/24740331.htm;   http:
//link.springer.de/link/

                                        service/series/0558/papers/
                                        2474/24740331.pdf.

Bubak:1994:FLG

[BMS94a] M. Bubak, J. Moscinski, and R. Slota. FHP lattice gas on networked workstations. In Gruber and Tomassini [GT94], pages 427–430. ISBN 2-88270-011-3. LCCN QC20.7.E4I58 1994.

Bubak:1994:IPL

[BMS94b] M. Bubak, J. Moscinski, and R. Slota. Implementation of parallel lattice gas program on workstations under PVM. In Dongarra and Wasniewski [DW94], pages 136–146. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1994. DM104.00.

Barthels:2017:DJA

[BMS^{+}17] Claude Barthels, Ingo Müller, Timo Schneider, Gustavo Alonso, and Torsten Hoefler. Distributed join algorithms on thousands of cores. Proceedings of the VLDB Endowment, 10(5): 517–528, January 2017. CO-DEN ???? ISSN 2150-8097.

Boschetti:2019:MOD

[BMS19] Marco Antonio Boschetti, Vittorio Maniezzo, and Francesco Strappaveccia. Membership overlay design optimization with resource constraints (accelerated on GPU). Journal of Parallel and Distributed Computing, 133(??): 286–296, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731518304908.

Berrendorf:2000:PCO

[BN00] Rudolf Berrendorf and Guido Nieken. Performance characteristics for OpenMP constructs on different parallel computer architectures. Concurrency: practice and experience, 12(12):1261– 1273, October 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76500355/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500355&PLACEBO=IE.
pdf.

Bawidamann:2012:ETO

[BN12] Uwe Bawidamann and Marco Nehmeier. Expression templates and OpenCL. Lecture Notes in Computer Science, 7204:71–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-31500-8_

Bull:2001:MSO

[BO01] J. Mark Bull and Darragh O’Neill. A microbenchmark suite for OpenMP 2.0. ACM SIGARCH Computer Architecture News, 29(5):41–48, December 2001. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Bubak:2000:IOB

[BoFBW00] Marian Bubak, W. odzimierz Funika, Bartosz Balis, and Roland Wismüller. Interoperability of OCM-based on-line tools. Lecture Notes in Computer Science, 1908:242–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1908/19080242.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1908/19080242.
                                        pdf.

Boisvert:1997:QNS

[Boi97] R. F. Boisvert, editor. Quality of numerical software: assessment and enhancement / proceedings of the IFIP TC2/WG2.5 Working Conference on the Quality of Numerical Software, Assessment and Enhancement, Oxford, United Kingdom, 8– 12 July 1996. Chapman and Hall, Ltd., London, UK, 1997. ISBN 0-412-80530-8. LCCN QA297 .I35 1996.

Bonnet:1996:UPW

[Bon96] C. Bonnet. Using PVM in wireless network environments. In Bode et al. [BDLS96], pages 296–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Booth:2001:OML

[Boo01] Stephen Booth. Optimising the MPI library for the T3E. Lecture Notes in Computer Science, 2150:80–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2150/21500080.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2150/21500080.
pdf.

Borkowski:1999:LVC

[Bor99] J. Borkowski. On line visualization or combining the standard ORNL PVM with a vendor PVM implementation. In Dongarra et al. [DLM99], pages 157–164. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Boszormenyi:1996:PCT

[Bos96] Laszlo Boszormenyi, editor. Parallel computation: Third International ACPC Conference with special emphasis on parallel databases and parallel I/O, Klagenfurt, Austria, September 23–25, 1996: proceedings, volume 1127 of Lecture notes in computer science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61695-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1127.

Brebbia:1993:ASE

[BP93] C. A. Brebbia and H. Power, editors. Applications of Supercomputers in Engineering III, 27–29 September 1993, Bath, UK. Computational Mechanics Publication, London, UK, 1993. ISBN 1-85312-236-X. LCCN TA345.I556 1993.

Berthou:1998:PHM

[BP98] J.-Y. Berthou and L. Plagne. Parallel HPF-MPI implementation of the TBSCM Poisson solver. Lecture Notes in Computer Science, 1401: 252–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbosa:1999:ADM

[BP99] J. Barbosa and A. Padilha. Algorithm-dependant method to determine the optimal number of computers in parallel virtual machines. Lecture Notes in Computer Science, 1573:508–521, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beletsky:1994:OPV

[BPC94] V. Beletsky, T. Popova, and A. Chemeris. Organization of a parallel virtual machine. In Horiguchi et al. [HHK94], pages 421– 426. ISBN 0-8186-6507-6 (case), 0-8186-6506-8 (microfiche). LCCN QA76.58 .I5673 1994 Bar. IEEE catalog number 94TH0697-3.

Becks:1994:NCT

[BPG94] K.-H. Becks and D. Perret-Gallix, editors. New computing techniques in physics research III: proceedings of the Third International Workshop on Software Engineering, Artificial Intelligence and Expert Systems for High Energy and Nuclear Physics: October 4– 8, 1993, Oberammergau, Germany. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1994. ISBN 981-02-1699-8. LCCN QC793.47.E4I58 1993.

Bouhrour:2022:TLC

[BPJ22] Stephane Bouhrour, Thibaut Pepin, and Julien Jaeger. Towards leveraging collective performance with the support of MPI 4.0 features in MPC. Parallel Computing, 109(??):??, March 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819121001034.

Barbosa:1997:EUW

[BPMN97] J. G. Barbosa, A. J. Padilha, J.-P. Madier, and T. Neubert. Experiments on using WPVM for industrial visual inspection problems. Lecture Notes in Computer Science, 1300:828–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baptista:2001:IOS

[BPS01] Tiago Baptista, Hernani Pedroso, and João Gabriel Silva. The implementation of one-sided communications for WMPI II. Lecture Notes in Computer Science, 2131:61–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2131/21310061.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2131/21310061.
                                        pdf.

Balou:1991:DIV

[BR91] A. T. Balou and A. N. Refenes. The design and implementation of VOOM: a parallel virtual object oriented machine. Microprocessing and Microprogramming, 32(1-5):289–296, August 1991. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic).

Burrer:1994:RRB

[BR94] C. Burrer and P. Remy. RU-BIS: a runtime basic interface software on TELMAT T9000 TN series. In de Gloria et al. [dGJM94], pages 63–78. ISBN ???? LCCN ????

Bernardi:1995:CCE

[BR95a] Francesco Bernardi and Jean-Louis Rivail, editors. Computational chemistry: 1st European conference on computational chemistry (May 1994, Nancy, France), number 330 in AIP Conference Proceedings. American Institute of Physics, Woodbury, NY, USA, 1995. ISBN 1-56396-457-0. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. LCCN QD39.3.E46 E15 1995.

Bernaschi:1995:PEI

[BR95b] M. Bernaschi and G. Richelli.

PVMe: an enhanced implementation of PVM for the IBM 9076 SP2. In Hertzberger and Serazzi [HS95a], pages 461–471. ISBN 3-540-59393-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Bernaschi:1995:DRP

[BR95c] Massimo Bernaschi and Giorgio Richelli. Development and results of PVMe on the IBM 9076 SP1. Journal of Parallel and Distributed Computing, 29(1): 75–83, August 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/doi/10.1006/jpdc.
                                        1995.1107/production;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        jpdc.1995.1107/production/
                                        pdf.

Bane:2002:EOA

[BR02] M. K. Bane and G. D. Riley. Extended overhead analysis for OpenMP (research note). Lecture Notes in Computer Science, 2400:162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2400/24000162.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2400/24000162.
                                        pdf.

Boeres:2004:ETF

[BR04] Cristina Boeres and Vinod E. F. Rebello. EasyGrid: towards a framework for the automatic Grid enabling of legacy MPI applications. Concurrency and Computation: Practice and Experience, 16(5): 425–432, April 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bergstrom:2012:NDP

[BR12] Lars Bergstrom and John Reppy. Nested dataparallelism on the GPU. ACM SIGPLAN Notices, 47(9):247–258, September 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bramley:1997:TNR

[Bra97] Randall Bramley. Technology news & reviews: Chemkin software; OpenMP Fortran Standard; ODE toolbox for Matlab; Java products; Scientific Work-Place 3.0. IEEE Computational Science & Engineering, 4(4):75–78, October/December 1997. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1997/pdf/c4075.
pdf.

Briscolini:1995:PID

[Bri95] M. Briscolini. A parallel implementation of a 3-D pseudospectral based code on the IBM 9076 scalable POWER parallel system. Parallel Computing, 21(11):1849– 1862, November 29, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=21&
issue=11&aid=1027.

Brieger:2000:HOO

[Bri00] Leesa Brieger. HPF to OpenMP on the Origin2000: a case study. Concurrency: practice and experience, 12(12):1147–1154, October 2000. CODEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76500351/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500351&PLACEBO=IE.
pdf.

Brightwell:2002:RMR

[Bri02] Ron Brightwell. Readymode receive: An optimized receive function for MPI. Lecture Notes in Computer Science, 2474:385–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/
service/series/0558/bibs/
2474/24740385.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740385.pdf.

Brightwell:2010:EDA

[Bri10] Ron Brightwell. Exploiting direct access shared memory for MPI on multi-core processors. The International Journal of High Performance Computing Applications, 24(1):69–77, February 2010. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/24/
1/69.full.pdf+html.

Brightwell:2003:DIP

[BRM03] Ron Brightwell, Rolf Riesen, and Arthur B. Maccabe. Design, implementation, and performance of MPI on Portals 3.0. The International Journal of High Performance Computing Applications, 17(1):7–20, Spring 2003. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Boudet:1999:PIH

[BRR99] V. Boudet, F. Rastello, and Y. Robert. PVM implementation of heterogeneous ScaLAPACK dense linear solvers. In Dongarra et al. [DLM99], pages 333–340. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Benzoni:1992:CLF

[BRS92] A. Benzoni, G. Richelli, and V. S. Sunderam. Concurrent LU factorization on workstation networks. In Evans et al. [EJL92], pages 159– 166. ISBN 0-444-89212-5. LCCN QA76.58.I545 1991.

Briley:1994:NNH

[BRST94] W. R. Briley, D. S. Reese, A. Skjellum, and L. H. Turcotte. NHPDCC: The National High Performance Distributed Computing Consortium. In IEEE [IEE94f], pages 2–9. ISBN 0-8186-4980-1. LCCN QA76.58.S34 1993.

Bruck:1995:EMPa

[Bru95] Jehoshua Bruck. Efficient message passing interface (MPI) for parallel computing on clusters of workstations. Research report RJ 9925 (87305), IBM T. J. Watson Research Center, Yorktown Heights, NY, USA, 1995. 31 pp.

Brightwell:2005:AIO

[BRU05] Ron Brightwell, Rolf Riesen, and Keith D. Underwood. Analyzing the impact of overlap, offload, and independent progress for Message Passing Interface applications. The International Journal of High Performance Computing Applications, 19(2):103–117, Summer 2005. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/19/
2/103.full.pdf+html.

Bruning:2012:MFT

[Brü12] Ulrich Brüning. MPI functions and their impact on interconnect hardware. Lecture Notes in Computer Science, 7490:10, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/accesspage/
chapter/10.1007/978-3-
642-33518-1_2.

Barth:1993:CNM

[BS93] N. H. Barth and S. L. Smith.

Coupling numerical models of the atmosphere and ocean using the parallel virtual machine (PVM) package. In Sincovec [Sin93], pages 71–

75. ISBN 0-89871-315-3. LCCN QA 76.58 S55 1993. Two volumes.

Bolding:1994:PCR

[BS94] Kevin Bolding and Lawrence Snyder, editors. Parallel computer routing and communication: first international workshop, PCRCW ’94, Seattle, Washington, USA, May 16–18, 1994: proceedings, number 853 in Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58429-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.P39 1994.

Beguelin:1996:TMD

[BS96a] A. Beguelin and V. Sunderam. Tools for monitoring, debugging, and programming in PVM. In Bode et al. [BDLS96], pages 7–13. ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Brightwell:1996:DIM

[BS96b] R. Brightwell and L. Shuler. Design and implementation of MPI on Puma portals. In IEEE [IEE96i], pages 18–25. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Blikberg:2001:NPA

[BS01] Ragnhild Blikberg and Tor Sørevik. Nested parallelism: Allocation of threads to tasks and OpenMP implementation. Scientific Programming, 9(2–3):185–194, Spring–Summer 2001. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

                                        iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=7pab6qgbaf8vxg991rwy%
                                        26referrer=parent%26backto=
                                        issue%2C11%2C11%3Bjournal%
                                        2C1%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Blikberg:2005:LBO

[BS05] R. Blikberg and T. Sørevik.

Load balancing and OpenMP implementation of nested parallelism. Parallel Computing, 31(10–12):984–998, October/December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Brown:2007:HSP

[BS07] Russell Brown and Ilya Sharapov. High-scalability parallelization of a molecular modeling application: Performance and productivity comparison between OpenMP and MPI implementations. International Journal of Parallel Programming, 35(5):441– 458, October 2007. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=5&spage=441.

Betcke:2021:DHP

[BS21] Timo Betcke and Matthew W.

Scroggs. Designing a highperformance boundary element library with OpenCL and Numba. Computing in Science and Engineering, 23(4):18–28, July/ August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bassomo:1999:PGE

[BSC99] P. Bassomo, I. Sakho, and A. Corbel. Porting generalized eigenvalue software on distributed memory machines using systolic model principles. In Dongarra et al. [DLM99], pages 396– 403. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Bolton:2000:MPL

[BSG00] Hermanus P. J. Bolton, Jaco F. Schutte, and Albert A. Groenwold. Multiple parallel local searches in global optimization. Lecture Notes in Computer Science, 1908:88–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1908/19080088.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1908/19080088.
                                        pdf.

Bukata:2015:SRC

[BSH15] Libor Bukata, Premysl Sucha, and Zdenek Hanzálek. Solving the resource constrained project scheduling problem using the parallel tabu search designed for the CUDA platform. Journal of Parallel and Distributed Computing, 77(??):

58–68, March 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731514002226.

Bakhtiari:1995:APL

[BSN95] S. Bakhtiari and R. Safavi-Naini. Application of PVM to linear cryptanalysis. In Gray and Naghdy [GN95], pages 278–279. ISBN ???? LCCN ????

Bai:2013:SLA

[BST^{+}13] Mingze Bai, Shixin Sun, Hong Tang, Yusheng Dou, and Glenn V. Lo. An SPMDlike algorithm for parallelizing molecular dynamics using OpenMP. Computing in Science and Engineering, 15(4):48–56, July/ August 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Benzoni:1991:MFR

[BSvdG91] A. Benzoni, V. S. Sunderam, and R. van de Guijn. Matrix factorization on a RISC workstation network. In Durand and El Dabaghi [DE91], pages 207–218. ISBN 0-444-89224-9. LCCN QA75.5.I585 1991.

Blaszczyk:1996:EPI

[BT96] A. Blaszczyk and C. Trinitis. Experience with PVM in an industrial environment. In Bode et al.

[BDLS96], pages 174–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

biewski:2001:MOS

[bT01a] Maciej Go biewski and Jesper Larsson Träff. MPI-2 one-sided communications on a Giganet SMP cluster. Lecture Notes in Computer Science, 2131:16–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310016.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310016.
pdf.

Bu:2001:PAC

[BT01b] Libor Bu and Pavel Tvrdík. A parallel algorithm for connected components on distributed memory machines. Lecture Notes in Computer Science, 2131:280–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310280.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310280.
pdf.

Bonelli:2017:MCA

[BTC^{+}17] Francesco Bonelli, Michele Tuttafesta, Gianpiero Colonna, Luigi Cutrone, and Giuseppe Pascazio. An MPI–CUDA approach for hypersonic flows with detailed stateto-state air kinetics using a GPU cluster. Computer Physics Communications, 219(??):178–195, October 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465517301613.

Badia:1999:SIT

[BV99] J. M. Badia and A. M. Vidal. Solving the inverse Toeplitz eigenproblem using ScaLAPACK and MPI. In Dongarra et al. [DLM99], pages 372–379. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Baltas:1994:CPC

[BvdB94] N. D. Baltas and C. S.

van den Berghe. Comparison of the porting of a computational fluid dynamics application to SIMD and MIMD computers. In Dekker et al. [DSZ94], pages 761–767. ISBN 0-444-81784-

0. LCCN QA76.58.E98 1994.

Berendsen:1995:GMP

[BvdSvD95] H. J. C. Berendsen, D. van der Spoel, and R. van Drunen. GROMACS: a messagepassing parallel molecular dynamics implementation. Computer Physics Communications, 91(1-3):43–56, September 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Baskaran:2012:ACO

[BVML12] Muthu Manikandan Baskaran, Nicolas Vasilache, Benoit Meister, and Richard Lethin. Automatic communication optimizations through memory reuse strategies. ACM SIGPLAN Notices, 47(8): 277–278, August 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Berg:2012:FCL

[BW12] Bernd A. Berg and Hao Wu. Fortran code for SU(3) lattice gauge theory with and without MPI checkerboard parallelization. Computer Physics Communications, 183(10):2145– 2157, October 2012. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465512001269.

Blum:1996:PIP

[BWT96] J. M. Blum, T. M. Warschko, and W. F. Tichy. PSPVM: implementing PVM on a high-speed interconnect for workstation clusters. In Bode et al. [BDLS96], pages 235–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Bureddy:2012:OGM

[BWV^{+}12] D. Bureddy, H. Wang, A. Venkatesh, S. Potluri, and D. K. Panda. OMB-GPU: a micro-benchmark suite for evaluating MPI libraries on GPU clusters. Lecture Notes in Computer Science, 7490: 110–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-33518-1_
16/.

Bihari:2012:CIT

[BWW^{+}12] Barna L. Bihari, Michael Wong, Amy Wang, Bronis R. de Supinski, and Wang Chen. A case for including transactions in OpenMP II: Hardware transactional memory. Lecture Notes in Computer Science, 7312: 44–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-30961-8_

Blattner:2012:PSC

[BY12] Timothy Blattner and Shiming Yang. Performance study on CUDA GPUs for parallelizing the local ensemble transformed Kalman filter algorithm. Concurrency and Computation: Practice and Experience, 24(2): 167–177, February 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bendtsen:1997:RLS

[BZ97] C. Bendtsen and Z. Zlatev. Running large-scale air pollution models on message passing machines. Lecture Notes in Computer Science, 1332:417–426, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carpen-Amarie:2017:EOC

[CAHT17] Alexandra Carpen-Amarie, Sascha Hunold, and Jesper Larsson Träff. On expected and observed communication performance with MPI derived datatypes. Parallel Computing, 69(??):98– 117, November 2017. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819117301217.

Calmet:1994:RWC

[Cal94] J. Calmet, editor. Rhine workshop on computer algebra — March 22–24, 1994, Karlsruhe, Germany. Universität Karlsruhe, Karlsruhe, Germany, 1994. ISBN ???? LCCN ????

Cabarle:2012:SNP

[CAM12] Francis George C. Cabarle, Henry Adorna, and Miguel A. Martínez. A spiking neural P system simulator based on CUDA. Lecture Notes in Computer Science, 7184: 87–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-28024-5_
8/.

Carbajal:2007:PTD

[Car07] Santiago Garcia Carbajal.

Parallelizing three dimensional cellular automata with OpenMP. Parallel Processing Letters, 17(4):349– 361, December 2007. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Campanoni:2010:HFP

[CARB10] Simone Campanoni, Giovanni Agosta, Stefano Crespi Reghizzi, and Andrea Di Biagio. A highly flexible, parallel virtual machine: design and experience of ILD-JIT. Software—Practice and Experience, 40(2):177–207, February ??, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cavender:1993:APV

[Cav93] Mark Edward Cavender. Asynchronous parallel virtual machine. M.s. thesis, University of Texas at San Antonio. Division of Mathematics and Computer Science and Statistics, San Antonio, TX, USA, 1993. vi + 228 pp.

Chabbi:2017:EAL

[CAWL17] Milind Chabbi, Abdelhalim Amer, Shasha Wen, and Xu Liu. An efficient abortable-locking protocol for multi-level NUMA systems. ACM SIGPLAN Notices, 52(8):61–74, August 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cartwright:2000:AOE

[CB00] Keith L. Cartwright and Joseph D. Blahovec. Adding OpenMP to an existing MPI code: Will it be beneficial? In ACM [ACM00],

                                        page 145. URL http://www.
                                        sc2000.org/proceedings/
                                        info/fp.pdf.

Czapinski:2011:TST

[CB11] Michal Czapiński and Stuart Barnes. Tabu Search with two approaches to parallel flowshop evaluation on CUDA platform. Journal of Parallel and Distributed Computing, 71(6): 802–811, June 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731511000384.

Creech:2016:TSS

[CB16] Timothy Creech and Rajeev Barua. Transparently space sharing a multicore among multiple processes. ACM Transactions on Parallel Computing (TOPC), 3(3):17:1–17:??, December 2016. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Cesarini:2020:CSR

[CBB^{+}20] D. Cesarini, A. Bartolini, A. Borghesi, C. Cavazzoni, M. Luisier, and L. Benini. Countdown slack: a runtime library to reduce energy footprint in large-scale MPI applications. IEEE Transactions on Parallel and Distributed Systems, 31(11): 2696–2709, 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cesarini:2021:CRT

[CBB^{+}21] D. Cesarini, A. Bartolini, P. Bonfà, C. Cavazzoni, and L. Benini. COUNT-DOWN: a run-time library for performance-neutral energy saving in MPI applications. IEEE Transactions on Computers, 70(5): 682–695, 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cooper:1994:CHF

[CBHH94] M. D. Cooper, N. A. Burton, R. J. Hall, and I. H. Hillier. Combined Hartree– Fock and density functional theory: a distributed memory parallel implementation. Journal of molecular structure. Theochem, 121:97–107, December 1994. CODEN THEODJ. ISSN 0166-1280 (print), 1872-7999 (electronic).

Coronado-Barrientos:2019:ANF

[CBIGL19]  E.  Coronado-Barrientos,

G. Indalecio, and A. García-Loureiro. AXC: a new format to perform the SpMV oriented to Intel Xeon Phi architecture in OpenCL. Concurrency and Computation: Practice and Experience, 31(1):e4864:1– e4864:??, January 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Casas:2010:APD

[CBL10] Marc Casas, Rosa M. Badia, and Jesús Labarta. Automatic phase detection and structure extraction of MPI applications. The International Journal of High Performance Computing Applications, 24(3):335–360, August 2010. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/24/
3/335.full.pdf+html.

Che:2008:PSG

[CBM^{+}08] Shuai Che, Michael Boyer, Jiayuan Meng, David Tarjan, Jeremy W. Sheaffer, and Kevin Skadron. A performance study of generalpurpose applications on graphics processors using CUDA. Journal of Parallel and Distributed Computing, 68(10):1370–1380, October 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chapman:2002:APU

[CBPP02] B. Chapman, F. Bregier, A. Patil, and A. Prabhakar. Achieving performance under OpenMP on ccNUMA and software distributed shared memory systems. Concurrency and Computation: Practice and Experience, 14(8–9):713–739, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/95016122/
START;   http://www3.interscience.

                                        wiley.com/cgi-bin/fulltext?
                                        ID=95016122{\&}PLACEBO=
                                        IE.pdf.

Cowles:2018:ISB

[CBS18] Mary Kathryn Cowles, Stephen Bonett, and Michael Seedorff. Independent sampling for Bayesian normal conditional autoregressive models with OpenCL acceleration. Computational Statistics, 33(1):159– 177, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s00180-017-0752-0.

Clay:2018:GAP

[CBYG18] M. P. Clay, D. Buaria, P. K. Yeung, and T. Gotoh. GPU acceleration of a petascale application for turbulent mixing at high Schmidt number using OpenMP 4.5. Computer Physics Communications, 228(??):100–114, July 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465518300596.

Chapple:1995:PUL

[CC95] S. R. Chapple and L. J. Clarke. The Parallel Utilities Library. In IEEE [IEE95j], pages 21–30. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Cormen:1999:PBP

[CC99] Thomas H. Cormen and James C. Clippinger. Performing BMMC permutations efficiently on distributedmemory multiprocessors with MPI. Algorithmica, 24 (3–4):349–370, August 1999. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:/

/link.springer.de/link/
service/journals/00453/
bibs/24n3p349.html;   http:
//www.springerlink.com/
openurl.asp?genre=article&
issn=0178-4617&volume=
24&issue=3&spage=349.

Ciaccio:2000:GMG

[CC00a] Giuseppe Ciaccio and Giovanni Chiola. GAMMA and MPI/GAMMA on gigabit Ethernet. Lecture Notes in Computer Science, 1908:129–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080129.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080129.
pdf.

Couturier:2000:PMD

[CC00b] Raphaël Couturier and Christophe Chipot. Parallel molecular dynamics using OpenMP on a shared memory machine. Computer Physics Communications, 124(1):49–59, January 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465599004324.

Cardoso:2010:MSO

[CC10] M. C. Cardoso and F. M. Costa. MPI support on opportunistic grids based on the InteGrade middleware. Concurrency and Computation: Practice and Experience, 22(3):343–357, March 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2017:AAG

[CC17] Jian Chen and Russell M. Clapp. Astro: Autogeneration of synthetic traces using scaling pattern recognition for MPI workloads. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2159–2171, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/08/07809142-
abs.html.

Chen:2000:MCO

[CCA00] Hsiang Ann Chen, Yvette O. Carrasco, and Amy W.

Apon. MPI collective operations over IP multicast. Lecture Notes in Computer Science, 1800:51–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1800/18000051.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1800/18000051.
pdf.

Couder-Castaneda:2015:PCM

[CCBPGA15] C. Couder-Castañeda, H. Barrios-Piña, I. Gitler, and M. Arroyo. Performance of a code migration for the simulation of supersonic ejector flow to SMP, MIC, and GPU using OpenMP, OpenMP+LEO, and OpenACC directives. Scientific Programming, 2015(??): 739107:1–739107:20, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/739107/

.

Castagnera:1994:NEP

[CCF^{+}94] K. Castagnera, D. Cheng, R. Fatoohi, E. Hook, B. Kramer, C. Manning, J. Musch, C. Niggley, W. Saphir, D. Sheppard, M. Smith, I. Stockdale, S. Welch, R. Williams, and D. Yip. NAS experiences with a prototype cluster of workstations. In IEEE [IEE94h], pages 410–419. ISBN 0-8186-6607-2, 0-8186-6605-6, 0-8186-6606-4. ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Cooperman:2003:UTC

[CCHW03] Gene Cooperman, Henri Casanova, Jim Hayes, and Thomas Witzel. Using TOP-C and AMPIC to port large parallel applications to the Computational Grid. Future Generation Computer Systems, 19(4):587–596, May 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Casas:1995:MMT

[CCK^{+}95] Jeremy Casas, Dan L. Clark, Ravi Konuru, Steve W. Otto, Robert M. Prouty, and Jonathan Walpole. MPVM: a migration transparent version of PVM. Computing systems: the journal of the USENIX Association, 8(2): 171–216, Spring 1995. CO-DEN CMSYE2. ISSN 0895-6340.

Collingbourne:2012:STO

[CCK12] Peter Collingbourne, Cristian Cadar, and Paul H. J. Kelly. Symbolic testing of OpenCL code. Lecture Notes in Computer Science, 7261: 203–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.

springer.com/chapter/10.
1007/978-3-642-34188-5_
18/.

Costa:2006:ROA

[CCM^{+}06] J. J. Costa, T. Cortes, X. Martorell, E. Ayguade, and J. Labarta. Running OpenMP applications efficiently on an everythingshared SDSM. Journal of Parallel and Distributed Computing, 66(5):647–658, May 2006. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:2012:PUA

[CCM12] Yifeng Chen, Xiang Cui, and Hong Mei. PARRAY: a unifying array representation for heterogeneous parallelism. ACM SIGPLAN Notices, 47(8):171–180, August 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Ciglaric:2019:OLP

[CCS19] Tadej Ciglaric, Rok Cesnovar, and Erik Strumbelj. An OpenCL library for parallel random number generators. The Journal of Supercomputing, 75 (7):3866–3881, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Clematis:1997:DNL

[CCSM97] A. Clematis, A. Coda, M. Spagnuolo, and M. Mineter. Developing non-local iterative parallel algorithms for GIS on Cray T3D using MPI. Lecture Notes in Computer Science, 1332: 435–442, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chamaret:1995:PFE

[CCU95] B. Chamaret, H. Cherefi, and S. Ubeda. Parallel filter estimation maximisation algorithm for segmentation on a LAN of workstation. In Bailey et al. [BBG^{+}95], pages 68–69. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Coulaud:1996:EIP

[CD96] O. Coulaud and E. Dillon. Early implementation of Para++ with MPI-2. In IEEE [IEE96i], pages 95– 101. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Cunha:1998:MPP

[CD98] J. C. Cunha and V. Duarte. Monitoring PVM programs using the DAMS approach. Lecture Notes in Computer Science, 1497:273–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cotronis:2001:RAP

[CD01] Yiannis Cotronis and J. J.

Dongarra, editors. Recent advances in parallel virtual machine and message passing interface: 8th European PVM/MPI Users’ Group Meeting, Santorini/Thera, Greece, September 23–26, 2001: proceedings, volume 2131 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42609-4 (paperback). LCCN QA76.58 E975 2001; QA267.A1 L43 no.2131. URL http:/

/link.springer-ny.com/
link/service/series/0558/
tocs/t2131.htm.

Clemencon:1996:THM

[CDD^{+}96] C. Clemencon, K. M.

Decker, V. R. Deshpande, A. Endo, J. Fritscher, P. A. R. Lorenzo, N. Masuda, A. Muller, R. Ruhl, W. Sawyer, B. J. N. Wylie, and F. Zimmermann. Toolssupported HPF and MPI parallelization of the NAS parallel benchmarks. In IEEE [IEE96c], pages 309– 318. ISBN 0-8186-7551-9. LCCN QA76.58 .S95 1996. IEEE catalog number 96TB100062.

Cao:2013:CHP

[CDD^{+}13] Chongxiao Cao, Jack Dongarra, Peng Du, Mark Gates, Piotr Luszczek, and Stanimire Tomov. clMAGMA: High performance dense linear algebra with OpenCL. LAPACK Working Note 275, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, March 2013. URL http:/

                                        /www.netlib.org/lapack/
                                        lawnspdf/lawn275.pdf.

Conforti:1996:PIA

[CdGM96] D. Conforti, L. de Luca, L. Grandinetti, and R. Musmanno. A parallel implementation of automatic differentiation for partially separable functions using PVM. Parallel Computing, 22(5): 643–656, August 8, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=22&
                                        issue=5&aid=1065.

Cownie:1994:PPP

[CDH^{+}94] J. Cownie, A. Dunlop, S. Hellberg, A. J. G. Hey, and D. Pritchard. Portable parallel programming environments-the ES-PRIT PPPE project. In Dekker et al. [DSZ94], pages 135–142. ISBN 0-444-81784-

0. LCCN QA76.58.E98 1994.

Chang:1995:EPCb

[CDH^{+}95] Sheue-Ling Chang, David Hung-Chang Du, Jenwei Hsieh, Rose P. Tsang, and Mengjou Lin. Enhanced PVM communications over a High-Speed LAN. IEEE parallel and distributed technology: systems and applications, 3(3):20–32, Fall 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Chang:1995:EPCa

[CDHL95] S.-L. Chang, D. H. C. Du, J. Hsieh, and M. Lin. Enhanced PVM communications over a high-speed local area network. In Alnuweiri and Hamdi [AH95], pages 37–46. ISBN 0-8186-7124-6. LCCN TK5105.5 .H56 1995.

Casanova:1995:PPM

[CDJ95] Henri Casanova, Jack Dongarra, and Weicheng Jiang. The performance of PVM on MPP systems. Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, August

1995. URL http://www.
netlib.org/utk/papers/
pvmmpp.ps;   http://www.
netlib.org/utk/papers/
pvmmpp/pvmmpp.html;   http:
//www.netlib.org/utk/people/
JackDongarra/pdf/pvmmpp.
pdf.

Chandra:2001:PPO

[CDK^{+}01] Rohit Chandra, Leonardo Dagum, David Kohr, Dror Maydan, Jeff McDonald, and Ramesh Menon. Parallel Programming in OpenMP. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2001. ISBN 1-55860-671-8. xvi + 230 pp. LCCN QA76.642 .P38 2001. US$39.95.

                                        URL http://www.mkp.com/
                                        books_catalog/catalog.
                                        asp?ISBN=1-55860-671-8.

Colombet:1993:SMI

[CDM93] L. Colombet, L. Desbat, and F. Menard. Star modeling on IBM RS6000 networks using PVM. In IEEE [IEE93c], pages 121–128. ISBN 0-8186-3900-8, 0-8186-3901-6. LCCN QA76.9.D5I593 1993. IEEE catalog no. 93TH0550-4.

Casanova:2015:SMA

[CDMS15] Henri Casanova, Frédéric Desprez, George S. Markomanolis, and Frédéric Suter. Simulation of MPI applications with time-independent traces. Concurrency and Computation: Practice and Experience, 27(5):1145–1168, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cotronis:2011:RAM

[CDND11] Yiannis Cotronis, Anthony Danalis, Dimitrios S. Nikolopoulos, and Jack Dongarra, editors. Recent Advances in the Message Passing Interface: 18th European MPI Users’ Group Meeting, EuroMPI 2011, Santorini, Greece, September 18–21, 2011. Proceedings, volume 6960 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CO-DEN LNCSD9. ISBN 3-642-24448-3 (print), 3-642-24449-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:/

/www.springerlink.com/
content/978-3-642-24449-
0.

Cabral:2020:EMO

[CdOO^{+}20] Frederico L. Cabral, Sanderson L. Gonzaga de Oliveira, Carla Osthoff, Gabriel P. Costa, Diego N. Brandão, and Mauricio Kischinhevsky. An evaluation of MPI and OpenMP paradigms in finite-difference explicit methods for PDEs on shared-memory multi- and manycore systems. Concurrency and Computation: Practice and Experience, 32 (20):e5642:1–e5642:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chaussumier:1999:ACM

[CDP99] F. Chaussumier, F. Desprez, and L. Prylli. Asynchronous communications in MPI — the BIP/Myrinet approach. In Dongarra et al. [DLM99], pages 485–492. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Coll:2003:SHB

[CDPM03] Salvador Coll, Jose Duato, Fabrizio Petrini, and Francisco J. Mora. Scalable hardware-based multicast trees. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN ???? URL http:/

                                        /www.sc-conference.org/
                                        sc2003/inter_cal/inter_
                                        cal_detail.php?eventid=
                                        10702#2;   http://www.
                                        sc-conference.org/sc2003/
                                        paperpdfs/pap300.pdf.

Chan:2005:CCI

[CDT05] Albert Chan, Frank Dehne, and Ryan Taylor. CGM-GRAPH/CGMLIB: Implementing and testing CGM graph algorithms on PC clusters and shared memory machines. The International Journal of High Performance Computing Applications, 19(1):81–97, Spring 2005. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

                                        tronic). URL http://hpc.

sagepub.com/content/19/
1/81.full.pdf+html.

Ceron:1998:PID

[CDZ^{+}98] C. Ceron, J. Dopazo, E. L.

Zapata, J. M. Carazo, and O. Trelles. Parallel implementation of DNAml program on message-passing architectures. Parallel Computing, 24(5–6):701–716, June 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/24/
5-6/1279.pdf.

Cappello:2000:MVM

[CE00] Franck Cappello and Daniel Etiemble. MPI versus MPI+OpenMP on the IBM SP for the NAS Benchmarks. In ACM [ACM00],

page 51. URL http://www.
sc2000.org/proceedings/
techpapr/papers/pap214.
pdf.

Clemencon:1995:AEP

[CEF^{+}95] C. Clemencon, A. Endo, J. Fritscher, A. Muller, R. Ruhl, and B. J. N. Wylie. The ’annai’ environment for portable distributed parallel programming. In El-Rewini and Shriver [ERS95], pages 242–251 (vol. 2). ISBN 0-8186-6935-7. LCCN ????

Chau:2007:MIP

[CEGS07] Ming Chau, Didier El Baz, Ronan Guivarch, and Pierre Spiteri. MPI implementation of parallel subdomain methods for linear and nonlinear convection–diffusion problems. Journal of Parallel and Distributed Computing, 67(5):581–591, May 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cerin:1999:DMP

[Cer99] C. Cerin. Differentiating message passing interface and bulk synchronous parallel computation models. Lecture Notes in Computer Science, 1662:477–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:2001:FFT

[CF01] Qun Chen and Michael C. Ferris. FATCOP: a fault tolerant Condor–PVM mixed integer programming solver. SIAM Journal on Optimization, 11(4):1019–1036, March/May 2001. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/35391.

Crivellini:2019:OPS

[CF19] Andrea Crivellini and Matteo Franciolini. OpenMP parallelization strategies for a discontinuous Galerkin solver. International Journal of Parallel Programming, 47(5–6):838–873, December 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/s10766-
018-0589-3.

Chen:2001:TMK

[CFDL01] Yu Chen, Qian Fang, Zhihui Du, and Sanli Li. TH-MPI: OS kernel integrated fault tolerant MPI. Lecture Notes in Computer Science, 2131:75–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310075.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310075.
pdf.

Choudhary:1994:LCR

[CFF^{+}94] Alok Choudhary, Ian Foster, Geoffrey Fox, Ken Kennedy, Carl Kesselman, Charles Koelbel, Joel Saltz, and Marc Snir. Languages, compilers, and runtime systems support for parallel inputoutput, 1994. URL http:

//www.ccsf.caltech.edu/
SIO/SIO.html.  Scalable

I/O Initiative Working Paper Number 3. On WWW at

http://www.ccsf.caltech.
edu/SIO/SIO.html.

Corbett:1996:OMP

[CFF^{+}96] P. Corbett, D. Feitelson, S. Fineberg, Yarsun Hsu, B. Nitzberg, J.-P. Prost, M. Snir, B. Traversat, and Parkson Wong. Overview of the MPI-IO parallel I/O interface. In Jain et al. [JWB96], pages 127–146. ISBN 0-7923-9735-5. LCCN QA76.58.I485 1996.

Clauser:2019:FFO

[CFF19] C. F. Clauser, R. Farengo, and H. E. Ferrari. FO-CUS: a full-orbit CUDA solver for particle simulations in magnetized plasmas. Computer Physics Communications, 234(??): 126–136, January 2019. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465518302753.

Carpenter:2000:OSM

[CFKL00] Bryan Carpenter, Geoffrey Fox, Sung Hoon Ko, and Sang Lim. Object serialization for marshaling data in a Java interface to MPI. Concurrency: practice and experience, 12(7): 539–553, May 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/72516217/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516217&PLACEBO=IE.
pdf.

Clemencon:1995:IRD

[CFMR95] C. Clemencon, J. Fritscher, M. J. Meehan, and R. Rühl. An implementation of race detection and deterministic replay with MPI. In Haridi et al. [HAM95b], pages 155–166. ISBN 3-540-60247-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I553 1995.

Cotronis:1996:ECP

[CFP96] J. Y. Cotronis, E. Floros, and N. Papazis. Efficient composition of PVM programs. In Liddell et al. [LCHS96], pages 919–?? ISBN 3-540-61142-8 (paperback). LCCN QA76.88 .H52 1996.

Clematis:1995:PPH

[CFPS95] A. Clematis, B. Falcidieno, D. F. Prieto, and M. Spagnuolo. Parallel processing on heterogeneous networks for GIS applications. In Hertzberger and Serazzi [HS95a], pages 67–72. ISBN 3-540-59393-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Chandrasekharan:1993:RTB

[CG93] N. Chandrasekharan and V. Goel. Ray tracing and binary tree computations using PVM. In Mudge et al. [MMH93], pages 104–105 (vol. 2). ISBN 0-8186-3230-5. LCCN ???? Four volumes. IEEE catalog number 93TH0501-7.

Clematis:1996:CEP

[CG96] A. Clematis and V. Gianuzzi. CPVM — extending PVM for consistent checkpointing. In IEEE [IEE96g], pages 67–76. ISBN 0-8186-7376-1. LCCN QA76.58 .E97 1996. IEEE order number PR07376.

Clematis:1999:EPC

[CG99a] A. Clematis and V. Gianuzzi. Extending PVM with consistent cut capabilities: Application aspects and implementation strategies. In Dongarra et al. [DLM99], pages 101–108. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Cownie:1999:SID

[CG99b] J. Cownie and W. Gropp. A standard interface for debugger access to message queue information in MPI. In Dongarra et al. [DLM99], pages 51–58. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Chaudhuri:2010:PIC

[CGB^{+}10] Pranay Chaudhuri, Sukumar Ghosh, Raj Kumar Buyya, Jian-Nong Cao, and Oeepak Oahiya, editors. Proceedings of the 2010 1st International Conference on Parallel Distributed and Grid Computing (PDGC), Jaypee University of Information Technology Waknaghat, Solan, HP, India, 28–30 October, 2010. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-7675-5. LCCN ????

Carretero:2015:AMM

[CGBS^{+}15] Jesus Carretero, Javier Garcia-Blas, David E. Singh, Florin Isaila, Alexey Lastovetsky, Thomas Fahringer, Radu Prodan, Peter Zangerl, Christi Symeonidou, Afshin Fassihi, and Horacio Pérez-Sánchez. Acceleration of MPI mechanisms for sustainable HPC applications. Supercomputing Frontiers and Innovations, 2(2):28–45, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic). URL http:/

/superfri.org/superfri/
article/view/35.

Calderon:2002:IMI

[CGC^{+}02] Alejandro Calderón, Félix García, Jesús Carretero, Jose M. Pérez, and Javier Fernández. An implementation of MPI-IO on expand: a parallel file system based on NFS servers. Lecture Notes in Computer Science, 2474:306–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

                                        /link.springer.de/link/
                                        service/series/0558/bibs/
                                        2474/24740306.htm;   http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2474/24740306.pdf.

Camp:2011:SIU

[CGC^{+}11] David Camp, Christoph Garth, Hank Childs, Dave Pugmire, and Kenneth I. Joy. Streamline integration using MPI-hybrid parallelism on a large multicore architecture. IEEE Transactions on Visualization and Computer Graphics, 17(11): 1702–1713, November 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Carter:2010:PLN

[CGG10] John D. Carter, William B. Gardner, and Gary Grewal. The Pilot library for novice MPI programmers. ACM SIGPLAN Notices, 45(5): 351–352, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clarke:1994:MMP

[CGH94] L. Clarke, I. Glendinning, and R. Hempel. The MPI Message Passing Interface Standard. In Decker and Rehmann [DR94], pages 213–218. ISBN 0-8176-5090-3 (Boston), 3-7643-5090-3 (Basel). LCCN QA76.58.P767 1994.

Cunningham:2014:RXE

[CGH^{+}14] David Cunningham, David Grove, Benjamin Herta, Arun Iyengar, Kiyokuni Kawachiya, Hiroki Murata, Vijay Saraswat, Mikio Takeuchi, and Olivier Tardieu. Resilient X10: efficient failure-aware programming. ACM SIGPLAN Notices, 49(8):67–80, August 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carpenter:2000:MML

[CGJ^{+}00] Bryan Carpenter, Vladimir Getov, Glenn Judd, Anthony Skjellum, and Geoffrey Fox. MPJ: MPIlike message passing for Java. Concurrency: practice and experience, 12(11):1019– 1038, September 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76000188/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76000188&PLACEBO=IE.
pdf.

Catanzaro:2011:CCE

[CGK11] Bryan Catanzaro, Michael Garland, and Kurt Keutzer. Copperhead: compiling an embedded data parallel language. ACM SIGPLAN Notices, 46(8):47–56, August 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’11 Conference proceedings.

Calore:2016:PPA

[CGK^{+}16] Enrico Calore, Alessandro Gabbana, Jiri Kraus, Sebastiano Fabio Schifano, and Raffaele Tripiccione. Performance and portability of accelerated lattice Boltzmann applications with OpenACC. Concurrency and Computation: Practice and Experience, 28(12):3485–3502, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chapman:2011:OPE

[CGKM11] Barbara M. Chapman, William D. Gropp, Kalyan Kumaran, and Matthias S. Müller, editors. OpenMP in the Petascale Era: 7th International Workshop on OpenMP, IWOMP 2011, Chicago, IL, USA, June 13– 15, 2011. Proceedings, volume 6665 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CO-DEN LNCSD9. ISBN 3-642-21486-X (print), 3-642-21487-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:/

/www.springerlink.com/
content/978-3-642-21487-
5.

Chatterjee:1993:GLA

[CGL^{+}93] S. Chatterjee, J. R. Gilbert, F. J. E. Long, R. Schreiber, and S.-H. Teng. Generating local addresses and communication sets for dataparallel programs. ACM SIGPLAN Notices, 28(7): 149–158, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Caubet:2001:DTM

[CGLD01] Jordi Caubet, Judit Gimenez, Jesus Labarta, and Luiz DeRose. A dynamic tracing mechanism for performance analysis of OpenMP applications. Lecture Notes in Computer Science, 2104: 53–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2104/21040053.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2104/21040053.
pdf.

Chan:1998:PCT

[CGPR98] K. J. Chan, A. M. Gibbons, M. Pias, and W. Rytter. On the PVM computations of transitive closure and algebraic path problems. Lecture Notes in Computer Science, 1497:338–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casanova:2015:TMS

[CGS15] Henri Casanova, Anshul Gupta, and Frédéric Suter. Toward more scalable offline simulations of MPI applications. Parallel Processing Letters, 25(3):1541002, September 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Cecilia:2012:CSC

[CGU12] José María Cecilia, José Manuel García, and Manuel Ujaldón. CUDA 2D stencil computations for the Jacobi method. Lecture Notes in Computer Science, 7133: 173–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-28151-8_
17/.

Chen:2013:IRM

[CGZQ13] Zhezhe Chen, Qi Gao, Wenbin Zhang, and Feng Qin. Improving the reliability of MPI libraries via message flow checking. IEEE Transactions on Parallel and Distributed Systems, 24(3):535– 549, March 2013. CODEN ITDSEO. ISSN 1045-9219.

Cheng:1994:PDP

[CH94] D. Cheng and R. Hood. A portable debugger for parallel and distributed programs. In IEEE [IEE94h], pages 723–732. ISBN 0-8186-6607-2, 0-8186-6605-6, 0-8186-6606-4. ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Ciancarini:1996:CLM

[CH96] Paolo Ciancarini and Chris Hankin, editors. Coordination languages and models: First International Conference COORDINATION ’96, Cesena, Italy, April 15–17, 1996: proceedings, number 1061 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61052-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I52 1996.

Charny:1996:MPV

[Cha96] B. Charny. Matrix partitioning on a virtual shared memory parallel machine. IEEE Transactions on Parallel and Distributed Systems, 7(4):343–355, April 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chapman:2002:PAD

[Cha02] Barbara Chapman. Parallel application development with the hybrid MPI + OpenMP programming model. Lecture Notes in Computer Science, 2474: 13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.de/link/
service/series/0558/bibs/
2474/24740013.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740013.pdf.

Chapman:2005:SMP

[Cha05] Barbara M. Chapman, editor. Shared memory parallel programming with OpenMP: 5th International Workshop on OpenMP Applications and Tools, WOMPAT 2004, Houston, TX, USA, May 17–18, 2004: Revised selected papers, volume 3349 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CO-DEN LNCSD9. ISBN 3-540-24560-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76 .A1 L42 NO.3349. URL http:

//www.springerlink.com/
openurl.asp?genre=issue&
issn=0302-9743&volume=
3349;   http://www.springerlink.
com/openurl.asp?genre=

volume&id=doi:10.1007/
b105895.

Cappello:2007:RAP

[CHD07] Franck Cappello, Thomas Herault, and Jack Dongarra, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 14th European PVM/ MPI User’s Group Meeting, Paris, France, September 30 — October 3, 2007. Proceedings, volume 4757 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CO-DEN LNCSD9. ISBN 3-540-75415-6 (print), 3-540-75416-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:/

/www.springerlink.com/
content/978-3-540-75416-
9.

Cappello:2009:FSI

[CHD09] Franck Cappello, Thomas Herault, and Jack Dongarra. Foreword: Special issue: selected papers from the 14th European PVM/ MPI Users Group Meeting. Parallel Computing, 35(12): 571, 2009. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Held in Paris, September 30–October 3,

Chergui:1999:UPP

[Che99] J. Chergui. Using PMD to parallel solve large-scale Navier–Stokes equations. performance analysis on SGI/CRAY-T3E machine. In Dongarra et al. [DLM99], pages 341–348. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Cheng:2010:BRBb

[Che10]  Jie Cheng.  Book re-

view: CUDA by Example: An Introduction to General-Purpose GPU Programming, by Jason Sanders and Edward Kandrot, ISBN-13 978-0-13-138768-3. Scalable Computing: Practice and Experience, 11 (4):401, December 2010. CODEN ???? ISSN 1895-1767. URL http://

                                        www.scpe.org/index.php/
                                        scpe/article/view/663.

See [SK10].

Cho:2015:OAO

[CHKK15] Myeongjin Cho, Youngsun Han, Minseong Kim, and Seon Wook Kim. O2WebCL: an automatic OpenCL-to-WebCL translator for high performance web computing. The Journal of Supercomputing, 71(6):2050– 2065, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/s11227-014-1260-4.

Chapman:2001:PDE

[CHPP01] B. Chapman, O. Hernandez, A. Patil, and A. Prabhakar. Program development environment for OpenMP programs on cc-NUMA architectures. Lecture Notes in Computer Science, 2179:210–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2179/21790210.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2179/21790210.
pdf.

Cho:2010:OPP

[CIJ^{+}10] S. M. Cho, D. W. Im, O. Y.

Jang, H. J. Song, B. D. Paulovicks, V. Sheinin, and H. Yeo. OpenCL and parallel primitives for digital TV applications. IBM Journal of Research and Development, 54(5):7:1–7:14, ???? 2010. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Cook:1995:TAS

[CJNW95] B. M. Cook, M. R. Jane, P. Nixon, and P. M. Welch, editors. Transputer Applications and Systems ’95. Proceedings of the 1995 World Transputer Congress, 4–6 September 1995, Harrogate, North Yorkshire, UK. IOS Press, Postal Drawer 10558, Burke, VA 2209-0558, USA, 1995. ISBN 90-5199-235-1 (IOS Press), 4-274-90062-2 (Ohmsha). LCCN ????

Cadenelli:2019:CUO

[CJPC19] Nicola Cadenelli, Zoran Jakšić, Jordà Polo, and David Carrera. Considerations in using OpenCL on GPUs and FPGAs for throughput-oriented genomics workloads. Future Generation Computer Systems, 94(??):148–159, May 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314183.

Chapman:2008:UOP

[CJvdP08] Barbara Chapman, Gabriele Jost, and Ruud van der Pas. Using OpenMP: portable shared memory parallel programming. Scientific and engineering computation. MIT Press, Cambridge, MA, USA, 2008. ISBN 0-262-03377-1 (hardcover), 0-262-53302-2 (paperback). xxii + 353 pp. LCCN QA76.642 .C49 2008. URL http://

www.loc.gov/catdir/toc/
ecip0721/2007026656.html.

Czarnul:1999:DAP

[CK99] P. Czarnul and H. Krawczyk.

Dynamic assignment with process migration in distributed environments. In Dongarra et al. [DLM99], pages 509–516. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Chang:2016:DLD

[CKmWH16] Li-Wen Chang, Hee-Seok Kim, and Wen mei W. Hwu. DySel: Lightweight dynamic selection for kernelbased data-parallel programming model. ACM SIG-PLAN Notices, 51(4):667– 680, April 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Casas:1994:ALM

[CKO^{+}94] J. Casas, R. Konuru, S. W. Otto, R. Prouty, and J. Walpole. Adaptive load migration systems for PVM. In IEEE [IEE94h], pages 390–399. ISBN 0-8186-6607-2, 0-8186-6605-6, 0-8186-6606-4. ISSN 1063-9535. LCCN QA76.5 .S894 1994. URL http:

//sc94.ameslab.gov/AP/
contents.html. IEEE cata-

log number 94CH34819.

Culler:1993:LTR

[CKP^{+}93] David E. Culler, Richard M.

Karp, David A. Patterson, Abhijit Sahay, Klaus E. Schauser, Eunice Santos, Ramesh Subramonian, and Thorsten von Eicken. LogP: towards a realistic model of parallel computation. ACM SIGPLAN Notices, 28(7): 1–12, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Castro-Leon:1993:MCP

[CL93] E. Castro-Leon. A model of computation with parallel solvers. In Anonymous [Ano93g], pages 189– 198. ISBN ???? LCCN ????

Clark:1998:FOP

[Cla98]  David Clark.  Focus:

OpenMP: a parallel standard for the masses. IEEE Concurrency, 6(1):10–12, January/March 1998. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

                                        //dlib.computer.org/pd/
                                        books/pd1998/pdf/p1010.
                                        pdf.

Chikin:2019:MAA

[CLA^{+}19] Artem Chikin, Taylor Lloyd, José Nelson Amaral, Ettore Tiotto, and Muhammad Usman. Memory-access-aware safety and profitability analysis for transformation of accelerator-bound OpenMP loops. ACM Transactions on Architecture and Code Optimization, 16(3): 30:1–30:??, July 2019. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Cornelis:2017:HAV

[CLBS17] Jan G. Cornelis, Jan Lemeire, Tim Bruylants, and Peter Schelkens. Heterogeneous acceleration of volumetric JPEG 2000 using OpenCL. The International Journal of High Performance Computing Applications, 31(3):229–245, 2017. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

//journals.sagepub.com/
doi/full/10.1177/1094342016646438.

Chabbi:2015:BEP

[CLdJ^{+}15] Milind Chabbi, Wim Lavrijsen, Wibe de Jong, Koushik Sen, John Mellor-Crummey, and Costin Iancu. Barrier elision for production parallel programs. ACM SIG-PLAN Notices, 50(8):109– 119, August 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chakraborty:2020:ESE

[CLE^{+}20] Sourav Chakraborty, Ignacio Laguna, Murali Emani, Kathryn Mohror, Dhabaleswar K. Panda, Martin Schulz, and Hari Subramoni. ER einit: Scalable and efficient fault-tolerance for bulk-synchronous MPI applications. Concurrency and Computation: Practice and Experience, 32(3):e4863:1– e4863:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2003:GMD

[CLL03] L. Chen, C. LiWang, and F. C. M. Lau. A grid middleware for distributed Java computing with MPI binding and process migration supports. Journal of computer science and technology, 18(4):505–514, 2003. CO-DEN JCTEEM. ISSN 1000-9000.

Corbacho-Lozano:1999:EDD

[CLLASPDP99] J. Corbacho-Lozano, O.- I. Lepe-Aldama, J. Sole-Pareta, and J. Domingo-Pascual. Experiences deploying a distributed parallel processing environment over a broadband multiservice network. In Dongarra et al. [DLM99], pages 477– 484. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Cantoni:1995:CCA

[CLM^{+}95] Virginio Cantoni, L. Lombardi, M. Mosconi, M. Savini, and A. Setti, editors. CAMP ’95, computer architectures for machine perception: proceedings, September 18–20, 1995, Como, Italy. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7134-3. LCCN QA76.9.A73W675 1995. IEEE catalog no. 95TB8093.

Chen:2018:FOB

[CLOL18] Cen Chen, Kenli Li, Aijia Ouyang, and Keqin Li. FlinkCL: An OpenCLbased in-memory computing architecture on heterogeneous CPU–GPU clusters for big data. IEEE Transactions on Computers, 67 (12):1765–1779, ???? 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/
document/8362980/.

Chien:1999:DEH

[CLP^{+}99] A. Chien, M. Lauria, R. Pennington, M. Showerman, G. Iannello, M. Buchanan, K. Connelly, L. Giannini, G. Koenig, S. Krishnamurthy, Q. Liu, S. Pakin, and G. Sampemane. Design and evaluation of an HPVMbased Windows NT supercomputer. The International Journal of High Performance Computing Applications, 13(3):201–219, Fall 1999. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Chandra:2007:ESP

[CLSP07] Sumir Chandra, Xiaolin Li, Taher Saif, and Manish Parashar. Enabling scalable parallel implementations of structured adaptive mesh refinement applications. The Journal of Supercomputing, 39(2): 177–203, February 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0920-8542&volume=
                                        39&issue=2&spage=177.

Chang:2016:APC

[CLYC16] Chih-Hung Chang, Chih-Wei Lu, Chao-Tung Yang, and Tzu-Chieh Chang. An approach of performance comparisons with OpenMP and CUDA parallel programming on multicore systems. Concurrency and Computation: Practice and Experience, 28(16):4230– 4245, November 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chapman:1998:OHI

[CM98] B. Chapman and P. Mehrotra. OpenMP and HPF: Integrating two paradigms. Lecture Notes in Computer Science, 1470:650–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chapman:2005:O

[CM05] Barbara M. Chapman and Federico Massaioli. OpenMP. Parallel Computing, 31 (10–12):957–959, October/ December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Claver:1999:PCS

[CMH99] J. M. Claver, M. Mollar, and V. Hernandez. Parallel computation of the SVD of a matrix product. In Dongarra et al. [DLM99], pages 388–395. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Cahir:2000:PMM

[CMK00] Margaret Cahir, Robert Moench, and Alice E. Koniges. Programming models and methods. In Koniges [Kon00], chapter 3, pages 27–54. ISBN 1-55860-540-1. LCCN QA76.58 .I483 2000. Discusses PVM, MPI, SHMEM, High-Performance Fortran, and POSIX threads.

Corbalan:2004:PMD

[CML04] Julita Corbalan, Xavier Martorell, and Jesus Labarta. Page migration with dynamic space-sharing scheduling policies: The case of the SGI O2000. International Journal of Parallel Programming, 32(4):263– 288, August 2004. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
32&issue=4&spage=263.

Carson:2003:CGU

[CMM03] Brett Carson, Robert Murison, and Ian A. Mason. Computational gains using RPVM on a Beowulf cluster. R News: the Newsletter of the R Project, 3(1): 21–26, June 2003. CO-DEN ???? ISSN 1609-

3631. URL http://CRAN.R-
project.org/doc/Rnews/.

Chapman:2012:OHW

[CMMR12] Barbara M. Chapman, Federico Massaioli, Matthias S. Müller, and Marco Rorro, editors. OpenMP in a Heterogeneous World: 8th International Workshop on OpenMP, IWOMP 2012, Rome, Italy, June 11–13, 2012. Proceedings, volume 7312 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CO-DEN LNCSD9. ISBN 3-642-30960-7 (print), 3-642-30961-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:/

/www.springerlink.com/
content/978-3-642-30961-
8.

Campanai:1994:EAS

[CMV^{+}94] M. Campanai, O. Morales, S. Viti, R. Trotta, P. Viliani, and M. Lo Moro. Experiences assessing software testing activities: the adoption of PVM, a prediction and validation model. In Anonymous [Ano94i], pages 491–500. ISBN 3-7281-2153-3. LCCN ????

Chapman:1999:EOF

[CMZ99] B. Chapman, P. Mehrotra, and H. Zima. Enhancing OpenMP with features for locality control. In ????, editor, Proceedings of Eighth ECMWF Workshop on the Use of Parallel Processors in Meteorology. Towards Teracomputing, pages 301–313. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1999.

Chou:2010:CMI

[CNC10] Yu-Cheng Chou, Stephen S.

Nestinger, and Harry H. Cheng. Ch MPI: Interpretive parallel computing in C. Computing in Science and Engineering, 12(2): 54–67, March/April 2010. CODEN CSENFA. ISSN 0740-7475 (print), 1558-1918 (electronic).

Chalkidis:2011:HPH

[CNM11] Georgios Chalkidis, Masao Nagasaki, and Satoru Miyano. High performance hybrid functional Petri net simulations of biological pathway models on CUDA. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 8(6): 1545–1556, November 2011. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Coelho:1994:EHC

[Coe94] F. Coelho. Experiments with HPF compilation for a network of workstations. In Gentzsch and Harms [GH94], pages 423– 428. ISBN 0-387-57981-8 (New York), 3-540-57981-8 (Berlin). LCCN QA76.88.I57 1994. DM96.00. Two volumes.

Cho:2020:PMP

[COE20] Y. Cho, S. Oh, and B. Egger. Performance modeling of parallel loops on multi-socket platforms using queueing systems. IEEE Transactions on Parallel and Distributed Systems, 31(2): 318–331, February 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cooperman:1995:SBP

[Coo95a] G. Cooperman. STAR/MPI: binding a parallel library to interactive symbolic algebra systems. In Levelt [Lev95], pages 126–132. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995.

Cooperman:1995:SMB

[Coo95b] Gene Cooperman. STAR/ MPI: Binding a parallel library to interactive symbolic algebra systems. In Levelt [Lev95], pages 126–132. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995.

Cotronis:1997:MPP

[Cot97] J. Y. Cotronis. Messagepassing program development by ensemble. Lecture Notes in Computer Science, 1332:242–249, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cotronis:1998:DMP

[Cot98]  Y. Cotronis.  Develop-

ing message-passing applications on MPICH under ensemble. Lecture Notes in Computer Science, 1497: 145–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cotronis:2004:CMP

[Cot04] Yiannis Cotronis. Composition of Message Passing Interface applications over MPICH-G2. The International Journal of High Performance Computing Applications, 18(3):327–339, Fall 2004. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.

                                        sagepub.com/content/18/
                                        3/327.full.pdf+html.

Coussement:1993:PMO

[Cou93] G. Coussement. Parallelization of a mesh optimization code on a RS/ 6000 cluster. In Anonymous [Ano93f], pages 185–212. ISBN ???? ISSN 0254-6213. LCCN ????

Carvalho:1997:PCC

[CP97] L. M. R. Carvalho and J. M. L. M. Palma. Parallelization of a CFD code using PVM and domain decomposition techniques. Lecture Notes in Computer Science, 1215:247–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carissimi:1998:AEM

[CP98] A. Carissimi and M. Pasin. Athapascan: An experience on mixing MPI communications and threads. Lecture Notes in Computer Science, 1497:137–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cercos-Pita:2015:ANF

[CP15] J. L. Cercos-Pita. AQUAgpusph, a new free 3D SPH solver accelerated with OpenCL. Computer Physics Communications, 192(??): 295–312, July 2015. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465515000909.

Celik:2017:BET

[CPKG17] Ahmet Celik, Sreepathi Pai, Sarfraz Khurshid, and Milos Gligoric. Bounded exhaustive test-input generation on GPUs. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):94:1–94:??, October 2017. CODEN ???? ISSN 2475-1421.

Castello:2018:EIR

[CPM^{+}18] Adrián Castelló, Antonio J.

Peña, Rafael Mayo, Judit Planas, Enrique S. Quintana-Ortí, and Pavan Balaji. Exploring the interoperability of remote GPGPU virtualization using rCUDA and directive-based programming models. The Journal of Supercomputing, 74(11):5628–5642, November 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Corno:1995:PTA

[CPR^{+}95] F. Corno, P. Prinetto, M. Rebaudengo, M. Sonza Reorda, and E. Veiluva. A PVM tool for automatic test generation on parallel and distributed systems. In Hertzberger and Serazzi [HS95a], pages 39–44. ISBN 3-540-59393-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

ChassindeKergommeaux:1999:MER

[CRD99] J. Chassin de Kergommeaux, M. Ronsse, and K. De Bosschere. MPL0*: Efficient record/replay of nondeterministic features of message passing libraries. In Dongarra et al. [DLM99], pages 141–148. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Cappello:1999:PNB

[CRE99] F. Cappello, O. Richard, and D. Etiemble. Performance of the NAS benchmarks on a cluster of SMP PCs using a parallelization of the MPI programs with OpenMP. Lecture Notes in Computer Science, 1662:339–350, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cappello:2001:UPS

[CRE01] Franck Cappello, Olivier Richard, and Daniel Etiemble. Understanding performance of SMP clusters running MPI programs. Future Generation Computer Systems, 17(6):711– 720, April 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

19/19/45/33/30/abstract.
html.

Creel:2016:NJM

[Cre16]  Michael Creel.  A note

on Julia and MPI, with code examples. Computational Economics, 48 (3):??, 2016. CODEN CNOMEL. URL http://

link.springer.com/article/
10.1007/s10614-015-9516-
5.

Cores:2014:FAM

[CRGM14] Iván Cores, Gabriel Rodríguez, Patricia González, and María J. Martín. Failure avoidance in MPI applications using an applicationlevel approach. The Computer Journal, 57(1):100– 114, January 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/1/100.full.
pdf+html.

Cores:2016:ROM

[CRGM16] Iván Cores, Mónica Rodríguez, Patricia González, and María J. Martín. Reducing the overhead of an MPI application-level migration approach. Parallel Computing, 54(??):72–82, May 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819116000429.

Cores:2014:MAL

[CRM14] Iván Cores, Gabriel Rodríguez, and María J. Martín. Inmemory application-level checkpoint-based migration for MPI programs. The Journal of Supercomputing, 70(2):660–670, November 2014. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s11227-014-1120-2.

Ciampolini:1996:EPM

[CS96] A. Ciampolini and C. Stefanelli. Extending PVM to a massively parallel architecture. Future Generation Computer Systems, 12(1): 13–23, May 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Coole:2014:FFH

[CS14] James Coole and Greg Stitt. Fast, flexible highlevel synthesis from OpenCL using reconfiguration contexts. IEEE Micro, 34(1): 42–53, January/February 2014. CODEN IEMIDZ. ISSN 0272-1732.

Chen:2019:STG

[CS19] Yong Chen and Weijia Shang. Supernode transformation on GPGPUs. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 34(2):181–202, 2019. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Chetlur:1998:ALE

[CSAGR98] M. Chetlur, G. D. Sharma, N. Abu-Ghazaleh, and U. K. V. Rajasekaran. An active layer extension to MPI. Lecture Notes in Computer Science, 1497:97–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clement:1996:NPM

[CSC96] Mark J. Clement, Michael R.

Steed, and Phyllis E. Crandall. Network performance modeling for PVM clusters. In ACM [ACM96c], page ?? ISBN 0-89791-854-1. LCCN QA 76.88 S8573 1996. URL http://

www.supercomp.org/sc96/
proceedings/SC96PROC/CLEMENT/
INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Cavenaghi:1996:UPS

[CSPM^{+}96] M. A. Cavenaghi, R. Spolon, J. E. M. Perea-Martins, S. G. Domingues, and A. Garcia Neto. Using PVM in the simulation of a hybrid dataflow architecture. In Bode et al. [BDLS96], pages 343–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Carreira:1995:DEL

[CSS95] J. Carreira, L. Silva, and J. G. Silva. On the design of Eilean: a Linda-like library for MPI. In IEEE [IEE95j], pages 175–184. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Chevitarese:2012:STN

[CSV12] Daniel Salles Chevitarese, Dilza Szwarcman, and Marley Vellasco. Speeding up the training of neural networks with CUDA technology. Lecture Notes in Computer Science, 7267: 30–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/10.
                                        1007/978-3-642-29347-4_
                                        4/.

Ciegis:1997:NID

[CSW97] R. Ciegis, R. Sablinskas, and J. Wasniewski. Numerical integration on distributedmemory parallel systems. Lecture Notes in Computer Science, 1332:329–336, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciegis:1999:HDA

[CSW99] R. Ciegis, R. Sablinskas, and J. Wasniewski. Hyperrectangle distribution algorithm for parallel multidimensional numerical integration. In Dongarra et al. [DLM99], pages 275–282. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Calotoiu:2012:PID

[CSW12] Alexandru Calotoiu, Christian Siebert, and Felix Wolf. Pattern-independent detection of manual collectives in MPI programs. Lecture Notes in Computer Science, 7484:28–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-32820-6_
5/.

Cote:1994:PSA

[CT94a] J. Cote and S. J. Thomas.

Parallel semi-Lagrangian advection on the sphere using PVM. In Pierce and Regnier [PR94b], pages 470–477. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.58.S32 1994. IEEE catalog no. 94TH0637-9.

Cote:1994:PSL

[CT94b] J. Cote and S. J. Thomas.

Parallel semi-Lagrangian advection on the sphere using PVM. In Dekker et al. [DSZ94], pages 801–808. ISBN 0-444-81784-0. LCCN QA76.58.E98 1994.

Cotronis:2002:MMP

[CT02] Yiannis Cotronis and Zacharias Tsiatsoulis. Modular MPI and PVM components. Lecture Notes in Computer Science, 2474:252–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

                                        /link.springer.de/link/
                                        service/series/0558/bibs/
                                        2474/24740252.htm;   http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2474/24740252.pdf.

Chang:2013:PDS

[CT13] Yao-Lin Chang and I-Lun Tseng. A parallel dualscanline algorithm for partitioning parameterized 45degree polygons. ACM Transactions on Design Automation of Electronic Systems, 18(4):59:1–59:??, October 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chapp:2021:IDS

[CTBT21] Dylan Chapp, Nigel Tan, Sanjukta Bhowmick, and Michela Taufer. Identifying degree and sources of non-determinism in MPI applications via graph kernels. IEEE Transactions on Parallel and Distributed Systems, 32(12):2936–2952, December 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cotronis:2000:CMP

[CTK00] J. Y. Cotronis, Z. Tsiatsoulis, and C. Kouniakis. Composition of message passing applications on-demand. Lecture Notes in Computer Science, 1908: 192–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080192.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080192.
pdf.

Czarnul:2001:DPD

[CTK01] Pawel Czarnul, Karen Tomko, and Henryk Krawczyk. Dynamic partitioning of the divide-and-conquer scheme with migration in PVM environment. Lecture Notes in Computer Science, 2131: 174–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310174.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310174.
pdf.

Candel:2019:EMC

[CVPS19] F. Candel, A. Valero, S. Petit, and J. Sahuquillo. Efficient management of cache accesses to boost GPGPU memory subsystem performance. IEEE Transactions on Computers, 68(10): 1442–1454, October 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cao:2011:OMM

[CwCW^{+}11] Chao Cao, Yun wen Chen, Yuning Wu, Erik Deumens, and Hai-Ping Cheng. OPAL: a multiscale multicenter simulation package based on MPI-2 protocol. International Journal of Quantum Chemistry, 111(15):4020– 4029, December 2011. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chang:2020:ADI

[CWL^{+}20] Tyler H. Chang, Layne T. Watson, Thomas C. H. Lux, Ali R. Butt, Kirk W. Cameron, and Yili Hong. Algorithm 1012: DELAU-NAYSPARSE: Interpolation via a sparse subset of the Delaunay triangulation in medium to high dimensions. ACM Transactions on Mathematical Software, 46(4): 38:1–38:20, November 2020. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

dl.acm.org/doi/10.1145/
3422818.

Cui:2012:OOB

[CXB^{+}12] Zheng Cui, Lei Xia, Patrick G.

Bridges, Peter A. Dinda, and John R. Lange. Optimizing overlay-based virtual networking through optimistic interrupts and cutthrough forwarding. In Hollingsworth [Hol12], pages 99:1–99:?? ISBN 1-4673-0804-8. URL http:

//conferences.computer.
org/sc/2012/papers/1000a029.
pdf.

Cavender:1995:APN

[CZ95a] M. E. Cavender and Xiaodong Zhang. Asynchronous PVM network computing. In Bailey et al. [BBG^{+}95], pages 772–773. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Cavender:1995:SSA

[CZ95b] Mark E. Cavender and Xiaodong Zhang. Software support for asynchronous computing across networks. In IEEE [IEE95l], pages 376–382. CODEN PSICD2. ISBN 0-8186-7119-X. ISSN 0730-6512. LCCN QA 76.6 C6295 1995. IEEE catalog number 95CB35838.

Chengqing:1996:WIP

[CZ96] Ye Chengqing and Cui Zhenqian. The ways of improving parallel computing efficiency in PVM. Mini-Micro Systems, 17(4):12–16, April 1996. CODEN XWJXEH. ISSN 1000-1220.

Czarnul:2002:DTI

[Cza02] Pawel Czarnul. Development and tuning of irregular divide-and-conquer applications in DAMPVM/DAC. Lecture Notes in Computer Science, 2474:208–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/
service/series/0558/bibs/
2474/24740208.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740208.pdf.

Czarnul:2003:PTA

[Cza03] Pawel Czarnul. Programming, tuning and automatic parallelization of irregular divide-and-conquer applications in DAMPVM/DAC. The International Journal of High Performance Computing Applications, 17(1):77– 93, Spring 2003. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Czapinski:2013:EPM

[Cza13] Michal Czapiński. An effective Parallel Multistart Tabu Search for Quadratic Assignment Problem on CUDA platform. Journal of Parallel and Distributed Computing, 73(11): 1461–1468, November 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S074373151200175X.

Czech:2016:IPC

[Cze16] Zbigniew J. Czech. Introduction to Parallel Computing. Cambridge University Press, Cambridge, UK, 2016. ISBN 1-107-17439-2 (hardcover), 1-316-79583-7 (e-book). xvii + 354 pp. LCCN QA76.58 .C975 2016.

Chapman:2008:PPM

[CZG^{+}08] Barbara Chapman, Weiming Zheng, Guang R. Gao, Mitsuhisa Sato, Eduard Ayguadé, and Dongsheng Wang, editors. A Practical Programming Model for the Multi-Core Era: 3rd International Workshop on OpenMP, IWOMP 2007, Beijing, China, June 3–7, 2007 Proceedings, volume 4935 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CO-DEN LNCSD9. ISBN 3-540-69302-5 (print), 3-540-69303-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:/

/www.springerlink.com/

                                        content/978-3-540-69303-
                                        1.

Chen:2021:CCR

[CZP21] Genlang Chen, Jiajian Zhang, and Chaoyi Pang. CRState: checkpoint/restart of OpenCL program for inkernel applications. The Journal of Supercomputing, 77(6):5426–5467, June 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s11227-
                                        020-03460-2.

Dongarra:1991:UGP

[D^{+}91] Jack Dongarra et al. A Users’ Guide to PVM Parallel Virtual Machine. Oak Ridge National Laboratory, Knoxville, TN, USA, July 1991.

Dongarra:1995:HPC

[D^{+}95] J. J. Dongarra et al., editors. High performance computing: technology, methods, and applications (Advanced workshop, June 1994, Cetraro, Italy), volume 10 of Advances in Parallel Computing. Elsevier, Amsterdam, The Netherlands, 1995. ISBN 0-444-82163-5. ISSN 0927-5452. LCCN QA76.88.H55 1995.

Daberdaku:2019:ACT

[Dab19] Sebastian Daberdaku. Accelerating the computation of triangulated molecular surfaces with OpenMP. The Journal of Supercomputing, 75(7):3426–3470, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Dieguez:2019:TPR

[DAD19] Adrián P. Diéguez, Margarita Amor, and Ramón Doallo. Tree partitioning reduction: A new parallel partition method for solving tridiagonal systems. ACM Transactions on Mathematical Software, 45(3):31:1– 31:26, August 2019. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/citation.
cfm?id=3328731.

Dimov:1998:IMC

[DAK98] I. Dimov, V. Alexandrov, and A. Karaivanova. Implementation of Monte Carlo algorithms for eigenvalue problem using MPI. Lecture Notes in Computer Science, 1497:346–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dieguez:2018:SLP

[DALD18] Adrián Pérez Diéguez, Margarita Amor, Jacobo Lobeiras, and Ramón Doallo. Solving large problem sizes of indexdigit algorithms on GPU: FFT and tridiagonal system solvers. IEEE Transactions on Computers, 67 (1):86–101, January 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

                                        //ieeexplore.ieee.org/
                                        document/7970194/.

Danalis:2012:MCT

[Dan12] Anthony Danalis. MPI and compiler technology: a lovehate relationship. Lecture Notes in Computer Science, 7490:12–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/accesspage/
                                        chapter/10.1007/978-3-
                                        642-33518-1_4.

Darema:2001:SMP

[Dar01]  Frederica Darema.  The

SPMD model: Past, present and future. Lecture Notes in Computer Science, 2131: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2131/21310001.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2131/21310001.
                                        pdf.

Demidov:2013:PCO

[DARG13] Denis Demidov, Karsten Ahnert, Karl Rupp, and Peter Gottschling. Programming CUDA and OpenCL: a case study using modern C++ libraries. SIAM Journal on Scientific Computing, 35 (5):C453–C472, ???? 2013. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

deAndrade:2017:OFH

[dAT17] Douglas Coimbra de Andrade and Luís Gonzaga Trabasso. An OpenCL framework for high performance extraction of image features. Journal of Parallel and Distributed Computing, 109(??):75–88, November 2017. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0743731517301624.

Demuynck:1997:DOD

[DBA97] K. Demuynck, J. Broeckhove, and F. Arickx. Dynamic optimization of a distributed VR system by network-balancing. Lecture Notes in Computer Science, 1332:443–450, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dinan:2016:IEM

[DBB^{+}16] James Dinan, Pavan Balaji, Darius Buntinas, David Goodell, William Gropp, and Rajeev Thakur. An implementation and evaluation of the MPI 3.0 onesided communication interface. Concurrency and Computation: Practice and Experience, 28(17):4385–4404, December 10, 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dursun:2009:MPM

[DBK^{+}09] Hikmet Dursun, Kevin J. Barker, Darren J. Kerbyson, Scott Pakin, Richard Seymour, Rajiv K. Kalia, Aiichiro Nakano, and Priya Vashishta. An MPI performance monitoring interface for cell based compute nodes. Parallel Processing Letters, 19(4):535–552, December 2009. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Dotsenko:2011:ATF

[DBLG11] Yuri Dotsenko, Sara S. Baghsorkhi, Brandon Lloyd, and Naga K. Govindaraju. Auto-tuning of Fast Fourier Transform on graphics processors. ACM SIGPLAN Notices, 46(8):257–266, August 2011. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’11 Conference proceedings.

DiMartino:2001:WDS

[DBVF01] Beniamino Di Martino, Sergio Briguglio, Gregorio Vlad, and Giuliana Fogaccia. Workload decomposition strategies for shared memory parallel systems with OpenMP. Scientific Programming, 9(2–3):109–122, Spring–Summer 2001. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C5%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

DAgostino:2014:CAM

[DCD^{+}14] Daniele D’Agostino, Andrea Clematis, Sergio Decherchi, Walter Rocchia, Luciano Milanesi, and Ivan Merelli. CUDA accelerated molecular surface generation. Concurrency and Computation: Practice and Experience, 26 (10):1819–1831, July 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

daCunha:1993:PLA

[dCH93] R. D. da Cunha and T. Hopkins. Porting linear algebra subroutines from transputers to clusters of workstations. In Grebe et al. [GHH^{+}93], pages 660–667. ISBN 90-5199-140-1. LCCN ????

Dow:2002:CMA

[DCH02] Chyi-Ren Dow, Jong-Shin Chen, and Min-Chang Hsieh.

Checkpointing MPI applications on symmetric multiprocessor machines using SMPCkpt. The Journal of Systems and Software, 63(2): 137–150, August 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Didelot:2012:IMC

[DCPJ12] Sylvain Didelot, Patrick Carribault, Marc Pérache, and William Jalby. Improving MPI communication overlap with collaborative polling. Lecture Notes in Computer Science, 7490:37–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-33518-1_
9/.

Didelot:2014:IMC

[DCPJ14] Sylvain Didelot, Patrick Carribault, Marc Pérache, and William Jalby. Improving MPI communication overlap with collaborative polling. Computing, 96(4): 263–278, April 2014. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

link.springer.com/article/
10.1007/s00607-013-0327-
z.

delCuvillo:2006:LOC

[dCZG06] Juan del Cuvillo, Weirong Zhu, and Guang Gao. Landing OpenMP on Cyclops-64: an efficient mapping of OpenMP to a manycore system-on-a-chip. In ACM [ACM06b], pages 41– 50. ISBN 1-59593-302-6. ACM order number 104060.

Dozsa:2000:THL

[DDL00] Gábor Dózsa, Dániel Drótos, and Róbert Lovas. Translation of a high-level graphical code to message-passing primitives in the GRADE programming environment. Lecture Notes in Computer Science, 1908:258–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080258.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080258.
pdf.

Decker:1995:TDU

[DDLM95] T. Decker, R. Diekmann, R. Luling, and B. Monien. Towards developing universal dynamic mapping algorithms. In IEEE [IEE95g], pages 456–459. ISBN 0-8186-7195-5. LCCN QA 76.58 I42 1995. IEEE catalog number 95TB8131.

Dichev:2022:PLR

[DDN^{+}22] Kiril Dichev, Daniele De Sensi, Dimitrios S. Nikolopoulos, Kirk W. Cameron, and Ivor Spence. Power Log n Roll: Power-efficient localized rollback for MPI applications using message logging protocols. IEEE Transactions on Parallel and Distributed Systems, 33 (6):1276–1288, June 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Deveci:2019:GMT

[DDP^{+}19] M. Deveci, K. D. Devine, K. Pedretti, M. A. Taylor, S. Rajamanickam, and Ü. V. Çatalyürek. Geometric mapping of tasks to processors on parallel computers with mesh or torus networks. IEEE Transactions on Parallel and Distributed Systems, 30(9):2018–2032, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dongarra:1997:BCA

[DDPR97] J. J. Dongarra, F. Desprez, A. Petitet, and C. Randriamaro. Block-cyclic array redistribution on networks of workstations. Lecture Notes in Computer Science, 1332:343–350, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dean:1994:CPV

[DDS^{+}94] C. E. Dean, R. C. Denny, P. C. Stephenson, G. J. Milne, and E. Pantos. Computing with parallel virtual machines. Journal de physique. IV, Colloque, 4 (C9):C9/445–448, November 1994. CODEN JPICEI. ISSN 1155-4339.

Dan:1999:QAM

[DDYM99] Pei Dan, Wang Dongsheng, Zhang Youhui, and Shen Meiming. Quasiasynchronous migration: a novel migration protocol for PVM tasks. Operating Systems Review, 33(2):5–14, April 1999. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Durand:1991:HPC

[DE91] M. Durand and F. El Dabaghi, editors. High performance computing, II: proceedings of the Second Symposium on High Performance Computing, Montpellier, France, 7–9 October, 1991. North-Holland, Amsterdam, The Netherlands, 1991. ISBN 0-444-89224-9. LCCN QA75.5.I585 1991.

Demaine:1996:FCC

[Dem96] E. Demaine. First class communication in MPI. In IEEE [IEE96i], pages 189– 194. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

DePasquale:2003:UJU

[DeP03] C. J. DePasquale. Using the JVMPI to understand the behavior of Java classes during the development process. Cmg, 2(??):821–832, 2003. CODEN ????

Dehne:2001:CPD

[DERC01] Frank Dehne, Todd Eavis, and Andrew Rau-Chaplin. Computing partial data cubes for parallel data warehousing applications. Lecture Notes in Computer Science, 2131:319–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310319.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310319.
pdf.

Dashti:2017:AMM

[DF17] Mohammad Dashti and Alexandra Fedorova. Analyzing memory management methods on integrated CPU–GPU systems. ACM SIGPLAN Notices, 52(9): 59–69, September 2017. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dalcin:2021:MSU

[DF21] Lisandro Dalcin and Yao-Lung L. Fang. mpi4py: Status update after 12 years of development. Computing in Science and Engineering, 23(4):47–54, July/ August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Duran:2009:PEO

[DFA^{+}09] Alejandro Duran, Roger Ferrer, Eduard Ayguadé, Rosa M. Badia, and Jesus Labarta. A proposal to extend the OpenMP tasking model with dependent tasks. International Journal of Parallel Programming, 37(3):292–305, June 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=3&spage=292.

Duran:2007:PEH

[DFC^{+}07] Alejandro Duran, Roger Ferrer, Juan José Costa, Marc Gonzàlez, Xavier Martorell, Eduard Ayguadé, and Jesús Labarta. A proposal for error handling in OpenMP. International Journal of Parallel Programming, 35 (4):393–416, August 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=4&spage=393.

Figueiredo:2019:MOP

[dFdOSR^{+}19] Marco Antonio C. de Figueiredo, Jr., Edans F. de Oliveira Sandes, Genaina N. Rodrigues, George L. M. Teodoro, and Alba Cristina M. A. de Melo. MASA-OpenCL: Parallel pruned comparison of long DNA sequences with OpenCL. Concurrency and Computation: Practice and Experience, 31(11):e5039:1– e5039:??, June 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Demaine:2001:GCM

[DFKS01] E. D. Demaine, I. Foster, C. Kesselman, and M. Snir. Generalized communicators in the message passing interface. IEEE Transactions on Parallel and Distributed Systems, 12(6):610– 616, June 2001. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://dlib.
computer.org/td/books/
td2001/pdf/l0610.pdf;
http://www.computer.org/
tpds/td2001/l0610abs.htm.

Deshpande:1994:ADN

[DFMD94] Manish Deshpande, Jinzhang Feng, Charles L. Merkle, and Ashish Deshpande. Application of a distributed network in computational fluid dynamic simulations. The International Journal of Supercomputer Applications, 8 (1):64–67, Spring 1994. CO-DEN IJSAE9. ISSN 0890-2720.

Diaz:2012:CCF

[DFN12] M. J. Castro Díaz and E. Fernández-Nieto. A class of computationally fast first order finite volume solvers: PVM methods. SIAM Journal on Scientific Computing, 34(4): A2173–A2196, ???? 2012. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Diaz:2019:AOO

[DFP^{+}19] Jose Monsalve Diaz, Kyle Friedline, Swaroop Pophale, Oscar Hernandez, David E. Bernholdt, and Sunita Chandrasekaran. Analysis of OpenMP 4.5 offloading in implementations: Correctness and overhead. Parallel Computing, 89(??):Article 102546, November 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819119301371.

Doulis:2019:CMP

[DFSW19] Georgios Doulis, Jörg Frauendiener, Chris Stevens, and Ben Whale. COFFEE — an MPI-parallelized Python package for the numerical evolution of differential equations. SoftwareX, 10 (??):Article 100283, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2352711019300950.

DAmbra:1995:CBC

[DG95] P. D’Ambra and G. Giunta. Concurrent banded Cholesky factorization on workstation networks using PVM. Parallel Computing, 21(3): 487–494, March 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dinan:2014:ECC

[DGB^{+}14] James Dinan, Ryan E. Grant, Pavan Balaji, David Goodell, Douglas Miller, Marc Snir, and Rajeev Thakur. Enabling communication concurrency through flexible MPI endpoints. The International Journal of High Performance Computing Applications, 28(4): 390–405, November 2014. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

                                        http://hpc.sagepub.com/
                                        content/28/4/390.

DiNapoli:1997:DCA

[DGF97] C. Di Napoli, M. Giordano, and M. M. Furnari. Distributed and cooperative applications in PVM. Lecture Notes in Computer Science, 1332:83–90, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dinan:2012:EMC

[DGG^{+}12] James Dinan, David Goodell, William Gropp, Rajeev Thakur, and Pavan Balaji. Efficient multithreaded context ID allocation in MPI. Lecture Notes in Computer Science, 7490:57–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-33518-1_
11/.

Dongarra:2019:PPL

[DGH^{+}19] Jack Dongarra, Mark Gates, Azzam Haidar, Jakub Kurzak, Piotr Luszczek, Panruo Wu, Ichitaro Yamazaki, Asim Yarkhan, Maksims Abalenkovs, Negin Bagherpour, Sven Hammarling, Jakub Sístek, David Stevens, Mawussi Zounon, and Samuel D. Relton. PLASMA: Parallel linear algebra software for multicore using OpenMP. ACM Transactions on Mathematical Software, 45(2): 16:1–16:35, April 2019. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/citation.
cfm?id=3264491.

deGloria:1994:TAS

[dGJM94] A. de Gloria, M. R. Jane, and D. Marini, editors.

Transputer Applications and Systems ’94. Proceedings of the 1994 World Transputer Congress. IOS Press, Postal Drawer 10558, Burke, VA 2209-0558, USA, 1994. ISBN ???? LCCN ????

Dongarra:1993:UPR

[DGMJ93] J. J. Dongarra, A. Geist, R. Manchek, and W. Jiang. Using PVM 3.0 to run grand challenge applications on a heterogeneous network of parallel computers. In Sincovec [Sin93], pages 873–877. ISBN 0-89871-315-3. LCCN QA 76.58 S55 1993. Two volumes.

Dongarra:1993:IPF

[DGMS93] Jack Dongarra, G. A. Geist, Robert Manchek, and V. S. Sunderam. Integrated PVM framework supports heterogeneous network computing. Computers in Physics, 7 (2):166–174 (or 166–175??), March–April 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

daCunha:1994:PIR

[dH94] Rudnei Dias da Cunha and Tim Hopkins. A parallel implementation of the restarted GMRES iterative algorithm for nonsymmetric systems of linear equations. Advances in computational mathematics, 2(3): 261–277, ???? 1994. CO-DEN ACMHEX. ISSN 1019-7168.

Dongarra:1995:PBC

[DH95] J. J. Dongarra and T. Hey.

The ParkBench benchmark collection. Supercomputer, 11(2-3):94–114, June 1995. CODEN SPCOEL. ISSN 0168-7875.

Du:2022:MPO

[DH22] Qi Du and Hui Huang.

MPI parameter optimization during debugging phase of HPC system. The Journal of Supercomputing, 78(2): 1696–1711, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

https://link.springer.
com/article/10.1007/s11227-
021-03939-6.

Dongarra:1992:PUL

[DHHW92] Jack J. Dongarra, Rolf Hempel, Anthony J. G. Hey, and David W. Walker. A proposal for a userlevel message-passing interface in a distributed memory environment. Technical Report TM-12231, Oak Ridge National Laboratory, Knoxville, TN, USA, October 1992.

Dongarra:1993:PUM

[DHHW93a] J. Dongarra, R. Hempel, A. Hay, and D. Walker. A proposal for a user-level message passing interface in a distributed memory environment. Technical Report ORNL/TM-12231, Oak Ridge National Laboratory, Knoxville, TN, USA, February 1993.

Dongarra:1993:DSM

[DHHW93b] J. J. Dongarra, R. Hempel, A. J. G. Hey, and D. W. Walker. A draft standard for message passing in a distributed memory environment. In Hoffmann and Kauranne [HK93], pages 465– 481. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Derakhshan:1997:PEP

[DHK97] M. Derakhshan, S. Hammarling, and A. Krommer. PINEAPL: a European project on Parallel Industrial Numerical Applications and Portable Libraries. Lecture Notes in Computer Science, 1332:337–342, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1997:CSD

[DHP97] J. J. Dongarra, S. Hammarling, and A. Petitet. Case studies on the development of ScaLAPACK and the NAG numerical PVM library. In Boisvert [Boi97], pages 236– 248. ISBN 0-412-80530-8. LCCN QA297 .I35 1996.

URL http://www.netlib.
org/utk/papers/woco96/

woco96.html;   http://
www.netlib.org/utk/papers/
woco96/woco96.ps;   http:
//www.netlib.org/utk/people/
JackDongarra/pdf/woco96.
pdf.

Dongarra:1996:SRP

[DHS96] J. J. Dongarra, T. Hey, and E. Strohmaier. Selected results from the PARK-BENCH benchmark. In Bouge et al. [BFMR96], pages 251–254. ISBN 3-540-61626-8 (vol. 1), 3-540-61627-6 (vol. 2). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I554 1996, QA267.A1 L43 no.1123-1124. Two volumes.

DiPierro:2014:PPP

[Di 14] Massimo Di Pierro. Portable parallel programs with Python and OpenCL. Computing in Science and Engineering, 16(1):34–40, January/ February 2014. CODEN CSENFA. ISSN 1521-9615.

DiSerio:2002:ENN

[DI02] Angela Di Serio and María B.

Ibáñez. Evaluation of a nearest-neighbor load balancing strategy for parallel molecular simulations in MPI environment. Lecture Notes in Computer Science, 2474:226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/

                                        service/series/0558/bibs/
                                        2474/24740226.htm;   http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2474/24740226.pdf.

DiNucci:1996:CDS

[DiN96]  D. C. DiNucci.  Co-

operative data sharing: a layered approach to an architecture-independent Message-Passing Interface. In IEEE [IEE96i], pages 58–

65. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Denis:2019:SPT

[DJJ^{+}19] Alexandre Denis, Julien Jaeger, Emmanuel Jeannot, Marc Pérache, and Hugo Taboada. Study on progress threads placement and dedicated cores for overlapping MPI nonblocking collectives on manycore processor. The International Journal of High Performance Computing Applications, 33(6):1240–1254, November 1, 2019. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:/

                                        /journals.sagepub.com/
                                        doi/full/10.1177/1094342019860184.

Karniadakis:2002:DLP

[DK02] Suchuan Dong and George Em. Karniadakis. Dual-level parallelism for deterministic and stochastic CFD problems. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL

http://www.sc-2002.org/
paperpdfs/pap.pap137.pdf.

Drosinos:2006:EPT

[DK06] Nikolaos Drosinos and Nectarios Koziris. The effect of process topology and load balancing on parallel programming models for SMP clusters and iterative algorithms. The Journal of Supercomputing, 35(1):65–91, January 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
35&issue=1&spage=65.

Deo:2013:PSA

[DK13] Mrinal Deo and Sean Keely.

Parallel suffix array and least common prefix for the GPU. ACM SIGPLAN Notices, 48 (8):197–206, August 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’13 Conference proceedings.

Davydov:2020:ADS

[DK20] Denis Davydov and Martin Kronbichler. Algorithms and data structures for matrix-free finite element operators with MPIparallel sparse multi-vectors. ACM Transactions on Parallel Computing (TOPC), 7 (3):20:1–20:30, August 2020. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3399736.

Diener:2020:HCO

[DKB20] Matthias Diener, Laxmikant V. Kale, and Daniel J. Bodony. Heterogeneous computing with OpenMP and Hydra. Concurrency and Computation: Practice and Experience, 32(20):e5728:1– e5728:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DiMartino:2005:RAP

[DKD05] Beniamino Di Martino, Dieter Kranzlmüller, and J. J. Dongarra, editors. Recent advances in parallel virtual machine and message passing interface: 12th European PVM/MPI User’s Group Meeting, Sorrento, Italy, September 18–21, 2005: proceedings, volume 3666 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CO-DEN LNCSD9. ISBN 3-540-29009-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E973 2005. URL http://

springerlink.metapress.
com/openurl.asp?genre=

issue&issn=0302-9743&volume=
3666.

DiMartino:2007:SIS

[DKD07] Beniamino Di Martino, Dieter Kranzlmüller, and Jack Dongarra. Special issue on selected papers from the EuroPVM/MPI 2005 Conference, Sorrento, Italy, 18-21 September 2005 — preface. The International Journal of High Performance Computing Applications, 21(2):129– 131, Summer 2007. ISSN 1094-3420 (print), 1741-2846 (electronic).

DiMartino:2008:SSG

[DKD08] Beniamino Di Martino, Dieter Kranzlmüller, and Jack Dongarra. Special section: Grid computing and the Message Passing Interface. Future Generation Computer Systems, 24 (2):119–120, February 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Damodaran-Kamal:1993:NTD

[DKF93] S. K. Damodaran-Kamal and J. M. Francioni. Nondeterminacy: testing and debugging in message passing parallel programs. ACM SIGPLAN Notices, 28(12): 118–128, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Damodaran-Kamal:1994:MSR

[DKF94a] S. K. Damodaran-Kamal and J. M. Francioni. mdb: a semantic race detection tool for PVM. In Pierce and Regnier [PR94b], pages 702–709. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.58.S32 1994. IEEE catalog no. 94TH0637-9.

Damodaran-Kamal:1994:TRP

[DKF94b] S. K. Damodaran-Kamal and J. M. Francioni. Testing races in parallel programs with an OtOt strategy. In Ostrand [Ost94]. CODEN SFENDP. ISBN 0-89791-683-2. ISSN 0163-5948. LCCN QA76.76.T48 I58 1994.

Dongarra:1992:PFS

[DKM^{+}92] J. Dongarra, P. Kennedy, P. Messina, D. C. Sorensen, and R. G. Voigt, editors. Proceedings of the Fifth SIAM Conference on Parallel Processing for Scientific Computing, 25–27 March 1991, Houston, TX, USA. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1992. ISBN 0-89871-303-X. LCCN QA76.58.P76 1992.

Dongarra:2000:RAP

[DKP00] J. J. Dongarra, Peter Kacsuk, and Norbert Podhorszki, editors. Recent advances in parallel virtual machine and message passing interface: 7th European PVM/MPI Users’ Group Meeting, Balatonfured, Hungary, September 10–13, 2000: proceedings, volume 1908 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-41010-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

Dickens:2010:HPI

[DL10] Phillip M. Dickens and Jeremy Logan. A high performance implementation of MPI-IO for a Lustre file system environment. Concurrency and Computation: Practice and Experience, 22 (11):1433–1449, August 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

delaAsuncion:2011:SOL

[dlAMC11] Marc de la Asunción, José M. Mantas, and Manuel J. Castro. Simulation of one-layer shallow water systems on multicore and CUDA architectures. The Journal of Supercomputing, 58(2):206– 214, November 2011. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&

                                        issn=0920-8542&volume=
                                        58&issue=2&spage=206.

delaAsuncion:2012:MCI

[dlAMCFN12] Marc de la Asunción, José M. Mantas, Manuel J. Castro, and E. D. Fernández-Nieto. An MPI-CUDA implementation of an improved Roe method for two-layer shallow water systems. Journal of Parallel and Distributed Computing, 72(9): 1065–1072, September 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S074373151100147X.

Desai:2007:CEM

[DLB07] Narayan Desai, Ewing Lusk, and Rick Bradshaw. A composition environment for MPI programs. The International Journal of High Performance Computing Applications, 21(2):166–173, May 2007. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

                                        tronic). URL http://hpc.
                                        sagepub.com/content/21/
                                        2/166.full.pdf+html.

Marcos:2002:DDP

[dlFMBdlFM02] Carlos de la Fuente Marcos, Pierre Barge, and Raúl de la Fuente Marcos. Dust dynamics in protoplanetary disks: Parallel computing with PVM. Journal of Computational Physics, 176(2):

276–294, March 1, 2002. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0021999101969785.

Deng:2019:CBV

[DLLZ19] Y. Deng, T. Li, Y. Luo, and X. Zhao. CUDA-based volume rendering and inspection for time-varying ultrasonic testing datasets. Computing in Science and Engineering, 21(5):76–86, September/October 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic). See corrections [DLLZ20].

Deng:2020:CCB

[DLLZ20] Y. Deng, T. Li, Y. Luo, and X. Zhao. Corrections to “CUDA-Based Volume Rendering and Inspection for Time-Varying Ultrasonic Testing Datasets”. Computing in Science and Engineering, 22(1):4, January/ February 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). See [DLLZ19].

Dongarra:1999:RAP

[DLM99] J. J. Dongarra, E. Luque, and Tomas Margalef, editors. Recent advances in parallel virtual machine and message passing interface: 6th European PVM/ MPI Users’ Group Meeting, Barcelona, Spain, September 26–29, 1999: proceedings, volume 1697 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Degomme:2017:SMA

[DLM^{+}17] Augustin Degomme, Arnaud Legrand, George S. Markomanolis, Martin Quinson, Mark Stillwell, and Frederic Suter. Simulating MPI applications: The SMPI approach. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2387–2400, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/08/07855780-
abs.html.

Dongarra:2003:RAP

[DLO03] Jack Dongarra, Domenico Laforenza, and Salvatore Orlando, editors. Recent advances in parallel virtual machine and message passing interface: 10th European PVM/MPI User’s group Meeting, Venice, Italy, September 29–October 2, 2003: Proceedings, volume 2840 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CO-DEN LNCSD9. ISBN 3-540-20149-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E973 2003. URL http:

//link.springer-ny.com/
link/service/series/0558/
tocs/t2840.htm.

DeKeyser:1994:RTL

[DLR94] J. DeKeyser, K. Lust, and D. Roose. Run-time load balancing support for a parallel multiblock Euler/ Navier–Stokes code with adaptive refinement on distributed memory computers. Parallel Computing, 20 (8):1069–1088, August 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lu:2004:AFS

[dLR04] Charng da Lu and Daniel A.

Reed. Assessing fault sensitivity in MPI applications. In ACM [ACM04], page 37. ISBN 0-7695-2153-3. LCCN ????

DeSande:1999:NBS

[DLRR99] F. De Sande, C. Leon, C. Rodriguez, and J. Roda. Nested bulk synchronous parallel computing. In Dongarra et al. [DLM99], pages 189– 198. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

DiPietro:2016:CLD

[DLV16] Roberto Di Pietro, Flavio Lombardi, and Antonio Villani. CUDA leaks: a detailed hack for CUDA and a (partial) fix. ACM Transactions on Embedded Computing Systems, 15(1): 15:1–15:??, February 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Despons:1993:CCP

[DM93] R. Despons and T. Muntean. Constructing correct protocols for a diffusion virtual machine in message passing parallel architectures. In Grebe et al. [GHH^{+}93], pages 465–480. ISBN 90-5199-140-1. LCCN ????

Davies:1995:NSP

[DM95a] G. Davies and N. Matloff. Network-specific performance enhancements for PVM. In IEEE [IEE95k], pages 205–210. ISBN 0-8186-7088-6. LCCN QA76.9.D5 I328 1995. IEEE catalog no. 95TB8075.

Davies:1995:NPE

[DM95b] Gregory Davies and Norman Matloff. Network-specific performance enhancements for PVM. In IEEE [IEE95k], pages 205–210. ISBN 0-8186-7088-6. LCCN QA76.9.D5 I328 1995. IEEE catalog no. 95TB8075.

Dagum:1998:OIS

[DM98] Leonardo Dagum and Ramesh Menon. OpenMP: An industry-standard API for shared-memory programming. IEEE Computational Science & Engineering, 5(1):46–55, January/ March 1998. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (elec-

tronic). URL http://dlib.
computer.org/cs/books/
cs1998/pdf/c1046.pdf;
http://www.computer.org/
cse/cs1998/c1046abs.htm.

Dziubak:2012:OOI

[DM12] Tomasz Dziubak and Jacek Matulewski. An objectoriented implementation of a solver of the time-dependent Schrödinger equation using the CUDA technology. Computer Physics Communications, 183(3): 800–812, March 2012. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465511003948.

Dathathri:2016:CAL

[DMB16] Roshan Dathathri, Ravi Teja Mullapudi, and Uday Bondhugula. Compiling affine loop nests for a dynamic scheduling runtime on shared and distributed memory.

ACM Transactions on Parallel Computing (TOPC), 3 (2):12:1–12:??, August 2016. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Dalcin:2019:FPM

[DMK19] Lisandro Dalcin, Mikael Mortensen, and David E. Keyes. Fast parallel multidimensional FFT using advanced MPI. Journal of Parallel and Distributed Computing, 128(??): 137–150, June 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S074373151830306X.

DiMartino:1997:IPD

[DMMV97] B. Di Martino, A. Mazzeo, N. Mazzocca, and U. Villano. Interaction patterns detection in PVM programs to support simulation. Lecture Notes in Computer Science, 1332:250–256, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1996:APC

[DMW96] Jack J. Dongarra, Kay Madsen, and Jerzy Wasniewski, editors. Applied parallel computing: computations in physics, chemistry, and engineering science: second international workshop, PARA ’95, Lyngby, Denmark, August 21–24, 1995: proceedings, volume 1041 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-60902-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.P35 1995.

Dinda:1996:PIA

[DO96] P. A. Dinda and D. R.

O’Hallaron. The performance impact of address relation caching. In Szymanski and Sinharoy [SS96], pages 213–226. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Donev:2006:ICF

[Don06] Aleksander Donev. Interoperability with C in Fortran 2003. ACM Fortran Forum, 25(1):8–12, April 2006. ISSN 1061-7264 (print), 1931-1311 (electronic).

Sandes:2016:CIS

[dOSMM^{+}16] Edans Flavius de Oliveira Sandes, Guillermo Miranda, Xavier Martorell, Eduard Ayguade, George Teodoro, and Alba Cristina Magalhaes Melo. CUDAlign 4.0: Incremental speculative traceback for exact chromosome-wide alignment in GPU clusters. IEEE Transactions on Parallel and Distributed Systems, 27(10): 2838–2850, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/10/07374729-
abs.html.

Dongarra:1995:IMS

[DOSW95] Jack Dongarra, Steve W. Otto, Marc Snir, and David Walker. An introduction to the MPI Standard. Technical report CS-95-274, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, January 1995. URL

http://www.netlib.org/
tennessee/ut-cs-95-274.
ps;   http://www.netlib.
org/utk/papers/intro-mpi/
intro-mpi.html;   http:
//www.netlib.org/utk/people/
JackDongarra/pdf/ut-cs-
95-274.pdf. Appears in

CACM [DOSW96].

Dongarra:1996:MPS

[DOSW96] Jack J. Dongarra, Steve W. Otto, Marc Snir, and David Walker. A message passing standard for MPP and workstations. Communications of the ACM, 39(7): 84–90, July 1996. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org/pubs/toc/
Abstracts/cacm/234000.
html.

DeRoeck:1994:CFP

[DP94] Y. H. De Roeck and R. E. Plessix. Combining F90 and PVM to construct synthetic seismograms by raytracing. In IEEE [IEE94c], pages II–653–II–658. ISBN 0-7803-2057-3, 0-7803-2056-5, 0-7803-2058-1. ISSN 0197-7385. LCCN TC 1505 O33197 1994. Three volumes. IEEE catalog no. 94CH3472-8.

Diep:2019:TSS

[DPFT19] Thanh-Dang Diep, Kien Trung Pham, Karl Fürlinger, and Nam Thoai. A timestamping system to detect memory consistency errors in MPI one-sided applications. Parallel Computing, 86(??):36–44, August 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819118303235.

Denis:2001:THP

[DPP01] Alexandre Denis, Christian Pérez, and Thierry Priol. Towards high performance CORBA and MPI middlewares for grid computing. Lecture Notes in Computer Science, 2242:14–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2242/22420014.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2242/22420014.
pdf.

Dalcin:2005:MP

[DPS05] Lisandro Dalcín, Rodrigo Paz, and Mario Storti. MPI for Python. Journal of Parallel and Distributed Computing, 65(9): 1108–1115, September 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dalcin:2008:MPP

[DPSD08] Lisandro Dalcín, Rodrigo Paz, Mario Storti, and Jorge D’Elía. MPI for Python: Performance improvements and MPI-2 extensions. Journal of Parallel and Distributed Computing, 68(5): 655–662, May 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dou:1997:ISV

[DPZ97] Yong Dou, Zhengbing Pang, and Xingming Zhou. Implementing a software virtual shared memory on PVM. In IEEE [IEE97a]. ISBN 0-8186-7876-3 (paperback and case), 0-8186-7878-X (microfiche). LCCN QA76.58 .A4 1997.

Decker:1994:PEM

[DR94] K. M. (Karsten M.) Decker and R. M. (Rene M.) Rehmann, editors. Programming environments for massively parallel distributed systems: working conference of the IFIP WG10.3, April 25–29, 1994, Ascona, Italy. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1994. ISBN 0-8176-5090-3 (Boston), 3-7643-5090-3 (Basel). LCCN QA76.58.P767 1994.

Dowaji:1995:LBS

[DR95] S. Dowaji and C. Roucairol.

Load balancing strategy and priority of tasks in distributed environments. In IEEE [IEE95b], pages 15–

22. ISBN 0-7803-2493-5, 0-7803-2492-7, 0-7803-2494-

3. LCCN TK7885.A1 I567 1995. IEEE catalog no. 95CH35751.

DiMartino:1997:MDH

[DR97] V. Di Martino and G. Ruocco.

Molecular dynamics on hybrid memory machines. Lecture Notes in Computer Science, 1332:451–456, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davina:2018:MCP

[DR18] A. Lamas Daviña and J. E.

Roman. MPI-CUDA parallel linear solvers for blocktridiagonal matrices in the context of SLEPc’s eigensolvers. Parallel Computing, 74(??):118–135, ???? 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819117301874.

Deuzeman:2012:LMP

[DRUE12] Albert Deuzeman, Siebren Reker, Carsten Urbach, and ETM Collaboration. Lemon: An MPI parallel I/O library for data encapsulation using LIME. Computer Physics Communications, 183(6): 1321–1335, June 2012. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465512000318.

Deshpande:1996:MIBb

[DS96a] V. Deshpande and W. Sawyer. An MPI implementation of the BLACS. In IEEE [IEE96a], pages 463–468. ISBN 0-8186-7557-8. LCCN QA76.88.I575 1996. IEEE catalog number 96TB100074.

Djordjevic:1996:ICI

[DS96b] G. L. Djordjevic and M. K. Stojcev. An interprocessor communication interface for message passing via shared memory modules-design and performances. Computers and Artificial Intelligence = Vychislitel’nye mashiny i iskusstvennyi intellekt, 15 (1):1–34, ???? 1996. CO-DEN CARIDY. ISSN 0232-0274.

Dang:2013:CES

[DS13] Hoang-Vu Dang and Bertil Schmidt. CUDA-enabled sparse matrix-vector multiplication on GPUs using atomic operations. Parallel Computing, 39(11): 737–750, November 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819113001178.

Deniz:2016:MGM

[DS16] Etem Deniz and Alper Sen.

MINIME-GPU: Multicore benchmark synthesizer for GPUs. ACM Transactions on Architecture and Code Optimization, 12(4):34:1– 34:??, January 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Delmas:2022:MGI

[DS22] Vincent Delmas and Azzedine Soulaïmani. Multi-GPU implementation of a time-explicit finite volume solver using CUDA and a CUDA-aware version of OpenMPI with application to shallow water flows. Computer Physics Communications, 271(??):Article 108190, February 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465521003027.

Duran:2005:RAP

[DSCL05] A. Duran, R. Silvera, J. Corbalan, and J. Labarta. Runtime adjustment of parallel nested loops. Lecture Notes in Computer Science, 3349: 137–??, 2005.

Dang:2017:ECB

[DSG17] Hoang-Vu Dang, Marc Snir, and William Gropp. Eliminating contention bottlenecks in multithreaded MPI. Parallel Computing, 69(??): 1–23, November 2017. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819117301187.

Dietrich:2017:CBA

[DSGS17] Robert Dietrich, Felix Schmitt, Alexander Grund, and Jonas Stolle. Critical-blame analysis for OpenMP 4.0 offloading on Intel Xeon Phi. The Journal of Systems and Software, 125(??):381– 388, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL /

                                        /www.sciencedirect.com/
                                        science/article/pii/S0164121215002940.

Davidor:1994:PPS

[DSM94] Yuval Davidor, Hans-Paul Schwefel, and Reinhard Manner, editors. Parallel problem solving from nature — PPSN III: International Conference on Evolutionary Computation, the Third Conference on Parallel Problem Solving from Nature, Jerusalem, Israel, October 9–14, 1994: proceedings, number 866 in Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58484-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I535 1994.

Dohi:2011:GIO

[DSOF11] Keisuke Dohi, Yuichiro Shibata, Kiyoshi Oguri, and Takafumi Fujimoto. GPU implementation and optimization of electromagnetic simulation using the FDTD method for antenna designing. ACM SIGARCH Computer Architecture News, 39 (4):26–31, September 2011. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Domokos:2000:PRC

[DSS00] Gábor Domokos, Imre Szeberényi, and Paul H. Steen. Parallel, recursive computation of global stability charts for liquid bridges. Lecture Notes in Computer Science, 1908:64–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080064.htm;
 http://link.springer-
ny.com/link/service/series/

                                        0558/papers/1908/19080064.
                                        pdf.

Daleiden:2020:GPP

[DSU20] Patrick Daleiden, Andreas Stefik, and Philip Merlin Uesbeck. GPU programming productivity in different abstraction paradigms: a randomized controlled trial comparing CUDA and thrust. ACM Transactions on Computing Education, 20(4):27:1–27:27, November 2020. CODEN ???? ISSN 1946-6226. URL https://

                                        dl.acm.org/doi/10.1145/
                                        3418301.

Deshpande:1996:MIBa

[DSW96] V. Deshpande, W. Sawyer, and D. W. Walker. An MPI implementation of the BLACS. In IEEE [IEE96i], pages 195–198. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Dekker:1994:MPP

[DSZ94] L. (Leendert) Dekker, W. Smit, and J. C. Zuidervaart, editors. Massively parallel processing applications and development: proceedings of the 1994 EUROSIM Conference on Massively Parallel Processing Applications and Development, Delft, The Netherlands, 21– 23 June 1994. Elsevier, Amsterdam, The Netherlands, 1994. ISBN 0-444-81784-0. LCCN QA76.58.E98 1994.

Dongarra:1994:PSW

[DT94] Jack J. Dongarra and Bernard Tourancheau, editors. Proceedings of the Second Workshop on Environments and Tools for Parallel Scientific Computing: Townsend, TN, USA, 25–27 May 1994. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-343-9. LCCN QA76.58.I568 1994.

Diavastos:2017:SLR

[DT17] Andreas Diavastos and Pedro Trancoso. SWITCHES: a lightweight runtime for dataflow execution of tasks on many-cores. ACM Transactions on Architecture and Code Optimization, 14(3): 31:1–31:??, September 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Duval:1992:TPP

[Duv92] D. Duval. Trends in parallel programming models for high performance computers. In Ferenczi [Fer92], page 33. ISBN ???? LCCN ????

Dikken:1994:DDL

[DvdLVS94] L. Dikken, F. van der Linden, J. Vesseur, and P. Sloot. DynamicPVM: Dynamic load balancing on parallel systems. In Gentzsch and Harms [GH94], pages 273–277. ISBN 0-387-57981-8 (New York), 3-540-57981-8 (Berlin). LCCN QA76.88.I57 1994. DM96.00. Two volumes.

Dongarra:1994:PSC

[DW94] Jack Dongarra and Jerzy Wasniewski, editors. Parallel scientific computing: First International Workshop, PARA ’94, Lyngby, Denmark, June 20–23, 1994: proceedings, volume 879 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1994. DM104.00.

DeRose:2002:CCG

[DW02] L. DeRose and F. Wolf. CATCH — a call-graph based automatic tool for capture of hardware performance metrics for MPI and OpenMP applications. Lecture Notes in Computer Science, 2400:167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2400/24000167.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2400/24000167.
pdf.

Du:2010:COT

[DWL^{+}10] Peng Du, Rick Weber, Piotr Luszczek, Stanimire Tomov, Gregory Peterson, and Jack Dongarra. From CUDA to OpenCL: Towards a performance-portable solution for multi-platform GPU programming. LA-PACK Working Note 228, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, September 6, 2010. URL http:/

/www.netlib.org/lapack/
lawnspdf/lawn228.pdf. UT-

CS-10-656.

Du:2012:COT

[DWL^{+}12] Peng Du, Rick Weber, Piotr Luszczek, Stanimire Tomov, Gregory Peterson, and Jack Dongarra. From CUDA to OpenCL: Towards a performance-portable solution for multi-platform GPU programming. Parallel Computing, 38(8):391– 407, August 2012. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819111001335.

Deshpande:2012:AGC

[DWM12] Vivek Deshpande, Xing Wu, and Frank Mueller. Autogeneration of communication benchmark traces. ACM SIGMETRICS Performance Evaluation Review, 40(2):

99–105, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dosanjh:2021:IEM

[DWS^{+}21] Matthew G. F. Dosanjh, Andrew Worley, Derek Schafer, Prema Soundararajan, Sheikh Ghafoor, Anthony Skjellum, Purushotham V. Bangalore, and Ryan E. Grant. Implementation and evaluation of MPI 4.0 partitioned communication libraries. Parallel Computing, 108(??):??, December 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819121000752.

Dong:1996:SPL

[DXB96] Li Dong, Li Xiaoming, and Fang Binxing. The study on the parallel library based on MPI. Mini-Micro Systems, 17(12):17–19, 1996. CODEN XWJXEH. ISSN 1000-1220.

Deng:2006:PIK

[DYN^{+}06] Junjun Deng, Hengyong Yu, Jun Ni, Tao He, Shiying Zhao, Lihe Wang, and Ge Wang. A parallel implementation of the katsevich algorithm for 3-D CT image reconstruction. The Journal of Supercomputing, 38(1):35–47, October 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
38&issue=1&spage=35.

Dantas:1996:ILB

[DZ96] M. A. R. Dantas and E. J.

Zaluska. Improving load balancing in an MPI environment with resource management. In Liddell et al. [LCHS96], pages 959–960. ISBN 3-540-61142-8 (paperback). LCCN QA76.88 .H52 1996.

Dantas:1998:ESM

[DZ98a] M. A. R. Dantas and E. J.

Zaluska. Efficient scheduling of MPI applications on networks of workstations. Future Generation Computer Systems, 13(6):489– 499, May 20, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/20/21/abstract.
html.

Delves:1998:HPF

[DZ98b] M. Delves and H. Zima.

High Performance Fortran: a status report or: Are we ready to give up MPI? Lecture Notes in Computer Science, 1497:161–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dragovitsch:1995:PPS

[DZDR95] P. Dragovitsch, X. Zhao, L. C. Dennis, and G. A. Riccardi. PVMGeant — a parallel simulation code for the CLAS detector at CE-BAF. International Journal of Supercomputer Applications and High Performance Computing, 9(2):128– 137, Summer 1995. CODEN IJSCFG. ISSN 1078-3482.

Dykes:1994:CCP

[DZZY94] S. G. Dykes, Xiaodong Zhang, Yan Zhou, and Haixu Yang. Communication and computation patterns of large scale image convolutions on parallel architectures. In Siegal [Sie94], pages 926–931. ISBN 0-8186-5602-6. LCCN QA76.58.I58 1994. IEEE catalog no. 94CH34819.

Edmonds:2019:HAS

[EADT19] Mark Edmonds, Tanvir Atahary, Scott Douglass, and Tarek Taha. Hardware accelerated semantic declarative memory systems through CUDA and MapReduce. IEEE Transactions on Parallel and Distributed Systems, 30(3):601– 614, March 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/03/08444694-
abs.html.

Edjlali:1995:DPP

[EASS95] G. Edjlali, G. Agrawal, A. Sussman, and J. Saltz. Data parallel programming in an adaptive environment. In IEEE [IEE95f], pages 827–832. ISBN 0-8186-7074-6. LCCN QA 76.58 I56 1995. IEEE catalog no. 95TH8052.

Eichenberger:2020:HCG

[EBB^{+}20] A. E. Eichenberger, G.-T.

Bercea, A. Bataev, L. Grinberg, and J. K. O’Brien. Hybrid CPU/GPU tasks optimized for concurrency in OpenMP. IBM Journal of Research and Development, 64(3/4):13:1–13:14, May/July 2020. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Elwasif:2001:AMT

[EBKG01] Wael R. Elwasif, David E.

Bernholdt, James A. Kohl, and G. A. Geist. An architecture for a multi-threaded harness kernel. Lecture Notes in Computer Science, 2131:126–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310126.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310126.
pdf.

Eppstein:1994:CSP

[ED94] M. J. Eppstein and D. E. Dougherty. A comparative study of PVM workstation cluster implementations of a two-phase subsurface flow model. Advances in water resources, 17(3):181–??, ???? 1994. CODEN AWREDI. ISSN 0309-1708 (print), 1872-9657 (electronic).

Eddelbuettel:2018:BRN

[Edd18] Dirk Eddelbuettel. Book review: Norman Matloff. Parallel Computing for Data Science: With Examples in R, C++, and CUDA. Boca Raton: CRC Press. Biometrics, 74(2):770, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eigenmann:2008:ONE

[EdS08] Rudolf Eigenmann and Bronis R. de Supinski, editors. OpenMP in a New Era of Parallelism: 4th International Workshop, IWOMP 2008 West Lafayette, IN, USA, May 12–14, 2008 Proceedings, volume 5004 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CO-DEN LNCSD9. ISBN 3-540-79560-X (print), 3-540-79561-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:/

/www.springerlink.com/
content/978-3-540-79561-
2.

ElMaghraoui:2009:MIM

[EDSV09] K. El Maghraoui, Travis J.

Desell, Boleslaw K. Szymanski, and Carlos A. Varela. Malleable iterative MPI applications. Concurrency and Computation: Practice and Experience, 21(3): 393–413, March 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Eleftheriou:2005:SFF

[EFR^{+}05] M. Eleftheriou, B. G. Fitch, A. Rayshubskiy, T. J. C. Ward, and R. S. Germain. Scalable framework for 3D FFTs on the Blue Gene/L supercomputer: Implementation and early performance measurements. IBM Journal of Research and Development, 49(2/3):457– 464, ???? 2005. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/eleftheriou.
pdf.

El-Ghazawi:2002:UPP

[EGC02] Tarek El-Ghazawi and François Cantonnet. UPC performance and potential: a NPB experimental study. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN

???? URL http://www.sc-
2002.org/paperpdfs/pap.
pap316.pdf.

Eppstein:1992:PGC

[EGDK92] Margaret J. Eppstein, Joseph F. Guarnaccia, David Emery Dougherty, and Robert S. Kerr. Parallel groundwater computations using PVM. In Russell et al. [R^{+}92], pages 713–720. ISBN 1-85166-871-3 (set), 1-85312-169-X (set: Computational Mechanics Publications, Southampton), 1-56252-098-9 (set: Computational Mechanics Publications, Boston), 1-85166-791-1 (v. 1: Elsevier Applied Science), 1-85312-197-5 (v. 1: Computational Mechanics Publications, Southampton), 1-56252-123-3 (v. 1: Computational Mechanics Publications, New York), 1-85166-870-5 (v. 2), 1-85312-198-3 (v. 2), 1-56252-124-1 (v. 2). LCCN GB656.2.E42 C65 1992 v.1-2 (c1992). Two volumes.

Eickermann:1999:PID

[EGH99] T. Eickermann, H. Grund, and J. Henrichs. Performance issues of distributed MPI applications in a German gigabit testbed. In Dongarra et al. [DLM99], pages 3–10. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Erhel:2014:DDM

[EGH^{+}14] Jocelyne Erhel, Martin J.

Gander, Laurence Halpern, Géraldine Pichot, Taoufik Sassi, and Olof Widlund, editors. Domain Decomposition Methods in Science and Engineering XXI, volume 98 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-05788-X (paperback), 3-319-05789-8 (e-book). ISSN 1439-7358 (print), 2197-7100 (electronic). LCCN QA71-

90. URL http://0-dx.doi.
org.fama.us.es/10.1007/
978-3-319-05789-7.

Ebrahimirad:2015:EAS

[EGR15] Vahid Ebrahimirad, Maziar Goudarzi, and Aboozar Rajabi. Energy-aware scheduling for precedenceconstrained parallel virtual machines in virtualized data centers. Journal of Grid Computing, 13(2): 233–253, June 2015. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10723-015-9327-x.

Evans:1992:PCP

[EJL92] D. J. Evans, G. R. Joubert, and H. Liddell, editors. Parallel computing ’91: proceedings of the International Conference on Parallel Computing ’91, London, UK, 3–6 September 1991, volume 4 of Advances in parallel computing. North-Holland, Amsterdam, The Netherlands, 1992. ISBN 0-444-89212-5. LCCN QA76.58.I545 1991.

Exbrayat:1997:OPS

[EK97] M. Exbrayat and H. Kosch. Offering parallelism to a sequential database management system on a network of workstations using PVM. Lecture Notes in Computer Science, 1332:457–435, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eberl:1999:PCP

[EKTB99] M. Eberl, W. Karl, C. Trinitis, and A. Blaszczyk. Parallel computing on PC clusters — an alternative to supercomputers for industrial applications. In Dongarra et al. [DLM99], pages 493– 498. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Elamvazuthi:1994:OPA

[EM94] C. Elamvazuthi and G. A. Manson. Occam, PVM and the alternative construct. In Miles and Chalmers [MC94], pages 56–68. ISBN 90-5199-163-0. LCCN ????

Eigenmann:2000:TMPa

[EM00a] Rudolf Eigenmann and Tim Mattson. Tutorial M6A: Parallel programming with OpenMP: Part I. In ACM [ACM00], page 21. URL

http://www.sc2000.org/
proceedings/info/fp.pdf.

Eigenmann:2000:TMPb

[EM00b] Rudolf Eigenmann and Tim Mattson. Tutorial M6B: Parallel programming with OpenMP: Part II. In ACM [ACM00], page 23. URL

http://www.sc2000.org/
proceedings/info/fp.pdf.

Espenica:2002:PPA

[EM02] Roberto Espenica and Pedro Medeiros. Porting PVM to the VIA architecture using a fast communication library. Lecture Notes in Computer Science, 2474:341–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/
service/series/0558/bibs/
2474/24740341.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740341.pdf.

Espinosa:1998:ADP

[EML98] A. Espinosa, T. Margalef, and E. Luque. Automatic detection of PVM program performance problems.

Lecture Notes in Computer Science, 1497:19–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Espinosa:2000:APA

[EML00] Antonio Espinosa, Tomas Margalef, and Emilio Luque. Automatic performance analysis of master/worker PVM applications with Kpi. Lecture Notes in Computer Science, 1908:47–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080047.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080047.
pdf.

Ewing:1993:DCW

[EMO^{+}93] R. E. Ewing, D. Mitchum, P. O’Leary, R. C. Sharpley, and J. S. Sochacki. Distributed computation of wave propagation models using PVM. In IEEE [IEE93e], pages 22–31. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Engquist:2000:SVG

[Eng00] Björn Engquist, editor. Simulation and visualization on the grid: Parallelldatorcentrum, Kungl. Tekniska Högskolan, seventh annual conference, Stockholm, Sweden, December 1999: proceedings, volume 13 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67264-8. ISSN 1439-7358. LCCN QA76.9.C65 S535 2000.

Emani:2015:CDM

[EO15] Murali Krishna Emani and Michael O’Boyle. Celebrating diversity: a mixture of experts approach for runtime mapping in dynamic environments. ACM SIG-PLAN Notices, 50(6):499– 508, June 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ebner:1996:TFP

[EP96] R. Ebner and A. Pfaffinger. Transformation of functional programs into data flow graphs implemented with PVM. In Bode et al. [BDLS96], pages 251–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Espinosa:1999:REB

[EPML99] A. Espinosa, F. Parcerisa, T. Margalef, and E. Luque.

Relating the execution behaviour with the structure of the application. In Dongarra et al. [DLM99], pages 91–100. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Eizenberg:2017:BBL

[EPP^{+}17] Ariel Eizenberg, Yuanfeng Peng, Toma Pigli, William Mansky, and Joseph Devietti. BARRACUDA: binarylevel analysis of runtime RAces in CUDA programs. ACM SIGPLAN Notices, 52(6):126–140, June 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ElZein:2012:GOC

[ER12] Ahmed H. El Zein and Alistair P. Rendell. Generating optimal CUDA sparse matrix–vector product implementations for evolving GPU hardware. Concurrency and Computation: Practice and Experience, 24(1):3–13, January 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

El-Rewini:1995:PTE

[ERS95] H. El-Rewini and B. D. Shriver, editors. Proceedings of the Twenty-Eighth Hawaii International Conference on System Sciences.

IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-6935-7. LCCN ????

El-Rewini:1996:PTN

[ERS96] Hesham El-Rewini and Bruce D. Shriver, editors. Proceedings of the Twenty-Ninth Hawaii International Conference on System Sciences (HICSS-29): Wailea, HI, USA, 3–6 January 1996. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7324-9. ISSN 1060-3425. LCCN ???? Five volumes.

Ewedafe:2011:PID

[ES11] Simon Uzezi Ewedafe and Rio Hirowati Shariffudin. Parallel implementation of 2-D telegraphic equation on MPI/PVM cluster. International Journal of Parallel Programming, 39(2): 202–231, April 2011. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=2&spage=202.

Ellingson:2013:SNU

[ESB13] Sally R. Ellingson, Jeremy C.

Smith, and Jerome Baudry. Software news and updates: VinaMPI: Facilitating multiple receptor highthroughput virtual docking on high-performance computers. Journal of Computational Chemistry, 34 (25):2212–2221, September 30, 2013. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ewing:1994:DCW

[ESM^{+}94] Richard E. Ewing, Robert C. Sharpley, Derek Mitchum, P. O’Leary, and J. S. Sochacki. Distributed computation of wave propagation models using PVM. IEEE parallel and distributed technology: systems and applications, 2(1):26– 31, Spring 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Escaig:1994:PMD

[ETV94] Y. Escaig, G. Touzot, and M. Vayssade. Parallelization of a multilevel domain decomposition method. Computing systems in engineering: an international journal, 5(3):253–263, June 1994. CODEN COSEEO. ISSN 0956-0521.

Eichenberger:2012:DOT

[ETWaM12] Alexandre E. Eichenberger, Christian Terboven, Michael Wong, and Dieter an Mey. The design of OpenMP thread affinity. Lecture Notes in Computer Science, 7312:15–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-30961-8_
2/.

Eigenmann:2001:OSM

[EV01] Rudolf Eigenmann and Michael J. Voss, editors. OpenMP shared memory parallel programming: International Workshop on OpenMP Applications and Tools, WOMPAT 2001, West Lafayette, IN, USA, July 30–31, 2001: Proceedings, volume 2104 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CO-DEN LNCSD9. ISBN 3-540-42346-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.642 .I589 2001; QA267.A1 L43 no.2104. URL http:

//link.springer-ny.com/
link/service/series/0558/
tocs/t2104.htm.

Eichstadt:2020:CSM

[EVMP20] Jan Eichstädt, Martin Vymazal, David Moxey, and Joaquim Peiró. A comparison of the shared-memory parallel programming models OpenMP, OpenACC and Kokkos in the context of implicit solvers for highorder FEM. Computer Physics Communications, 255(??):Article 107245, October 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0010465520300746.

Elis:2020:QNG

[EYP^{+}20] Bengisu Elis, Dai Yang, Olga Pearce, Kathryn Mohror, and Martin Schulz. QMPI: a next generation MPI profiling interface for modern HPC platforms. Parallel Computing, 96(??): Article 102635, August 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819120300284.

Eckert:2016:HAL

[EZBA16] C. H. J. Eckert, E. Zenker, M. Bussmann, and D. Albach. HASEonGPU — an adaptive, load-balanced MPI/GPU-code for calculating the amplified spontaneous emission in high power laser media. Computer Physics Communications, 207(??):362–374, October 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0010465516301436.

Faraji:2018:DCG

[FA18] Iman Faraji and Ahmad Afsahi. Design considerations for GPU-aware collective communications in MPI. Concurrency and Computation: Practice and Experience, 30(17):e4667:1– e4667:??, September 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fabeiro:2016:WPP

[FAF16] Jorge F. Fabeiro, Diego Andrade, and Basilio B. Fraguela. Writing a performanceportable matrix multiplication. Parallel Computing, 52 (??):65–77, February 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819115001611.

Fabeiro:2015:AGO

[FAFD15] Jorge F. Fabeiro, Diego Andrade, Basilio B. Fraguela, and Ramón Doallo. Automatic generation of optimized OpenCL codes using OCLoptimizer. The Computer Journal, 58(11): 3057–3073, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Fang:1998:DDL

[Fan98] Niandong Fang. Distributed data library and tools for an MPI programming environment, volume 1 of Research reports in computer science. Shaker, Aachen, Germany, 1998. ISBN 3-8265-4101-4. xx + 195 pp. LCCN ???? Also published as dissertation of the University of Basel.

Freeman:1994:SMM

[FB94] T. L. Freeman and J. M. Bull. Shared memory and message passing implementations of parallel algorithms for numerical integration. Lecture Notes in Computer Science, 879:219–228, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fang:1995:PMS

[FB95] Niandong Fang and H. Burkhart. PEMPI — from MPI standard to programming environment. In IEEE [IEE95j], pages 31–38. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Fang:1996:SPP

[FB96] N. Fang and H. Burkhart. Structured parallel programming using MPI. In Liddell et al. [LCHS96], pages 840–847. ISBN 3-540-61142-8 (paperback). LCCN QA76.88 .H52 1996.

Fang:1997:MDD

[FB97] Niandong Fang and Helmar Burkhart. MPI-DDL: a distributed-data library for MPI. Future Generation Computer Systems, 12(5):407–419, April 1, 1997. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/17/23/abstract.
html.

Fagg:2001:FTM

[FBD01a] Graham E. Fagg, Antonin Bukovsky, and Jack J. Dongarra. Fault tolerant MPI for the HARNESS metacomputing system. Lecture Notes in Computer Science, 2073:355–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2073/20730355.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2073/20730355.
pdf.

Fagg:2001:HFT

[FBD01b] Graham E. Fagg, Antonin Bukovsky, and Jack J. Dongarra. HARNESS and fault tolerant MPI. Parallel Computing, 27(11):1479–1495, October 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
35/21/47/41/32/abstract.
html;   http://www.elsevier.

                                        nl/gej-ng/10/35/21/47/
                                        41/32/article.pdf;   http:
                                        //www.netlib.org/utk/people/
                                        JackDongarra/PAPERS/harness-
                                        ftmpi-pc.pdf.

Friedel:2001:HMC

[FBSN01] Peter Friedel, Jörg Bergmann, Stephan Seidl, and Wolfgang E. Nagel. An hierarchical MPI communication model for the parallelized solution of multiple integrals. Lecture Notes in Computer Science, 2110:474–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2110/21100474.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2110/21100474.
                                        pdf.

Fagg:2002:FTM

[FBVD02] Graham E. Fagg, Antonin Bukovsky, Sathish Vadhiyar, and Jack J. Dongarra. Fault tolerant MPI for the HARNESS MetaComputing system. Technical report ????, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 2002. 14 pp.

                                        URL http://www.netlib.
                                        org/netlib/utk/people/
                                        JackDongarra/PAPERS/ft-
                                        mpi-iccs-gef.pdf.

Floros:2005:TGS

[FC05] Evangelos Floros and Yiannis Cotronis. Towards a Grid services based framework for the virtualization, execution and composition of MPI applications. Parallel Processing Letters, 15(1/2):85–98, March/June 2005. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Falzone:2007:PMF

[FCLG07] Christopher Falzone, Anthony Chan, Ewing Lusk, and William Gropp. A portable method for finding user errors in the usage of MPI collective operations. The International Journal of High Performance Computing Applications, 21(2):155– 165, May 2007. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/21/
2/155.full.pdf+html.

Ferschweiler:2001:CDP

[FCP^{+}01] Ken Ferschweiler, Mariacarla Calzarossa, Cherri Pancake, Daniele Tessera, and Dylan Keon. A community databank for performance tracefiles. Lecture Notes in Computer Science, 2131:233–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310233.htm;

                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2131/21310233.
                                        pdf.

Filgueira:2012:DCD

[FCS^{+}12] Rosa Filgueira, Jesús Carretero, David E. Singh, Alejandro Calderón, and Alberto Núñez. Dynamic– CoMPI: dynamic optimization techniques for MPI parallel applications. The Journal of Supercomputing, 59 (1):361–391, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0920-8542&volume=
                                        59&issue=1&spage=361.

Fujita:2019:EIM

[FCS^{+}19] Hajime Fujita, Chongxiao Cao, Sayantan Sur, Charles Archer, Erik Paulson, and Maria Garzaran. Efficient implementation of MPI-3 RMA over openFabrics interfaces. Parallel Computing, 87(??):1–10, September 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167819118303843.

Fagg:1996:PIP

[FD96] Graham Fagg and Jack Dongarra. PVMPI: An integration of PVM and MPI systems. Calculateurs Parallèles, 8(2):151– 166, 1996. CODEN ???? ISSN 1260-3198. URL

http://www.netlib.org/
utk/papers/pvmpi/paper.
html;   http://www.netlib.
org/utk/papers/pvmpi/pvmpi.
ps;   http://www.netlib.
org/utk/people/JackDongarra/
pdf/pvmpi.pdf.

Fischer:1997:AAP

[FD97] Markus Fischer and Jack Dongarra. Another architecture: PVM on Windows 95/ NT. In ????, editor, Concurrent Computing Conference, Atlanta, GA, March 10–11, 1994, page ?? ????, ????,

1997. URL http://www.
netlib.org/utk/people/
JackDongarra/PAPERS/nt-
paper.ps;   http://www.
netlib.org/utk/people/
JackDongarra/pdf/nt-paper.
pdf.

Fagg:2000:FMF

[FD00] Graham E. Fagg and Jack J.

Dongarra. FT-MPI: Fault Tolerant MPI, supporting dynamic applications in a dynamic world. Lecture Notes in Computer Science, 1908:346–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080346.htm;
 http://link.springer-

                                        ny.com/link/service/series/
                                        0558/papers/1908/19080346.
                                        pdf.

Fagg:2002:HFTa

[FD02a] Graham E. Fagg and Jack J. Dongarra. HARNESS fault tolerant MPI design, usage and performance issues. Technical report ????, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 2002.

                                        URL http://www.netlib.
                                        org/netlib/utk/people/
                                        JackDongarra/PAPERS/ft-
                                        mpi-fgcs-grid-se.pdf.

Fagg:2002:HFTb

[FD02b] Graham E. Fagg and Jack J. Dongarra. HARNESS fault tolerant MPI design, usage and performance issues. Future Generation Computer Systems, 18(8): 1127–1142, October 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fagg:2004:BUF

[FD04] Graham E. Fagg and Jack J. Dongarra. Building and using a fault-tolerant MPI implementation. The International Journal of High Performance Computing Applications, 18(3):353–361, Fall 2004. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

                                        tronic). URL http://hpc.

sagepub.com/content/18/
3/353.full.pdf+html.

Fagg:1997:HMAa

[FDG97a] G. Fagg, J. Dongarra, and A. Geist. Heterogeneous MPI application interoperation and process management under PVMPI. Technical report CS-97- ???, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, June

1997. URL http://www.
netlib.org/utk/papers/
pvmmpi97.ps;   http://
www.netlib.org/utk/people/
JackDongarra/pdf/pvmmpi97.
pdf.

Fagg:1997:HMAb

[FDG97b] G. E. Fagg, J. J. Dongarra, and A. Geist. Heterogeneous MPI application interoperation and process management under PVMPI. Lecture Notes in Computer Science, 1332:91–98, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faict:2019:MGI

[FDG19] Thomas Faict, Erik H.

D’Hollander, and Bart Goossens. Mapping a guided image filter on the HARP reconfigurable architecture using OpenCL. Algorithms (Basel), 12(8), August 2019. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://

                                        www.mdpi.com/1999-4893/
                                        12/8/149.

Falch:2017:MLB

[FE17a] Thomas L. Falch and Anne C. Elster. Machine learning-based auto-tuning for enhanced performance portability of OpenCL applications. Concurrency and Computation: Practice and Experience, 29(8): ??, April 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Falch:2017:RAM

[FE17b] Thomas L. Falch and Anne C. Elster. Machine learning-based auto-tuning for enhanced performance portability of OpenCL applications. Concurrency and Computation: Practice and Experience, 29(8): ??, April 25, 2017. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ferenczi:1992:AHW

[Fer92] S. Ferenczi, editor. 1st Austrian-Hungarian Workshop on Transporter Applications. Proceedings. Hungarian Acad.of Sci, Budapest, Hungary, 1992. ISBN ???? LCCN ????

Ferrari:1998:JNPb

[Fer98a]  Adam Ferrari.  JPVM:

network parallel computing in Java. Concurrency: practice and experience, 10(11–13):985–992, September 1998. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract?ID=10050413;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050413&PLACEBO=IE.

pdf. Special Issue: Java for High-performance Network Computing.

Ferrari:1998:JNPa

[Fer98b] Adam J. Ferrari. JPVM:

Network parallel computing in Java. In ACM [ACM98a], page ?? ISBN ???? LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/
papers/jpvm.pdf;   http:
//www.cs.ucsb.edu/conferences/
java98/papers/jpvm.ps.

Possibly unpublished, except electronically.

Fernando:2004:GGP

[Fer04] Randima Fernando, editor.

GPU gems: programming techniques, tips, and tricks for real-time graphics, volume 1 of GPU gems. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-22832-4. xvv + 765 pp. LCCN T385 .G6879 2004. US$45.99.

FerreiradaSilva:2010:PBC

[Fer10] Adelino Ferreira da Silva.

cudaBayesreg:  Bayesian

computation in CUDA. The R Journal, 2(2):48– 55, December 2010. CO-DEN ???? ISSN 2073-4859. URL http://

                                        journal.r-project.org/
                                        archive/2010-2/RJournal_
                                        2010-2_Ferreira~da-Silva.
                                        pdf.

Fritzson:1995:PPA

[FF95] Peter Fritzson and Leif Finmo, editors. Parallel programming and applications: proceedings of the Workshop on Parallel Programming and Computation (ZEUS ’95) and the 4th Nordic Transputer Conference (NTUG ’95): Linköping, Sweden. IOS Press, Postal Drawer 10558, Burke, VA 2209-0558, USA, 1995. ISBN 90-5199-229-7 (IOS Press), 4-274-90056-8 (Ohmsha). LCCN ????

Fava:1999:MPI

[FFB99] A. Fava, M. Fava, and M. Bertozzi. MPIPOV: a parallel implementation of POV-Ray based on MPI. In Dongarra et al. [DLM99], pages 426–433. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Frugoli:1999:DCH

[FFFC99] G. Frugoli, A. Fava, E. Fava, and G. Conte. Distributed collision handling for particle-based simulation. In Dongarra et al. [DLM99], pages 410–417. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Fousek:2011:AFC

[FFM11] Jan Fousek, Jiři Filipovič, and Matuš Madzin. Automatic fusions of CUDA– GPU kernels for parallel map. ACM SIGARCH Computer Architecture News, 39 (4):98–99, September 2011. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Fernandez:2003:BMN

[FFP03] Juan Fernandez, Eitan Frachtenberg, and Fabrizio Petrini. BCS-MPI: a new approach in the system software design for large-scale parallel computers. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN ???? URL http:/

/www.sc-conference.org/
sc2003/inter_cal/inter_
cal_detail.php?eventid=
10716#1;   http://www.
sc-conference.org/sc2003/
paperpdfs/pap306.pdf.

Foster:1998:WAI

[FGG^{+}98] Ian Foster, Jonathan Geisler, William Gropp, Nicholas Karonis, Ewing Lusk, George Thiruvathukal, and Steven Tuecke. Wide-area implementation of the Message Passing Interface. Parallel Computing, 24(12– 13):1735–1749, November 1, 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/parco/sub/1998/24/
12-13/1352.pdf.

Foster:1997:MMC

[FGKT97] Ian Foster, Jonathan Geisler, Carl Kesselman, and Steven Tuecke. Managing multiple communication methods in high-performance networked computing systems. Journal of Parallel and Distributed Computing, 40(1): 35–48, January 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.idealibrary.com/
links/doi/10.1006/jpdc.
1996.1266/production;
http://www.idealibrary.
com/links/doi/10.1006/
jpdc.1996.1266/production/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/
jpdc.1996.1266/production/
ref.

Ferreira:2020:HMM

[FGL^{+}20] Kurt Ferreira, Ryan E. Grant, Michael J. Levenhagen, Scott Levy, and Taylor Groves. Hardware MPI message matching: Insights into MPI matching behavior to inform design. Concurrency and Computation: Practice and Experience, 32 (3):e5150:1–e5150:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fagg:2001:PIS

[FGRD01] Graham E. Fagg, Edgar Gabriel, Michael Resch, and Jack J. Dongarra. Parallel IO support for meta-computing applications: MPI Connect IO applied to PACX–MPI. Lecture Notes in Computer Science, 2131:135–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310135.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310135.
pdf.

Fahringer:2000:FOP

[FGRT00] Thomas Fahringer, Michael Gerndt, Graham Riley, and Jesper Larsson Träff. Formalizing OpenMP performance properties with ASL. Lecture Notes in Computer Science, 1940:428–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/

                                        bibs/1940/19400428.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1940/19400428.
                                        pdf.

Foster:1996:MIW

[FGT96] I. Foster, J. Geisler, and S. Tuecke. MPI on the I-WAY: a wide-area, multimethod implementation of the Message Passing Interface. In IEEE [IEE96i], pages 10–17. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Fan:1995:DMP

[FH95] W. C. Fan and J. A. Halbleib, Sr. Distributed multitasking ITS with PVM. Transactions of the American Nuclear Society, 72 (????):146–147, ???? 1995. CODEN TANSAO. ISSN 0003-018X.

Fachat:1997:IEB

[FH97] André Fachat and Karl Heinz Hoffmann. Implementation of Ensemble-Based Simulated Annealing with dynamic load balancing under MPI. Computer Physics Communications, 107(1–3): 49–53, December 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465597000969.

Andre:1998:BVN

[FH98] André Fachat and Karl Heinz Hoffmann. Blocking vs. non-blocking communication under MPI on a masterworkerproblem. Preprint-Reihe des Chemnitzer SFB 393 Sonderforschungsbereich NumerischeSimulation auf Massiv Parallelen Rechnern 98,18, Universität Chemnitz-Zwickau, Chemnitz, Germany, 1998.

Friedley:2013:OPE

[FHB^{+}13] Andrew Friedley, Torsten Hoefler, Greg Bronevetsky, Andrew Lumsdaine, and Ching-Chen Ma. Ownership passing: efficient distributed memory programming on multi-core systems. ACM SIGPLAN Notices, 48 (8):177–186, August 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’13 Conference proceedings.

Franke:1995:AAV

[FHC^{+}95] E. A. Franke, S. D. Huffman, W. M. Carter, J. P. Baumgartner, and D. J. Wenzel. AVTP — an architecture for visualization using remote parallel/distributed computing. In Grinstein and Erbacher [GE95], pages 230–237. CODEN PSISDG. ISBN 0-8194-1757-2. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.2410.

Field:2001:RTF

[FHK01] Antony J. Field, Thomas L. Hansen, and Paul H. J. Kelly. Run-time fusion of MPI calls in a parallel C++ library. Lecture Notes in Computer Science, 2017:363–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2017/20170363.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2017/20170363.
                                        pdf.

Franke:1994:MMP

[FHP^{+}94] H. Franke, P. Hochschild, P. Pattnaik, J.-P. Prost, and M. Snir. MPI-F: an MPI prototype implementation on IBM SP1. In Dongarra and Tourancheau [DT94], pages 43–55. ISBN 0-89871-343-9. LCCN QA76.58.I568 1994.

Franke:1995:MIS

[FHP^{+}95] H. Franke, P. Hochschild, P. Pattnaik, J.-P. Prost, and M. Snir. MPI on IBM SP1/SP2: current status and future directions. In IEEE [IEE95j], pages 39–48. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Franke:1994:EIM

[FHPS94a] H. Franke, P. Hochschild, P. Pattnaik, and M. Snir. An efficient implementation of MPI. In Decker and Rehmann [DR94], pages 219–230. ISBN 0-8176-5090-3 (Boston), 3-7643-5090-3 (Basel). LCCN QA76.58.P767 1994.

Franke:1994:MEI

[FHPS94b] H. Franke, P. Hochschild, P. Pattnaik, and M. Snir. MPI-F: An efficient implementation of MPI on IBM-SP1. In Agrawal et al. [ATC94], pages III–197–III–

201. ISBN 0-8493-2496-3, 0-8493-2495-5. ISSN 0190-3918. LCCN QA 76.58 I55 1994. Three volumes.

Fang:1999:PMD

[FHSO99] Zhiwu Fang, A. D. J.

Haymet, Wataru Shinoda, and Susumu Okazaki. Parallel molecular dynamics simulation: Implementation of PVM for a lipid membrane. Computer Physics Communications, 116(2–3): 295–310, February 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465598000897.

Fineberg:1994:IMM

[Fin94] S. A. Fineberg. Implementing multidisciplinary and multi-zonal applications using MPI. In IEEE [IEE94a], pages 496–503. ISBN 0-8186-6965-9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

Fineberg:1995:IMM

[Fin95] Samuel A. Fineberg. Implementing multidisciplinary and multi-zonal applications using MPI. Frontiers of Massively Parallel Computation — Conference Proceedings, pages 496–503, ???? 1995. IEEE catalog number 95TH8024.

Fin:1997:CPM

[Fin97] Torsten Fin. Comparing the performance of MPI, PVM, and CORBA on Ethernet LANs. Berichte zur Rechnerarchitektur 3(4), Institut für Informatik, Lehrstuhl für Rechnerarchitektur und -kommunikation, Friedrich-Schiller-Universität Jena, Jena, Germany, 1997. 12 pp.

Fink:2000:IMC

[Fin00] Torsten Fink. Integrating MPI components into metacomputing applications. Lecture Notes in Computer Science, 1908:208–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080208.htm;
 http://link.springer-
ny.com/link/service/series/

0558/papers/1908/19080208.
pdf.

Fischer:2001:SAN

[Fis01] Markus Fischer. System area network extensions to the parallel virtual machine. Lecture Notes in Computer Science, 2131:98–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310098.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310098.
pdf.

Fernandez:2000:UPM

[FJBB^{+}00] Gustavo J. Fernández, Julio Jacobo-Berlles, Patricia Borensztejn, Marisa Bauzá, and Marta Mejail. Use of PVM for MAP image restoration: a parallel implementation of the ARTUR algorithm. Lecture Notes in Computer Science, 1908:113–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080113.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080113.
pdf.

Forejt:2017:PPA

[FJK^{+}17] Vojtăch Forejt, Saurabh Joshi, Daniel Kroening, Ganesh Narayanaswamy, and Subodh Sharma. Precise predictive analysis for discovering communication deadlocks in MPI programs. ACM Transactions on Programming Languages and Systems, 39(4):15:1–15:??, September 2017. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Feng:2014:SBS

[FJZ^{+}14] Xiaowen Feng, Hai Jin, Ran Zheng, Zhiyuan Shao, and Lei Zhu. A segment-based sparse matrix–vector multiplication on CUDA. Concurrency and Computation: Practice and Experience, 26 (1):271–286, January 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Flower:1994:EJM

[FK94] Jon Flower and Adam Kolawa. Express is not just a message passing system: current and future directions in Express. Parallel Computing, 20(4):597– 614, April 31, 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=20&
issue=4&aid=860.

Ferenczi:1995:PAH

[FK95] Szabolcs Ferenczi and Peter Kacsuk, editors. Proceedings of the 2nd Austrian-Hungarian Workshop on Transputer Applications: September 29–October 1, 1994, Budapest, Hungary. Hungarian Academy of Sciences, Central Research Intitute for Physics, Budapest, Hungary, 1995. ISBN ???? LCCN ???? Technical report KFKI-1995-2/M,N.

Fischer:2001:DNM

[FK01] Markus Fischer and Peter Kemper. Distributed numerical Markov chain analysis. Lecture Notes in Computer Science, 2131:272–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310272.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310272.
pdf.

Field:2002:OSR

[FKH02] A. J. Field, P. H. J. Kelly, and T. L. Hansen. Optimising shared reduction variables in MPI programs. Lecture Notes in Computer Science, 2400:630–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/

                                        link/service/series/0558/
                                        bibs/2400/24000630.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2400/24000630.
                                        pdf.

Foster:1996:MCL

[FKK96a] I. T. Foster, D. R. Kohr, Jr., and R. Krishnaiyer. MPI as a coordination layer for communicating HPF tasks. In IEEE [IEE96i], pages 68–78. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Foster:1996:CDT

[FKK^{+}96b] I. T. Foster, D. R. Kohr, Jr., R. Krishnaiyer, Choudhary, and A. Communicating data-parallel tasks: an MPI library for HPF. In IEEE [IEE96a], pages 433–

438. ISBN 0-8186-7557-

8. LCCN QA76.88.I575 1996. IEEE catalog number 96TB100074.

Foster:1996:DSB

[FKKC96] Ian Foster, David R. Kohr, Jr., Rakesh Krishnaiyer, and Alok Choudhary. Double standards: Bringing task parallelism to HPF via the message passing interface. In ACM [ACM96c], page ?? ISBN 0-89791-854-1. LCCN QA 76.88 S8573 1996. URL http://

                                        www.supercomp.org/sc96/
                                        proceedings/SC96PROC/FOSTER2/
                                        INDEX.HTM. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Freeh:2008:JTD

[FKLB08] Vincent W. Freeh, Nandini Kappiah, David K. Lowenthal, and Tyler K. Bletsch. Just-in-time dynamic voltage scaling: Exploiting internode slack to save energy in MPI programs. Journal of Parallel and Distributed Computing, 68(9): 1175–1185, September 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Foster:1996:GCM

[FKS96] I. Foster, C. Kesselman, and M. Snir. Generalized communicators in the Message Passing Interface. In IEEE [IEE96i], pages 42–49. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Ferreira:2021:EMR

[FL21] Kurt B. Ferreira and Scott Levy. Evaluating MPI resource usage summary statistics. Parallel Computing, 108(??):??, December 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167819121000739.

Florez:2005:LMM

[FLB^{+}05] German Florez, Zhen Liu, Susan M. Bridges, Anthony Skjellum, and Rayford B. Vaughn. Lightweight monitoring of MPI programs in real time. Concurrency and Computation: Practice and Experience, 17(13): 1547–1578, November 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fagg:1996:TGR

[FLD96] G. E. Fagg, K. S. London, and J. J. Dongarra. Taskers and general resource managers: PVM supporting DCE process management. In Bode et al. [BDLS96], pages 180–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Fagg:1998:MMH

[FLD98] G. E. Fagg, K. S. London, and J. J. Dongarra. MPI-Connect: Managing heterogeneous MPI applications interoperation and process control. Lecture Notes in Computer Science, 1497:93– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fachada:2017:CCF

[FLMR17] Nuno Fachada, Vitor V. Lopes, Rui C. Martins, and Agostinho C. Rosa. cf4ocl: a C framework for OpenCL. Science of Computer Programming, 143(??):

9–19, September 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167642317300540.

Ferreira:2018:CMM

[FLPG18] Kurt B. Ferreira, Scott Levy, Kevin Pedretti, and Ryan E. Grant. Characterizing MPI matching via trace-based simulation. Parallel Computing, 77(??):57– 83, September 2018. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819118301467.

Fan:2020:ALC

[FLS20] Q. Fan, D. J. Lilja, and S. S. Sapatnekar. Adaptivelength coding of image data for low-cost approximate storage. IEEE Transactions on Computers, 69(2): 239–252, February 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Feeley:1990:PVM

[FM90] Marc Feeley and James S.

Miller. A parallel virtual machine for efficient Scheme compilation. In ACM [ACM90], pages 119–

130. ISBN 0-89791-368-X. LCCN QA 76.73 L23 A24

1990. URL http://www.
acm.org/pubs/citations/

                                        proceedings/lfp/91556/
                                        p119-feeley/. ACM order

no. 552900.

Furlinger:2009:CAE

[FM09] Karl Fürlinger and Shirley Moore. Capturing and analyzing the execution control flow of OpenMP applications. International Journal of Parallel Programming, 37(3):266–276, June 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        37&issue=3&spage=266.

Fabero:1996:DLB

[FMBM96] J. C. Fabero, I. Martin, A. Bautista, and S. Molina. Dynamic load balancing in a heterogeneous environment under PVM. In IEEE [IEE96g], pages 414–419. ISBN 0-8186-7376-1. LCCN QA76.58 .E97 1996. IEEE order number PR07376.

Fiala:2012:DCS

[FME^{+}12] David Fiala, Frank Mueller, Christian Engelmann, Rolf Riesen, Kurt Ferreira, and Ron Brightwell. Detection and correction of silent data corruption for largescale high-performance computing. In Hollingsworth [Hol12], pages 78:1–78:?? ISBN 1-4673-0804-8. URL

                                        http://conferences.computer.

org/sc/2012/papers/1000a046.
pdf.

Filipovic:2015:OCC

[FMFM15] Jirí Filipovic, Matús Madzin, Jan Fousek, and Ludek Matyska. Optimizing CUDA code by kernel fusion: application on BLAS. The Journal of Supercomputing, 71(10):3934–3957, October 2015. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11227-015-1483-z.

Ferretti:2015:MCH

[FMS15] Marco Ferretti, Mirto Musci, and Luigi Santangelo. MPI– CMS: a hybrid parallel approach to geometrical motif search in proteins. Concurrency and Computation: Practice and Experience, 27(18):5500–5516, December 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fan:2017:SEE

[FMSG17] Xing Fan, Mostafa Mehrabi, Oliver Sinnen, and Nasser Giacaman. Supporting enhanced exception handling with OpenMP in object– oriented languages. International Journal of Parallel Programming, 45(6): 1366–1389, December 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Ferenc:1999:VMK

[FNSW99] D. Ferenc, J. Nabrzyski, M. Stroinski, and P. Wierzejewski. Visual MPI, a knowledge-based system for writing efficient MPI applications. In Dongarra et al. [DLM99], pages 257–

266. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Femminella:1994:PBP

[FO94] A. Femminella and A. Omodeo. PVM-based parallel computing: a case study on power plant simulation. Microprocessing and Microprogramming, 40(10-12):875– 878, December 1994. CO-DEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic).

Ford:1995:NNN

[For95]  Brian Ford.  The new

NAG numerical PVM library (or A new parallel numerical library based on PVM). In IFIP Working Group 2.5 [IFI95], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        nsc.liu.se/~boein/ifip/
                                        kyoto/workshop-info/proceedings/
                                        ford/ford1.html.

Foster:1998:GEM

[Fos98] Ian Foster. A grid-enabled MPI: Message passing in heterogeneous distributed computing systems. In ACM [ACM98b], page ?? ISBN ???? LCCN

????  URL http://
www.supercomp.org/sc98/
papers/.

Freeman:1992:PNA

[FP92] T. L. (Len) Freeman and C. (Christopher) Phillips. Parallel numerical algorithms. Prentice Hall International Series in Computer Science. Prentice-Hall International, Englewood Cliffs, NJ 07632, USA, 1992. ISBN 0-13-651597-5. xii + 315 pp. LCCN QA76.9.A43 F74 1992. US$40.00. Chapter 5 discusses HPF and PVM.

Faraj:2008:SPA

[FPY08] Ahmad Faraj, Pitch Patarasuk, and Xin Yuan. A study of process arrival patterns for MPI collective operations. International Journal of Parallel Programming, 36 (6):543–570, December 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=6&spage=543.

Ferreira:1995:PAI

[FR95] Afonso Ferreira and Jose Rolim, editors. Parallel algorithms for irregularly structured problems: second international workshop, IRREGULAR 95, Lyon, France, September, 4–6, 1995: proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-60321-2. LCCN QA76.642.I59 1995.

Franke:1995:MPEa

[Fra95] Hubertus Franke. MPI programming environment for IBM SP1/SP2. Research report RC 19991 (88480), IBM T. J. Watson Research Center, Yorktown Heights, NY, USA, 1995. 9 pp.

Fritscher:1993:PDC

[FS93] J. F. Fritscher and F. Sukup. 93SC038 parallel distributed computing using PVM. In Anonymous [Ano93a], pages 221–228. ISBN 0-947719-62-8. LCCN ????

Ferrari:1995:TDC

[FS95] A. J. Ferrari and V. S. Sunderam. TPVM: distributed concurrent computing with lightweight processes. In IEEE [IEE95k], pages 211– 218. ISBN 0-8186-7088-6. LCCN QA76.9.D5 I328 1995. IEEE catalog no. 95TB8075.

Fischer:1997:ESP

[FS97] M. Fischer and J. Simon. Embedding SCI into PVM.

Lecture Notes in Computer Science, 1332:177–184, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrari:1998:MDC

[FS98] Adam Ferrari and V. S.

Sunderam. Multiparadigm distributed computing with TPVM. Concurrency: practice and experience, 10(3): 199–228, March 1998. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract?ID=5374;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5374&PLACEBO=IE.pdf.

Filgueira:2011:ACE

[FSC^{+}11] Rosa Filgueira, David E.

Singh, Jesús Carretero, Alejandro Calderón, and Félix García. Adaptive-CoMPI: Enhancing MPI-based applications’ performance and scalability by using adaptive compression. The International Journal of High Performance Computing Applications, 25(1):93– 114, February 2011. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/25/
1/93.full.pdf+html.

Fan:2019:BPA

[FSG19a] Xing Fan, Oliver Sinnen, and Nasser Giacaman. Balancing parallelization and asynchronization in event-driven programs with OpenMP. Concurrency and Computation: Practice and Experience, 31 (4):e4959:1–e4959:??, February 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fan:2019:SAO

[FSG19b] Xing Fan, Oliver Sinnen, and Nasser Giacaman. Supporting asynchronization in OpenMP for event-driven programming. Parallel Computing, 82(??):57–74, ???? 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167819118300838.

Fuerle:1998:IPC

[FSLS98] T. Fuerle, E. Schikuta, C. Loeffelhardt, and K. Stockinger. On the implementation of a portable, client-server based MPI-IO interface. Lecture Notes in Computer Science, 1497:172–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fumero:2017:JTG

[FSSD17] Juan Fumero, Michel Steuwer, Lukas Stadler, and Christophe Dubach. Just-in-time GPU compilation for interpreted languages with partial evaluation. ACM SIGPLAN Notices, 52(7):60–73, July 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Folino:1998:EMC

[FST98a] G. Folino, G. Spezzano, and D. Talia. Evaluating and modeling communication overhead of MPI primitives on the Meiko CS-2. Lecture Notes in Computer Science, 1497:27–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Folino:1998:PEM

[FST98b] G. Folino, G. Spezzano, and D. Talia. Performance evaluation and modelling of MPI communications on the Meiko CS-2. Lecture Notes in Computer Science, 1401:932–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1999:PGP

[FSTG99] F. Fernandez, J. M. Sanchez, M. Tomassini, and J. A. Gomez. A parallel genetic programming tool based on PVM. In Dongarra et al. [DLM99], pages 241–248. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Fang:2014:API

[FSV14] Jianbin Fang, Henk Sips, and Ana Lucia Varbanescu. Aristotle: A performance impact indicator for the OpenCL kernels using local memory. Scientific Programming, 22(3):239–257, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Feng:2014:MSP

[FSXZ14] Chunsheng Feng, Shi Shu, Jinchao Xu, and Chen-Song Zhang. A multi-stage preconditioner for the black oil model and its OpenMP implementation. In Erhel et al. [EGH^{+}14], pages 141–153. ISBN 3-319-05788-X (paperback), 3-319-05789-8 (e-book). ISSN 1439-7358 (print), 2197-7100 (electronic). LCCN QA71-

90. URL http://link.
springer.com/chapter/10.
1007/978-3-319-05789-7_
11/.

Fernandez:2000:DCE

[FTVB00] Francisco Fernández, Marco Tomassini, Leonardo Vanneschi, and Laurent Bucher. A distributed computing environment for genetic programming using MPI. Lecture Notes in Computer Science, 1908:322–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080322.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080322.
pdf.

Fujimoto:2008:DMV

[Fuj08] Noriyuki Fujimoto. Dense matrix-vector multiplication on the CUDA architecture. Parallel Processing Letters, 18(4):511–530, December 2008. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Fagg:2000:AAC

[FVD00] Graham E. Fagg, Sathish S.

Vadhiyar, and Jack J. Dongarra. ACCT: Automatic Collective Communications Tuning. Lecture Notes in Computer Science, 1908:354–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080354.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080354.
pdf.

Fang:2015:EVD

[FVLS15] Jianbin Fang, Ana Lucia Varbanescu, Xiangke Liao, and Henk Sips. Evaluating vector data type usage in OpenCL kernels. Concurrency and Computation: Practice and Experience, 27(17):4586–4602, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fineberg:1996:PPI

[FWNK96] S. A. Fineberg, P. Wong, B. Nitzberg, and C. Kuszmaul. PMPIO-a portable implementation of MPI-IO. In IEEE [IEE96c], pages 188–195. ISBN 0-8186-7551-9. LCCN QA76.58 .S95 1996. IEEE catalog number 96TB100062.

Franke:1995:MPEb

[FWR^{+}95] Hubertus Franke, C. Eric Wu, Michel Riviere, Pratap Pattnaik, and Marc Snir. MPI programming environment for IBM SP1/SP2. In IEEE [IEE95i], pages 127– 135. ISBN 0-8186-7025-8. LCCN ???? IEEE catalog number 95CH35784.

Frust:2017:RDP

[FWS^{+}17] Tobias Frust, Michael Wagner, Jan Stephan, Guido Juckeland, and André Bieberle. Rapid data processing for ultrafast X-ray computed tomography using scalable and modular CUDA based pipelines. Computer Physics Communications, 219(??): 353–360, October 2017. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465517301674.

Grangeat:1996:PTI

[GA96] Pierre Grangeat and Jean-Louis Amans, editors. Proceedings of the Third International Meeting on Fully Three-Dimensional Image Reconstruction in Radiology and Nuclear Medicine, held July 4–6, 1995 at Domaine d’Aix-Marlioz, Aixles-Bains, France. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-4129-5. LCCN R857.T47 T485 1996.

Gutierrez:2020:MAP

[GADM20] Samuel K. Gutiérrez, Dorian C. Arnold, Kei Davis, and Patrick McCormick. On the memory attribution problem: a solution and case study using MPI. Concurrency and Computation: Practice and Experience, 32 (3):e5159:1–e5159:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Galibert:1997:YCL

[Gal97] O. Galibert. YLC, A C++ Linda system on top of PVM. Lecture Notes in Computer Science, 1332: 99–106, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalez:2000:NSF

[GAM^{+}00] Marc Gonzàlez, Eduard Ayguadé, Xavier Martorell, Jesús Labarta, Nacho Navarro, and José Oliver. NanosCompiler: supporting flexible multilevel parallelism exploitation in OpenMP. Concurrency: practice and experience, 12(12):1205–1218, October 2000. CODEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76500358/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500358&PLACEBO=IE.
pdf.

Gonzalez:2002:DLP

[GAM^{+}02] Marc González, Eduard Ayguadé, Xavier Martorell, Jesús Labarta, and Phu V. Luong. Dual-level parallelism exploitation with OpenMP in coastal ocean circulation modeling. Lecture Notes in Computer Science, 2327:469–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2327/23270469.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2327/23270469.
pdf.

Gonzalez:2001:DSP

[GAML01] M. Gonzalez, E. Ayguadé, X. Martorell, and J. Labarta. Defining and supporting pipelined executions in OpenMP. Lecture Notes in Computer Science, 2104: 155–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2104/21040155.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2104/21040155.
pdf.

Gonzalez:2000:PAM

[GAMR00] Daniel González, Francisco Almeida, Luz Marina Moreno, and Casiano Rodríguez. Pipeline algorithms on MPI: Optimal mapping of the path planing problem. Lecture Notes in Computer Science, 1908:104–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080104.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080104.
pdf.

Gao:2003:LSP

[Gao03] Shiwu Gao. Linear-scaling parallelization of the WIEN package with MPI. Computer Physics Communications, 153(2):190–198, June 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465503002248.

Galaktionov:1997:MST

[GAP97] A. S. Galaktionov, P. D. Anderson, and G. W. M. Peters. Mixing simulations: Tracking strongly deforming fluid volumes in 3D flows. Lecture Notes in Computer Science, 1332:436–469, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gates:1995:PFI

[Gat95] W. Lawrence (William Lawrence) Gates, editor. Proceedings of the First International AMIP Scientific Conference: Monterey, California, USA, 15–19 May 1995, number 732 in World Meteorological Organization — Publications — WMO TD 1995. World Meteorological Organization, Geneva, Switzerland, 1995. ISBN ???? LCCN SIO 1 WO326 v.92.

Gonzalez-Alvarez:2017:HMO

[GÁVRRL17] David L. González-Álvarez, Miguel A. Vega-Rodríguez, and Álvaro Rubio-Largo. A hybrid MPI/OpenMP parallel implementation of NSGA–II for finding patterns in protein sequences. The Journal of Supercomputing, 73(6):2285–2312, June 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gupta:1994:CTE

[GB94] M. Gupta and P. Banerjee.

Compile-time estimation of communication costs of programs. Journal of Programming Languages, 2(3):191– 225, September 1994. CO-DEN JPLAER. ISSN 0963-9306.

Ghosh:1996:ELM

[GB96] K. Ghosh and S. Breit. Evaluating the limits of message passing via the shared attraction memory on CC-COMA machines: Experiences with TCGMSG and PVM. In ACM [ACM96b], pages 173–180. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Gorlatch:1998:GMI

[GB98] Sergei Gorlatch and Holger Bischof. A generic MPI implementation for a data-parallel skeleton: Formal derivation and application to FFT. Parallel Processing Letters, 8(4):447–??, December 1998. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Geist:1994:PPV

[GBD^{+}94] Al Geist, Adam Beguelin, Jack Dongarra, Weicheng Jiang, Robert Manchek, and Vaidyalingam S. Sunderam. PVM: Parallel Virtual Machine: a Users’ Guide and Tutorial for Networked Parallel Computing. Scientific and engineering computation. MIT Press, Cambridge, MA, USA, 1994. ISBN 0-262-57108-0 (paperback). xvii + 279 pp. LCCN QA76.58 .P85 1994. US$27.50. URL http:/

                                        /www.mitpress.com/book-
                                        home.tcl?isbn=0262571080.

Gentzsch:1995:STP

[GBF95] W. Gentzsch, U. Block, and F. Ferstl. Software tools for parallel computers and workstation clusters. In Ferenczi and Kacsuk [FK95], pages 23–42. ISBN ???? LCCN ???? Technical report KFKI-1995-2/M,N.

Golebiewski:1999:HPI

[GBH99] M. Golebiewski, M. Baum, and R. Hempel. High performance implementation of MPI for Myrinet. Lecture Notes in Computer Science, 1557:510–521, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerstenberger:2014:EHS

[GBH14] Robert Gerstenberger, Maciej Besta, and Torsten Hoefler. Enabling highlyscalable remote memory access programming with MPI-3 One Sided. Scientific Programming, 22(2):75–91, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gerstenberger:2018:EHS

[GBH18] Robert Gerstenberger, Maciej Besta, and Torsten Hoefler. Enabling highly scalable remote memory access programming with MPI-3 one sided. Communications of the ACM, 61 (10):106–113, October 2018. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

cacm.acm.org/magazines/
2018/10/231375/fulltext.

Gabriel:1997:EMU

[GBR97] Edgar Gabriel, Thomas Beisel, and Michael Resch. Erweiterung einer MPI-Umgebung zur Interoperabilität verteilter MPP-Systeme. (German) [Extension of an MPI environment for interoperability with distributed MPI systems]. Studienarbeit angewandte Informatik RUS 37, Rechenzentrum Universität Stuttgart, Stuttgart, Germany, 1997.

Garain:2015:CCF

[GBR15] Sudip Garain, Dinshaw S.

Balsara, and John Reid. Comparing Coarray Fortran (CAF) with MPI for several structured mesh PDE applications. Journal of Computational Physics, 297(??):237–253, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002199911500354X.

Graham:2007:OMH

[GBS^{+}07] Richard L. Graham, Brian W. Barrett, Galen M. Shipman, Timothy S. Woodall, and George Bosilca. Open MPI: a high performance, flexible implementation of MPI point-to-point communications. Parallel Processing Letters, 17(1):79– 88, March 2007. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Grove:2005:CBP

[GC05] D. A. Grove and P. D. Coddington. Communication benchmarking and performance modelling of MPI programs on cluster computers. The Journal of Supercomputing, 34(2): 201–217, November 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&

issn=0920-8542&volume=
34&issue=2&spage=201.

Garcia:2012:DLB

[GCBL12] Marta Garcia, Julita Corbalan, Rosa Maria Badia, and Jesus Labarta. A dynamic load balancing approach with SMPSuperscalar and MPI. Lecture Notes in Computer Science, 7174:10–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-30397-5_
2/.

GarciaSalcines:1997:PRR

[GCBM97] E. Garcia Salcines, G. Cerruela Garcia, J. I. Benavides Benitez, and F. Muñoz Garcia. Parallel rendering of radiance on distributed memory system by PVM. Lecture Notes in Computer Science, 1332:502–507, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia:1999:MMI

[GCC99] F. Garcia, A. Calderon, and J. Carretero. MiMPI: a multithread-safe implementation of MPI. In Dongarra et al. [DLM99], pages 207– 214. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Garcia-Consuegra:1998:DGR

[GCGS98] J. D. Garcia-Consuegra, J. A. Gallud, and G. Sebastian. Distributed georeferring of remotely sensed Landsat-TM imagery using MPI. Lecture Notes in Computer Science, 1541: 161–166, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gelado:2010:ADS

[GCN^{+}10] Isaac Gelado, Javier Cabezas, Nacho Navarro, John E. Stone, Sanjay Patel, and Wen mei W. Hwu. An asymmetric distributed shared memory model for heterogeneous parallel systems. ACM SIGPLAN Notices, 45 (3):347–358, March 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gao:2013:GGA

[GCN^{+}13] Mingcen Gao, Thanh-Tung Cao, Ashwin Nanjappa, Tiow-Seng Tan, and Zhiyong Huang. gHull: a GPU algorithm for 3D convex hull. ACM Transactions on Mathematical Software, 40(1): 3:1–3:19, September 2013. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Geist:1993:PTW

[GDB^{+}93] A. Geist, J. Dongarra, A. Beguelin, B. Manchek, and Weicheng Jiang. PVM takes over the world. In IEEE [IEE93e], page 618. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Galizia:2015:MCL

[GDC15] Antonella Galizia, Daniele D’Agostino, and Andrea Clematis. An MPI–CUDA library for image processing on HPC architectures. Journal of Computational and Applied Mathematics, 273(??):414–427, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042714002374.

Ghose:2017:FOT

[GDDM17] Anirban Ghose, Lokesh Dokara, Soumyajit Dey, and Pabitra Mitra. A framework for OpenCL task scheduling on heterogeneous multicores. Parallel Processing Letters, 27(3–4):1750008, 2017. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Gonzalez-Dominguez:2020:CJA

[GDEBC20] Jorge González-Domínguez, Roberto R. Expósito, and Verónica Bolón-Canedo. CUDA-JMI: Acceleration of feature selection on heterogeneous systems. Future Generation Computer Systems, 102(??): 426–436, January 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167739X19312968.

Gonzalez-Dominguez:2018:MPC

[GDM18] Jorge Gonzalez-Dominguez and Maria J. Martin. MPI-GeneNet: Parallel calculation of gene co-expression networks on multicore clusters. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 15(5):1732–1737, September 2018. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Gonzalez-Dominguez:2022:MDP

[GDMME22] Jorge González-Domínguez, José M. Martín-Martínez, and Roberto R. Expósito.

MPI-dot2dot: A parallel

tool to find DNA tandem repeats on multicore clusters. The Journal of Supercomputing, 78(3):4217– 4235, February 2022. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

https://link.springer.
com/article/10.1007/s11227-
021-04025-7.

Gawande:2020:SDL

[GDS^{+}20] Nitin A. Gawande, Jeff A.

Daily, Charles Siegel, Nathan R. Tallent, and Abhinav Vishnu. Scaling deep learning workloads: NVIDIA DGX-1/ Pascal and Intel Knights Landing. Future Generation Computer Systems, 108(??): 1162–1172, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167739X17318599.

Grinstein:1995:VDE

[GE95] Georges G. Grinstein and Robert F. Erbacher, editors. Visual data exploration and analysis II: 8–10 February 1995, San Jose, California, volume 2410 of Proceedings of the SPIE — The International Society for Optical Engineering. Society of Photooptical Instrumentation Engineers (SPIE), Bellingham, WA, USA, 1995. CODEN PSISDG. ISBN 0-8194-1757-2. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.2410.

Grinstein:1996:VDE

[GE96] Georges G. Grinstein and Robert F. Erbacher, editors. Visual data exploration and analysis III: 31 January–2 February, 1996, San Jose, California, volume 2421 (or 2656??) of Proceedings of the SPIE — The International Society for Optical Engineering. Society of Photooptical Instrumentation Engineers (SPIE), Bellingham, WA, USA, 1996. CODEN PSISDG. ISBN 0-8194-2030-1. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.2656.

Geist:1993:ILP

[Gei93a] G. A. Geist. Invited lecture: PVM 3 beyond network computing. In Volkert [Vol93], pages 194– 203. ISBN 3-540-57314-3 (Berlin), 0-387-57314-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.734. DM58.00.

Geist:1993:PBN

[Gei93b] G. A. Geist. PVM 3 beyond network computing. In Volkert [Vol93], pages 194– 203. ISBN 3-540-57314-3 (Berlin), 0-387-57314-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.734. DM58.00.

Geist:1994:CCW

[Gei94] G. A. Geist. Cluster computing: the wave of the future? In Dongarra and Wasniewski [DW94], pages 236–246. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1994. DM104.00.

Geist:1996:APP

[Gei96] G. A. Geist. Advanced programming in PVM. In Bode et al. [BDLS96], pages 1–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Geist:1997:ACP

[Gei97] G. A. Geist. Advanced capabilities in PVM 3.4. Lecture Notes in Computer Science, 1332:107–115, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geist:1998:HNG

[Gei98]  G. A. Geist.  Harness:

The next generation beyond PVM. Lecture Notes in Computer Science, 1497:74– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geist:2000:PMW

[Gei00] Al Geist. PVM and MPI:

What else is needed for cluster computing? Lecture Notes in Computer Science, 1908:1–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080001.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080001.
pdf.

Geist:2001:BFN

[Gei01] G. Al Geist. Building a foundation for the next PVM: Petascale Virtual Machines. Lecture Notes in Computer Science, 2131:2–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310002.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310002.
pdf.

Gerbessiotis:2018:SIS

[Ger18] Alexandros V. Gerbessiotis. A study of integer sorting on multicores. Parallel Processing Letters, 28(04):??, December 2018. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.
com/doi/10.1142/S0129626418500147.

Grabowsky:1998:NMP

[GEW98] Lothar Grabowsky, Thomas Ermer, and Jörg Werner. Nutzung von MPI für parallele FEM-Systeme. (German) [Use of MPI for parallel FEM systems]. Preprint-Reihe des Chemnitzer SFB 393 Sonderforschungsbereich NumerischeSimulation auf Massiv Parallelen Rechnern 97,08; RA-TR 02-97, Universität Chemnitz-Zwickau, Chemnitz, Germany, 1998.

Gabriel:2003:FTC

[GFB^{+}03] Edgar Gabriel, Graham E.

Fagg, Antonin Bukovsky, Thara Angskun, and Jack J. Dongarra. A fault-tolerant communication library for Grid environments. In ????, editor, 17th Annual ACM International Conference on Supercomputing (ICS’03) International Workshop on Grid Computing and e-Science, June 21, 2003, San Francisco, page ?? ????, ????, 2003. ISBN ???? LCCN ????

URL http://www.netlib.
org/netlib/utk/people/
JackDongarra/PAPERS/FTMPI-
SF-gabriel.pdf.

Gonina:2014:SMC

[GFB^{+}14] Ekaterina Gonina, Gerald Friedland, Eric Battenberg, Penporn Koanantakool, Michael Driscoll, Evangelos Georganas, and Kurt Keutzer. Scalable multimedia content analysis on parallel platforms using Python. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(2):18:1–18:??, February 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gabriel:2003:EPM

[GFD03] Edgar Gabriel, Graham Fagg, and Jack Dongarra. Evaluating the performance of MPI-2 dynamic communicators and one-sided communication. In Dongarra et al. [DLO03], page ?? CO-DEN LNCSD9. ISBN 3-540-20149-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E973 2003. URL http:/

                                        /www.netlib.org/netlib/
                                        utk/people/JackDongarra/
                                        PAPERS/europvm-mpi-2003-
                                        mpi2.pdf.

Gabriel:2005:EDC

[GFD05] Edgar Gabriel, Graham E. Fagg, and Jack J. Dongarra. Evaluating dynamic communicators and onesided operations for current MPI libraries. The International Journal of High Performance Computing Applications, 19(1):67– 79, Spring 2005. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

                                        tronic). URL http://hpc.
                                        sagepub.com/content/19/
                                        1/67.full.pdf+html.

Gomez-Folgar:2018:MPA

[GFIS^{+}18] F. Gomez-Folgar, G. Indalecio, N. Seoane, T. F. Pena, and A. J. Garcia-Loureiro. MPI-Performance-Aware-Reallocation: method to optimize the mapping of processes applied to a cloud infrastructure. Computing, 100(2):211–226, February 2018. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gueunet:2019:TBA

[GFJT19] C. Gueunet, P. Fortin, J. Jomier, and J. Tierny. Task-based augmented contour trees with Fibonacci heaps. IEEE Transactions on Parallel and Distributed Systems, 30(8):1889–1905, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gravvanis:2012:SFD

[GFPG12] G. A. Gravvanis, C. K.

Filelis-Papadopoulos, and K. M. Giannoutakis. Solving finite difference linear systems on GPUs: CUDA based parallel explicit preconditioned biconjugate conjugate gradient type methods. The Journal of Supercomputing, 61(3):590– 604, September 2012. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
61&issue=3&spage=590.

Giordano:1999:IBP

[GFV99] M. Giordano, M. M. Furnari, and F. Vitobello. Interaction between PVM parameters and communication performances on ATM networks. Lecture Notes in Computer Science, 1557:586–587, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garzon:1999:PIE

[GG99] E. M. Garzon and I. Garcia. A parallel implementation of the eigenproblem for large, symmetric and sparse matrices. In Dongarra et al. [DLM99], pages 380–387. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999. Giannoutakis:2009:DIP

[GG09] Konstantinos M. Giannoutakis and George A. Gravvanis. Design and implementation of parallel approximate inverse classes using OpenMP. Concurrency and Computation: Practice and Experience, 21(2): 115–131, February 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Giannoutakis:2007:MHP

[GGC^{+}07] K. M. Giannoutakis, G. A. Gravvanis, B. Clayton, A. Patil, T. Enright, and J. P. Morrison. Matching high performance approximate inverse preconditioning to architectural platforms. The Journal of Supercomputing, 42(2): 145–163, November 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
42&issue=2&spage=145.

Gallud:2001:EDF

[GGCGO01] J. A. Gallud, J. García-Consuegra, J. M. García, and L. Orozco. Evaluating the DIPORSI framework: Distributed processing of remotely sensed imagery. Lecture Notes in Computer Science, 2131:401–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310401.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310401.
pdf.

Gallud:1999:DPR

[GGCM99] J. A. Gallud, J. Garcia-Consuegra, and A. Martinez. Distributed processing of remotely sensed Landsat-TM imagery using MPI. Parallel and Distributed Computing Practices, 2(2):??, ???? 1999. CODEN ???? ISSN 1097-

                                        2803. URL http://www.cs.
                                        okstate.edu/~pdcp/vols/
                                        vol02/vol02no2abs.html#
                                        gallud.

Gallud:1999:CCU

[GGGC99] J. A. Gallud, J. M. Garcia, and J. Garcia-Consuegra. Cluster computing using MPI and Windows NT to solve the processing of remotely sensed imagery. In Dongarra et al. [DLM99], pages 442–449. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Godlevsky:1999:PSA

[GGH99] A. Godlevsky, M. Gazak, and L. Hluchy. Parallelizing of sequential annotated programs in PVM environment. In Dongarra et al. [DLM99], pages 517–524. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Geist:1996:MEM

[GGHL^{+}96] A. Geist, W. Gropp, S. Huss-Lederman, A. Lumsdaine, E. Lusk, W. Saphir, T. Skjellum, and M. Snir. MPI-2: extending the Message-Passing Interface. In Bouge et al. [BFMR96], pages 128–135. ISBN 3-540-61626-8 (vol. 1), 3-540-61627-6 (vol. 2). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I554 1996, QA267.A1 L43 no.1123-1124. Two volumes.

Gawman:1993:PCT

[GGK^{+}93] Ann Gawman, W. Morven Gentleman, E. Kidd, Per-Åke Larson, and J. Slonim, editors. Proceedings CAS-CON ’93: Toronto, Ontario, Canada, 24–28 October 1993. Nat. Res. Council of Canada, Ottawa, Ont., Canada, 1993. ISBN ???? LCCN QA76.76.S64 C378 1993 v.1-2. Two volumes.

Genaud:2008:EPC

[GGL^{+}08] Stéphane Genaud, Pierre Gançarski, Guillaume Latu, Alexandre Blansché, Choopan Rattanapoka, and Damien Vouriot. Exploitation of a parallel clustering algorithm on commodity hardware with P2P-MPI. The Journal of Supercomputing, 43(1):21–41, January 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
43&issue=1&spage=21.

Getov:1999:MJM

[GGS99] Vladimir Getov, Paul Gray, and Vaidy Sunderam. MPI and Java-MPI: Contrasts and comparisons of lowlevel communication performance. In ACM [ACM99], page ??

Gao:2020:MES

[GGZ^{+}20] T. Gao, Y. Guo, B. Zhang, P. Cicotti, Y. Lu, P. Balaji, and M. Taufer. Memoryefficient and skew-tolerant MapReduce over MPI for supercomputing systems. IEEE Transactions on Parallel and Distributed Systems, 31(12):2734–2748, 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gentzsch:1994:HPC

[GH94] Wolfgang Gentzsch and Uwe Harms, editors. Highperformance computing and networking: international conference and exhibition, Munich, Germany, April 18–20, 1994: proceedings, volume 797 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-57981-8 (New York), 3-540-57981-8 (Berlin). LCCN QA76.88.I57 1994. DM96.00. Two volumes.

Ghosh:2012:RAA

[GHD12] Sudeep Ghosh, Jason Hiser, and Jack W. Davidson. Replacement attacks against VM-protected applications. ACM SIGPLAN Notices, 47(7):203–214, July 2012.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Grebe:1993:TAS

[GHH^{+}93] R. Grebe, J. Hektor, S. C.

Hilton, M. R. Jane, and P. H. Welch, editors. Transputer applications and systems ’93: proceedings of the 1993 World Transputer Congress, 20–22 September 1993, Aachen, Germany. IOS Press, Postal Drawer 10558, Burke, VA 2209-0558, USA, 1993. ISBN 90-5199-140-1. LCCN ????

Goumopoulos:1997:PCS

[GHL97] C. Goumopoulos, E. Housos, and O. Liljenzin. Parallel crew scheduling on workstation networks using PVM. Lecture Notes in Computer Science, 1332:470–477, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gropp:1998:MCR

[GHLL^{+}98] William Gropp, Steven Huss-Lederman, Andrew Lumsdaine, Ewing Lusk, Bill Nitzberg, William Saphir, and Marc Snir. MPI: The Complete Reference. Volume 2, The MPI-2 Extensions. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, second edition, 1998. ISBN 0-262-57123-4 (vol. 2), 0-262-69216-3 (set). 350 pp. LCCN QA76.642 .M65 1998. US$30 (paperback).

                                        URL http://mitpress.
                                        mit.edu/book-home.tcl?
                                        isbn=0262571234. See also

volume 1 [SOHL^{+}98].

Gong:2012:OCN

[GHZ12] Yifan Gong, Bingsheng He, and Jianlong Zhong. An overview of CMPI: network performance aware MPI in the cloud. ACM SIG-PLAN Notices, 47(8):297– 298, August 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Garcia:2011:KRR

[GJLT11] Saturnino Garcia, Donghwan Jeon, Christopher M. Louie, and Michael Bedford Taylor. Kremlin: rethinking and rebooting gprof for the multicore age. ACM SIG-PLAN Notices, 46(6):458– 469, June 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goglin:2018:HTM

[GJMM18] Brice Goglin, Emmanuel Jeannot, Farouk Mansouri, and Guillaume Mercier. Hardware topology management in MPI applications through hierarchical communicators. Parallel Computing, 76(??):70–90, August 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167819118301480.

Grecki:1997:MPE

[GJN97] M. Grecki, G. Jablonski, and A. Napieralski. MOPS

— parallel environment for simulation of electronic circuits using physical models of semiconductor devices. Lecture Notes in Computer Science, 1332:478–485, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerlach:2001:IOJ

[GJP01] Jens Gerlach, Zheng-Yu Jiang, and Hans-Werner Pohl. Integrating OpenMP into Janus. Lecture Notes in Computer Science, 2104: 101–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2104/21040101.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2104/21040101.
pdf.

Genaud:2009:FMP

[GJR09] Stéphane Genaud, Emmanuel Jeannot, and Choopan Rattanapoka. Fault-management in P2P-MPI. International Journal of Parallel Programming, 37(5):433– 461, October 2009. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=5&spage=433.

Gillett:1997:UMC

[GK97] Richard Gillett and Richard Kaufmann. Using the Memory Channel Network — using a cluster of standard PCI-based servers with a low-cost network to improve communication performance. IEEE Micro, 17(1): 19–25, January/February 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Granat:2010:PSS

[GK10] Robert Granat and Bo Kagstrom. Parallel solvers for Sylvestertype matrix equations with applications in condition estimation, Part I: Theory and algorithms. ACM Transactions on Mathematical Software, 37(3):32:1– 32:32, September 2010. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Grasso:2013:APS

[GKCF13] Ivan Grasso, Klaus Kofler, Biagio Cosenza, and Thomas Fahringer. Automatic problem size sensitive task partitioning on heterogeneous parallel systems. ACM SIG-PLAN Notices, 48(8):281– 282, August 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’13 Conference proceedings.

Gianinazzi:2018:CAP

[GKD^{+}18] Lukas Gianinazzi, Pavel Kalvoda, Alessandro De Palma, Maciej Besta, and Torsten Hoefler. Communicationavoiding parallel minimum cuts and connected components. ACM SIGPLAN Notices, 53(1):219–232, January 2018. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Granat:2009:NPQ

[GKK09] Robert Granat, Bo Kågström, and Daniel Kressner. A novel parallel QR algorithm for hybrid distributed memory HPC systems. LAPACK Working Note 216, Department of Computing Science and HPC2N, Umeå University, S-901 Umeå, Sweden, April 2009. URL http:/

/www.netlib.org/lapack/
lawnspdf/lawn216.pdf.

Gropp:1995:MGX

[GKL95] W. Gropp, E. Karrels, and E. Lusk. MPE graphicsscalable X11 graphics in MPI. In IEEE [IEE95j], pages 49–54. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Guan:1997:PDI

[GkLyCY97] Huiwei Guan, Chi kwong Li, To yat Cheung, and Songnian Yu. Parallel design and implementation of SOM neural computing model in PVM environment of a distributed system. In IEEE [IEE97a], pages 26–

31. ISBN 0-8186-7876-3 (paperback and case), 0-8186-7878-X (microfiche). LCCN QA76.58 .A4 1997.

Geist:1996:VDP

[GKP96] G. A. Geist, James Kohn, and Philip Papadopoulos. Visualization, debugging, and performance in PVM. Technical report, Oak Ridge National Laboratory, Knoxville, TN, USA, 1996.

11 pp. URL http://www.
epm.ornl.gov/~geist/CapeCod.
ps.

Geist:1997:CPF

[GKP97] G. A. Geist, II, James Arthur Kohl, and Philip M. Papadopoulos. CUMULVS: Providing fault tolerance, visualization, and steering of parallel applications. International Journal of Supercomputer Applications and High Performance Computing, 11(3):224–235, Fall 1997. CODEN IJSCFG. ISSN 1078-3482.

Geist:1997:BPW

[GKPS97] G. A. Geist, J. A. Kohl, P. M. Papadopoulos, and S. L. Scott. Beyond PVM 3.4: What we’ve learned, what’s next, and why. Lecture Notes in Computer Science, 1332:116–126, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gopalakrishnan:2011:FAM

[GKS^{+}11]  Ganesh  Gopalakrishnan,

Robert M. Kirby, Stephen Siegel, Rajeev Thakur, William Gropp, Ewing Lusk, Bronis R. De Supinski, Martin Schulz, and Greg Bronevetsky. Formal analysis of MPI-based parallel programs. Communications of the ACM, 54(12):82–91, December 2011. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Garland:2012:DUP

[GKZ12] Michael Garland, Manjunath Kudlur, and Yili Zheng. Designing a unified programming model for heterogeneous machines. In Hollingsworth [Hol12], pages 67:1–67:?? ISBN 1-4673-0804-8. URL http:

//conferences.computer.
org/sc/2012/papers/1000a064.
pdf.

Gropp:1992:TIM

[GL92] Bill Gropp and Ewing Lusk. A test implementation of the MPI draft message-passing standard. Technical report, Mathematics and Computer Science Division, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1992.

Gropp:1994:MCL

[GL94] W. Gropp and E. Lusk. The MPI communication library: its design and a portable implementation. In IEEE [IEE94f], pages 160– 165. ISBN 0-8186-4980-1. LCCN QA76.58.S34 1993.

Gropp:1995:DPM

[GL95a] W. Gropp and E. Lusk. Dynamic process management in an MPI setting. In IEEE [IEE95g], pages 530– 533. CODEN PSPDF8. ISBN 0-8186-7195-5. ISSN 1063-6374. LCCN QA 76.58 I42 1995. IEEE catalog number 95TB8131.

Gropp:1995:IMM

[GL95b] W. Gropp and E. Lusk. Implementing MPI: the 1994 MPI Implementors’ Workshop. In IEEE [IEE95j], pages 55–59. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Gropp:1995:MMI

[GL95c] W. Gropp and E. Lusk. The MPI message-passing interface standard: Overview and status. In Dongarra et al. [D^{+}95], pages 265– 270. ISBN 0-444-82163-5. ISSN 0927-5452. LCCN QA76.88.H55 1995.

Gropp:1995:EIS

[GL95d] W. D. Gropp and E. Lusk.

Experiences with the IBM SP1. IBM Systems Journal, 34(2):249–262, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http:

//www.research.ibm.com/
journal/sj34-2.html#seven.

Gropp:1996:HPM

[GL96] W. Gropp and E. Lusk. A high-performance MPI implementation on a sharedmemory vector supercomputer. Parallel Computing, 22(11):1513–??, ???? 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gropp:1997:SMC

[GL97a] W. Gropp and E. Lusk. Sowing MPICH: a case study in the dissemination of a portable environment for parallel scientific computing. International Journal of Supercomputer Applications and High Performance Computing, 11(2):103–114, Summer 1997. CODEN IJSCFG. ISSN 1078-3482.

Gropp:1997:WPM

[GL97b] W. Gropp and E. Lusk. Why are PVM and MPI so different? Lecture Notes in Computer Science, 1332:3–10, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gropp:1997:HPM

[GL97c] William Gropp and Ewing Lusk. A high-performance MPI implementation on a shared-memory vector supercomputer. Parallel Computing, 22(11):1513–1526, January 26, 1997. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=22&
                                        issue=11&aid=1113.

Gropp:1999:RMM

[GL99] W. Gropp and E. Lusk. Reproducible measurements of MPI performance characteristics. In Dongarra et al. [DLM99], pages 11– 18. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Gropp:2002:MG

[GL02] William Gropp and Ewing Lusk. MPI on the Grid. Lecture Notes in Computer Science, 2474:12–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/
service/series/0558/bibs/
2474/24740012.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740012.pdf.

Gropp:2004:FTM

[GL04] William Gropp and Ewing Lusk. Fault tolerance in Message Passing Interface programs. The International Journal of High Performance Computing Applications, 18(3):363–372, Fall 2004. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/18/
3/363.full.pdf+html.

Girona:2000:VDC

[GLB00] Sergi Girona, Jesús Labarta, and Rosa M. Badia. Validation of dimemas communication model for MPI collective operations. Lecture Notes in Computer Science, 1908:39–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080039.htm;
 http://link.springer-
ny.com/link/service/series/

                                        0558/papers/1908/19080039.
                                        pdf.

Gropp:1996:HPP

[GLDS96] William Gropp, Ewing Lusk, Nathan Doss, and Anthony Skjellum. High-performance, portable implementation of the MPI Message Passing Interface Standard. Parallel Computing, 22(6):789– 828, September 20, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=22&
                                        issue=6&aid=1075.

Glendinning:1993:MMP

[Gle93] I. Glendinning. 93SC041 the MPI message passing interface. In Anonymous [Ano93a], pages 229–236. ISBN 0-947719-62-8. LCCN ????

Gregoretti:2008:MGE

[GLM^{+}08] F. Gregoretti, G. Laccetti, A. Murli, G. Oliva, and U. Scafuri. MGF: a gridenabled MPI library. Future Generation Computer Systems, 24(2):158–165, February 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Garland:2008:PCE

[GLN^{+}08] Michael Garland, Scott Le Grand, John Nickolls, Joshua Anderson, Jim Hardwick, Scott Morton, Everett Phillips, Yao Zhang, and Vasily Volkov. Parallel computing experiences with CUDA. IEEE Micro, 28(4): 13–27, July/August 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gonzalez:2000:TSN

[GLP^{+}00] J. A. González, C. León, F. Piccoli, M. Printista, J. L. Roda, C. Rodríguez, and F. Sande. Towards standard nested parallelism. Lecture Notes in Computer Science, 1908:96–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080096.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080096.
pdf.

Gonzalez:2001:MIM

[GLRS01] J. A. González, C. León, C. Rodríguez, and F. Sande. A model to integrate message passing and shared memory programming. Lecture Notes in Computer Science, 2131:114–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/

                                        bibs/2131/21310114.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2131/21310114.
                                        pdf.

Gropp:1994:UMP

[GLS94] William Gropp, Ewing Lusk, and Anthony Skjellum. Using MPI: Portable Parallel Programming with the Message-Passing Interface. Scientific and engineering computation. MIT Press, Cambridge, MA, USA, 1994. ISBN 0-262-57104-8. xx + 307 pp. LCCN QA76.642 G76 1994. US$24.95. URL http:/

                                        /www.mitpress.com/book-
                                        home.tcl?isbn=0262571048.

Gropp:1999:UMP

[GLS99] William Gropp, Ewing Lusk, and Anthony Skjellum. Using MPI: Portable Parallel Programming with the Message Passing Interface. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, second edition, November 1999. ISBN 0-262-57132-3 (vol. 1), 0-262-57134-X (set). 350 pp. LCCN QA76.642.G76 1999. US$32.50. URL http:/

                                        /www.mitpress.com/book-
                                        home.tcl?isbn=0262571323.

Gropp:1999:UMA

[GLT99] William Gropp, Ewing Lusk, and Rajeev Thakur. Using MPI-2: Advanced Features of the Message Passing Interface. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, November 1999. ISBN 0-262-57133-1. 275 pp. LCCN QA76.642 .G762 1999. US$32.50. URL http:

//www.mitpress.com/book-
home.tcl?isbn=0262571331.

Gropp:2000:UMA

[GLT00a] William Gropp, Ewing Lusk, and Rajeev Thakur. Using MPI-2: Advanced Features of the Message Passing Interface. Scientific and engineering computation. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-57133-1. xxi + 382 pp. LCCN QA76.642 .G762 1999.

Gropp:2000:TSU

[GLT00b] William Gropp, Ewing (Rusty)

Lusk, and Rajeev S. Thakur. Tutorial S1: Using MPI-2: a tutorial on advanced features of the message-passing interface. In ACM [ACM00],

page 11. URL http://www.
sc2000.org/proceedings/
info/fp.pdf.

Gropp:2012:AMI

[GLT12] William Gropp, Ewing Lusk, and Rajeev Thakur. Advanced MPI including new MPI-3 features. Lecture Notes in Computer Science, 7490:14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.

                                        springer.com/accesspage/
                                        chapter/10.1007/978-3-
                                        642-33518-1_5.

Gajecki:1994:NAT

[GM94] M. Gajecki and J. Moscinski. A new algorithm for the traveling salesman problem on networked workstations. In Dongarra and Wasniewski [DW94], pages 229–235. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1994. DM104.00.

Gianuzzi:1995:UPI

[GM95] V. Gianuzzi and F. Merani. Using PVM to implement a distributed dependable simulation system. In IEEE [IEE95h], pages 529–535. ISBN 0-8186-7031-2, 0-8186-7032-0. LCCN QA76.58 .E97 1995.

Goglin:2013:KGS

[GM13] Brice Goglin and Stéphanie Moreaud. KNEM: a generic and scalable kernel-assisted intra-node MPI communication framework. Journal of Parallel and Distributed Computing, 73(2): 176–188, February 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0743731512002316.

Gupta:2018:ALQ

[GM18] Sourendu Gupta and Pushan Majumdar. Accelerating lattice QCD simulations with 2 flavors of staggered fermions on multiple GPUs using OpenACC — a first attempt. Computer Physics Communications, 228(??): 44–53, July 2018. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465518300808.

Ghazimirsaeed:2020:CAM

[GMA20] S. Mahdieh Ghazimirsaeed, Seyed H. Mirsadeghi, and Ahmad Afsahi. Communicationaware message matching in MPI. Concurrency and Computation: Practice and Experience, 32(3):e4862:1– e4862:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gu:2007:IPC

[GMdMBD^{+}07] Feng Long Gu, Hyacinthe Nzigou M., Guilherme de Melo Baptista Domingues, Takeshi Nanri, and Kazuaki Murakami. Investigating the performance of collective communications on SMP clusters: a case for

MPI Allgather. In Simos

and Maroulis [SM07], pages 52–56. ISBN 0-7354-0476-3 (set), 0-7354-0477-1 (vol. 1), 0-7354-0478-X (vol. 2). ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. LCCN Q183.9 .I524 2007.

                                        URL http://proceedings.
                                        aip.org/getpdf/servlet/
                                        GetPDFServlet?filetype=
                                        pdf&amp;   id=APCPCS000963000002000052000001&
                                        amp; idtype=cvips.

Gong:2016:NPG

[GML^{+}16] Jing Gong, Stefano Markidis, Erwin Laure, Matthew Otten, Paul Fischer, and Misun Min. Nekbone performance on GPUs with OpenACC and CUDA Fortran implementations. The Journal of Supercomputing, 72(11): 4160–4180, November 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Goujon:1998:AAT

[GMPD98] D. S. Goujon, M. Michel, J. Peeters, and J. E. Devaney. AutoMap and AutoLink: Tools for communicating complex and dynamic data-structures using MPI. Lecture Notes in Computer Science, 1362:98–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guan:1995:SCC

[GMU95] Xiaojun Guan, Richard J. Mural, and Edward C. Uberbacher. Sequence comparison on a cluster of workstations using the PVM system. In IEEE [IEE95f], pages 190–195. CODEN PSPDF8. ISBN 0-8186-7074-6. ISSN 1063-6374. LCCN QA 76.58 I56 1995. IEEE catalog no. 95TH8052.

Gray:1995:PCT

[GN95] J. P. Gray and F. Naghdy, editors. Parallel Computing: Technology and Practice. PCAT-94. Proceedings of the 7th Australian Transputer and Occam User Group Conference: Woollongong, NSW, Australia, 8– 9 November 1994. IOS Press, Postal Drawer 10558, Burke, VA 2209-0558, USA, 1995. ISBN ???? LCCN ????

Gloster:2019:CBP

[GNP19] Andrew Gloster, Lennon Ó Náraigh, and Khang Ee Pang. cuPentBatch — a batched pentadiagonal solver for NVIDIA GPUs. Computer Physics Communications, 241(??):113– 121, August 2019. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465519300979.

Gloster:2019:CCF

[GÓ19] Andrew Gloster and Lennon Ó Náraigh. cuSten — CUDA finite difference and stencil library. SoftwareX, 10 (??):Article 100337, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2352711019300561.

Goedecker:2002:OPF

[Goe02] Stefan Goedecker. Optimization and parallelization of a force field for silicon using OpenMP. Computer Physics Communications, 148(1): 124–135, October 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465502004666.

Gonzalez:2001:OET

[GOM^{+}01] Marc Gonzalez, Jose Oliver, Xavier Martorell, Eduard Ayguade, Jesus Labarta, and Nacho Navarro. OpenMP extensions for thread groups and their run-time support. Lecture Notes in Computer Science, 2017:324–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2017/20170324.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2017/20170324.
                                        pdf.

Gorzig:2001:CCP

[Gör01] Steffen Görzig. CPPvm — C++ and PVM. Lecture Notes in Computer Science, 2131:83–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310083.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310083.
pdf.

Guarracino:1995:PMB

[GP95] M. R. Guarracino and F. Perla. A parallel modified block Lanczos algorithm for distributed memory architectures. In IEEE [IEE95h], pages 424–431. ISBN 0-8186-7031-2, 0-8186-7032-0. LCCN QA76.58 .E97 1995.

Grosset:2017:TTT

[GPC^{+}17] A. V. Pascal Grosset, Manasa Prasad, Cameron Christensen, Aaron Knoll, and Charles Hansen. TODtree: Task-overlapped direct send tree image compositing for hybrid MPI parallelism and GPUs. IEEE Transactions on Visualization and Computer Graphics, 23(6):1677–1690, June 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-

9306.  URL https://
www.computer.org/csdl/
trans/tg/2017/06/07433468-
abs.html.

Govindan:1996:OMP

[GPL^{+}96] V. Govindan, Y. Park, X. Li, S. Crear, and O. Johnson.

An overview of a MPI profiling environment for the NEC Cenju-3. In IEEE [IEE96i], pages 185–188. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Gillich:1995:FPP

[GR95] S. Gillich and B. Ries. Flexible, portable performance analysis for PARMACS and MPI. In Hertzberger and Serazzi [HS95a], pages 937–?? ISBN 3-540-59393-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Genaud:2007:PMP

[GR07] Stéphane Genaud and Choopan Rattanapoka. P2P–MPI: a peer-to-peer framework for robust execution of message passing parallel programs on Grids. Journal of Grid Computing, 5(1):27–42, March 2007. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=1570-7873&volume=
5&issue=1&spage=27.

Grabowsky:1997:MBK

[Gra97] Lothar Grabowsky. MPIbasierte Koppelrandkommunikation und Einflüß der Partitionierung im 3D-Fall. (German) [MPI-based coupled edge communication and influence of partitioning in 3D-Fall].

Preprint-Reihe des Chemnitzer SFB 393 97,17, Universität Chemnitz-Zwickau, Chemnitz, Germany, 1997. 13 pp.

Gravvanis:2009:OBP

[Gra09] George A. Gravvanis. OpenMP based parallel normalized direct methods for sparse finite element linear systems. The Journal of Supercomputing, 47(1):44–52, January 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
47&issue=1&spage=44.

Grengbondai:1994:CPU

[Gre94] Jules Crephat Grengbondai.

Concurrent processing under parallel virtual machine (PVM). M.s. thesis, Department of Computer Science, Southern Illinois University at Carbondale, Carbondale, IL, USA, 1994. vi + 97 pp.

Greenfield:1995:OPS

[Gre95] J. Greenfield. An overview of the PVM software system. In IEEE [IEE95d], pages 17– 23. ISBN ???? LCCN ????

Gropp:2000:RCD

[Gro00] William D. Gropp. Runtime checking of datatype signatures in MPI. Lecture Notes in Computer Science, 1908:160–??, 2000.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080160.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080160.
pdf.

Gropp:2001:CSA

[Gro01a] William D. Gropp. Challenges and successes in achieving the potential of MPI. Lecture Notes in Computer Science, 2131: 7–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310007.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310007.
pdf.

Gropp:2001:LSM

[Gro01b] William D. Gropp. Learning from the success of MPI. Lecture Notes in Computer Science, 2228:81–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2228/22280081.htm;
 http://link.springer-
ny.com/link/service/series/

0558/papers/2228/22280081.
pdf.

Gropp:2002:BLC

[Gro02a]  William Gropp.  Build-

ing library components that can use any MPI implementation. Lecture Notes in Computer Science, 2474: 280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.de/link/
service/series/0558/bibs/
2474/24740280.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740280.pdf.

Gropp:2002:MNS

[Gro02b] William Gropp. MPICH2:

a new start for MPI implementations. Lecture Notes in Computer Science, 2474:7–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/
service/series/0558/bibs/
2474/24740007.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740007.pdf.

Gropp:2012:MBW

[Gro12] William Gropp. MPI 3 and beyond: Why MPI is successful and what challenges it faces. Lecture Notes in Computer Science, 7490:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-33518-1_
1/.

Gropp:2019:UNS

[Gro19] William D. Gropp. Using node and socket information to implement MPI Cartesian topologies. Parallel Computing, 85(??): 98–108, July 2019. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819118303156.

Gonzalez:1999:PPM

[GRRM99] J. A. Gonzalez, C. Rodriguez, J. L. Roda, and D. G. Morales. Performance and predictability of MPI and BSP programs on the CRAY T3E. In Dongarra et al. [DLM99], pages 27–34. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Gutierrez:2010:QCS

[GRTZ10] Eladio Gutiérrez, Sergio Romero, María A. Trenas, and Emilio L. Zapata. Quantum computer simulation using the CUDA programming model. Computer Physics Communications, 181(2): 283–300, February 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465509003117.

Gaito:2001:ADC

[GRV01] A. Gaito, M. Rak, and U. Villano. Adding dynamic coscheduling support to PVM. Lecture Notes in Computer Science, 2131:106–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310106.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310106.
pdf.

Gittens:2019:AAS

[GRW^{+}19] Alex Gittens, Kai Rothauge, Shusen Wang, Michael W. Mahoney, Jey Kottalam, Lisa Gerhardt, Prabhat, Michael Ringenburg, and Kristyn Maschhoff. Alchemist: an Apache Spark ↔ MPI interface. Concurrency and Computation: Practice and Experience, 31 (16):e5026:1–e5026:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Geist:1991:ENB

[GS91a] G. A. Geist and V. S.

Sunderam. Experiences with network based concurrent computing on the PVM system. Technical Report ORNL/TM-11760, Oak Ridge National Laboratory, Knoxville, TN, USA, January 1991.

Geist:1991:PSS

[GS91b] G. A. Geist and V. S. Sunderam. The PVM system: Supercomputer level concurrent computation on a heterogeneous network of workstations. In Stout and Wolfe [SW91], pages 258– 261. ISBN 0-8186-2291-1. LCCN QA76.5 .D58 1991.

Geist:1992:NBC

[GS92] G. A. Geist and V. S. Sunderam. Network-based concurrent computing on the PVM system. Concurrency: practice and experience, 4(4):293–312 (or 293– 311??), June 1992. CODEN CPEXEI. ISSN 1040-3108.

Geist:1993:EPC

[GS93] G. A. Geist and V. S. Sunderam. The evolution of the PVM concurrent computing system. In IEEE [IEE93a], pages 549–557. ISBN 0-8186-3400-6. LCCN QA75.5.C58 1993. IEEE catalog no. 93CH3251-6.

Gropp:1994:SEP

[GS94] W. Gropp and B. Smith. Scalable, extensible, and portable numerical libraries.

In IEEE [IEE94f], pages 87–

93. ISBN 0-8186-4980-1. LCCN QA76.58.S34 1993.

Gold:1996:UAL

[GS96] C. Gold and T. Schnekenburger. Using the ALDY load distribution system for PVM applications. In Bode et al. [BDLS96], pages 278–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Geist:19xx:NBC

[GSxx] G. A. Geist and V. S. Sunderam. Network based concurrent computing on the PVM system. Technical report, Oak Ridge National Laboratory and Emory University, Knoxville, TN, USA and Atlanta, GA, USA, 19xx.

Garg:2002:TOA

[GS02] Rajat P. Garg and Ilya Sharapov. Techniques for optimizing applications: high performance computing. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2002. ISBN 0-13-093476-3. xliii + 616 pp. LCCN QA76.88 .G37

2002. URL http://www.
sun.com/books/catalog/
garg.html/index.html;
http://www.sun.com/solutions/
blueprints/tools/.

Gao:2008:GEI

[GSA08] Guang R. Gao, Mitsuhisa Sato, and Eduard Ayguadé. Guest Editors introduction: Special issue on OpenMP. International Journal of Parallel Programming, 36 (3):287–288, June 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=3&spage=287.

Gardner:2013:CCE

[GScFM13] Mark Gardner, Paul Sathre, Wu chun Feng, and Gabriel Martinez. Characterizing the challenges and evaluating the efficacy of a CUDAto-OpenCL translator. Parallel Computing, 39(12): 769–786, December 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819113001075.

Gine:2002:ALT

[GSHL02] Francesc Giné, Francesc Solsona, Porfidio Hernández, and Emilio Luque. Adjusting the lengths of time slices when scheduling PVM jobs with high memory requirements. Lecture Notes in Computer Science, 2474: 156–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.de/link/
service/series/0558/bibs/
2474/24740156.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740156.pdf.

Gerlach:1997:ECS

[GSI97] J. Gerlach, M. Sato, and Y. Ishikawa. Experiences with the C++ standard template library and MPI for a parallel particle simulation method. Lecture Notes in Computer Science, 1225: 961–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalez:2000:AIT

[GSM^{+}00] M. Gonzalez, A. Serra, X. Martorell, J. Oliver, E. Ayguade, J. Labarta, and N. Navarro. Applying interposition techniques for performance analysis of OpenMP parallel applications. In ????, editor, Proceedings 14th International Parallel and Distributed Processing Symposium. IPDPS 2000, pages 235–240. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000.

Germanas:2017:HUP

[GSMK17] D. Germanas, A. Stepsys, S. Mickevicius, and R. K.

Kalinauskas. HOTB update: Parallel code for calculation of three- and fourparticle harmonic oscillator transformation brackets and their matrices using OpenMP. Computer Physics Communications, 215(??):259–264, June 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465517300401.

Gine:2001:MMM

[GSN^{+}01] Francesc Giné, Francesc Solsona, Xavi Navarro, Porfidio Hernández, and Emilio Luque. MemTo: a memory monitoring tool for a Linux cluster. Lecture Notes in Computer Science, 2131:225–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310225.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310225.
pdf.

Gu:2013:PCI

[GSY^{+}13] Zheng Gu, Matthew Small, Xin Yuan, Aniruddha Marathe, and David K. Lowenthal. Protocol customization for improving MPI performance on RDMA-enabled clusters. International Journal of Parallel Programming, 41(5):682–703, October 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10766-013-0242-0.

Gong:2021:TDG

[GSYT21] Dunwei Gong, Baicai Sun, Xiangjuan Yao, and Tian Tian. Test data generation for path coverage of MPI programs using SAEO. ACM Transactions on Software Engineering and Methodology, 30(2):17:1–17:37, March 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL

https://dl.acm.org/doi/
10.1145/3423132.

Gruber:1994:PJE

[GT94] Ralf Gruber and Marco Tomassini, editors. Proceedings of the 6th Joint EPS-APS International Conference on Physics Computing: Physics Computing ’94, Palazzo dei Congressi, Lugano, Switzerland, 22–26 August 1994. European Physical Society, Geneva, Switzerland, 1994. ISBN 2-88270-011-3. LCCN QC20.7.E4I58 1994.

Golbiewski:2001:MOS

[GT01] Maciej Gołbiewski and Jesper Larsson Träff. MPI-2 one-sided communications on a Giganet SMP cluster. Lecture Notes in Computer Science, 2131:16–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2131/21310016.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2131/21310016.
                                        pdf.

Gropp:2007:TSM

[GT07] William Gropp and Rajeev Thakur. Thread-safety in an MPI implementation: Requirements and analysis. Parallel Computing, 33(9): 595–604, September 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gropp:2019:GEI

[GT19] William Gropp and Rajeev Thakur. Guest Editor’s introduction: Special issue on best papers from EuroMPI/ USA 2017. Parallel Computing, 84(??):62, May 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819119300560.

Gennart:1996:CAG

[GTH96] B. A. Gennart, J. Tarraga Gimenez, and R. D. Hersch. Computer-assisted generation of PVM/C++ programs using CAP. Lecture Notes in Computer Science, 1156:259–269, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gidra:2015:NGC

[GTS^{+}15] Lokesh Gidra, Gaël Thomas, Julien Sopena, Marc Shapiro, and Nhan Nguyen. NumaGiC: a garbage collector for big data on big NUMA machines. ACM SIGARCH Computer Architecture News, 43(1):661– 673, March 2015. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Guang:2016:NMN

[Gua16] Suo Guang. NR-MPI: A non-stop and fault resilient MPI supporting programmer defined data backup and restore for E-scale super computing systems. Supercomputing Frontiers and Innovations, 3(1):4–21, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic). URL http:/

/superfri.org/superfri/
article/view/89.

Gallardo:2018:EMM

[GVF^{+}18] Esthela Gallardo, Jérôme Vienne, Leonardo Fialho, Patricia Teller, and James Browne. Employing MPI T in MPI advisor to optimize application performance.

The International Journal of High Performance Computing Applications, 32(6): 882–896, November 1, 2018. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:

                                        //journals.sagepub.com/
                                        doi/full/10.1177/1094342016684005.

Ge:1995:DHA

[GWC95] Yuzhen Ge, L. T. Watson, and E. G. Collins, Jr. Distributed homotopy algorithms for H^{2}/H^{∞} controller synthesis. In Bailey et al. [BBG^{+}95], pages 84–89. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Guerrero:2014:PCM

[GWVP^{+}14] Ginés D. Guerrero, Richard M. Wallace, José L. Vázquez-Poletti, José M. Cecilia, José M. García, Daniel Mozos, and Horacio Pérez-Sánchez. A performance/ cost model for a CUDA drug discovery application on physical and public cloud infrastructures. Concurrency and Computation: Practice and Experience, 26 (10):1787–1798, July 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hadjidoukas:2010:NOP

[HA10] Panagiotis E. Hadjidoukas and Laurent Amsaleg. Nested OpenMP parallelization of a hierarchical data clustering algorithm. Parallel Processing Letters, 20(2):187– 208, June 2010. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Han:2011:HHL

[HA11] Tianyi David Han and Tarek S. Abdelrahman. hiCUDA: High-level GPGPU programming. IEEE Transactions on Parallel and Distributed Systems, 22(1):78– 90, January 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hussain:2011:PIA

[HAA^{+}11] Masroor Hussain, Muhammad Abid, Mushtaq Ahmad, Ashfaq Khokhar, and Arif Masud. A parallel implementation of ALE moving mesh technique for FSI problems using OpenMP. International Journal of Parallel Programming, 39(6): 717–745, December 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=6&spage=717.

Hoeflinger:2001:PSP

[HAJK01] Jay Hoeflinger, Prasad Alavilli, Thomas Jackson, and Bob Kuhn. Producing scalable performance with OpenMP: Experiments with two CFD applications. Parallel Computing, 27(4):391– 413, March 2001. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/35/21/47/28/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/28/26/article.
pdf.

Hamza:1995:PII

[Ham95a] M. H. Hamza, editor. Proceedings of the IASTED International Conference. Modelling and Simulation: Pittsburgh, PA, USA, 27–29 April 1995. IASTEC-Acta Press, Anaheim, CA, USA, 1995. ISBN 0-88986-218-

4. LCCN QA76.9.C65 I295 1995.

Haridi:1995:EPP

[HAM95b] Seif Haridi, Khayri Ali, and Peter Magnusson, editors. EURO-PAR ’95 parallel processing: First International EURO PAR Conference, Stockholm, Sweden, August 29–31, 1995: proceedings, number 966 in Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-60247-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I553 1995.

Hansen:1998:EMP

[Han98] Per Brinch Hansen. An evaluation of the Message-Passing Interface. ACM SIGPLAN Notices, 33(3): 65–72, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). The author criticizes MPI, and remarks “MPI … lack[s] the elegance and security that can only by checked by a parallel programming language.”.

Hardwick:1994:PVL

[Har94] Jonathan C. Hardwick.

Porting a vector library: a comparison of MPI, Paris, CMMD and PVM (or, “I’ll never have to port CVL again”). Research paper CMU-CS-94-200, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, 1994. 16 pp.

Hardwick:1995:PVL

[Har95] J. C. Hardwick. Porting a vector library: a comparison of MPI, Paris, CMMD and PVM. In IEEE [IEE95j], pages 68–77. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Hassanzadeh:1995:MMG

[Has95] Siamak Hassanzadeh, editor. Mathematical methods in geophysical imaging III: 12–13 July 1995, San Diego, California, volume 2571 of Proceedings of the SPIE — The International Society for Optical Engineering. Society of Photooptical Instrumentation Engineers (SPIE), Bellingham, WA, USA, 1995. CODEN PSISDG. ISBN 0-8194-1930-3. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.2571.

Hisley:2000:PPE

[HASnP00] Dixie Hisley, Gagan Agrawal, Punyam Satya-narayana, and Lori Pollock. Porting and performance evaluation of irregular codes using OpenMP. Concurrency: practice and experience, 12(12):1241–1259, October 2000. CODEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76500349/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500349&PLACEBO=IE.
pdf.

Hatazaki:1998:RRS

[Hat98]  T. Hatazaki.  Rank re-

ordering strategy for MPI topology creation functions. Lecture Notes in Computer Science, 1497:188–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hachler:1996:IAC

[HB96a] G. Hachler and H. Burkhart.

Implementing the ALWAN communication and data distribution library using PVM. In Bode et al. [BDLS96], pages 243–250. ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Haechler:1996:IAC

[HB96b] G. Haechler and H. Burkhart.

Implementing the ALWAN communication and data distribution library using PVM. Lecture Notes in Computer Science, 1156:243–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hausner:1995:EIP

[HBT95] M. Hausner, M. Burrows, and C. A. Thekkath. Efficient implementation of PVM on the AN2 ATM network. In Hertzberger and Serazzi [HS95a], pages 562–569. ISBN 3-540-59393-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Huang:2006:ECS

[HC06] Jih-Woei Huang and Chih-Ping Chu. An efficient communication scheduling method for the processor mapping technique applied data redistribution. The Journal of Supercomputing, 37(3):297–318, September 2006. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
37&issue=3&spage=297.

Huang:2008:FPM

[HC08] Jih-Woei Huang and Chih-Ping Chu. A flexible processor mapping technique toward data localization for block-cyclic data redistribution. The Journal of Supercomputing, 45(2):151– 172, August 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
45&issue=2&spage=151.

Hamid:2010:CMB

[HC10] Nor Asilah Wati Abdul Hamid and Paul Coddington. Comparison of MPI benchmark programs on shared memory and distributed memory machines (point-to-point communication). The International Journal of High Performance Computing Applications, 24(4):469–483, November 2010. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/24/
4/469.full.pdf+html.

Han:2017:SLS

[HC17] Yiming Han and Anthony T.

Chronopoulos. Scalable loop self-scheduling schemes for large-scale clusters and cloud systems. International Journal of Parallel Programming, 45(3):595–611, June 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Hunold:2016:RMB

[HCA16] Sascha Hunold and Alexandra Carpen-Amarie. Reproducible MPI benchmarking is still not as easy as you think. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3617–3630, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07426807-
abs.html.

Hashmi:2020:FXZ

[HCC^{+}20] Jahanzeb Maqbool Hashmi, Ching-Hsiang Chu, Sourav Chakraborty, Mohammadreza Bayatpour, Hari Subramoni, and Dhabaleswar K. Panda. FALCON-X: Zero-copy MPI derived datatype processing on modern CPU and GPU architectures. Journal of Parallel and Distributed Computing, 144(??):

1–13, October 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0743731520302872.

Hurwitz:2005:AMP

[HcF05] Justin (Gus) Hurwitz and Wu chun Feng. Analyzing MPI performance over 10-gigabit Ethernet. Journal of Parallel and Distributed Computing, 65(10): 1253–1260, October 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Huang:2005:TME

[HCL05] Lei Huang, Barbara Chapman, and Zhenying Liu. Towards a more efficient implementation of OpenMP for clusters via translation to global arrays. Parallel Computing, 31(10–12):1114– 1139, October/December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hu:2016:CLG

[HCZ16] Liang Hu, Xilong Che, and Si-Qing Zheng. A closer look at GPGPU. ACM Computing Surveys, 48(4): 60:1–60:??, May 2016. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

He:2000:UAA

[HD00a] Yun He and Chris H. Q.

Ding. Using accurate arithmetics to improve numerical reproducibility and stability in parallel applications. In Reynders and Veidenbaum [RV00], pages 225– 234. ISBN 1-58113-270-0. LCCN QA76.88 .I573 2000.

URL https://dl.acm.org/
doi/abs/10.1145/335231.
335253.

He:2000:PAA

[HD00b] Yun (Helen) He and Chris H. Q. Ding. Platforms: An accurate arithmetics approach. In ACM [ACM00],

page 150. URL http://www.
sc2000.org/proceedings/
info/fp.pdf.

Ding:2002:MOP

[HD02a] Yun He and Chris H. Q.

Ding. MPI and OpenMP paradigms on cluster of SMP architectures. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN

???? URL http://www.sc-
2002.org/paperpdfs/pap.
pap325.pdf.

He:2002:MOP

[HD02b] Yun He and Chris H. Q.

Ding. MPI and OpenMP paradigms on cluster of SMP architectures: The vacancy tracking algorithm for multidimensional array transposition. Parallel and Distributed Computing Practices, 5(2):117–128, June 2002. CODEN ???? ISSN 1097-2803.

Harvey:2011:STP

[HD11] M. J. Harvey and G. De Fabritiis. Swan: a tool for porting CUDA programs to OpenCL. Computer Physics Communications, 182(4): 1093–1099, April 2011. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465511000117.

Hoefler:2012:LMO

[HDB^{+}12] Torsten Hoefler, James Dinan, Darius Buntinas, Pavan Balaji, and Brian W. Barrett. Leveraging MPI’s onesided communication interface for shared-memory programming. Lecture Notes in Computer Science, 7490: 132–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-33518-1_
18/.

Hoefler:2013:MMN

[HDB^{+}13] Torsten Hoefler, James Dinan, Darius Buntinas, Pavan Balaji, and Brian Barrett …. MPI + MPI: a new hybrid approach to parallel programming with MPI plus shared memory. Computing, 95(12):1121–1136, December 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00607-013-0324-2.

Hadjidoukas:2009:HPF

[HDDG09] P. E. Hadjidoukas, V. V.

Dimakopoulos, M. Delakis, and C. Garcia. A highperformance face detection system using OpenMP. Concurrency and Computation: Practice and Experience, 21(15):1819–1837, October 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hoefler:2015:RMA

[HDT^{+}15] Torsten Hoefler, James Dinan, Rajeev Thakur, Brian Barrett, Pavan Balaji, William Gropp, and Keith Underwood. Remote memory access programming in MPI-3. ACM Transactions on Parallel Computing (TOPC), 2(2):9:1–9:??, July 2015. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Ho:2021:GFD

[HDW21] Nhut-Minh Ho, Himeshi De Silva, and Weng-Fai Wong. GRAM: a framework for dynamically mixing precisions in GPU applications. ACM Transactions on Architecture and Code Optimization, 18(2):19:1–19:24, March 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3441830.

He:2020:SMO

[HDZ^{+}20] Feng He, Xiaoshe Dong, Nianjun Zou, Weiguo Wu, and Xingjun Zhang. Structured mesh-oriented framework design and optimization for a coarse-grained parallel CFD solver based on hybrid MPI/OpenMP programming. The Journal of Supercomputing, 76 (4):2815–2841, April 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Heikonen:2002:ILB

[HE02] Jussi Heikonen and Kalle Eerola. Improving load balance in a weather code: Asynchronous output in HIRLAM with MPI. Lecture Notes in Computer Science, 2367:567–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2367/23670567.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2367/23670567.
                                        pdf.

Hadi:2013:CFA

[HE13] Mohammed F. Hadi and Seyed A. Esmaeili. CUDA Fortran acceleration for the finite-difference time-domain method. Computer Physics Communications, 184(5): 1395–1400, May 2013. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465513000118.

Havran:2015:EBT

[HE15] Vlastimil Havran and Petr Egert. Extensions to bidirectional texture function compression with multilevel vector quantization in OpenCL. Computers and Graphics, 48(??): 1–10, May 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000060.

Hebeker:1993:CPC

[Heb93] F.-K. Hebeker. On a coarsegrained parallel code to simulate reactive flows on an IBM RS/ 6000 workstationcluster. In Brebbia and Power [BP93], pages 253–

262. ISBN 1-85312-236-X. LCCN TA345.I556 1993.

Herland:1998:CML

[HEH98] B. G. Herland, M. Eberl, and H. Hellwagner. A common messaging layer for MPI and PVM over SCI. Lecture Notes in Computer Science, 1401:576–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:2009:EGO

[HEHC09] Lei Huang, Deepak Eachempati, Marcus W. Hervey, and Barbara Chapman. Exploiting global optimizations for OpenMP programs in the OpenUH compiler. ACM SIGPLAN Notices, 44(4):289–290, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hempel:1994:MSM

[Hem94]  R. Hempel.  The MPI

Standard for Message Passing. In Gentzsch and Harms [GH94], pages 247– 252. ISBN 0-387-57981-8 (New York), 3-540-57981-8 (Berlin). LCCN QA76.88.I57 1994. DM96.00. Two volumes.

Hempel:1996:SMM

[Hem96] R. Hempel. The status of the MPI message-passing standard and its relation to PVM. In Bode et al. [BDLS96], pages 14–21. ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Holmen:2014:ASI

[HF14a] John K. Holmen and David L. Foster. Accelerating single iteration performance of CUDA–based 3D reaction– diffusion simulations. International Journal of Parallel Programming, 42(2):343– 363, April 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10766-013-0251-z.

See erratum [HF14b].

Holmen:2014:EAS

[HF14b] John K. Holmen and David L.

Foster. Erratum to: Accelerating single iteration performance of CUDA–based 3D reaction–diffusion simulations. International Journal of Parallel Programming, 42 (2):364, April 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.
springer.com/content/pdf/
10.1007/s10766-014-0305-

x.pdf. See [HF14a].

Hahne:2021:APP

[HFB21] Jens Hahne, Stephanie Friedhoff, and Matthias Bolten. Algorithm 1016: PyMGRIT: a Python package for the parallel-in-time method MGRIT. ACM Transactions on Mathematical Software, 47(2):19:1– 19:22, April 2021. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3446979.

Hursey:2012:AFA

[HG12] Joshua Hursey and Richard L. Graham. Analyzing fault aware collective performance in a process fault tolerant MPI. Parallel Computing, 38(1–2):15–25, January/February 2012. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819111001414.

Hermanns:2012:SDM

[HGMW12] Marc-André Hermanns, Markus Geimer, Bernd Mohr, and Felix Wolf. Scalable detection of MPI-2 remote memory access inefficiency patterns. The International Journal of High Performance Computing Applications, 26(3):227–236, August 2012. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

                                        tronic). URL http://hpc.
                                        sagepub.com/content/26/
                                        3/227.full.pdf+html.

Haghi:2022:RSH

[HGX^{+}22] Pouya Haghi, Anqi Guo, Qingqing Xiong, Chen Yang, Tong Geng, Justin T. Broaddus, Ryan Marshall, Derek Schafer, Anthony Skjellum, and Martin C. Herbordt. Reconfigurable switches for high performance and flexible MPI collectives. Concurrency and Computation: Practice and Experience, 34(6):e6769:1– e6769:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hong:1995:PNP

[HH95] Lin Hong and Chen Huaping. PVM and network parallel computing. Mini-Micro Systems, 16(2):53–58, February 1995. CODEN XWJXEH. ISSN 1000-1220.

Hanson:2014:NCM

[HH14] Richard J. Hanson and Tim Hopkins. Numerical computing with modern Fortran. Applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2014. ISBN 1-61197-311-2 (paperback), 1-61197-312-0 (e-book). xv + 244 pp. LCCN QA76.73.F25 H367 2013.

Huckelheim:2022:SSA

[HH22] Jan Hückelheim and Laurent Hascoët. Source-to-source automatic differentiation of OpenMP parallel loops. ACM Transactions on Mathematical Software, 48(1):7:1–7:32, March 2022. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

dl.acm.org/doi/10.1145/
3472796.

Hui:1995:SPS

[HHA95] Chi-Chung Hui, Mounir Hamdi, and Ishfaq Ahmad. Software platform for solving PDEs on distributed systems: Implementation issues and performance prediction. In IEEE [IEE95l], pages 383–388. CODEN PSICD2. ISBN 0-8186-7119-X. ISSN 0730-6512. LCCN QA 76.6 C6295 1995. IEEE catalog number 95CB35838.

Huang:2018:ACO

[HHC^{+}18] Kai Huang, Biao Hu, Long Chen, Alois Knoll, and Zhihua Wang. Adas on Cots with OpenCL: A case study with lane detection. IEEE Transactions on Computers, 67(4):559–565, ???? 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

                                        //ieeexplore.ieee.org/
                                        document/8057795/.

Horiguchi:1994:ISP

[HHK94] S. Horiguchi, D. Frank Hsu, and M. Kimura, editors. International Symposium on Parallel Architectures, Algorithms, and Networks (ISPAN): proceedings of the 1994, December 14–16, 1994, Kanazawa, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6507-6 (case), 0-8186-6506-8 (microfiche). LCCN QA76.58 .I5673 1994 Bar. IEEE catalog number 94TH0697-3.

Hermanns:2019:MEI

[HHK^{+}19] Marc-André Hermanns, Nathan T.

Hjelm, Michael Knobloch, Kathryn Mohror, and Martin Schulz. The MPI T events interface: an early evaluation and overview of the interface. Parallel Computing, 85(??):119– 130, July 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167819118303314.

Halver:2018:FPM

[HHS18] Rene Halver, Wilhelm Homberg, and Godehard Sutmann. Function portability of molecular dynamics on heterogeneous parallel architectures with OpenCL. The Journal of Supercomputing, 74(4):1522–1533, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Huckelheim:2019:RMA

[HHSM19] Jan Hückelheim, Paul Hovland, Michelle Mills Strout, and Jens-Dominik Müller. Reverse-mode algorithmic differentiation of an OpenMPparallel compressible flow solver. The International Journal of High Performance Computing Applications, 33(1):140–154, January 1, 2019. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:/

                                        /journals.sagepub.com/
                                        doi/full/10.1177/1094342017712060.

Haque:2016:ACV

[HIIG16] Syed Arefinul Haque, Salekul Islam, Md. Jahidul Islam, and Jean-Charles Grégoire. An architecture for client virtualization: a case study. Computer Networks (Amsterdam, Netherlands: 1999), 100(??):75–89, May 8, 2016. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1389128616300421.

Hinde:2011:QMD

[Hin11] Robert J. Hinde. QSATS: MPI-driven quantum simulations of atomic solids at zero temperature. Computer Physics Communications, 182(11):2339–2349, November 2011. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0010465511001615.

Huttunen:2002:MCC

[HIP02] Pentti Huttunen, Jouni Ikonen, and Jari Porras. MPIT — communication/ computation paradigm for networks of SMP workstations. Lecture Notes in Computer Science, 2367: 160–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2367/23670160.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2367/23670160.
pdf.

Haimes:1998:UPM

[HJ98] R. Haimes and K. E. Jordan. Using PVM and MPI for co-processed, distributed and parallel scientific visualization. Lecture Notes in Computer Science, 1388:1098–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hori:2021:ISM

[HJB^{+}21] Atsushi Hori, Emmanuel Jeannot, George Bosilca, Takahiro Ogura, Balazs Gerofi, Jie Yin, and Yutaka Ishikawa. An international survey on MPI users. Parallel Computing, 108(??):??, December 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819121000983.

Hall:2014:MMC

[HJBB14] Clifford Hall, Weixiao Ji, and Estela Blaisten-Barojas. The Metropolis Monte Carlo method with CUDA enabled Graphic Processing Units. Journal of Computational Physics, 258(??): 871–879, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0021999113007626.

Huang:2010:ELA

[HJYC10] Lei Huang, Haoqiang Jin, Liqi Yi, and Barbara Chapman. Enabling localityaware computations in OpenMP. Scientific Programming, 18 (3–4):169–181, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hoffmann:1993:PFE

[HK93] Geerd-R. Hoffmann and Tuomo Kauranne, editors. Proceedings of the Fifth ECMWF Workshop on the Use of Parallel Processors in Meteorology. Parallel Supercomputing in Atmospheric Science. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1993. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Henriksen:1994:PCF

[HK94] P. Henriksen and R. Keunings. Parallel computation of the flow of integral viscoelastic fluids on a heterogeneous network of workstations. International Journal for Numerical Methods in Fluids, 18(12):1167– 1183, June 1994. CODEN IJNFDW. ISSN 0271-2091.

Hoffmann:1995:CAP

[HK95] Geerd-R. Hoffmann and Norbert Kreitz, editors. Coming of age: proceedings of the Sixth ECMWF Workshop on the Use of Parallel Processors in Meteorology, Reading, UK, November 21–25, 1994. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1995. ISBN 981-02-2211-4. LCCN QC866.E26 1994.

Hong:2009:AMG

[HK09] Sunpyo Hong and Hyesoon Kim. An analytical model for a GPU architecture with memory-level and threadlevel parallelism awareness. ACM SIGARCH Computer Architecture News, 37(3): 152–163, June 2009. CO-DEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Hong:2010:IGP

[HK10] Sunpyo Hong and Hyesoon Kim. An integrated GPU power and performance model. ACM SIGARCH Computer Architecture News, 38(3):280– 289, June 2010. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Hiranandani:1994:CTB

[HKMCS94] S. Hiranandani, K. Kennedy, J. Mellor-Crummey, and A. Sethi. Compilation techniques for block-cyclic distributions. In ACM [ACM94], pages 392–403. ISBN 0-89791-665-4. LCCN

                                        ????  URL http://www.
                                        acm.org/pubs/contents/
                                        proceedings/supercomputing/
                                        181181/.

Hoeflinger:2001:IPV

[HKN^{+}01] Jay Hoeflinger, Bob Kuhn, Wolfgang Nagel, Paul Petersen, Hrabri Rajic, Sanjiv Shah, Jeff Vetter, Michael Voss, and Renee Woo. An integrated performance visualizer for MPI/OpenMP programs. Lecture Notes in Computer Science, 2104: 40–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2104/21040040.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2104/21040040.
                                        pdf.

Hong:2011:ACG

[HKOO11] Sungpack Hong, Sang Kyun Kim, Tayo Oguntebi, and Kunle Olukotun. Accelerating CUDA graph algorithms at maximum warp. ACM SIGPLAN Notices, 46 (8):267–276, August 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’11 Conference proceedings.

Hori:2012:EKL

[HKT^{+}12] Atsushi Hori, Toyohisa Kameyama, Yuichi Tsujita, Mitaro Namiki, and Yutaka Ishikawa. An efficient kernellevel blocking MPI implementation. Lecture Notes in Computer Science, 7490: 153–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-33518-1_
20/.

Hasanov:2017:HRC

[HL17] Khalid Hasanov and Alexey Lastovetsky. Hierarchical redesign of classic MPI reduction algorithms. The Journal of Supercomputing, 73 (2):713–725, February 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hu:2000:ONS

[HLCZ00] Y. Charlie Hu, Honghui Lu, Alan L. Cox, and Willy Zwaenepoel. OpenMP for networks of SMPs. Journal of Parallel and Distributed Computing, 60(12):1512– 1530, December 1, 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/
jpdc.2000.1658;   http:
//www.idealibrary.com/
links/doi/10.1006/jpdc.
2000.1658/pdf;   http:
//www.idealibrary.com/
links/doi/10.1006/jpdc.
2000.1658/ref.

Hagedorn:2020:AHP

[HLK^{+}20] Bastian Hagedorn, Johannes Lenfers, Thomas Kœhler, Xueying Qin, Sergei Gorlatch, and Michel Steuwer. Achieving high-performance the functional way: a functional pearl on expressing high-performance optimizations as rewrite strategies. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):92:1– 92:29, August 2020. URL

https://dl.acm.org/doi/
10.1145/3408974.

Haque:2017:CCL

[HLM^{+}17] S. Anisul Haque, X. Li, F. Mansouri, M. Moreno Maza, D. Mohajerani, and W. Pan. CUMODP: a CUDA library for modular polynomial computation. ACM Communications in Computer Algebra, 51(3):89–91, September 2017. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Hung:2016:EBP

[HLO^{+}16] Che-Lun Hung, Chun-Yuan Lin, Chia-Shin Ou, Yuan-Hong Tseng, Po-Yen Hung, Ship-Peng Li, and Chun-Ting Fu. Efficient bitparallel subcircuit extraction using CUDA. Concurrency and Computation: Practice and Experience, 28(16): 4326–4338, November 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hong:1996:RDM

[HLOC96] Chul-Eui Hong, Bum-Sik Lee, Gi-Won On, and Dong-Hae Chi. Replay for debugging MPI parallel programs. In IEEE [IEE96i], pages 156–160. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Hawick:2010:PGC

[HLP10] K. A. Hawick, A. Leist, and D. P. Playne. Parallel graph component labelling with GPUs and CUDA. Parallel Computing, 36(12): 655–678, December 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hawick:2011:RLS

[HLP11] K. A. Hawick, A. Leist, and D. P. Playne. Regular lattice and small-world spin model simulations using CUDA and GPUs. International Journal of Parallel Programming, 39(2): 183–201, April 2011. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=2&spage=183.

Huband:2001:DTB

[HM01] Simon Huband and Chris McDonald. DEPICT: a topology-based debugger for MPI programs. Lecture Notes in Computer Science, 2026:109–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2026/20260109.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2026/20260109.
pdf.

Hilbrich:2009:MCC

[HMK09] Tobias Hilbrich, Matthias S. Müller, and Bettina Krammer. MPI correctness checking for OpenMP/MPI applications. International Journal of Parallel Programming, 37(3):277–291, June 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=3&spage=277.

Hajihassani:2019:FAI

[HMKG19] O. Hajihassani, S. K. Monfared, S. H. Khasteh, and S. Gorgin. Fast AES implementation: A highthroughput bitsliced approach. IEEE Transactions on Parallel and Distributed Systems, 30(10): 2211–2222, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hakula:1994:FEM

[HMKV94] H. Hakula, J. Malinen, P. Kallberg, and P. Valve. The finite element method applied to the exterior Helmholtz problem on the IBM SP-1. In Dongarra and Wasniewski [DW94], pages 262–269. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1994. DM104.00.

Holmes:2019:PPE

[HMS^{+}19] Daniel J. Holmes, Bradley Morgan, Anthony Skjellum, Purushotham V. Bangalore, and Srinivas Sridharan. Planning for performance: Enhancing achievable performance for MPI through persistent collective operations. Parallel Computing, 81(??):32–57, January 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819118302412.

Haynes:2014:MOA

[HO14] Ronald D. Haynes and Benjamin W. Ong. MPI– OpenMP algorithms for the parallel space-time solution of time dependent PDEs. In Erhel et al. [EGH^{+}14], pages 179–187. ISBN 3-319-05788-X (paperback), 3-319-05789-8 (e-book). ISSN 1439-7358 (print), 2197-7100 (electronic). LCCN QA71-

                                        90. URL http://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-05789-7_
                                        14/.

Hogg:2013:FDT

[Hog13]  J. D. Hogg.  A fast

dense triangular solve in CUDA. SIAM Journal on Scientific Computing, 35(3): C303–C322, ???? 2013. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Hollerbach:1995:FDA

[Hol95] Rainer Hollerbach. Fast dynamo action in spherical geometry: Numerical calculations using parallel virtual machines. Computers in Physics, 9(4):460– ??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/
                                        10.1063/1.168547.

Hollingsworth:2012:SPI

[Hol12] Jeffrey Hollingsworth, editor. SC ’12: Proceedings of the International Conference on High Performance Computing, Networking, Storage and Analysis, Salt Lake Convention Center, Salt Lake City, UT, USA, November 10–16, 2012. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-0804-8.

Hosking:2012:CHL

[Hos12] Tony Hosking. Compiling a high-level language for GPUs: (via language support for architectures and compilers). ACM SIGPLAN Notices, 47(6):1–12, June 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PLDI ’12 proceedings.

Hadjidoukas:2005:OEM

[HP05] P. E. Hadjidoukas and T. S.

Papatheodorou. OpenMP extensions for master-slave message passing computing. Parallel Computing, 31(10–12):1155–1167, October/December 2005. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hawick:2011:HSL

[HP11] K. A. Hawick and D. P.

Playne. Hypercubic storage layout and transforms in arbitrary dimensions using GPUs and CUDA. Concurrency and Computation: Practice and Experience, 23 (10):1027–1050, July 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hidalgo:1999:MMP

[HPLT99] J. I. Hidalgo, M. Prieto, J. Lanchares, and F. Tirado. A method for model parameter identification using parallel genetic algorithms. In Dongarra et al. [DLM99], pages 291–298. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Hadjidoukas:2002:MOI

[HPP02] Panagiotis E. Hadjidoukas, Eleftherios D. Polychronopoulos, and Theodore S. Papatheodorou. A modular OpenMP implementation for clusters of multiprocessors. Parallel and Distributed Computing Practices, 5(2):153–168, June 2002. CODEN ???? ISSN 1097-2803.

Hariri:1995:STE

[HPR^{+}95] S. Hariri, Sung-Yong Park, R. Reddy, M. Subramanyan, R. Yadav, G. C. Fox, and M. Parashar. Software tool evaluation methodology. In IEEE [IEE95i], pages 3–10. ISBN 0-8186-7025-8. LCCN ???? IEEE catalog number 95CH35784.

Hondroudakis:1995:PEV

[HPS95] A. Hondroudakis, R. Procter, and K. Shanmugam. Performance evaluation and visualization with VISPAT. In Malyshkin [Mal95], pages 180–185. ISBN 3-540-60222-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I547 1995.

Heckathorn:1996:SSP

[HPS^{+}96] H. Heckathorn, B. Popp, W. Smith, D. Conklin, D. A. Newman, and F. Wieland. SSGM: from serial to parallel processing using PVM. Proceedings of the SPIE — The International Society for Optical Engineering, 2741:267– 277, ???? 1996. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Hilbrich:2012:MRE

[HPS^{+}12] Tobias Hilbrich, Joachim Protze, Martin Schulz, Bronis R. de Supinski, and Matthias S. Müller. MPI runtime error detection with MUST: advances in deadlock detection. In Hollingsworth [Hol12], pages 30:1–30:?? ISBN 1-4673-0804-8. URL http:

//conferences.computer.
org/sc/2012/papers/1000a010.
pdf.

Hilbrich:2013:MRE

[HPS^{+}13] Tobias Hilbrich, Joachim Protze, Martin Schulz, Bronis R. de Supinski, and Matthias S. Müller. MPI runtime error detection with MUST: Advances in deadlock detection. Scientific Programming, 21(3–4):109– 121, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hariri:1993:MPI

[HPY^{+}93] S. Hariri, J. B. Park, F.- K. Yu, M. Parashar, and G. C. Fox. A message passing interface for parallel and distributed computing. In IEEE [IEE93c], pages 84–91. ISBN 0-8186-3900-8, 0-8186-3901-6. LCCN QA76.9.D5I593 1993. IEEE catalog no. 93TH0550-4.

Hoefler:2011:SPT

[HRR^{+}11] Torsten Hoefler, Rolf Rabenseifner, Hubert Ritzdorf, Bronis R. de Supinski, Rajeev Thakur, and Jesper Larsson Träff. The scalable process topology interface of MPI 2.2. Concurrency and Computation: Practice and Experience, 23 (4):293–310, March 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hoyos-Rivera:1997:UPB

[HRSA97] G. J. Hoyos-Rivera and V. G. Sanchez-Arias. Using PVM to build an interface to support cooperative work in a distributed systems environment. Lecture Notes in Computer Science, 1332:127–134, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hempel:1997:IMN

[HRZ97] R. Hempel, H. Ritzdorf, and F. Zimmermann. Implementation of MPI on NEC’s SX-4 multi-node architecture. Lecture Notes in Computer Science, 1332:185–193, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartley:1993:CPS

[HS93] C. L. Hartley and V. S. Sunderam. Concurrent programming with shared objects in networked environments. In IEEE [IEE93b], pages 471– 478. ISBN 0-8186-3442-1. LCCN QA 76.58 I56 1993. IEEE catalog no. 93TH0513-2.

Hesham:1994:PTS

[HS94] E.-R. Hesham and B. D.

Shriver, editors. Proceedings of the Twenty-Seventh Hawaii International Conference on System Sciences. Vol. II: Software Technology, January 4–7, 1994, Wailea, HI, USA, volume 27. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5060-5. ISSN 1060-3425. LCCN ???? IEEE catalog no. 94TH0607-2.

Hertzberger:1995:HPM

[HS95a] Bob Hertzberger and Giuseppe Serazzi, editors. High-Performance computing and networking: International Conference and Exhibition, Milan, Italy, May 3–5, 1995: proceedings, number 919 in Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-59393-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Hungenahally:1995:PIQ

[HS95b] A. Hungenahally and A. Suresh. PVM implementation of quadtree building algorithms on SIMD hypercube system. IEEE International Conference on Algorithms and Architectures for Parallel Processing, 2:855–858, ???? 1995. IEEE catalog number 95TH0682-5.

Hoefler:2012:OPC

[HS12] Torsten Hoefler and Timo Schneider. Optimization principles for collective neighborhood communications. In Hollingsworth [Hol12], pages 98:1–98:?? ISBN 1-4673-0804-8. URL

http://conferences.computer.
org/sc/2012/papers/1000a028.
pdf.

Henriksen:2017:FPF

[HSE^{+}17] Troels Henriksen, Niels G. W. Serup, Martin Elsman, Fritz Henglein, and Cosmin E. Oancea. Futhark: purely functional GPUprogramming with nested parallelism and in-place array updates. ACM SIG-PLAN Notices, 52(6):556– 571, June 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Haeuser:1994:RNS

[HSMW94] J. Haeuser, M. Spel, J. Muylaert, and R. D. Williams. Results for the Navier– Stokes solver ParNSS on workstation clusters and IBM SP1 using PVM. In Wagner et al. [WPH94], pages 432–442. ISBN 0-471-95063-7. LCCN QA911.E95 1994.

Halbiniak:2021:EOH

[HSO^{+}21] Kamil Halbiniak, Lukasz Szustak, Tomasz Olas, Roman Wyrzykowski, and Pawel Gepner. Exploration of OpenCL heterogeneous programming for porting solidification modeling to CPU-GPU platforms. Concurrency and Computation: Practice and Experience, 33(4):e6011:1– e6011:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Heimel:2013:HOP

[HSP^{+}13] Max Heimel, Michael Saecker, Holger Pirk, Stefan Manegold, and Volker Markl. Hardware-oblivious parallelism for in-memory columnstores. Proceedings of the VLDB Endowment, 6(9): 709–720, July 2013. CODEN ???? ISSN 2150-8097.

Hormati:2012:SPS

[HSW^{+}12] Amir H. Hormati, Mehrzad Samadi, Mark Woh, Trevor Mudge, and Scott Mahlke. Sponge: portable stream programming on graphics engines. ACM SIG-PLAN Notices, 47(4):381– 392, April 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hu:2001:PCC

[HT01] Hong Hu and Edward L. Turner. Parallel CFD computing using shared memory OpenMP. Lecture Notes in Computer Science, 2073:1137–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2073/20731137.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2073/20731137.
pdf.

Howes:2008:U

[HT08] L. Howes and D. B. Thomas.

Efficient random number generation and application using CUDA. In Nguyen [Ngu08], chapter 37, pages 805–830. ISBN 0-321-51526-9. LCCN T385 .G6882 2008. URL http://

www.loc.gov/catdir/toc/
ecip0720/2007023985.html.

Ha:2008:NBP

[HTA08] Phuong Hoai Ha, Philippas Tsigas, and Otto J. Anshus. Non-blocking programming on multi-core graphics processors: (extended abstract). ACM SIGARCH Computer Architecture News, 36(5):19– 28, December 2008. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Hluchy:1999:GWF

[HTHD99] L. Hluchy, V. D. Tran, L. Halada, and M. Dobrucky. Ground water flow modelling in PVM. In Dongarra et al. [DLM99], pages 450–460. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Hariri:2016:PPA

[HTJ^{+}16] F. Hariri, T. M. Tran, A. Jocksch, E. Lanti, J. Progsch, P. Messmer, S. Brunner, C. Gheller, and L. Villard. A portable platform for accelerated PIC codes and its application to GPUs using OpenACC. Computer Physics Communications, 207(??): 69–82, October 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465516301242.

Huckle:1996:PIS

[Huc96] T. Huckle. PVM-implementation of sparse approximate inverse preconditioners for solving large sparse linear equations. Lecture Notes in Computer Science, 1156:166–173, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Humphres:1995:LBE

[Hum95] Christopher Wade Humphres. A load balancing extension for the PVM software system. M.e.e. thesis, Department of Electrical Engineering, University of Alabama, Tuscaloosa, AL, USA, 1995. viii + 98 pp.

Husbands:1998:MSD

[Hus98] Parry J. Husbands. MPI-StarT: Delivering network performance to numerical applications. In ACM [ACM98b], page ?? ISBN ???? LCCN

????  URL http://
www.supercomp.org/sc98/
papers/.

Huse:1999:CCD

[Hus99]  L. P. Huse.  Collective

communication on dedicated clusters of workstations. In Dongarra et al. [DLM99], pages 469–476. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Huse:2000:MOS

[Hus00] Lars Paul Huse. MPI optimization for SMP based clusters interconnected with SCI. Lecture Notes in Computer Science, 1908: 56–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080056.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080056.
pdf.

Huse:2001:LST

[Hus01] Lars Paul Huse. Layering SHMEM on top of MPI. Lecture Notes in Computer Science, 2131:44–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310044.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310044.
pdf.

Hamidouche:2016:CAO

[HVA^{+}16] Khaled Hamidouche, Akshay Venkatesh, Ammar Ahmad Awan, Hari Subramoni, Ching-Hsiang Chu, and Dhabaleswar K. Panda. CUDA-aware OpenSHMEM: Extensions and designs for high performance OpenSH-MEM on GPU clusters. Parallel Computing, 58(??):27– 36, October 2016. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819116300345.

Houzeaux:2011:HMO

[HVSC11] G. Houzeaux, M. Vázquez, X. Sáez, and J. M. Cela. Hybrid MPI–OpenMP performance in massively parallel computational fluid dynamics. In Tromeur-Dervout et al. [TDBEE11], pages 293–297. CO-DEN LNCSA6. ISBN 3-642-14437-3 (print), 3-642-14438-1 (e-book). ISSN 1439-7358. LCCN ???? URL

http://link.springer.com/
content/pdf/10.1007/978-

3-642-14438-7_31. Pro-

ceedings of the twentieth meeting, Parallel CFD 2008, held May 19–22, 2008 in Lyon, France.

Hoekstra:1995:CPP

[HVSH95] A. G. Hoekstra, F. Van der Linden, P. M. A. Sloot, and L. O. Hertzberger. Comparing the Parix and PVM parallel programming environments. In Fritzson and Finmo [FF95], pages 288– 292. ISBN 90-5199-229-7 (IOS Press), 4-274-90056-8 (Ohmsha). LCCN ????

Hager:2011:IHP

[HW11] Georg Hager and Gerhard Wellein. Introduction to high performance computing for scientists and engineers, volume 7 of Chapman and Hall/CRC computational science series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-4398-1192-X. xxv + 330 + 4 pp. LCCN QA76.88 .H34 2011.

Huang:2002:DDD

[HWM02] Wei Huang, Zhe Wang, and Jie Ma. Design of DMPI on DAWNING-3000. Lecture Notes in Computer Science, 2474:314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/

                                        service/series/0558/bibs/
                                        2474/24740314.htm;   http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2474/24740314.pdf.

He:2009:AVS

[HWS09] Jian He, Layne T. Watson, and Masha Sosonkina. Algorithm 897: VT-DIRECT95: Serial and parallel codes for the global optimization algorithm direct. ACM Transactions on Mathematical Software, 36(3):17:1–17:24, July 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remark [SWH15].

Hwang:1997:EMC

[HWW97] Kai Hwang, Choming Wang, and Cho-Li Wang. Evaluating MPI collective communication on the SP2, T3D, and Paragon multicomputers. In IEEE [IEE97c], pages 106–115. ISBN 0-8186-7764-3. LCCN QA76.9.A73I566 1997. IEEE catalog number 97TB100094.

Huang:2021:LBI

[HWW21] Liang-Tsung Huang, Kai-Cheng Wei, and Jian-An Wang. A lightweight BLASTP and its implementation on CUDA GPUs. The Journal of Supercomputing, 77(1):322–342, January 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

https://link.springer.
com/article/10.1007/s11227-
020-03267-1.

Huang:2013:ACM

[HWX^{+}13] Libo Huang, Zhiying Wang, Nong Xiao, Yongwen Wang, and Qiang Dou. Adaptive communication mechanism for accelerating MPI functions in NoC-based multicore processors. ACM Transactions on Architecture and Code Optimization, 10(3):18:1–18:??, September 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Huang:2020:POL

[HY20] Ming Hsiang Huang and Wuu Yang. PFACC: an OpenACC-like programming model for irregular nested parallelism. Software—Practice and Experience, 50(10): 1877–1904, October 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hellberg:1994:PPP

[HZ94] S. A. Hellberg and E. Zaluska. A portable parallel programming environment based around PCTE. Information and Software Technology, 36(7):419–425, July 1994. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

Hempel:1996:APT

[HZ96] R. Hempel and F. Zimmermann. On the automatic PARMACS-to-MPI transformation in application programs. In Liddell et al. [LCHS96], pages 1033–1034. ISBN 3-540-61142-8 (paperback). LCCN QA76.88 .H52 1996.

Hempel:1999:AMP

[HZ99] Rolf Hempel and Falk Zimmermann. Automatic migration from PARMACS to MPI in parallel Fortran applications. Scientific Programming, 7(1):39– 46, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

                                        iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=64cr5a4mg33tuhcbdr02%
                                        26referrer=parent%26backto=
                                        issue%2C3%2C7%3Bjournal%
                                        2C8%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Hou:2008:BBS

[HZG08] Qiming Hou, Kun Zhou, and Baining Guo. BSGP: bulk-synchronous GPU programming. ACM Transactions on Graphics, 27(3): 19:1–19:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Izadpanah:2019:PAP

[IADB19] Ramin Izadpanah, Benjamin A. Allan, Damian Dechev, and Jim Brandt. Production application performance data streaming for system monitoring. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(2): 8:1–8:??, June 2019. CO-DEN ???? ISSN 2376-3639.

URL https://dl.acm.org/
citation.cfm?id=3319498.

Isaila:2010:SMP

[IBC^{+}10] Florin Isaila, Francisco Javier Garcia Blas, Jesús Carretero, Wei keng Liao, and Alok Choudhary. A scalable Message Passing Interface implementation of an ad-hoc parallel I/O system. The International Journal of High Performance Computing Applications, 24(2):164– 184, May 2010. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/24/
2/164.full.pdf+html.

Isabel:2002:CMO

[ICC02] Dorta Isabel, León Coromoto, and Rodríguez Casiano. Comparing MPI and OpenMP implementations of the 0-1 knapsack problem. Parallel and Distributed Computing Practices, 5(2):129–137, June 2002. CODEN ???? ISSN 1097-2803.

Issman:1994:PME

[IDD94] E. Issman, G. Degrez, and J. De Keyser. A parallel multiblock Euler/Navier– Stokes solver on a cluster of workstations using PVM. In Gentzsch and Harms [GH94], pages 157–162. ISBN 0-387-57981-8 (New York), 3-540-57981-8 (Berlin). LCCN QA76.88.I57 1994. DM96.00. Two volumes.

Ibanez:2016:HMT

[IDS16] Dan Ibanez, Ian Dunn, and Mark S. Shephard. Hybrid MPI-thread parallelization of adaptive mesh operations. Parallel Computing, 52(??): 133–143, February 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819116000041.

IEEE:1991:PSA

[IEE91] IEEE, editor. Proceedings, Supercomputing ’91: Albuquerque, New Mexico, November 18–22, 1991. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-9158-1 (IEEE: case), 0-8186-2158-3 (IEEE: paper), 0-8186-6158-5 (IEEE: microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. IEEE catalog no. 91CH3058-5.

IEEE:1992:PSH

[IEE92]  IEEE, editor.  Proceed-

ings / Scalable High Performance Computing Conference, SHPCC-92, April 26–29, 1992, Williamsburg, Virginia. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2775-1. LCCN QA76.76.A65S33 1992. IEEE catalog no. 92TH0432-5.

IEEE:1993:DPC

[IEE93a] IEEE, editor. Digest of papers: Compcon spring ’93, San Francisco, California, February 22–26, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-3400-6. LCCN QA75.5.C58 1993. IEEE catalog no. 93CH3251-6.

IEEE:1993:PSI

[IEE93b]  IEEE, editor.  Proceed-

ings / Seventh International Parallel Processing Symposium, April 13–16, 1993, Newport Beach, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-3442-1. LCCN QA 76.58 I56 1993. IEEE catalog no. 93TH0513-2.

IEEE:1993:PIS

[IEE93c] IEEE, editor. Proceedings of the 2nd International Symposium on High Performance Distributed Computing, July 20–23, 1993, Spokane, Washington, Cavanaugh’s Inn at the Park, Proceedings of the International Symposium on High Performance Distributed Computing 2nd. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-3900-8, 0-8186-3901-6. LCCN QA76.9.D5I593 1993. IEEE catalog no. 93TH0550-4.

IEEE:1993:PFW

[IEE93d]  IEEE, editor.  Proceed-

ings of the Fourth Workshop on Future Trends of Distributed Computing Systems, September 22–24, 1993, Lisbon, Portugal. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4430-3. LCCN QA76.9.D5I335 1993. IEEE catalog no. 93TH0574-4.

IEEE:1993:PSP

[IEE93e] IEEE, editor. Proceedings, Supercomputing ’93: Portland, Oregon, November 15– 19, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

IEEE:1993:WHP

[IEE93f] IEEE, editor. Workshop on Heterogeneous Processing (1992: Beverly Hills, Calif.) Proceedings / Workshop on Heterogeneous Processing, March 23, 1992, Beverly Hills, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-2702-6. LCCN QA76.58 .W654 1992.

IEEE:1994:FSF

[IEE94a] IEEE, editor. Frontiers’95, the 5th Symposium on the Frontiers of Massively Parallel Computation: proceedings, February 6–9, 1995, McLean, Virginia. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6965-9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

IEEE:1994:IPN

[IEE94b] IEEE, editor. ICIP ’94:

proceedings, November 13– 16, 1994, Austin Convention Center, Austin, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6952-7 (casebound), 0-8186-6950-0 (paperback), 0-8186-6951-9 (microfiche). LCCN TA1637.I25 1994. Three volumes. IEEE catalog no. 94CH35708.

IEEE:1994:OOE

[IEE94c] IEEE, editor. Oceans 94: Oceans engineering for today’s technology and tomorrow’s preservation: proceedings, 13–16 September 13– 16, 1994, Brest, France, Oceans. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-2057-3, 0-7803-2056-5, 0-7803-2058-1. ISSN 0197-7385. LCCN TC 1505 O33197 1994. Three volumes. IEEE catalog no. 94CH3472-8.

IEEE:1994:PSI

[IEE94d] IEEE, editor. Proceedings / Second International Workshop on Configurable Distributed Systems, March 21– 23, 1994, Carnegie Mellon University, Pittsburgh, Pennsylvania. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5390-6. LCCN QA76.9.D5I595 1994. IEEE catalog no. 94TH0651-0.

IEEE:1994:PIF

[IEE94e] IEEE, editor. Proceedings of the 1994 IEEE Frequency Control Symposium (the 48th annual symposium), 1–3 June 1994, Westin Hotel-Copley Place, Boston, Massachusetts, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1945-1. LCCN TK 7872 O7 I34 1994. IEEE catalog no. 94CH3446-2.

IEEE:1994:PSP

[IEE94f] IEEE, editor. Proceedings of the Scalable Parallel Libraries Conference, October 6–8, 1993, Mississippi State, Mississippi. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-4980-1. LCCN QA76.58.S34 1993.

IEEE:1994:PTI

[IEE94g]  IEEE, editor.  Proceed-

ings of the Third IEEE International Symposium on High Performance Distributed Computing, August 2–5, 1994, San Francisco, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6395-2. LCCN QA76.9.D5I328 1994. IEEE catalog no. 94TH0667-

IEEE:1994:PSW

[IEE94h]  IEEE, editor.  Proceed-

ings, Supercomputing ’94: Washington, DC, November 14–18, 1994, Supercomputing. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6607-2, 0-8186-6605-6, 0-8186-6606-4. ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

IEEE:1995:IIC

[IEE95a] IEEE, editor. 1995 IEEE International Conference on Systems, Man, and Cybernetics: intelligent systems for the 21st century: Vancouver, British Columbia, Canada, October 22–25, 1995. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2559-1. LCCN TA168.I19 1995. Five volumes. IEEE catalog no. 95CH3576-7.

IEEE:1995:CPI

[IEE95b] IEEE, editor. Conference proceedings of the 1995 IEEE Fourteenth Annual International Phoenix Conference on Computers and Communications: Scottsdale, Arizona, USA, March 28–31, 1995, volume 14. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2493-5, 0-7803-2492-7, 0-7803-2494-3. LCCN TK7885.A1 I567 1995. IEEE catalog no. 95CH35751.

IEEE:1995:DPT

[IEE95c] IEEE, editor. Digest of papers / the Twenty-fifth International Symposium on Fault-Tolerant Computing, June 27–30, 1995, Pasadena, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7079-7. LCCN QA 76.9 F38 I57 1995. IEEE catalog no. 95CB35823.

IEEE:1995:ISE

[IEE95d] IEEE, editor. Ideas in Science and Electronics Exposition and Symposium. Proceedings: Albuquerque, NM, USA, 9–11 May 1995, volume 17 of Annual Ideas in Science and Electronics Exposition and Symposium Conference. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN ???? LCCN ????

IEEE:1995:IPR

[IEE95e] IEEE, editor. IEEE Pacific Rim Conference on Communications, Computers, and Signal Processing: proceedings / May 17–19, 1995, Victoria Conference Centre, Victoria, British Columbia, Canada. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2553-2. LCCN TK 5101 A1 I34 1995. IEEE catalog no. 95CH35765.

IEEE:1995:PIP

[IEE95f] IEEE, editor. Proceedings / 9th International Parallel Processing Symposium, April 25–28, 1995, Santa Barbara, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7074-6. LCCN QA 76.58 I56 1995. IEEE catalog no. 95TH8052.

IEEE:1995:PSI

[IEE95g]  IEEE, editor.  Proceed-

ings / Seventh IEEE Symposium on Parallel and Distributed Processing, October 25–28, 1995, San Antonio, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7195-5. LCCN QA 76.58 I42 1995. IEEE catalog number 95TB8131.

IEEE:1995:PEW

[IEE95h]  IEEE, editor.  Proceed-

ings: Euromicro Workshop on Parallel and Distributed Processing, San Remo, Italy, January 25–27, 1995, Euromicro Workshop on Parallel and Distributed Processing 1995; 3rd. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7031-2, 0-8186-7032-0. LCCN QA76.58 .E97 1995.

IEEE:1995:PIC

[IEE95i] IEEE, editor. Proceedings of the 15th International Conference on Distributed Computing Systems: Vancouver, BC, Canada, 30 May– 2 June 1995. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7025-8. LCCN ???? IEEE catalog number 95CH35784.

IEEE:1995:PSP

[IEE95j] IEEE, editor. Proceedings of the 1994 Scalable Parallel Libraries Conference: October 12–14, 1994, Mississippi State University, Mississippi. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

IEEE:1995:PFI

[IEE95k]  IEEE, editor.  Proceed-

ings of the Fourth IEEE International Symposium on High Performance Distributed Computing, August 2–4, 1995, Washington, DC, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7088-6. LCCN QA76.9.D5 I328 1995. IEEE catalog no. 95TB8075.

IEEE:1995:PNA

[IEE95l] IEEE, editor. Proceedings: the nineteenth annual International Computer Software and Applications Conference (COMPSAC ’95): August 9–11, 1995, Dallas, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7119-X. LCCN QA 76.6 C6295 1995. IEEE catalog no. 95CB35838.

IEEE:1996:ICH

[IEE96a]  IEEE, editor.  3rd In-

ternational Conference on High Performance Computing: proceedings, December 19–22, 1996, Trivandrum, India. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7557-8. LCCN QA76.88.I575 1996. IEEE catalog number 96TB100074.

IEEE:1996:EIS

[IEE96b] IEEE, editor. Eighth IEEE Symposium on Parallel and Distributed Processing: October 23–26, 1996, New Orleans, Louisiana. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7683-3, 0-8186-7685-X (microfiche). LCCN QA76.58 .I42 1996. IEEE Computer Society Press order number PR07683. IEEE Order Plan catalog number 96TB100088.

IEEE:1996:FSS

[IEE96c] IEEE, editor. Frontiers’96, the Sixth Symposium on the Frontiers of Massively Parallel Computation: October 27–31, 1996, Annapolis, Maryland: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7551-9. LCCN QA76.58 .S95 1996. IEEE catalog number 96TB100062.

IEEE:1996:PIS

[IEE96d] IEEE, editor. Proceedings of 1996 IEEE Second International Conference on Algorithms and Architectures for Parallel Processing, ICA PP ’96: June 11–13, 1996, Singapore. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3529-5 (softbound), 0-7803-3530-9 (microfiche). LCCN QA76.58.I33 1996. IEEE catalog number 96TH8204.

IEEE:1996:PII

[IEE96e]  IEEE, editor.  Proceed-

ings of IPPS ’96. The 10th International Parallel Processing Symposium: Honolulu, HI, USA, 15–19 April 1996. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7255-2. LCCN QA76.58 .I565 1996. IEEE catalog number 96TB100038. IEEE Computer Society Press order number PR07255.

IEEE:1996:PFI

[IEE96f] IEEE, editor. Proceedings of the Fifth IEEE International Symposium on High Performance Distributed Computing, Syracuse, NY, USA, 6– 9 August 1996. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7582-9. LCCN QA 76.88 I52 1996. IEEE catalog number TB100069.

IEEE:1996:PFE

[IEE96g]  IEEE, editor.  Proceed-

ings of the fourth Euromicro Workshop on Parallel and Distributed Processing (PDP ’96): January 24–26, 1996, Braga, Portugal. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7376-1. LCCN QA76.58 .E97 1996. IEEE order number PR07376.

IEEE:1996:PSI

[IEE96h] IEEE, editor. Proceedings of the Seventh Israeli Conference on Computer Systems and Software Engineering: June 12–13, 1996, Herzliya, Israel. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7536-5. LCCN QA75.5 .I75 1996. IEEE Computer Society Press Order Number PR07536.

IEEE:1996:PSM

[IEE96i]  IEEE, editor.  Proceed-

ings. Second MPI Developer’s Conference: Notre Dame, IN, USA, 1–2 July 1996. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

IEEE:1997:APD

[IEE97a] IEEE, editor. Advances in parallel and distributed computing: March 19–21, 1997, Shanghai, China: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7876-3 (paperback and case), 0-8186-7878-X (microfiche). LCCN QA76.58 .A4 1997.

IEEE:1997:PIP

[IEE97b] IEEE, editor. Proceedings. 11th International Parallel Processing Symposium, April 1–5, 1997, Geneva, Switzerland. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7793-7. LCCN QA76.58 .I56 1997. IEEE catalog number 97TB100107. IEEE Computer Society Press order number PR07792.

IEEE:1997:TIS

[IEE97c] IEEE, editor. Third International Symposium on High-Performance Computer Architecture: proceedings, February 1–5, 1997, San Antonio, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7764-3. LCCN QA76.9.A73I566 1997. IEEE catalog number 97TB100094.

IEEE:2002:STI

[IEE02] IEEE, editor. SC2002: From Terabytes to Insight. Proceedings of the IEEE ACM SC 2002 Conference, November 16–22, 2002, Baltimore, MD, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1524-X. LCCN ????

IEEE:2005:IPD

[IEE05] IEEE, editor. 19th International Parallel and Distributed Processing Symposium: proceedings: April 4–8, 2005, Denver, Colorado. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7695-2312-9. LCCN ???? IEEE Computer Society Order Number P2312.

Iida:2016:GET

[IFA^{+}16] Yuki Iida, Yusuke Fujii, Takuya Azumi, Nobuhiko Nishio, and Shinpei Kato. GPUrpc: Exploring transparent access to remote GPUs. ACM Transactions on Embedded Computing Systems, 16(1):17:1– 17:??, November 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

IFIP:1995:KWC

[IFI95] IFIP Working Group 2.5, editor. Kyoto Workshop 1995: Current Directions in Numerical Software and High Performance Computing, 19–20 October 1995, Kyoto, Japan. ????, ????, 1995. ISBN ???? LCCN

                                        ????  URL http://www.
                                        nsc.liu.se/~boein/ifip/
                                        kyoto/kyoto.html#reid;
                                         http://www.nsc.liu.se/
                                        ~boein/ifip/kyoto/workshop-
                                        info/proceedings/.

Iwasaki:2004:NPS

[IH04] Hideya Iwasaki and Zhenjiang Hu. A new parallel skeleton for general accumulative computations. International Journal of Parallel Programming, 32 (5):389–414, October 2004. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        32&issue=5&spage=389.

Izaguirre:2005:PMS

[IHM05] Jesús A. Izaguirre, Scott S. Hampton, and Thierry Matthey. Parallel multigrid summation for the N-body problem. Journal of Parallel and Distributed Computing, 65(8):949–962, August 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Iskra:2000:PMD

[IHvA^{+}00] K. A. Iskra, Z. W. Hendrikse, G. D. van Albada, B. J. Overeinder, and P. M. A. Sloot. Performance measurements on Dynamite/DPVM. Lecture Notes in Computer Science, 1908:27–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080027.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080027.
pdf.

Ierotheou:2005:GOC

[IJM^{+}05] C. S. Ierotheou, H. Jin, G. Matthews, S. P. Johnson, and R. Hood. Generating OpenMP code using an interactive parallelization environment. Parallel Computing, 31(10–12):999–1012, October/December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Iwama:2001:PLS

[IKM^{+}01] Kazuo Iwama, Daisuke Kawai, Shuichi Miyazaki, Yasuo Okabe, and Jun Umemoto. Parallelizing local search for CNF satisfiability using vectorization and PVM. Lecture Notes in Computer Science, 1982:123–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1982/19820123.htm;
 http://link.springer-
ny.com/link/service/series/

                                        0558/papers/1982/19820123.
                                        pdf.

Iwama:2002:PLS

[IKM^{+}02] Kazuo Iwama, Daisuke Kawai, Shuichi Miyazaki, Yasuo Okabe, and Jun Umemoto. Parallelizing local search for CNF satisfiability using vectorization and PVM. ACM Journal of Experimental Algorithmics, 7: 2, ???? 2002. CODEN ???? ISSN 1084-6654.

Iwashita:1994:IPE

[IM94] S. Iwashita and K. Murakami. Implementation and performances evaluation of KU PVM3/AP1000. Engineering Sciences Reports, Kyushu University, 16(3): 345–352, December 1994. CODEN SRKHEK. ISSN 0388-1717.

Ingle:1995:MAS

[IM95] N. K. Ingle and T. J. Mountziaris. A multifrontal algorithm for the solution of large systems of equations using network-based parallel computing. Computers & Chemical Engineering, 19(6-7):671–681, June–July 1995. CODEN CCENDW. ISSN 0098-1354.

Islam:2016:EMT

[IMS16] Tanzima Islam, Kathryn Mohror, and Martin Schulz. Exploring the MPI tool information interface: features and capabilities. The International Journal of High Performance Computing Applications, 30(2):212– 222, 2016. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:/

/journals.sagepub.com/
doi/full/10.1177/1094342015600507.

Ishizaka:2000:CGT

[IOK00] Kazuhisa Ishizaka, Motoki Obata, and Hironori Kasahara. Coarse-grain task parallel processing using the OpenMP backend of the OSCAR multigrain parallelizing compiler. Lecture Notes in Computer Science, 1940:457–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1940/19400457.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1940/19400457.
pdf.

Imbernon:2018:ELS

[IPG^{+}18] Baldomero Imbernón, Javier Prades, Domingo Giménez, José M. Cecilia, and Federico Silla. Enhancing largescale docking simulation on heterogeneous systems: An MPI vs rCUDA study. Future Generation Computer Systems, 79 (part 1)(??): 26–37, 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167739X17309974.

Ilroy:2001:IMP

[IRU01] Jonathan Ilroy, Cyrille Randriamaro, and Gil Utard. Improving MPI-I/O performance on PVFS. Lecture Notes in Computer Science, 2150:911–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2150/21500911.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2150/21500911.
                                        pdf.

Ilie:2016:AEC

[IS16] Silvana Ilie and Arne Storjohann. Abstracts of the 2015 East Coast Computer Algebra Day. ACM Communications in Computer Algebra, 50(1):35–39, March 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Satake:2012:OGA

[iSYS12] Shin ichi Satake, Hajime Yoshimori, and Takayuki Suzuki. Optimizations of a GPU accelerated heat conduction equation by a programming of CUDA Fortran from an analysis of a PTX file. Computer Physics Communications, 183(11): 2376–2385, November 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465512002068.

Imamura:2000:ASM

[ITKT00] Toshiyuki Imamura, Yuichi Tsujita, Hiroshi Koide, and Hiroshi Takemiya. An architecture of Stampi: MPI library on a cluster of parallel computers. Lecture Notes in Computer Science, 1908:200–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080200.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080200.
pdf.

Ishihara:1999:VBS

[ITT99] S. Ishihara, S. Tani, and A. Takahara. Virtual BUS: a simple implementation of an effortless networking system based on PVM. In Dongarra et al. [DLM99], pages 461–468. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Islam:2002:IAC

[ITT02] Mohammad Towhidul Islam, Parimala Thulasiraman, and Ruppa K. Thulasiram. Implementation of ant colony optimization algorithm for mobile ad hoc network applications: OpenMP experiences. Parallel and Distributed Computing Practices, 5(2):177–191, June 2002. CODEN ???? ISSN 1097-2803.

Iskra:2000:IDE

[IvdLH^{+}00] K. A. Iskra, F. van der Linden, Z. W. Hendrikse, B. J. Overeinder, G. D. van Albada, and P. M. A. Sloot. The implementation of dynamite: an environment for migrating PVM tasks. Operating Systems Review, 34(3): 40–55, July 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jatala:2017:SSG

[JAK17] Vishwesh Jatala, Jayvant Anantpur, and Amey Karkare. Scratchpad sharing in GPUs. ACM Transactions on Architecture and Code Optimization, 14(2):15:1–15:??, July 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Jabbarzadeh:1997:PSS

[JAT97] A. Jabbarzadeh, J. D. Atkinson, and R. I. Tanner. Parallel simulation of shear flow of polymers between structured walls by molecular dynamics simulation on PVM. Computer Physics Communications, 107(1–3): 123–136, December 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S001046559700088X.

Jacoby:1996:ADA

[JB96] G. H. (George H.) Jacoby and Jeannette V. Barnes, editors. Astronomical data analysis software and systems V: meeting held at Tucson, Arizona, 23–25 October 1995, volume 101 of Astronomical Society of the Pacific Conference Series. Astronomical Society of the Pacific, San Francisco, CA, USA, 1996. ISBN ???? ISSN 1080-7926. LCCN QB51.3.E43 A87 1995.

Juhasz:1996:PIP

[JC96] Z. Juhasz and D. Crookes.

A PVM implementation of a portable parallel image processing library. In Bode et al. [BDLS96], pages 188–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Jarzabek:2017:PEU

[JC17] Lukasz Jarzabek and Pawel Czarnul. Performance evaluation of unified memory and dynamic parallelism for selected parallel CUDA applications. The Journal of Supercomputing, 73(12):5378–5401, December 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/content/pdf/
10.1007/s11227-017-2091-
x.pdf.

Jin:2008:PEM

[JCH^{+}08] Haoqiang Jin, Barbara Chapman, Lei Huang, Dieter an Mey, and Thomas Reichstein. Performance evaluation of a multizone application in different OpenMP approaches. International Journal of Parallel Programming, 36(3): 312–325, June 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=3&spage=312.

Jaeger:2015:FGD

[JCP15] Julien Jaeger, Patrick Carribault, and Marc Pérache. Fine-grain data management directory for OpenMP 4.0 and OpenACC. Concurrency and Computation: Practice and Experience, 27(6):1528–1539, April 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jaksic:2020:HPF

[JCP^{+}20] Zoran Jakšić, Nicola Cadenelli, David Buchaca Prats, Jordà Polo, Josep Lluís Berral Garcia, and David Carrera Perez. A highly parameterizable framework for conditional restricted Boltzmann machine based workloads accelerated with FPGAs and OpenCL. Future Generation Computer Systems, 104(??):201– 211, March 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167739X19313676.

Jenkins:2014:PMD

[JDB^{+}14] John Jenkins, James Dinan, Pavan Balaji, Tom Peterka, Nagiza F. Samatova, and Rajeev Thakur. Processing MPI derived datatypes on noncontiguous GPUresident data. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2627–2637, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06600679-
abs.html.

Jeremiassen:1995:RFS

[JE95] T. E. Jeremiassen and S. J.

Eggers. Reducing false sharing on shared memory multiprocessors through compile time data transformations. ACM SIGPLAN Notices, 30(8):179–188, August 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jesshope:1993:LRV

[Jes93a] C. Jesshope. Latency reduction in VLSI routers. Parallel Processing Letters, 3 (4):485–494, December 1993. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Jesshope:1993:MCA

[Jes93b] C. Jesshope. The MPI chip and its applications. In Anonymous [Ano93c], pages 47–54. ISBN ???? LCCN ????

Jann:1995:AMP

[JF95] Joefon Jann and Hubertus Franke. Analysis of an MPI program using UTE on the IBM SP2. Research report RC 20085 (88832), IBM T. J. Watson Research Center, Yorktown Heights, NY, USA, 1995. 11 pp.

Johnson:2012:FOL

[JFGRF12] Tim Johnson, Pierre Fite-Georgel, Rahul Raguram, and Jan-Michael Frahm. Fast organization of large photo collections using CUDA. Lecture Notes in Computer Science, 6554: 463–476, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/content/pdf/
10.1007/978-3-642-35740-
4_36.

Jin:2000:AGO

[JFY00] Haoqiang Jin, Michael Frumkin, and Jerry Yan. Automatic generation of OpenMP directives and its application to computational fluid dynamics codes. Lecture Notes in Computer Science, 1940:440–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1940/19400440.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1940/19400440.
pdf.

Jackson:1997:SYE

[JH97] D. J. Jackson and C. W.

Humphres. A simple yet effective load balancing extension to the PVM software system. Parallel Computing, 22(12):1647–1660, February 21, 1997. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=22&
issue=12&aid=1112.

Jin:2011:HPC

[JJM^{+}11] Haoqiang Jin, Dennis Jespersen, Piyush Mehrotra, Rupak Biswas, Lei Huang, and Barbara Chapman. High performance computing using MPI and OpenMP on multi-core parallel systems. Parallel Computing, 37(9):562–575, September 2011. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167819111000159.

Jo:2017:PMA

[JJPL17] Gangwon Jo, Jaehoon Jung, Jiyoung Park, and Jaejin Lee. Poster: MAPA: an automatic memory access pattern analyzer for GPU applications. ACM SIGPLAN Notices, 52(8):443–444, August 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jin:2003:AMP

[JJY^{+}03] Haoqiang Jin, Gabriele Jost, Jerry Yan, et al. Automatic multilevel parallelization using OpenMP. Scientific Programming, 11(2): 177–190, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Januszewski:2010:ANS

[JK10] M. Januszewski and M. Kostur. Accelerating numerical solution of stochastic differential equations with CUDA. Computer Physics Communications, 181(1): 183–188, January 2010. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465509002999.

Jeun:2008:OPB

[JKHK08] Woo-Chul Jeun, Yang-Suk Kee, Soonhoi Ha, and Changdon Kee. Overcoming performance bottlenecks in using OpenMP on SMP clusters. Parallel Computing, 34 (10):570–592, October 2008. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Jan:2017:ITF

[JKM^{+}17] Bilal Jan, Fiaz Gul Khan, Bartolomeo Montrucchio, Anthony Theodore Chronopoulos, Shahaboddin Shamshirband, and Abdul Nasir Khan. Introducing ToPe– FFT: An OpenCL-based FFT library targeting GPUs. Concurrency and Computation: Practice and Experience, 29(21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jog:2013:OCT

[JKN^{+}13] Adwait Jog, Onur Kayiran, Nachiappan Chidambaram Nachiappan, Asit K. Mishra, Mahmut T. Kandemir, Onur Mutlu, Ravishankar Iyer, and Chita R. Das. OWL: cooperative thread array aware scheduling techniques for improving GPGPU performance. ACM SIGPLAN Notices, 48(4):395–406, April 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jani:2022:HST

[JKN22] Kunal Jani, Ankit Kumar, and Ronak Nahata. Hpcfolder: a simple tool used to parallelize algorithms using the message passing interface (MPI). The Journal of Supercomputing, 78(1):258–278, January 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

https://link.springer.
com/article/10.1007/s11227-
021-03896-0.

Jambunathan:2018:COB

[JL18] Revathi Jambunathan and Deborah A. Levin. CHAOS: an octree-based PIC–DSMC code for modeling of electron kinetic properties in a plasma plume using MPI–CUDA parallelization. Journal of Computational Physics, 373(??):571–604, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0021999118304601.

Jost:2005:WMP

[JLG05] G. Jost, J. Labarta, and J. Gimenez. What multilevel parallel programs do when you are not watching: a performance analysis case study comparing MPI/ OpenMP, MLP, and Nested OpenMP. Lecture Notes in Computer Science, 3349:29– ??, 2005.

Jie:2014:ASP

[JLS^{+}14] Liang Jie, KenLi Li, Lin Shi, RangSu Liu, and Jing Mei. Accelerating solidification process simulation for large-sized system of liquid metal atoms using GPU with CUDA. Journal of Computational Physics, 257(??): 521–535, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0021999113006803.

Julian-Moreno:2017:FPA

[JMdVG^{+}17] Guillermo Julián-Moreno, Jorge E. López de Vergara, Iván González, Luis de Pedro, Javier Royuela del Val, and Federico Simmross-Wattenberg. Fast parallel α-stable distribution function evaluation and parameter estimation using OpenCL in GPGPUs. Statistics and Computing, 27(5): 1365–1382, September 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Jorba:2001:SFF

[JML01] Josep Jorba, Tomàs Margalef, and Emilio Luque. Simulation of forest fire propagation on parallel & distributed PVM platforms. Lecture Notes in Computer Science, 2131:386–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2131/21310386.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2131/21310386.
                                        pdf.

Jung:2014:MCM

[JMS14] Jaewoon Jung, Takaharu Mori, and Yuji Sugita. Midpoint cell method for hybrid (MPI + OpenMP) parallelization of molecular dynamics simulations. Journal of Computational Chemistry, 35(14):1064– 1072, May 30, 2014. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Jo:2015:ALM

[JNL^{+}15] Gangwon Jo, Jeongho Nah, Jun Lee, Jungwon Kim, and Jaejin Lee. Accelerating LINPACK with MPI-OpenCL on clusters of multi-GPU nodes. IEEE Transactions on Parallel and Distributed Systems, 26 (7):1814–1825, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06846313-
abs.html.

Jones:1996:LLM

[Jon96] Chris R. Jones. Low latency MPI for Meiko CS/2 and ATM clusters. Thesis (m.a.), Department of Computer Science, University of California, Santa Barbara, Santa Barbara, CA, USA, 1996.

Joubert:1994:PAL

[Jou94] A. Joubert. Parallel algorithms for linear and nonlinear equations derived from networks. In Joubert et al. [JPTE94], pages 145–152. ISBN 0-444-81841-3. LCCN QA76.58 .P3794 1993.

Janssen:2022:GPU

[JPL22] Dylan M. Janssen, Wayne Pullan, and Alan Wee-Chung Liew. Graphics processing unit acceleration of the island model genetic algorithm using the CUDA programming platform. Concurrency and Computation: Practice and Experience, 34(2):e6286:1– e6286:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2012:OSP

[JPOJ12] Lei Jiang, Pragneshkumar B. Patel, George Ostrouchov, and Ferdinand Jamitzky. OpenMP-style parallelism in data-centered multicore computing with R. ACM SIGPLAN Notices, 47 (8):335–336, August 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Juric:1995:UPV

[JPP95] M. Juric, W. D. Potter, and M. Plaksin. Using the Parallel Virtual Machine for hunting snake-in-the-box codes. In Arabnia [Ara95], pages 97–102. ISBN 90-5199-187-8 (IOS Press), 4-274-90017-7 (Ohmsha). ISSN 0925-4986. LCCN ????

Joldes:2014:SSH

[JPT14] Mioara Joldes, Valentina Popescu, and Warwick Tucker. Searching for sinks for the Hénon map using a multiple-precision GPU arithmetic library. ACM SIGARCH Computer Architecture News, 42(4):63–68, September 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Joubert:1994:PCT

[JPTE94] G. R. Joubert, F. J. Peters, D. Trystram, and D. J. Evans, editors. Parallel computing: trends and applications: proceedings of the international conference ParCo93, Grenoble, France, 7–10 September 1993, volume 9 of Advances in parallel computing. North-Holland, Amsterdam, The Netherlands, 1994. ISBN 0-444-81841-3. LCCN QA76.58 .P3794 1993.

Jost:2010:EUH

[JR10] Gabriele Jost and Bob Robins. Experiences using hybrid MPI/OpenMP in the real world: Parallelization of a 3D CFD solver for multicore node clusters. Scientific Programming, 18(3–4):127– 138, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jimenez:2013:BCA

[JR13] Jesús Jiménez and Juan Ruiz de Miras. Box-counting algorithm on GPU and multi-core CPU: an OpenCL cross-platform study. The Journal of Supercomputing, 65(3):1327–1352, September 2013. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s11227-013-0885-z.

Jalowiecki:2021:BFS

[JRG21] Konrad Jałowiecki, Marek M. Rams, and Bartłomiej Gardas. Brute-forcing spinglass problems with CUDA. Computer Physics Communications, 260(??):Article 107728, March 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S001046552030360X.

Judd:1994:PIV

[JRM^{+}94] D. Judd, N. K. Ratha, P. K. McKinley, J. Weng, and A. K. Jain. Parallel implementation of vision algorithms on workstation clusters. In IEEE [IEE94e], pages 317–321 (vol. 3). ISBN 0-7803-1945-1. LCCN TK 7872 O7 I34 1994. IEEE catalog no. 94CH3446-2.

Jin:2013:PCU

[JS13] Hui Jin and Xian-He Sun. Performance comparison under failures of MPI and MapReduce: an analytical approach. Future Generation Computer Systems, 29(7):1808–1815, September 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13000290.

Jung:2005:DIM

[JSH^{+}05] Hyungsoo Jung, Dongin Shin, Hyuck Han, Jai W. Kim, Heon Y. Yeom, and Jongsuk Lee. Design and implementation of multiple fault-tolerant MPI over Myrinet (M^{3}). In ACM [ACM05], page 32. ISBN 1-59593-061-2. LCCN ????

Jaaskelainen:2015:PPP

[JSS^{+}15] Pekka Jääskeläinen, Carlos Sánchez de La Lama, Erik Schnetter, Kalle Raiskila, Jarmo Takala, and Heikki Berg. pocl: A performanceportable OpenCL implementation. International Journal of Parallel Programming, 43(5):752–785, October 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10766-014-0320-y.

Ju:1996:SPT

[JW96] Jiubin Ju and Yong Wang.

Scheduling PVM tasks. Operating Systems Review, 30 (3):22–31, July 1996. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jain:1996:IOP

[JWB96] Ravi Jain, John Werth, and James C. Browne, editors. Input/output and parallel and distributed computer systems. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-9735-5. LCCN QA76.58.I485 1996.

Jin:1995:LTP

[JY95] Lan Jin and Lan Yang. A laboratory for teaching parallel computing on parallel structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1):71– 75, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kumar:1995:MWD

[KA95] S. Kumar and H. Adeli. Minimum weight design of large structures on a network of workstations. Microcomputers in Civil Engineering, 10(6):423–432, November 1995. CODEN MCENE7. ISSN 0885-9507.

Kepner:2004:M

[KA04] Jeremy Kepner and Stan Ahalt. MatlabMPI. Journal of Parallel and Distributed Computing, 64(8):997–1005, August 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kumar:2013:GAI

[KA13] Piyush Kumar and Anupam Agrawal. Gpu-accelerated interactive visualization of 3D volumetric data using CUDA. International Journal of Image and Graphics (IJIG), 13(2):??, April 2013. CODEN ???? ISSN 0219-4678. URL http:

//doi.acm.org/10.1142/
S0219467813400032.

Krawezik:2002:SOV

[KAC02] Géraud Krawezik, Guillaume Alléon, and Franck Cappello. SPMD OpenMP versus MPI on a IBM SMP for 3 kernels of the NAS benchmarks. Lecture Notes in Computer Science, 2327:425–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2327/23270425.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2327/23270425.
pdf.

Krone:1996:ICF

[KAHS96] O. Krone, M. Aguilar, B. Hirsbrunner, and V. Sunderam. Integrating coordination features in PVM. In Ciancarini and Hankin [CH96], pages 432–435. ISBN 3-540-61052-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I52 1996.

Kapinos:2010:PPP

[KaM10] Paul Kapinos and Dieter an Mey. Productivity and performance portability of the OpenMP 3.0 tasking concept when applied to an engineering code written in Fortran 95. International Journal of Parallel Programming, 38(5– 6):379–395, October 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=5&spage=379.

Khan:2017:RCS

[KAMAMA17] Ayaz H. Khan, Mayez Al-Mouhamed, Muhammed Al-Mulhem, and Adel F. Ahmed. RT-CUDA: A software tool for CUDA code restructuring. International Journal of Parallel Programming, 45(3):551– 594, June 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Kanal:2012:PAI

[Kan12] M. E. Kanal. Parallel algorithm on inversion for adjacent pentadiagonal matrices with MPI. The Journal of Supercomputing, 59(2): 1071–1078, February 2012.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
59&issue=2&spage=1071.

Katamneni:1993:PPE

[Kat93] Sreevenu Katamneni. Parallel processing extensions to Verilog HDL using the PVM environment. M.s.e.e. thesis, Department of Electrical Engineering, University of Alabama, Tuscaloosa, AL, USA, 1993. viii + 108 pp.

Karlsson:1998:CCC

[KB98] S. Karlsson and M. Brorsson. A comparative characterization of communication patterns in applications using MPI and shared memory on an IBM SP2. Lecture Notes in Computer Science, 1362:189–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaiser:2001:OCC

[KB01] Timothy H. Kaiser and Scott B. Baden. Overlapping communication and computation with OpenMP and MPI. Scientific Programming, 9(2–3):73–81, Spring– Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/

app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C2%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Kruzel:2013:VOI

[KB13] Filip Kruzel and Krzysztof Banaś. Vectorized OpenCL implementation of numerical integration for higher order finite elements. Computers and Mathematics with Applications, 66(10):2030– 2044, December 2013. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S089812211300521X.

Kim:2021:GRP

[KB21] Mingyu Kim and Nakhoon Baek. A 3D graphics rendering pipeline implementation based on the openCL massively parallel processing. The Journal of Supercomputing, 77(7):7351– 7367, July 2021. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

https://link.springer.
com/article/10.1007/s11227-
020-03581-8.

Kabir:2002:DIS

[KBA02] Yacine Kabir and A. Belhadj-Aissa. Distributed image segmentation system by a multi-agents approach (under PVM environment). Lecture Notes in Computer Science, 2474:138–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/
service/series/0558/bibs/
2474/24740138.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740138.pdf.

Klemm:2009:RTM

[KBG^{+}09] Michael Klemm, Matthias Bezold, Stefan Gabriel, Ronald Veldema, and Michael Philippsen. Reparallelization techniques for migrating OpenMP codes in computational grids. Concurrency and Computation: Practice and Experience, 21(3): 281–299, March 10, 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kulkarni:2016:HAP

[KBG16] Kedar Kulkarni, Shreeya Badhe, and Geetanjali Gadre. HCA aware parallel communication library: A feasibility study for offloading MPI requirements. Supercomputing Frontiers and Innovations, 3(3):56– 60, ???? 2016. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic). URL http:/

/superfri.org/superfri/
article/view/109.

Knies:1994:SLL

[KBHA94] A. D. Knies, F. R. Barriuso, W. J. Harrod, and G. B. Adams, III. SLICC: a low latency interface for collective communications. In IEEE [IEE94h], pages 89–96. ISBN 0-8186-6607-2, 0-8186-6605-6, 0-8186-6606-4. ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Kitowski:1997:CPM

[KBM97] J. Kitowski, K. Boryczko, and J. Moscinski. Comparison of PVM and MPI performance in short-range molecular dynamics simulation. Lecture Notes in Computer Science, 1332:11–16, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kannan:2016:HPP

[KBP16] Ramakrishnan Kannan, Grey Ballard, and Haesun Park. A high-performance parallel algorithm for nonnegative matrix factorization. ACM SIGPLAN Notices, 51 (8):9:1–9:??, August 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ke:2004:RCM

[KBS04] Jian Ke, Martin Burtscher, and Evan Speight. Runtime compression of MPI messages to improve the performance and scalability of parallel applications. In ACM [ACM04], page 59. ISBN 0-7695-2153-3. LCCN ????

Klemm:2007:JIO

[KBVP07] Michael Klemm, Matthias Bezold, Ronald Veldema, and Michael Philippsen. JaMP: an implementation of OpenMP for a Java DSM. Concurrency and Computation: Practice and Experience, 19(18):2333–2352, December 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karamcheti:1994:SOM

[KC94] Vijay Karamcheti and Andrew A. Chien. Software overhead in messaging layers: where does the time go? ACM SIGPLAN Notices, 29(11):51–60, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/
citations/proceedings/
asplos/195473/p51-karamcheti/

.

Krawezik:2006:PCM

[KC06] Géraud Krawezik and Franck Cappello. Performance comparison of MPI and OpenMP on shared memory multiprocessors. Concurrency and Computation: Practice and Experience, 18(1):29– 61, January 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Knap:2019:PEU

[KC19] Marcin Knap and Paweł Czarnul. Performance evaluation of unified memory with prefetching and oversubscription for selected parallel CUDA applications on NVIDIA Pascal and Volta GPUs. The Journal of Supercomputing, 75(11):7625–7645, November 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.
                                        springer.com/content/pdf/
                                        10.1007/s11227-019-02966-
                                        8.pdf.

Kacsuk:1997:GDD

[KCD^{+}97] Peter Kacsuk, Jose C. Cunha, Gabor Dozsa, Joao Lourenco, Tibor Fadgyas, and Tiago Antao. A graphical development and debugging environment for parallel programs. Parallel Computing, 22(13):1747– 1770, February 28, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/parco/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=22&
                                        issue=13&aid=1126.

Konuru:1994:ULP

[KCP^{+}94a] R. Konuru, J. Casas, R. Prouty, S. Otto, and J. Walpole. A user-level process package for PVM. In Pierce and Regnier [PR94b], pages 48–55. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.58.S32 1994. IEEE catalog no. 94TH0637-9.

Konuru:1994:UPP

[KCP^{+}94b] R. Konuru, J. Casas, R. Prouty, S. Otto, and J. Walpole. A user-level process package for PVM. In Pierce and Regnier [PR94b], pages 48–55. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.58.S32 1994. IEEE catalog no. 94TH0637-9.

Kotselidis:2017:HMR

[KCR^{+}17] Christos Kotselidis, James Clarkson, Andrey Rodchenko, Andy Nisbet, John Mawer, and Mikel Luján. Heterogeneous managed runtime systems: a computer vision case study. ACM SIGPLAN Notices, 52(7): 74–82, July 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kanal:2012:MMC

[KD12] M. E. Kanal and M. Demiralp. A modified method of calculating High Dimensional Model Representation (HDMR) Terms for parallelization with MPI and CUDA. The Journal of Supercomputing, 62 (1):199–213, October 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0920-8542&volume=
                                        62&issue=1&spage=199.

Krotkiewski:2013:ESC

[KD13] Marcin Krotkiewski and Marcin Dabrowski. Efficient 3D stencil computations using CUDA. Parallel Computing, 39(10):533– 548, October 2013. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016781911300094X.

Kang:2018:PRS

[KDHZ18] Zhijiang Kang, Ze Deng, Wei Han, and Dongmei Zhang. Parallel reservoir simulation with OpenACC and domain decomposition. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

                                        URL https://www.mdpi.
                                        com/1999-4893/11/12/213.

Klingebiel:1995:COD

[KDL^{+}95a] P. Klingebiel, R. Diekmann, U. Lefarth, M. Fischer, and J. Seuss. CAMeL/PVM: an open, distributed CAE environment for modelling and simulating mechatronic systems. In Breitenecker and Husinsky [BH95], pages 645– 650. ISBN 0-444-82241-0. LCCN A76.9.C65E966 1995.

Klingebiel:1995:CPO

[KDL^{+}95b] P. Klingebiel, R. Diekmann, U. Lefarth, M. Fischer, and J. Seuss. CAMeL/PVM: An open, distributed CAE environment for modelling and simulating mechatronic systems. In Breitenecker and Husinsky [BH95], pages 645– 650. ISBN 0-444-82241-0. LCCN A76.9.C65E966 1995.

Kakimoto:2012:PCG

[KDSO12] Takeshi Kakimoto, Keisuke Dohi, Yuichiro Shibata, and Kiyoshi Oguri. Performance comparison of GPU programming frameworks with the striped Smith– Waterman algorithm. ACM SIGARCH Computer Architecture News, 40(5):70–75, December 2012. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). HEART ’12 conference proceedings.

Klemm:2012:EOV

[KDT^{+}12] Michael Klemm, Alejandro Duran, Xinmin Tian, Hideki Saito, and Diego Caballero. Extending OpenMP* with vector constructs for modern multicore SIMD architectures. Lecture Notes in Computer Science, 7312: 59–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/10.
                                        1007/978-3-642-30961-8_
                                        5/.

Komatitsch:2010:HOF

[KEGM10] Dimitri Komatitsch, Gordon Erlebacher, Dominik Göddeke, and David Michéa. High-order finite-element seismic wave propagation modeling with MPI on a large GPU cluster. Journal of Computational Physics, 229(20):7692–7714, October 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0021999110003396.

Kepner:2005:PPM

[Kep05] Jeremy Kepner. Parallel programming with MatlabMPI. World-Wide Web site., 2005.

                                        URL http://www.ll.mit.
                                        edu/MatlabMPI/.

Koitka:2016:NGA

[KF16] Sven Koitka and Christoph M. Friedrich. nmfgpu4R: GPUaccelerated computation of the non-negative matrix factorization (NMF) using CUDA capable hardware. The R Journal, 8(2):382–392, December 2016. ISSN 2073-4859.

                                        URL https://journal.
                                        r-project.org/archive/
                                        2016/RJ-2016-053.

Kale:1996:PMD

[KFA96] R. P. Kale, M. E. Fleharty, and P. M. Alsing. Parallel molecular dynamics visualization using MPI with MPE graphics. In IEEE [IEE96i], pages 104–110. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Kappiah:2005:JTD

[KFL05] Nandini Kappiah, Vincent W. Freeh, and David K. Lowenthal. Just in time dynamic voltage scaling: Exploiting inter-node slack to save energy in MPI programs. In ACM [ACM05], page 33. ISBN 1-59593-061-2. LCCN ????

Kramer-Fuhrmann:1994:TGP

[KFSS94] O. Kramer-Fuhrmann, L. Schafers, and C. Scheidler. TRAP-PER — a graphical programming environment for parallel systems. In Becks and Perret-Gallix [BPG94], pages 3–15. ISBN 981-02-1699-8. LCCN QC793.47.E4I58 1993.

Kowalik:1993:SPC

[KG93] Janusz S. Kowalik and Lucio Grandinetti, editors. Software for parallel computation: Proceedings of the NATO Advanced Workshop on Software for Parallel Computation, held at Cetraro, Cosenza, Italy, June 22–26, 1992, volume 106 of NATO ASI series. Series F, Computer and systems sciences. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-56451-9 (Berlin), 0-387-56451-9 (New York). LCCN QA76.58 .S629 1993.

Kohl:1996:PTF

[KG96] J. A. Kohl and G. A. Geist. The PVM 3.4 tracing facility and XPVM 1.1. In El-Rewini and Shriver [ERS96], pages 290–299. ISBN 0-8186-7324-9. ISSN 1060-3425. LCCN ???? Five volumes.

Kainz:2009:RCM

[KGB^{+}09] Bernhard Kainz, Markus Grabner, Alexander Bornik, Stefan Hauswiesner, Judith Muehl, and Dieter Schmalstieg. Ray casting of multiple volumetric datasets with polyhedral boundaries on manycore GPUs. ACM Transactions on Graphics, 28(5):152:1–152:9, December 2009. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Keller:2003:TEE

[KGK^{+}03] Rainer Keller, Edgar Gabriel, Bettina Krammer, Matthias S. Müller, and Michael M. Resch. Towards efficient execution of MPI applications on the Grid: Porting and optimization issues. Journal of Grid Computing, 1(2):133–149, ????

2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http://

ipsapp008.kluweronline.
com/IPS/content/ext/x/
J/6160/I/4/A/4/abstract.
htm.

Keller:2010:RAM

[KGRD10] Rainer Keller, Edgar Gabriel, Michael Resch, and Jack Dongarra, editors. Recent Advances in the Message Passing Interface: 17th European MPI Users’ Group Meeting, EuroMPI 2010, Stuttgart, Germany, September 12–15, 2010. Proceedings, volume 6305 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CO-DEN LNCSD9. ISBN 3-642-15645-2 (print), 3-642-15646-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:/

/www.springerlink.com/
content/978-3-642-15646-
5.

Kafura:1996:CCC

[KH96] D. Kafura and L. Huang.

Collective communication and communicators in mpi++. In IEEE [IEE96i], pages 79– 86. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Kwon:2010:SPC

[KH10] Seongnam Kwon and Soonhoi Ha. Serialized parallel code generation framework for MPSoC. ACM Transactions on Design Automation of Electronic Systems, 15(2): 11:1–11:??, February 2010. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Karrenberg:2012:IPO

[KH12] Ralf Karrenberg and Sebastian Hack. Improving performance of OpenCL on CPUs. Lecture Notes in Computer Science, 7210: 1–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/10.
                                        1007/978-3-642-28652-0_
                                        1/.

Kramer:2015:SET

[KH15] Stephan C. Kramer and Johannes Hagemann. SciPAL: Expression templates and composition closure objects for high performance computational physics with CUDA and OpenMP. ACM Transactions on Parallel Computing (TOPC), 1(2):15:1– 15:??, January 2015. CO-DEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Khanna:2013:HPN

[Kha13]  Gaurav Khanna.  High-

precision numerical simulations on a CUDA GPU:

Kerr black hole tails. Journal of Scientific Computing, 56(2):366–380, August 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10915-012-9679-3;
 http://link.springer.
com/content/pdf/10.1007/
s10915-012-9679-3.pdf.

Kielmann:1999:MMC

[KHB^{+}99] Thilo Kielmann, Rutger F. H. Hofman, Henri E. Bal, Aske Plaat, and Raoul A. F. Bhoedjang. Mag-PIe: MPI’s collective communication operations for clustered wide area systems. ACM SIGPLAN Notices, 34 (8):131–140, August 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
proceedings/ppopp/301104/
p131-kielmann/.

Kallenborn:2019:MPC

[KHBS19] Felix Kallenborn, Christian Hundt, Sebastian Böser, and Bertil Schmidt. Massively parallel computation of atmospheric neutrino oscillations on CUDAenabled accelerators. Computer Physics Communications, 234(??):235–244, January 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0010465518302790.

Kucukboyaci:2001:PPT

[KHS01] Vefa Kucukboyaci, Alireza Haghighat, and Glenn E. Sjoden. Performance of PENTRAN TM 3-D parallel particle transport code on the IBM SP2 and PC-TRAN cluster. Lecture Notes in Computer Science, 2131:36–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2131/21310036.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2131/21310036.
                                        pdf.

Kjolstad:2012:ADG

[KHS12] Fredrik Kjolstad, Torsten Hoefler, and Marc Snir. Automatic datatype generation and optimization. ACM SIGPLAN Notices, 47 (8):327–328, August 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Kojima:2017:HLG

[KI17] Kensuke Kojima and Atsushi Igarashi. A Hoare logic for GPU kernels. ACM Transactions on Computational Logic, 18(1):3:1– 3:??, April 2017. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic).

Kikuchi:1993:PAS

[Kik93]  S. Kikuchi.  Paralleliza-

tion assist system. Joho-Shori (J. Information Processing Soc. Japan), 34(9): 1158–1169, September 1993. CODEN JOSHA4. ISSN 0447-8053.

Kranz:1993:IMP

[KJA^{+}93] David Kranz, Kirk L. Johnson, Anant Agarwal, John Kubiatowicz, and Beng-Hong Lim. Integrating message-passing and sharedmemory: early experience. ACM SIGPLAN Notices, 28 (7):54–63, July 1993. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kwon:2012:HAO

[KJEM12] Okwan Kwon, Fahed Jubair, Rudolf Eigenmann, and Samuel Midkiff. A hybrid approach of OpenMP for clusters. ACM SIGPLAN Notices, 47(8):75–84, August 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Kim:2016:DOF

[KJJ^{+}16] Junghyun Kim, Gangwon Jo, Jaehoon Jung, Jungwon Kim, and Jaejin Lee. A distributed OpenCL framework using redundant computation and data replication. ACM SIGPLAN Notices, 51(6):553–569, June 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kemelmakher:1998:SAR

[KK98] M. Kemelmakher and O. Kremien. Scalable and adaptive resource sharing in PVM. Lecture Notes in Computer Science, 1497:196–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karniadakis:2002:PSC

[KK02a] George Em Karniadakis and Robert M. Kirby. Parallel Scientific Computing in C++ and MPI: a Seamless Approach to Parallel Algorithms. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-52080-0 (paperback), 0-521-81754-4 (hardcover). xi + 616 pp. LCCN QA76.58 .K37 2003. US$50.00 (paperback), US$130.00 (hardcover). URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/cam031/2002034805.

html;   http://www.loc.
gov/catdir/samples/cam033/
2002034805.html;   http:
//www.loc.gov/catdir/toc/
cam031/2002034805.html.

Krysztop:2002:IFP

[KK02b] Bartosz Krysztop and Henryk Krawczyk. Improving flexibility and performance of PVM applications by distributed partial evaluation. Lecture Notes in Computer Science, 2474:376–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/
service/series/0558/bibs/
2474/24740376.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740376.pdf.

Kronbichler:2019:FMF

[KK19] Martin Kronbichler and Katharina Kormann. Fast matrix-free evaluation of discontinuous Galerkin finite element operators. ACM Transactions on Mathematical Software, 45(3):29:1– 29:40, August 2019. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/citation.
cfm?id=3325864.

Kohnke:2021:CFM

[KKB^{+}21] Bartosz Kohnke, Carsten Kutzner, Andreas Beckmann, Gert Lube, Ivo Kabadshow, Holger Dachsel, and Helmut Grubmüller. A CUDA fast multipole method with highly efficient M2L far field evaluation. The International Journal of High Performance Computing Applications, 35(1): 97–117, January 1, 2021. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:

//journals.sagepub.com/
doi/full/10.1177/1094342020964857.

Kranzlmuller:2004:RAP

[KKD04] Dieter Kranzlmüller, Péter Kacsuk, and Jack J. Dongarra, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 11th European PVM/MPI Users’ Group Meeting, Budapest, Hungary, September 19–22, 2004: proceedings, volume 3241 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CO-DEN LNCSD9. ISBN 3-540-23163-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E973 2004. URL http:

//www.springerlink.com/
openurl.asp?genre=issue&
issn=0302-9743&volume=
3241;   http://www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b100820.

Kranzlmuller:2005:RAP

[KKD05] Dieter Kranzlmüller, Peter Kacsuk, and Jack Dongarra. Recent advances in Parallel Virtual Machine and Message Passing Interface. The International Journal of High Performance Computing Applications, 19(2):99– 101, Summer 2005. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/19/
2/99.full.pdf+html.

Kranzlmuller:2003:RAP

[KKDV03] Dieter Kranzlmüller, Peter Kacsuk, Jack Dongarra, and Jens Volkert. Recent advances in parallel virtual machine and message passing interface (select papers from the EuroPVMMPI 2002 Conference). The International Journal of High Performance Computing Applications, 17(1):3–5, Spring 2003. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Kee:2003:POP

[KKH03] Yang-Suk Kee, Jin-Soo Kim, and Soonhoi Ha. ParADE: An OpenMP programming environment for SMP cluster systems. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN ???? URL http:/

/www.sc-conference.org/
sc2003/inter_cal/inter_
cal_detail.php?eventid=

                                        10708#0;   http://www.
                                        sc-conference.org/sc2003/
                                        paperpdfs/pap130.pdf.

Kwon:2008:RPP

[KKJ^{+}08] Seongnam Kwon, Yongjoo Kim, Woo-Chul Jeun, Soonhoi Ha, and Yunheung Paek. A retargetable parallelprogramming framework for MPSoC. ACM Transactions on Design Automation of Electronic Systems, 13(3):39:1–39:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kim:2011:ASC

[KKLL11] Jungwon Kim, Honggyu Kim, Joo Hwan Lee, and Jaejin Lee. Achieving a single compute device image in OpenCL for multiple GPUs. ACM SIGPLAN Notices, 46 (8):277–288, August 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’11 Conference proceedings.

Karami:2015:SPA

[KKM15] Ali Karami, Farshad Khunjush, and Seyyed Ali Mirsoleimani. A statistical performance analyzer framework for OpenCL kernels on Nvidia GPUs. The Journal of Supercomputing, 71(8):2900–2921, August 2015. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11227-014-1338-z.

Konstantinou:2001:TTO

[KKP01] Dimitris Konstantinou, Nectarios Koziris, and George Papakonstantinou. TOP-PER: a tool for optimizing the performance of parallel applications. Lecture Notes in Computer Science, 2131:148–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310148.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310148.
pdf.

Kobler:2001:DOP

[KKV01] Rene Kobler, Dieter Kranzlmüller, and Jens Volkert. Debugging OpenMP programs using event manipulation. Lecture Notes in Computer Science, 2104: 81–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2104/21040081.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2104/21040081.
pdf.

Karrels:1994:PAM

[KL94] E. Karrels and E. Lusk. Performance analysis of MPI programs. In Dongarra and Tourancheau [DT94], pages 195–200. ISBN 0-89871-343-9. LCCN QA76.58.I568 1994.

Kofakis:1995:DPI

[KL95] P. Kofakis and J. Louis. Distributed parallel implementation of seismic algorithms. In Hassanzadeh [Has95], pages 229–238. CO-DEN PSISDG. ISBN 0-8194-1930-3. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.2571.

Liao:2011:DEM

[kL11] Wei keng Liao. Design and evaluation of MPI file domain partitioning methods under extent-based file locking protocol. IEEE Transactions on Parallel and Distributed Systems, 22(2):260– 272, February 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liao:2006:SDI

[kLCC^{+}06] Wei keng Liao, Kenin Coloma, Alok Choudhary, Lee Ward, Eric Russell, and Neil Pundit. Scalable design and implementations for MPI parallel overlapping I/O. IEEE Transactions on Parallel and Distributed Systems, 17(11): 1264–1276, November 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liao:2007:CCS

[kLCCW07] Wei keng Liao, Kenin Coloma, Alok Choudhary, and Lee Ward. Cooperative client-side file caching for MPI applications. The International Journal of High Performance Computing Applications, 21(2):144– 154, May 2007. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/21/
2/144.full.pdf+html.

Kang:2020:IMC

[KLH^{+}20] Q. Kang, S. Lee, K. Hou, R. Ross, A. Agrawal, A. Choudhary, and W. Liao. Improving MPI collective I/O for high volume noncontiguous requests with intra-node aggregation. IEEE Transactions on Parallel and Distributed Systems, 31(11): 2682–2695, 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kumar:2019:FOP

[KLM^{+}19] Ramavarmaraja Kishor Kumar, Vladimir Loncar, Paulsamy Muruganandam, Sadhan K. Adhikari, and Antun Balaz. C and Fortran OpenMP programs for rotating Bose–Einstein condensates. Computer Physics Communications, 240(??): 74–82, July 2019. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465519300827.

Klawonn:2015:HMO

[KLR^{+}15] Axel Klawonn, Martin Lanser, Oliver Rheinbach, Holger Stengel, and Gerhard Wellein. Hybrid MPI/OpenMP parallelization in FETI–DP methods. In Mehl et al. [MBS15], pages 67–84. ISBN 3-319-22996-6, 3-319-22997-4 (ebook). LCCN QA71-90;

                                        TA329. URL http://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-22997-3_
                                        4/.

Kutyniok:2016:SFD

[KLR16] Gitta Kutyniok, Wang-Q Lim, and Rafael Reisenhofer. ShearLab 3D: Faithful digital shearlet transforms based on compactly supported shearlets. ACM Transactions on Mathematical Software, 42(1):5:1–5:42, February 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kim:2015:OBU

[KLV15] Jungwon Kim, Seyong Lee, and Jeffrey S. Vetter. An OpenACC-based unified programming model for multi-accelerator systems. ACM SIGPLAN Notices, 50 (8):257–258, August 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Khanna:2010:NMG

[KM10] Gaurav Khanna and Justin McKennon. Numerical modeling of gravitational wave sources accelerated by OpenCL. Computer Physics Communications, 181(9): 1605–1611, September 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465510001682.

Kormicki:1996:PLS

[KMC96] M. Kormicki, A. Mahmood, and B. S. Carlson. Parallel logic simulation on a network of workstations using PVM. In IEEE [IEE96b], pages 2–9. ISBN 0-8186-7683-3, 0-8186-7685-X (microfiche). LCCN QA76.58 .I42 1996. IEEE Computer Society Press order number PR07683. IEEE Order Plan catalog number 96TB100088.

Kormicki:1997:PLS

[KMC97] Maciek Kormicki, Ausif Mahmood, and Bradley S.

Carlson. Parallel logic simulation on a network of workstations using parallel virtual machine. ACM Transactions on Design Automation of Electronic Systems, 2(2):123–134, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

tronic). URL http://www.
acm.org/pubs/articles/
journals/todaes/1997-2-
2/p123-kormicki/p123-kormicki.
pdf;   http://www.acm.
org/pubs/citations/journals/
todaes/1997-2-2/p123-kormicki/

.

Komatitsch:2009:PHO

[KME09] Dimitri Komatitsch, David Michéa, and Gordon Erlebacher. Porting a highorder finite-element earthquake modeling application to NVIDIA graphics cards using CUDA. Journal of Parallel and Distributed Computing, 69(5): 451–460, May 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Koholka:1999:MPR

[KMG99] R. Koholka, H. Mayer, and A. Goller. MPI-parallelized radiance on SGI CoW and SMP. Lecture Notes in Computer Science, 1557: 549–558, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kumar:2014:OMC

[KMH^{+}14] Sameer Kumar, Amith Mamidala, Philip Heidelberger, Dong Chen, and Daniel Faraj. Optimization of MPI collective operations on the IBM Blue Gene/Q supercomputer. The International Journal of High Performance Computing Applications, 28(4):450–464, November 2014. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/28/
4/450.

Kobayashi:2016:HSV

[KMK16] Ryohei Kobayashi, Tomohiro Misono, and Kenji Kise. A high-speed Verilog HDL simulation method using a lightweight translator. ACM SIGARCH Computer Architecture News, 44 (4):26–31, September 2016. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Kouzinopoulos:2015:MSM

[KMM15] Charalampos S. Kouzinopoulos, Panagiotis D. Michailidis, and Konstantinos G. Margaritis. Multiple string matching on a GPU using CUDAs. Scalable Computing: Practice and Experience, 16(2):121–138, ???? 2015. CODEN ???? ISSN 1895-1767. URL https://

                                        www.scpe.org/index.php/
                                        scpe/article/view/1085.

Kirk:2010:PMP

[KmWH10] David B. Kirk and Wen mei W. Hwu. Programming Massively Parallel Processors: a Hands-on Approach. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2010. ISBN 0-12-381472-3. xviii + 258 pp. LCCN QA76.642 .K57 2010. Chapter 7 (pages 125–140) discusses GPU floating-point considerations.

Kalns:1995:DPD

[KN95] E. T. Kalns and L. M. Ni. DaReL: a portable data redistribution library for distributed-memory machines. In IEEE [IEE95j], pages 78–87. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Katouda:2017:MOH

[KN17] Michio Katouda and Takahito Nakajima. MPI/OpenMP hybrid parallel algorithm for resolution of identity second-order Møller–Plesset perturbation calculation of analytical energy gradient for massively parallel multicore supercomputers. Journal of Computational Chemistry, 38(8):489–507, March 30, 2017. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kono:2018:EOW

[KNH^{+}18] Fumiya Kono, Naohito Nakasato, Kensaku Hayashi, Alexander Vazhenin, and Stanislav Sedukhin. Evaluations of OpenCL-written tsunami simulation on FPGA and comparison with GPU implementation. The Journal of Supercomputing, 74 (6):2747–2775, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kasprzyk:2002:APV

[KNT02] Leszek Kasprzyk, Ryszard Nawrowski, and Andrzej Tomczewski. Application of a parallel virtual machine for the analysis of a luminous field. Lecture Notes in Computer Science, 2474:122–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/
service/series/0558/bibs/
2474/24740122.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740122.pdf.

Komura:2014:CPG

[KO14] Yukihiro Komura and Yutaka Okabe. CUDA programs for the GPU computing of the Swendsen– Wang multi-cluster spin flip algorithm: 2D and 3D Ising, Potts, and XY models. Computer Physics Communications, 185(3):1038– 1043, March 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465513003743.

Kambites:2001:OLI

[KOB01] M. E. Kambites, J. Obdržálek, and J. M. Bull. An OpenMP-like interface for parallel programming in Java. Concurrency and Computation: Practice and Experience, 13(8–9):793– 814, July/August 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/84503220/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=84503220&PLACEBO=IE.
pdf.

Kasahara:2001:ACG

[KOI01] Hironori Kasahara, Motoki Obata, and Kazuhisa Ishizaka. Automatic coarse grain task parallel processing on SMP using OpenMP. Lecture Notes in Computer Science, 2017:189–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2017/20170189.htm;
 http://link.springer-
ny.com/link/service/series/

0558/papers/2017/20170189.
pdf.

Komura:2015:OPS

[Kom15] Yukihiro Komura. OpenACC programs of the Swendsen–Wang multi-cluster spin flip algorithm. Computer Physics Communications, 197(??):298–303, December 2015. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465515003197.

Koniges:2000:ISP

[Kon00] Alice E. Koniges, editor.

Industrial Strength Parallel Computing. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 1-55860-540-1. xxv + 597 pp. LCCN QA76.58 .I483 2000.

Kauranne:1995:OHM

[KOS^{+}95a] T. Kauranne, J. Oinonen, S. Saarinen, O. Serimaa, and J. Hietaniemi. The operational HIRLAM 2 model on parallel computers (weather forecasting). In Hoffmann and Kreitz [HK95], pages 63–74. ISBN 981-02-2211-4. LCCN QC866.E26 1994.

Koski:1995:STL

[Kos95b] Kimmo Koski. A step towards large scale parallelism: Building a parallel computing environment from heterogeneous resources. Future Generation Computer Systems, 11(4–5):491–498, August 1995. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Konuru:1997:MUL

[KOW97] Ravi B. Konuru, Steve W. Otto, and Jonathan Walpole. A migratable user-level process package for PVM. Journal of Parallel and Distributed Computing, 40(1): 81–102, January 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.idealibrary.com/
links/doi/10.1006/jpdc.
1996.1270/production;
http://www.idealibrary.
com/links/doi/10.1006/
jpdc.1996.1270/production/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/
jpdc.1996.1270/production/
ref.

Kermarrec:1996:PDS

[KP96] Y. Kermarrec and L. Pautet. Programming distributed systems with both Ada 95 and PVM. In Toussaint [Tou96], pages 206–216. ISBN 3-540-60757-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A35I57 1995.

Kuckuk:2013:IPD

[KPK13] Sebastian Kuckuk, Tobias Preclik, and Harald Köstler. Interactive particle dynamics using OpenCL and Kinect. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 28(6):519– 536, 2013.

Klockner:2012:PPS

[KPL^{+}12] Andreas Klöckner, Nicolas Pinto, Yunsup Lee, Bryan Catanzaro, Paul Ivanov, and Ahmed Fasih. Py-CUDA and PyOpenCL: a scripting-based approach to GPU run-time code generation. Parallel Computing, 38(3):157–174, March 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819111001281.

Kolesnichenko:2016:CBG

[KPNM16] Alexey Kolesnichenko, Christopher M. Poskitt, Sebastian Nanz, and Bertrand Meyer. Contract-based general-purpose GPU programming. ACM SIG-PLAN Notices, 51(3):75– 84, March 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kuhn:2000:OVT

[KPO00] Bob Kuhn, Paul Petersen, and Eamonn O’Toole. OpenMP versus threading in C/C++. Concurrency: practice and experience, 12(12):1165– 1176, October 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76500354/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500354&PLACEBO=IE.
pdf.

Kamal:2005:SVT

[KPW05] Humaira Kamal, Brad Penoff, and Alan Wagner. SCTP versus TCP for MPI. In ACM [ACM05], page 30. ISBN 1-59593-061-2. LCCN ????

Klemm:2021:OAH

[KQT^{+}21] Michael Klemm, Eduardo Quiñones, Tucker Taft, Dirk Ziegenbein, and Sara Royuela. The OpenMP API for high integrity systems: Moving responsibility from users to vendors. ACM SIGADA Ada Letters, 40(2):48–50, April 2021. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463478.3463480.

Klimach:2009:PCH

[KR09] Harald Klimach and Sabine P. Roller. Parallel coupling of heterogeneous domains with KOP3D using PACX-MPI. In Tuncer et al. [TGEM09], pages 339–345. CODEN LNCSA6. ISBN 3-540-92743-3 (print), 3-540-92744-1 (e-book). ISSN 1439-7358. LCCN ???? URL

http://link.springer.com/
content/pdf/10.1007/978-
3-540-92744-0_42. Parallel

CFD 2007 was held in Antalya, Turkey, from May 21 to 24, 2007.

Kranzlmuller:2002:RAP

[Kra02] Dieter Kranzlmuller, editor.

Recent advances in parallel virtual machine and message passing interface: 9th European PVM/MPI Users’ Group Meeting, Linz, Austria, September 29–October 2, 2002: proceedings, volume 2474 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-44296-0 (softcover). LCCN QA76.58 .E975 2002. Also available via the World Wide Web.

Kouetcha:2017:USP

[KRC17] Daniella Nguemalieu Kouetcha, Hamidréza Ramézani, and Nathalie Cohaut. Ultrafast scalable parallel algorithm for the radial distribution function histogramming using MPI maps. The Journal of Supercomputing, 73(4):1629–1653, April 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kunaseth:2013:ASD

[KRG13] Manaschai Kunaseth, David F.

Richards, and James N.

Glosli. Analysis of scalable data-privatization threading algorithms for hybrid MPI/OpenMP parallelization of molecular dynamics. The Journal of Supercomputing, 66(1):406–430, October 2013. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s11227-013-0915-x.

Kalentev:2011:CCL

[KRKS11] Oleksandr Kalentev, Abha Rai, Stefan Kemnitz, and Ralf Schneider. Connected component labeling on a 2D grid using CUDA. Journal of Parallel and Distributed Computing, 71(4): 615–620, April 2011. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kranzlmueller:1999:MOM

[KRS99] D. Kranzlmueller, R. Reussner, and C. Schaubschlaeger. Monitor overhead measurement with SKaMPI. In Dongarra et al. [DLM99], pages 43–50. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Kotsis:1996:EEP

[KS96] G. Kotsis and F. Sukup. Efficiency evaluation of PVM 2.X, PVM 3.X, P4, EX-PRESS and LINDA on a workstation cluster using the NAS parallel benchmarks. In Zaky and Lewis [ZL96], pages 149–171. ISBN 0-7923-9675-8. LCCN QA76.58.T65 1996.

Krantz:1997:CSC

[KS97] A. T. Krantz and V. S.

Sunderam. Client server computing on message passing systems: Experiences with PVM-RPC. Lecture Notes in Computer Science, 1300:110–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krawczyk:2001:PIM

[KS01] Henryk Krawczyk and Jamil Saif. Parallel image matching on PC cluster. Lecture Notes in Computer Science, 2131:312–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310312.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310312.
pdf.

Kim:2013:MPE

[KS13] Yooseong Kim and Aviral Shrivastava. Memory performance estimation of CUDA programs. ACM Transactions on Embedded Computing Systems, 13(2):

21:1–21:??, September 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kaliman:2015:SNU

[KS15a] Ilya A. Kaliman and Lyudmila V. Slipchenko. Software news and updates: Hybrid MPI/OpenMP parallelization of the effective fragment potential method in the libefp software library. Journal of Computational Chemistry, 36(2):129–135, January 15, 2015. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kovanen:2015:TAC

[KS15b] Janne Kovanen and Tapani Sarjakoski. Tilewise accumulated cost surface computation with graphics processing units. ACM Transactions on Spatial Algorithms and Systems (TSAS), 1(2): 8:1–8:27, November 2015. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic). URL http:

//dl.acm.org/citation.
cfm?id=2803172.

Klinkenberg:2020:CRL

[KSB^{+}20] Jannis Klinkenberg, Philipp Samfass, Michael Bader, Christian Terboven, and Matthias S. Müller. CHAMELEON: Reactive load balancing for hybrid MPI + OpenMP task-parallel applications.

Journal of Parallel and Distributed Computing, 138(??): 55–64, April 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731519305180.

Knight:2019:TES

[KSC^{+}19] Louise Knight, Polona Stefanic, Matej Cigale, Andrew C. Jones, and Ian Taylor. Towards extending the SWITCH platform for time-critical, cloud-based CUDA applications: Job scheduling parameters influencing performance. Future Generation Computer Systems, 100(??):542–556, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18311014.

Kegel:2013:DTU

[KSG13] Philipp Kegel, Michel Steuwer, and Sergei Gorlatch. dOpenCL: Towards uniform programming of distributed heterogeneous multi-/many-core systems. Journal of Parallel and Distributed Computing, 73(12):1639–1648, December 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0743731513001597.

Kusano:2001:OOC

[KSHS01] Kazuhiro Kusano, Mitsuhisa Sato, Takeo Hosomi, and Yoshiki Seo. The Omni OpenMP compiler on the distributed shared memory of Cenju-4. Lecture Notes in Computer Science, 2104:20–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2104/21040020.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2104/21040020.
                                        pdf.

Katkere:1995:VBW

[KSJ95] A. Katkere, J. Schlenzig, and R. Jain. VRML-Based WWW interface to MPI video. In Nadeau and Moreland [NM95], pages 25–31,

137. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Katkere:1996:VWI

[KSJ96] A. Katkere, J. Schlenzig, and R. Jain. VRMLbased WWW interface to MPI video. In ACM [ACM96a], pages 25–31, 137. ISBN 0-89791-818-5. LCCN

                                        ????  URL http://www.
                                        acm.org/pubs/contents/
                                        proceedings/graph/217306/

.

Kim:2014:VVF

[KSJ14] Young-Joo Kim, Sejun Song, and Yong-Kee Jun. VORD: A versatile on-the-fly race detection tool in OpenMP programs. International Journal of Parallel Programming, 42(6):900–930, December 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10766-013-0257-6.

Kim:2012:OUP

[KSL^{+}12] Jungwon Kim, Sangmin Seo, Jun Lee, Jeongho Nah, Gangwon Jo, and Jaejin Lee. OpenCL as a unified programming model for heterogeneous CPU/GPU clusters. ACM SIGPLAN Notices, 47 (8):299–300, August 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Kusano:2000:PEO

[KSS00] Kazuhiro Kusano, Shigehisa Satoh, and Mitsuhisa Sato. Performance evaluation of the omni OpenMP compiler. Lecture Notes in Computer Science, 1940:403–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1940/19400403.htm;

                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1940/19400403.
                                        pdf.

Kotsifakou:2018:HHP

[KSS^{+}18] Maria Kotsifakou, Prakalp Srivastava, Matthew D. Sinclair, Rakesh Komuravelli, Vikram Adve, and Sarita Adve. HPVM: heterogeneous parallel virtual machine. ACM SIGPLAN Notices, 53(1):68–80, January 2018. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kurzyniec:2007:UCA

[KSSS07] Dawid Kurzyniec, Magdalena Slawińska, Jaroslaw Slawiński, and Vaidy Sunderam. Unibus: a contrarian approach to Grid computing. The Journal of Supercomputing, 42 (1):125–144, October 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0920-8542&volume=
                                        42&issue=1&spage=125.

Kirkham:2020:FEM

[KSTM20] Jake Kirkham, Tyler Sorensen, Esin Tureci, and Margaret Martonosi. Foundations of empirical memory consistency testing. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA): 226:1–226:29, November

2020.  URL https://
dl.acm.org/doi/10.1145/
3428294.

Kranzlmuller:2001:IRM

[KSV01] Dieter Kranzlmüller, Christian Schaubschläger, and Jens Volkert. An integrated record&replay mechanism for nondeterministic message passing programs. Lecture Notes in Computer Science, 2131:192–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310192.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310192.
pdf.

Keppens:2002:OPM

[KT02] R. Keppens and G. Tóth.

OpenMP parallelism for multi-dimensional grid-adaptive magnetohydrodynamic simulations. Lecture Notes in Computer Science, 2329: 940–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2329/23290940.htm;
 http://link.springer-
ny.com/link/service/series/

                                        0558/papers/2329/23290940.
                                        pdf.

Koval:2010:USB

[KT10] Peter Koval and J. D. Talman. Update of spherical Bessel transform: FFTW and OpenMP. Computer Physics Communications, 181(12):2212–2213, December 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0010465510003188.

Kang:2019:SAM

[KTAB^{+}19] Qiao Kang, Jesper Larsson Träff, Reda Al-Bahrani, Ankit Agrawal, Alok Choudhary, and Wei keng Liao. Scalable algorithms for MPI intergroup Allgather and Allgatherv. Parallel Computing, 85(??):220–230, July 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S016781911830320X.

Karonis:2003:MGG

[KTF03] Nicholas T. Karonis, Brian Toonen, and Ian Foster. MPICH-G2: a Grid-enabled implementation of the Message Passing Interface. Journal of Parallel and Distributed Computing, 63(5): 551–563, May 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Komatitsch:2003:BDF

[KTJT03] Dimitri Komatitsch, Seiji Tsuboi, Chen Ji, and Jeroen Tromp. A 14.6 billion degrees of freedom, 5 teraflops, 2.5 terabyte earthquake simulation on the Earth Simulator. In ACM [ACM03], page ?? ISBN 1-58113-695-

1. LCCN ???? URL http:/

/www.sc-conference.org/
sc2003/inter_cal/inter_
cal_detail.php?eventid=
10711#1;   http://www.
sc-conference.org/sc2003/
paperpdfs/pap124.pdf.

Keppens:2021:MAP

[KTXP21] Rony Keppens, Jannis Teunissen, Chun Xia, and Oliver Porth. MPI-AMRVAC: a parallel, grid-adaptive PDE toolkit. Computers and Mathematics with Applications, 81(??):316–333, January 1, 2021. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122120301279.

Kuhn:1998:FFW

[Kuh98] Bob Kuhn. Fortran Futures: Workshop: OpenMP for parallel Fortran applications. ACM Fortran Forum, 17(3):22, December 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Kumar:1994:PPI

[Kum94] V. K. Prasanna Kumar, editor. Parallel processing: 1st IWWP: proceedings of the First International Workshop on Parallel Processing (IWPP-94), December 26– 31, 1994, Bangalore, India. Tata McGraw-Hill Pub. Co, New Delhi, India, 1994. ISBN 0-07-462332-X. LCCN QA 76.58 I587 1994.

Kranzlmueller:1998:DPP

[KV98] D. Kranzlmueller and J. Volkert. Debugging point-to-point communication in MPI and PVM. Lecture Notes in Computer Science, 1497: 265–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolonias:2011:DIE

[KVGH11] Vasileios Kolonias, Artemios G. Voyiatzis, George Goulas, and Efthymios Housos. Design and implementation of an efficient integer count sort in CUDA GPUs. Concurrency and Computation: Practice and Experience, 23(18):2365–2381, December 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Krotz-Vogel:1997:PPP

[KVH97] W. Krotz-Vogel and H.- C. Hoppe. The PALLAS parallel programming environment. Lecture Notes in Computer Science, 1332: 257–266, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamal:2014:IFG

[KW14] Humaira Kamal and Alan Wagner. An integrated finegrain runtime system for MPI. Computing, 96(4):293– 309, April 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00607-013-0329-x.

Korch:2020:ILE

[KW20] Matthias Korch and Tim Werner. Improving locality of explicit one-step methods on GPUs by tiling across stages and time steps. Future Generation Computer Systems, 102(??):889– 901, January 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167739X19307186.

Kamburugamuve:2018:AML

[KWEF18] Supun Kamburugamuve, Pulasthi Wickramasinghe, Saliya Ekanayake, and Geoffrey C. Fox. Anatomy of machine learning algorithm implementations in MPI, Spark, and Flink. The International Journal of High Performance Computing Applications, 32(1):61– 73, January 2018. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Kamal:2010:EIN

[KY10] A. A. Kamal and A. M. Youssef. Enhanced implementation of the NTRUEncrypt algorithm using graphics cards. In Chaudhuri et al. [CGB^{+}10], pages 168–174. ISBN 1-4244-7675-5. LCCN ????

Karwande:2003:CMC

[KYL03] Amit Karwande, Xin Yuan, and David K. Lowenthal. CC–MPI: a compiled communication capable MPI prototype for Ethernet switched clusters. ACM SIGPLAN Notices, 38 (10):95–106, October 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Karwande:2005:MPC

[KYL05] Amit Karwande, Xin Yuan, and David K. Lowenthal. An MPI prototype for compiled communication on Ethernet switched clusters. Journal of Parallel and Distributed Computing, 65(10):

1123–1133, October 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Krantz:1996:RFP

[KZCS96] A. T. Krantz, A. Zadroga, S. E. Chodrow, and V. S. Sunderam. An RPC facility for PVM. In Liddell et al. [LCHS96], pages 798– ?? ISBN 3-540-61142-8 (paperback). LCCN QA76.88 .H52 1996.

Lopez:2002:ESM

[LA02] Félix César García López and Nieves Luz Frías Arrocha. Expanding the synchronization model for OpenMP. Parallel and Distributed Computing Practices, 5(2):169–175, June 2002. CODEN ???? ISSN 1097-2803.

Lopez:2006:ESM

[LA06] F. C. García López and N. L. Frías Arrocha. An efficient synchronization model for OpenMP. Journal of Parallel and Distributed Computing, 66(11):1359– 1365, November 2006. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ladd:2004:GPP

[Lad04] Scott Ladd. Guide to Parallel Programming. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 0-387-40577-1. 465 (est.) pp. LCCN ???? Includes CD-ROM.

Lobeiras:2016:DEI

[LAD16] Jacobo Lobeiras, Margarita Amor, and Ramon Doallo. Designing efficient indexdigit algorithms for CUDA GPU architectures. IEEE Transactions on Parallel and Distributed Systems, 27 (5):1331–1343, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07138631-
abs.html.

Laguna:2015:DPF

[LAdS^{+}15] Ignacio Laguna, Dong H. Ahn, Bronis R. de Supinski, Saurabh Bagchi, and Todd Gamblin. Diagnosis of performance faults in LargeScale MPI applications via probabilistic progressdependence inference. IEEE Transactions on Parallel and Distributed Systems, 26 (5):1280–1289, May 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06803050-
abs.html.

Laforenza:2001:PHP

[Laf01] Domenico Laforenza. Programming high performance applications in grid environments. Lecture Notes in Computer Science, 2131: 8–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310008.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310008.
pdf.

Lorentz:2015:AMS

[LAFA15] Istvan Lorentz, Razvan Andonie, and Levente Fabry-Asztalos. Accelerating molecular structure determination based on inter-atomic distances using OpenCL. IEEE Transactions on Parallel and Distributed Systems, 26(12):3250–3263, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/06995963-
abs.html.

Langdon:2009:FHQ

[Lan09] W. B. Langdon. A fast high quality pseudo random number generator for nVidia CUDA. In Franz Rothlauf, editor, GECCO ’09 Proceedings of the 11th Annual Conference Companion on Genetic and Evolutionary Computation Conference: Late Breaking Papers, pages 2511–2513. ACM Press, New York, NY 10036, USA, 2009. ISBN 1-60558-505-X. LCCN ???? URL http://

                                        www.cs.ucl.ac.uk/staff/
                                        W.Langdon/ftp/gp-code/
                                        random-numbers/cuda_park-
                                        miller.tar.gz.

Loos:1996:MPS

[LB96] T. Loos and R. Bramley. MPI performance on the SGI Power Challenge. In IEEE [IEE96i], pages 203– 206. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Lavi:1998:IPD

[LB98] R. Lavi and A. Barak. Improving the PVM daemon network performance by direct network access. Lecture Notes in Computer Science, 1497:44–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lashgar:2016:ESM

[LB16] Ahmad Lashgar and Amirali Baniasadi. Employing software-managed caches in OpenACC: Opportunities and benefits. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 1(1): 2:1–2:34, March 2016. CO-DEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http:

                                        //dl.acm.org/citation.
                                        cfm?id=2798724.

Loncar:2016:CPS

[LBB^{+}16] Vladimir Loncar, Antun Balaz, Aleksandar Bogojević, Srdjan Skrbić, Paulsamy Muruganandam, and Sadhan K. Adhikari. CUDA programs for solving the time-dependent dipolar Gross–Pitaevskii equation in an anisotropic trap. Computer Physics Communications, 200(??):406– 410, March 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465515004361.

Losada:2019:LRR

[LBB^{+}19] Nuria Losada, George Bosilca, Aurélien Bouteiller, Patricia González, and María J. Martín. Local rollback for resilient MPI applications with application-level checkpointing and message logging. Future Generation Computer Systems, 91(??):450– 464, February 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0167739X18303443.

Liu:2021:BMN

[LBB^{+}21] Feilong Liu, Claude Barthels, Spyros Blanas, Hideaki Kimura, and Garret Swart. Beyond MPI: New communication interfaces for database systems and dataintensive applications. SIG-MOD Record (ACM Special Interest Group on Management of Data), 49(4):12– 17, March 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://

dl.acm.org/doi/10.1145/
3456859.3456862.

Lawton:1996:BHP

[LBD^{+}96] J. V. Lawton, J. J. Brosnan, M. P. Doyle, S. D. O. Riordain, and T. G. Reddin. Building a high-performance message-passing system for MEMORY CHANNEL clusters. Digital Technical Journal of Digital Equipment Corporation, 8(2):96–116, October 1996. CODEN DTJOEL. ISSN 0898-901X.

URL http://www.digital.
com:80/DTJM08/DTJM08P8.
PS.

Larrea:2020:EPM

[LBG^{+}20] Verónica G. Vergara Larrea, Reuben D. Budiardja, Rahulkumar Gayatri, Christopher Daley, Oscar Hernandez, and Wayne Joubert. Experiences in porting mini-applications to OpenACC and OpenMP on heterogeneous systems. Concurrency and Computation: Practice and Experience, 32 (20):e5780:1–e5780:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ling:2012:HPP

[LBH12] Cheng Ling, Khaled Benkrid, and Tsuyoshi Hamada. High performance phylogenetic analysis on CUDAcompatible GPUs. ACM SIGARCH Computer Architecture News, 40(5):52–57, December 2012. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). HEART ’12 conference proceedings.

Lewis:1993:PCP

[LC93] M. J. Lewis and R. E.

Cline, Jr. PVM communication performance in a switched FDDI heterogeneous distributed computing environment. In Bhargava [Bha93], pages 13–19. ISBN 0-8186-5250-0, 0-8186-5251-9. LCCN QA76.58.I444 1993.

Lauria:1997:MFH

[LC97a] Mario Lauria and Andrew Chien. MPI-FM: High performance MPI on workstation clusters. Journal of Parallel and Distributed Computing, 40(1): 4–18, January 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.idealibrary.com/
links/doi/10.1006/jpdc.
1996.1264/production;

http://www.idealibrary.
com/links/doi/10.1006/
jpdc.1996.1264/production/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/
jpdc.1996.1264/production/
ref.

Luecke:1997:HPF

[LC97b] G. R. Luecke and J. J. Coyle. High Performance Fortran versus explicit message passing on the IBM SP-2 for the parallel LU, QR, and Cholesky factorizations. Supercomputer, 13(2): 4–14, ???? 1997. CODEN SPCOEL. ISSN 0168-7875.

Li:2007:DIV

[LC07] Kuan-Ching Li and Hsun-Chang Chang. The design and implementation of visual performance monitoring and analysis toolkit for cluster and Grid environments. The Journal of Supercomputing, 40(3):299–317, June 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
40&issue=3&spage=299.

Luecke:2003:MCT

[LCC^{+}03] Glenn Luecke, Hua Chen, James Coyle, Jim Hoekstra, Marina Kraeva, and Yan Zou. MPI-CHECK: a tool for checking Fortran 90 MPI programs. Concurrency and Computation:

Practice and Experience, 15 (2):93–100, February 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:CDC

[LCH^{+}22] Wentao Li, Zhiwen Chen, Xin He, Guoyun Duan, Jianhua Sun, and Hao Chen. CVFuzz: Detecting complexity vulnerabilities in OpenCL kernels via automated pathological input generation. Future Generation Computer Systems, 127(??):384–395, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003526.

Liddell:1996:HPC

[LCHS96] Heather Mary Liddell, A. Colbrook, B. Hertzberger, and P. Sloot, editors. Highperformance computing and networking: international conference and exhibition, HPCN EUROPE 1966, Brussels, Belgium, April 15– 19, 1996: proceedings, volume 1067 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61142-8 (paperback). LCCN QA76.88 .H52 1996.

Lathrop:2011:SPI

[LCK11] Scott Lathrop, Jim Costa, and William Kramer, editors. SC’11: Proceedings of 2011 International Conference for High Performance Computing, Networking, Storage and Analysis, Seattle, WA, November 12–18 2011. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4503-0771-X. LCCN ????

Lashuk:2012:MPA

[LCL^{+}12] Ilya Lashuk, Aparna Chandramowlishwaran, Harper Langston, Tuan-Anh Nguyen, Rahul Sampath, Aashay Shringarpure, Richard Vuduc, Lexing Ying, Denis Zorin, and George Biros. A massively parallel adaptive fast multipole method on heterogeneous architectures. Communications of the ACM, 55(5):101–109, May 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Losada:2017:RMA

[LCMG17] Nuria Losada, Iván Cores, María J. Martín, and Patricia González. Resilient MPI applications using an application-level checkpointing framework and ULFM. The Journal of Supercomputing, 73(1):100–113, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lonsdale:1994:CRP

[LCVD94a] G. Lonsdale, J. Clinckemaillie, S. Vlachoutsis, and J. Dubois. Communication requirements in parallel crashworthiness simulation. In Gentzsch and Harms [GH94], pages 55– 61. ISBN 0-387-57981-8 (New York), 3-540-57981-8 (Berlin). LCCN QA76.88.I57 1994. DM96.00. Two volumes.

Lonsdale:1994:CMH

[LCVD94b] G. Lonsdale, J. Clinckemaillie, S. Vlachoutsis, and J. Dubois. Crash-simulation migration to HPC systems. In Dekker et al. [DSZ94], pages 439–446. ISBN 0-444-81784-0. LCCN QA76.58.E98 1994.

Liu:2003:PCM

[LCW^{+}03] Jiuxing Liu, Balasubramanian Chandrasekaran, Jiesheng Wu, Weihang Jiang, Sushmitha Kini, Weikuan Yu, Darius Buntinas, Pete Wyckoff, and D. K. Panda. Performance comparison of MPI implementations over InfiniBand, Myrinet and Quadrics. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN ???? URL http:/

/www.sc-conference.org/

sc2003/inter_cal/inter_
cal_detail.php?eventid=
10696#0;   http://www.
sc-conference.org/sc2003/
paperpdfs/pap310.pdf.

Liu:1996:BMP

[LCY96] L. T. Liu, D. E. Culler, and C. Yoshikawa. Benchmarking message passing performance using MPI. In Reeves [Ree96], pages 101– 110. ISBN 0-8186-7623-X. LCCN QA76.58 .I34 1996. Three volumes.

Liu:2019:MML

[LCY19] Qixiao Liu, Zhifeng Chen, and Zhibin Yu. MiC: Multi-level characterization and optimization of GPGPU kernels. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(3):25:1– 25:??, June 2019. CO-DEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3304108.

Lee:2001:APT

[LD01] D. J. Lee and T. J. Downar. The application of POSIX threads and OpenMP to the U.S. NRC neutron kinetics code PARCS. Lecture Notes in Computer Science, 2104:90–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/

bibs/2104/21040090.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2104/21040090.
pdf.

Lu:1997:QPD

[LDCZ97] Honghui Lu, Sandhya Dwarkadas, Alan L. Cox, and Willy Zwaenepoel. Quantifying the performance differences between PVM and Tread-Marks. Journal of Parallel and Distributed Computing, 43(2):65–78, June 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.idealibrary.com/
links/doi/10.1006/jpdc.
1997.1332/production;
http://www.idealibrary.
com/links/doi/10.1006/
jpdc.1997.1332/production/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/
jpdc.1997.1332/production/
ref.

Liu:2013:DLO

[LDJK13] Jun Liu, Wei Ding, Ohyoung Jang, and Mahmut Kandemir. Data layout optimization for GPGPU architectures. ACM SIG-PLAN Notices, 48(8):283– 284, August 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’13 Conference proceedings.

Lorenzon:2019:ASO

[LdSB19] A. F. Lorenzon, C. C. de Oliveira, J. D. Souza, and A. C. S. Beck. Aurora: Seamless optimization of OpenMP applications. IEEE Transactions on Parallel and Distributed Systems, 30 (5):1007–1021, May 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2006:PT

[Lee06] Edward A. Lee. The problem with threads. Computer, 39 (5):33–42, May 2006. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:2012:SMO

[Lee12] Jaejin Lee. SnuCL and an MPI + OpenCL implementation of HPL on heterogeneous CPU/GPU clusters. In ????, editor, ATIP ’12: Proceedings of the ATIP/ A*CRC Workshop on Accelerator Technologies for High-Performance Computing: Does Asia Lead the Way?, page ?? ACM Press, New York, NY 10036, USA, 2012. ISBN 1-4503-1644-1. LCCN ????

Levelt:1995:IIS

[Lev95] A. H. M. Levelt, editor. ISSAC ’95: International symposium on symbolic and algebraic computation — July 10–12, 1995, Montréal, Canada, ISSAC — Proceedings. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995.

Law:1993:EDM

[LF^{+}93a] K. H. Law, R. E. Fulton, et al., editors. Engineering data management: key to success in a global market: proceedings of the 1993 ASME International Computers in Engineering Conference and Exposition, August 8–12, San Diego, California, COM-PUTERS IN ENGINEER-ING VOL COM. American Society Mech. Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, USA, 1993. ISBN 0-7918-1169-7. LCCN TA345.A86 1993.

Levesque:1993:SAA

[LF93b] J. M. Levesque and R. Friedman. The state of the art in automatic parallelisation. In Anonymous [Ano93g], pages 95–107. ISBN ???? LCCN ????

Lim:2011:ATC

[LFL11] Min Yeol Lim, Vincent W.

Freeh, and David K. Lowenthal. Adaptive, transparent CPU scaling algorithms leveraging internode MPI communication regions. Parallel Computing, 37(10–11):667–683, October/November 2011. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819111000871.

Leon:1992:FP

[LFS92] Juan Leon, Allan L. Fisher, and Peter Steenkiste. Failsafe PVM. In SCRI WCC’92 [SCR92], page ?? ISBN ???? LCCN ???? Proceedings available via anonymous

                                        ftp from ftp.scri.fsu.edu
                                        in directory pub/parallel-
                                        workshop.92.

Leon:1993:FPA

[LFS93a] J. Leon, A. L. Fisher, and P. Steenkiste. Failsafe PVM: a portable package for distributed programming with transparent recovery. Technical Report CMU-CS-93-124, Carnegie-Mellon University, Department of Computer Science, 1993.

Leon:1993:FPP

[LFS93b] Juan Leon, Allan L. Fisher, and Peter Alfons Steenkiste. Fail-safe PVM: a portable package for distributed programming with transparent recovery. Technical report, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, 1993. 22 pp.

Levy:2019:USE

[LFS^{+}19] Scott Levy, Kurt B. Ferreira, Whit Schonbein, Ryan E.

Grant, and Matthew G. F. Dosanjh. Using simulation to examine the effect of MPI message matching costs on application performance. Parallel Computing, 84(??):63–74, May 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819118303272.

Levy:2020:UVA

[LFW20] Scott Levy, Kurt B. Ferreira, and Patrick Widener. The unexpected virtue of almost: Exploiting MPI collective operations to approximately coordinate checkpoints. Concurrency and Computation: Practice and Experience, 32(3):e4890:1– e4890:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Loyot:1993:VVM

[LG93] E. C. Loyot, Jr. and A. S.

Grimshaw. VMPP: a virtual machine for parallel processing. In IEEE [IEE93b], pages 735–740. ISBN 0-8186-3442-

1. LCCN QA 76.58 I56 1993. IEEE catalog no. 93TH0513-2.

Lee:1999:PEJ

[LGCH99] Bu-Sung Lee, Yan Gu, Wentong Cai, and Alfred Heng. Performance evaluation of JPVM. Parallel Processing Letters, 9(3):401–??, September 1999. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Liu:2016:MBM

[LGG16] Weifeng Liu, Michael Gerndt, and Bin Gong. Modelbased MPI-IO tuning with Periscope tuning framework. Concurrency and Computation: Practice and Experience, 28(1):3–20, January 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2010:SVC

[LGKQ10] Guodong Li, Ganesh Gopalakrishnan, Robert M. Kirby, and Dan Quinlan. A symbolic verifier for CUDA programs. ACM SIGPLAN Notices, 45(5):357–358, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lassous:2000:HGA

[LGM00] Isabelle Guérin Lassous, Jens Gustedt, and Michel Morvan. Handling graphs according to a coarse grained approach: Experiments with PVM and MPI. Lecture Notes in Computer Science, 1908:72–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/

bibs/1908/19080072.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080072.
pdf.

Losada:2020:FTM

[LGM^{+}20] Nuria Losada, Patricia González, María J. Martín, George Bosilca, Aurélien Bouteiller, and Keita Teranishi. Fault tolerance of MPI applications in exascale systems: the ULFM solution. Future Generation Computer Systems, 106(??): 467–481, May 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167739X1930860X.

Lopez-Gomez:2019:ESP

[LGMdRA^{+}19] Javier López-Gómez, Javier Fernández Muñoz, David del Rio Astorga, Manuel F. Dolz, and J. Daniel Garcia. Exploring stream parallel patterns in distributed MPI environments. Parallel Computing, 84(??):24–36, May 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819118303442.

Leung:1995:EPE

[LH95] K.-C. Leung and M. Hamdi.

Evaluating PVM and Express on various network clusters. In Alnuweiri and Hamdi [AH95], pages 57–66. ISBN 0-8186-7124-6. LCCN TK5105.5 .H56 1995.

Leung:1998:PAN

[LH98] Ka-Cheong Leung and Mounir Hamdi. Performance assessment of network protocols and parallel programming tools for distributed computing systems. International Journal of Computer Systems Science and Engineering, 13(1):67–80, January 1998. CODEN CSSEEI. ISSN 0267-6192.

Liao:2007:OOP

[LHC^{+}07] Chunhua Liao, Oscar Hernandez, Barbara Chapman, Wenguang Chen, and Weimin Zheng. OpenUH: an optimizing, portable OpenMP compiler. Concurrency and Computation: Practice and Experience, 19(18):2317–2332, December 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:1996:TSF

[LHCT96] Bu-Sung Lee, A. Heng, W. Cai, and Tai-Ann Tan. Task scheduling facility for PVM. Parallel Processing Letters, 6(4):563–574, December 1996. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Liu:2005:EIO

[LHCW05] Z. Liu, L. Huang, B. Chapman, and T. Weng. Efficient implementation of OpenMP for clusters with implicit data distribution. Lecture Notes in Computer Science, 3349:121–??, 2005.

Lin:1994:DNC

[LHD^{+}94] Mengjou Lin, Jehwei Hsieh, D. H. C. Du, J. P. Thomas, and J. A. MacDonald. Distributed network computing over local ATM networks. In IEEE [IEE94h], pages 154– 163. ISBN 0-8186-6607-2, 0-8186-6605-6, 0-8186-6606-4. ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Lin:1995:DNC

[LHD^{+}95] Mengjou Lin, J. Hsieh, D. H. C. Du, J. P. Thomas, and J. A. MacDonald. Distributed network computing over local ATM networks. IEEE Journal on Selected Areas in Communications, 13(4):733–748, May 1995. CODEN ISACEM. ISSN 0733-8716 (print), 1558-0008 (electronic).

Li:1996:PSI

[LHHM96] G.-J. Li, D. F. Hsu, S. Horiguchi, and B. Maggs, editors. Proceedings. Second International Symposium on Parallel Architectures, Algorithms, and Networks (I-SPAN ’96): June 12–14, 1996, Beijing, China. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7460-1. LCCN QA76.58.I5673 1996. IEEE catalog number 96TB100044.

Liu:2010:RTC

[LHLK10] Fuchang Liu, Takahiro Harada, Youngeun Lee, and Young J. Kim. Real-time collision culling of a million bodies on graphics processing units. ACM Transactions on Graphics, 29 (6):154:1–154:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:1997:PIO

[LHZ97] Wei Li, Xiaohu Huang, and Nanning Zheng. Parallel implementing OpenGL on PVM. Parallel Computing, 23(12):1839–1850, December 15, 1997. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1997&volume=23&
issue=12&aid=1248.

Lu:1998:ONW

[LHZ98] Honghui Lu, Y. Charlie Hu, and Willy Zwaenepoel. OpenMP on networks of workstations. In ACM [ACM98b], page ?? ISBN ???? LCCN ???? URL

http://www.supercomp.org/
sc98/TechPapers/sc98_FullAbstracts/
Lu1105/index.htm.

Liang:2020:AMD

[LHZ^{+}20] Jianguo Liang, Rong Hua, Hao Zhang, Wenqiang Zhu, and You Fu. Accelerated molecular dynamics simulation of silicon crystals on TaihuLight using OpenACC. Parallel Computing, 99(??):Article 102667, November 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167819120300600.

Li:1996:SIS

[Li96] Guo-Jie Li, editor. Second International Symposium on Parallel Architectures, Algorithms, and Networks (I-SPAN ’96): proceedings, June 12–14, 1996, Beijing, China. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7460-

1. LCCN QA76.58.I565 1996. IEEE catalog number 94TH0697-3.

Liu:1995:WCD

[Liu95] Xiaomao Liu. Workstations cluster for distributed supercomputing. Mini-Micro Systems, 16(2):45–52, February 1995. CODEN XWJXEH. ISSN 1000-1220.

Livny:2000:MYW

[Liv00]  Miron Livny.  Manag-

ing your workforce on a computational grid. Lecture Notes in Computer Science, 1908:3–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1908/19080003.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1908/19080003.
                                        pdf.

Lastovetsky:2010:RAP

[LK10] Alexey Lastovetsky and Tahar Kechadi. Recent advances in Parallel Virtual Machine and Message Passing Interface. The International Journal of High Performance Computing Applications, 24(1):3–4, February 2010. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

                                        tronic). URL http://hpc.
                                        sagepub.com/content/24/
                                        1/3.full.pdf+html.

LaSalle:2014:MBD

[LK14] Dominique LaSalle and George Karypis. MPI for big data: New tricks for an old dog. Parallel Computing, 40(10):754–767, December 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167819114000830.

Li:2020:SLF

[LK20] Qinbo Li and Nima Khademi Kalantari. Synthesizing light field from a single image with variable MPI and two network fusion. ACM Transactions on Graphics, 39(6):229:1–229:10, November 2020. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https://

dl.acm.org/doi/10.1145/
3414685.3417785.

Lastovetsky:2008:RAP

[LKD08] Alexey Lastovetsky, Tahar Kechadi, and Jack Dongarra, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 15th European PVM/MPI Users’ Group Meeting, Dublin, Ireland, September 7–10, 2008. Proceedings, volume 5205 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CO-DEN LNCSD9. ISBN 3-540-87474-7 (print), 3-540-87475-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:/

/www.springerlink.com/

                                        content/978-3-540-87475-
                                        1.

Luecke:2003:CPM

[LKJ03] Glenn R. Luecke, Marina Kraeva, and Lili Ju. Comparing the performance of MPICH with Cray’s MPI and with SGI’s MPI. Concurrency and Computation: Practice and Experience, 15(9):779–802, August 10, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liang:1996:AEO

[LKL96] Wen-Yew Liang, Chun-Ta King, and Feipei Lai. Adsmith: an efficient objectbased distributed shared memory system on PVM. In Li [Li96]. ISBN 0-8186-7460-1. LCCN QA76.58.I565 1996. IEEE catalog number 94TH0697-3.

Li:2003:PNH

[LkLC^{+}03] Jianwei Li, Wei keng Liao, Alok Choudhary, Robert Ross, Rajeev Thakur, William Gropp, Rob Latham, Andrew Siegel, Brad Gallagher, and Michael Zingale. Parallel netCDF: a highperformance scientific I/O interface. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN ???? URL http:/

                                        /www.sc-conference.org/
                                        sc2003/inter_cal/inter_
                                        cal_detail.php?eventid=
                                        10722#1;   http://www.

sc-conference.org/sc2003/
paperpdfs/pap258.pdf.

Luecke:2004:PSM

[LKYS04] Glenn R. Luecke, Marina Kraeva, Jing Yuan, and Silvia Spanoyannis. Performance and scalability of MPI on PC clusters. Concurrency and Computation: Practice and Experience, 16 (1):79–107, January 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ludwig:1995:PPF

[LL95] T. Ludwig and S. Lamberts. PFSLib — a parallel file system for workstation clusters. In Malyshkin [Mal95], pages 246–251. ISBN 3-540-60222-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I547 1995.

Luecke:2001:SPO

[LL01] Glenn R. Luecke and Wei-Hua Lin. Scalability and performance of OpenMP and MPI on a 128-processor SGI Origin 2000. Concurrency and Computation: Practice and Experience, 13(10):905–928, August 25, 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/85007180/
START;   http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=85007180&PLACEBO=IE.
pdf.

Lin:2016:VDF

[LL16] Yu-Te Lin and Jenq-Kuen Lee. Vector data flow analysis for SIMD optimizations on OpenCL programs. Concurrency and Computation: Practice and Experience, 28(5):1629–1654, April 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2013:COM

[LLC13] Hung-Fu Li, Tyng-Yeu Liang, and Jun-Yao Chiu. A compound OpenMP/ MPI program development toolkit for hybrid CPU/ GPU clusters. The Journal of Supercomputing, 66 (1):381–405, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

link.springer.com/article/
10.1007/s11227-013-0912-
0.

Lidbury:2015:MCC

[LLCD15] Christopher Lidbury, Andrei Lascu, Nathan Chong, and Alastair F. Donaldson. Many-core compiler fuzzing. ACM SIGPLAN Notices, 50(6):65–76, June 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2012:PFA

[LLG12] Peng Li, Guodong Li, and Ganesh Gopalakrishnan. Parametric flows: automated behavior equivalencing for symbolic analysis of races in CUDA programs. In Hollingsworth [Hol12], pages 29:1–29:?? ISBN 1-4673-0804-8. URL http:

//conferences.computer.
org/sc/2012/papers/1000a009.
pdf.

Luo:2014:ISM

[LLH^{+}14] Miao Luo, Xiaoyi Lu, Khaled Hamidouche, Krishna Kandalla, and Dhabaleswar K. Panda. Initial study of multi-endpoint runtime for MPI + OpenMP hybrid programming model on multi-core systems. ACM SIGPLAN Notices, 49(8): 395–396, August 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lyu:2021:FFA

[lLLmH^{+}21] Xing long Lyu, Tiexiang Li, Tsung ming Huang, Jia wei Lin, Wen wei Lin, and Sheng Wang. FAME: Fast algorithms for Maxwell’s equations for three-dimensional photonic crystals. ACM Transactions on Mathematical Software, 47(3):26:1– 26:24, June 2021. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3446329.

Langlais:2002:SSM

[LLRS02] M. Langlais, G. Latu, J. Roman, and P. Silan. Stochastic simulation of a marine host-parasite system using a hybrid MPI/ OpenMP programming. Lecture Notes in Computer Science, 2400:436–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2400/24000436.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2400/24000436.
                                        pdf.

Lambert:2021:OOFa

[LLVM21a] Jacob Lambert, Seyong Lee, Jeffrey S. Vetter, and Allen D. Malony. Optimization with the OpenACC-to-FPGA framework on the Arria 10 and Stratix 10 FP-GAs. Parallel Computing, 104–105(??):??, July 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819121000417.

Lambert:2021:OOFb

[LLVM21b] Jacob Lambert, Seyong Lee, Jeffrey S. Vetter, and Allen D. Malony. Optimization with the OpenACC-to-FPGA framework on the arria 10 and stratix 10 FP-GAs. Parallel Computing, 104–105(??):??, July 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819121000417.

Li:1993:SLL

[LLY93] Q. Li, J.-C. Liu, and T. G.

Yip. Solving large linear equations using PVM system. In Law et al. [LF^{+}93a], pages 685–690. ISBN 0-7918-1169-7. LCCN TA345.A86 1993.

Loh:1994:ISR

[LM94] B. C. Loh and G. A. Manson.

Incorporating software reuse into the PCSC methodology. In de Gloria et al. [dGJM94], pages 929–941. ISBN ???? LCCN ????

Larsen:1999:SPG

[LM99] M. Larsen and P. Madsen. A scalable parallel Gauss–Seidel and Jacobi solver for animal genetics. In Dongarra et al. [DLM99], pages 356–363. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Lu:2013:MLP

[LM13] Ligang Lu and Karen Magerlein. Multi-level parallel computing of reverse time migration for seismic imaging on Blue Gene/Q. ACM SIGPLAN Notices, 48 (8):291–292, August 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’13 Conference proceedings.

Lee:2009:OGC

[LME09] Seyong Lee, Seung-Jai Min, and Rudolf Eigenmann. OpenMP to GPGPU: a compiler framework for automatic translation and optimization. ACM SIG-PLAN Notices, 44(4):101– 110, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Losada:2017:ARV

[LMG17] Nuria Losada, María J. Martín, and Patricia González. Assessing resilient versus stop-and-restart fault-tolerant solutions in MPI applications. The Journal of Supercomputing, 73(1):316–329, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2019:TBH

[LML^{+}19] Bing Li, Mengjie Mao, Xiaoxiao Liu, Tao Liu, Zihao Liu, Wujie Wen, Yiran Chen, and Hai (Helen) Li. Thread batching for highperformance energy-efficient GPU memory design. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(4):39:1– 39:??, December 2019. CO-DEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3330152.

Lopez:2015:PBV

[LMM^{+}15] Hugo A. López, Eduardo R. B. Marques, Francisco Martins, Nicholas Ng, César Santos, Vasco Thudichum Vasconcelos, and Nobuko Yoshida. Protocol-based verification of message-passing parallel programs. ACM SIGPLAN Notices, 50(10): 280–298, October 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Losada:2014:EAL

[LMRG14] N. Losada, M. J. Martín, G. Rodríguez, and P. González. Extending an applicationlevel checkpointing tool to provide fault tolerance support to OpenMP applications. J.UCS: Journal of Universal Computer Science, 20(9):1351– ??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_9/
extending_an_application_
level.

Lee:2015:OPE

[LNK^{+}15] Joo Hwan Lee, Nimit Nigania, Hyesoon Kim, Kaushik Patel, and Hyojong Kim. OpenCL performance evaluation on modern multicore CPUs. Scientific Programming, 2015(??):859491:1– 859491:20, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL https://

                                        www.hindawi.com/journals/
                                        sp/2015/859491/.

Louca:2000:MFP

[LNLE00] S. Louca, N. Neophytou, A. Lachanas, and P. Evripidou. MPI-FT: Portable fault tolerance scheme for MPI. Parallel Processing Letters, 10(4):371–??, December 2000. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL http://

                                        ejournals.wspc.com.sg/
                                        ppl/10/1004/S0129626400000342.
                                        html.

Lima:2012:PEO

[LNW^{+}12] Antonio M. Lima, Marco A. S. Netto, Thais Webber, Ricardo M. Czekster, Cesar A. F. De Rose, and Paulo Fernandes. Performance evaluation of OpenMPbased algorithms for handling Kronecker descriptors. Journal of Parallel and Distributed Computing, 72(5):678–692, May 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731512000354.

Lu:1996:PIF

[LO96] E. J.-L. Lu and D. I. Okunbor. Parallel implementation of 3D FMA using MPI. In IEEE [IEE96i], pages 119–

124. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Labarta:2001:NOD

[LOHA01] J. Labarta, J. Oliver, D. S. Henty, and Eduard Ayguadé. New OpenMP directives for irregular data access loops. Scientific Programming, 9(2–3):175–183, Spring–Summer 2001. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C10%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Lou:1995:PIN

[Lou95] J. Z. Lou. A parallel incompressible Navier–Stokes solver with multigrid iterations. In Bailey et al. [BBG^{+}95], pages 167–168. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Landman:2000:PLR

[LP00] Joseph Landman and Piotr Piecuch. Parallelization of a legacy research program using OpenMP. ACM Fortran Forum, 19(2):16–23, August 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Li:2011:FSM

[LPD^{+}11] Guodong Li, Robert Palmer, Michael DeLisi, Ganesh Gopalakrishnan, and Robert M. Kirby. Formal specification of MPI 2.0: Case study in specifying a practical concurrent programming API. Science of Computer Programming, 76(2): 65–81, February 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lockey:1998:CRM

[LPJ98] P. Lockey, R. Proctor, and I. D. James. Characterization of I/O requirements in a massively parallel shelf sea model. The International Journal of High Performance Computing Applications, 12(3):320–332, Fall 1998. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:/

/journals.sagepub.com/
doi/pdf/10.1177/109434209801200302.

Li:2001:PCS

[LR01] Michael Na Li and A. J. Rossini. RPVM: Cluster statistical computing in R.

R News: the Newsletter of the R Project, 1(3):4– 7, September 2001. CO-DEN ???? ISSN 1609-

3631. URL http://CRAN.R-
project.org/doc/Rnews/.

Lastovetsky:2006:HTM

[LR06a] Alexey Lastovetsky and Ravi Reddy. HeteroMPI: Towards a message-passing library for heterogeneous networks of computers. Journal of Parallel and Distributed Computing, 66(2): 197–220, February 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Le:2006:DMC

[LR06b] Thuy T. Le and Jalel Rejeb.

A detailed MPI communication model for distributed systems. Future Generation Computer Systems, 22 (3):269–278, February 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lotfi:2015:AAC

[LRBG15] Atieh Lotfi, Abbas Rahimi, Luca Benini, and Rajesh K. Gupta. Aging-aware compilation for GP-GPUs. ACM Transactions on Architecture and Code Optimization, 12(2):24:1–24:??, July 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lee:2014:BCA

[LRG14] Changmin Lee, Won Woo Ro, and Jean-Luc Gaudiot. Boosting CUDA applications with CPU–GPU hybrid computing. International Journal of Parallel Programming, 42(2):384– 404, April 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10766-013-0252-y.

Laguna:2016:EEU

[LRG^{+}16] Ignacio Laguna, David F. Richards, Todd Gamblin, Martin Schulz, Bronis R. de Supinski, Kathryn Mohror, and Howard Pritchard. Evaluating and extending userlevel fault tolerance in MPI applications. The International Journal of High Performance Computing Applications, 30(3):305–319, 2016. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

//journals.sagepub.com/
doi/full/10.1177/1094342015623623.

Lima:2019:PEA

[LRLG19] João Vicente Ferreira Lima, Issam Raïs, Laurent Lefèvre, and Thierry Gautier. Performance and energy analysis of OpenMP runtime systems with dense linear algebra algorithms. The International Journal of High Performance Computing Applications, 33(3):431–443, May 1, 2019. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:/

/journals.sagepub.com/
doi/full/10.1177/1094342018792079.

Luo:2001:PDE

[LRQ01] Jun Luo, Sanguthevar Rajasekaran, and Chenxia Qiu. Parallizing 1-dimensional estuarine model. Lecture Notes in Computer Science, 2131:257–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310257.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310257.
pdf.

Latham:2007:IMI

[LRT07] Robert Latham, Robert Ross, and Rajeev Thakur. Implementing MPI-IO atomic mode and shared file pointers using MPI one-sided communication. The International Journal of High Performance Computing Applications, 21(2):132–143, May 2007. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/21/
2/132.full.pdf+html.

Li:2001:WMB

[LRW01] Maozhen Li, Omer F. Rana, and David W. Walker. Wrapping MPI-based legacy codes as Java/CORBA components. Future Generation Computer Systems, 18(2): 213–223, October 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/19/19/60/31/29/abstract.
                                        html.

Luckow:2008:MFT

[LS08] André Luckow and Bettina Schnor. Migol: a fault-tolerant service framework for MPI applications in the Grid. Future Generation Computer Systems, 24 (2):142–152, February 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lin:2010:TLS

[LS10] Paul T. Lin and John N. Shadid. Towards large-scale multi-socket, multicore parallel simulations: Performance of an MPI-only semiconductor device simulator. Journal of Computational Physics, 229(19):6804–6818, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0021999110002846.

Lashgar:2015:CSR

[LSB15] Ahmad Lashgar, Ebad Salehi, and Amirali Baniasadi. A case study in reverse engineering GPG-PUs: Outstanding memory handling resources. ACM SIGARCH Computer Architecture News, 43(4):15–21, September 2015. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Li:2018:CER

[LSB^{+}18] Xiangbo Li, Mohsen Amini Salehi, Magdy Bayoumi, Nian-Feng Tzeng, and Rajkumar Buyya. Cost-efficient and robust on-demand video transcoding using heterogeneous cloud services. IEEE Transactions on Parallel and Distributed Systems, 29 (3):556–571, ???? 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//ieeexplore.ieee.org/
document/8081853/.

Li:2020:OOS

[LSB^{+}20] Ting Li, Lawrence V. Stanislawski, Tyler Brockmeyer, Shaowen Wang, and Ethan Shavers. OpenCLC: an opensource software tool for similarity assessment of linear hydrographic features. SoftwareX, 11(??):Article 100401, January/June 2020. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711018302747.

Levesque:2012:HEA

[LSG12] John M. Levesque, Ramanan Sankaran, and Ray Grout. Hybridizing S3D into an exascale application using OpenACC: an approach for moving to multi-petaflops and beyond. In Hollingsworth [Hol12], pages 15:1–15:?? ISBN 1-4673-0804-8. URL http:

                                        //conferences.computer.
                                        org/sc/2012/papers/1000a040.
                                        pdf.

Luecke:2004:PSS

[LSK04] Glenn R. Luecke, Silvia Spanoyannis, and Marina Kraeva. The performance and scalability of SHMEM and MPI-2 one-sided routines on a SGI Origin 2000 and a Cray T3E-600. Concurrency and Computation: Practice and Experience, 16 (10):1037–1060, August 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2018:CHM

[LSM^{+}18] Han Lin, Zhichao Su, Xiandong Meng, Xu Jin, Zhong Wang, Wenting Han, Hong An, Mengxian Chi, and Zheng Wu. Combining Hadoop with MPI to solve metagenomics problems that are both dataand compute-intensive. International Journal of Parallel Programming, 46(4):762– 775, August 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Liu:2011:CBA

[LSMW11] Weiguo Liu, Bertil Schmidt, and Wolfgang Muller-Wittig. CUDA-BLASTP: Accelerating BLASTP on CUDAenabled graphics hardware. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 8(6): 1678–1684, November 2011. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Lumsdaine:1995:WIM

[LSR95] A. Lumsdaine, J. M. Squyres, and M. W. Reichelt. Waveform iterative methods for parallel solution of initial value problems. In IEEE [IEE95j], pages 88–97. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Li:2015:AMR

[LSSZ15] Jiansen Li, Jianqi Sun, Ying Song, and Jun Zhao. Accelerating MRI reconstruction via three-dimensional dual-dictionary learning using CUDA. The Journal of Supercomputing, 71(7):2381– 2396, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s11227-015-1386-z.

Liu:2008:AMD

[LSVMW08] Weiguo Liu, Bertil Schmidt, Gerrit Voss, and Wolfgang Müller-Wittig. Accelerating molecular dynamics simulations using graphics processing units with CUDA. Computer Physics Communications, 179(9):634–641, November 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465508002191.

Lazzarino:2002:PBP

[LSZL02] Oscar Lazzarino, Andrea Sanna, Claudio Zunino, and Fabrizio Lamberti. A PVMbased parallel implementation of the REYES image rendering architecture. Lecture Notes in Computer Science, 2474:165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

                                        /link.springer.de/link/
                                        service/series/0558/bibs/
                                        2474/24740165.htm;   http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2474/24740165.pdf.

Langr:2014:APP

[LTDD14] Daniel Langr, Pavel Tvrdík, Tomás Dytrych, and Jerry P. Draayer. Algorithm 947:

Paraperm — parallel generation of random permutations with MPI. ACM Transactions on Mathematical Software, 41(1):5:1–5:26, October 2014. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lazar:1994:SRE

[LTLC94] A. A. Lazar, K. H. Tseng, Koon Seng Lim, and W. Choe. A scalable and reusable emulator for evaluating the performance of SS7 networks. IEEE Journal on Selected Areas in Communications, 12(3):395–404, April 1994. CODEN ISACEM. ISSN 0733-8716 (print), 1558-0008 (electronic).

Laohawee:2000:PDT

[LTR00] P. Laohawee, A. Tangpong, and A. Rungsawang. Parallel DSIR text indexing system: Using multiple master/slave concept. Lecture Notes in Computer Science, 1908:297–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080297.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080297.
pdf.

Lee:2002:IPC

[LTRA02] Nung Kion Lee, David Taniar, J. Wenny Rahayu, and Mafruz Zaman Ashrafi. Implementation of parallel collection equi-join using MPI. Lecture Notes in Computer Science, 2367: 217–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2367/23670217.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2367/23670217.
                                        pdf.

Langr:2016:ASM

[LTS16] Daniel Langr, Pavel Tvrdik, and Ivan Simecek. AQsort: Scalable multi-array in-place sorting with OpenMP. Scalable Computing: Practice and Experience, 17(4):369– 391, ???? 2016. CO-DEN ???? ISSN 1895-

                                        1767.  URL https://
                                        www.scpe.org/index.php/
                                        scpe/article/view/1207.

Luo:1999:SMV

[Luo99] Yong Luo. Shared memory vs. message passing: The COMOPS benchmark experiment. The Journal of Supercomputing, 13(3): 283–301, May 1999. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
13&issue=3&spage=283;
http://www.wkap.nl/oasis.
htm/206582.

Lusk:2000:IIC

[Lus00]  Ewing Lusk.  Isolating

and interfacing the components of a parallel computing environment. Lecture Notes in Computer Science, 1908:5–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080005.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080005.
pdf.

Lee:2012:EED

[LV12] Seyong Lee and Jeffrey S.

Vetter. Early evaluation of directive-based GPU programming models for productive exascale computing. In Hollingsworth [Hol12], pages 23:1–23:?? ISBN 1-4673-0804-8. URL http:

//conferences.computer.
org/sc/2012/papers/1000a051.
pdf.

Liu:2004:BMI

[LVP04] Jiuxing Liu, Abhinav Vishnu, and Dhabaleswar K. Panda. Building multirail Infini-Band clusters: MPI-level design and performance evaluation. In ACM [ACM04], page 33. ISBN 0-7695-2153-

3. LCCN ????

Li:1995:CPP

[LW95] Liwei Li and Paul S. Wang. The CL-PVM package. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29 (3–4):2–8, December 1995. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ludwig:1997:OUI

[LW97] T. Ludwig and R. Wismueller. OMIS 2.0 — a universal interface for monitoring systems. Lecture Notes in Computer Science, 1332:267–276, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:2020:GTD

[LW20] Huanxin Lin and Cho-Li Wang. On-GPU threaddata remapping for nested branch divergence. Journal of Parallel and Distributed Computing, 139(??): 75–86, May 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0743731518308967.

Lee:2015:GCE

[LWKA15] J. Lee, D. H. Woo, H. Kim, and M. Azimi. GREEN cache: Exploiting the disciplined memory model of OpenCL on GPUs. IEEE Transactions on Computers, 64(11):3167–3180, November 2015. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2004:HPR

[LWP04] Jiuxing Liu, Jiesheng Wu, and Dhabaleswar K. Panda. High performance RDMAbased MPI implementation over InfiniBand. International Journal of Parallel Programming, 32(3): 167–198, June 2004. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
32&issue=3&spage=167.

Laguna:2019:GPD

[LWSB19] Ignacio Laguna, Paul C.

Wood, Ranvijay Singh, and Saurabh Bagchi. GPUMixer: Performance-driven floatingpoint tuning for GPU scientific applications. Report, Lawrence Livermore National Laboratory, Livermore CA 94550, USA, 2019. URL http://

lagunaresearch.org/docs/
isc-2019.pdf;   https:/

                                        /www.hpcwire.com/2019/
                                        08/05/llnl-purdue-researchers-
                                        harness-gpu-mixed-precision-
                                        for-accuracy-performance-
                                        tradeoff/.

Liang:2018:FMP

[LWZ18] Yun Liang, Shuo Wang, and Wei Zhang. FlexCL: A model of performance and power for OpenCL workloads on FPGAs. IEEE Transactions on Computers, 67(12):1750–1764, ???? 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8365849/.

Li:1993:MSU

[LY93] Q. Li and T. G. Yip. Monitoring systems using PVM. In Law et al. [LF^{+}93a], pages 781–785. ISBN 0-7918-1169-

7. LCCN TA345.A86 1993.

Lu:2020:GQO

[LYGG20] Q. Lu, J. Yao, H. Guan, and P. Gao. gQoS: a QoSoriented GPU virtualization with adaptive capacity sharing. IEEE Transactions on Parallel and Distributed Systems, 31(4):843– 855, April 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lopes:2019:FBD

[LYIP19] Paulo A. C. Lopes, Satyendra Singh Yadav, Aleksandar Ilic, and Sarat Kumar Patra. Fast block distributed CUDA implementation of the Hungarian algorithm. Journal of Parallel and Distributed Computing, 130(??):50–62, August 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731519302254.

Loncar:2016:OOM

[LYSS^{+}16] Vladimir Loncar, Luis E.

Young-S., Srdjan Skrbić, Paulsamy Muruganandam, Sadhan K. Adhikari, and Antun Balaz. OpenMP, OpenMP/MPI, and CUDA/ MPI C programs for solving the time-dependent dipolar Gross–Pitaevskii equation. Computer Physics Communications, 209(??): 190–196, December 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465516302272.

Lu:2013:WGA

[LYZ13] Xiangwen Lu, Jiabin Yuan, and Weiwei Zhang. Workflow of the Grover algorithm simulation incorporating CUDA and GPGPU. Computer Physics Communications, 184(9):2035– 2041, September 2013. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465513001148.

Li:1997:EHC

[LZ97] Konming Gary Li and Nabil M. Zamel. An evaluation of HPF compilers and the implementation of a parallel linear equation solver using HPF and MPI. In ACM [ACM97b], page ?? ISBN 0-89791-985-8. LCCN QA76.9.A25 A265 1997. URL http://

                                        www.supercomp.org/sc97/
                                        proceedings/TECH/LI/INDEX.

HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Luecke:2002:DDM

[LZC^{+}02] Glenn R. Luecke, Yan Zou, James Coyle, Jim Hoekstra, and Marina Kraeva. Deadlock detection in MPI programs. Concurrency and Computation: Practice and Experience, 14(11):911–932, August 25, 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

                                        tronic). URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/97519209/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=97519209{\&}PLACEBO=
                                        IE.pdf.

Lin:2020:EAM

[LZC^{+}20] Bo Lin, Chijie Zhuang, Zhenning Cai, Rong Zeng, and Weizhu Bao. An efficient and accurate MPIbased parallel simulator for streamer discharges in three dimensions. Journal of Computational Physics, 401 (??):Article 109026, January 15, 2020. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0021999119307326.

Li:2017:PCO

[LZH17] Shigang Li, Yunquan Zhang, and Torsten Hoefler. Poster: Cache-oblivious MPI allto-all communications on many-core architectures. ACM SIGPLAN Notices, 52 (8):445–446, August 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2018:COM

[LZH18] Shigang Li, Yunquan Zhang, and Torsten Hoefler. Cacheoblivious MPI all-to-all communications based on Morton order. IEEE Transactions on Parallel and Distributed Systems, 29(3):542– 555, ???? 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//ieeexplore.ieee.org/
document/8091010/.

Lu:2019:PMM

[LZHY19] Gangzhao Lu, Weizhe Zhang, Hui He, and Laurence T. Yang. Performance modeling for MPI applications with low overhead fine-grained profiling. Future Generation Computer Systems, 90(??): 317–326, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167739X18308252.

Liao:2020:DCS

[LZZ^{+}20] Xiaofei Liao, Long Zheng, Binsheng Zhang, Yu Zhang, Hai Jin, Xuanhua Shi, and Yi Lin. Dynamic cluster strategy for hierarchical rollback-recovery protocols in MPI HPC applications. Concurrency and Computation: Practice and Experience, 32(3):e4173:1– e4173:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2009:CRS

[MA09] Wenjing Ma and Gagan Agrawal. A compiler and runtime system for enabling data mining applications on GPUs. ACM SIG-PLAN Notices, 44(4):287– 288, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mavriplis:2005:HRAa

[MAB05] Dimitri J. Mavriplis, Michael J.

Aftosmis, and Marsha Berger. High resolution aerospace applications using the NASA Columbia Supercomputer. In ACM [ACM05], page 61. ISBN 1-59593-061-2. LCCN ????

Miguel:1996:APN

[MABG96] Jose Miguel, Agustin Arruabarrena, Ramon Beivide, and Jose Angel Gregorio. Assessing the performance of the new IBM SP2 communication subsystem. IEEE parallel and distributed technology: systems and applications, 4(4):12– 22, Winter 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Maffeis:1994:SSD

[Maf94] S. Maffeis. System support for distributed computing. In Gentzsch and Harms [GH94], pages 293– 301. ISBN 0-387-57981-8 (New York), 3-540-57981-8 (Berlin). LCCN QA76.88.I57 1994. DM96.00. Two volumes.

Moreno:2001:AEP

[MAGR01] Luz Marina Moreno, Francisco Almeida, Daniel González, and Casiano Rodríguez. Adaptive execution of pipelines. Lecture Notes in Computer Science, 2131:217–??, 2001.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310217.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310217.
pdf.

Mainland:2012:EHM

[Mai12] Geoffrey Mainland. Explicitly heterogeneous metaprogramming with MetaHaskell. ACM SIGPLAN Notices, 47(9):311–322, September 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Molero-Armenta:2014:OOI

[MAIVAH14] M. Molero-Armenta, Ursula Iturrarán-Viveros, S. Aparicio, and M. G. Hernández. Optimized OpenCL implementation of the Elastodynamic Finite Integration Technique for viscoelastic media. Computer Physics Communications, 185(10): 2683–2696, October 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465514001702.

Malyshkin:1995:PCT

[Mal95] Victor Malyshkin, editor. Parallel computing technologies: third international conference, PaCT-95, St. Petersburg, Russia, September 12–25, 1995: proceedings, number 964 in Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-60222-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I547 1995.

Malfetti:2001:AOW

[Mal01]  Paolo Malfetti.  Appli-

cation of OpenMP to weather, wave and ocean codes. Scientific Programming, 9(2–3):99–107, Spring–Summer 2001. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C4%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Mirvis:1995:HML

[MALM95] Y. Mirvis, F. Abdi, B. Lajevardi, and P. Murthy. Hierarchical multi-level optimization solution for massive parallel simulation of composite system. AIAA/ASME/ASCE/AHS Structures, Structural Dynamics & Materials Conference — Collection of Technical Papers, 4, ???? 1995.

CODEN CPSCDO. ISSN 0273-4508.

Manchek:1994:DIP

[Man94] Robert J. Manchek. Design and implementation of PVM version 3. M.s. thesis, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 1994. viii + 81 pp.

Mans:1998:PDP

[Man98] Bernard Mans. Portable distributed priority queues with MPI. Concurrency: practice and experience, 10(3): 175–198, March 1998. CO-DEN CPEXEI. ISSN 1040-

                                        3108. URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract?ID=5373;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=5373&PLACEBO=IE.pdf.

Manis:2001:PNP

[Man01] G. Manis. Persistent and non-persistent data objects on top of PVM and MPI. Lecture Notes in Computer Science, 2131:91–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2131/21310091.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2131/21310091.
                                        pdf.

Miguel-Alonso:2009:INS

[MANR09] J. Miguel-Alonso, J. Navaridas, and F. J. Ridruejo. Interconnection network simulation using traces of MPI applications. International Journal of Parallel Programming, 37(2): 153–174, April 2009. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=2&spage=153.

Marowka:2002:ISI

[Mar02] Ami Marowka. Introduction to the special issue: OpenMP: Experiences, implementations and applications. Parallel and Distributed Computing Practices, 5(2):v, June 2002. CO-DEN ???? ISSN 1097-2803.

Marowka:2003:EOT

[Mar03] Ami Marowka. Extending OpenMP for task parallelism. Parallel Processing Letters, 13(3):341–??, September 2003. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Marowka:2005:EMT

[Mar05] Ami Marowka. Execution model of three parallel languages: OpenMP, UPC and CAF. Scientific Programming, 13(2):127–135, ????

2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Marowka:2006:BRP

[Mar06] Ami Marowka. Book review: Parallel Scientific Computation: A Structured Approach using BSP and MPI. Scalable Computing: Practice and Experience, 7(2): 107–108, June 2006. CO-DEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.
                                        org/vols/vol07/no2/vol07no2bookreview.
                                        html.

Marowka:2007:PCD

[Mar07] Ami Marowka. Parallel computing on any desktop. Communications of the ACM, 50 (9):74–78, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Marowka:2009:BCT

[Mar09] Ami Marowka. BSP2OMP: a compiler for translating BSP programs to OpenMP. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 24(4):293–310, 2009. CO-DEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Mehta:2006:MSG

[MAS06] Paras Mehta, José Nelson Amaral, and Duane Szafron. Is MPI suitable for a generative design-pattern system?

Parallel Computing, 32(7– 8):616–626, September 2006. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mattson:1994:PEP

[Mat94] T. G. Mattson. Programming environments for parallel computing: a comparison of CPS, Linda, P4, PVM, POSYBL, and TCGMSG. In Hesham and Shriver [HS94], pages 586–594. ISBN 0-8186-5060-5. ISSN 1060-3425. LCCN ???? IEEE catalog no. 94TH0607-2.

Mattson:1995:PEP

[Mat95] Timothy G. Mattson. Programming environments for parallel and distributed computing: a comparison of P4, PVM, Linda, and TCGMSG. International Journal of Supercomputer Applications and High Performance Computing, 9(2): 138–161, Summer 1995. CO-DEN IJSCFG. ISSN 1078-3482.

Mattson:2000:BOF

[Mat00a]  Tim Mattson.  BOF:

OpenMP and its future developments. In ACM [ACM00], page 106. URL

http://www.sc2000.org/
proceedings/info/fp.pdf.

Mattson:2000:IO

[Mat00b] Timothy G. Mattson. An introduction to OpenMP 2.0.

Lecture Notes in Computer Science, 1940:384–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1940/19400384.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1940/19400384.
pdf.

Mattson:2001:EO

[Mat01a] Timothy Mattson. The evolution of OpenMP. Lecture Notes in Computer Science, 1947:19–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1947/19470019.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1947/19470019.
pdf.

Matuszek:2001:APS

[Mat01b] Mariusz R. Matuszek. Assessment of PVM suitability to testbed client-agent-server applications. Lecture Notes in Computer Science, 2131:69–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310069.htm;
 http://link.springer-

ny.com/link/service/series/
0558/papers/2131/21310069.
pdf.

Mattson:2003:HGO

[Mat03] Timothy G. Mattson. How good is OpenMP? Scientific Programming, 11(2): 81–93, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Matloff:2016:PCD

[Mat16] Norman S. Matloff. Parallel Computing for Data Science: with Examples in R, C++ and CUDA, volume 28 of Chapman and Hall/CRC: The R Series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2016. ISBN 1-4665-8701-6 (hardcover). xxiii + 324 pp. LCCN QA76.642 M37 2016. URL http:

//www.tandf.net/books/
details/9781466587014.

Mourao:2000:SSC

[MB00] Elson Mourão and Stephen Booth. Single sided communications in multi-protocol MPI. Lecture Notes in Computer Science, 1908: 176–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080176.htm;
 http://link.springer-

                                        ny.com/link/service/series/
                                        0558/papers/1908/19080176.
                                        pdf.

Marongiu:2012:OCE

[MB12] Andrea Marongiu and Luca Benini. An OpenMP compiler for efficient use of distributed scratchpad memory in MPSoCs. IEEE Transactions on Computers, 61 (2):222–236, February 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Maleki:2018:AHP

[MB18] Sepideh Maleki and Martin Burtscher. Automatic hierarchical parallelization of linear recurrences. ACM SIGPLAN Notices, 53(2): 128–138, February 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Margolin:2021:TBF

[MB21] Alexander Margolin and Amnon Barak. Treebased fault-tolerant collective operations for MPI. Concurrency and Computation: Practice and Experience, 33(14):e5826:1– e5826:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Muruganandam:2021:OSR

[MBA21] Paulsamy Muruganandam, Antun Balaz, and Sadhan K. Adhikari. OpenMP solver for rotating spin-1 spin-orbit- and Rabicoupled Bose–Einstein condensates. Computer Physics Communications, 264(??): Article 107926, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465521000618.

Muller:2012:SOA

[MBB^{+}12] Matthias S. Müller, John Baron, William C. Brantley, Huiyu Feng, and Daniel Hackenberg. SPEC OMP2012— an application benchmark suite for parallel systems using OpenMP. Lecture Notes in Computer Science, 7312:223– 236, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-30961-8_
17/.

Ma:2013:KAT

[MBBD13] Teng Ma, George Bosilca, Aurelien Bouteiller, and Jack J. Dongarra. Kernelassisted and topology-aware MPI collective communications on multicore/manycore platforms. Journal of Parallel and Distributed Computing, 73(7): 1000–1010, July 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731513000166.

Min:2003:OOP

[MBE03] Seung-Jai Min, Ayon Basumallik, and Rudolf Eigenmann. Optimizing OpenMP programs on software distributed shared memory systems. International Journal of Parallel Programming, 31(3):225–249, June 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL /ips/frames/
Refs/referenceskapmain.
asp?J=4773&I=33&A=5&LK=
NM;   http://ipsapp007.
kluweronline.com/content/
getfile/4773/33/5/abstract.
htm;   http://ipsapp007.
kluweronline.com/content/
getfile/4773/33/5/fulltext.
pdf.

McKenzie:1994:CIM

[MBES94] N. R. McKenzie, K. Bolding, C. Ebeling, and L. Snyder. CRANIUM: An interface for message passing on adaptive packet routing networks. In Bolding and Snyder [BS94], pages 266–280. ISBN 3-540-58429-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.P39 1994.

Malits:2012:ELG

[MBKM12] Roman Malits, Evgeny Bolotin, Avinoam Kolodny, and Avi Mendelson. Exploring the limits of GPGPU scheduling in control flow bound applications. ACM Transactions on Architecture and Code Optimization, 8(4):29:1–29:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mehl:2015:RTC

[MBS15] Miriam Mehl, Manfred Bischoff, and Michael Schäfer, editors. Recent Trends in Computational Engineering — CE2014: Optimization, Uncertainty, Parallel Algorithms, Coupled and Complex Problems, volume 105 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-22996-6, 3-319-22997-4 (e-book). 317 (est.) pp. LCCN QA71-90; TA329. URL http:

//www.springerlink.com/
content/978-3-319-22997-
3.

Miles:1994:PTO

[MC94] Roger Miles and Alan Chalmers, editors. Progress in Transputer and occam Research, WoTUG-17 Proceedings of the 17th World occam and Transputer User Group Technical Meeting, April 10– 13, 1994, Bristol, UK, volume 38 of Transputer and Occam Engineering Series. IOS Press, Postal Drawer 10558, Burke, VA 2209-0558, USA, 1994. ISBN 90-5199-163-0. LCCN ????

Medeiros:1998:IPM

[MC98] P. D. Medeiros and J. C. Cunha. Interconnecting PVM and MPI applications. Lecture Notes in Computer Science, 1497:105–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morrison:1999:FPP

[MC99] J. P. Morrison and R. W. Connolly. Facilitating parallel programming in PVM using condensed graphs. In Dongarra et al. [DLM99], pages 181–188. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Maier:2017:OLD

[MC17] Andrew J. Maier and Bruce F. Cockburn. Optimization of low-density parity check decoder performance for OpenCL designs synthesized to FPGAs. Journal of Parallel and Distributed Computing, 107(??): 134–145, September 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731517301004.

Malinowski:2018:SIP

[MC18] Artur Malinowski and Pawel Czarnul. A solution to image processing with parallel MPI I/O and distributed NVRAM cache. Scalable Computing: Practice and Experience, 19(1):1–14, ???? 2018. CODEN ???? ISSN 1895-1767. URL https://

www.scpe.org/index.php/
scpe/article/view/1389.

Massaioli:2005:OPA

[MCB05] Federico Massaioli, Filippo Castiglione, and Massimo Bernaschi. OpenMP parallelization of agent-based models. Parallel Computing, 31(10–12):1066–1081, October/December 2005. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

McDonald:1996:NNP

[McD96] K. McDonald. The NAG Numerical PVM Library. In Dongarra et al. [DMW96], pages 419–428. ISBN 3-540-60902-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.P35 1995.

Mueller:2008:OSM

[MCdS^{+}08] Matthias S. Mueller, Barbara M. Chapman, Bronis R. de Supinski, Allen D. Malony, and Michael Voss, editors. OpenMP Shared Memory Parallel Programming: International Workshops, IWOMP 2005 and IWOMP 2006, Eugene, OR, USA, June 1–4, 2005, Reims, France, June 12–15, 2006. Proceedings, volume 4315 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CO-DEN LNCSD9. ISBN 3-540-68554-5 (print), 3-540-68555-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:/

/www.springerlink.com/
content/978-3-540-68555-
5.

McKinney:1994:PGU

[McK94] G. W. McKinney. A practical guide to using MCNP with PVM. Transactions of the American Nuclear Society, 71(????):397–398, ???? 1994. CODEN TANSAO. ISSN 0003-018X.

Moore:2001:RPA

[MCLD01] Shirley Moore, David Cronk, Kevin London, and Jack Dongarra. Review of performance analysis tools for MPI parallel programs. Lecture Notes in Computer Science, 2131:241–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310241.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310241.
pdf.

Moreira:2017:FCR

[MCP17] Rubens E. A. Moreira, Sylvain Collange, and Fernando Magno Quintão Pereira. Function call re-vectorization. ACM SIGPLAN Notices, 52 (8):313–326, August 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McRae:1992:VC

[McR92] S. J. McRae. VM communications. In Anonymous [Ano92], pages 439–453.

Mierendorff:2000:WMB

[MCS00] Hermann Mierendorff, Kläre Cassirer, and Helmut Schwamborn. Working with MPI benchmarking suites on ccNUMA architectures. Lecture Notes in Computer Science, 1908: 18–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080018.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080018.
pdf.

Marin:2017:ERF

[MDM17] Manuel Marin, David Defour, and Federico Milano. An efficient representation format for fuzzy intervals based on symmetric membership functions. ACM Transactions on Mathematical Software, 43(3):23:1– 23:??, January 2017. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/citation.
cfm?id=2939364.

Monteiro:2018:EGC

[MdSAS^{+}18] Felipe R. Monteiro, Erickson H. da S. Alves, Isabela S. Silva, Hussama I. Ismail, Lucas C. Cordeiro, and Eddie B. de Lima Filho. ESBMC-GPU: a contextbounded model checking tool to verify CUDA programs. Science of Computer Programming, 152(??): 63–69, January 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167642317301934.

Muller:2009:EOA

[MdSC09] Matthias S. Müller, Bronis R. de Supinski, and Barbara M. Chapman, editors. Evolving OpenMP in an Age of Extreme Parallelism: 5th International Workshop on OpenMP, IWOMP 2009 Dresden, Germany, June 3– 5, 2009 Proceedings, volume 5568 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CO-DEN LNCSD9. ISBN 3-642-02284-7 (print), 3-642-02303-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:/

/www.springerlink.com/
content/978-3-642-02303-
3.

Matheou:2017:DDC

[ME17] George Matheou and Paraskevas Evripidou. Data-driven concurrency for high performance computing. ACM Transactions on Architecture and Code Optimization, 14(4):53:1–53:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Megson:1998:CRH

[MFC98] G. M. Megson, R. S. Fish, and D. N. J. Clarke. Creation of reconfigurable hardware objects in PVM environments. Lecture Notes in Computer Science, 1497: 215–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milovanovic:2008:NEE

[MFG^{+}08] Milos Milovanović, Roger Ferrer, Vladimir Gajinov, Osman S. Unsal, Adrian Cristal, Eduard Ayguadé, and Mateo Valero. Nebelung: Execution environment for transactional OpenMP. International Journal of Parallel Programming, 36(3): 326–346, June 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=3&spage=326.

Moody:2003:SNB

[MFPP03] Adam Moody, Juan Fernandez, Fabrizio Petrini, and Dhabaleswar K. Panda. Scalable NIC-based reduction on large-scale clusters. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN ???? URL http:/

/www.sc-conference.org/
sc2003/inter_cal/inter_
cal_detail.php?eventid=
10716#2;   http://www.
sc-conference.org/sc2003/
paperpdfs/pap316.pdf.

Martin:1995:DPC

[MFTB95] I. Martin, J. C. Fabero, F. Tirado, and A. Bautista. Distributed parallel computers versus PVM on a workstation cluster in the simulation of time dependent partial differential equations. In IEEE [IEE95h], pages 20–26. ISBN 0-8186-7031-2, 0-8186-7032-0. LCCN QA76.58 .E97 1995.

Mintchev:1997:TPM

[MG97] S. Mintchev and V. Getov.

Towards portable message passing in Java: Binding MPI. Lecture Notes in Computer Science, 1332: 135–142, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehta:2015:MTP

[MG15] Kshitij Mehta and Edgar Gabriel. Multi-threaded parallel I/O for OpenMP applications. International Journal of Parallel Programming, 43(2):286–309, April 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10766-014-0306-9.

Mendonca:2017:DAA

[MGA^{+}17] Gleison Mendonça, Breno Guimarães, Péricles Alves, Márcio Pereira, Guido Araújo, and Fernando Magno Quintão Pereira. DawnCC: Automatic annotation for data parallelism and offloading. ACM Transactions on Architecture and Code Optimization, 14(2):13:1–13:??, July 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mehta:2012:SPE

[MGC12] Kshitij Mehta, Edgar Gabriel, and Barbara Chapman. Specification and performance evaluation of parallel I/O interfaces for OpenMP. Lecture Notes in Computer Science, 7312: 1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-30961-8_
1/.

Muralidharan:2015:COP

[MGC^{+}15] Saurav Muralidharan, Michael Garland, Bryan Catanzaro, Albert Sidelnik, and Mary Hall. A collection-oriented programming model for performance portability. ACM SIGPLAN Notices, 50(8): 263–264, August 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Medvedev:2005:OMA

[MGG05] Dmitry M. Medvedev, Evelyn M. Goldfield, and Stephen K. Gray. An OpenMP/MPI approach to the parallelization of iterative four-atom quantum mechanics. Computer Physics Communications, 166(2):94– 108, March 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465504005260.

Montella:2017:VCB

[MGL^{+}17] Raffaele Montella, Giulio Giunta, Giuliano Laccetti, Marco Lapegna, Carlo Palmieri, Carmine Ferraro, Valentina Pelliccia, Cheol-Ho Hong, Ivor Spence, and Dimitrios S. Nikolopoulos. On the virtualization of CUDA based GPU remoting on ARM and x86 machines in the GVirtuS framework. International Journal of Parallel Programming, 45(5):1142–1163, October 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Mazzariol:1997:PCS

[MGMH97] M. Mazzariol, B. A. Gennart, V. Messerli, and R. D. Hersch. Performance of CAP-specified linear algebra algorithms. Lecture Notes in Computer Science, 1332: 351–358, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Markidis:2015:OAN

[MGS^{+}15] Stefano Markidis, Jing Gong, Michael Schliephake, Erwin Laure, Alistair Hart, David Henty, Katherine Heisey, and Paul Fischer. OpenACC acceleration of the Nek5000 spectral element code. The International Journal of High Performance Computing Applications, 29(3):311–319, August 2015. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Matthey:2001:EMO

[MH01] T. Matthey and J. P. Hansen. Evaluation of MPI’s one-sided communication mechanism for shortrange molecular dynamics on the Origin2000. Lecture Notes in Computer Science, 1947:356–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1947/19470356.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1947/19470356.
                                        pdf.

Hwu:2012:GCG

[mH12] Wen mei Hwu, editor. GPU computing gems. Applications of GPU computing series. Morgan Kaufmann, Boston, MA, jade edition, 2012. ISBN 0-12-385963-8 (hardback). xvi + 541 + 16 pp. LCCN T385 .G6875 2012.

Moll:2018:PCF

[MH18] Simon Moll and Sebastian Hack. Partial control-flow linearization. ACM SIG-PLAN Notices, 53(4):543– 556, April 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muller:2021:MAE

[MH21] Stefan K. Muller and Jan Hoffmann. Modeling and analyzing evaluation cost of CUDA kernels. Proceedings of the ACM on Programming Languages (PACMPL), 5 (POPL):25:1–25:31, January 2021. URL https://

dl.acm.org/doi/10.1145/
3434306.

Miller:1994:PPP

[MHC94a] B. P. Miller, J. K. Hollingsworth, and M. D. Callaghan. The Paradyn parallel performance tools and PVM. In Dongarra and Tourancheau [DT94], pages 201–210. ISBN 0-89871-343-9. LCCN QA76.58.I568 1994.

Miller:1994:PPT

[MHC94b] B. P. Miller, J. K. Hollingworth, and M. D. Callaghan. The Paradyn performance tools and PVM. In Dongarra and Tourancheau [DT94], pages 201–210. ISBN 0-89871-343-9. LCCN QA76.58.I568 1994.

Munshi:2016:OCS

[MHSK16] Aaftab Munshi, Lee Howes, Bartosz Sochacki, and Khronos OpenCL Working Group. The OpenCL C specification version: 2.0 document revision: 33. Web document., April 13, 2016. URL https:

//www.khronos.org/registry/
OpenCL/specs/opencl-2.
0-openclc.pdf.

Michielse:1993:PMU

[Mic93] P. Michielse. Parallel multigrid using PVM. Supercomputer, 10(6):10–23, ???? 1993. CODEN SPCOEL. ISSN 0168-7875.

Michielse:1995:PMU

[Mic95]  Peter Michielse.  Paral-

lel multigrid using PVM. Applied Numerical Mathematics: Transactions of IMACS, 19(1-2):63–69, November 1995. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Muddukrishna:2015:LAT

[MJB15] Ananya Muddukrishna, Peter A. Jonsson, and Mats Brorsson. Locality-aware task scheduling and data distribution for OpenMP programs on NUMA systems and manycore processors. Scientific Programming, 2015(??):981759:1– 981759:16, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL https://

www.hindawi.com/journals/
sp/2015/981759/.

Mittal:2012:CAS

[MJG^{+}12] Anshul Mittal, Nikhil Jain, Thomas George, Yogish Sabharwal, and Sameer Kumar. Collective algorithms for sub-communicators. ACM SIGPLAN Notices, 47(8): 315–316, August 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Muddukrishna:2016:GGO

[MJPB16] Ananya Muddukrishna, Peter A. Jonsson, Artur Podobas, and Mats Brorsson. Grain graphs: OpenMP performance analysis made easy. ACM SIGPLAN Notices, 51(8):28:1–28:??, August 2016. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Matyska:1994:DCS

[MK94] Ludĕk Matyska and Jaroslav Koča. D-CICADA: a software for conformational PES elucidation on network of workstations. Journal of Computational Chemistry, 15(9):937–946, September 1994. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

McDonald:1997:IPT

[MK97] Chris McDonald and Kamran Kazemi. Improving the PVM teaching environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):219– 223, March 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McDonald:2000:TPA

[MK00] Chris McDonald and Kamran Kazemi. Teaching parallel algorithm with process topologies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):70– 74, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mohror:2004:PTS

[MK04] Kathryn Mohror and Karen L. Karavanic. Performance tool support for MPI-2 on Linux. In ACM [ACM04], page 28. ISBN 0-7695-2153-3. LCCN ????

Manwade:2017:DFA

[MK17] Karveer B. Manwade and Dinesh B. Kulkarni. Data flow analysis of MPI program using dynamic analysis technique with partial execution. Scalable Computing: Practice and Experience, 18(4):375–385, ???? 2017. CODEN ???? ISSN 1895-1767. URL https://

www.scpe.org/index.php/
scpe/article/view/1335.

Maheo:2012:AOL

[MKC^{+}12] Aurèle Mahéo, Souad Koliaï, Patrick Carribault, Marc Pérache, and William Jalby. Adaptive OpenMP for large NUMA nodes. Lecture Notes in Computer Science, 7312: 254–257, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-30961-8_
20/.

Munch:2021:HDE

[MKK21] Peter Munch, Katharina Kormann, and Martin Kronbichler. hyper.deal: an efficient, matrix-free finiteelement library for highdimensional partial differential equations. ACM Transactions on Mathematical Software, 47(4):33:1– 33:34, December 2021. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

dl.acm.org/doi/10.1145/
3469720.

Markus:1996:PEM

[MKP^{+}96] S. Markus, S. B. Kim, K. Pantazopoulos, A. L. Ocken, E. N. Houstis, P. Wu, S. Weerawarana, and D. Maharry. Performance evaluation of MPI implementations and MPI based Parallel ELLPACK solvers. In IEEE [IEE96i], pages 162– 169. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Meyer:2022:DFA

[MKP22] Marius Meyer, Tobias Kenter, and Christian Plessl. In-depth FPGA accelerator performance evaluation with single node benchmarks from the HPC challenge benchmark suite for Intel and Xilinx FPGAs using OpenCL. Journal of Parallel and Distributed Computing, 160 (??):79–89, February 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0743731521002057.

Min:2001:PCO

[MKV^{+}01] Seung Jai Min, Seon Wook Kim, Michael Voss, Sang Ik Lee, and Rudolf Eigenmann. Portable compilers for OpenMP. Lecture Notes in Computer Science, 2104:11–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2104/21040011.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2104/21040011.
                                        pdf.

Mokbel:2011:ASR

[MKW11] Mohammed F. Mokbel, Robert D. Kent, and Michael Wong. An abstract semantically rich compiler collocative and interpretative model for OpenMP programs. The Computer Journal, 54(8):1325–1343, August 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/8/1325.
full.pdf+html.

Mitra:2014:AAP

[MLA^{+}14] Subrata Mitra, Ignacio Laguna, Dong H. Ahn, Saurabh Bagchi, Martin Schulz, and Todd Gamblin. Accurate application progress analysis for large-scale parallel debugging. ACM SIG-PLAN Notices, 49(6):193– 203, June 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marjanovic:2010:ECC

[MLAV10] Vladimir Marjanovic, Jesús Labarta, Eduard Ayguadé, and Mateo Valero. Effective communication and computation overlap with hybrid MPI/SMPSs. ACM SIG-PLAN Notices, 45(5):337– 338, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marowka:2004:OOA

[MLC04] Ami Marowka, Zhenying Liu, and Barbara Chapman. OpenMP-oriented applications for distributed shared memory architectures. Concurrency and Computation: Practice and Experience, 16 (4):371–384, April 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Malakhov:2018:CMT

[MLGW18] Anton Malakhov, David Liu, Anton Gorshkov, and Terry Wilmarth. Composable multi-threading and multiprocessing for numeric libraries. In Fatih Akici, David Lippa, Dillon Niederhut, and M. Pacer, editors, Proceedings of the 17th Python in Science Conference, Austin, TX, 9–15 July 2018, pages 15–21. ????, ????, 2018. URL http:

                                        //conference.scipy.org/
                                        proceedings/scipy2018/
                                        anton_malakhov.html.

Marendic:2016:NMR

[MLVS16] P. Marendic, J. Lemeire, D. Vucinic, and P. Schelkens. A novel MPI reduction algorithm resilient to imbalances in process arrival times. The Journal of Supercomputing, 72(5):1973– 2013, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/s11227-016-1707-x.

Majumdar:1992:PPC

[MM92] A. Majumdar and W. R.

Martin. Parallel preconditioned conjugate gradient algorithm applied to neutron diffusion problem. Transactions of the American Nuclear Society, 65:209–210, 1992. CODEN TANSAO. ISSN 0003-018X.

Mantovani:1995:HPS

[MM95] M. L. Mantovani and M. Malagoli. Highly parallel SCF calculation: the SYSMO program. In IEEE [IEE95h], pages 502–507. ISBN 0-8186-7031-2, 0-8186-7032-0. LCCN QA76.58 .E97 1995.

Michailidis:2001:TSH

[MM01] Panagiotis D. Michailidis and Konstantinos G. Margaritis. Text searching on a heterogeneous cluster of workstations. Lecture Notes in Computer Science, 2131:378–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310378.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310378.
pdf.

Michailidis:2002:PSL

[MM02] Panagiotis D. Michailidis and Konstantinos G. Margaritis. A performance study of load balancing strategies for approximate string matching on an MPI heterogeneous system environment. Lecture Notes in Computer Science, 2474: 432–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.de/link/
service/series/0558/bibs/
2474/24740432.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740432.pdf.

Michailidis:2003:PEL

[MM03] Panagiotis D. Michailidis and Konstantinos G. Margaritis. Performance evaluation of load balancing strategies for approximate string matching application on an MPI cluster of heterogeneous workstations. Future Generation Computer Systems, 19(7):1075–1104, October 2003. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Marathe:2007:SCC

[MM07] Jaydeep Marathe and Frank Mueller. Source-code-correlated cache coherence characterization of OpenMP benchmarks. IEEE Transactions on Parallel and Distributed Systems, 18(6):818– 834, June 2007. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Michailidis:2011:PDM

[MM11] Panagiotis D. Michailidis and Konstantinos G. Margaritis. Parallel direct methods for solving the system of linear equations with pipelining on a multicore using OpenMP. Journal of Computational and Applied Mathematics, 236(3): 326–341, September 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711004183.

Morishima:2014:PEG

[MM14] Shin Morishima and Hiroki Matsutani. Performance evaluations of graph database using CUDA and OpenMP compatible libraries. ACM SIGARCH Computer Architecture News, 42(4):75–80, 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Mofrad:2020:GNA

[MMAH20]  Mohammad  Hasanzadeh

Mofrad, Rami Melhem, Yousuf Ahmad, and Mohammad Hammoud. Graphite: a NUMA-aware HPC system for graph analytics based on a new MPI * X parallelism model. Proceedings of the VLDB Endowment, 13 (6):783–797, February 2020. CODEN ???? ISSN 2150-8097. URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        14778/3380750.3380751.

Malony:1994:PAP

[MMB^{+}94] A. Malony, B. Mohr, P. Beckman, D. Gannon, S. Yang, and F. Bodin. Performance analysis of pC++: a portable data-parallel programming system for scalable parallel computers. In Siegal [Sie94], pages 75–84. ISBN 0-8186-5602-6. LCCN QA76.58.I58 1994. IEEE catalog no. 94CH34819.

Mackay:1998:SPF

[MMD98] David Mackay, G. Mahinthakumar, and Ed D’Azevedo. A study of I/O in a parallel finite element groundwater transport code. The International Journal of High Performance Computing Applications, 12(3): 307–319, Fall 1998. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

                                        //journals.sagepub.com/
                                        doi/pdf/10.1177/109434209801200301.

Mironov:2019:EMO

[MMDA19] Vladimir Mironov, Alexander Moskovsky, Michael D’Mello, and Yuri Alexeev. An efficient MPI/ OpenMP parallelization of the Hartree–Fock–Roothaan method for the first generation of Intel(R) Xeon Phi^{T M} processor architecture. The International Journal of High Performance Computing Applications, 33(1): 212–224, January 1, 2019. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:

//journals.sagepub.com/
doi/full/10.1177/1094342017732628.

Mudge:1993:PTS

[MMH93] T. N. Mudge, V. Milutinovic, and L. Hunter, editors. Proceedings of the Twenty-Sixth Hawaii International Conference on System Science (HICSS-26), held in Wailea, Hawaii in January 5–8, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-3230-5. LCCN ???? Four volumes. IEEE catalog number 93TH0501-7.

Morimoto:1998:IMM

[MMH98] K. Morimoto, T. Matsumoto, and K. Hiraki. Implementing MPI with the memory-based communication facilities on the SSS-CORE operating system. Lecture Notes in Computer Science, 1497:223–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morimoto:1999:PEM

[MMH99] K. Morimoto, T. Matsumoto, and K. Hiraki. Performance evaluation of the MPI/MBCF with the NAS parallel benchmarks. In Dongarra et al. [DLM99], pages 19–26. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Mohamed:2013:MMM

[MMM13] Hisham Mohamed and Stéphane Marchand-Maillet. MRO-MPI: MapReduce overlapping using MPI and an optimized data exchange policy. Parallel Computing, 39(12): 851–866, December 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819113001026.

Manca:2016:CQI

[MMO^{+}16] Emanuele Manca, Andrea Manconi, Alessandro Orro, Giuliano Armano, and Luciano Milanesi. CUDAquicksort: an improved GPU-based implementation of quicksort. Concurrency and Computation: Practice and Experience, 28(1):21– 43, January 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

MacFarlane:1999:PPI

[MMR99] A. MacFarlane, J. A. Mc-Cann, and S. E. Robertson. PLIERS: a parallel information retrieval system using MPI. In Dongarra et al. [DLM99], pages 317– 324. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Morris:2007:SNO

[MMS07] Alan Morris, Allen D. Malony, and Sameer S. Shende. Supporting nested OpenMP parallelism in the TAU performance system. International Journal of Parallel Programming, 35(4):417– 436, August 2007. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=4&spage=417.

Mohr:2002:DPP

[MMSW02] Bernd Mohr, Allen D. Malony, Sameer Shende, and Felix Wolf. Design and prototype of a performance tool interface for OpenMP. The Journal of Supercomputing, 23(1):105–128, August 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp008.kluweronline.
com/content/getfile/5189/

                                        37/8/abstract.htm;   http:
                                        //ipsapp008.kluweronline.
                                        com/content/getfile/5189/
                                        37/8/fulltext.pdf.

Matuszek:1999:BPG

[MMU99] M. R. Matuszek, A. Mazurkiewicz, and P. W. Uminski. Benchmarking the PVM group communication efficiency. In Dongarra et al. [DLM99], pages 499–508. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Martin:1996:WTW

[MMW96] D. E. Martin, T. J. McBrayer, and P. A. Wilsey. WARPED: a time warp simulation kernel for analysis and application development. In H. El-Rewini and B. D. Shriver, editors, Proceedings of the Twenty-Ninth Hawaii International Conference on System Sciences, volume 1, pages 5–?? ????, ????, 1996. ISBN 0-8186-7324-9. LCCN ????

Meleshchuk:1991:IPP

[MN91] S. B. Meleshchuk and A. N. Nedumov. Implementation of a protocol for parallel database access with virtual machine communications facilities. Programmirovanie, 17(1):35–42, January/February 1991. CO-DEN PCSODA. ISSN 0132-3474, 0361-7688. English translation in Programming and Computer Software, vol. 17, no. 1, pp. 27–32, November 1991.

Martinez-Noriega:2021:COE

[MNYN21] Edgar Josafat Martinez-Noriega, Syunji Yazaki, and Tetsu Narumi. CUDA offloading for energy-efficient and high-frame-rate simulations using tablets. Concurrency and Computation: Practice and Experience, 33 (2):e5488:1–e5488:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Midorikawa:2005:PNM

[MOL05] Edson Toshimi Midorikawa, Helio Marci Oliveira, and Jean Marcos Laine. PEM-PIs: a new methodology for modeling and prediction of MPI programs performance. International Journal of Parallel Programming, 33 (5):499–527, October 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=5&spage=499.

Mork:1995:DPP

[Mor95] P. Mork. Debugging parallel programs with execution tracing. In Ferenczi and Kacsuk [FK95], pages 176– 183. ISBN ???? LCCN ???? Technical report KFKI-1995-2/M,N.

Manke:1995:MPP

[MP95] J. W. Manke and J. C. Patterson. Message passing performance of Intel Paragon, IBM SP1 and CRAY T3D using PVM. In Bailey et al. [BBG^{+}95], pages 768–769. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Martin:2004:HPA

[MPD04] María J. Martín, Marta Parada, and Ramón Doallo. High performance air pollution simulation using OpenMP. The Journal of Supercomputing, 28(3):311– 321, June 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

                                        ipsapp008.kluweronline.
                                        com/IPS/content/ext/x/
                                        J/5189/I/54/A/5/abstract.
                                        htm.

MF:1998:SIM

[MPI98a] MPI Forum. Special issue: MPI2: a messagepassing interface standard. The International Journal of High Performance Computing Applications, 12(1–2):1– 299, Spring–Summer 1998. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

MPIForum:1998:SIM

[MPI98b] MPI Forum. Special issue: MPI2: a message-passing interface standard. International Journal of Supercomputer Applications and High Performance Computing, 12(1–2):1–299, Spring– Summer 1998. CODEN IJSCFG. ISSN 1078-3482.

Mena:2020:GAS

[MPS20] Hermann Mena, Lena-Maria Pfurtscheller, and Tony Stillfjord. GPU acceleration of splitting schemes applied to differential matrix equations. Numerical Algorithms, 83(1):395– 419, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.
springer.com/content/pdf/
10.1007/s11075-019-00687-
w.pdf.

Meyer:2021:IBH

[MPZ21] Bruno Henrique Meyer, Aurora Trinidad Ramirez Pozo, and Wagner M. Nunan Zola. Improving Barnes–Hut t-SNE algorithm in modern GPU architectures with random forest KNN and simulated wide-warp. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):53:1– 53:26, October 2021. CO-DEN ???? ISSN 1550-4832.

URL https://dl.acm.org/
doi/10.1145/3447779.

Muller:1996:CDI

[MR96] A. Muller and R. Ruhl.

Communication-buffers for data-parallel, irregular computations. In Szymanski and Sinharoy [SS96], pages 295– 298. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Martins:2012:PDC

[MR12] Wellington S. Martins and Thiago F. Rangel. Phylogenetic distance computation using CUDA. Lecture Notes in Computer Science, 7409: 168–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/10.
                                        1007/978-3-642-31927-3_
                                        15/.

Meister:2017:PME

[MRB17] Oliver Meister, Kaveh Rahnema, and Michael Bader. Parallel memory-efficient adaptive mesh refinement on structured triangular meshes with billions of grid cells. ACM Transactions on Mathematical Software, 43(3): 19:1–19:27, January 2017. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

                                        //dl.acm.org/citation.
                                        cfm?id=2947668.

Mo:1996:IOP

[MRH^{+}96] J. Mo, F. Romelfanger, R. J. Hanisch, D. Redding, S. Sirlin, and A. Boden. Implementation of an optical prescription retrieval code using PVM (parallel virtual machine) in a mixed architecture network. In Jacoby and Barnes [JB96], pages 100–

103. ISBN ???? ISSN 1080-7926. LCCN QB51.3.E43 A87 1995.

Mininni:2011:HMO

[MRRP11] Pablo D. Mininni, Duane Rosenberg, Raghu Reddy, and Annick Pouquet. A hybrid MPI–OpenMP scheme for scalable parallel pseudospectral computations for fluid turbulence. Parallel Computing, 37(6–7): 316–326, June/July 2011. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819111000512.

Mazzocca:2000:TPP

[MRV00] N. Mazzocca, M. Rak, and U. Villano. The transition from a PVM program simulator to a heterogeneous system simulator: The HeSSE project. Lecture Notes in Computer Science, 1908:266–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080266.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080266.
pdf.

Morinishi:1995:PIB

[MS95] K. Morinishi and N. Satofuka. Parallel implementation of the Boltzmann equation solvers using PVM. In Satofuka et al. [SPE95], pages 339–346. ISBN 0-444-82317-4. LCCN QA911 .P35 1994.

McMahon:1996:EEE

[MS96a] T. P. McMahon and A. Skjellum. eMPI/eMPICH: embedding MPI. In IEEE [IEE96i], pages 180–184. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Menden:1996:PPP

[MS96b] J. Menden and G. Stellner. Proving properties of PVM applications — a case study with CoCheck. In Bode et al. [BDLS96], pages 134–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Marinho:1998:WMP

[MS98] J. Marinho and J. G. Silva. WMPI — message passing interface for Win32 clusters. Lecture Notes in Computer Science, 1497:113–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mierendorff:1999:PMB

[MS99a] H. Mierendorff and H. Schwamborn. Performance modeling based on PVM. In Dongarra et al. [DLM99], pages 75– 82. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Migliardi:1999:PEH

[MS99b] M. Migliardi and V. Sunderam. PVM emulation in the harness metacomputing system: a plugin based approach. In Dongarra et al. [DLM99], pages 117–124. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Mourao:1999:IMO

[MS99c] F. E. Mourao and J. G.

Silva. Implementing MPI’s one-sided communications for WMPI. In Dongarra et al. [DLM99], pages 231– 240. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Macias:2002:SEA

[MS02a] Elsa M. Macías and Alvaro Suárez. Solving engineering applications with LAM-GAC over MPI-2. Lecture Notes in Computer Science, 2474:130–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/
service/series/0558/bibs/
2474/24740130.htm;   http:
//link.springer.de/link/

                                        service/series/0558/papers/
                                        2474/24740130.pdf.

Mahinthakumar:2002:HMO

[MS02b] G. Mahinthakumar and F. Saied. A hybrid MPI-OpenMP implementation of an implicit finite-element code on parallel architectures. The International Journal of High Performance Computing Applications, 16(4):371–393, Winter 2002. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Mertens:2004:CCP

[MS04] Stephan Mertens and Alexander Schinner. Cluster Computing: Praktische Einführung in das wissenschaftliche Rechnen auf Workstation-Clustern. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-42299-4. 300 (est.) pp. LCCN ???? Includes CD-ROM.

Mysliwiec:1997:IPS

[MSB97] G. Mysliwiec, J. Sipowicz, and H. Burkhart. Implementing parallel SBS-type linear solvers using ALWAN. Lecture Notes in Computer Science, 1332:359–366, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matise:1995:PCG

[MSCW95] T. C. Matise, M. D.

Schroeder, D. M. Chiarulli, and D. E. Weeks. Parallel computation of genetic likelihoods using CRI-MAP, PVM, and a network of distributed workstations. Human heredity, 45(2):103– ??, ???? 1995. CODEN HUHEAS. ISSN 0001-5652.

Migliardi:2000:SFT

[MSF00] Mauro Migliardi, Vaidy Sunderam, and Arrigo Frisiani. A simple, fault tolerant naming space for the HARNESS metacomputing system. Lecture Notes in Computer Science, 1908:152–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080152.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080152.
pdf.

McCandless:1996:OOM

[MSL96] B. C. McCandless, J. M.

Squyres, and A. Lumsdaine. Object oriented MPI (OOMPI): a class library for the Message Passing Interface. In IEEE [IEE96i], pages 87–94. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Massetto:2012:NSB

[MSL12] Francisco Isidro Massetto, Liria Matsumoto Sato, and Kuan-Ching Li. A novel strategy for building interoperable MPI environment in heterogeneous high performance systems. The Journal of Supercomputing, 60(1):87–116, April 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
60&issue=1&spage=87.

Mattson:2005:PPP

[MSM05] Timothy G. Mattson, Beverly A. Sanders, and Berna Massingill. Patterns for Parallel Programming. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-321-22811-1 (hardcover). xiii + 355 pp. LCCN QA76.642 .M38 2005. URL http://

www.loc.gov/catdir/toc/
ecip0418/2004013240.html.

Martin:2015:EPM

[MSMC15] Gonzalo Martín, David E. Singh, Maria-Cristina Marinescu, and Jesús Carretero. Enhancing the performance of malleable MPI applications by using performanceaware dynamic reconfiguration. Parallel Computing, 46(??):60–77, July 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819115000642.

Molnar:2010:APM

[MSML10] F. Molnár, Jr., T. Szakály, R. Mészáros, and I. Lagzi. Air pollution modelling using a Graphics Processing Unit with CUDA. Computer Physics Communications, 181(1):105–112, January 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465509002872.

Macias:2001:PPA

[MSOGR01] Elsa M. Macías, Alvaro Suárez, C. N. Ojeda-Guerra, and E. Robayna. Programming parallel applications with LAMGAC in a LAN–WLAN environment. Lecture Notes in Computer Science, 2131:158–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310158.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310158.
pdf.

Matrone:1993:LPC

[MSP93] A. Matrone, P. Schiano, and V. Puoti. LINDA and PVM: a comparison between two environments for parallel programming. Parallel Computing, 19(8):949– 957, August 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mysliwiec:1997:CAM

[MSS97] G. Mysliwiec, J. Sipowicz, and R. Schaefer. Control activities in message passing environment. Lecture Notes in Computer Science, 1332:143–150, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martins:1998:JIW

[MSS98] P. Martins, L. M. Silva, and J. Silva. A Java interface for WMPI. Lecture Notes in Computer Science, 1497:121–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martorell:2005:BGP

[MSW^{+}05] X. Martorell, N. Smeds, R. Walkup, J. R. Brunheroto, G. Almási, J. A. Gunnels, L. DeRose, J. Labarta, F. Escalé, J. Giménez, H. Servat, and J. E. Moreira. Blue Gene/L performance tools. IBM Journal of Research and Development, 49(2/3):407–424, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/martorell.
pdf.

Mossaiby:2017:OIH

[MSZG17] F. Mossaiby, A. Shojaei, M. Zaccariotto, and U. Galvanetto. OpenCL implementation of a high performance 3D peridynamic model on graphics accelerators. Computers and Mathematics with Applications, 74(8):1856– 1870, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122117304030.

Miei:1996:IER

[MT96] T. Miei and N. Takahashi.

Implementation and evaluation of a replay-based debugger for PVM programs. Transactions of the Information Processing Society of Japan, 37(7):1308–1319, July 1996. CODEN JS-GRD5. ISSN 0387-5806.

Mallon:2016:MUB

[MTK16] Damián A. Mallón, Guillermo L.

Taboada, and Lars Koesterke. MPI and UPC broadcast, scatter and gather algorithms in Xeon Phi. Concurrency and Computation: Practice and Experience, 28(8):2322–2340, June 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marin:1994:GAL

[MTSS94] F. J. Marin, O. Trelles-Salazar, and F. Sandoval. Genetic algorithms on LAN-Message passing architectures using PVM: Application to the routing problem. In Davidor et al. [DSM94], pages 534–545 (or 534–543??). ISBN 3-540-58484-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I535 1994.

Momeni:2015:EEO

[MTU^{+}15] Amir Momeni, Hamed Tabkhi, Yash Ukidave, Gunar Schirner, and David Kaeli. Exploring the efficiency of the OpenCL pipe semantic on an FPGA. ACM SIGARCH Computer Architecture News, 43(4):52–57, September 2015. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Mohr:2007:SPE

[MTW07] Bernd Mohr, Jesper Larsson Träff, and Joachim Worringen. Selected papers from EuroPVM/MPI 2006. Parallel Computing, 33(9): 593–594, September 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mohr:2006:RAP

[MTWD06] Bernd Mohr, Jesper Larsson Träff, Joachim Worringen, and Jack Dongarra, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 13th European PVM/MPI User’s Group Meeting Bonn, Germany, September 17–20, 2006 Proceedings, volume 4192 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CO-DEN LNCSD9. ISBN 3-540-39110-X (print), 3-540-39112-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:/

/www.springerlink.com/
content/978-3-540-39112-
8.

Muller:2001:SSO

[Mül01]  Matthias Müller.  Some

simple OpenMP optimization techniques. Lecture Notes in Computer Science, 2104:31–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2104/21040031.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2104/21040031.
pdf.

Muller:2002:SMB

[Mül02]  Matthias S. Müller.  A

shared memory benchmark in OpenMP. Lecture Notes in Computer Science, 2327: 380–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2327/23270380.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2327/23270380.
                                        pdf.

Muller:2003:OCB

[Mül03] Matthias S. Müller. An OpenMP compiler benchmark. Scientific Programming, 11(2):125–131, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Malakar:2017:DMO

[MV17] Preeti Malakar and Venkatram Vishwanath. Data movement optimizations for independent MPI I/O on the Blue Gene/Q. Parallel Computing, 61(??):35– 51, January 2017. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016781911630062X.

Mantas:2020:HOC

[MV20] José M. Mantas and Francesco Vecil. Hybrid OpenMP– CUDA parallel implementation of a deterministic solver for ultrashort DG-MOSFETs. The International Journal of High Performance Computing Applications, 34(1):81–102, January 1, 2020. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:/

/journals.sagepub.com/
doi/full/10.1177/1094342019879985.

Manis:1996:EPT

[MVTP96] G. Manis, C. Voliotis, P. Tsanakas, and G. Papakonstantinou. Enhancing PVM with threads in distributed programming. In Liddell et al. [LCHS96], pages 1013–?? ISBN 3-540-61142-8 (paperback). LCCN QA76.88 .H52 1996.

Muller:2010:SMA

[MvWL^{+}10] Matthias S. Müller, Matthijs van Waveren, Ron Lieberman, Brian Whitney, Hideki Saito, Kalyan Kumaran, John Baron, William C. Brantley, Chris Parrott, Tom Elken, Huiyu Feng, and Carl Ponder. SPEC MPI2007 — an application benchmark suite for parallel systems using MPI. Concurrency and Computation: Practice and Experience, 22 (2):191–205, February 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mehra:1995:AIM

[MVY95] P. Mehra, B. Van Voorst, and J. Yan. Automated instrumentation, monitoring and visualization of PVM programs. In Bailey et al. [BBG^{+}95], pages 832–837. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

McKinney:1993:MMI

[MW93] G. W. McKinney and J. T. West. Multiprocessing MCNP on an IBM RS/ 6000 cluster. Transactions of the American Nuclear Society, 68(pt.A):212–214, 1993. CO-DEN TANSAO. ISSN 0003-018X.

Mamontov:1998:AES

[MW98] Y. V. Mamontov and M. Willander. An algorithm to evaluate spectral densities of high-dimensional stationary diffusion stochastic processes with non-linear coefficients: The general scheme and issues on implementation with PVM. Lecture Notes in Computer Science, 1541:315–321, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matwiejew:2021:QFP

[MW21] Edric Matwiejew and Jingbo Wang. QSW MPI: a framework for parallel simulation of quantum stochastic walks. Computer Physics Communications, 260(??): Article 107724, March 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465520303581.

Manegold:1997:QBM

[MWG97] S. Manegold, F. Waas, and D. Gudlat. In quest of the bottleneck — monitoring parallel database systems. Lecture Notes in Computer Science, 1332:277–284, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morton:1995:LLP

[MWO95] Don Morton, Kefei Wang, and David O. Ogbe. Lessons learned in porting Fortran/ PVM code to the Cray T3D. IEEE parallel and distributed technology: systems and applications, 3(1): 4–11, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Maleki:2016:HOT

[MYB16] Sepideh Maleki, Annie Yang, and Martin Burtscher. Higher-order and tuplebased massively-parallel prefix sums. ACM SIG-PLAN Notices, 51(6):539– 552, June 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mercan:2019:CCH

[MYK19] H. Mercan, C. Yilmaz, and K. Kaya. CHiP: A configurable hybrid parallel covering array constructor. IEEE Transactions on Software Engineering, 45(12): 1270–1291, December 2019. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Ma:2021:CSB

[MYL21] Wenpeng Ma, Wu Yuan, and Xiazhen Liu. A comparative study of block incomplete sparse approximate inverses preconditioning on Tesla K20 and V100 GPUs. Algorithms (Basel), 14(7), July 2021. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/
14/7/204.

Maly:1993:DCP

[MZK93] K. Maly, M. Zubair, and S. Kelbar. Distributed computing with parallel networking. In IEEE [IEE93d], pages 375–379. ISBN 0-8186-4430-3. LCCN QA76.9.D5I335 1993. IEEE catalog no. 93TH0574-4.

Mu:2020:OOB

[MZLS20] Jiandong Mu, Wei Zhang, Hao Liang, and Sharad Sinha. Optimizing OpenCLbased CNN design on FPGA with comprehensive design space exploration and collaborative performance modeling. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 13(3):13:1–13:28, September 2020. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https://

dl.acm.org/doi/10.1145/
3397514.

Nikolopoulos:2001:SID

[NA01] Dimitrios S. Nikolopoulos and Eduard Ayguadé. A study of implicit data distribution methods for OpenMP using the SPEC benchmarks. Lecture Notes in Computer Science, 2104: 115–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2104/21040115.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2104/21040115.
pdf.

Nikolopoulos:2001:EMA

[NAAL01] D. S. Nikolopoulos, E. Artiaga, E. Ayguadé, and J. Labarta. Exploiting memory affinity in OpenMP through schedule reuse. ACM SIGARCH Computer Architecture News, 29(5):49– 55, December 2001. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Nagle:2005:BRM

[Nag05] Dan Nagle. Book review:

MPI — The Complete Reference, Vol. 1, The MPI Core, 2nd ed., Scientific and Engineering Computation Series, by Marc Snir, Steve Otto, Steven Huss– Lederman, David Walker and Jack Dongarra. Scientific Programming, 13(1):57– 63, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nicolescu:1999:PWA

[NAJ99] C. Nicolescu, B. Albers, and P. Jonker. Parallel watershed algorithm on images from cranial CT-scans using PVM and MPI on a distributed memory system. In Dongarra et al. [DLM99], pages 418–425. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Nakajima:2003:PIS

[Nak03] Kengo Nakajima. Parallel iterative solvers of GeoFEM with selective blocking preconditioning for nonlinear contact problems on the Earth Simulator. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://
www.sc-conference.org/
sc2003/inter_cal/inter_
cal_detail.php?eventid=
10703#1;   http://www.
sc-conference.org/sc2003/
paperpdfs/pap155.pdf.

Nakajima:2005:PIS

[Nak05a] Kengo Nakajima. Parallel iterative solvers for finiteelement methods using an OpenMP/MPI hybrid programming model on the Earth Simulator. Parallel Computing, 31(10–12):1048– 1065, October/December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nakajima:2005:TLH

[Nak05b] Kengo Nakajima. Threelevel hybrid vs. flat MPI on the Earth Simulator: Parallel iterative solvers for finite-element method. Applied Numerical Mathematics: Transactions of IMACS, 54(2):237–255, July 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Narashimhan:1995:IIF

[Nar95] V. L. Narashimhan, editor.

ICAPP 95. IEEE First International Conference on Algorithms and Architectures for Parallel Processing, Brisbane, Australia, 19– 21 April, 1995. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2018-2 (paperback), 0-7803-2019-0 (microfiche). LCCN QA76.6.I15 1995. Two volumes. IEEE catalog no. 95TH0682-5.

Nagel:1996:VVA

[NAW^{+}96] W. E. Nagel, A. Arnold, M. Weber, H. C. Hoppe, and K. Solchenbach. VAMPIR: Visualization and analysis of MPI resources. Supercomputer, 12(1):69–80, January 1996. CODEN SPCOEL. ISSN 0168-7875.

NicCanna:1996:LGS

[NB96] C. Nic Canna and C. J. Bean. Larger grids and shorter wall-clock times on a parallel virtual machine (PVM) — an example using a finite difference wave simulation algorithm. In Abrahart [Abr96], pages 2– ?? ISBN ???? LCCN ????

Nickolls:2008:SPP

[NBGS08] John Nickolls, Ian Buck, Michael Garland, and Kevin Skadron. Scalable parallel programming with CUDA. ACM Queue: Tomorrow’s Computing Today, 6(2):40– 53, March 2008. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Neyman:1999:ERP

[NBK99] M. Neyman, M. Bukowski, and P. Kuzora. Efficient replay of PVM programs. In Dongarra et al. [DLM99], pages 83–90. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Nguyen:2012:BTM

[NCB^{+}12] Tan Nguyen, Pietro Cicotti, Eric Bylaska, Dan Quinlan, and Scott B. Baden. Bamboo: translating MPI applications to a latencytolerant, data-driven form. In Hollingsworth [Hol12], pages 39:1–39:?? ISBN 1-4673-0804-8. URL http:

//conferences.computer.
org/sc/2012/papers/1000a032.
pdf.

Nguyen:2017:ATM

[NCB^{+}17] Tan Nguyen, Pietro Cicotti, Eric Bylaska, Dan Quinlan, and Scott Baden. Automatic translation of MPI source into a latencytolerant, data-driven form. Journal of Parallel and Distributed Computing, 106(??): 1–13, August 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731517300771.

Nobari:2012:SPM

[NCKB12] Sadegh Nobari, Thanh-Tung Cao, Panagiotis Karras, and Stéphane Bressan. Scalable parallel minimum spanning forest computation. ACM SIGPLAN Notices, 47 (8):205–214, August 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Neophytou:1998:NDJ

[NE98] N. Neophytou and P. Evripidou. Net-dbx: a Java powered tool for interactive debugging of MPI programs across the Internet. Lecture Notes in Computer Science, 1470:181–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neophytou:2001:NDW

[NE01] Neophytos Neophytou and Paraskevas Evripidou. Netdbx: a Web-based debugger of MPI programs over low-bandwidth lines. IEEE Transactions on Parallel and Distributed Systems, 12(9):986–995, September 2001. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/td/books/
                                        td2001/pdf/l0986.pdf;
                                        http://www.computer.org/
                                        tpds/td2001/l0986abs.htm.

Nelson:1993:PPP

[Nel93] M. L. Nelson. PVM provides power in the public domain. Parallelogram, 53:20– 21, May–June 1993. CO-DEN PRALEH. ISSN 0953-7252.

Neugebauer:2017:PAR

[NEM17] Olaf Neugebauer, Michael Engel, and Peter Marwedel. A parallelization approach for resource-restricted embedded heterogeneous MP-SoCs inspired by OpenMP. The Journal of Systems and Software, 125(??):439– 448, March 2017. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL /

/www.sciencedirect.com/
science/article/pii/S0164121216301534.

Nesterov:2010:SPT

[Nes10] Oleksandr Nesterov. A simple parallelization technique with MPI for ocean circulation models. Journal of Parallel and Distributed Computing, 70(1):35–44, January 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Neun:1994:UPB

[Neu94] W. Neun. Using PVM based software for parallel computation in computer algebra. In Calmet [Cal94], pages 46–

51. ISBN ???? LCCN ????

Neyman:2000:CDA

[Ney00]  Marcin Neyman.  Com-

parison of different approaches to trace PVM program execution. Lecture Notes in Computer Science, 1908:274–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080274.htm;

                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1908/19080274.
                                        pdf.

Nordling:1994:SOD

[NF94] P. Nordling and P. Fritzson. Solving ordinary differential equations on parallel computers — applied to dynamic rolling bearings simulation. In Dongarra and Wasniewski [DW94], pages 397–415. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1994. DM104.00.

Nunez:2010:NTS

[NFG^{+}10] Alberto Núñez, Javier Fernández, Jose D. Garcia, Félix Garcia, and Jesús Carretero. New techniques for simulating high performance MPI applications on large storage networks. The Journal of Supercomputing, 51 (1):40–57, January 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0920-8542&volume=
                                        51&issue=1&spage=40.

Nieplocha:1998:CHP

[NFK98] Jarek Nieplocha, Ian Foster, and Rick A. Kendall. ChemIO: High performance parallel I/O for computational chemistry applications. The International Journal of High Performance Computing Applications, 12(3):345–363, Fall 1998. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:/

/journals.sagepub.com/
doi/pdf/10.1177/109434209801200304.

Nguyen:2008:GG

[Ngu08] Hubert Nguyen, editor.

GPU gems 3, volume 3 of GPU gems. Addison-Wes-ley, Reading, MA, USA, 2008. ISBN 0-321-51526-9. l + 942 pp. LCCN T385 .G6882 2008. URL http://

www.loc.gov/catdir/toc/
ecip0720/2007023985.html.

Nguyen:1995:SPI

[NH95] D. Nguyen and B. Hillberg. Simulations of pinhole imaging for AXAF: Distributed processing using the MPI standard. In Shaw et al. [SPH95], pages 361– 366 (or 361–363??). ISBN 0-937707-96-1. ISSN 1080-7926. LCCN QB51.3.E43 A87 1994.

Norden:2002:OVM

[NHT02] M. Nordén, S. Holmgren, and M. Thuné. OpenMP versus MPI for PDE solvers based on regular sparse numerical operators. Lecture Notes in Computer Science, 2331:681–??, 2002.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2331/23310681.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2331/23310681.
                                        pdf.

Norden:2006:OVM

[NHT06] Markus Nordén, Sverker Holmgren, and Michael Thuné. OpenMP versus MPI for PDE solvers based on regular sparse numerical operators. Future Generation Computer Systems, 22(1– 2):194–203, January 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nakano:2002:SCG

[NIO^{+}02] Hirofumi Nakano, Kazuhisa Ishizaka, Motoki Obata, Keiji Kimura, and Hironori Kasahara. Static coarse grain task scheduling with cache optimization using OpenMP. Lecture Notes in Computer Science, 2327:479–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2327/23270479.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/

0558/papers/2327/23270479.
pdf.

Nakano:2003:SCG

[NIO^{+}03] Hirofumi Nakano, Kazuhisa Ishizaka, Motoki Obata, Keiji Kimura, and Hironori Kasahara. Static coarse grain task scheduling with cache optimization using OpenMP. International Journal of Parallel Programming, 31(3):211– 223, June 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL /ips/frames/
Refs/referenceskapmain.
asp?J=4773&I=33&A=4&LK=
NM;   http://ipsapp007.
kluweronline.com/content/
getfile/4773/33/4/abstract.
htm;   http://ipsapp007.
kluweronline.com/content/
getfile/4773/33/4/fulltext.
pdf.

Nitsche:2000:TCM

[Nit00] Thomas Nitsche. Thread communication over MPI. Lecture Notes in Computer Science, 1908:145–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080145.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080145.
pdf.

Nicolescu:2001:DTP

[NJ01] Cristina Nicolescu and Pieter Jonker. A data and task parallel image processing environment. Lecture Notes in Computer Science, 2131:393–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2131/21310393.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2131/21310393.
                                        pdf.

Norden:2007:DDM

[NLRH07] Markus Nordén, Henrik Löf, Jarmo Rantakokko, and Sverker Holmgren. Dynamic data migration for structured AMR solvers. International Journal of Parallel Programming, 35(5): 477–491, October 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        35&issue=5&spage=477.

Nadeau:1995:SVR

[NM95] David R. Nadeau and John L. Moreland, editors. 1995 Symposium on the Virtual Reality Modeling Language, VRML ’95, San Diego, California, December 14–15, 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Novotny:1995:BRA

[NMC95] Mark Novotny, Susan McKay, and Wolfgang Christian. Book review: Al Geist, Adam Beguelin, Jack Dongarra, Weicheng Jiang, Robert Manchek, and Vaidy Sunderam, PVM — Parallel Virtual Machine: a Users’ Guide and Tutorial for Networked Parallel Computing. Computers in Physics, 9 (6):607–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4823450.

Nomura:2014:PAM

[NMS^{+}14] Shimpei Nomura, Takuji Mitsuishi, Jun Suzuki, Yuki Hayashi, Masaki Kan, and Hideharu Amano. Performance analysis of the multi-GPU system with ExpEther. ACM SIGARCH Computer Architecture News, 42(4):9– 14, September 2014. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Nanayakkara:1993:PIR

[NMW93] A. Nanayakkara, D. Moncrieff, and S. Wilson. Performance of IBM RISC System/6000 workstation clusters in a quantum chemical application. Parallel Computing, 19(9):1053– 1062, September 1993. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nupairoj:1995:PES

[NN95] N. Nupairoj and L. M. Ni. Performance evaluation of some MPI implementations on workstation clusters. In IEEE [IEE95j], pages 98– 105. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Nishitani:2000:IEO

[NNON00] Yasunori Nishitani, Kiyoshi Negishi, Hiroshi Ohta, and Eiji Nunohiro. Implementation and evaluation of OpenMP for Hitachi SR8000. Lecture Notes in Computer Science, 1940: 391–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1940/19400391.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1940/19400391.
                                        pdf.

Nakajima:2002:PISb

[NO02a] Kengo Nakajima and Hiroshi Okuda. Parallel iterative solvers for unstructured grids using a directive/MPI hybrid programming model for the GeoFEM platform on SMP cluster architectures. Concurrency and Computation: Practice and Experience, 14(6–7):411–429, May/June 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/94515747/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=94515747{\&}PLACEBO=
IE.pdf.

Nakajima:2002:PISa

[NO02b] Kengo Nakajima and Hiroshi Okuda. Parallel iterative solvers for unstructured grids using an OpenMP/MPI hybrid programming model for the GeoFEM platform on SMP cluster architectures. Lecture Notes in Computer Science, 2327:437–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2327/23270437.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2327/23270437.
pdf.

Noble:2008:GMY

[Nob08] Michael S. Noble. Getting more from your multicore: exploiting OpenMP from an open-source numerical scripting language. Concurrency and Computation: Practice and Experience, 20(16):1877–1891, November 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Novotny:1995:BPP

[Nov95] Mark Novotny. BOOKS: PVM — parallel virtual machine: a users’ guide and tutorial for networked parallel computing. Computers in Physics, 9(6):607– ??, ???? 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Nemer-Preece:1994:LBH

[NP94] Nicole Anne Nemer-Preece. Load balancing the heat equation in a heterogeneous environment with PVM. M.s. thesis, University of Missouri, Rolla, Rolla, MO, USA, 1994. viii + 52 pp.

Nguyen:2012:SCS

[NP12] Donald Nguyen and Keshav Pingali. Synthesizing concurrent schedulers for irregular algorithms. ACM SIG-PLAN Notices, 47(4):333– 344, April 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nikolopoulos:2000:TRD

[NPP^{+}00a] Dimitrios S. Nikolopoulos, Theodore S. Papatheodorou, Constantine D. Polychronopoulos, et al. A transparent runtime data distribution engine for OpenMP. Scientific Programming, 8 (3):143–162, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nikolopoulos:2000:DDN

[NPP^{+}00b] Dimitrios S. Nikolopoulos, Theodore S. Papatheodorou, Constantine D. Polychronopoulos, Jesus Labarta, and Eduard Ayguade. Is data distribution necessary in OpenMP? In ACM [ACM00], page 68. URL

http://www.sc2000.org/
proceedings/techpapr/papers/
pap192.pdf.

Nikolopoulos:2000:LTD

[NPP^{+}00c] Dimitrios S. Nikolopoulos, Theodore S. Papatheodorou, Constantine D. Polychronopoulos, Jesús Labarta, and Eduard Ayguadé. Leveraging transparent data distribution in OpenMP via user-level dynamic page migration. Lecture Notes in Computer Science, 1940: 415–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/

                                        bibs/1940/19400415.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1940/19400415.
                                        pdf.

Nikolopoulos:2000:ULR

[NPP^{+}00d] Dimitrios S. Nikolopoulos, Theodore S. Papatheodorou, Constantine D. Polychronopoulos, Jesús Labarta, and Eduard Ayguadé. UPM LIB: a runtime system for tuning the memory performance of OpenMP programs on scalable sharedmemory multiprocessors. Lecture Notes in Computer Science, 1915:85–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1915/19150085.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1915/19150085.
                                        pdf.

Notz:2012:GBS

[NPS12] Patrick K. Notz, Roger P. Pawlowski, and James C. Sutherland. Graph-based software design for managing complexity and enabling concurrency in multiphysics PDE software. ACM Transactions on Mathematical Software, 39(1):1:1–1:21, November 2012. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Naranjo:2020:ASC

[NRdA^{+}20] Diana M. Naranjo, Sebastián Risco, Carlos de Alfonso, Alfonso Pérez, Ignacio Blanquer, and Germán Moltó. Accelerated serverless computing based on GPU virtualization. Journal of Parallel and Distributed Computing, 139(??): 32–42, May 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731519303533.

Nagaraj:1991:MHL

[NS91] U. Nagaraj and U. S. Shukla.

MK: a high level interface for message passing. In Bhavsar and Gujar [BG91], pages 493–502. ISBN 0-920114-14-

8. LCCN QA76.88.S87 1991.

Naumenko:2016:ACT

[NS16] Mikhail A. Naumenko and Vyacheslav V. Samarin. Application of CUDA technology to calculation of ground states of few-body nuclei by Feynman’s continual integrals method. Supercomputing Frontiers and Innovations, 3(2):80–95, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic). URL http:/

/superfri.org/superfri/
article/view/102.

Nandal:2020:NSG

[NS20] P. Nandal and R. P. Sharma. Numerical simulation on GPUs with CUDA to study nonlinear dynamics of whistler wave and its turbulent spectrum in radiation belts. Computer Physics Communications, 254(??):Article 107214, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465520300497.

Nascimento:2007:DDS

[NSBR07] Aline P. Nascimento, Alexandre C. Sena, Cristina Boeres, and Vinod E. F. Rebello. Distributed and dynamic self-scheduling of parallel MPI Grid applications. Concurrency and Computation: Practice and Experience, 19(14):1955–1974, September 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nadal-Serrano:2016:PSC

[NSLV16] Jose M. Nadal-Serrano and Marisa Lopez-Vallejo. A performance study of CUDA UVM versus manual optimizations in a realworld setup: Application to a Monte Carlo waveparticle event-based interaction model. IEEE Transactions on Parallel and Distributed Systems, 27(6): 1579–1588, June 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07175058-
abs.html.

Nukada:2012:SMG

[NSM12] Akira Nukada, Kento Sato, and Satoshi Matsuoka. Scalable multi-GPU 3-D FFT for TSUBAME 2.0 supercomputer. In Hollingsworth [Hol12], pages 44:1–44:?? ISBN 1-4673-0804-8. URL

http://conferences.computer.
org/sc/2012/papers/1000a044.
pdf.

Neuberger:2012:MIS

[NSS12] John M. Neuberger, Nándor Sieben, and James W. Swift. An MPI implementation of a self-submitting parallel job queue. International Journal of Parallel Programming, 40(4):443–464, August 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=4&spage=443.

Nandivada:2013:TFO

[NSZS13] V. Krishna Nandivada, Jun Shirako, Jisheng Zhao, and Vivek Sarkar. A transformation framework for optimizing task-parallel programs. ACM Transactions on Programming Languages and Systems, 35(1): 3:1–3:??, April 2013. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Nogueira:2016:BBW

[NTR16] David Nogueira, Pedro Tomas, and Nuno Roma. BowMapCL: Burrows–Wheeler mapping on multiple heterogeneous accelerators. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 13(5):926–938, September 2016. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Norcen:2005:HPJ

[NU05] Roland Norcen and Andreas Uhl. High performance JPEG 2000 and MPEG-4 VTC on SMPs using OpenMP. Parallel Computing, 31(10–12):1082–1098, October/December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nitsche:1998:FMP

[NW98] T. Nitsche and W. Webers. Functional message passing with OPAL-MPI. Lecture Notes in Computer Science, 1497:281–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:2021:EMA

[NWT21] Truong Thao Nguyen, Mohamed Wahib, and Ryousei Takano. Efficient MPI-AllReduce for largescale deep learning on GPUclusters. Concurrency and Computation: Practice and Experience, 33(12):e5574:1– e5574:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ng:2012:STT

[NYNT12] Nicholas Ng, Nobuko Yoshida, Xin Yu Niu, and Kuen Hung Tsoi. Session types: towards safe and fast reconfigurable programming. ACM SIGARCH Computer Architecture News, 40(5):22–27, December 2012. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). HEART ’12 conference proceedings.

Nguyen:1994:DCE

[NZZ94] S. T. Nguyen, B. J.

Zook, and Xiaodong Zhang. Distributed computation of electromagnetic scattering problems using finitedifference time-domain decompositions. In IEEE [IEE94g], pages 85–93. ISBN 0-8186-6395-2. LCCN QA76.9.D5I328 1994. IEEE catalog no. 94TH0667-6.

Omar:2017:PSF

[OA17] Cyrus Omar and Jonathan Aldrich. Programmable semantic fragments: the design and implementation of typy. ACM SIGPLAN Notices, 52(3):81–92, March 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oberhuber:1996:MNP

[Obe96]  M. Oberhuber.  Manag-

ing nondeterminism in PVM programs. In Bode et al. [BDLS96], pages 347–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Orr:2015:SUR

[OCY^{+}15] Marc S. Orr, Shuai Che, Ayse Yilmazer, Bradford M. Beckmann, Mark D. Hill, and David A. Wood. Synchronization using remotescope promotion. ACM SIGARCH Computer Architecture News, 43(1):73– 86, March 2015. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Okulicka-Dluzewska:2001:PFE

[OD01] Felicja Okulicka-Dłuzewska. Parallelization of finite element package by MPI library. Lecture Notes in Computer Science, 2131:

427–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310427.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310427.
pdf.

Olivier:2012:CMW

[OdSSP12] Stephen L. Olivier, Bronis R. de Supinski, Martin Schulz, and Jan F. Prins. Characterizing and mitigating work time inflation in task parallel programs. In Hollingsworth [Hol12], pages 65:1–65:?? ISBN 1-4673-0804-8. URL http:

//conferences.computer.
org/sc/2012/papers/1000a066.
pdf.

Oed:1993:CRM

[Oed93] Wilfried Oed. The Cray Research massively parallel processor system CRAY T3D. Technical report, Cray Research GmbH, München, Germany, November 15 1993.

Ong:2000:PCL

[OF00] Hong Ong and Paul A. Farrell. Performance comparison of LAM/MPI, MPICH, and MVICH on a Linux cluster connected by a Gigabit Ethernet network. In USENIX [USE00], page ??

ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
ong.html.

Owaida:2015:EDS

[OFA^{+}15] Muhsen Owaida, Gabriel Falcao, Joao Andrade, Christos Antonopoulos, Nikolaos Bellas, Madhura Purnaprajna, David Novo, Georgios Karakonstantis, Andreas Burg, and Paolo Ienne. Enhancing design space exploration by extending CPU/GPU specifications onto FPGAs. ACM Transactions on Embedded Computing Systems, 14(2):33:1– 33:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Otten:2016:MOI

[OGM^{+}16] Matthew Otten, Jing Gong, Azamat Mametjanov, Aaron Vose, John Levesque, Paul Fischer, and Misun Min. An MPI/OpenACC implementation of a high-order electromagnetics solver with GPUDirect communication. The International Journal of High Performance Computing Applications, 30(3):320– 334, August 2016. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Otero:2019:OAA

[OGM^{+}19] Evelyn Otero, Jing Gong, Misun Min, Paul Fischer, Philipp Schlatter, and Erwin Laure. OpenACC acceleration for the P_{N}–P_{N−2} algorithm in Nek5000. Journal of Parallel and Distributed Computing, 132 (??):69–78, October 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731518305549.

Ortega:2019:CAC

[OHG19] G. Ortega, E. M. T. Hendrix, and I. García. A CUDA approach to compute perishable inventory control policies using value iteration. The Journal of Supercomputing, 75(3):1580– 1593, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/content/pdf/
10.1007/s11227-018-2692-
z.pdf.

Okitsu:2010:HPC

[OIH10] Yusuke Okitsu, Fumihiko Ino, and Kenichi Hagihara. High-performance cone beam reconstruction using CUDA compatible GPUs. Parallel Computing, 36(2–3):129–141, February/March 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ohara:2006:MMP

[OIS^{+}06] M. Ohara, H. Inoue, Y. Sohda, H. Komatsu, and T. Nakatani. MPI microtask for programming the Cell Broadband Engine^{T M} processor. IBM Systems Journal, 45(1):85–102, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http:

                                        //www.research.ibm.com/
                                        journal/sj/451/ohara.html.

Oh:2012:MOO

[OKM12] Kwang Jin Oh, Ji Hoon Kang, and Hun Joo Myung. mm par2.0: An objectoriented molecular dynamics simulation program parallelized using a hierarchical scheme with MPI and OPENMP. Computer Physics Communications, 183(2):440–441, February 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0010465511003407.

Oakley:1995:ADR

[OKW95] D. R. Oakley, N. F. Knight, Jr., and D. D. Warner. Adaptive dynamic relaxation algorithm for nonlinear hyperelastic structures. III. Parallel implementation. Computer Methods in Applied Mechanics and Engineering, 126(1-2):111–129, September 1995. CODEN CMMECC. ISSN 0045-7825, 0374-2830.

Orlando:2005:PSP

[OL05] Salvatore Orlando and Domenico Laforenza. Preface: Selected papers from the EU-ROPVM/MPI 2003 Conference, Venice, Italy, 29 September–2 October 2003. The International Journal of High Performance Computing Applications, 19(1): 47, Spring 2005. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/19/
1/47.full.pdf+html.

Oldehoeft:2002:SIS

[Old02] Rod Oldehoeft, editor. Special issue on software for high-performance systems: papers from the symposium of the Los Alamos Computer Science Institute, held in Santa Fe, NM, USA on October 15–18, 2001, volume 23(1) of The journal of supercomputing. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ong:2001:SUC

[OLG01] Emil Ong, Ewing Lusk, and William Gropp. Scalable Unix commands for parallel processors: a highperformance implementation. Lecture Notes in Computer Science, 2131: 410–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310410.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310410.
pdf.

Oger:2016:DMM

[OLG^{+}16] G. Oger, D. Le Touzé, D. Guibert, M. de Leffe, J. Biddiscombe, J. Soumagne, and J.-G. Piccinali. On distributed memory MPI-based parallelization of SPH codes in massive HPC context. Computer Physics Communications, 200(??):1–14, March 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465515003070.

Olszewski:1995:TCC

[Ols95] Luke Olszewski. A timing comparison of the conjugate gradient and Gauss– Seidel parallel algorithms in a one-dimensional flow equation using PVM. In ACM [ACM95a], pages 205–212.

ISBN 0-89791-747-2. LCCN ????

Olukotun:2014:BPP

[Olu14] Kunle Olukotun. Beyond parallel programming with domain specific languages. ACM SIGPLAN Notices, 49 (8):179–180, August 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ogawa:1996:OOM

[OM96] Hirotaka Ogawa and Satoshi Matsuoka. OMPI: Optimizing MPI programs using partial evaluation. In ACM [ACM96c], page ?? ISBN 0-89791-854-1. LCCN QA 76.88 S8573 1996. URL

http://www.supercomp.org/
sc96/proceedings/SC96PROC/
OGAWA/INDEX.HTM. ACM

Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Ozgun:2009:PCB

[OMK09] Ozlem Ozgun, Raj Mittra, and Mustafa Kuzuoğlu. Parallelized characteristic basis finite element method (CBFEM–MPI) — a noniterative domain decomposition algorithm for electromagnetic scattering problems. Journal of Computational Physics, 228(6): 2225–2238, April 1, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0021999108006293.

OBroin:2012:OIS

[ÓN12] Cathal Ó Broin and L. A. A. Nikolopoulos. An OpenCL implementation for the solution of the time-dependent Schrödinger equation on GPUs and CPUs. Computer Physics Communications, 183(10):2071–2080, October 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0010465512001774.

Ong:2002:MRS

[Ong02] Emil Ong. MPI Ruby: Scripting in a parallel environment. Computing in Science and Engineering, 4(4):78–82, July/ August 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/04/c4078abs.htm;
                                         http://csdl.computer.
                                        org/dl/mags/cs/2002/04/
                                        c4078.htm;   http://csdl.
                                        computer.org/dl/mags/cs/
                                        2002/04/c4078.pdf.

OBrien:2008:SOC

[OOS^{+}08] Kevin OB́rien, Kathryn OB́rien, Zehra Sura, Tong Chen, and Tao Zhang. Supporting OpenMP on Cell. International Journal of Parallel Programming, 36(3):289–311, June 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=3&spage=289.

Orlando:1998:MBR

[OP98] S. Orlando and R. Perego.

An MPI-based run-time support to coordinate HPF tasks. Lecture Notes in Computer Science, 1497:289–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olivier:2010:COO

[OP10] Stephen L. Olivier and Jan F. Prins. Comparison of OpenMP 3.0 and other task parallel frameworks on unbalanced task graphs. International Journal of Parallel Programming, 38(5– 6):341–360, October 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=5&spage=341.

Oh:2019:HPT

[OPJ^{+}19] S. Oh, N. Park, J. Jang, L. Sael, and U. Kang. Highperformance Tucker factorization on heterogeneous platforms. IEEE Transactions on Parallel and Distributed Systems, 30(10): 2237–2248, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

ODowd:2006:WGM

[OPM06] Padraig J. O’Dowd, Adarsh Patil, and John P. Morrison. WebCom-G and MPICH-G2 jobs. Scalable Computing: Practice and Experience, 7(3):75–86, September 2006. CODEN ???? ISSN 1895-1767.

                                        URL http://www.scpe.
                                        org/vols/vol07/no3/SCPE_
                                        7_3_07.pdf;   http://www.
                                        scpe.org/vols/vol07/no3/
                                        SCPE_7_3_07.zip.

Orlando:2000:MDT

[OPP00] S. Orlando, P. Palmerini, and R. Perego. Mixed data and task parallelism with HPF and PVM. Cluster Computing, 3(3):201–213, 2000. CODEN ???? ISSN 1386-7857.

Olivier:2012:OTS

[OPW^{+}12] Stephen L. Olivier, Allan K. Porterfield, Kyle B. Wheeler, Michael Spiegel, and Jan F. Prins. OpenMP task scheduling strategies for multicore NUMA systems. The International Journal of High Performance Computing Applications, 26(2):110– 124, May 2012. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/26/
2/110.full.pdf+html.

Oliveira:2012:CCO

[ORA12] Rafael Sachetto Oliveira, Bernardo Martins Rocha, and Ronan Mendonça Amorim. Comparing CUDA, OpenCL and OpenGL implementations of the cardiac monodomain equations. Lecture Notes in Computer Science, 7204:111–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/
10.1007/978-3-642-31500-
8_12/.

Overeinder:1997:BCD

[OS97] B. J. Overeinder and P. M. A.

Sloot. Breaking the curse of dynamics by task migration: Pilot experiments in the Polder Metacomputer. Lecture Notes in Computer Science, 1332:194–207, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ostrand:1994:PIS

[Ost94] Thomas Ostrand, editor.

Proceedings of the 1994 International Symposium on Software Testing and Analysis (ISSTA): August 17– 19, 1994, Seattle, Washington, USA, ACM SIGSOFT Software Engineering Notes. ACM Press, New York, NY 10036, USA, 1994. CO-DEN SFENDP. ISBN 0-89791-683-2. ISSN 0163-5948. LCCN QA76.76.T48 I58 1994.

Obrecht:2015:PEO

[OTK15] Christian Obrecht, Bernard Tourancheau, and Frédéric Kuznik. Performance evaluation of an OpenCL implementation of the Lattice Boltzmann Method on the Intel Xeon Phi. Parallel Processing Letters, 25(3): 1541001, September 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Otto:1993:PAC

[Ott93] S. W. Otto. Parallel array classes and lightweight sharing mechanisms. Scientific Programming, 2(4):203– 216, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Otto:1994:PVM

[Ott94] S. W. Otto. Processor virtualization and migration for PVM. In Dongarra and Tourancheau [DT94], pages 66–75. ISBN 0-89871-343-9. LCCN QA76.58.I568 1994.

Otto:1992:MAP

[OW92] S. W. Otto and M. Wolfe. The MetaMP approach to parallel programming. In Siegel [Sie92a], pages 562–

565. ISBN 0-8186-2772-

7. LCCN QA76.58.S95 1992. IEEE catalog no. 92CH3185-6.

Oldfield:1998:EPS

[OWO98] Ron A. Oldfield, David E.

Womble, and Curtis C. Ober. Efficient parallel I/O in seismic processing. The International Journal of High Performance Computing Applications, 12(3): 333–344, Fall 1998. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

//journals.sagepub.com/
doi/pdf/10.1177/109434209801200303.

Ouenes:1995:PRA

[OWSA95] A. Ouenes, W. W. Weiss, J. A. Sultan, and J. Anwar. Parallel reservoir automatic history matching using a network of workstations and PVM. In Anonymous [Ano95d], pages 125–

134. ISBN ???? LCCN ????

Pacheco:1997:PPM

[Pac97] Peter S. Pacheco. Parallel programming with MPI. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1997. ISBN 1-55860-339-5. xxii + 418 pp. LCCN QA76.642 .P3 1997.

Pereira:2017:SBC

[PAdS^{+}17] Phillipe Pereira, Higo Albuquerque, Isabela da Silva, Hendrio Marques, Felipe Monteiro, Ricardo Ferreira, and Lucas Cordeiro. SMTbased context-bounded model checking for CUDA programs. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Panda:1995:GRW

[Pan95a] D. K. Panda. Global reduction in wormhole k-ary n-cube networks with multidestination exchange worms. In IEEE [IEE95f], pages 652–659. ISBN 0-8186-7074-6. LCCN QA 76.58 I56 1995. IEEE catalog no. 95TH8052.

Panda:1995:IDE

[Pan95b] D. K. Panda. Issues in designing efficient and practical algorithms for collective communication on wormhole-routed systems. In Agrawal [Agr95a], pages 8– 15. ISBN 0-8493-2618-4. LCCN QA76.58.I34 1995.

Panda:2014:GAM

[Pan14] Dhabaleswar K. Panda. GPU-aware MPI on RDMAenabled clusters: Design, implementation and evaluation. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2595–2605, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06587715-
abs.html.

Parsons:1993:EDC

[Par93] I. Parsons. Evaluation of distributed communication systems. In Gawman et al. [GGK^{+}93], pages 956–970 vol.2. ISBN ???? LCCN QA76.76.S64 C378 1993 v.1-

2. Two volumes.

Pal:2014:PMH

[PARB14] Anirban Pal, Abhishek Agarwala, Soumyendu Raha, and Baidurya Bhattacharya. Performance metrics in a hybrid MPI–OpenMP based molecular dynamics simulation with short-range interactions. Journal of Parallel and Distributed Computing, 74(3):2203–2214, March 2014. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0743731513002505.

Patterson:1993:PPE

[Pat93] Christopher S. Patterson.

Parametric positron emission tomographic imaging using parallel virtual machine: with an example using myocardial blood flow analysis. M.s. thesis, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 1993. x + 132 pp.

Puzniakowski:2012:TOI

[PB12] Tadeusz Puźniakowski and Marek A. Bednarczyk. Towards an OpenCL implementation of ‘genetic algorithms’ on GPUs. Lecture Notes in Computer Science, 7053:190–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer.com/chapter/
                                        10.1007/978-3-642-25261-
                                        7_15/.

Pringle:2001:TPF

[PBC^{+}01] Gavin J. Pringle, Steven P. Booth, Hugh M. P. Couchman, Frazer R. Pearce, and Alan D. Simpson. Towards a portable, fast parallel AP^{3}M-SPH code: HY-DRA MPI. Lecture Notes in Computer Science, 2131: 360–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2131/21310360.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2131/21310360.
                                        pdf.

Pingali:1995:LCP

[PBG^{+}95] K. Pingali, U. Banerjee, D. Gelernter, A. Nicolau, and D. Padua, editors. Languages and compilers for parallel computing: 7th International Workshop, Ithaca, NY, USA, August 8– 10, 1994: proceedings, volume 892 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-58868-X. LCCN QA76.58 .W656 1994.

Plazek:1999:IIC

[PBK99] J. Plazek, K. Banas, and J. Kitowski. Implementation issues of computational fluid dynamics algorithms on parallel computers. In Dongarra et al. [DLM99], pages 349–355. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Plazek:2000:SCC

[PBK00] Joanna Płazek, Krzysztof Banas, and Jacek Kitowski. Scalable CFD computations using message-passing and distributed shared memory algorithms. Lecture Notes in Computer Science, 1908:282–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080282.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080282.
pdf.

Prasanna:1995:FIP

[PBPT95] Viktor K. Prasanna, V. P. Bhatkar, L. M. Patnaik, and S. K. Tripathi, editors. First IWPP parallel processing: proceedings of the First International Workshop on Parallel Processing (IWPP-94): December 26–31, 1994, Bangalore, India. Taka McGraw-Hill Pub. Co, New Delhi; New York, 1995. ISBN 0-07-462332-X. LCCN QA 76.58 I587 1994.

Puthukattukaran:1994:DIP

[PCS94] J. Puthukattukaran, S. Chalasani, and P. Senapathy. Design and implementation of parallel algorithms for gene-finding. In IEEE [IEE94g], pages 186–193. ISBN 0-8186-6395-2. LCCN QA76.9.D5I328 1994. IEEE catalog no. 94TH0667-6.

Peng:2014:IDI

[PCY14] Yi Peng, Li Chen, and Jun-Hai Yong. Importancedriven isosurface decimation for visualization of large simulation data based on OpenCL. Computing in Science and Engineering, 16(1): 24–32, January/February 2014. CODEN CSENFA. ISSN 1521-9615.

Poggi:1998:UPD

[PD98] Agostino Poggi and Giulio Destri. Using PVM to develop a distributed objectoriented language for heterogeneous processing. The Journal of Systems and Software, 40(2):139–150, February 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Plimpton:2011:MML

[PD11] Steven J. Plimpton and Karen D. Devine. MapReduce in MPI for large-scale graph algorithms. Parallel Computing, 37(9):610– 632, September 2011. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819111000172.

Pawliczek:2014:VED

[PDY14] Piotr Pawliczek, Witold Dzwinel, and David A. Yuen. Visual exploration of data by using multidimensional scaling on multicore CPU, GPU, and MPI cluster. Concurrency and Computation: Practice and Experience, 26 (3):662–682, March 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pennington:1995:DHC

[Pen95] R. L. Pennington. Distributed and heterogeneous computing. In Vandoni and Verkerk [VV95], pages 25–

57. ISBN 92-9083-069-7. CERN report 95-01.

Pernice:1996:RPP

[Per96] Michael Pernice. Review of “PVM: Parallel Virtual Machine. A User’s Guide and Tutorial for Networked Parallel Computing”. IEEE parallel and distributed technology: systems and applications, 4(1):84, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http:

                                        //dlib.computer.org/pd/
                                        books/pd1996/pdf/p1084.
                                        pdf.

Pernice:1997:BRM

[Per97]  Michael Pernice.  Book

review: MPI: The Complete Reference. IEEE Concurrency, 5(1):80–81, January/March 1997. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

                                        //dlib.computer.org/pd/
                                        books/pd1997/pdf/p1080.
                                        pdf.

Pereira:1999:PBI

[Per99] N. S. A. Pereira. A parallel N-body integrator using MPI. Lecture Notes in Computer Science, 1573: 627–639, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perepu:2021:OIP

[Per21] Pavan Kumar Perepu. OpenMP implementation of parallel longest common subsequence algorithm for mathematical expression retrieval. Parallel Processing Letters, 31(02):??, June 2021. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.
com/doi/10.1142/S0129626421500079.

Papagapiou:1999:NWD

[PES99] A. Papagapiou, P. Evripidou, and G. Samaras. Net-Console: a Web-based development environment for MPI programs. In Dongarra et al. [DLM99], pages 249–

256. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Petcu:1997:ISM

[Pet97]  D. Petcu.  Implementa-

tion of some multiprocessor algorithms for ODEs using PVM. Lecture Notes in Computer Science, 1332: 375–382, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petcu:2000:PDAa

[Pet00a] Dana Petcu. PVMaple: a distributed approach to cooperative work of Maple processes. Technical report, Westers University of Timisoara, Timisoara, Romania, May 2000. URL

                                        http://www.risc.uni-linz.
                                        ac.at/software/distmaple/
                                        misc/PVMaple.ps.gz.

Petcu:2000:PDAb

[Pet00b] Dana Petcu. PVMaple: a distributed approach to cooperative work of Maple processes. Lecture Notes in Computer Science, 1908: 216–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1908/19080216.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1908/19080216.
                                        pdf.

Petcu:2001:WMM

[Pet01] Dana Petcu. Working with multiple Maple kernels connected by Distributed Maple or PVMaple. Technical report, Westers University of Timisoara, Timisoara, Romania, March 2001.

                                        URL http://www.risc.
                                        uni-linz.ac.at/software/
                                        distmaple/misc/petcu2001.
                                        ps.gz.

Pharr:2005:GGP

[PF05] Matt Pharr and Randima Fernando, editors. GPU gems 2: programming techniques for high-performance graphics and general-purpose computation, volume 2 of GPU gems. Addison-Wes-ley, Reading, MA, USA, 2005. ISBN 0-321-33559-7 (hardcover). xlix + 814 pp. LCCN T385 .G688 2005.

URL http://www-docs.tu-
cottbus.de/bibliothek/
public/katalog/420569.
PDF;   http://www.loc.
gov/catdir/toc/ecip055/
2004030181.html.

Piernas:1997:APM

[PFG97] J. Piernas, A. Flores, and J. M. Garcia. Analyzing the performance of MPI in a cluster of workstations based on Fast Ethernet. Lecture Notes in Computer Science, 1332:17–24, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prabhu:2018:DRC

[PG18] Tarun Prabhu and William Gropp. DAME: Runtimecompilation for data movement. The International Journal of High Performance Computing Applications, 32(5):760–774, 2018. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

//journals.sagepub.com/
doi/full/10.1177/1094342017695444.

Pjesivac-Grbovic:2005:PAM

[PGAB^{+}05] J. Pjesivac-Grbovic, T. Angskun, G. Bosilca, G. E. Fagg, E. Gabriel, and J. J. Dongarra. Performance analysis of MPI collective operations. In IEEE [IEE05], pages 272a–272a. ISBN 0-7695-2312-9. LCCN ???? IEEE Computer Society Order Number P2312. Pjesivac-Grbovic:2007:PAM

[PGAB^{+}07]  Jelena  Pjesivac-Grbovic,

Thara Angskun, George Bosilca, Graham E. Fagg, Edgar Gabriel, and Jack J. Dongarra. Performance analysis of MPI collective operations. The Journal of Networks, Software Tools, and Cluster Computing, 10 (2):127–143, ???? 2007. ISSN 1386-7857.

Pjesivac-Grbovic:2007:MCA

[PGBF^{+}07]  Jelena  Pješivac-Grbović,

George Bosilca, Graham E. Fagg, Thara Angskun, and Jack J. Dongarra. MPI collective algorithm selection and quadtree encoding. Parallel Computing, 33(9): 613–623, September 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Prabhakar:2002:PCB

[PGC02] Achal Prabhakar, Vladimir Getov, and Barbara Chapman. Performance comparisons of basic OpenMP constructs. Lecture Notes in Computer Science, 2327: 413–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/

bibs/2327/23270413.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2327/23270413.
pdf.

Peng:2018:CDC

[PGD18] Yuanfeng Peng, Vinod Grover, and Joseph Devietti. CURD: a dynamic CUDA race detector. ACM SIG-PLAN Notices, 53(4):390– 403, April 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pessoa:2018:GAB

[PGdCJ^{+}18] Tiago Carneiro Pessoa, Jan Gmys, Francisco Heron de Carvalho Júnior, Nouredine Melab, and Daniel Tuyttens. GPU-accelerated backtracking using CUDA Dynamic Parallelism. Concurrency and Computation: Practice and Experience, 30 (9), May 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https:

//onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.
4374.

Poirier:2018:DAB

[PGF18] Carl Poirier, Benoit Gosselin, and Paul Fortier. DNA assembly with de Bruijn graphs using an FPGA platform. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 15 (3):1003–1009, May 2018. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Pervez:2010:FMA

[PGK^{+}10] Salman Pervez, Ganesh Gopalakrishnan, Robert M. Kirby, Rajeev Thakur, and William Gropp. Formal methods applied to high-performance computing software design: a case study of MPI one-sided communication-based locking. Software—Practice and Experience, 40(1):23–43, January ??, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pimentel-Garcia:2021:EIP

[PGPCK21] Ernesto Pimentel-García, Carlos Parés, Manuel J. Castro, and Julian Koellermeier. On the efficient implementation of PVM methods and simple Riemann solvers. Application to the Roe method for large hyperbolic systems. Applied Mathematics and Computation, 388(??):Article 125544, January 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300320305002.

Papakonstantinou:2013:ECC

[PGS^{+}13] Alexandros Papakonstantinou, Karthik Gururaj, John A. Stratton, Deming Chen, Jason Cong, and Wen-Mei W. Hwu. Efficient compilation of CUDA kernels for high-performance computing on FPGAs. ACM Transactions on Embedded Computing Systems, 13(2):25:1– 25:??, September 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pan:2010:CPS

[PHA10] Heidi Pan, Benjamin Hindman, and Krste Asanović. Composing parallel software efficiently with Lithe. ACM SIGPLAN Notices, 45(6):376–387, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pennycook:2011:PAH

[PHJM11] S. J. Pennycook, S. D. Hammond, S. A. Jarvis, and G. R. Mudalige. Performance analysis of a hybrid MPI/CUDA implementation of the NASLU benchmark. ACM SIGMETRICS Performance Evaluation Review, 38(4):23–29, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Protze:2022:MDT

[PHM^{+}22] Joachim Protze, Marc-André Hermanns, Matthias S.

Müller, Van Man Nguyen, Julien Jaeger, Emmanuelle Saillard, Patrick Carribault, and Denis Barthou. MPI detach — towards automatic asynchronous local completion. Parallel Computing, 109(??):??, March 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819121001022.

Power:2015:GGH

[PHO^{+}15] Jason Power, Joel Hestness, Marc S. Orr, Mark D. Hill, and David A. Wood. gem5gpu: A heterogeneous CPU– GPU simulator. IEEE Computer Architecture Letters, 14(1):34–36, January/June 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Pennycook:2013:IPP

[PHW^{+}13] S. J. Pennycook, S. D. Hammond, S. A. Wright, J. A. Herdman, I. Miller, and S. A. Jarvis. An investigation of the performance portability of OpenCL. Journal of Parallel and Distributed Computing, 73(11): 1439–1450, November 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731512001669.

Pierce:1994:NMP

[Pie94] P. Pierce. The NX message passing interface. Parallel Computing, 20(4):463– 480, April 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Prades:2020:MRU

[PIR^{+}20] Javier Prades, Baldomero Imbernón, Carlos Reaño, Jorge Peña-García, Jose Pedro Cerón-Carrasco, Federico Silla, and Horacio Pérez-Sánchez. Maximizing resource usage in multifold molecular dynamics with rCUDA. The International Journal of High Performance Computing Applications, 34(1):5–19, January 1, 2020. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:/

/journals.sagepub.com/
doi/full/10.1177/1094342019857131.

Papadopoulos:1998:DVS

[PK98] P. M. Papadopoulos and J. A. Kohl. Dynamic visualization and steering using PVM and MPI. Lecture Notes in Computer Science, 1497:297–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:2005:SOA

[PK05] Inho Park and Seon Wook Kim. Study of OpenMP applications on the InfiniBandbased software distributed shared-memory system. Parallel Computing, 31(10– 12):1099–1113, October/ December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Papadopoulos:2001:NRC

[PKB01] Philip M. Papadopoulos, Mason J. Katz, and Greg Bruno. NPACI rocks clusters: Tools for easily deploying and maintaining manageable high-performance Linux clusters. Lecture Notes in Computer Science, 2131:10–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310010.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310010.
pdf.

Paul:2006:TLF

[PKB06] Jerome L. Paul, Michal Kouril, and Kenneth A. Berman. A template library to facilitate teaching message passing parallel computing. In ACM [ACM06a], pages 464–468. ISBN 1-59593-259-3. ACM order number 457060.

Prabhakar:2016:GCH

[PKB^{+}16] Raghu Prabhakar, David Koeplinger, Kevin J. Brown, HyoukJoong Lee, Christopher De Sa, Christos Kozyrakis, and Kunle Olukotun. Generating configurable hardware from parallel patterns. ACM SIGPLAN Notices, 51(4):651–665, April 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Plank:1995:ADC

[PKD95] J. S. Plank, Youngbae Kim, and J. J. Dongarra. Algorithm-based diskless checkpointing for fault tolerant matrix operations. In IEEE [IEE95c], pages 351– 360. ISBN 0-8186-7079-7. LCCN QA 76.9 F38 I57 1995. IEEE catalog no. 95CB35823.

Preissl:2010:OCC

[PKE^{+}10] Robert Preissl, Alice Koniges, Stephan Ethier, Weixing Wang, and Nathan Wichmann. Overlapping communication with computation using OpenMP tasks on the GTS magnetic fusion code. Scientific Programming, 18 (3–4):139–151, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Periyathamby:1995:NSG

[PKYW95] U. Periyathamby, B. C.

Khoo, K. S. Yeo, and Q. X. Wang. A numerical simulation of the growth and collapse of vapour cavity near a free surface on distributed computing through PVM. In Bilger [Bil95], pages 815– 818. ISBN 0-86934-034-4. LCCN ????

Pruyne:1996:ICP

[PL96] Jim Pruyne and Miron Livny. Interfacing Condor and PVM to harness the cycles of workstation clusters. Future Generation Computer Systems, 12(1): 67–85, May 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Plachetka:2002:QTS

[Pla02] Tomas Plachetka. (quasi-) thread-safe PVM and (quasi- ) thread-safe MPI without active polling. Lecture Notes in Computer Science, 2474:296–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

                                        /link.springer.de/link/
                                        service/series/0558/bibs/
                                        2474/24740296.htm;   http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2474/24740296.pdf.

Park:2004:DID

[PLK^{+}04] K.-L. Park, H.-J. Lee, O.- Y. Kwon, S.-Y. Park, H.-W. Park, and S.-D. Kim. Design and implementation of a dynamic communication MPI library for the grid. International Journal of Computers and Applications, 26 (3):1–8, 2004. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:

//www.tandfonline.com/
doi/full/10.1080/1206212X.
2004.11441738.

Piriyakumar:2002:EFI

[PLR02] Douglas Antony Louis Piriyakumar, Paul Levi, and Rolf Rabenseifner. Enhanced file interoperability with parallel MPI file-I/O in image processing. Lecture Notes in Computer Science, 2474:174–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/
service/series/0558/bibs/
2474/24740174.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740174.pdf.

Pfenning:1995:OCP

[PM95] Jörg-Thomas Pfenning and Christoph Moll. Optimized communication patterns on workstation clusters. Parallel Computing, 21 (3):373–388, March 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=21&
issue=3&aid=964.

Piscaglia:1995:DOC

[PMM95] P. Piscaglia, B. Macq, and P. Maes. Distributed optimization of codebooks. Signal Processing: Image Communication, 7(3):211–223, September 1995. CODEN SPICEF. ISSN 0923-5965 (print), 1879-2677 (electronic).

Poulson:2013:ENF

[PMvdG^{+}13] Jack Poulson, Bryan Marker, Robert A. van de Geijn, Jeff R. Hammond, and Nichols A. Romero. Elemental: a new framework for distributed memory dense matrix computations. ACM Transactions on Mathematical Software, 39(2):13:1– 13:24, February 2013. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Pirk:2016:VVA

[PMZM16] Holger Pirk, Oscar Moll, Matei Zaharia, and Sam Madden. Voodoo — a vector algebra for portable database performance on modern hardware. Proceedings of the VLDB Endowment, 9(14):1707–1718, October 2016. CODEN ???? ISSN 2150-8097.

Plagianakos:2001:LCP

[PNV01] V. P. Plagianakos, N. K. Nousis, and M. N. Vrahatis. Locating and computing in parallel all the simple roots of special functions using PVM. Journal of Computational and Applied Mathematics, 133(1– 2):545–554, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042700006750.

Pokorny:1996:CMP

[Pok96] S. Pokorny. A comparison of message-passing parallelization to shared-memory parallelization. Lecture Notes in Computer Science, 1156: 22–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parrilia:1999:UPD

[POL99] L. Parrilia, J. Ortega, and A. Lloris. Using PVM for distributed logic minimization in a network of computers. In Dongarra et al. [DLM99], pages 541–548. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Pai:2016:CTO

[PP16] Sreepathi Pai and Keshav Pingali. A compiler for throughput optimization of graph algorithms on GPUs. ACM SIGPLAN Notices, 51(10):1–19, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Poplawski:1989:MPP

[PPF89] D. A. Poplawski, S. Pahwa, and J. M. Francioni. Models of parallel program behavior. In Anonymous [Ano89], pages 857–860 (vol. 2). LCCN QA76.5.C619215 1989. Two volumes.

Park:2001:CSL

[PPJ01] So-Hee Park, Mi-Young Park, and Yong-Kee Jun. A comparison of scalable labeling schemes for detecting races in OpenMP programs. Lecture Notes in Computer Science, 2104:68–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2104/21040068.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2104/21040068.
                                        pdf.

Pagourtzis:2001:PCT

[PPR01] Aris Pagourtzis, Igor Potapov, and Wojciech Rytter. PVM computation of the transitive closure: The dependency graph approach. Lecture Notes in Computer Science, 2131:249–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310249.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310249.
pdf.

Papakostas:1996:PSP

[PPT96a] N. Papakostas, G. Papakonstantinou, and P. Tsanakas. PPARDB / PVM: a portable PVM based parallel database management system. Lecture Notes in Computer Science, 1127:219–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papakostas:1996:PPP

[PPT96b] N. Papakostas, G. Papakonstantinou, and P. Tsanakas. PPARDB/PVM: a portable PVM based parallel database management system. In Boszormenyi [Bos96]. ISBN 3-540-61695-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1127.

Papakostas:1996:UPI

[PPT96c] N. Papakostas, G. Papakonstantinou, and P. Tsanakas. Using PVM to implement PPARDB/PVM, a portable parallel database management system. In Bode et al. [BDLS96], pages 108–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Pedicini:2007:PPE

[PQ07] Marco Pedicini and Francesco Quaglia. PELCR: Parallel environment for optimal lambda-calculus reduction. ACM Transactions on Computational Logic, 8(3): 14:1–14:??, July 2007. CO-DEN ???? ISSN 1529-3785 (print), 1557-945X (electronic).

Pinho:2018:CTM

[PQR18] Luis Miguel Pinho, Eduardo Quiñones, and Sara Royuela. Combining the tasklet model with OpenMP. ACM SIGADA Ada Letters, 38(1):14–18, June 2018. CO-DEN AALEE5. ISSN 0736-721X.

Pierce:1994:PIN

[PR94a] P. Pierce and G. Regnier. The Paragon implementation of the NX message passing interface. In Proceedings of the Scalable High-Performance Computing Conference, May 23– 25, 1994, Knoxville, Tennessee [PR94b], pages 184– 190. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.58.S32 1994. IEEE catalog no. 94TH0637-9.

Pierce:1994:PSH

[PR94b] P. Pierce and G. Regnier, editors. Proceedings of the Scalable High-Performance Computing Conference, May 23–25, 1994, Knoxville, Tennessee. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.58.S32 1994. IEEE catalog no. 94TH0637-9.

Pozo:1994:FTE

[PR94c] R. Pozo and K. Remington. Fast three-dimensional elliptic solvers on distributed network clusters. In Joubert et al. [JPTE94], pages 201–208. ISBN 0-444-81841-

3. LCCN QA76.58 .P3794 1993.

Priimak:2014:FDN

[Pri14] Dmitri Priimak. Finite difference numerical method for the superlattice Boltzmann transport equation and case comparison of CPU(C) and GPU(CUDA) implementations. Journal of Computational Physics, 278(??):182–192, December 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0021999114005828.

Proficz:2021:AGA

[Pro21] Jerzy Proficz. All-gather algorithms resilient to imbalanced process arrival patterns. ACM Transactions on Architecture and Code Optimization, 18(4):41:1–41:22, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https://

dl.acm.org/doi/10.1145/
3460122.

Pinho:2021:RTI

[PRQ21] Luis Miguel Pinho, Sara Royuela, and Eduardo Quiñones. Real-time issues in the Ada parallel model with OpenMP. ACM SIGADA Ada Letters, 40(2):96–102, April 2021. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463478.3463491.

Pena:2014:CEC

[PRS^{+}14] Antonio J. Peña, Carlos Reaño, Federico Silla, Rafael Mayo, Enrique S. Quintana-Ortí, and José Duato. A complete and efficient CUDA-sharing solution for HPC clusters. Parallel Computing, 40(10): 574–588, December 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819114001227.

Prades:2016:CAX

[PRS16] Javier Prades, Carlos Reaño, and Federico Silla. CUDA acceleration for Xen virtual machines in InfiniBand clusters with rCUDA. ACM SIGPLAN Notices, 51(8):

35:1–35:??, August 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pedroso:2000:MPC

[PS00a] Hernâni Pedroso and João Gabriel Silva. MPI-2 process creation & management implementation for NT clusters. Lecture Notes in Computer Science, 1908:184–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080184.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080184.
pdf.

Protopopov:2000:SMC

[PS00b] Boris V. Protopopov and Anthony Skjellum. Sharedmemory communication approaches for an MPI messagepassing library. Concurrency: practice and experience, 12(9):799–820, August 10, 2000. CODEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/72516482/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516482&PLACEBO=IE.
pdf.

Pedroso:2001:WLE

[PS01a] Hernâni Pedroso and João Gabriel Silva. The WMPI library evolution: Experience with MPI development for Windows environments. Lecture Notes in Computer Science, 1900:1157–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1900/19001157.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1900/19001157.
                                        pdf.

Protopopov:2001:MMP

[PS01b] Boris V. Protopopov and Anthony Skjellum. A multithreaded Message Passing Interface (MPI) architecture: Performance and program issues. Journal of Parallel and Distributed Computing, 61(4):449–466, April 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        jpdc.2000.1674;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/jpdc.
                                        2000.1674/pdf;   http:
                                        //www.idealibrary.com/
                                        links/doi/10.1006/jpdc.
                                        2000.1674/ref.

Pandey:2007:SCM

[PS07] Nirved Pandey and G. K.

Sharma. Startup comparison for message passing libraries with DTM on Linux clusters. The Journal of Supercomputing, 39 (1):59–72, January 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
39&issue=1&spage=59.

Park:2019:DBO

[PS19a] Sanghyun Park and Taeweon Suh. DQN-based OpenCL workload partition for performance optimization. The Journal of Supercomputing, 75(8):4875–4893, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Prades:2019:GJM

[PS19b] J. Prades and F. Silla. GPUjob migration: The rCUDA case. IEEE Transactions on Parallel and Distributed Systems, 30(12):2718–2729, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pehrson:1994:IPP

[PSB^{+}94] Björn Pehrson, Imre Simon, Klaus Brunnstein, Eckart Raubold, Karen Duncan, and Karl Krueger, editors. Information processing ’94: proceedings of the IFIP 13th World Computer Congress, Hamburg, Germany, 28 August–2 September, 1994, volume A-51, A-52, A-53 of IFIP Transactions. A. Computer Science and Technology. North-Holland, Amsterdam, The Netherlands, 1994. CODEN ITATEC. ISBN 0-444-81990-8, 0-444-81989-4. ISSN 0926-5473. LCCN QA75.5.I3785 1994. Three volumes.

Perez:2019:ATO

[PSB^{+}19] B. Pérez, E. Stafford, J. L. Bosque, R. Beivide, S. Mateo, X. Teruel, X. Martorell, and E. Ayguadé. Auto-tuned OpenCL kernel co-execution in OmpSs for heterogeneous systems. Journal of Parallel and Distributed Computing, 125(??):45–57, March 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731518308189.

Petrovic:2020:BSH

[PSH^{+}20] Filip Petrovič, David Střelák, Jana Hozzová, Jaroslav Ol’ha, Richard Trembecký, Siegfried Benkner, and Jiří Filipovič. A benchmark set of highly-efficient CUDA and OpenCL kernels and its dynamic autotuning with Kernel Tuning Toolkit. Future Generation Computer Systems, 108(??):161–177, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19327360.

Peters:2011:FPC

[PSHL11] Hagen Peters, Ole Schulz-Hildebrandt, and Norbert Luttenberger. Fast in-place, comparison-based sorting with CUDA: a study with bitonic sort. Concurrency and Computation: Practice and Experience, 23(7):681– 693, May 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Patrick:2008:CEO

[PSK08] Christina M. Patrick, Seung-Woo Son, and Mahmut Kandemir. Comparative evaluation of overlap strategies with study of I/O overlap in MPI-IO. Operating Systems Review, 42(6):43– 49, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Preissl:2010:TMS

[PSK^{+}10] Robert Preissl, Martin Schulz, Dieter Kranzlmüller, Bronis R. de Supinski, and Daniel J. Quinlan. Transforming MPI source code based on communication patterns. Future Generation Computer Systems, 26 (1):147–154, January 2010.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Prieto:1999:PRM

[PSLT99] M. Prieto, R. Santiago, I. M. Llorente, and F. Tirado. A parallel robust multigrid algorithm based on semicoarsening. In Dongarra et al. [DLM99], pages 307– 316. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Peng:2014:BAH

[PSM^{+}14] Yuanxi Peng, Manuel Saldaña, Christopher A. Madill, Xiaofeng Zou, and Paul Chow. Benefits of adding hardware support for broadcast and reduce operations in MP-SoC applications. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 7(3):17:1– 17:??, August 2014. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Plunkett:2001:AMD

[PSSS01] Craig L. Plunkett, Alfred G. Striz, and J. Sobieszczanski-Sobieski. Application of MPI in displacement based multilevel structural optimization. Lecture Notes in Computer Science, 2131: 335–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310335.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310335.
pdf.

Pikle:2019:AFE

[PSV19] Nileshchandra K. Pikle, Shailesh R. Sathe, and Arvind Y. Vyavahare. Accelerating the finite element analysis of functionally graded materials using fixedgrid strategy on CUDAenabled GPUs. Concurrency and Computation: Practice and Experience, 31(17): e5207:1–e5207:??, September 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Payrits:2000:UPC

[PSZÉ00] Szabolcs Payrits, Zoltán Szatmáry, László Zalányi, and Péter Érdi. Use of parallel computers in neurocomputing. Lecture Notes in Computer Science, 1908:313–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080313.htm;
 http://link.springer-
ny.com/link/service/series/

                                        0558/papers/1908/19080313.
                                        pdf.

Pears:2001:DLB

[PT01] Arnold N. Pears and Nicola Thong. A dynamic load balancing architecture for PDES using PVM on clusters. Lecture Notes in Computer Science, 2131: 166–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2131/21310166.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2131/21310166.
                                        pdf.

Pai:2013:IGC

[PTG13] Sreepathi Pai, Matthew J. Thazhuthaveetil, and R. Govindarajan. Improving GPGPU concurrency with elastic kernels. ACM SIGPLAN Notices, 48(4):407–418, April 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prost:2001:MIG

[PTH^{+}01a] Jean-Pierre Prost, Richard Treumann, Richard Hedges, Bin Jia, and Alice Koniges. MPI-IO/GPFS, an optimized implementation of MPI-IO on top of GPFS. In ACM [ACM01], page ?? ISBN 1-58113-293-X. LCCN

????  URL http://www.
sc2001.org/papers/pap.
pap186.pdf.

Prost:2001:THP

[PTH^{+}01b] Jean-Pierre Prost, Richard Treumann, Richard Hedges, Alice Koniges, and Alison White. Towards a highperformance implementation of MPI–IO on top of GPFS. Lecture Notes in Computer Science, 1900:1253–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1900/19001253.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/1900/19001253.
pdf.

Peraza:2016:PGQ

[PTL^{+}16] Joshua Peraza, Ananta Tiwari, Michael Laurenzano, Laura Carrington, and Allan Snavely. PMaC’s green queue: a framework for selecting energy optimal DVFS configurations in large scale MPI applications. Concurrency and Computation: Practice and Experience, 28 (2):211–231, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pierro:2018:SFP

[PTMF18] Vincenzo Pierro, Luigi Troiano, Elena Mejuto, and Giovanni Filatrella. Stochastic first passage time accelerated with CUDA. Journal of Computational Physics, 361 (??):136–149, May 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0021999118300494.

Phan-Thien:1994:CDL

[PTT94] N. Phan-Thien and D. Tullock. Completed double layer boundary element method in elasticity and Stokes flow: Distributed computing through PVM. Computational mechanics, 14(4):370–383, July 1994. CODEN CMMEEE. ISSN 0178-7675.

Prylli:1999:DHP

[PTW99] L. Prylli, B. Tourancheau, and R. Westrelin. The design for a high performance MPI implementation on the Myrinet network. In Dongarra et al. [DLM99], pages 223–230. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Puskas:1995:LBW

[Pus95] Z. Puskas. Load balancing on workstation clusters using PVM. In Ferenczi and Kacsuk [FK95], pages 112–

123. ISBN ???? LCCN ???? Technical report KFKI-1995-2/M,N.

Peinado:1997:HPC

[PV97] M. Peinado and R. Venkatesan. Highly parallel cryptographic attacks. Lecture Notes in Computer Science, 1332:367–374, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:2001:PPE

[PVKE01] Insung Park, Michael J.

Voss, Seon Wook Kim, and Rudolf Eigenmann. Parallel programming environment for OpenMP. Scientific Programming, 9(2–3):143–161, Spring–Summer 2001. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C8%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Pahl:1995:CCB

[PW95] Peter Jan Pahl and Heinrich Werner, editors. Computing in civil and building engineering: 6th International conference — July 1995, Berlin, Computing in Civil and Building Engineering 6th. A. A. Balkema, Brookfield, VT, USA, 1995. ISBN 90-5410-556-9, 90-5410-557-

7. LCCN TA345 .I565 1995 v.1-2. Two volumes.

Preissl:2012:CSS

[PWD^{+}12] Robert Preissl, Theodore M. Wong, Pallab Datta, Myron Flickner, Raghavendra Singh, Steven K. Esser, William P. Risk, Horst D. Simon, and Dharmendra S. Modha. Compass: a scalable simulator for an architecture for cognitive computing. In Hollingsworth [Hol12], pages 54:1–54:?? ISBN 1-4673-0804-8. URL http:

//conferences.computer.
org/sc/2012/papers/1000a085.
pdf.

Pang:2016:MKR

[PWP^{+}16] Yeyong Pang, Shaojun Wang, Yu Peng, Xiyuan Peng, Nicholas J. Fraser, and Philip H. W. Leong. A microcoded kernel recursive least squares processor using FPGA technology. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(1):5:1– 5:??, December 2016. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Pirkelbauer:2019:BTF

[PWPD19] Peter Pirkelbauer, Amalee Wilson, Christina Peterson, and Damian Dechev. Blaze-Tasks: a framework for computing parallel reductions over tasks. ACM Transactions on Architecture and Code Optimization, 15(4): 66:1–66:??, January 2019.

CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Prasad:1995:PPB

[PY95] S. K. Prasad and K. M.

Yu. Performance of a PVMbased optimistic simulation testbed on different parallel architectures. In Hamza [Ham95a], pages 511–514. ISBN 0-88986-218-4. LCCN QA76.9.C65 I295 1995.

Perla:2012:PAH

[PZ12] Francesca Perla and Paolo Zanetti. Performance analysis of an hybrid MPI/ OpenMP ALM software for life insurance policies on multi-core architectures. Lecture Notes in Computer Science, 7312: 250–253, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-30961-8_
19/.

Phillips:2002:NBS

[PZKK02] James C. Phillips, Gengbin Zheng, Sameer Kumar, and Laxmikant V. Kalé. NAMD: Biomolecular simulation on thousands of processors. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN

???? URL http://www.sc-
2002.org/paperpdfs/pap.
pap277.pdf.

Qiu:2012:PWM

[QB12] Judy Qiu and Seung-Hee Bae. Performance of windows multicore systems on threading and MPI. Concurrency and Computation: Practice and Experience, 24 (1):14–28, January 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qawasmeh:2017:PPR

[QHCC17] Ahmad Qawasmeh, Maxime R. Hugues, Henri Calandra, and Barbara M. Chapman. Performance portability in reverse time migration and seismic modelling via OpenACC. The International Journal of High Performance Computing Applications, 31(5):422–440, September 2017. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Quaranta:2021:NMH

[QM21] Lionel Quaranta and Lalith Maddegedara. A novel MPI+MPI hybrid approach combining MPI-3 shared memory windows and C11/ C++11 memory model. Journal of Parallel and Distributed Computing, 157(??): 125–144, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S074373152100143X.

Quoy:2000:PNN

[QMGR00] Mathias Quoy, Sorin Moga, Philippe Gaussier, and Arnaud Revel. Parallelization of neural networks using PVM. Lecture Notes in Computer Science, 1908: 289–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080289.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080289.
pdf.

Qaddouri:1995:MFS

[QRG95] A. Qaddouri, R. Roy, and B. Goulard. Multigroup flux solvers using PVM [Parallel Virtual Machine]. In ANS [ANS95], pages 1554–1562. ISBN 0-89448-198-3. LCCN TK9006.M37 1995. Two volumes.

Qaddouri:1996:CPC

[QRMG96] A. Qaddouri, R. Roy, M. Mayrand, and B. Goulard. Collision probability calculation and multigroup flux solvers using PVM. Nuclear Science and Engineering, 123(3):392–402, July 1996. CODEN NSENAO. ISSN 0029-5639.

Qu:1995:FAS

[Qu95]  Su Qu.  Feature-driven

area-based stereo matching method on PVM. M.s. thesis, University of Georgia, Athens, GA, USA, 1995. x + 110 pp. Directed by Hamid R. Arabnia.

Quinn:2003:PPC

[Qui03] Michael J. (Michael Jay) Quinn. Parallel programming in C with MPI and OpenMP. McGraw-Hill, New York, NY, USA, 2003. ISBN 0-07-123265-6, 0-07-282256-2. xiv + 529 pp. LCCN QA76.73.C15 Q55 2003; QA76.73 .C15 Q55 2003.

Russell:1992:CMW

[R^{+}92] Thomas F. Russell et al., editors. Computational methods in water resources IX: Proceedings of the Ninth International Conference on Computational Methods in Water Resources, held at the University of Colorado, Denver, in June 1992. Elsevier Applied Science, London, UK, 1992. ISBN 1-85166-871-3 (set), 1-85312-169-X (set: Computational Mechanics Publications, Southampton), 1-56252-098-9 (set: Computational Mechanics Publications, Boston), 1-85166-791-1 (v. 1: Elsevier Applied Science), 1-85312-197-5 (v. 1: Computational Mechanics Publications, Southampton), 1-56252-123-3 (v. 1: Computational Mechanics Publications, New York), 1-85166-870-5 (v. 2), 1-85312-198-3 (v. 2), 1-56252-124-1 (v. 2). LCCN GB656.2.E42 C65 1992 v.1-2 (c1992). Two volumes.

Rashti:2009:SAM

[RA09] Mohammad J. Rashti and Ahmad Afsahi. A speculative and adaptive MPI rendezvous protocol over RDMA-enabled interconnects. International Journal of Parallel Programming, 37(2):223–246, April 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=2&spage=223.

Rabenseifner:1998:MGI

[Rab98]  R. Rabenseifner.  MPI-

GLUE: Interoperable highperformance MPI combining different vendor’s MPI worlds. Lecture Notes in Computer Science, 1470: 563–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rabenseifner:1999:APM

[Rab99]  R. Rabenseifner.  Auto-

matic profiling of MPI applications with hardware performance counters. In Dongarra et al. [DLM99], pages 35–42. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Ragg:1996:PEN

[Rag96] T. Ragg. Parallelization of an evolutionary neural network optimizer based on PVM. In Bode et al. [BDLS96], pages 351–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Ratha:1995:DED

[RAGJ95] N. K. Ratha, T. Acar, M. Gokmen, and A. K. Jain. A distributed edge detection and surface reconstruction algorithm. In Cantoni et al. [CLM^{+}95], pages 149–154. ISBN 0-8186-7134-3. LCCN QA76.9.A73W675 1995. IEEE catalog no. 95TB8093.

Ramadan:2007:TDM

[Ram07] Omar Ramadan. Three dimensional MPI parallel implementation of the PML algorithm for truncating finite-difference time-domain Grids. Parallel Computing, 33(2):109–115, March 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rantakokko:2005:DMO

[Ran05] Jarmo Rantakokko. A dynamic MPI–OpenMP model for structured adaptive mesh refinement. Parallel Processing Letters, 15(1/2):37–47, March/June 2005. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Rehman:2016:VMJ

[RAS16] Waqas Ur Rehman, Muhammad Sohaib Ayub, and Junaid Haroon Siddiqui. Verification of MPI Java programs using software model checking. ACM SIGPLAN Notices, 51(8):55:1–55:??, August 2016. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roussos:2001:BMB

[RB01] George Roussos and B. J. C.

Baxter. Biharmonic many body calculations for fast evaluation of radial basis function interpolants in cluster environments. Lecture Notes in Computer Science, 2131:288–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310288.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310288.
pdf.

Rufai:2005:MPO

[RBAA05] Raimi Rufai, Muslim Bozyigit, Jaralla Alghamdi, and Moataz Ahmed. Multithreaded parallelism with OpenMP. Parallel Processing Letters, 15 (4):367–378, December 2005. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Rejitha:2017:EPC

[RBAI17] R. S. Rejitha, Shajulin Benedict, Suja A. Alex, and Shany Infanto. Energy prediction of CUDA application instances using dynamic regression models. Computing, 99(8):765–790, August 2017. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Resch:1997:CMP

[RBB97a] M. Resch, H. Berger, and T. Boenisch. A comparison of MPI performance on different MPPs. Lecture Notes in Computer Science, 1332:25–32, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Resch:1997:PM

[RBB97b] Michael Resch, Thomas Beisel, and Holger Berger. PACX-MPI. BI: Informationen für Nutzer des Rechenzentrums 1997,11/12, Universität Stuttgart, Zentrale Universitätseinrichtung, Stuttgart, Germany, 1997.

Resch:1997:PMC

[RBB97c] Michael Resch, Holger Berger, and Thomas Bönisch. Performance of MPI on a Cray T3E-512. BI: Informationen für Nutzer des Rechenzentrums 1997,5/6, Universität Stuttgart, Zentrale Universitätseinrichtung, Stuttgart, Germany, 1997. ?? pp. Third European CRAY-SGI MPP Workshop.

Rodriguez:2015:OPI

[RBB15] Marcos Rodríguez, Fernando Blesa, and Roberto Barrio. OpenCL parallel integration of ordinary differential equations: Applications in computational dynamics. Computer Physics Communications, 192(??): 228–236, July 2015. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465515000703.

Russo:2017:MPG

[RBB17] Igor L. S. Russo, Heder S.

Bernardino, and Helio J. C. Barbosa. A massively parallel grammatical evolution technique with OpenCL. Journal of Parallel and Distributed Computing, 109(??): 333–349, November 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S074373151730206X.

Reis:2020:CMC

[RBC20] Luís Reis, João Bispo, and João M. P. Cardoso. Compilation of MATLAB computations to CPU/GPU via C/ OpenCL generation. Concurrency and Computation: Practice and Experience, 32 (22):e5854:1–e5854:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramachandran:2021:PPB

[RBP^{+}21] Prabhu Ramachandran, Aditya Bhosale, Kunal Puri, Pawan Negi, Abhinav Muta, A. Dinesh, Dileep Menon, Rahul Govind, Suraj Sanka, Amal S. Sebastian, Ananyo Sen, Rohan Kaushik, Anshuman Kumar, Vikas Kurapati, Mrinalgouda Patil, Deep Tavker, Pankaj Pandey, Chandrashekhar Kaushik, Arkopal Dutt, and Arpit Agarwal. PySPH: a Pythonbased framework for smoothed particle hydrodynamics. ACM Transactions on Mathematical Software, 47(4):34:1– 34:38, December 2021. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3460773.

Reale:1994:PCU

[RBS94] F. Reale, F. Bocchino, and S. Sciortino. Parallel computing on Unix workstation arrays. Computer Physics Communications, 83(2-3): 130–140, December 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Rasch:2020:DHL

[RBW^{+}20] Ari Rasch, Julian Bigge, Martin Wrodarczyk, Richard Schulze, and Sergei Gorlatch. dOCAL: highlevel distributed programming with OpenCL and CUDA. The Journal of Supercomputing, 76(7):5117– 5138, July 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Reinhard:1997:MHP

[RC97] E. Reinhard and A. Chalmers.

Message handling in parallel radiance. Lecture Notes in Computer Science, 1332: 486–493, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reimann:1996:CBT

[RCFS96] D. A. Reimann, V. Chaudhary, M. J. Flynn, and I. K. Sethi. Cone beam tomography using MPI on heterogeneous workstation clusters. In IEEE [IEE96i], pages 142–148. ISBN 0-8186-7533-0. LCCN QA76.642 .M67

Ross:1995:DCM

[RCG95] D. L. Ross, J. S. Collins, and J. H. George. A dynamic capacity model using concurrent processing. Neural, Parallel and Scientific Computations, 3(2):249–262, June 1995. CODEN NPACEM. ISSN 1061-5369.

Royuela:2012:ASO

[RDLQ12] Sara Royuela, Alejandro Duran, Chunhua Liao, and Daniel J. Quinlan. Auto-scoping for OpenMP tasks. Lecture Notes in Computer Science, 7312: 29–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/10.
                                        1007/978-3-642-30961-8_
                                        3/.

Radhakrishna:1999:MBP

[RDMB99] H. Radhakrishna, S. Divakar, N. Magotra, and S. R. J. Brueck. MPIbased parallel implementation of a lithography pattern simulation algorithm. Lecture Notes in Computer Science, 1593:109–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reeves:1996:PIC

[Ree96] A. Reeves, editor. Proceedings of the 1996 International Conference on Challenges for Parallel Processing, Ithaca, NY, USA, August 12, 1996, volume 1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7623-X. LCCN QA76.58 .I34 1996. Three volumes.

Reinefeld:2001:CDI

[Rei01] Alexander Reinefeld. Clusters for data-intensive applications in the grid. Lecture Notes in Computer Science, 2131:12–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310012.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310012.
pdf.

Reussner:2001:SSK

[Reu01] Ralf H. Reussner. SKaMPI:

the special Karlsruher MPIbenchmark: user manual. Interner Bericht 99,02, Fakultät für Informatik, Universität Karlsruhe, Karlsruhe, Germany, 2001. 78 pp.

Reussner:2003:USD

[Reu03]  Ralf H. Reussner.  Us-

ing SKaMPI for developing high-performance MPI programs with performance portability. Future Generation Computer Systems, 19(5):749–759, July 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Roy:2000:MGQ

[RFG^{+}00] Alain J. Roy, Ian Foster, William Gropp, Nicholas Karonis, Volker Sander, and Brian Toonen. MPICH-GQ: Quality-of-service for message passing programs. In ACM [ACM00], page 54.

                                        URL http://www.sc2000.
                                        org/proceedings/techpapr/
                                        papers/pap234.pdf.

Reynders:1995:OOO

[RFH^{+}95] John V. W. Reynders, David W. Forslund, Paul J. Hinker, Marydell Tholburn, David G. Kilman, and William F. Humphrey. OOPS: an object-oriented particle simulation class library for distributed architectures. Computer Physics Communications, 87(1–2): 212–224, May 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/001046559400172X.

Russ:1996:HAT

[RFRH96] S. H. Russ, B. Flachs, J. Robinson, and B. Heckel. Hector: automated task allocation for MPI. In IEEE [IEE96e], pages 344–

348. ISBN 0-8186-7255-

2. LCCN QA76.58 .I565 1996. IEEE catalog number 96TB100038. IEEE Computer Society Press order number PR07255.

Rasch:2018:MDH

[RG18] Ari Rasch and Sergei Gorlatch. Multi-dimensional homomorphisms and their implementation in OpenCL. International Journal of Parallel Programming, 46 (1):101–119, February 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Rucci:2018:OOS

[RGB^{+}18] Enzo Rucci, Carlos Garcia, Guillermo Botella, Armando E. De Giusti, Marcelo Naiouf, and Manuel Prieto-Matias. OSWALD: OpenCL Smith–Waterman on Altera’s FPGA for large protein databases. The International Journal of High Performance Computing Applications, 32(3):337– 350, May 2018. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Rough:1997:PRD

[RGD97] J. Rough, A. Goscinski, and D. De Paoli. PVM on the RHODOS distributed operating system. Lecture Notes in Computer Science, 1332:208–218, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodrigues:2013:MAA

[RGD13] A. Wendell O. Rodrigues, Frédéric Guyomarc’h, and Jean-Luc Dekeyser. An MDE approach for automatic code generation from UML/MARTE to OpenCL. Computing in Science and Engineering, 15(1): 46–55, January/February 2013. CODEN CSENFA. ISSN 1521-9615.

Rico-Gallego:2015:ILM

[RGDM15] Juan-Antonio Rico-Gallego and Juan-Carlos Díaz-Martín. τ-Lop: Modeling performance of shared memory MPI. Parallel Computing, 46(??):14–31, July 2015. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819115000447.

Rico-Gallego:2016:EIL

[RGDML16] Juan-Antonio Rico-Gallego, Juan-Carlos Díaz-Martín, and Alexey L. Lastovetsky. Extending τ-lop to model concurrent MPI communications in multicore clusters. Future Generation Computer Systems, 61(??): 66–82, August 2016. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167739X16300346.

Rivas-Gomez:2018:MWS

[RGGP^{+}18] Sergio Rivas-Gomez, Roberto Gioiosa, Ivy Bo Peng, Gokcen Kestor, Sai Narasimhamurthy, Erwin Laure, and Stefano Markidis. MPI windows on storage for HPC applications. Parallel Computing, 77(??):38–56, September 2018. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167819118301571.

Rocco:2022:LFR

[RGP22] Roberto Rocco, Davide Gadioli, and Gianluca Palermo. Legio: fault resiliency for embarrassingly parallel MPI applications. The Journal of Supercomputing, 78(2): 2175–2195, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

https://link.springer.
com/article/10.1007/s11227-
021-03951-w.

Reussner:2001:APP

[RH01] Ralf Reussner and Gunnar Hunzelmann. Achieving performance portability with SKaMPI for highperformance MPI programs. Lecture Notes in Computer Science, 2074:841–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/

                                        link/service/series/0558/
                                        bibs/2074/20740841.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2074/20740841.
                                        pdf.

Roda:1996:PEI

[RHG^{+}96] J. Roda, J. Herrera, J. Gonzalez, C. Rodriguez, F. Almeida, and D. Gonzalez. Practical experiments to improve PVM algorithms. In Bode et al. [BDLS96], pages 30–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Rathgeber:2017:FAF

[RHM^{+}17] Florian Rathgeber, David A. Ham, Lawrence Mitchell, Michael Lange, Fabio Luporini, Andrew T. T. Mcrae, Gheorghe-Teodor Bercea, Graham R. Markall, and Paul H. J. Kelly. Firedrake: Automating the finite element method by composing abstractions. ACM Transactions on Mathematical Software, 43(3):24:1– 24:??, January 2017. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

                                        //dl.acm.org/citation.
                                        cfm?id=2998441.

Rizzardi:2017:ATS

[Riz17] Mariarosaria Rizzardi. Algorithm 981: Talbot Suite DE: Application of modified Talbot’s method to solve differential problems. ACM Transactions on Mathematical Software, 44(2): 18:1–18:23, September 2017. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http:

//dl.acm.org/citation.
cfm?id=3089248.

Ramroach:2021:ADP

[RJ21] Sterling Ramroach and Ajay Joshi. Accelerating data-parallel neural network training with weightedaveraging reparameterisation. Parallel Processing Letters, 31(02):??, June 2021. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.
com/doi/10.1142/S0129626421500092.

Ratha:1995:CUC

[RJC95] N. K. Ratha, A. K. Jain, and M. J. Chung. Clustering using a coarse-grained parallel genetic algorithm: a preliminary study. In Cantoni et al. [CLM^{+}95], pages 331–338. ISBN 0-8186-7134-

3. LCCN QA76.9.A73W675 1995. IEEE catalog no. 95TB8093.

Rodrigues:2014:TPS

[RJDH14] Christopher Rodrigues, Thomas Jablin, Abdul Dakkak, and Wen-Mei Hwu. Triolet: a programming system that unifies algorithmic skeleton interfaces for highperformance cluster computing. ACM SIGPLAN Notices, 49(8):247–258, August 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Renaud:2020:IMS

[RJH^{+}20] Nicolas Renaud, Yong Jung, Vasant Honavar, Cunliang Geng, Alexandre M. J. J. Bonvin, and Li C. Xue. iScore: an MPI supported software for ranking protein-protein docking models based on a random walk graph kernel and support vector machines. SoftwareX, 11(??):Article 100462, January/June 2020. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711019303061.

Robinson:1993:ECD

[RJMC93] D. F. Robinson, D. Judd, P. K. McKinely, and B. H. C. Cheng. Efficient collective data distribution in allport wormhole-routed hypercubes. Proceedings of the Supercomputing Conference, pages 792–801, ???? 1993. CODEN ???? ISBN 0-8186-4340-4. ISSN 1063-9535.

Rabenseifner:2001:ECF

[RK01] Rolf Rabenseifner and Alice E. Koniges. Effective communication and file-I/O bandwidth benchmarks. Lecture Notes in Computer Science, 2131:24–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310024.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310024.
pdf.

Ragan-Kelley:2013:HLC

[RKBA^{+}13] Jonathan Ragan-Kelley, Connelly Barnes, Andrew Adams, Sylvain Paris, Frédo Durand, and Saman Amarasinghe. Halide: a language and compiler for optimizing parallelism, locality, and recomputation in image processing pipelines. ACM SIG-PLAN Notices, 48(6):519– 530, June 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reyes:2013:PEO

[RLFdS13] Ruymán Reyes, Iván López, Juan J. Fumero, and Francisco de Sande. A preliminary evaluation of OpenACC implementations. The Journal of Supercomputing, 65(3):1063–1075, September 2013. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11227-012-0853-z.

Rungsawang:2001:LCP

[RLL01] A. Rungsawang, A. Laohakanniyom, and M. Lertprasertkune. Low-cost parallel text retrieval using PC-cluster. Lecture Notes in Computer Science, 2131:419–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2131/21310419.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2131/21310419.
                                        pdf.

Rubio-Largo:2012:UMO

[RLVRGP12] Álvaro Rubio-Largo, Miguel A. Vega-Rodríguez, and Juan A. Gómez-Pulido. Using a multiobjective OpenMP+MPI DE for the static RWA problem. Lecture Notes in Computer Science, 6927: 224–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/content/pdf/
                                        10.1007/978-3-642-27549-
                                        4_29.

Roe:1999:PMI

[RM99] Kevin Roe and Piyush Mehrotra. Parallelization of a multigrid incompressible viscous cavity flow solver using openMP. NASA contractor report NASA/CR-1999-209551, NASA Langley Research Center, Hampton, VA, USA, 1999. ???? pp. Also ICASE report 99-36.

Rietmann:2012:FAS

[RMNM^{+}12] Max Rietmann, Peter Messmer, Tarje Nissen-Meyer, Daniel Peter, Piero Basini, Dimitri Komatitsch, Olaf Schenk, Jeroen Tromp, Lapo Boschi, and Domenico Giardini. Forward and adjoint simulations of seismic wave propagation on emerging large-scale GPU architectures. In Hollingsworth [Hol12], pages 38:1–38:?? ISBN 1-4673-0804-8. URL

http://conferences.computer.
org/sc/2012/papers/1000a104.
pdf.

Ramesh:2018:MPE

[RMS^{+}18] Srinivasan Ramesh, Aurèle Mahéo, Sameer Shende, Allen D. Malony, Hari Subramoni, Amit Ruhela, and Dhabaleswar K. (DK) Panda. MPI performance engineering with the MPI tool interface: the integration of MVAPICH and TAU. Parallel Computing, 77(??): 19–37, September 2018. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819118301479.

Rodrigues:2013:POM

[RNPM13] Eduardo R. Rodrigues, Philippe O. A. Navaux, Jairo Panetta, and Celso L. Mendes. Preserving the original MPI semantics in a virtualized processor environment. Science of Computer Programming, 78(4): 412–421, April 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167642312001335.

Rohrl:2000:PPS

[Röh00] Armin Röhrl. Parallel processing in statistical computation: BSP, FPGas and MPI for the S-language. Thèse sciences, EPF Lausanne, Lausanne, Switzerland, 2000. 137 pp.

Rolfe:1994:PAP

[Rol94] T. J. Rolfe. PVM: An affordable parallel processing environment. In Anonymous [Ano94h], pages 118–

125. ISBN ???? LCCN ????

Rolfe:2008:PFO

[Rol08a] Timothy J. Rolfe. Perverse and foolish oft I strayed. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2): 52–55, June 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:/

/ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2008.bib.

Rolfe:2008:SMA

[Rol08b] Timothy J. Rolfe. A specimen MPI application: Nqueens in parallel. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4): 42–45, December 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rosen:2013:PVA

[Ros13] Paul Rosen. Performance:

A visual approach to investigating shared and global memory behavior of CUDA kernels. Computer Graphics Forum, 32(3pt2):161– 170, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Roth:2019:AOC

[Rót19] Ágoston Róth. Algorithm 992: An OpenGL- and C++-based function library for curve and surface modeling in a large class of extended Chebyshev spaces. ACM Transactions on Mathematical Software, 45(1): 13:1–13:32, March 2019. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/citation.
cfm?id=3284979.

Ramon:1995:PKV

[RP95] J. Ramon and P. Pena.

Parallelization of KENO-Va Monte Carlo code. Computer Physics Communications, 88(1):76–82, July 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
001046559500025B.

Rodriguez:2008:FTS

[RPM^{+}08] Gabriel Rodríguez, Xoán C. Pardo, María J. Martín, Patricia González, and Daniel Díaz. A fault tolerance solution for sequential and MPI applications on the Grid. Scalable Computing: Practice and Experience, 9(2): 101–109, June 2008. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.
org/vols/vol09/no2/SCPE_
9_2_03.pdf;   http://www.
scpe.org/vols/vol09/no2/
SCPE_9_2_03.zip.

Reano:2019:APP

[RPS19] Carlos Reaño, Javier Prades, and Federico Silla. Analyzing the performance/ power tradeoff of the rCUDA middleware for future exascale systems. Journal of Parallel and Distributed Computing, 132(??):344– 362, October 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731519303491.

Rabaea:2000:EPM

[RR00] Adrian Rabaea and Monica Rabaea. Experiments with parallel Monte Carlo simulation for pricing options using PVM. Lecture Notes in Computer Science, 1908:330–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080330.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080330.
pdf.

Rageb:2001:CEM

[RR01] Khaled Rageb and Wolfgang Rehm. CHEMPI: efficient MPI for VIA/SCI. Preprint-Reihe des Chemnitzer SFB 393, Technische Universität Chemnitz, Chemnitz, Germany, 2001. 12 pp.

Rauber:2002:LSH

[RR02] Thomas Rauber and Gudula Rünger. Library support for hierarchical multiprocessor tasks. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN

???? URL http://www.sc-
2002.org/paperpdfs/pap.
pap176.pdf.

Roda:1997:PPI

[RRAGM97] J. L. Roda, C. Rodriguez, F. Almeida, and D. Gonzalez-Morales. Predicting the performance of injection communication patterns on PVM. Lecture Notes in Computer Science, 1332:33– 40, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roig:2001:EMM

[RRBL01] Concepció Roig, Ana Ripoll, Javier Borrás, and Emilio Luque. Efficient mapping for message-passing applications using the TTIG model: a case study in image processing. Lecture Notes in Computer Science, 2131:370–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310370.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310370.
pdf.

Robinson:1996:TMI

[RRFH96] J. Robinson, S. H. Russ, B. Flachs, and B. Heckel. A task migration implementation of the Message-Passing Interface. In IEEE [IEE96f], pages 61–68. ISBN 0-8186-7582-9. LCCN QA 76.88 I52 1996. IEEE catalog number TB100069.

Russ:1999:UHR

[RRG^{+}99] Samuel H. Russ, Jonathan Robinson, Matt Gleeson, Brad Meyers, Laxman Rajagopalan, and Chun-Heong Tan. Using Hector to run MPI programs over networked workstations. Concurrency: practice and experience, 11(4):189–204, April 10, 1999. CODEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract?ID=61004080;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=61004080&PLACEBO=IE.

pdf. Special Issue: Applications of Distributed Computing Environments.

Russek:2020:SLC

[RRJ^{+}20] Paweł Russek, Paweł Russek, Ernest Jamro, Agnieszka Dabrowska-Boruch, and Kazimierz Wiatr. A study of the loops control for reconfigurable computing with OpenCL in the LABS local search problem. The International Journal of High Performance Computing Applications, 34(1): 103–114, January 1, 2020. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:

//journals.sagepub.com/
doi/full/10.1177/1094342019868515.

Rabenseifner:1993:CDR

[RS93] R. Rabenseifner and A. Schuch.

Comparison of DCE RPC, DFN-RPC, ONC and PVM. In Schill [Sch93], pages 39–

46. ISBN 3-540-57306-2, 0-387-57306-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C55I58 1993.

Reinefeld:1995:PVE

[RS95] A. Reinefeld and V. Schnecke. Portability versus efficiency? parallel applications on PVM and Parix. In Fritzson and Finmo [FF95], pages 35–49. ISBN 90-5199-229-7 (IOS Press), 4-274-90056-8 (Ohmsha). LCCN ????

Roy:1997:PNT

[RS97] R. Roy and Z. Stankovski. Parallelization of neutron transport solvers. Lecture Notes in Computer Science, 1332:494–501, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reano:2019:SIN

[RS19] Carlos Reaño and Federico Silla. On the support of inter-node P2P GPU memory copies in rCUDA. Journal of Parallel and Distributed Computing, 127(??): 28–43, May 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0743731519300255.

Reano:2021:RRC

[RS21] Carlos Reaño and Federico Silla. Redesigning the rCUDA communication layer for a better adaptation to the underlying hardware. Concurrency and Computation: Practice and Experience, 33(14):e5481:1– e5481:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Raskovalov:2022:AMD

[RS22] Anton Raskovalov and Platon Surkov. azTotMD 2.0: Molecular dynamics with the radiative thermostat and temperature-dependent force field (CUDA version). SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711022000127.

Rambu:1995:DSS

[RSBT95] N. Rambu, S. Stefan, D. Borsan, and S. Talpos. A diagnostic study of some meteorological fields simulated with UKMO and MPI atmospheric general circulation models. In Gates [Gat95], pages 493–498. ISBN ???? LCCN SIO 1 WO326 v.92.

Reano:2015:IUE

[RSC^{+}15] Carlos Reaño, Federico Silla, Adrián Castelló, Antonio J. Peña, Rafael Mayo, Enrique S. Quintana-Ortí, and José Duato. Improving the user experience of the rCUDA remote GPU virtualization framework. Concurrency and Computation: Practice and Experience, 27(14):3746–3770, September 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ruhela:2019:EDM

[RSC^{+}19] Amit Ruhela, Hari Subramoni, Sourav Chakraborty, Mohammadreza Bayatpour, Pouya Kousha, and Dhabaleswar K. (DK) Panda. Efficient design for MPI asynchronous progress without dedicated resources. Parallel Computing, 85(??): 13–26, July 2019. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819118303302.

Reussner:1998:SDA

[RSPM98] R. Reussner, P. Sanders, L. Prechelt, and M. Mueller. SKaMPI: a detailed, accurate MPI benchmark. Lecture Notes in Computer Science, 1497:52–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reussner:2002:SCB

[RST02] Ralf Reussner, Peter Sanders, and Jesper Larsson Träff. SKaMPI: a comprehensive benchmark for public benchmarking of MPI. Scientific Programming, 10(1): 55–65, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=9ejnuvwuvby9737jte27%
26referrer=parent%26backto=
issue%2C6%2C9%3Bjournal%
2C2%2C12%3Blinkingpublicationresults%
2C1%2C1.

Rozman:2006:CPL

[RšT06] Igor Rozman, Marjan šterk, and Roman Trobec. Communication performance of LAM/MPI and MPICH on a Linux cluster. Parallel Processing Letters, 16(3): 323–334, September 2006. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Roberti:2005:PIL

[RSV^{+}05] Debora R. Roberti, Roberto P.

Souto, Haroldo F. Campos Velho, Gervasio A. Degrazia, and Domenico Anfossi. Parallel implementation of a Lagrangian stochastic model for pollutant dispersion. International Journal of Parallel Programming, 33(5): 485–498, October 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=5&spage=485.

Reussner:2000:BMD

[RTH00] Ralf Reussner, Jesper Larsson Träff, and Gunnar Hunzelmann. A benchmark for MPI derived datatypes. Lecture Notes in Computer Science, 1908:10–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1908/19080010.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1908/19080010.
                                        pdf.

Rungsawang:1999:PDT

[RTL99] A. Rungsawang, A. Tangpong, and P. Laohawee. Parallel DSIR text retrieval system. In Dongarra et al. [DLM99], pages 325–332. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Rundo:2021:CPM

[RTN21] Leonardo Rundo, Andrea Tangherloni, and Marco S. Nobile. A CUDA-powered method for the feature extraction and unsupervised analysis of medical images. The Journal of Supercomputing, 77(8):8514– 8531, August 2021. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

https://link.springer.
com/article/10.1007/s11227-
020-03565-8.

Rycerz:2007:IBS

[RTRG^{+}07] Katarzyna Rycerz, Alfredo Tirado-Ramos, Alessia Gualandris, Simon F. Portegies Zwart, Marian Bubak, and Peter M. A. Sloot. Interactive N-body simulations on the Grid: HLA versus MPI. The International Journal of High Performance Computing Applications, 21(2):210– 221, May 2007. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/21/
2/210.full.pdf+html.

Reynders:2000:IPI

[RV00] John Reynders and Alexander V. Veidenbaum, editors. ICS ’00: Proceedings of the 14th international conference on Supercomputing: Santa Fe, New Mexico, USA, May 8– 11, 2000. ACM Press, New York, NY 10036, USA, 2000. ISBN 1-58113-270-0. LCCN QA76.88 .I573 2000. URL

https://dl.acm.org/doi/
proceedings/10.1145/335231.

Riebler:2018:ACA

[RVKP18] Heinrich Riebler, Gavin Vaz, Tobias Kenter, and Christian Plessl. Automated code acceleration targeting heterogeneous OpenCL devices. ACM SIGPLAN Notices, 53(1):417–418, January 2018. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Riebler:2019:TAH

[RVKP19] Heinrich Riebler, Gavin Vaz, Tobias Kenter, and Christian Plessl. Transparent acceleration for heterogeneous platforms with compilation to OpenCL. ACM Transactions on Architecture and Code Optimization, 16(2):14:1–14:??, May 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ropo:2009:RAP

[RWD09] Matti Ropo, Jan Westerholm, and Jack Dongarra, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 16th European PVM/ MPI Users’ Group Meeting, Espoo, Finland, September 7–10, 2009. Proceedings, volume 5759 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CO-DEN LNCSD9. ISBN 3-642-03769-0 (print), 3-642-03770-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:/

/www.springerlink.com/

content/978-3-642-03770-
2.

Simonsen:1993:DMD

[SA93] H. H. Simonsen and J. Amundsen. Distributed molecular dynamics using the PVM system. In Sincovec [Sin93], pages 183–186. ISBN 0-89871-315-3. LCCN QA 76.58 S55 1993. Two volumes.

Saarinen:1994:EES

[Saa94] S. Saarinen. EASYPVM — an enhanced subroutine library for PVM. In Gentzsch and Harms [GH94], pages 267–272. ISBN 0-387-57981-8 (New York), 3-540-57981-8 (Berlin). LCCN QA76.88.I57 1994. DM96.00. Two volumes.

Sainio:2010:CGA

[Sai10]  J. Sainio.  CUDAEASY

— a GPU accelerated cosmological lattice program. Computer Physics Communications, 181(5):906– 912, May 2010. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465510000159.

Sato:2017:NIT

[SAL^{+}17] Kento Sato, Dong H. Ahn, Ignacio Laguna, Gregory L. Lee, Martin Schulz, and Christopher M. Chambreau. Noise injection techniques to expose subtle and unintended message races. ACM SIGPLAN Notices, 52(8): 89–101, August 2017. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Saphir:1997:SMI

[Sap97] William Saphir. A survey of MPI implementations. NHSE Review, 2(1):??, November 1997.

Soldado:2016:ECM

[SAP16] Fábio Soldado, Fernando Alexandre, and Hervé Paulino. Execution of compound multi-kernel OpenCL computations in multi-CPU/ multi-GPU environments. Concurrency and Computation: Practice and Experience, 28(3):768–787, March 10, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sahimi:2001:AAS

[SAS01] Mohd Salleh Sahimi, Norma Alias, and Elankovan Sundararajan. The AGEB algorithm for solving the heat equation in three space dimensions and its parallelization using PVM. Lecture Notes in Computer Science, 2073:918–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2073/20730918.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2073/20730918.
pdf.

Schuster:1995:CSM

[SB95] G. Schuster and F. Breitenecker. Coupling simulators with the model interconnection concept and PVM. In Breitenecker and Husinsky [BH95], pages 321–326. ISBN 0-444-82241-0. LCCN A76.9.C65E966 1995.

Smith:2001:DMM

[SB01] Lorna Smith and Mark Bull. Development of mixed mode MPI/OpenMP applications. Scientific Programming, 9(2–3):83–98, Spring– Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C3%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Spiliotis:2020:PII

[SBB20] Iraklis M. Spiliotis, Michael P.

Bekakos, and Yiannis S. Boutalis. Parallel implementation of the Image Block Representation using OpenMP. Journal of Parallel and Distributed Computing, 137 (??):134–147, March 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0743731519307622.

Seyfarth:1994:GEE

[SBF94] B. R. Seyfarth, J. L. Bickham, and M. R. Fernandez. Glenda: an environment for easy parallel programming. In Pierce and Regnier [PR94b], pages 637–

641. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.58.S32 1994. IEEE catalog no. 94TH0637-9.

Schulz:2004:IES

[SBF^{+}04] Martin Schulz, Greg Bronevetsky, Rohit Fernandes, Daniel Marques, Keshav Pingali, and Paul Stodghill. Implementation and evaluation of a scalable application-level checkpoint-recovery scheme for MPI programs. In ACM [ACM04], page 38. ISBN 0-7695-2153-3. LCCN ????

Selikhov:2002:MCC

[SBG^{+}02] Anton Selikhov, George Bosilca, Cecile Germain, Gilles Fedak, and Franck Cappello. MPICH-CM: a communication library design for a P2P MPI implementation. Lecture Notes in Computer Science, 2474: 323–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.de/link/
service/series/0558/bibs/
2474/24740323.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740323.pdf.

Schindewolf:2012:WSA

[SBG^{+}12] Martin Schindewolf, Barna Bihari, John Gyllenhaal, Martin Schulz, Amy Wang, and Wolfgang Karl. What scientific applications can benefit from hardware transactional memory? In Hollingsworth [Hol12], pages 90:1–90:?? ISBN 1-4673-0804-8. URL http:

//conferences.computer.
org/sc/2012/papers/1000a073.
pdf.

Skjellum:2020:FSI

[SBG20] Anthony Skjellum, Purushotham V. Bangalore, and Ryan E. Grant. Foreword to the special issue of the Workshop on Exascale MPI (ExaMPI 2017). Concurrency and Computation: Practice and Experience, 32 (3):e5459:1–e5459:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sojoodi:2021:IGG

[SBK21] Amir Hossein Sojoodi, Majid Salimi Beni, and Farshad Khunjush. Igniteg-GPU: a GPU-enabled inmemory computing architecture on clusters. The Journal of Supercomputing, 77 (3):3165–3192, March 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

https://link.springer.
com/article/10.1007/s11227-
020-03390-z.

Sani:2014:PDF

[SBQZ14] Ardalan Amiri Sani, Kevin Boos, Shaopu Qin, and Lin Zhong. I/O paravirtualization at the device file boundary. ACM SIGARCH Computer Architecture News, 42 (1):319–332, March 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Smith:1995:CRC

[SBR95] K. A. Smith, A. J. Baratta, and G. E. Robinson. Coupled RELAP5 and CON-TAIN accident analysis using PVM. Nuclear safety, 36(1):94–108, January–June 1995. CODEN NUSAAZ. ISSN 0029-5604.

Smith:2004:SIP

[SBT04] Kevin B. Smith, Aart J. C. Bik, and Xinmin Tian. Support for the Intel(R) Pentium(R) 4 processor with hyper-threading technology in Intel(R) 8.0 compilers. Intel Technology Journal, 8(1): 19–31, February 2004. ISSN 1535-766X. URL http:

//developer.intel.com/
technology/itj/2004/volume08issue01/
art02_compilers/p01_abstract.
htm.

Saltz:1991:MRT

[SBW91] J. Saltz, H. Berryman, and J. Wu. Multiprocessors and run-time compilation. Concurrency: practice and experience, 3(6):573–592, December 1991. CODEN CPEXEI. ISSN 1040-3108.

Stubbs:1995:ICE

[SC95] S. S. Stubbs and D. L.

Carver. IPCC++: a C++ extension for interprocess communication with objects. In IEEE [IEE95l], pages 205–210. ISBN 0-8186-7119-X. LCCN QA 76.6 C6295 1995. IEEE catalog no. 95CB35838.

Smith:1996:UWC

[SC96a] N. P. G. Smith and C. Christopoulos. Utilising workstation clusters with PVM for the solution of large TLM problems. In Silvester [Sil96], pages 3–11. ISBN 1-85312-395-1. LCCN TK5.I59 1996.

Steed:1996:PPP

[SC96b] M. R. Steed and M. J.

Clement. Performance prediction of PVM programs. In IEEE [IEE96e], pages 803–807. ISBN 0-8186-7255-2. LCCN QA76.58 .I565 1996. IEEE catalog number 96TB100038. IEEE Computer Society Press order number PR07255.

Sievert:2004:SMP

[SC04] Otto Sievert and Henri Casanova. A simple MPI process swapping architecture for iterative applications. The International Journal of High Performance Computing Applications, 18(3):341–352, Fall 2004. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

http://hpc.sagepub.com/
content/18/3/341.full.
pdf+html.

Shterenlikht:2019:MVF

[SC19] Anton Shterenlikht and Luis Cebamanos. MPI vs Fortran coarrays beyond 100k cores: 3D cellular automata. Parallel Computing, 84(??): 37–49, May 2019. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819118303181.

Saillard:2014:PCS

[SCB14] Emmanuelle Saillard, Patrick Carribault, and Denis Barthou. PARCOACH: Combining static and dynamic validation of MPI collective communications. The International Journal of High Performance Computing Applications, 28(4):425–434, November 2014. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/28/
4/425.

Saillard:2015:SDV

[SCB15] Emmanuelle Saillard, Patrick Carribault, and Denis Barthou. Static/dynamic validation of MPI collective communications in multi-threaded context. ACM SIGPLAN Notices, 50(8):279–280, August 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stagg:1995:IPN

[SCC95] A. K. Stagg, D. D. Cline, and G. F. Carey. Implementing a parabolized Navier– Stokes flow solver on the Cray T3D. In Bailey et al. [BBG^{+}95], pages 143–148. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Shyu:1996:ILQ

[SCC96] Shyong Jian Shyu, H. K.- C. Chang, and K.-C. Chou. Implementation of a linear quadtree coding scheme on the parallel virtual machine. International Journal of High Speed Computing, 8 (1):65–79, March 1996. CO-DEN IHSCEZ. ISSN 0129-0533.

Schill:1993:DOD

[Sch93] Alexander Schill, editor.

DCE — the OSF distributed computing environment: client/server model and beyond: International DCE Workshop, Karlsruhe, Germany, October 7–8, 1993: proceedings, number 731 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-57306-2, 0-387-57306-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C55I58 1993.

Schneenman:1994:DSS

[Sch94] Richard D. Schneenman. Distributed supercomputing software: experiences with the parallel virtual machine — PVM. Technical Report NISTIR 5381, U.S. Dept. of Commerce, National Institute of Standards and Technology, Gaithersburg, MD, USA, 1994. vi + 18 pp.

Schuele:1996:PLA

[Sch96a] J. Schuele. Parallel Lanczos algorithm on a CRAY-T3D combining PVM and SHMEM routines. Lecture Notes in Computer Science, 1156:158–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schule:1996:PLA

[Sch96b] J. Schule. Parallel Lanczos algorithm on a CRAY-T3D combining PVM and SHMEM routines. In Bode et al. [BDLS96], pages 158–165. ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Schuele:1999:HAP

[Sch99] J. Schuele. Heading for an asynchronous parallel ocean model. In Dongarra et al. [DLM99], pages 404–409. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Schevtschenko:2001:PAS

[Sch01] I. V. Schevtschenko. A parallel ADI and steepest descent methods. Lecture Notes in Computer Science, 2131:265–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310265.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310265.
pdf.

Searles:2019:MOA

[SCJH19] Robert Searles, Sunita Chandrasekaran, Wayne Joubert, and Oscar Hernandez. MPI + OpenACC: Accelerating radiation transport mini-application, minisweep, on heterogeneous systems. Computer Physics Communications, 236(??): 176–187, March 2019. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465518303552.

Song:1997:ALL

[SCL97] Jianjian Song, Heng Kek Choo, and Kuok Ming Lee. Application-level load migration and its implementation on top of PVM. Concurrency: practice and experience, 9(1):1–19, January 1997. CODEN CPEXEI. ISSN 1040-3108.

Suppi:2000:IOP

[SCL00] Remo Suppi, Fernando Cores, and Emilio Luque. Improving optimistic PDES in PVM environments. Lecture Notes in Computer Science, 1908:304–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1908/19080304.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1908/19080304.
                                        pdf.

Suppi:2001:PCS

[SCL01] Remo Suppi, Fernando Cores, and Emilio Luque. PDES: a case study using the switch time warp.

Lecture Notes in Computer Science, 2131:327–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310327.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310327.
pdf.

Santos:1997:ECP

[SCP97] L. P. Santos, V. Castro, and A. Proenca. Evaluation of the communication performance on a parallel processing system. Lecture Notes in Computer Science, 1332:41–48, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

SCRI:1992:PWC

[SCR92] Proceedings of the Workshop on Cluster Computing. Supercomputing Computations Research Institute, Florida State University, Tallahassee, FL, USA, December 1992. ISBN ???? LCCN ???? Proceedings available via anonymous

ftp from ftp.scri.fsu.edu
in directory pub/parallel-
workshop.92.

Shi:2012:VGA

[SCSL12] Lin Shi, Hao Chen, Jianhua Sun, and Kenli Li. vCUDA: GPU-accelerated high-performance computing in virtual machines. IEEE Transactions on Computers, 61(6):804–816, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Szeberenyi:1999:SGB

[SD99] I. Szeberenyi and G. Domokos. Solving generalized boundary value problems with distributed computing and recursive programming. In Dongarra et al. [DLM99], pages 267–274. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

SM-D:2013:BRC

[SD13] SM-D. Book review: CUDA Programming, Shane Cook. Morgan Kaufmann. ISBN 978-0-12-415933-4. Network Security, 2013(1):4, January 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S1353485813700151.

Sorensen:2016:EER

[SD16] Tyler Sorensen and Alastair F. Donaldson. Exposing errors related to weak memory in GPU applications. ACM SIGPLAN Notices, 51(6):100–113, June 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Skjellum:1994:WLM

[SDB94] A. Skjellum, N. E. Doss, and P. V. Bangalore. Writing libraries in MPI. In IEEE [IEE94f], pages 166– 173. ISBN 0-8186-4980-1. LCCN QA76.58.S34 1993.

Sorensen:2016:PIW

[SDB^{+}16] Tyler Sorensen, Alastair F.

Donaldson, Mark Batty, Ganesh Gopalakrishnan, and Zvonimir Rakamarić. Portable inter-workgroup barrier synchronisation for GPUs. ACM SIGPLAN Notices, 51(10):39–58, October 2016. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schmitt:2017:SCP

[SDJ17] Felix Schmitt, Robert Dietrich, and Guido Juckeland. Scalable critical-path analysis and optimization guidance for hybrid MPI–CUDA applications. The International Journal of High Performance Computing Applications, 31(6):485–498, November 2017. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Sandes:2010:CUG

[SdM10] Edans Flavius O. Sandes and Alba Cristina M. A. de Melo. CUDAlign: using GPU to accelerate the comparison of megabase genomic sequences. ACM SIG-PLAN Notices, 45(5):137– 146, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sistare:1999:MSP

[SDN99] Steve Sistare, Erica Dorenkamp, and Nick Nevin. MPI support in the Prism programming environment. In ACM [ACM99], page ??

Schwarzrock:2021:RNI

[SdR^{+}21] J. Schwarzrock, C. C. de Oliveira, M. Ritt, A. F. Lorenzon, and A. C. S. Beck. A runtime and non-intrusive approach to optimize EDP by tuning threads and CPU frequency for OpenMP applications. IEEE Transactions on Parallel and Distributed Systems, 32(7):1713–1724, 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sampaio:2013:DA

[SdSCP13] Diogo Sampaio, Rafael Martins de Souza, Sylvain Collange, and Fernando Magno Quintão Pereira. Divergence analysis. ACM Transactions on Programming Languages and Systems, 35(4):13:1–13:??, December 2013. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Skjellum:1995:EMP

[SDV^{+}95] A. Skjellum, N. E. Doss, K. Viswanathan, A. Chowdappa, and P. V. Bangalore. Extending the message passing interface (MPI). In IEEE [IEE95j], pages 106– 118. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Sack:2002:FMB

[SE02] Paul Sack and Anne C.

Elster. Fast MPI broadcasts through reliable multicasting. Lecture Notes in Computer Science, 2367: 445–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2367/23670445.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2367/23670445.
pdf.

Spencer:2015:DLN

[SEC15] Matt Spencer, Jesse Eickholt, and Jianlin Cheng. A deep learning network approach to ab initio protein secondary structure prediction. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 12 (1):103–112, January 2015. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Schenck:2016:EPM

[SEF^{+}16] Wolfram Schenck, Salem El Sayed, Maciej Foszczynski, Wilhelm Homberg, and Dirk Pleiter. Evaluation and performance modeling of a burst buffer solution. Operating Systems Review, 50(3):12– 26, December 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Segovia:2010:PPN

[Seg10] Alejandro Segovia. Parallel programming with NVIDIA CUDA. Linux Journal, 2010(200):2:1–2:??, December 2010. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Seifert:1999:ESI

[Sei99] Friedrich Seifert. Entwicklung von Systemsoftware zur Integration der Virtual InterfaceArchitecture (VIA) in den Linux Betriebssystemkern für optimiertes MessagePassing. (German) [Development of system software for integration of the Virtual InterfaceArchitecture (VIA) in the Linux operating system for optimized message passing]. Diplomarbeit, Technische Universität Chemnitz-Zwickau, Chemnitz, Germany, 1999. 115 pp.

Sept:1993:DIP

[Sep93] Doug Sept. The design, implementation and performance of a queue manager for PVM. M.s. thesis, Computer Science Department, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 1993. viii + 45 pp.

Serot:1997:EPF

[Ser97] J. Serot. Embodying parallel functional skeletons: An experimental implementation on top of MPI. Lecture Notes in Computer Science, 1300:629–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sevenich:1998:PPU

[Sev98] Richard Sevenich. Parallel processing using PVM. Linux Journal, 45:??, January 1998. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Scott:1998:PWN

[SFG98] S. L. Scott, M. Fischer, and A. Geist. PVM on Windows and NT clusters. Lecture Notes in Computer Science, 1497:231–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoinas:1994:FGA

[SFL^{+}94] Ioannis Schoinas, Babak Falsafi, Alvin R. Lebeck, Steven K. Reinhardt, James R. Larus, and David A. Wood. Fine-grain access control for distributed shared memory. ACM SIGPLAN Notices, 29(11):297–306, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/
citations/proceedings/
asplos/195473/p297-schoinas/

.

Steuwer:2015:GPP

[SFLD15] Michel Steuwer, Christian Fensch, Sam Lindley, and Christophe Dubach. Generating performance portable code using rewrite rules: from high-level functional expressions to high-performance OpenCL code. ACM SIG-PLAN Notices, 50(9):205– 217, September 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Siegelin:1995:BPW

[SFO95] C. Siegelin, U. Finger, and C. O’Donnell. Boosting the performance of workstations through WARPmemory. In Haridi et al. [HAM95b], pages 703–706. ISBN 3-540-60247-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I553 1995.

Shen:2013:ACE

[SFSV13] Jie Shen, Jianbin Fang, Henk Sips, and Ana Lucia Varbanescu. An applicationcentric evaluation of OpenCL on multi-core CPUs. Parallel Computing, 39(12): 834–850, December 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819113001014.

Selikhov:2005:CMB

[SG05] A. Selikhov and C. Germain.

A Channel Memory based fault tolerance for MPI applications. Future Generation Computer Systems, 21(5):709–715, May 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sharma:2012:SRP

[SG12] Subodh Sharma and Ganesh Gopalakrishnan. A sound reduction of persistent-sets for deadlock detection in MPI applications. Lecture Notes in Computer Science, 7498: 194–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-33296-8_

Steuwer:2014:SHL

[SG14] Michel Steuwer and Sergei Gorlatch. SkelCL: a highlevel extension of OpenCL for multi-GPU systems. The Journal of Supercomputing, 69(1):25–33, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

link.springer.com/article/
10.1007/s11227-014-1213-
y.

Sack:2015:CAM

[SG15] Paul Sack and William Gropp. Collective algorithms for multiported torus networks. ACM Transactions on Parallel Computing (TOPC), 1(2):12:1– 12:??, January 2015. CO-DEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Sunderam:1994:PCC

[SGDM94] V. S. Sunderam, G. A. Geist, J. Dongarra, and R. Manchek. The PVM concurrent computing system: Evolution, experiences, and trends. Parallel Computing, 20(4):531– 545, March 31, 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=20&
issue=4&aid=861.

Schneider:2012:MAC

[SGH12] Timo Schneider, Robert Gerstenberger, and Torsten Hoefler. Micro-applications for communication data access patterns and MPI datatypes. Lecture Notes in Computer Science, 7490: 121–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-33518-1_
17/.

Solsona:2001:IEI

[SGHL01] Francesc Solsona, Francesc Giné, Porfidio Hernández, and Emilio Luque. Implementing explicit and implicit coscheduling in a PVM environment (research note). Lecture Notes in Computer Science, 1900:1165–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1900/19001165.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/1900/19001165.
pdf.

Saito:2003:LSP

[SGJ^{+}03] Hideki Saito, Greg Gaertner, Wesley Jones, Rudolf Eigenmann, Hidetoshi Iwashita, Ron Lieberman, Matthijs van Waveren, and Brian Whitney. Large system performance of SPEC OMP benchmark suites. International Journal of Parallel Programming, 31(3):197– 209, June 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL /ips/frames/
Refs/referenceskapmain.
asp?J=4773&I=33&A=3&LK=
NM;   http://ipsapp007.
kluweronline.com/content/
getfile/4773/33/3/abstract.
htm;   http://ipsapp007.
kluweronline.com/content/
getfile/4773/33/3/fulltext.
pdf.

Solsona:2000:MCM

[SGL^{+}00] Francesc Solsona, Francesc Giné, Josep Lérida, Porfidio Hernández, and Emilio Luque. Monito: a communication monitoring tool for a PVM–Linux environment. Lecture Notes in Computer Science, 1908:233–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080233.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080233.
pdf.

Sun:2020:RTS

[SGL^{+}20] J. Sun, N. Guan, F. Li, H. Gao, C. Shi, and W. Yi. Real-time scheduling and analysis of OpenMP DAG tasks supporting nested parallelism. IEEE Transactions on Computers, 69(9): 1335–1348, September 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sekharan:1995:LBM

[SGS95] Chandra N. Sekharan, Vineet Goel, and R. Sridhar. Load balancing methods for ray tracing and binary tree computing using PVM. Parallel Computing, 21(12):1963–1978, December 12, 1995. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1995&volume=21&
issue=12&aid=1028.

Stone:2010:OPP

[SGS10] John E. Stone, David Gohara, and Guochun Shi. OpenCL: a parallel programming standard for heterogeneous computing systems. Computing in Science and Engineering, 12 (3):66–73, May/June 2010. CODEN CSENFA. ISSN 0740-7475 (print), 1558-1918 (electronic).

Sun:2021:ACW

[SGS^{+}21] J. Sun, N. Guan, J. Sun, X. Zhang, Y. Chi, and F. Li. Algorithms for computing the WCRT bound of OpenMP task systems with conditional branches. IEEE Transactions on Computers, 70(1):57–71, January 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Scherer:2000:APO

[SGZ00] Alex Scherer, Thomas Gross, and Willy Zwaenepoel. Adaptive parallelism for OpenMP task parallel programs. Lecture Notes in Computer Science, 1915: 113–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1915/19150113.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1915/19150113.
                                        pdf.

Schmidt:1994:IAP

[SH94] M. Schmidt and R. Hanisch. Implementation of an air pollution transport model on parallel hardware. In Dekker et al. [DSZ94], pages 277– 284. ISBN 0-444-81784-0. LCCN QA76.58.E98 1994.

Sitsky:1996:MLW

[SH96] D. Sitsky and E. Hayashi. An MPI library which uses polling, interrupts and remote copying for the Fujitsu AP1000+. In Li et al. [LHHM96], pages 43– 49. ISBN 0-8186-7460-1. LCCN QA76.58.I5673 1996. IEEE catalog number 96TB100044.

Song:2014:DAT

[SH14] Sukhyun Song and Jeffrey K.

Hollingsworth. Designing and auto-tuning parallel 3-D FFT for computationcommunication overlap. ACM SIGPLAN Notices, 49(8): 181–192, August 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shen:1995:PSM

[She95] H. Shen. Parallel k-set mutual range-join in hypercubes. Microprocessing and Microprogramming, 41(7): 443–448, November 1995. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic).

Sloot:1994:CIO

[SHH94a] P. M. A. Sloot, A. G. Hoekstra, and L. O. Hertzberger. A comparison of the Iserver-Occam, Parix, Express, and PVM programming environments on a Parsytec GCel. In Gentzsch and Harms [GH94], pages 253– 259. ISBN 0-387-57981-8 (New York), 3-540-57981-8 (Berlin). LCCN QA76.88.I57 1994. DM96.00. Two volumes.

Sloot:1994:CIP

[SHH94b] P. M. A. Sloot, A. G. Hoekstra, and L. O. Hertzberger. A comparison of the Iserver-Occam, Parix, Express, and PVM programming environments on a Parsytec GCel. In Gentzsch and Harms [GH94], pages 253–

259. ISBN 0-387-57981-8 (New York), 3-540-57981-8 (Berlin). LCCN QA76.88.I57 1994. DM96.00. Two volumes.

Sojka:2018:IEM

[SHHC18] Radim Sojka, David Horák, Václav Hapla, and Martin Cermák. The impact of enabling multiple subdomains per MPI process in the TFETI domain decomposition method. Applied Mathematics and Computation, 319(??):586–597, February 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300317304927.

Sato:2001:CEO

[SHHI01] Mitsuhisa Sato, Hiroshi Harada, Atsushi Hasegawa, and Yutaka Ishikawa. Clusterenabled OpenMP: An OpenMP compiler for the SCASH software distributed shared memory system. Scientific Programming, 9(2–3):123– 130, Spring–Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

http://iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C6%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Shing:1994:UPC

[Shi94] C.-C. Shing. Use PVM on computation of analysis of repeated measurement designs. In Sall and Lehman [SL94a], pages 139–

142. ISBN 1-886658-00-5. LCCN QA276.4.S95 1994.

Samadi:2014:LGU

[SHLM14] Mehrzad Samadi, Amir Hormati, Janghaeng Lee, and Scott Mahlke. Leveraging GPUs using cooperative loop speculation. ACM Transactions on Architecture and Code Optimization, 11(1): 3:1–3:??, February 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sato:2010:BLL

[SHM^{+}10] Mitsuhisa Sato, Toshihiro Hanawa, Matthias S. Müller, Barbara M. Chapman, and Bronis R. de Supinski, editors. Beyond Loop Level Parallelism in OpenMP: Accelerators, Tasking and More: 6th International Workshop on OpenMP, IWOMP 2010, Tsukuba, Japan, June 14–16, 2010 Proceedings, volume 6132 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CO-DEN LNCSD9. ISBN 3-642-13216-2 (print), 3-642-13217-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:/

/www.springerlink.com/
content/978-3-642-13217-
9.

Samadi:2012:AIA

[SHM^{+}12] Mehrzad Samadi, Amir Hormati, Mojtaba Mehrara, Janghaeng Lee, and Scott Mahlke. Adaptive inputaware compilation for graphics engines. ACM SIGPLAN Notices, 47(6):13–22, June 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PLDI ’12 proceedings.

Shah:2000:FCS

[SHPT00] Sanjiv Shah, Grant Haab, Paul Petersen, and Joe Throop. Flexible control structures for parallelism in OpenMP. Concurrency: practice and experience, 12(12):1219–1239, October 2000. CODEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76500348/

START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500348&PLACEBO=IE.
pdf.

Sato:2001:OGR

[SHTS01] Mitsuhisa Sato, Motonari Hirano, Yoshio Tanaka, and Satoshi Sekiguchi. OmniRPC: a Grid RPC facility for cluster and global computing in OpenMP. Lecture Notes in Computer Science, 2104:130–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2104/21040130.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2104/21040130.
pdf.

Simmendinger:2019:ISG

[SIC^{+}19] Christian Simmendinger, Roman Iakymchuk, Luis Cebamanos, Dana Akhmetova, Valeria Bartsch, Tiberiu Rotaru, Mirko Rahn, Erwin Laure, and Stefano Markidis. Interoperability strategies for GASPI and MPI in large-scale scientific applications. The International Journal of High Performance Computing Applications, 33(3):554–568, May 1, 2019. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:/

                                        /journals.sagepub.com/
                                        doi/full/10.1177/1094342018808359.

Siegel:1992:FFS

[Sie92a] H. J. Siegel, editor. Frontiers ’92, the Fourth Symposium on the Frontiers of Massive Parallel Computation, October 19–21, 1992, McLean, Virginia. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2772-

7. LCCN QA76.58.S95 1992. IEEE catalog no. 92CH3185-6.

Siegel:1992:FSF

[Sie92b] H. J. Siegel, editor. The Fourth Symposium on the Frontiers of Massively Parallel Computation: Frontiers ’92 / October 19– 21, 1992, McLean Virginia. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2772-7. LCCN QA76.58.S95 1992. IEEE catalog number 92CH3185-6.

Siegal:1994:PEI

[Sie94] Howard Jay Siegal, editor. Proceedings / Eighth International Parallel Processing Symposium, April 26– 29, 1994, Cancun, Mexico. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994.

ISBN 0-8186-5602-6. LCCN QA76.58.I58 1994. IEEE catalog no. 94CH34819.

Silvester:1996:SEE

[Sil96] P. P. Silvester, editor. Software for electrical engineering analysis and design: Third International Conference on Software for Electrical Engineering Analysis and Design, Electrosoft ’96, Pisa, Italy. Computational Mechanics Publications, Boston, MA, USA, 1996. ISBN 1-85312-395-1. LCCN TK5.I59 1996.

Sincovec:1993:SCP

[Sin93] Richard F. Sincovec, editor.

SIAM Conference on Parallel Processing for Scientific Computing (6th: 1993: Norfolk, VA, USA). Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1993. ISBN 0-89871-315-3. LCCN QA 76.58 S55 1993. Two volumes.

Silla:2017:BRG

[SIRP17] Federico Silla, Sergio Iserte, Carlos Reaño, and Javier Prades. On the benefits of the remote GPU virtualization mechanism: The rCUDA case. Concurrency and Computation: Practice and Experience, 29(13), July 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sharma:2017:PDR

[SIS17] Prateek Sharma, David Irwin, and Prashant Shenoy. Portfolio-driven resource management for transient cloud servers. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1): 5:1–5:??, June 2017. CO-DEN ???? ISSN 2476-1249.

                                        URL http://dl.acm.org/
                                        citation.cfm?id=3084442.

Sistare:2002:UHP

[SJ02] Steven J. Sistare and Christopher J. Jackson. Ultrahigh performance communication with MPI and the Sun Fire(^{T M}) link interconnect. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN

                                        ???? URL http://www.sc-
                                        2002.org/paperpdfs/pap.
                                        pap142.pdf.

Szo:2017:PET

[SJK^{+}17a] Máté Szőke, Tamás István Józsa, Ádám Koleszár, Irene Moulitsas, and László Könözsy. Performance evaluation of a two-dimensional lattice Boltzmann solver using CUDA and PGAS UPC based parallelisation. ACM Transactions on Mathematical Software, 44(1): 8:1–8:??, July 2017. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

                                        //dl.acm.org/citation.
                                        cfm?id=3085590.

Szoke:2017:PET

[SJK^{+}17b] Máté Szőke, Tamás István Józsa, Ádám Koleszár, Irene Moulitsas, and László Könözsy. Performance evaluation of a two-dimensional lattice Boltzmann solver using CUDA and PGAS UPC based parallelisation. ACM Transactions on Mathematical Software, 44(1):8:1–8:22, July 2017. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Samadi:2014:PPB

[SJLM14] Mehrzad Samadi, Davoud Anoushe Jamshidi, Janghaeng Lee, and Scott Mahlke. Paraprox: pattern-based approximation for data parallel applications. ACM SIGARCH Computer Architecture News, 42(1):35– 50, March 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Shen:1992:VTD

[SK92] S. Shen and L. Kleinrock. The virtual-time data-parallel machine. In Siegel [Sie92b], pages 46–53. ISBN 0-8186-2772-7. LCCN QA76.58.S95 1992. IEEE catalog number 92CH3185-6.

Smith:2000:DPM

[SK00] Lorna Smith and Paul Kent.

Development and performance of a mixed OpenMP/ MPI quantum Monte Carlo code. Concurrency: practice and experience, 12(12):1121– 1129, October 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76500350/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500350&PLACEBO=IE.
pdf.

Sanders:2010:CEI

[SK10] Jason Sanders and Edward Kandrot. CUDA by Example: an Introduction to General-purpose GPU Programming. Addison-Wes-ley, Reading, MA, USA, 2010. ISBN 0-13-138768-

5. xix + 290 pp. LCCN QA76.76.A65.

Steinberger:2014:WTB

[SKB^{+}14] Markus Steinberger, Michael Kenzel, Pedro Boechat, Bernhard Kerbl, Mark Dokter, and Dieter Schmalstieg. Whippletree: task-based scheduling of dynamic workloads on the GPU. ACM Transactions on Graphics, 33(6):228:1–228:??, November 2014. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Skjellum:2004:RTM

[SKD^{+}04] Anthony Skjellum, Arkady Kanevsky, Yoginder S. Dandass, Jerrell Watts, Steve Paavola, Dennis Cottel, Greg Henley, L. Shane Hebert, Zhenqian Cui, Anna Rounbehler, and The Real-Time Message Passing Interface (Mpi and Rt) Forum. The Real-Time Message Passing Interface Standard (MPI/RT-1.1). Concurrency and Computation: Practice and Experience, 16(S1):Si– S322, December 25, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Subramaniam:1996:CLU

[SKH96] Krishnan R. Subramaniam, Suraj C. Kothari, and Don Heller. A communication library using active messages to improve performance of PVM. Journal of Parallel and Distributed Computing, 39(2): 146–152, December 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.idealibrary.com/
links/doi/10.1006/jpdc.
1996.0162/production;
http://www.idealibrary.
com/links/doi/10.1006/
jpdc.1996.0162/production/
pdf.

Skjellum:1993:SLH

[Skj93] A. Skjellum. Scalable libraries in a heterogeneous environment. In IEEE [IEE93c], pages 13–20. ISBN 0-8186-3900-8, 0-8186-3901-

6. LCCN QA76.9.D5I593 1993. IEEE catalog no. 93TH0550-4.

Steinberger:2012:SDS

[SKK^{+}12] Markus Steinberger, Bernhard Kainz, Bernhard Kerbl, Stefan Hauswiesner, Michael Kenzel, and Dieter Schmalstieg. Softshell: dynamic scheduling on GPUs. ACM Transactions on Graphics, 31(6):161:1–161:??, November 2012. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Spiechowicz:2015:GAM

[SKM15] J. Spiechowicz, M. Kostur, and L. Machura. GPU accelerated Monte Carlo simulation of Brownian motors dynamics with CUDA. Computer Physics Communications, 191(??):140– 149, June 2015. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465515000417.

Satoh:2001:COT

[SKS01] Shigehisa Satoh, Kazuhiro Kusano, and Mitsuhisa Sato. Compiler optimization techniques for OpenMP programs. Scientific Programming, 9(2–3):131–142, Spring–Summer 2001. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C7%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Sall:1994:CIS

[SL94a] J. Sall and A. Lehman, editors. Computational intensive statistical methods: 26th Symposium on the interface

— June 15-18, 1994, Research Triangle Park, NC, USA, volume 26 of Computing Science and Statistics Conference. Fairfax Station: Interface Foundation of North America, ????, 1994. ISBN 1-886658-00-5. LCCN QA276.4.S95 1994.

Scales:1994:DES

[SL94b] D. J. Scales and M. S. Lam.

The design and evaluation of a shared object system for distributed memory machines. In USENIX [USE94], pages 101–114. ISBN 1-880446-66-9. LCCN QA 76.76 O63 U87 1994.

Swanson:1995:PAP

[SL95] Eric Swanson and Terry P.

Lybrand. PVM-AMBER: a parallel implementation of the AMBER molecular mechanics package for workstation clusters. Journal of Computational Chemistry, 16(9):1131–1140, September 1995. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shyu:2000:APV

[SL00] Shyong-Jian Shyu and B. M. T. Lin. An application of parallel virtual machine framework to film production problem. Computers and Mathematics with Applications, 39(12): 53–62, June 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0898122100001292.

Skjellum:1995:EAM

[SLG95] Anthony Skjellum, Ewing Lusk, and William Gropp. Early applications in the Message-Passing Interface (MPI). International Journal of Supercomputer Applications and High Performance Computing, 9(2):79– 94, Summer 1995. CODEN IJSCFG. ISSN 1078-3482.

Scherer:1999:TAP

[SLGZ99] Alex Scherer, Honghui Lu, Thomas Gross, and Willy Zwaenepoel. Transparent adaptive parallelism on NOWs using OpenMP. ACM SIGPLAN Notices, 34 (8):96–106, August 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.

acm.org/pubs/citations/
proceedings/ppopp/301104/
p96-scherer/.

Samadi:2014:SPS

[SLJ^{+}14] Mehrzad Samadi, Janghaeng Lee, D. Anoushe Jamshidi, Scott Mahlke, and Amir Hormati. Scaling performance via self-tuning approximation for graphics engines. ACM Transactions on Computer Systems, 32(3): 7:1–7:??, September 2014. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Su:2012:CPB

[SLN^{+}12] ChunYi Su, Dong Li, Dimitrios S. Nikolopoulos, Matthew Grove, Kirk Cameron, and Bronis R. de Supinski. Critical pathbased thread placement for NUMA systems. ACM SIGMETRICS Performance Evaluation Review, 40(2): 106–112, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sloan:2005:HPL

[Slo05] Joseph D. (Joseph Donald) Sloan. High performance Linux clusters with OSCAR, Rocks, openMosix, and MPI. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00570-9. xv + 350 pp.

LCCN QA76.58; QA76.58 .S56 2005eb; QA76.58 .S56 2005; QA76.58 .S58 2005; QA76.58 .S595 2005. URL

http://www.oreilly.com/
catalog/9780596005702.

Squyres:1996:CBP

[SLS96] J. M. Squyres, A. Lumsdaine, and R. L. Stevenson. A cluster-based parallel image processing toolkit. In Grinstein and Erbacher [GE96], pages 228–239. CO-DEN PSISDG. ISBN 0-8194-2030-1. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.2656.

Shires:2002:EHM

[SM02] D. Shires and R. Mohan. An evaluation of HPF and MPI approaches and performance in unstructured finite element simulations. Journal of Mathematical Modelling and Algorithms, 1(3): 153–167, 2002. CODEN ???? ISSN 1570-1166.

Shires:2003:OPF

[SM03] Dale Shires and Ram Mohan. Optimization and performance of a Fortran 90 MPI-based unstructured code on large-scale parallel systems. The Journal of Supercomputing, 25(2): 131–141, June 2003. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp009.kluweronline.

com/content/getfile/5189/
44/4/abstract.htm;   http:
//ipsapp009.kluweronline.
com/content/getfile/5189/
44/4/fulltext.pdf.

Simos:2007:CMS

[SM07] Theodore E. Simos and George Maroulis, editors. Computation in Modern Science and Engineering: Proceedings of the [Fifth] International Conference on Computational Methods in Science and Engineering 2007 (ICCMSE 2007), Corfu, Greece, 25– 30 September 2007, volume 2A, 2B of AIP Conference Proceedings (#963). American Institute of Physics, Woodbury, NY, USA, 2007. ISBN 0-7354-0476-3 (set), 0-7354-0477-1 (vol. 1), 0-7354-0478-X (vol. 2). ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. LCCN Q183.9 .I524

2007. URL http://www.
springer.com/physics/atoms/
book/978-0-7354-0478-6.

Santos:2012:ICC

[SM12] Bruno F. L. Santos and Hendrik T. Macedo. Improving CUDA^{T M} C/C++ encoding readability to foster parallel application development. ACM SIGSOFT Software Engineering Notes, 37(1):1–5, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shea:2019:HSD

[SM19] Colin Shea and Tinoosh Mohsenin. Heterogeneous scheduling of deep neural networks for low-power real-time designs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(4):36:1– 36:??, December 2019. CO-DEN ???? ISSN 1550-

                                        4832. URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3358699.

Siegel:2008:CSE

[SMAC08] Stephen F. Siegel, Anastasia Mironova, George S. Avrunin, and Lori A. Clarke. Combining symbolic execution with model checking to verify parallel numerical programs. ACM Transactions on Software Engineering and Methodology, 17 (2):10:1–10:??, April 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Shterenlikht:2015:FC

[SMCH15] Anton Shterenlikht, Lee Margetts, Luis Cebamanos, and David Henty. Fortran 2008 coarrays. ACM Fortran Forum, 34(1):10–30, April 2015. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Smith:1993:MBA

[Smi93a]  K. A. Smith.  Multi-

processor based accident using PVM. In Sincovec [Sin93], pages 262–265. ISBN 0-89871-315-3. LCCN QA 76.58 S55 1993. Two volumes.

Smith:1993:DSI

[Smi93b]  S. L. Smith.  Dynamic

scheduling of irregularly structured parallel computations in heterogeneous distributed systems. ACM SIGPLAN Notices, 28(1): 86, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schardl:2017:TEF

[SML17] Tao B. Schardl, William S.

Moses, and Charles E. Leiserson. Tapir: Embedding fork-join parallelism into LLVM’s intermediate representation. ACM SIG-PLAN Notices, 52(8):249– 265, August 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schardl:2019:TER

[SML19] Tao B. Schardl, William S.

Moses, and Charles E. Leiserson. Tapir: Embedding recursive fork-join parallelism into LLVM’s intermediate representation. ACM Transactions on Parallel Computing (TOPC), 6(4):19:1– 19:??, December 2019. CO-DEN ???? ISSN 2329-4949 (print), 2329-4957 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3365655.

Sandes:2016:MMA

[SMM^{+}16] Edans F. De O. Sandes, Guillermo Miranda, Xavier Martorell, Eduard Ayguade, George Teodoro, and Alba C. M. A. De Melo. MASA: a multiplatform architecture for sequence aligners with block pruning. ACM Transactions on Parallel Computing (TOPC), 2(4):28:1– 28:??, March 2016. CO-DEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Sochacki:1993:DCW

[SMOE93] J. S. Sochacki, D. Mitchum, P. O’Leary, and R. E. Ewing. Distributed computation of wave propagation models using PVM. In IEEE [IEE93e], pages 22–33. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Silva:2000:HPC

[SMS00] Luís Moura Silva, Paulo Martins, and João Gabriel Silva. Heterogeneous parallel computing using Java and WMPI. Concurrency: practice and experience, 12(11):1077–1091, September 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76000189/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76000189&PLACEBO=IE.
pdf.

Su:2006:APP

[SMSW06] Hai-Jun Su, J. Michael McCarthy, Masha Sosonkina, and Layne T. Watson. Algorithm 857: POL-SYS GLP—a parallel general linear product homotopy code for solving polynomial systems of equations. ACM Transactions on Mathematical Software, 32 (4):561–579, December 2006. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sitsky:1996:IMU

[SMTW96] D. Sitsky, P. Mackerras, A. Tridgell, and D. Walsh. Implementing MPI under AP/Linux. In IEEE [IEE96i], pages 32–39. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Sunderam:2001:CAP

[SN01] Vaidy Sunderam and Zsolt Németh. A comparative analysis of PVM/MPI and computational Grids. Lecture Notes in Computer Science, 2131:14–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310014.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310014.
pdf.

Snir:2018:FMT

[Sni18]  Marc Snir.  The future

of MPI: technical perspective. Communications of the ACM, 61(10):105, October 2018. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

cacm.acm.org/magazines/
2018/10/231376/fulltext.

Suciu:2010:PIN

[SNMP10] A. Suciu, I. Nagy, K. Marton, and I. Pinca. Parallel implementation of the NIST Statistical Test Suite. In Ioan Alfred Letia, editor, Proceedings, 2010 IEEE 6th International Conference on Intelligent Computer Communication and Processing: Cluj-Napoca, Romania, August 26–28, 2010, pages 363– 368. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-8228-3 (print), 1-4244-8230-5 (electronic). LCCN QA76.76.E95. URL http:

//ieeexplore.ieee.org/
servlet/opac?punumber=

5598248. IEEE catalog number CFP1009D-ART.

Shekofteh:2019:MSG

[SNN^{+}19] S.-Kazem Shekofteh, Hamid Noori, Mahmoud Naghibzadeh, Hadi Sadoghi Yazdi, and Holger Fröning. Metric selection for GPU kernel classification. ACM Transactions on Architecture and Code Optimization, 15(4): 68:1–68:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Shekofteh:2020:CEC

[SNN^{+}20] S.-Kazen Shekofteh, Hamid Noori, Mahmoud Naghibzadeh, Holger Fröning, and Hadi Sadog Yazdi. cCUDA: Effective coscheduling of concurrent kernels on GPUs. IEEE Transactions on Parallel and Distributed Systems, 31(4):766– 778, April 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sintorn:2011:EAF

[SOA11] Erik Sintorn, Ola Olsson, and Ulf Assarsson. An efficient alias-free shadow algorithm for opaque and transparent objects using per-triangle shadow volumes. ACM Transactions on Graphics, 30(6):153:1– 153:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Snir:1996:MCR

[SOHL^{+}96] Marc Snir, Steve W. Otto, Steven Huss-Lederman, David W. Walker, and Jack Dongarra. MPI: the complete reference. MIT Press, Cambridge, MA, USA, 1996. ISBN 0-262-69184-1. xii + 336 pp. LCCN QA76.642.M65 1996. US$27.50.

Snir:1998:MCR

[SOHL^{+}98] Marc Snir, Steve W. Otto, Steven Huss-Lederman, David W. Walker, and Jack Dongarra. MPI: The Complete Reference. Volume 1, The MPI-1 Core. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, second edition, September 1998. ISBN 0-262-69215-5 (vol. 1), 0-262-69216-3 (set). 450 pp. LCCN QA76.642 .M65 1998. US$35 (paperback).

URL http://mitpress.
mit.edu/book-home.tcl?
isbn=0262692155. See also

volume 2 [GHLL^{+}98].

SousaPinto:2001:PEI

[Sou01] Jorge Sousa Pinto. Parallel evaluation of interaction nets with MPINE. Lecture Notes in Computer Science, 2051:353–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2051/20510353.htm;

http://link.springer-
ny.com/link/service/series/
0558/papers/2051/20510353.
pdf.

St-Onge:2019:ESS

[SOYHDD19] Guillaume St-Onge, Jean-Gabriel Young, Laurent Hébert-Dufresne, and Louis J. Dubé. Efficient sampling of spreading processes on complex networks using a composition and rejection algorithm. Computer Physics Communications, 240(??): 30–37, July 2019. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465519300608.

Sidonio:1999:PBI

[SP99] N. Sidonio and A. Pereira.

A parallel N-body integrator using MPI. Lecture Notes in Computer Science, 1573:627–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stpiczynski:2011:SKB

[SP11] Przemyslaw Stpiczyński and Joanna Potiopa. Solving a kind of boundaryvalue problem for ordinary differential equations using Fermi — the next generation CUDA computing architecture. Journal of Computational and Applied Mathematics, 236(3):384– 393, September 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042711004237.

Singh:2017:EER

[SPB^{+}17] Amit Kumar Singh, Alok Prakash, Karunakar Reddy Basireddy, Geoff V. Merrett, and Bashir M. Al-Hashimi. Energy-efficient run-time mapping and thread partitioning of concurrent OpenCL applications on CPU–GPU MPSoCs. ACM Transactions on Embedded Computing Systems, 16(5s):147:1– 147:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Silla:2020:IPP

[SPBR20] Federico Silla, Javier Prades, Elvira Baydal, and Carlos Reaño. Improving the performance of physics applications in atom-based clusters with rCUDA. Journal of Parallel and Distributed Computing, 137 (??):160–178, March 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731519304034.

Satofuka:1995:PCF

[SPE95] N. Satofuka, Jacques Periaux, and Akin Ecer, editors. Parallel computational fluid dynamics: new algorithms and applications: proceedings of the Parallel CFD ’94 Conference, Kyoto, Japan, 16–19 May 1994. Elsevier, Amsterdam, The Netherlands, 1995. ISBN 0-444-82317-4. LCCN QA911 .P35 1994.

Speck:2019:APP

[Spe19] Robert Speck. Algorithm 997: pySDC-prototyping spectral deferred corrections. ACM Transactions on Mathematical Software, 45(3): 35:1–35:23, August 2019. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/citation.
cfm?id=3310410.

Shaw:1995:ADA

[SPH95] R. A. (Richard A.) Shaw, H. E. (Harry E.) Payne, and J. J. E. (Jeffrey J. E.) Hayes, editors. Astronomical data analysis software and systems IV: meeting held at Baltimore, Maryland, 25–28 September 1994, volume 77 of Astronomical Society of the Pacific Conference Series. Astronomical Society of the Pacific, San Francisco, CA, USA, 1995. ISBN 0-937707-96-1. ISSN 1080-7926. LCCN QB51.3.E43 A87 1994.

Skjellum:1996:TTM

[SPH96] A. Skjellum, B. Protopopov, and S. Hebert. A thread taxonomy for MPI. In IEEE [IEE96i], pages 50–57. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Si:2018:DAA

[SPH^{+}18] Min Si, Antonio J. Pena, Jeff Hammond, Pavan Balaji, Masamichi Takagi, and Yutaka Ishikawa. Dynamic adaptable asynchronous progress model for MPI RMA multiphase applications. IEEE Transactions on Parallel and Distributed Systems, 29(9): 1975–1989, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/09/08315136-
                                        abs.html.

Sener:1996:DPP

[SPK96] C. Sener, Y. Paker, and A. Kiper. Data-parallel programming on Helios, parallel environment and PVM. In Yetongnon and Hariri [YH96], pages 2–?? ISBN ???? LCCN ????

Subramoni:2012:DSI

[SPK^{+}12] H. Subramoni, S. Potluri, K. Kandalla, B. Barth, J. Vienne, J. Keasler, K. Tomko, K. Schulz, A. Moody, and D. K. Panda. Design of a scalable InfiniBand topology service to enable network-topology-aware placement of processes. In Hollingsworth [Hol12], pages 70:1–70:?? ISBN 1-4673-0804-8. URL http:

//conferences.computer.
org/sc/2012/papers/1000a076.
pdf.

Silva:1999:DPP

[SPL99] F. Silva, H. Paulino, and L. Lopes. DipSystem: a parallel programming system for distributed memory architectures. In Dongarra et al. [DLM99], pages 525–

532. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Schmidl:2012:PAT

[SPL^{+}12] Dirk Schmidl, Peter Philippen, Daniel Lorenz, Christian Rössel, and Markus Geimer. Performance analysis techniques for taskbased OpenMP applications. Lecture Notes in Computer Science, 7312: 196–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-30961-8_
15/.

Saldana:2010:MPM

[SPM^{+}10] Manuel Saldaña, Arun Patel, Christopher Madill, Daniel Nunes, Danyao Wang, Paul Chow, Ralph Wittig, Henry Styles, and Andrew Putnam. MPI as a programming model for high-performance reconfigurable computers. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(4): 22:1–22:??, November 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Symeonidou:2014:DRB

[SPNB14] Christi Symeonidou, Polyvios Pratikakis, Dimitrios S. Nikolopoulos, and Angelos Bilas. Distributed regionbased memory allocation and synchronization. The International Journal of High Performance Computing Applications, 28(4):406– 414, 2014. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

                                        tronic). URL http://hpc.
                                        sagepub.com/content/28/
                                        4/406.

Squyres:2003:CAL

[Squ03]  Jeffrey M. Squyres.  A

component architecture for LAM/MPI (citation only). ACM SIGPLAN Notices, page ??, 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sivaraman:1995:PSP

[SR95] H. Sivaraman and C. S. Raghavendra. Parallelizing sequential programs to a cluster of workstations. In Agrawal [Agr95a], pages 38– 41. ISBN 0-8493-2618-4. LCCN QA76.58.I34 1995.

Sivaraman:1996:AAD

[SR96] H. Sivaraman and C. S.

Raghavendra. ADDT: Automatic data distribution tool for porting programs to PVM. In El-Rewini and Shriver [ERS96], pages 557–

564. ISBN 0-8186-7324-9. ISSN 1060-3425. LCCN ???? Five volumes.

Simitci:1998:CLP

[SR98] Huseyin Simitci and Daniel A.

Reed. A comparison of logical and physical parallel I/O patterns. The International Journal of High Performance Computing Applications, 12(3): 364–380, Fall 1998. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

//journals.sagepub.com/
doi/pdf/10.1177/109434209801200305.

Szalay:2011:FCD

[SR11] Zsófia Szalay and János Rohonczy. Fast calculation of DNMR spectra on CUDAenabled graphics card. Journal of Computational Chemistry, 32(7):1262–1270, May 2011. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Speck:2012:MST

[SRK^{+}12] R. Speck, D. Ruprecht, R. Krause, M. Emmett, M. Minion, M. Winkel, and P. Gibbon. A massively space-time parallel N-body solver. In Hollingsworth [Hol12], pages 92:1–92:?? ISBN 1-4673-0804-8. URL http:

//conferences.computer.
org/sc/2012/papers/1000a083.
pdf.

Sultana:2019:FRB

[SRS^{+}19] Nawrin Sultana, Martin Rüfenacht, Anthony Skjellum, Ignacio Laguna, and Kathryn Mohror. Failure recovery for bulk synchronous applications with MPI stages. Parallel Computing, 84(??):1–14, May 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167819118303260.

Schmidt:1994:EAO

[SS94] B. K. Schmidt and V. S. Sunderam. Empirical analysis of overheads in cluster environments. Concurrency: practice and experience, 6(1):1–32, February 1994. CODEN CPEXEI. ISSN 1040-3108.

Szymanski:1996:LCR

[SS96] Boleslaw K. Szymanski and Balaram Sinharoy, editors. Languages, Compilers and Run-Time Systems for Scalable Computers, 22–24 May 1995, Troy, NY, USA. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Silva:1999:IME

[SS99] P. Silva and J. G. Silva. Implementing MPI-2 extended collective operations. In Dongarra et al. [DLM99], pages 125–132. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Shan:2001:CMS

[SS01] Hongzhang Shan and Jaswinder Pal Singh. A comparison of MPI, SHMEM and cache-coherent shared address space programming models on a tightly-coupled multiprocessors. International Journal of Parallel Programming, 29(3):283–318, June 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//ipsapp009.lwwonline.
com/content/getfile/4773/
21/3/abstract.htm;   http:
//ipsapp009.lwwonline.
com/content/getfile/4773/
21/3/fulltext.pdf.

Schwarz:2009:GFG

[SS09] Michael Schwarz and Marc Stamminger. GPU: Fast GPU-based adaptive tessellation with CUDA. Computer Graphics Forum, 28(2): 365–374, April 2009. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shan:2012:OAA

[SSAS12] Hongzhang Shan, Erich Strohmaier, James Amundson, and Eric G. Stern. Optimizing the advanced accelerator simulation framework Synergia using OpenMP. Lecture Notes in Computer Science, 7312:140– 153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/10.
                                        1007/978-3-642-30961-8_
                                        11/.

Sankaran:2005:LMC

[SSB^{+}05] Sriram Sankaran, Jeffrey M. Squyres, Brian Barrett, Vishal Sahay, Andrew Lumsdaine, Jason Duell, Paul Hargrove, and Eric Roman. The LAM/MPI checkpoint/ restart framework: Systeminitiated checkpointing. The International Journal of High Performance Computing Applications, 19(4):479– 493, Winter 2005. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

                                        tronic). URL http://hpc.
                                        sagepub.com/content/19/
                                        4/479.full.pdf+html.

Sataric:2016:HOM

[SSB^{+}16] Bogdan Satarić, Vladimir Slavnić, Aleksandar Belić, Antun Balaz, Paulsamy Muruganandam, and Sadhan K. Adhikari. Hybrid OpenMP/MPI programs for solving the time-dependent Gross–Pitaevskii equation in a fully anisotropic trap. Computer Physics Communications, 200(??):411– 417, March 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465515004440.

Sotomayor:2017:ACG

[SSB^{+}17] Rafael Sotomayor, Luis Miguel Sanchez, Javier Garcia Blas, Javier Fernandez, and J. Daniel Garcia. Automatic CPU/GPU generation of multi-versioned OpenCL kernels for C++ scientific applications. International Journal of Parallel Programming, 45(2):262– 282, April 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10766-016-0425-6.

Spiliotis:2021:PCD

[SSB21] Iraklis M. Spiliotis, Charalampos Sitaridis, and Michael P. Bekakos. Parallel computation of discrete orthogonal moment on block represented images using OpenMP. International Journal of Parallel Programming, 49(3):440–462, June 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s10766-
                                        021-00713-2.

Silva:1996:IDS

[SSC96] L. M. Silva, J. G. Silva, and S. Chapple. Implementing distributed shared memory on top of MPI: the DSMPI library. In IEEE [IEE96g], pages 50–57. ISBN 0-8186-7376-1. LCCN QA76.58 .E97 1996. IEEE order number PR07376.

Silva:1997:IPD

[SSC97] Luis M. Silva, Joao Gabriel Silva, and Simon Chapple. Implementation and performance of DSMPI. Scientific Programming, 6(2):201– 214, Summer 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Silva:1995:PCR

[SSCC95] L. M. Silva, J. G. Silva, S. Chapple, and L. Clarke. Portable checkpointing and recovery. In IEEE [IEE95k], pages 188–195. ISBN 0-8186-7088-6. LCCN QA76.9.D5 I328 1995. IEEE catalog no. 95TB8075.

Skjellum:1994:DEZ

[SSD^{+}94] A. Skjellum, S. G. Smith, N. E. Doss, A. P. Leung, and M. Morari. The design and evolution of Zipcode. Parallel Computing, 20 (4):565–596, March 31, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Shen:2020:GPC

[SSD^{+}20] Qi Shen, Craig Sharp, Richard Davison, Gary Ushaw, Rajiv Ranjan, Albert Y. Zomaya, and Graham Morgan. A general purpose contention manager for software transactions on the GPU. Journal of Parallel and Distributed Computing, 139(??):1–17, May 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731519301376.

Sabne:2012:ECO

[SSE12] Amit Sabne, Putt Sakdhnagool, and Rudolf Eigenmann. Effects of compiler optimizations in OpenMP to CUDA translation. Lecture Notes in Computer Science, 7312:169–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/
10.1007/978-3-642-30961-
8_13/.

Stellner:1995:CMP

[SSG95] G. Stellner, M. Schumann, and M. Girnghuber. Comparing message-passing libraries with the SPY analysis environment. Informationstechnik und technische Informatik: IT + TI, 37(2): 46–52, April 1995. CODEN ITINEV. ISSN 0944-2774.

Sosa:2000:IQC

[SSGF00] C. P. Sosa, G. Scalmani, R. Gomperts, and M. J. Frisch. Ab initio quantum chemistry on a ccNUMA architecture using openMP. III. Parallel Computing, 26 (7–8):843–856, July 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/35/21/42/29/
                                        25/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/35/21/42/29/25/article.
                                        pdf.

Sala:2008:PHP

[SSH08] Marzio Sala, W. F. Spotz, and M. A. Heroux. PyTrilinos: High-performance distributed-memory solvers for Python. ACM Transactions on Mathematical Software, 34(2):7:1–7:33, March 2008. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sharif:2019:APC

[SSH^{+}19] Hashim Sharif, Prakalp Srivastava, Muhammad Huzaifa, Maria Kotsifakou, Keyur Joshi, Yasmin Sarita, Nathan Zhao, Vikram S. Adve, Sasa Misailovic, and Sarita Adve. ApproxHPVM: a portable compiler IR for accuracyaware optimizations. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA): 186:1–186:30, October 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3360612.

Schmitt:2018:RHG

[SSK^{+}18] Christian Schmitt, Moritz Schmid, Sebastian Kuckuk, Harald Köstler, Jürgen Teich, and Frank Hannig. Reconfigurable hardware generation of multigrid solvers with conjugate gradient coarse-grid solution. Parallel Processing Letters, 28(04):??, December 2018. ISSN 0129-6264 (print), 1793-642X (electronic). URL https:

//www.worldscientific.
com/doi/10.1142/S0129626418500160.

Schafers:1995:TGP

[SSKF95] L. Schafers, C. Scheidler, and O. Kramer-Fuhrmann. TRAPPER: a graphical programming environment for parallel systems. Future Generation Computer Systems, 11(4-5):351–361, August 1995. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Squyres:1997:DEM

[SSL97] J. M. Squyres, B. Saphir, and A. Lumsdaine. The design and evolution of the MPI-2 C++ interface. Lecture Notes in Computer Science, 1343:57–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shi:2010:PAE

[SSLMW10] Haixiang Shi, Bertil Schmidt, Weiguo Liu, and Wolfgang Müller-Wittig. A parallel algorithm for error correction in high-throughput short-read data on CUDAenabled graphics hardware. Journal of Computational Biology, 17(4):603– 615, April 2010. CODEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (elec-

                                        tronic). URL https://www.
                                        liebertpub.com/doi/abs/
                                        10.1089/cmb.2009.0062;
                                         https://www.liebertpub.
                                        com/doi/pdf/10.1089/cmb.
                                        2009.0062.

Stone:1994:PSO

[SSN94] L. C. Stone, S. B. Shukla, and B. Neta. Parallel satellite orbit prediction using a workstation cluster. Computers and Mathematics with Applications, 28(8): 1–8, October 1994. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Schuchart:2021:CBC

[SSN^{+}21] Joseph Schuchart, Philipp Samfass, Christoph Niethammer, José Gracia, and George Bosilca. Callbackbased completion notification using MPI continuations. Parallel Computing, 106(??):??, September 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819121000466.

Shelton:1994:FPS

[SSP^{+}94] W. A. Shelton, G. M. Stocks, F. J. Pinski, R. G. Jordan, Y. Liu, L. Qui, J. B. Staunton, D. D. Johnson, and B. Ginatempo. First principles simulation of materials properties. In Pierce and Regnier [PR94b], pages 103–110. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.58.S32 1994. IEEE catalog no. 94TH0637-9.

Sen:1999:PBD

[SSS99] Vikramaditya Sen, Mrinal K. Sen, and Paul L. Stoffa. PVM based 3-D Kirchhoff depth migration using dynamically computed travel-times: an application in seismic data processing. Parallel Computing, 25(3):231–248, March 22, 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.

                                        elsevier.com/cas/tree/
                                        store/parco/sub/1999/25/
                                        3/1389.pdf.

Santana:1996:PVM

[SSSS96] M. S. Santana, P. S. Souza, R. C. Santana, and S. S. Souzza. Parallel Virtual Machine for Windows95. In Bode et al. [BDLS96], pages 288–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Souza:1997:EPH

[SSSS97] P. S. Souza, L. J. Senger, M. J. Santana, and R. C. Santana. Evaluating personal high performance computing with PVM on Windows and LINUX environments. Lecture Notes in Computer Science, 1332:49– 56, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stellner:1997:LBB

[ST97] G. Stellner and J. Trinitis. Load balancing based on process migration for MPI. Lecture Notes in Computer Science, 1300:150–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smyk:2002:AMM

[ST02a] Adam Smyk and Marek Tudruj. Application of mixed MPI OpenMP programming in a multi SMP cluster computer. Lecture Notes in Computer Science, 2328:288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2328/23280288.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2328/23280288.
pdf.

Smyk:2002:OMP

[ST02b] Adam Smyk and Marek Tudruj. OpenMP / mpi programming in a multi-cluster system based on shared memory/message passing communication. Lecture Notes in Computer Science, 2326:241–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2326/23260241.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2326/23260241.
pdf.

Steele:2017:UBP

[ST17] Guy L. Steele, Jr. and Jean-Baptiste Tristan. Using butterfly-patterned partial sums to draw from discrete distributions. ACM SIG-PLAN Notices, 52(8):341– 355, August 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stals:1995:AMP

[Sta95a] L. Stals. Adaptive multigrid in parallel. In Bailey et al. [BBG^{+}95], pages 367–

372. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Stankovski:1995:MPA

[Sta95b] Z. Stankovski. A massively parallel algorithm for the collision probability calculations in the APOLLO-II code using the PVM library. In ANS [ANS95], pages 1573–1583. ISBN 0-89448-198-3. LCCN TK9006.M37 1995. Two volumes.

Salinas:2020:FEI

[STA20] Álvaro Salinas, Claudio Torres, and Orlando Ayala. A fast and efficient integration of boundary conditions into a unified CUDA kernel for a shallow water solver lattice Boltzmann method. Computer Physics Communications, 249(??): Article 107009, April 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465519303443.

Stephens:1994:PBT

[Ste94] R. Stephens. Parallel benchmarks on the Transtech Paramid supercomputer. In de Gloria et al. [dGJM94], pages 136–146. ISBN ???? LCCN ????

Stellner:1996:CCP

[Ste96]  G. Stellner.  CoCheck:

checkpointing and process migration for MPI. In IEEE [IEE96e], pages 526– 531. ISBN 0-8186-7255-2. LCCN QA76.58 .I565 1996. IEEE catalog number 96TB100038. IEEE Computer Society Press order number PR07255.

Sterling:2000:SCB

[Ste00] Thomas Sterling. Symbolic computing with Beowulfclass PC clusters. Lecture Notes in Computer Science, 1908:7–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080007.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080007.
pdf.

Smith:2022:PAM

[STH22] Matthew Smith, Arjen Tamerus, and Phil Hasnip. Portable acceleration of materials modeling software: CASTEP, GPUs, and OpenACC. Computing in Science and Engineering, 24(1):46–55, January/ February 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Still:1994:PPC

[Sti94] C. H. Still. Portable parallel computing via the MPI1 message-passing standard. Computers in Physics, 8(5): 533–536, 538–539, September– October 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Schmitz:2008:IIG

[STK08] Arne Schmitz, Markus Tavenrath, and Leif Kobbelt. Illumination: Interactive global illumination for deformable geometry in CUDA. Computer Graphics Forum, 27(7):1979–1986, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sunderam:1997:TAS

[STMK97] V. Sunderam, B. Topol, S. Moyer, and A. Krantz. Tools and auxiliary subsystems in PVM. Lecture Notes in Computer Science, 1332:285–294, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stockinger:1998:VPC

[Sto98] Kurt Stockinger. ViMPIOS — a portable, client-server based implementation of MPI-IO on ViPIOS. Diplom-Arbeit, Universität Wien, Vienna, Austria, 1998. 155 pp.

Stpiczynski:2002:PPO

[Stp02]  Przemyslaw  Stpiczynski.

Parallel Programming in OpenMP helps novices: a review of Parallel Programming in OpenMP by Rohit Chandra, Leonardo Dagum, Dave Kohr, Dror Maydan, Jeff McDonald, and Ramesh Menon. IEEE Distributed Systems Online, 3(8), 2002. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:/

/dsonline.computer.org/
0208/d/bks_a.htm.

Stpiczynski:2018:LBV

[Stp18]  Przemysław  Stpiczyński.

Language-based vectorization and parallelization using intrinsics, OpenMP, TBB and Cilk Plus. The Journal of Supercomputing, 74(4):1461–1472, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

link.springer.com/content/
pdf/10.1007/s11227-017-
2231-3.pdf.

Sala:2019:IBN

[STP^{+}19] Kevin Sala, Xavier Teruel, Josep M. Perez, Antonio J. Peña, Vicenç Beltran, and Jesus Labarta. Integrating blocking and non-blocking MPI primitives with taskbased programming models. Parallel Computing, 85(??):153–166, July 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819118303326.

Stpiczynski:2020:ALB

[Stp20] Przemysław Stpiczyński. Algorithmic and languagebased optimization of Marsa-LFIB4 pseudorandom number generator using OpenMP, OpenACC and CUDA. Journal of Parallel and Distributed Computing, 137 (??):238–245, March 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731519304885.

Strok:1994:NJI

[Str94] Dale C. Strok. In the news: Jupiter impacts: Resolution makes a big difference. supercomputer farming down under. HPF Forum welcomes comments. Smithsonian Awards honor computational scientists. lowlife computer viruses. PVM developers get R&D-100 award. the eyes have it. neural nets detect breast cancer. better cars through cooperation. parallel version of global climate model. Lockheed to run Idaho National Engineering Lab. publicprivate partners: new drugs, new software. IEEE Computational Science & Engineering, 1(3):88–90, Fall 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Strietzel:1996:PTS

[Str96] M. Strietzel. Parallel turbulence simulation based on MPI. In Liddell et al. [LCHS96], pages 283–289. ISBN 3-540-61142-8 (paperback). LCCN QA76.88 .H52 1996.

Strietzel:1997:PTS

[Str97] M. Strietzel. Parallel turbulence simulation: Resolving the inertial subrange of Kolmogorov’s spectra. Lecture Notes in Computer Science, 1332:508–516, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strzodka:2012:DLO

[Str12]  Robert Strzodka.  Data

layout optimization for multi-valued containers in OpenCL. Journal of Parallel and Distributed Computing, 72(9):1073–1082, September 2012. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0743731511002115.

Soch:1996:PCG

[STT96] M. Soch, J. Trdlicka, and P. Tvrdik. PVM, computational geometry, and parallel computing course. In Bode et al. [BDLS96], pages 38–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Soch:1997:PGP

[STV97] M. Soch, P. Tvrdik, and M. Volf. Parallel graphpartitioning using the mob heuristic. Lecture Notes in Computer Science, 1332: 383–389, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shen:1999:ATL

[STY99] Kai Shen, Hong Tang, and Tao Yang. Adaptive twolevel thread management for fast MPI execution on shared memory machines. In ACM [ACM99], page ??

Stone:1996:RNF

[SU96] J. Stone and M. Underwood. Rendering of numerical flow simulations using MPI. In IEEE [IEE96i], pages 138– 141. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Sumimoto:2012:MCL

[Sum12] Shinji Sumimoto. The MPI Communication Library for the K computer: Its design and implementation. Lecture Notes in Computer Science, 7490:11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/accesspage/
chapter/10.1007/978-3-
642-33518-1_3.

Sunderam:1990:PFPa

[Sun90a] V. S. Sunderam. PVM: a framework for parallel distributed computing. Technical Report ORNL/TM-11375, Dept. of Math and Computer Science, Emory University, Atlanta, GA, USA, February 1990. See also [Sun90b].

Sunderam:1990:PFPb

[Sun90b] V. S. Sunderam. PVM: a framework for parallel distributed computing. Concurrency: practice and experience, 2(4):315–339, December 1990. CODEN CPEXEI. ISSN 1040-3108. See also the earlier technical report [Sun90a].

Sunderam:1992:CCP

[Sun92] Vaidy Sunderam. Concurrent computing with PVM. In SCRI WCC’92 [SCR92], page ?? ISBN ???? LCCN ???? Proceedings available via anonymous

ftp from ftp.scri.fsu.edu
in directory pub/parallel-
workshop.92.

Sunderam:1993:PCC

[Sun93] V. Sunderam. The PVM concurrent computing system. In Anonymous [Ano93h], pages 20–84. ISBN ???? LCCN ????

Sunderam:1994:GPP

[Sun94a]  V. Sunderam.  General

purpose parallel computing with PVM. In Anonymous [Ano94f], pages 185– 198. ISBN ???? LCCN ????

Sunderam:1994:MSH

[Sun94b] V. S. Sunderam. Methodologies and systems for heterogeneous concurrent computing. In Joubert et al. [JPTE94], pages 29–45. ISBN 0-444-81841-3. LCCN QA76.58 .P3794 1993.

Sunderam:1995:RIH

[Sun95] V. S. Sunderam. Recent initiatives in heterogeneous parallel computing. In Gray and Naghdy [GN95], pages 1–16. ISBN ???? LCCN ????

Sunderam:1996:PSS

[Sun96] V. Sunderam. The PVM system: status, trends, and directions. In Bode et al. [BDLS96], pages 68–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Suresh:1995:IOP

[Sur95a] H. Suresh. Implementation of an optimal parallel algorithm for arithmetic expression parsing. In Narashimhan [Nar95], page 925 vol.2. ISBN 0-7803-2018-2 (paperback), 0-7803-2019-0 (microfiche). LCCN QA76.6.I15 1995. Two volumes. IEEE catalog no. 95TH0682-5.

Suresh:1995:PIQ

[Sur95b] H. Suresh. PVM implementation of quadtree building algorithms on SIMD hypercube system. In Narashimhan [Nar95], pages 855–858 (vol. 2). ISBN 0-7803-2018-2 (paperback), 0-7803-2019-0 (microfiche). LCCN QA76.6.I15 1995. Two volumes. IEEE catalog no. 95TH0682-5.

Suttner:1996:SPB

[Sut96] C. B. Suttner. SPTHEO — a PVM-based parallel theorem prover. Lecture Notes in Computer Science, 1156:116–125, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smelyanskiy:2011:HPL

[SVC^{+}11] Mikhail Smelyanskiy, Karthikeyan Vaidyanathan, Jee Choi, Bálint Joó, Jatin Chhugani, Michael A. Clark, and Pradeep Dubey. Highperformance lattice QCD for multi-core based parallel systems using a cache-friendly hybrid threaded-MPI approach. In Lathrop et al.

[LCK11], pages 69:1–69:11. ISBN 1-4503-0771-X. LCCN ????

Sistare:1999:OMC

[SvL99] Steve Sistare, Rolf vande-Vaart, and Eugene Loh. Optimization of MPI collectives on clusters of large-scale SMPs. In ACM [ACM99], page ??

Stout:1991:SDM

[SW91] Quentin F. Stout and Michael Joseph Wolfe, editors. The Sixth Distributed Memory Computing Conference proceedings April 28– May 1, 1991, Portland, Oregon. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-2291-1. LCCN QA76.5 .D58 1991.

Sehrish:2012:RFS

[SW12] Saba Sehrish and Jun Wang. Reduced Function Set Abstraction (RFSA) for MPI-IO. The Journal of Supercomputing, 59(1):131– 146, January 2012. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0920-8542&volume=
                                        59&issue=1&spage=131.

Swann:2001:SPC

[Swa01] Christopher A. Swann. Software for parallel computing: the LAM implementation of MPI. Journal of Applied Econometrics, 16(2): 185–194, March–April 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Samfass:2020:LTO

[SWCB20] Philipp Samfass, Tobias Weinzierl, Dominic E. Charrier, and Michael Bader. Lightweight task offloading exploiting MPI wait times for parallel adaptive mesh refinement. Concurrency and Computation: Practice and Experience, 32(24): e5916:1–e5916:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sosonkina:2015:RAV

[SWH15] Masha Sosonkina, Layne T.

Watson, and Jian He. Remark on algorithm 897: VT-DIRECT95: Serial and parallel codes for the global optimization algorithm DI-RECT. ACM Transactions on Mathematical Software, 41(3):22:1–22:2, June 2015. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [HWS09].

Santhanaraman:2005:DZC

[SWHP05] Gopalakrishnan Santhanaraman, Jiesheng Wu, Wei Huang, and Dhabaleswar K. Panda. Designing zerocopy Message Passing Interface derived datatype communication over Infiniband: Alternative approaches and performance evaluation. The International Journal of High Performance Computing Applications, 19(2):129– 142, Summer 2005. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/19/
2/129.full.pdf+html.

Sitsky:1995:IPM

[SWJ95] D. Sitsky, D. Walsh, and C. Johnson. Implementation and performance of the MPI message passing interface on the Fujitsu AP1000 multicomputer. Australian Computer Science Communications, 17(1):475–481, ???? 1995. CODEN ACSCDD. ISSN 0157-3055.

Skjellum:2001:OOA

[SWL^{+}01] Anthony Skjellum, Diane G. Wooley, Ziyang Lu, Michael Wolf, Purushotham V. Bangalore, Andrew Lumsdaine, Jeffrey M. Squyres, and Brian McCandless. Objectoriented analysis and design of the Message Passing Interface. Concurrency and Computation: Practice and Experience, 13(4):245–292, April 10, 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/

cgi-bin/abstract/78502300/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=78502300&PLACEBO=IE.
pdf.

Shan:2012:PEH

[SWS^{+}12] Hongzhang Shan, Nicholas J.

Wright, John Shalf, Katherine Yelick, Marcus Wagner, and Nathan Wichmann. A preliminary evaluation of the hardware acceleration of the Cray Gemini interconnect for PGAS languages and comparison with MPI. ACM SIGMETRICS Performance Evaluation Review, 40(2): 92–98, September 2012. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Shee:1994:DMA

[SWYC94] Jang Chung Shee, Chao Chin Wu, Lin Wen You, and Cheng Chen. Design of a multithread architecture and its parallel simulation and evaluation environment. In Anonymous [Ano94a], pages 69–76 (vol. 1). ISBN ???? LCCN ???? 2 vol.

Sotiriou-Xanthopoulos:2018:OBV

[SXMX^{+}18] Efstathios Sotiriou-Xanthopoulos, Leonard Masing, Sotirios Xydis, Kostas Siozios, JÜrgen Becker, and Dimitrios Soudris. OpenCL-based virtual prototyping and simulation of many-accelerator architectures. ACM Transactions on Embedded Computing Systems, 17(5):86:1–86:??, November 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3242179.

Stathopoulos:1995:DLB

[SY95] A. Stathopoulos and A. Ynnerman. Dynamic load balancing of atomic structure programs on a PVM cluster. In Hertzberger and Serazzi [HS95a], pages 384–391. ISBN 3-540-59393-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Sydow:1994:PSA

[Syd94] A. Sydow. Parallel simulation of air pollution. In Pehrson et al. [PSB^{+}94], pages 605–612. CODEN ITATEC. ISBN 0-444-81990-8, 0-444-81989-4. ISSN 0926-5473. LCCN QA75.5.I3785 1994. Three volumes.

Stathopoulos:1996:PIM

[SYF96] Andreas Stathopoulos, Anders B. Ynnerman, and Charlotte Froese Fischer. A PVM implementation of the MCHF atomic structure package. International Journal of Supercomputer Applications and High Performance Computing, 10(1):41– 61, Spring 1996. CODEN IJSCFG. ISSN 1078-3482.

Song:2019:PGA

[SYL19] You Song, Siyu Yang, and Jinzhi Lei. ParaCells: a GPU architecture for cellcentered models in computational biology. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 16(3):994–1006, May 2019. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Schneider:2009:CPM

[SYR^{+}09] Scott Schneider, Jae-Seung Yeom, Benjamin Rose, John C. Linford, Adrian Sandu, and Dimitrios S. Nikolopoulos. A comparison of programming models for multiprocessors with explicitly managed memory hierarchies. ACM SIG-PLAN Notices, 44(4):131– 140, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stankovic:1999:NVJ

[SZ99] N. Stankovic and K. Zhang.

Native versus Java message passing. In Dongarra et al. [DLM99], pages 165– 172. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Siegel:2011:AFV

[SZ11] Stephen F. Siegel and Timothy K. Zirkel. Automatic formal verification of MPIbased parallel programs. ACM SIGPLAN Notices, 46 (8):309–310, August 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’11 Conference proceedings.

Simmunovic:1995:MIP

[SZBS95a] S. Simmunovic, T. Zacharia, N. Baltas, and D. B. Spalding. MPI implementation of Phoenics: a general purpose computational fluid dynamics code. In Tentner [Ten95], pages 122–127. ISBN 1-56555-078-1. LCCN ????

Simunovic:1995:MIP

[SZBS95b] S. Simunovic, T. Zacharia, N. Baltas, and D. B. Spalding. MPI implementation of PHOENICS: a general purpose computational fluid dynamics code. In Tentner [Ten95], pages 122–127. ISBN 1-56555-078-1. LCCN ????

Thompson:2014:CIC

[TA14] Elizabeth A. Thompson and Timothy R. Anderson. A CUDA implementation of the Continuous Space Language Model. The Journal of Supercomputing, 68(1): 65–86, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11227-013-1023-7.

Taft:2021:LMA

[Taf21] S. Tucker Taft. A layered mapping of Ada 202X to OpenMP. ACM SIGADA Ada Letters, 40(2):55–58, April 2021. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463478.3463482.

Takeda:2001:AME

[TAH^{+}01] K. Takeda, N. K. Allsopp, J. C. Hardwick, P. C. Macey, D. A. Nicole, S. J. Cox, and D. J. Lancaster. An assessment of MPI environments for Windows NT. The Journal of Supercomputing, 19(3):315–323, July 2001. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/www.wkap.nl/oasis.htm/
338207.

Traff:2014:SPE

[TB14] Jesper Larsson Träff and Siegfried Benkner. Selected papers from EuroMPI 2012. Computing, 96(4):259– 261, April 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00607-013-0335-z.

Tao:2012:UGA

[TBB12] Jian Tao, Marek Blazewicz, and Steven R. Brandt. Using GPU’s to accelerate stencilbased computation kernels for the development of large scale scientific applications on heterogeneous systems. ACM SIGPLAN Notices, 47 (8):287–288, August 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Touhafi:1996:DPC

[TBD96] A. Touhafi, W. Brissinck, and E. F. Dirkx. Development of PVM code for a low latency switch based interconnect. In Bode et al. [BDLS96], pages 229–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Traff:2012:RAM

[TBD12] Jesper Larsson Träff, Siegfried Benkner, and Jack J. Dongarra, editors. Recent Advances in the Message Passing Interface: 19th European MPI Users’ Group Meeting, EuroMPI 2012, Vienna, Austria, September 23–26, 2012. Proceedings, volume 7490 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CO-DEN LNCSD9. ISBN 3-642-33517-9 (print), 3-642-33518-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:/

/www.springerlink.com/
content/978-3-642-33518-
1.

Tian:2002:IOC

[TBG^{+}02] Xinmin Tian, Aart Bik, Milind Girkar, Paul Grey, Hideki Saito, and Ernesto Su. Intel(R) OpenMP C++/Fortran compiler for hyper-threading technology: Implementation and performance. Intel Technology Journal, 6(1):36– 46, February 2002. ISSN 1535-766X. URL http:

//developer.intel.com/
technology/itj/2002/volume06issue01/
vol6iss1_hyper_threading_
technology.pdf.

Tahan:2012:ITC

[TBS12] Oussama Tahan, Mats Brorsson, and Mohamed Shawky. Introducing task cancellation to OpenMP. Lecture Notes in Computer Science, 7312:73–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/
10.1007/978-3-642-30961-
8_6/.

Thomas:1994:PSA

[TC94] S. J. Thomas and J. Cote.

Parallel Semi-Lagrangian advection using PVM. In Dekker et al. [DSZ94], pages 801–808. ISBN 0-444-81784-

0. LCCN QA76.58.E98 1994.

Tzannes:2010:LBS

[TCBV10] Alexandros Tzannes, George C. Caragea, Rajeev Barua, and Uzi Vishkin. Lazy binarysplitting: a run-time adaptive work-stealing scheduler. ACM SIGPLAN Notices, 45(5):179–190, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tagliavini:2018:UFG

[TCM18] Giuseppe Tagliavini, Daniele Cesarini, and Andrea Marongiu. Unleashing fine-grained parallelism on embedded manycore accelerators with lightweight OpenMP tasking. IEEE Transactions on Parallel and Distributed Systems, 29(9): 2150–2163, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/09/08314096-
                                        abs.html.

Thompson:2015:PCI

[TCP15] Elizabeth Thompson, Nathan Clem, and David A. Peter. Parallel CUDA implementation of conflict detection for application to airspace deconfliction. The Journal of Supercomputing, 71(10):3787–3810, October 2015. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/s11227-015-1467-z.

Tourino:1998:PBL

[TD98] J. Touriño and R. Doallo.

A PVM-based library for sparse matrix factorizations. Lecture Notes in Computer Science, 1497:304–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tourino:1999:MMC

[TD99] J. Touriño and R. Doallo.

Modeling MPI collective communications on the AP3000 Multicomputer. In Dongarra et al. [DLM99], pages 133–140. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Thiruvathukal:2000:JNW

[TDB00] George K. Thiruvathukal, Phillip M. Dickens, and Shahzad Bhatti. Java on networks of workstations (JavaNOW): a parallel computing framework inspired by Linda and the Message Passing Interface (MPI). Concurrency: practice and experience, 12(11):1093–1116, September 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76000187/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?

ID=76000187&PLACEBO=IE.
pdf.

Tromeur-Dervout:2011:PCF

[TDBEE11] Damien Tromeur-Dervout, Gunther Brenner, David R. Emerson, and Jocelyne Erhel, editors. Parallel Computational Fluid Dynamics 2008: Parallel Numerical Methods, Software Development and Applications, volume 74 of Lecture Notes in Computational Science and Engineering. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CO-DEN LNCSA6. ISBN 3-642-14437-3 (print), 3-642-14438-1 (e-book). ISSN 1439-7358. LCCN ???? URL

http://link.springer.com/
book/10.1007/978-3-642-
14438-7;   http://www.
springerlink.com/content/
978-3-642-14438-7. Pro-

ceedings of the twentieth meeting, Parallel CFD 2008, held May 19–22, 2008 in Lyon, France.

Totoni:2013:EFE

[TDG13] Ehsan Totoni, Mert Dikmen, and María Jesús Garzarán. Easy, fast, and energyefficient object detection on heterogeneous on-chip architectures. ACM Transactions on Architecture and Code Optimization, 10(4): 45:1–45:??, December 2013. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Tentner:1995:HPC

[Ten95] A. Tentner, editor. High Performance Computing Symposium 1995 ‘Grand Challenges in Computer Simulation’. Proceedings of the 1995 Simulation Multiconference: Phoenix, AZ, USA, 9–13 April 1995. Society for Computer Simulation, San Diego, CA, USA, 1995. ISBN 1-56555-078-1. LCCN ????

Truong:2002:PAM

[TFGM02] Hong-Linh Truong, Thomas Fahringer, Michael Geissler, and Georg Madsen. Performance analysis for MPI applications with SCALEA. Lecture Notes in Computer Science, 2474:421–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/
service/series/0558/bibs/
2474/24740421.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740421.pdf.

Tu:2012:PAO

[TFZZ12] Bibo Tu, Jianping Fan, Jianfeng Zhan, and Xiaofang Zhao. Performance analysis and optimization of MPI collective operations on multi-core clusters. The Journal of Supercomputing, 60(1):141–162, April 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0920-8542&volume=
                                        60&issue=1&spage=141.

Turchi:1994:SDA

[TG94] Patrice E. A. Turchi and Antonios Gonis, editors. Statics and dynamics of alloy phase transformations: Proceedings of a NATO Advanced Study Institute on Statics and Dynamics of Alloy Phase Transformations, held June 21–July 3, 1992, in Rhodes, Greece, volume 319 of NATO ASI Series B Physics. Plenum Press, New York, NY, USA, 1994. ISBN 0-306-44626-X. ISSN 0258-1221. LCCN TN690.S77 1994.

Thakur:2009:TSE

[TG09] Rajeev Thakur and William Gropp. Test suite for evaluating performance of multithreaded MPI communication. Parallel Computing, 35(12):608–617, December 2009. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tian:2005:PCT

[TGBS05] Xinmin Tian, Milind Girkar, Aart Bik, and Hideki Saito. Practical compiler techniques on efficient multithreaded code generation for OpenMP programs. The Computer Journal, 48(5): 588–601, September 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
48/5/588;   http://comjnl.
oxfordjournals.org/cgi/
reprint/48/5/588.

Tuncer:2009:PCF

[TGEM09] Ismail H. Tuncer, Ülgen Gülcat, David R. Emerson, and Kenichi Matsuno, editors. Parallel Computational Fluid Dynamics 2007: Implementations and Experiences on Large Scale and Grid Computing, volume 67 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CO-DEN LNCSA6. ISBN 3-540-92743-3 (print), 3-540-92744-1 (e-book). ISSN 1439-7358. LCCN ???? URL

http://link.springer.com/
book/10.1007/978-3-540-
92744-0;   http://www.
springerlink.com/content/
978-3-540-92744-0. Paral-

lel CFD 2007 was held in Antalya, Turkey, from May 21 to 24, 2007.

Tian:2019:GAB

[TGKL19] Tian Tian, Dunwei Gong, Fei-Ching Kuo, and Huai Liu. Genetic algorithm based test data generation for MPI parallel programs with blocking communication. The Journal of Systems and Software, 155(??): 130–144, September 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0164121219300810.

Thakur:2002:ONA

[TGL02] Rajeev Thakur, William Gropp, and Ewing Lusk. Optimizing noncontiguous accesses in MPI-IO. Parallel Computing, 28(1):83– 105, January 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        35/21/60/27/32/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/35/21/60/
                                        27/32/00001686.pdf.

Tsiolakis:2020:NPG

[TGS^{+}20] Vasileios Tsiolakis, Matteo Giacomini, Ruben Sevilla, Carsten Othmer, and Antonio Huerta. Nonintrusive proper generalised decomposition for parametrised incompressible flow problems in OpenFOAM. Computer Physics Communications, 249(??):Article 107013, April 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465519303479.

Thakur:2005:OSO

[TGT05] Rajeev Thakur, William Gropp, and Brian Toonen. Optimizing the synchronization operations in Message Passing Interface one-sided communication. The International Journal of High Performance Computing Applications, 19(2):119– 128, Summer 2005. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/19/
2/119.full.pdf+html.

Traff:2010:SCM

[TGT10] Jesper Larsson Traff, William D.

Gropp, and Rajeev Thakur. Self-consistent MPI performance guidelines. IEEE Transactions on Parallel and Distributed Systems, 21(5): 698–709, May 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Traff:2020:SIS

[TH20] Jesper Larsson Träff and Torsten Hoefler. Special issue: Selected papers from EuroMPI 2019. Parallel Computing, 99(??):Article 102695, November 2020.

CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819120300855.

Thakur:1998:CUM

[Tha98]  Rajeev S. Thakur.  A

case for using MPI’s derived datatypes to improve I/O performance. In ACM [ACM98b], page ?? ISBN ???? LCCN

                                        ????  URL http://
                                        www.supercomp.org/sc98/
                                        papers/.

Teijeiro:2019:OPS

[THDS19] Carlos Teijeiro, Thomas Hammerschmidt, Ralf Drautz, and Godehard Sutmann. Optimized parallel simulations of analytic bondorder potentials on hybrid shared/distributed memory with MPI and OpenMP. The International Journal of High Performance Computing Applications, 33(2): 227–241, March 1, 2019. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:

                                        //journals.sagepub.com/
                                        doi/full/10.1177/1094342017727060.

Tian:2005:CEN

[THH^{+}05] Xinmin Tian, Jay P. Hoeflinger, Grant Haab, Yen-Kuang Chen, Milind Girkar, and Sanjiv Shah. A compiler for exploiting nested parallelism in OpenMP programs. Parallel Computing, 31(10–12):960–983, October/December 2005. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Trefftz:1994:DPE

[THM^{+}94] C. Trefftz, C. C. Huang, P. K. McKinley, T. Y. Li, and Z. Zeng. Design and performance evaluation of a distributed eigenvalue solver on a workstation cluster. In IEEE [IEE94b], pages 608–615. ISBN 0-8186-6952-7 (casebound), 0-8186-6950-0 (paperback), 0-8186-6951-9 (microfiche). LCCN TA1637.I25 1994. Three volumes. IEEE catalog no. 94CH35708.

Traff:2021:MCC

[THMH21] Jesper Larsson Träff, Sascha Hunold, Guillaume Mercier, and Daniel J. Holmes. MPI collective communication through a single set of interfaces: a case for orthogonality. Parallel Computing, 107(??):??, October 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819121000740.

Tran:2000:PPM

[THN00] Viet D. Tran, Ladislav Hluchy, and Giang T. Nguyen. Parallel program model for distributed systems. Lecture Notes in Computer Science, 1908: 250–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1908/19080250.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1908/19080250.
                                        pdf.

Thomsen:1994:RTS

[Tho94]  P. G. Thomsen.  Real

time simulation in a cluster computing environment. In Dongarra and Wasniewski [DW94], pages 493– 497. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1994. DM104.00.

Throop:1999:SOS

[Thr99] Joe Throop. Standards: OpenMP: Shared-memory parallelism from the ashes. Computer, 32(5):108–109, May 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.
                                        computer.org/co/books/
                                        co1999/pdf/r5108.pdf.

Traeff:1999:FFE

[THRZ99] J. L. Traeff, R. Hempel, H. Ritzdoff, and F. Zimmermann. Flattening on the fly: Efficient handling of MPI derived datatypes. In Dongarra et al. [DLM99], pages 109–116. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Takizawa:2015:ODT

[THS^{+}15] Hiroyuki Takizawa, Shoichi Hirasawa, Makoto Sugawara, Isaac Gelado, Hiroaki Kobayashi, and Wen mei W. Hwu. Optimized data transfers based on the OpenCL event management mechanism. Scientific Programming, 2015(??): 576498:1–576498:16, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/576498/

.

Tabakin:2009:QPE

[TJD09] Frank Tabakin and Bruno Juliá-Díaz. QCMPI: a parallel environment for quantum computing. Computer Physics Communications, 180(6):948–964, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465508004141.

Thoman:2012:AOL

[TJPF12] Peter Thoman, Herbert Jordan, Simone Pellegrini, and Thomas Fahringer. Automatic OpenMP loop scheduling: a combined compiler and runtime approach. Lecture Notes in Computer Science, 7312: 88–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/10.
                                        1007/978-3-642-30961-8_
                                        7/.

Tang:2016:AKM

[TK16] Qing Y. Tang and Mohammed A. S. Khalid. Acceleration of k-means algorithm using Altera SDK for OpenCL. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(1):6:1–6:??, December 2016. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Teunissen:2019:GML

[TK19] J. Teunissen and R. Keppens. A geometric multigrid library for quadtree/ octree AMR grids coupled to MPI-AMRVAC. Computer Physics Communications, 245(??):Article 106866, December 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S001046551930253X.

Tennyson:2015:MOI

[TKP15] P. Gerald Tennyson, G. M. Karthik, and G. Phanikumar. MPI + OpenCL implementation of a phasefield method incorporating CALPHAD description of Gibbs energies on heterogeneous computing platforms. Computer Physics Communications, 186(??): 48–64, January 2015. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465514003208.

Tu:2019:AOS

[TL19] Chia-Heng Tu and Te-Sheng Lin. Augmenting operating systems with OpenCL accelerators. ACM Transactions on Design Automation of Electronic Systems, 24(3):30:1–30:29, June 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3315569.

Tallent:2009:EPM

[TMC09] Nathan R. Tallent and John M. Mellor-Crummey. Effective performance measurement and analysis of multithreaded applications. ACM SIGPLAN Notices, 44(4):229–240, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tampouratzis:2016:AIH

[TMP16] Nikolaos Tampouratzis, Pavlos M. Mattheakis, and Ioannis Papaefstathiou. Accelerating intercommunication in highly parallel systems. ACM Transactions on Architecture and Code Optimization, 13(4):40:1–40:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Trobec:2001:IEM

[TMPJ01] R. Trobec, M.Šterk, M. Praprotnik, and D. Janežič. Implementation and evaluation of MPI-based parallel MD program. International Journal of Quantum Chemistry, 84(1):23– 31, ???? 2001. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (elec-

                                        tronic). URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/84002438/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext/
                                        84002438/FILE?TPL=ftx_
                                        start;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=84002438&PLACEBO=IE.
                                        pdf.

Tiotto:2020:OCO

[TMT^{+}20] E. Tiotto, B. Mahjour, W. Tsang, X. Xue, T. Islam, and W. Chen. OpenMP 4.5 compiler optimization for GPU offloading. IBM Journal of Research and Development, 64(3/4):14:1–14:11, May/July 2020. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Theodoropoulos:1996:ESP

[TMTP96] P. Theodoropoulos, G. Manis, P. Tsanakas, and G. Papakonstantinou. Extending synchronization PVM mechanisms. In Bode et al. [BDLS96], pages 315–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Taylor:2017:AOO

[TMW17] Ben Taylor, Vicent Sanz Marco, and Zheng Wang. Adaptive optimization for OpenCL programs on embedded heterogeneous systems. ACM SIGPLAN Notices, 52(4):11–20, May 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Takafuji:2017:CCC

[TNIB17] Daisuke Takafuji, Koji Nakano, Yasuaki Ito, and Jacir Bordim. C2CU: a CUDA–C program generator for bulk execution of a sequential algorithm. Concurrency and Computation: Practice and Experience, 29 (17), September 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tracy:2018:CMC

[TOC18] Fred Thomas Tracy, Thomas C. Oppe, and Maureen K. Corcoran. A comparison of MPI and co-array FOR-TRAN for large finite element variably saturated flow simulations. Scalable Computing: Practice and Experience, 19(4):423–432, ???? 2018. CODEN ???? ISSN 1895-1767. URL https://

                                        www.scpe.org/index.php/
                                        scpe/article/view/1468.

Takahashi:1999:IEM

[TOTH99] T. Takahashi, F. O’Carroll, H. Tezuka, and A. Hori. Implementation and evaluation of MPI on an SMP cluster. Lecture Notes in Computer Science, 1586:1178–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toussaint:1996:AES

[Tou96] Marcel Toussaint, editor. Ada in Europe: Second International Eurospace-Ada-Europe Symposium, Frankfurt/Main, Germany, October 2–6, 1995: proceedings, number 1031 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-60757-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A35I57 1995.

Tourancheau:2000:HSN

[Tou00] Bernard Tourancheau. High speed networks for clusters, the BIP-Myrinet experience. Lecture Notes in Computer Science, 1908:9–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1908/19080009.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080009.
pdf.

Thebault:2015:SEI

[TPD15] Loïc Thébault, Eric Petit, and Quang Dinh. Scalable and efficient implementation of 3D unstructured meshes computation: a case study on matrix assembly. ACM SIGPLAN Notices, 50 (8):120–129, August 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tang:2019:QDL

[TPK^{+}19] Xulong Tang, Ashutosh Pattnaik, Onur Kayiran, Adwait Jog, Mahmut Taylan Kandemir, and Chita Das. Quantifying data locality in dynamic parallelism in GPUs. ACM SIGMETRICS Performance Evaluation Review, 47(1):25–26, December 2019. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3376930.3376947.

Tong:2018:FCM

[TPLY18] Zhou Tong, Scott Pakin, Michael Lang, and Xin Yuan. Fast classification of MPI applications using Lamport’s logical clocks. Journal of Parallel and Distributed Computing, 120 (??):77–88, October 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S074373151830340X.

Turchetto:2020:GDS

[TPV20] M. Turchetto, A. D. Palù, and R. Vacondio. A general design for a scalable MPI-GPU multi-resolution 2D numerical solver. IEEE Transactions on Parallel and Distributed Systems, 31 (5):1036–1047, May 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tinetti:2001:HNW

[TQDL01] Fernando Tinetti, Antonio Quijano, Armando De Giusti, and Emilio Luque. Heterogeneous networks of workstations and the parallel matrix multiplication.

Lecture Notes in Computer Science, 2131:296–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310296.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310296.
pdf.

Traeff:1998:PRL

[Tra98]  J. L. Traeff.  Portable

randomized list ranking on multiprocessors using MPI. Lecture Notes in Computer Science, 1497:395–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Traff:2002:IMP

[Tra02a] Jesper Larsson Traff. Implementing the MPI process topology mechanism. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN

???? URL http://www.sc-
2002.org/paperpdfs/pap.
pap122.pdf.

Traff:2002:IMA

[Trä02b] Jesper Larsson Träff. Improved MPI all-to-all communication on a Giganet SMP cluster. Lecture Notes in Computer Science, 2474: 392–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.de/link/
service/series/0558/bibs/
2474/24740392.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740392.pdf.

Traff:2012:AUE

[Trä12a] Jesper Larsson Träff. Alternative, uniformly expressive and more scalable interfaces for collective communication in MPI. Parallel Computing, 38(1–2):26– 36, January/February 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819111001402.

Traff:2012:MTM

[Trä12b] Jesper Larsson Träff. mpicroscope: Towards an MPI benchmark tool for performance guideline verification. Lecture Notes in Computer Science, 7490:100–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/
10.1007/978-3-642-33518-
1_15/.

Thakur:2005:OCC

[TRG05] Rajeev Thakur, Rolf Rabenseifner, and William Gropp. Optimization of collective communication operations in MPICH. The International Journal of High Performance Computing Applications, 19(1):49–66, Spring 2005. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/19/
1/49.full.pdf+html.

Traff:2000:IMO

[TRH00] Jesper Larsson Traff, Hubert Ritzdorf, and Rolf Hempel. The implementation of MPI-2 one-sided communication for the NEC SX-5. In ACM [ACM00], pages 45–

46.  URL http://www.
sc2000.org/proceedings/
techpapr/papers/pap181.
pdf.

Tahan:2012:UDT

[TS12a] Oussama Tahan and Mohamed Shawky. Using dynamic task level redundancy for OpenMP fault tolerance. Lecture Notes in Computer Science, 7179: 25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-28293-5_
3/.

Thibault:2012:AIF

[TS12b] Julien C. Thibault and Inanc Senocak. Accelerating incompressible flow computations with a Pthreads– CUDA implementation on small-footprint multi-GPU platforms. The Journal of Supercomputing, 59(2): 693–719, February 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
59&issue=2&spage=693.

Takahashi:2002:PEH

[TSB02] Daisuke Takahashi, Mitsuhisa Sato, and Taisuke Boku. Performance evaluation of the Hitachi SR8000 using OpenMP benchmarks. Lecture Notes in Computer Science, 2327:390–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2327/23270390.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2327/23270390.
pdf.

Takahashi:2003:PEH

[TSB03] Daisuke Takahashi, Mitsuhisa Sato, and Taisuke Boku. Performance evaluation of the Hitachi SR8000 using SPEC OMP2001 benchmarks. International Journal of Parallel Programming, 31(3):185– 196, June 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL /ips/frames/
Refs/referenceskapmain.

asp?J=4773&I=33&A=2&LK=
NM;   http://ipsapp007.
kluweronline.com/content/
getfile/4773/33/2/abstract.
htm;   http://ipsapp007.
kluweronline.com/content/
getfile/4773/33/2/fulltext.
pdf.

Terboven:2012:AOT

[TSCaM12] Christian Terboven, Dirk Schmidl, Tim Cramer, and Dieter an Mey. Assessing OpenMP tasking implementations on NUMA architectures. Lecture Notes in Computer Science, 7312: 182–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-30961-8_
14/.

Teixido:2014:MBI

[TSCS14] Ivan Teixidó, Francesc Sebé, Josep Conde, and Francesc Solsona. MPI-based implementation of an enhanced algorithm to solve the LPN problem in a memoryconstrained environment. Parallel Computing, 40(5– 6):100–112, May 2014. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819114000453.

Takizawa:2021:OLO

[TSEE21] Hiroyuki Takizawa, Shinji Shiotsuki, Naoki Ebata, and Ryusuke Egawa. OpenCLlike offloading with metaprogramming for SX-aurora TSUBASA. Parallel Computing, 102(??):Article 102754, May 2021. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167819121000144.

Tanaka:2021:NRP

[TSN21] Ushio Tanaka, Masami Saga, and Junji Nakano. NScluster: An R package for maximum palm likelihood estimation for cluster point process models using OpenMP. Journal of Statistical Software, 98 (??):??, ???? 2021. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL https://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v098i06;
                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v098i06/v98i06.pdf.

Ten:1995:TPE

[TSP95] S. V. Ten, V. V. Savchenko, and A. A. Pasko. Time performance evaluation of implicit surface polygonization on distributed systems. In Gray and Naghdy [GN95], pages 183–193. ISBN ???? LCCN ????

Topol:1998:PTV

[TSS98] Brad Topol, John T. Stasko, and Vaidy Sunderam. PVaniM: a tool for visualization in network computing environments. Concurrency: practice and experience, 10(14):1197–1222, December 10, 1998. CODEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract?ID=40005932;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=40005932&PLACEBO=IE.
pdf.

Tatebe:2000:IOO

[TSS00a] Osamu Tatebe, Mitsuhisa Sato, and Satoshi Sekiguchi. Impact of OpenMP optimizations for the MGCG method. Lecture Notes in Computer Science, 1940: 471–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1940/19400471.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1940/19400471.
pdf.

Tavora:2000:DCM

[TSS00b] Vítor N. Távora, Luís M.

Silva, and João Gabriel Silva. Distributed checkpointing mechanism for a parallel file system. Lecture Notes in Computer Science, 1908:137–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1908/19080137.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1908/19080137.
                                        pdf.

Tsunekawa:1995:EIE

[Tsu95] H. Tsunekawa. Effective implementation of EDEM workstation cluster using PVM. In Pahl and Werner [PW95], pages 503–508. ISBN 90-5410-556-9, 90-5410-557-7. LCCN TA345 .I565 1995 v.1-2. Two volumes.

Tsujita:2007:RMP

[Tsu07] Y. Tsujita. Remote MPI-I/O on a parallel virtual file system using a circular buffer for high throughput. International Journal of Computers and Applications, 29(3):291–299, 2007. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:

                                        //www.tandfonline.com/
                                        doi/full/10.1080/1206212X.
                                        2007.11441859.

Tsutsui:2012:AMG

[Tsu12] Shigeyoshi Tsutsui. ACO on multiple GPUs with CUDA for faster solution of QAPs. Lecture Notes in Computer Science, 7492: 174–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-32964-7_
18/.

Tang:1999:CRT

[TSY99] Hong Tang, Kai Shen, and Tao Yang. Compile/runtime support for threaded MPI execution on multiprogrammed shared memory machines. ACM SIG-PLAN Notices, 34(8):107– 118, August 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
proceedings/ppopp/301104/
p107-tang/.

Tang:2000:PTR

[TSY00] Hong Tang, Kai Shen, and Tao Yang. Program transformation and runtime support for threaded MPI execution on shared-memory machines. ACM Transactions on Programming Languages and Systems, 22(4): 673–700, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/toplas/2000-22-
4/p673-tang/.

Trelles-Salazar:1994:MSS

[TSZC94] O. Trelles-Salazar, E. L.

Zapata, and J.-M. Carazo.

Mapping strategies for sequential sequence comparison algorithms on LANbased message passing architectures. In Gentzsch and Harms [GH94], pages 197–202. ISBN 0-387-57981-8 (New York), 3-540-57981-8 (Berlin). LCCN QA76.88.I57 1994. DM96.00. Two volumes.

Theodoropoulos:1997:GSP

[TTP97] P. Theodoropoulos, P. Tsanakas, and G. Papakonstantinou. Global semaphores in a parallel programming environment. Lecture Notes in Computer Science, 1332: 151–158, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanaka:2000:PEO

[TTSY00] Yoshizumi Tanaka, Kenjiro Taura, Mitsuhisa Sato, and Akinori Yonezawa. Performance evaluation of OpenMP applications with nested parallelism. Lecture Notes in Computer Science, 1915:100–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1915/19150100.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1915/19150100.
pdf.

Tellez-Velazquez:2018:CSI

[TVCB18] Arturo Téllez-Velázquez and Raúl Cruz-Barbosa. A CUDA-streams inference machine for non-singleton fuzzy systems. Concurrency and Computation: Practice and Experience, 30 (8), April 25, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https:

//onlinelibrary.wiley.
com/doi/abs/10.1002/cpe.
4382.

Twerda:1996:PIT

[TVV96] A. Twerda, A. P. Van den Berg, and A. J. Van der Steen. Parallel implementation of time dependent Rayleigh-Benard convection. Supercomputer, 12(2):36–47, March 1996. CODEN SP-COEL. ISSN 0168-7875.

Tourancheau:2001:SMN

[TW01] Bernard Tourancheau and Roland Westrelin. Support for MPI at the network interface level. Lecture Notes in Computer Science, 2131:52–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310052.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310052.
pdf.

Thorson:2012:SUF

[TW12] Greg Thorson and Michael Woodacre. SGI UV2: a fused computation and data analysis machine. In Hollingsworth [Hol12], pages 105:1–105:?? ISBN 1-4673-0804-8. URL http:

                                        //conferences.computer.
                                        org/sc/2012/papers/1000a080.
                                        pdf.

Tournavitis:2009:THA

[TWFO09] Georgios Tournavitis, Zheng Wang, Björn Franke, and Michael F. P. O’Boyle. Towards a holistic approach to auto-parallelization: integrating profile-driven parallelism detection and machinelearning based mapping. ACM SIGPLAN Notices, 44(6):177–187, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tang:2019:MNT

[TWLL19] Yibin Tang, Ying Wang, Huawei Li, and Xiaowei Li. MV-Net: Toward realtime deep learning on mobile GPGPU systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(4):35:1– 35:??, December 2019. CO-DEN ???? ISSN 1550-

                                        4832. URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3358696.

Tien:2014:EOS

[TY14] Tsan-Rong Tien and Yi-Ping You. Enabling OpenCL support for GPGPU in kernel-based virtual machine. Software—Practice and Experience, 44(5):483– 510, May 2014. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Utterback:2017:POR

[UALK17] Robert Utterback, Kunal Agrawal, I-Ting Angelina Lee, and Milind Kulkarni. Processor-oblivious record and replay. ACM SIG-PLAN Notices, 52(8):145– 161, August 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Utterback:2019:POR

[UALK19] Robert Utterback, Kunal Agrawal, I-Ting Angelina Lee, and Milind Kulkarni. Processor-oblivious record and replay. ACM Transactions on Parallel Computing (TOPC), 6(4):20:1– 20:??, December 2019. CO-DEN ???? ISSN 2329-4949 (print), 2329-4957 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3365659.

Uselton:1995:PRS

[UCW95] Samuel P. Uselton, Michael Brian Cox, and Craig M. Wittenbrink, editors. 1995 Parallel Rendering Symposium (PRS 95): Atlanta, Georgia, October 30–31, 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-774-1 (softbound) [invalid checksum], 0-7803-3120-6 (microfiche). LCCN QA76.58.P3778 1995. ACM order number 428957. IEEE Computer Society Press order number 95TB8134.

Udupa:2009:SES

[UGT09] Abhishek Udupa, R. Govindarajan, and Matthew J. Thazhuthaveetil. Synergistic execution of stream programs on multicores with accelerators. ACM SIG-PLAN Notices, 44(7):99– 108, July 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Uhl:1996:PIC

[UH96] A. Uhl and J. Hammerle. Parallel image compression on a workstation cluster using PVM. In Bode et al. [BDLS96], pages 301–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Uhl:1994:PCC

[Uhl94] A. Uhl. Parallel compact coding of satellite images with wavelet packets using PVM. In Kumar [Kum94], pages 382–387. ISBN 0-07-462332-X. LCCN QA 76.58 I587 1994.

Uhl:1995:AWA

[Uhl95a] A. Uhl. Adapted wavelet analysis on moderate parallel distributed memory MIMD architectures. In Ferreira and Rolim [FR95], pages 275–283. ISBN 3-540-60321-2. LCCN QA76.642.I59 1995.

Uhl:1995:PCC

[Uhl95b] A. Uhl. Parallel compact coding of satellite images with wavelet packets using PVM. In Prasanna et al. [PBPT95], pages 382–387. ISBN 0-07-462332-X. LCCN QA 76.58 I587 1994.

Uhl:1995:VPW

[Uhl95c] A. Uhl. Vector and parallel wavelet transforms for the analysis of timevarying signals. In Bailey et al. [BBG^{+}95], pages 9–14. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Uminski:1997:EEP

[UMK97] P. W. Uminski, M. R. Matuszek, and H. Krawczyk. Experimental evaluation of PVM group communication. Lecture Notes in Computer Science, 1332:57–66, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uthayopas:2001:FSR

[UP01] Putchong Uthayopas and Sugree Phatanapherom. Fast and scalable real-time monitoring system for Beowulf clusters. Lecture Notes in Computer Science, 2131: 201–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310201.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310201.
pdf.

Urena:2012:IMI

[URKG12] Isaías A. Comprés Ureña, Michael Riepen, Michael Konow, and Michael Gerndt. Invasive MPI on Intel’s single-chip cloud computer. Lecture Notes in Computer Science, 7179:74–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/
10.1007/978-3-642-28293-
5_7/.

USENIX:1994:PFU

[USE94] USENIX, editor. Proceedings of the First USENIX Symposium on Operating Systems Design and Implementation (OSDI), November 14–17, 1994, Monterey, California, USA. USENIX, Berkeley, CA, USA, 1994. ISBN 1-880446-66-9. LCCN QA 76.76 O63 U87 1994.

USENIX:1995:PUT

[USE95] USENIX, editor. Proceedings of the 1995 USENIX Technical Conference, January 16–20, 1995, New Orleans, Louisiana, USA. USENIX, Berkeley, CA, USA, 1995. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

USENIX:2000:PAL

[USE00]  USENIX, editor.  Pro-

ceedings of the 4th Annual Linux Showcase and Conference, Atlanta, October 10– 14, 2000, Atlanta, Georgia, USA. USENIX, Berkeley, CA, USA, 2000. ISBN 1-880446-17-0. LCCN ????

URL http://www.usenix.
org/publications/library/
proceedings/als2000/.

Uehara:2002:MBP

[UTY02] Hitoshi Uehara, Masanori Tamura, and Mitsuo Yokokawa. An MPI benchmark program library and its application to the Earth simulator. Lecture Notes in Computer Science, 2327:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2327/23270219.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2327/23270219.
pdf.

Unat:2012:AFD

[UZC^{+}12] Didem Unat, Jun Zhou, Yifeng Cui, Scott B. Baden, and Xing Cai. Accelerating a 43D finite-difference earthquake simulation with a C-to-CUDA translator. Computing in Science and Engineering, 14(3):48–59, May/June 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

vanderPas:1993:PIG

[van93] R. van der Pas. The PVM implementation of a Generalized Red Black algorithm. Supercomputer, 10(4-5):72– 85, July–September 1993. CODEN SPCOEL. ISSN 0168-7875.

VanKatwijk:1995:AAC

[Van95] Jan Van Katwijk, editor. ACSCI ’95: 1st Annual conference — May 1995, Heijen, The Netherlands, Proceedings of the Annual Conference — Advanced School for Computing and Imaging, 1st. ASCI, Delft, The Netherlands, 1995. ISBN 90-90-08344-8. LCCN QA75.5 .A38x 1995.

vandeGeijn:1997:UPP

[van97] Robert A. van de Geijn. Using PLAPACK: Parallel Linear Algebra Package. MIT Press, Cambridge, MA, USA, 1997. ISBN 0-262-72026-4. xvii + 194 pp. LCCN QA185.D37 V36 1997. US$27.50. With contributions by Philip Alpatov and others.

Vlassov:1995:MEP

[VAT95] V. Vlassov, H. Ahmed, and L.-E. Thorelli. mEDA-2: An extension of PVM. In Malyshkin [Mal95], pages 288–293. ISBN 3-540-60222-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I547 1995.

Vazquez:1999:PNS

[VB99] G. E. Vazquez and N. B.

Brignole. Parallel NLP strategies using PVM on heterogeneous distributed environments. In Dongarra et al. [DLM99], pages 533– 540. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Villaverde:2018:PTI

[VBB18] Alejandro F. Villaverde, Kolja Becker, and Julio R. Banga. PREMER: a tool to infer biological networks. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 15(4):1193– 1202, July 2018. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

VanZee:2008:SPF

[VBLvdG08] Field G. Van Zee, Paolo Bientinesi, Tze Meng Low, and Robert A. van de Geijn. Scalable parallelization of FLAME code via the workqueuing model. ACM Transactions on Mathematical Software, 34(2):10:1– 10:29, March 2008. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Vapirev:2015:IRC

[VDL^{+}15] A. Vapirev, J. Deca, G. Lapenta, S. Markidis, I. Hur, and J.- L. Cambier. Initial results on computational performance of Intel many integrated core, Sandy Bridge, and graphical processing unit architectures: implementation of a 1D C++/OpenMP electrostatic particle-in-cell code. Concurrency and Computation: Practice and Experience, 27(3):581–593, March 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vanderLaan:2011:AWL

[vdLJR11] Wladimir J. van der Laan, Andrei C. Jalba, and Jos B. T. M. Roerdink. Accelerating wavelet lifting on graphics hardware using CUDA. IEEE Transactions on Parallel and Distributed Systems, 22(1):132– 146, January 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

vanderPas:2017:UON

[vdP17] Ruud van der Pas. Using OpenMP — the next step: affinity, accelerators, tasking, and SIMD. Scientific and engineering computation. MIT Press, Cambridge, MA, USA, 2017. ISBN 0-262-53478-9 (paperback). xxi + 365 pp. LCCN QA76.642 .P427 2017.

Vetter:2000:DST

[VdS00] Jeffrey S. Vetter and Bronis R. de Supinski. Dynamic software testing of MPI applications with Umpire. In ACM [ACM00],

page 70. URL http://www.
sc2000.org/proceedings/
techpapr/papers/pap208.
pdf.

Vetter:2002:DSP

[Vet02] Jeffrey Vetter. Dynamic statistical profiling of communication activity in distributed applications. ACM SIGMETRICS Performance Evaluation Review, 30(1): 240–250, June 2002. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vadhiyar:2002:PMS

[VFD02] Sathish S. Vadhiyar, Graham E. Fagg, and Jack J. Dongarra. Performance modeling for self adapting collective communications for MPI. In Oldehoeft [Old02], page ?? CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
netlib.org/utk/people/
JackDongarra/PAPERS/coll-
lacsi-2001.pdf.

Vitali:2019:EOO

[VGP^{+}19] Emanuele Vitali, Davide Gadioli, Gianluca Palermo, Andrea Beccari, Carlo Cavazzoni, and Cristina Silvano. Exploiting OpenMP and OpenACC to accelerate a geometric approach to molecular docking in heterogeneous HPC nodes. The Journal of Supercomputing, 75 (7):3374–3396, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Vega-Gisbert:2016:DIJ

[VGRS16] Oscar Vega-Gisbert, Jose E. Roman, and Jeffrey M. Squyres. Design and implementation of Java bindings in Open MPI. Parallel Computing, 59(??):1– 20, November 2016. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819116300758.

Vikas:2014:MGA

[VGS14] Vikas, Nasser Giacaman, and Oliver Sinnen. Multiprocessing with GUI-awareness using OpenMP-like directives in Java. Parallel Computing, 40(2):69–89, February 2014. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167819113001439.

vonHanxleden:1994:VDF

[vHKS94] R. von Hanxleden, K. Kennedy, and J. Saltz. Valuebased distributions in Fortran D. In Gentzsch and Harms [GH94], pages 434–440. ISBN 0-387-57981-8 (New York), 3-540-57981-8 (Berlin). LCCN QA76.88.I57 1994. DM96.00. Two volumes. Viswanathan:1995:PCM

[Vis95] Kishore Viswanathan. A parallel client-server model for distributed computing. M.s. thesis, Department of Computer Science, Mississippi State University, Starkville, MS, USA, 1995. vii + 79 pp.

Valero-Lara:2020:SFA

[VLCM^{+}20] Pedro Valero-Lara, Sandra Catalán, Xavier Martorell, Tetsuzo Usui, and Jesús Labarta. sLASs: a fully automatic auto-tuned linear algebra library based on OpenMP extensions implemented in OmpSs (LASs library). Journal of Parallel and Distributed Computing, 138(??):153–171, April 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731519303417.

Valero-Lara:2018:CCC

[VLMPS^{+}18] Pedro Valero-Lara, Ivan Martínez-Pérez, Raül Sirvent, Xavier Martorell, and Antonio J. Peña. cuThomas-Batch and cuThomasV-Batch, CUDA routines to compute batch of tridiagonal systems on NVIDIA GPUs. Concurrency and Computation: Practice and Experience, 30(24):e4909:1– e4909:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Valencia:2008:PPR

[VLO^{+}08] David Valencia, Alexey Lastovetsky, Maureen O’Flynn, Antonio Plaza, and Javier Plaza. Parallel processing of remotely sensed hyperspectral images on heterogeneous networks of workstations using HeteroMPI. The International Journal of High Performance Computing Applications, 22(4): 386–407, November 2008. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

http://hpc.sagepub.com/

content/22/4/386.full.
pdf+html.

Valero-Lara:2019:MTS

[VLSPL19] Pedro Valero-Lara, Raül Sirvent, Antonio J. Peña, and Jesús Labarta. MPI + OpenMP tasking scalability for multi-morphology simulations of the human brain. Parallel Computing, 84(??):50–61, May 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016781911830317X.

Varadarajan:1994:FDT

[VM94] V. Varadarajan and R. Mittra. Finite-difference timedomain (FDTD) analysis using distributed computing. IEEE Microwave and Guided Wave Letters, 4(5):144–145, September/ October 1994. CODEN IMGLE3. ISSN 1051-8207 (print), 1558-2329 (electronic).

Vincent:1995:HPP

[VM95] James J. Vincent and Kenneth M. Merz Jr. A highly portable parallel implementation of AMBER4 using the message passing interface standard. Journal of Computational Chemistry, 16(11):1420–1427, November 1995. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Vogel:2013:BWC

[Vog13] Thomas Vogel. All the Way to CUDA [book review]. Computing in Science and Engineering, 15 (5):6–8, September/October 2013. CODEN CSENFA. ISSN 1521-9615.

Volkert:1993:PCS

[Vol93] Jens Volkert, editor. Parallel computation: Second International ACPC Conference, Gmunden, Austria, October 4–6, 1993: proceedings, volume 734 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-57314-3 (Berlin), 0-387-57314-3 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.734. DM58.00.

Voss:2003:OSM

[Vos03] Michael J. Voss, editor. OpenMP shared memory parallel programming: International Workshop on OpenMP Applications and Tools, WOMPAT 2003, Toronto, Canada, June 26– 27, 2003: Proceedings, volume 2716 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNCSD9. ISBN 3-540-40435-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.642 .I589 2003. URL http:

//link.springer-ny.com/
link/service/series/0558/
tocs/t2716.htm;   http:
//www.springerlink.com/
openurl.asp?genre=issue&
issn=0302-9743&volume=
2716.

VidalMacia:2000:IPM

[VP00] Antonio Vidal Maciá and José Luis Pérez Gómez. Introducción a la programación en MPI. (Spanish) [Introduction to programming in MPI]. Technical report SPUPV-2000.209, Departamento de Sistemas Informáticos y Computación, Facultad de Informática, Universidad Politécnica de Valencia, Servicio de Publicaciones, Valencia, Spain, 2000. 78 pp.

Vargas-Perez:2017:HMO

[VPS17] Sandino Vargas-Perez and Fahad Saeed. A hybrid MPI–OpenMP strategy to speedup the compression of big next-generation sequencing datasets. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2760–2769, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

                                        trans/td/2017/10/07895161-
                                        abs.html.

Vrenios:2004:PPC

[Vre04] A. Vrenios. Parallel Programming in C with MPI and OpenMP [book review]. IEEE Distributed Systems Online, 5(1):7.1–7.3, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //ieeexplore.ieee.org/
                                        iel5/8968/28452/01270716.
                                        pdf?isnumber=28452&prod=
                                        JNL&arnumber=1270716&arSt=
                                        +7.1&ared=+7.3&arAuthor=
                                        Vrenios%2C+A.;   http:
                                        //ieeexplore.ieee.org/
                                        xpls/abs_all.jsp?isnumber=
                                        28452&arnumber=1270716&
                                        count=8&index=5.

Varin:2000:PAL

[VRS00] E. Varin, R. Roy, and G. Samba. Parallel algorithms for the least-squares finite element solution of the neutron transport equation. Lecture Notes in Computer Science, 1908:121–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1908/19080121.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1908/19080121.
                                        pdf.

VanVoorst:2000:CMI

[VS00] Brian Van Voorst and Steven Seidel. Comparison of MPI implementations on a shared memory machine. Lecture Notes in Computer Science, 1800:847–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/1800/18000847.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/1800/18000847.
pdf.

Vaughan:1994:MPM

[VSRC94] P. L. Vaughan, A. Skjellum, D. S. Reese, and Fei-Chen Cheng. Migrating from PVM to MPI. I. the Unify system. In IEEE [IEE94a], pages 488–495. ISBN 0-8186-6965-9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

Vaughan:1995:MPM

[VSRC95] Paula L. Vaughan, Anthony Skjellum, Donna S. Reese, and Fei-Chen Cheng. Migrating from PVM to MPI, part I: The Unify system. Frontiers of Massively Parallel Computation — Conference Proceedings, pages 488– 495, ???? 1995. IEEE catalog number 95TH8024.

Vaidya:2013:SDO

[VSW^{+}13] Aniruddha S. Vaidya, Anahita Shayesteh, Dong Hyuk Woo, Roy Saharoy, and Mani Azimi. SIMD divergence optimization through intrawarp compaction. ACM SIGARCH Computer Architecture News, 41(3):368– 379, June 2013. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ICSA ’13 conference proceedings.

Vlassov:1997:SSM

[VT97] V. Vlassov and L.-E. Thorelli. A synchronizing shared memory: Model and programming implementation. Lecture Notes in Computer Science, 1332:159–166, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vu:2019:FMT

[VT19] V. A. Vu and G. Tan. A framework for mesoscopic traffic simulation in GPU. IEEE Transactions on Parallel and Distributed Systems, 30(8):1691–1703, August 2019. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Vandoni:1995:CSC

[VV95] C. E. Vandoni and C. Verkerk, editors. 1994 CERN School of Computing: Sopron, Hungary, 28 August– 10 September 1994: proceedings. CERN, Geneva, Switzerland, 1995. ISBN 92-9083-069-7. CERN report 95-01.

Vo:2009:FVP

[VVD^{+}09] Anh Vo, Sarvani Vakkalanka, Michael DeLisi, Ganesh Gopalakrishnan, Robert M. Kirby, and Rajeev Thakur. Formal verification of practical MPI programs. ACM SIGPLAN Notices, 44(4): 261–270, April 2009. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Verkerk:1992:PIC

[VW92] C. Verkerk and W. Wojcik, editors. Proceedings of the International Conference on Computing in High Energy Physics ’92, Annecy, France, 21–25 September 1992. CERN, Geneve, Switzerland, 1992. ISBN 92-9083-049-2. LCCN QC783.3 C65 1992. CERN report 92-07.

Vetter:2002:EPE

[VY02] Jeffrey S. Vetter and Andy Yoo. An empirical performance evaluation of scalable scientific applications. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN

???? URL http://www.sc-
2002.org/paperpdfs/pap.
pap222.pdf.

Verschelde:2015:PHC

[VY15] Jan Verschelde and Xiangcheng Yu. Polynomial homotopy continuation on GPUs. ACM Communications in Computer Algebra, 49(4):130–133, December 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Vasilache:2019:NAL

[VZT^{+}19] Nicolas Vasilache, Oleksandr Zinenko, Theodoros Theodoridis, Priya Goyal, Zachary Devito, William S. Moses, Sven Verdoolaege, Andrew Adams, and Albert Cohen. The next 700 accelerated layers: From mathematical expressions of network computation graphs to accelerated GPU kernels, automatically. ACM Transactions on Architecture and Code Optimization, 16(4):38:1–38:??, October 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Wong:1999:BMM

[WADC99] F. C. Wong, A. C. Arpaci-Dusseau, and D. E. Culler. Building MPI for multiprogramming systems using implicit information. In Dongarra et al. [DLM99], pages 215–222. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Walker:1994:DSM

[Wal94a] David W. Walker. The design of a standard message passing interface for distributed memory concurrent computers. Parallel Computing, 20(4):657– 673, March 31, 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/parco/
cas_sub/browse/browse.
cgi?year=1994&volume=20&
issue=4&aid=865;   http:
//www.epm.ornl.gov/~walker/
mpi/papers/parcomp94.ps.

Z. See erratum [Wal94b].

Walker:1994:EDS

[Wal94b] David W. Walker. Erratum to: “The design of a standard message passing interface for distributed memory concurrent computers”. Parallel Computing, 20(8):1215, August 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). See [Wal94a].

Walker:1995:MVB

[Wal95] D. W. Walker. An MPI version of the BLACS. In IEEE [IEE95j], pages 129– 146. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Walker:1996:MFA

[Wal96a] David W. Walker. MPI:

from fundamentals to applications. Technical report, Oak Ridge National Laboratory, Knoxville, TN, USA,

1996. URL http://www.

                                        epm.ornl.gov/~walker/mpi/
                                        SLIDES/mpi-tutorial.html.

Walker:1996:MP

[Wal96b] David W. Walker. MPI2 proposals. World-Wide Web,

                                        1996. URL http://www.
                                        epm.ornl.gov/~walker/mpi/
                                        mpi2-proposals.html.

Wallcraft:2000:SOV

[Wal00] Alan J. Wallcraft. SPMD OpenMP versus MPI for ocean models. Concurrency: practice and experience, 12(12):1155–1164, October 2000. CODEN CPEXEI. ISSN 1040-

                                        3108. URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/76500353/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=76500353&PLACEBO=IE.
                                        pdf.

Walker:2001:DLB

[Wal01a] Reginald L. Walker. Dynamic load balancing model: Preliminary results for parallel pseudo-search engine indexers/crawler mechanisms using MPI and genetic programming. Lecture Notes in Computer Science, 1981: 61–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1981/19810061.htm;
                                         http://link.springer-

ny.com/link/service/series/
0558/papers/1981/19810061.
pdf.

Walker:2001:SEC

[Wal01b] Reginald L. Walker. Search engine case study: searching the Web using genetic programming and MPI. Parallel Computing, 27(1–2): 71–89, January 2001. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/35/21/47/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/35/21/47/25/25/article.
pdf.

Wallcraft:2002:CCA

[Wal02] Alan J. Wallcraft. A comparison of Co-Array Fortran and OpenMP Fortran for SPMD programming. The Journal of Supercomputing, 22(3):231–250, July 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://

ipsapp008.kluweronline.
com/content/getfile/5189/
36/1/abstract.htm;   http:
//ipsapp008.kluweronline.
com/content/getfile/5189/
36/1/fulltext.pdf.

Wang:1997:TPD

[Wan97] Paul S. Wang. Tools for parallel/distributed mathematical computation. In ACM [ACM97a], pages 188–195. ISBN ???? LCCN ????

Wang:2002:OPG

[Wan02] Ping Wang. OpenMP programming for a global inverse model. Scientific Programming, 10(3):253–261, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wasniowski:1995:NAP

[Was95a] R. A. Wasniowski. Nonlinear adaptive prediction algorithm and its parallel implementation. Informatica (Ljubljana, Slovenia), 19(3): 371–377, September 1995. CODEN INFOFF. ISSN 0350-5596.

White:1995:PNP

[WAS95b] S. White, A. Alund, and V. S. Sunderam. Performance of the NAS parallel benchmarks on PVM-Based networks. Journal of Parallel and Distributed Computing, 26(1):61–71, April 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

                                        //www.idealibrary.com/
                                        links/doi/10.1006/jpdc.
                                        1995.1048/production;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/
                                        jpdc.1995.1048/production/
                                        pdf.

Wasniewski:1996:APC

[Was96] Jerzy Wasniewski, editor. Applied parallel computing: industrial computation and optimization: Third International Workshop, PARA ’96, Lyngby, Denmark, August 18–21, 1996: proceedings, volume 1184 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-62095-8. LCCN QA76.58 .P35 1996.

Wolf:1996:CFS

[WB96] K. Wolf and E. Brakkee.

Coupling fluids and structures codes on MPI. In IEEE [IEE96i], pages 130– 137. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Wickerson:2015:RSP

[WBBD15] John Wickerson, Mark Batty, Bradford M. Beckmann, and Alastair F. Donaldson. Remote-scope promotion: clarified, rectified, and verified. ACM SIG-PLAN Notices, 50(10):731– 747, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wolf:1997:CMP

[WBH97] K. Wolf, E. Brakkee, and D. P. Ho. Communication in multi-physics applications. Lecture Notes in Computer Science, 1332:167–176, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wickerson:2017:ACM

[WBSC17] John Wickerson, Mark Batty, Tyler Sorensen, and George A. Constantinides. Automatically comparing memory consistency models. ACM SIGPLAN Notices, 52(1):190–204, January 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Walters:2009:RBF

[WC09] John Paul Walters and Vipin Chaudhary. Replicationbased fault tolerance for MPI applications. IEEE Transactions on Parallel and Distributed Systems, 20(7):997–1010, July 2009. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2015:AST

[WC15] Chun-Kun Wang and Peng-Sheng Chen. Automatic scoping of task clauses for the OpenMP tasking model. The Journal of Supercomputing, 71(3):808– 823, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s11227-014-1326-3.

Wang:2007:EAP

[WCC^{+}07] Perry H. Wang, Jamison D. Collins, Gautham N.

Chinya, Hong Jiang, Xinmin Tian, Milind Girkar, Nick Y. Yang, Guei-Yuan Lueh, and Hong Wang. EXOCHI: architecture and programming environment for a heterogeneous multi-core multithreaded system. ACM SIG-PLAN Notices, 42(6):156– 166, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2012:OVT

[WCC12] Cheng Wang, Sunita Chandrasekaran, and Barbara Chapman. An OpenMP 3.1 validation testsuite. Lecture Notes in Computer Science, 7312:237– 249, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-30961-8_
18/.

Wu:1999:JBD

[WCS99] X. Wu, Q. Chen, and X.- H. Sun. A Java-based distributed debbuger supporting MPI and PVM. Parallel and Distributed Computing Practices, 2(4): ??, ???? 1999. CO-DEN ???? ISSN 1097-

2803. URL http://www.cs.
okstate.edu/~pdcp/vols/
vol02/vol02no4abs.html#
wu.

Wang:2013:PMO

[WCS^{+}13] Cheng Wang, Sunita Chandrasekaran, Peng Sun, Barbara Chapman, and Jim Holt. Portable mapping of openMP to multicore embedded systems using MCA APIs. ACM SIG-PLAN Notices, 48(5):153– 162, May 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wedemeijer:1996:PSA

[WCVR96] H. Wedemeijer, H. L. H. Cox, D. J. Verschuur, and I. L. Ritsema. Parallelisation of seismic algorithms using PVM and FORGE. In Liddell et al. [LCHS96], pages 352–?? ISBN 3-540-61142-8 (paperback). LCCN QA76.88 .H52 1996.

Walker:1996:MSM

[WD96] D. W. Walker and J. J. Dongarra. MPI: a standard message passing interface. Supercomputer, 12(1): 56–68, January 1996. CO-DEN SPCOEL. ISSN 0168-7875.

Wozniak:2019:MJW

[WDR^{+}19] Justin M. Wozniak, Matthieu Dorier, Robert Ross, Tong Shu, Tahsin Kurc, Li Tang, Norbert Podhorszki, and Matthew Wolf. MPI jobs within MPI jobs: a practical way of enabling tasklevel fault-tolerance in HPC workflows. Future Generation Computer Systems, 101(??):576–589, December 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1830757X.

Welch:1994:PVM

[Wel94] L. R. Welch. A parallel virtual machine for programs composed of abstract data types. IEEE Transactions on Computers, 43(11): 1249–1261, November 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Werner:1995:UMP

[Wer95] Jörg Werner. Überblick zum Message-Passing-Interface Standard, MPI. (German) [Overview of the Message-Passing Interface Standard, MPI]. Parlab-Mitteilungen 04/95, Technische Universität Chemnitz-Zwickau, Chemnitz, Germany, 1995. 35 pp.

Weber:2017:MAL

[WG17] Nicolas Weber and Michael Goesele. MATOG: Array layout auto-tuning for CUDA. ACM Transactions on Architecture and Code Optimization, 14(3): 28:1–28:??, September 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Warren:2019:CBG

[WGG^{+}19] Craig Warren, Antonios Giannopoulos, Alan Gray, Iraklis Giannakis, Alan Patterson, Laura Wetter, and Andre Hamrah. A CUDA-based GPU engine for gprMax: Open source FDTD electromagnetic simulation software. Computer Physics Communications, 237(??):208–218, April 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465518303990.

Wark:1994:PIR

[WH94] P. Wark and J. Holt. PVM implementation of a repeated matching heuristic for vehicle routing. In Arnold et al. [ACDR94], pages 207–216 (or 207– 214??). ISBN 90-5199-149-5. LCCN ????

Wagner:1996:PMM

[WH96] J. C. Wagner and A. Haghighat. Parallel MCNP Monte Carlo transport calculations with MPI. Transactions of the American Nuclear Society, 75(??):338–339, ???? 1996. CODEN TANSAO. ISSN 0003-018X.

Wiese:2005:IPN

[WHDB05] Kay C. Wiese, Andrew Hendriks, Alain Deschenes, and Belgacem Ben Youssef. The impact of pseudorandom number quality on P-RnaPredict, a parallel genetic algorithm for RNA secondary structure prediction. In Beyer et al. [B^{+}05], pages 479–480. ISBN 1-59593-010-8 (paperback). LCCN QA76.623 .G44 2005. URL http://

www.cs.bham.ac.uk/~wbl/
biblio/gecco2005lbp/papers/
52-wiese.pdf. ACM order

number 910050.

White:1994:VVC

[Whi94] R. White. VCMON — the VM/ESA Connectivity Monitor. In Anonymous [Ano94g], pages 783–792. ISBN ???? LCCN ????

White:2004:CMM

[Whi04] R. E. (Robert E.) White.

Computational Mathematics: Models, Methods, and Analysis with MATLAB and MPI. Chapman and Hall/CRC, Boca Raton, FL, USA, 2004. ISBN 1-58488-364-2. xvi + 385 pp. LCCN QA297 .W495 2004.

Waidyasooriya:2019:OBD

[WHMO19] Hasitha Muthumala Waidyasooriya, Masanori Hariyama, Masamichi J. Miyama, and Masayuki Ohzeki. OpenCLbased design of an FPGA accelerator for quantum annealing simulation. The Journal of Supercomputing, 75(8):5019–5039, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wilkinson:1993:IFT

[Wil93] Timothy James Wilkinson. Implementing Fault Tolerance in a 64-bit Distributed Operating System. PhD thesis, Systems Architecture Research Centre, City University, London, UK, July 1993.

Wilhelms:1994:DAL

[Wil94] Gerhard Wilhelms. Dynamische adaptive Lastverteilung für PVM mittels unscharfer Benutzerprofile – PVM^{+} (English: Dynamic adaptive load distribution for PVM by blurred user profiles – PVM^{+} ). Dissertation, Math.- Naturwiss. Fakultät, Universität Augsburg, Augsburg, Germany, 1994. iv + 74 pp.

Wismueller:1996:SBV

[Wis96a] R. Wismueller. State based visualization of PVM applications. Lecture Notes in Computer Science, 1156:91– ??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wismuller:1996:SBV

[Wis96b] R. Wismuller. State based visualization of PVM applications. In Bode et al. [BDLS96]. ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Wismueller:1997:DMP

[Wis97] R. Wismueller. Debugging message passing programs using invisible message tags. Lecture Notes in Computer Science, 1332:295–304, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wismueller:1998:LMS

[Wis98]  R. Wismueller.  On-line

monitoring support in PVM and MPI. Lecture Notes in Computer Science, 1497: 312–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wismuller:2001:UMT

[Wis01]  Roland Wismüller.  Us-

ing monitoring techniques to support the cooperation of software components. Lecture Notes in Computer Science, 2131:183–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310183.htm;
http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310183.
pdf.

Witchel:2016:PPW

[Wit16] Emmett Witchel. Programmer productivity in a world of mushy interfaces: Challenges of the post-ISA reality. Operating Systems Review, 50(2):591, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wei:2012:OLL

[WJ12] Zheng Wei and Joseph Jaja. Optimization of linked list prefix computations on multithreaded GPUs using CUDA. Parallel Processing Letters, 22(4):1250012, December 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Wang:2019:MEM

[WJA^{+}19] L. Wang, M. Jahre, A. Adileh, Z. Wang, and L. Eeckhout. Modeling emerging memory-divergent GPU applications. IEEE Computer Architecture Letters, 18(2): 95–98, July 2019. ISSN 1556-6056 (print), 1556-6064 (electronic).

Wu:2014:OFB

[WJB14] Jing Wu, Joseph JaJa, and Elias Balaras. An optimized FFT-based direct Poisson solver on CUDA GPUs. IEEE Transactions on Parallel and Distributed Systems, 25(3):550– 559, March 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2021:PBS

[WJG^{+}21] Y. Wang, X. Jiang, N. Guan, Z. Guo, X. Liu, and W. Yi. Partitioning-based scheduling of OpenMP task systems with tied tasks. IEEE Transactions on Parallel and Distributed Systems, 32(6): 1322–1339, 2021. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wegiel:2008:MCVa

[WK08a] Michal Wegiel and Chandra Krintz. The mapping collector: virtual memory support for generational, parallel, and concurrent compaction. ACM SIGARCH Computer Architecture News, 36(1):91– 102, March 2008. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Wegiel:2008:MCVb

[WK08b] Michal Wegiel and Chandra Krintz. The Mapping Collector: virtual memory support for generational, parallel, and concurrent compaction. Operating Systems Review, 42(2):91–102, March 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wegiel:2008:MCVc

[WK08c] Michal Wegiel and Chandra Krintz. The mapping collector: virtual memory support for generational, parallel, and concurrent compaction. ACM SIGPLAN Notices, 43(3):91–102, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

White:2020:OPP

[WK20] Sam White and Laxmikant V. Kale. Optimizing pointto-point communication between adaptive MPI endpoints in shared memory. Concurrency and Computation: Practice and Experience, 32(3):e4467:1– e4467:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wittenbrink:2011:FGG

[WKP11] Craig M. Wittenbrink, Emmett Kilgariff, and Arjun Prabhu. Fermi GF100 GPU architecture. IEEE Micro, 31(2):50–59, March/April 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wagner:1996:GSG

[WKS96] T. Wagner, C. Kueblbeck, and C. Schittko. Genetic selection and generation of textural features with PVM. In Bode et al. [BDLS96], pages 305–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Lehman:1994:IZP

[wL94] Li wei Lehman. Integrating Zipcode and PVM: towards a higher-level messagepassing environment. Technical report MSSU-EIRS-ERC 94-2, Engineering Research Center for Computational Field Simulation, Mississippi State University, Starkville, MS, USA, 1994. 7 pp.

Wismueller:1996:TSI

[WL96a] R. Wismueller and T. Ludwig. The tool-set — an integrated tool environment for PVM. Lecture Notes in Computer Science, ?? (1067):1029–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wismuller:1996:TSI

[WL96b] R. Wismuller and T. Ludwig. The Tool Set — an integrated tool environment for PVM. In Liddell et al. [LCHS96]. ISBN 3-540-61142-8 (paperback). LCCN QA76.88 .H52 1996.

Wu:2007:IFR

[WLC07] C.-L. Wu, D.-C. Lou, and S.-Y. Chen. Integer factorization for RSA cryptosystem under a PVM environment. International Journal of Computer Systems Science and Engineering, 22(1– 2):??, January/March 2007. CODEN CSSEEI. ISSN 0267-6192.

Wolfe:2018:ODM

[WLK^{+}18] Michael Wolfe, Seyong Lee, Jungwon Kim, Xiaonan Tian, Rengan Xu, Barbara Chapman, and Sunita Chandrasekaran. The OpenACC data model: Preliminary study on its major challenges and implementations. Parallel Computing, 78(??):15–27, October 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819118302175.

Weatherly:2003:DMS

[WLNL03] D. Brent Weatherly, David K. Lowenthal, Mario Nakazawa, and Franklin Lowenthal. Dyn-MPI: Supporting MPI on non dedicated clusters. In ACM [ACM03], page ?? ISBN 1-58113-695-

1. LCCN ???? URL http:/

/www.sc-conference.org/
sc2003/inter_cal/inter_
cal_detail.php?eventid=
10708#1;   http://www.
sc-conference.org/sc2003/
paperpdfs/pap126.pdf.

Weatherly:2006:DMS

[WLNL06] D. Brent Weatherly, David K.

Lowenthal, Mario Nakazawa, and Franklin Lowenthal. Dyn-MPI: Supporting MPI on medium-scale, nondedicated clusters. Journal of Parallel and Distributed Computing, 66(6): 822–838, June 2006. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Willcock:2005:UMC

[WLR05] Jeremiah Willcock, Andrew Lumsdaine, and Arch Robison. Using MPI with C# and the Common Language Infrastructure. Concurrency and Computation: Practice and Experience, 17(7– 8):895–917, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2012:UHM

[WLYC12] Chao-Chin Wu, Lien-Fu Lai, Chao-Tung Yang, and Po-Hsun Chiu. Using hybrid MPI and OpenMP programming to optimize communications in parallel loop selfscheduling schemes for multicore PC clusters. The Journal of Supercomputing, 60(1):31–61, April 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
60&issue=1&spage=31.

Weng:2020:CMS

[WLYL20] Tien-Hsiung Weng, Kuan-Ching Li, Zhiliu Yang, and Chen Liu. On the code modernization of shared sampling alpha matting with OpenMP. Future Generation Computer Systems, 107(??): 177–191, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167739X19314116.

Wolf:2001:APA

[WM01] Felix Wolf and Bernd Mohr. Automatic performance analysis of MPI applications based on event traces. Lecture Notes in Computer Science, 1900: 123–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/1900/19000123.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1900/19000123.
                                        pdf.

Wolfe:2018:MLS

[WMC^{+}18] Noah Wolfe, Misbah Mubarak, Christopher D. Carothers, Robert B. Ross, and Philip H. Carns. Modeling largescale slim fly networks using parallel discrete-event simulation. ACM Transactions on Modeling and Computer Simulation, 28(4): 29:1–29:??, October 2018. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wende:2019:OVT

[WMK^{+}19] Florian Wende, Martijn Marsman, Jeongnim Kim, Fedor Vasilev, Zhengji Zhao, and Thomas Steinke. OpenMP in VASP: Threading and SIMD. International Journal of Quantum Chemistry, 119(12):e25851:1–e25851:??, June 15, 2019. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wu:2014:MAG

[WMP14] Xing Wu, Frank Mueller, and Scott Pakin. A methodology for automatic generation of executable communication specifications from parallel MPI applications. ACM Transactions on Parallel Computing (TOPC), 1(1):6:1–6:??, September 2014. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Winkler:2017:GSM

[WMRR17] Daniel Winkler, Michael Meister, Massoud Rezavand, and Wolfgang Rauch. gpuSPHASE — a shared memory caching implementation for 2D SPH using CUDA. Computer Physics Communications, 213(??):

165–180, April 2017. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465516303666.

Wendykier:2010:PCH

[WN10] Piotr Wendykier and James G. Nagy. Parallel Colt: a high-performance Java library for scientific computing and image processing. ACM Transactions on Mathematical Software, 37(3): 31:1–31:22, September 2010. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Walker:1995:RBD

[WO95] David W. Walker and Steve W. Otto. Redistribution of block-cyclic data distributions using MPI. Technical Report ORNL/TM-12999, Oak Ridge National Laboratory, Knoxville, TN, USA, June 1995. iii + 20

                                        pp. URL http://www.epm.
                                        ornl.gov/~walker/mpi/redistribution.
                                        ps.Z.

Walker:1996:RBC

[WO96] D. W. Walker and S. W. Otto. Redistribution of block-cyclic data distributions using MPI. Concurrency: practice and experience, 8(9):707–728, November 1996. CODEN CPEXEI. ISSN 1040-

                                        3108. URL http://www3.

interscience.wiley.com/
cgi-bin/abstract?ID=23305.

Winstanley:1997:PDP

[WO97] N. Winstanley and J. O’Donnell.

Parallel distributed programming with Haskell+PVM. Lecture Notes in Computer Science, 1300:670–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:2009:MPM

[WO09] Zheng Wang and Michael F. P. O’Boyle. Mapping parallelism to multicores: a machine learning based approach. ACM SIG-PLAN Notices, 44(4):75– 84, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wolbers:1992:SPP

[Wol92] S. Wolbers. Software for parallel processing applications. In Verkerk and Wojcik [VW92], pages 111–116. ISBN 92-9083-049-2. LCCN QC783.3 C65 1992. CERN report 92-07.

Worley:1996:MPE

[Wor96] P. H. Worley. MPI performance evaluation and characterization using a compact application benchmark code. In IEEE [IEE96i], pages 170–177. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Weng:2007:OIS

[WPC07] Tien-Hsiung Weng, Ruey-Kuen Perng, and Barbara Chapman. OpenMP implementation of SPICE3 circuit simulator. International Journal of Parallel Programming, 35(5):493– 505, October 2007. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=5&spage=493.

Wagner:1994:CFD

[WPH94] S. (Siegfried) Wagner, J. (Jacques) Periaux, and E. H. (Ernst-Heinrich) Hirschel, editors. Computational fluid dynamics ’94: proceedings of the Second European Computational Fluid Dynamics Conference, 5–8 September 1994, Stuttgart, Germany. Wiley, New York, NY, USA, 1994. ISBN 0-471-95063-7. LCCN QA911.E95 1994.

Wang:1995:PPG

[WPL95] Cho-Li Wang, V. K. Prasanna, and Young Won Lim. Parallelization of perceptual grouping on distributed memory machines. In Cantoni et al. [CLM^{+}95], pages 323–330. ISBN 0-8186-7134-3. LCCN QA76.9.A73W675 1995. IEEE catalog no. 95TB8093.

Wang:2020:EPE

[WQKH20] X. Wang, X. Qian, A. Knoll, and K. Huang. Efficient performance estimation and work-group size pruning for OpenCL kernels on GPUs. IEEE Transactions on Parallel and Distributed Systems, 31(5):1089–1106, May 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2001:PCS

[WR01] Guang Jun Wu and Robert Roy. Parallelization of characteristics solvers for 3D neutron transport. Lecture Notes in Computer Science, 2131:344–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2131/21310344.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310344.
pdf.

Worsch:2002:BCM

[WRA02] Thomas Worsch, Ralf Reussner, and Werner Augustin. On benchmarking collective MPI operations. Lecture Notes in Computer Science, 2474:271–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/
service/series/0558/bibs/

                                        2474/24740271.htm;   http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2474/24740271.pdf.

Winkler:2019:GSM

[WRMR19] Daniel Winkler, Massoud Rezavand, Michael Meister, and Wolfgang Rauch. gpuS-PHASE — a shared memory caching implementation for 2D SPH using CUDA (new version announcement). Computer Physics Communications, 235(??): 514–516, February 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465518303126.

Wang:2016:LLA

[WRSY16] Jin Wang, Norm Rubin, Albert Sidelnik, and Sudhakar Yalamanchili. LaPerm: locality aware scheduler for dynamic parallelism on GPUs. ACM SIGARCH Computer Architecture News, 44(3):583–595, June 2016. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Wisniewski:1999:SME

[WSN99] Len Wisniewski, Brad Smisloff, and Nils Nieuwejaar. Sun MPI I/O: Efficient I/O for parallel applications. In ACM [ACM99], page ??

West:1995:AVV

[WST95] J. E. West, M. M. Stephens, and L. H. Turcotte. Adaptation of volume visualization techniques to MIMD architectures using MPI. In IEEE [IEE95j], pages 147– 156. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Wu:2011:PCH

[WT11] Xingfu Wu and Valerie Taylor. Performance characteristics of hybrid MPI/ OpenMP implementations of NAS parallel benchmarks SP and BT on large-scale multicore supercomputers. ACM SIGMETRICS Performance Evaluation Review, 38(4):56–62, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wu:2012:PCH

[WT12] Xingfu Wu and Valerie Taylor. Performance characteristics of hybrid MPI/ OpenMP implementations of NAS Parallel Benchmarks SP and BT on large-scale multicore clusters. The Computer Journal, 55(2): 154–167, February 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/2/154.full.
pdf+html.

Wu:2013:PMH

[WT13] Xingfu Wu and Valerie Taylor. Performance modeling of hybrid MPI/OpenMP scientific applications on largescale multicore supercomputers. Journal of Computer and System Sciences, 79(8): 1256–1268, December 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0022000013000639.

Wang:2014:IPD

[WTFO14] Zheng Wang, Georgios Tournavitis, Björn Franke, and Michael F. P. O’boyle. Integrating profile-driven parallelism detection and machine-learning-based mapping. ACM Transactions on Architecture and Code Optimization, 11(1):2:1–2:??, February 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Worringen:2003:FPN

[WTR03] Joachim Worringen, Jesper Larson Traff, and Hubert Ritzdorf. Fast parallel non-contiguous file access. In ACM [ACM03], page ?? ISBN 1-58113-695-

1. LCCN ???? URL http:/

/www.sc-conference.org/
sc2003/inter_cal/inter_
cal_detail.php?eventid=
10722#0;   http://www.

sc-conference.org/sc2003/
paperpdfs/pap319.pdf.

Wang:2019:FBA

[WTS19] Haomiao Wang, Prabu Thiagaraj, and Oliver Sinnen. FPGA-based acceleration of FT convolution for pulsar search using OpenCL. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(4):24:1– 24:??, January 2019. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3268933.

Waidyasooriya:2017:OBF

[WTTH17] Hasitha Muthumala Waidyasooriya, Yasuhiro Takei, Shunsuke Tatsumi, and Masanori Hariyama. OpenCLbased FPGA-platform for stencil computation and its optimization methodology. IEEE Transactions on Parallel and Distributed Systems, 28(5):1390–1402, May 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/05/07582502-
abs.html.

Wu:1999:MCC

[Wu99] P.-Y. Wu. Minimum communication cost fractal image compression on PVM. In Dongarra et al. [DLM99], pages 434–441. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Wong:2011:EMS

[WWFT11] Hon-Cheng Wong, Un-Hong Wong, Xueshang Feng, and Zesheng Tang. Efficient magnetohydrodynamic simulations on graphics processing units with CUDA. Computer Physics Communications, 182(10):2132– 2160, October 2011. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465511001676.

Wilson:1996:SMS

[WWZ^{+}96] G. C. Wilson, T. H. Wood, J. L. Zyskind, J. W. Sulhoff, J. E. Johnson, T. Tanbun-Ek, and P. A. Morton. SBS and MPI suppression in analogue systems with integrated electroabsorption modulator/DFB laser transmitters. Electronics Letters, 32(16):1502–1504, ???? 1996. CODEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Wang:2021:PBD

[WYH^{+}21] Shao-Chung Wang, Lin-Ya Yu, Li-An Her, Yuan-Shin Hwang, and Jenq-Kuen Lee. Pointer-based divergence analysis for OpenCL 2.0 programs. ACM Transactions on Parallel Computing (TOPC), 8(4):20:1– 20:23, December 2021. CO-DEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic). URL https://

dl.acm.org/doi/10.1145/
3470644.

Wu:2012:DPL

[WYLC12] Chao-Chin Wu, Chao-Tung Yang, Kuan-Chou Lai, and Po-Hsun Chiu. Designing parallel loop self-scheduling schemes using the hybrid MPI and OpenMP programming model for multi-core grid systems. The Journal of Supercomputing, 59 (1):42–60, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
59&issue=1&spage=42.

Wu:2019:PMG

[WYZ^{+}19] J. Wu, X. Yang, Z. Zhang, G. Chen, and R. Mao. A performance model for GPU architectures that considers on-chip resources: Application to medical image registration. IEEE Transactions on Parallel and Distributed Systems, 30(9):1947–1961, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2016:MMF

[WZHZ16] Zeke Wang, Shuhao Zhang, Bingsheng He, and Wei Zhang. Melia: A MapReduce framework on OpenCLbased FPGAs. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3547–3560, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/12/07425227-
                                        abs.html.

Wang:2017:CEG

[WZM17] Siqi Wang, Guanwen Zhong, and Tulika Mitra. CGPredict: Embedded GPU performance estimation from single-threaded applications. ACM Transactions on Embedded Computing Systems, 16(5s):146:1–146:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2021:ATD

[WZW21] Farui Wang, Weizhe Zhang, and Zheng Wang. Automatic translation of data parallel programs for heterogeneous parallelism through OpenMP offloading. The Journal of Supercomputing, 77(5):4957–4987, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

                                        https://link.springer.

com/article/10.1007/s11227-
020-03452-2.

Wang:2008:PIM

[WZWS08] Kun Wang, Yu Zhang, Huayong Wang, and Xiaowei Shen. Parallelization of IBM Mambo system simulator in functional modes. Operating Systems Review, 42(1):71– 76, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Xu:1995:IPP

[XF95] H. Xu and T. W. Fisher. Improving PVM performance using ATOMIC user-level protocol. In Alnuweiri and Hamdi [AH95], pages 108– 117. ISBN 0-8186-7124-6. LCCN TK5105.5 .H56 1995.

Xu:1996:MCO

[XH96] Zhiwei Xu and Kai Hwang.

Modeling communication overhead: MPI and MPL performance on the IBM SP2. IEEE parallel and distributed technology: systems and applications, 4(1): 9–24, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Xue:2021:IFG

[XJR21] Weicheng Xue, Charles W.

Jackson, and Christoper J. Roy. An improved framework of GPU computing for CFD applications on structured grids using OpenACC.

Journal of Parallel and Distributed Computing, 156 (??):64–85, October 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731521001155.

Xue:2009:MSR

[XLW^{+}09] Ruini Xue, Xuezheng Liu, Ming Wu, Zhenyu Guo, Wenguang Chen, Weimin Zheng, Zheng Zhang, and Geoffrey Voelker. MPIWiz: subgroup reproducible replay of MPI applications. ACM SIGPLAN Notices, 44(4):251–260, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xue:2021:MGP

[XR21] Weicheng Xue and Christoper J. Roy. Multi-GPU performance optimization of a computational fluid dynamics code using OpenACC. Concurrency and Computation: Practice and Experience, 33(5):e6036:1– e6036:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiong:1996:BID

[XWZS96] Jianxin Xiong, Dingxing Wang, Weimin Zheng, and Meiming Shen. BUSTER: an integrated debugger for PVM. In IEEE [IEE96d]. ISBN 0-7803-3529-5 (softbound), 0-7803-3530-9 (microfiche). LCCN QA76.58.I33 1996. IEEE catalog number 96TH8204.

Xu:2013:PMO

[XXL13] Shiming Xu, Wei Xue, and Hai Xiang Lin. Performance modeling and optimization of sparse matrixvector multiplication on NVIDIA CUDA platform. The Journal of Supercomputing, 63(3):710–721, March 2013. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11227-011-0626-0;
 http://link.springer.
com/content/pdf/10.1007/
s11227-011-0626-0.

Yelon:1993:PTS

[Y^{+}93] W. B. Yelon et al., editors.

Proceedings of the Thirtyseventh Annual Conference on Magnetism and Magnetic Materials: December 1–4, 1992, Houston, Texas, volume 73(10) of Journal of Applied Physics. American Institute of Physics, Woodbury, NY, USA, May 1993. CODEN JAPIAU. ISBN 1-56396-212-8. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. LCCN QC753 .C748 1990. Two volumes.

Yazdanpanah:2015:PHR

[YÁJG^{+}15] Fahimeh Yazdanpanah, Carlos Álvarez, Daniel Jiménez-González, Rosa M. Badia, and Mateo Valero. Picos: a hardware runtime architecture support for OmpSs. Future Generation Computer Systems, 53(??):130– 139, December 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167739X14002702.

Yan:1994:PTA

[Yan94] J. C. Yan. Performance tuning with AIMS — an Automated Instrumentation and Monitoring System for multicomputers. In Hesham and Shriver [HS94], pages 625–

633. ISBN 0-8186-5060-5. ISSN 1060-3425. LCCN ???? IEEE catalog no. 94TH0607-2.

Yang:2014:PMI

[YBMCB14] Chaoran Yang, Wesley Bland, John Mellor-Crummey, and Pavan Balaji. Portable, MPI-interoperable Coarray Fortran. ACM SIGPLAN Notices, 49(8):81–92, August 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ying:2003:NPK

[YBZL03] Lexing Ying, George Biros, Denis Zorin, and Harper Langston. A new parallel kernel-independent fast multipole method. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://
www.sc-conference.org/
sc2003/inter_cal/inter_
cal_detail.php?eventid=
10707#2;   http://www.
sc-conference.org/sc2003/
paperpdfs/pap166.pdf.

Yalamanchilli:1998:CPJ

[YC98] Narendar Yalamanchilli and William Cohen. Communication performance of Java based Parallel Virtual Machines. In ACM [ACM98a], page ?? ISBN ???? LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/
papers/passing.pdf;   http:
//www.cs.ucsb.edu/conferences/
java98/papers/passing.

ps. Possibly unpublished, except electronically.

Yviquel:2018:CPU

[YCA18] Hervé Yviquel, Lauro Cruz, and Guido Araujo. Cluster programming using the OpenMP accelerator model. ACM Transactions on Architecture and Code Optimization, 15(3):35:1–35:??, October 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3226112.

Yang:2014:HPD

[YCL14] Luobin Yang, Steve C. Chiu, and Wei-Keng Liao. High performance data clustering: a comparative analysis of performance for GPU, RASC, MPI, and OpenMP implementations. The Journal of Supercomputing, 70(1):284–300, October 2014. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s11227-013-0906-y.

Yu:2013:AGA

[YEG^{+}13] Zhibin Yu, Lieven Eeckhout, Nilanjan Goswami, Tao Li, Lizy John, Hai Jin, and Chengzhong Xu. Accelerating GPGPU architecture simulation. ACM SIGMETRICS Performance Evaluation Review, 41(1): 331–332, June 2013. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yoon:1996:WBP

[YG96] D.-K. Yoon and J.-L. Gaudiot. Worker-based parallel computing on PVM. Lecture Notes in Computer Science, 1123:506–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:2014:IMP

[YGH^{+}14] Xu Yang, Deyuan Guo, Hu He, Haijing Tang, and Yanjun Zhang. An implementation of Message-Passing Interface over Vx-Works for real-time embedded multi-core systems. The Computer Journal, 57(11): 1756–1764, November 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/11/1756.

Yetongnon:1996:PII

[YH96] K. Yetongnon and S. Hariri, editors. Proceedings of the ISCA International Conference. Parallel and Distributed Computing Systems: Dijon, France, 25– 27 September 1996 (PDCS ’96: 9th). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN ???? LCCN ????

Yero:2001:JOO

[YHGL01] Eduardo J. H. Yero, Marco A. A. Henriques, Javier R. Garcïa, and Alina C. Leyva. JOINT: An object oriented message passing interface for parallel programming in Java. Lecture Notes in Computer Science, 2110:637–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer-ny.com/
                                        link/service/series/0558/
                                        bibs/2110/21100637.htm;
                                         http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2110/21100637.
                                        pdf.

Yang:2011:HCO

[YHL11] Chao-Tung Yang, Chih-Lin Huang, and Cheng-Fang Lin. Hybrid CUDA, OpenMP, and MPI parallel programming on multicore GPU clusters. Computer Physics Communications, 182(1): 266–269, January 2011. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465510002262.

Yuasa:1996:RPG

[YKI^{+}96] F. Yuasa, S. Kawabata, T. Ishikawa, D. Perret-Gallix, and T. Kaneko. Running PVM-GRACE on workstation clusters. Lecture Notes in Computer Science, 1156:335–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

YarKhan:2017:PPN

[YKLD17] Asim YarKhan, Jakub Kurzak, Piotr Luszczek, and Jack Dongarra. Porting the PLASMA numerical library to the OpenMP standard. International Journal of Parallel Programming, 45(3):612–633, June 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Yamazaki:2018:SIL

[YKW^{+}18] Ichitaro Yamazaki, Jakub Kurzak, Panruo Wu, Mawussi Zounon, and Jack Dongarra. Symmetric indefinite linear solver using OpenMP task on multicore architectures. IEEE Transactions on Parallel and Distributed Systems, 29(8): 1879–1892, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/08/08301559-
abs.html.

Yang:2009:DBM

[YL09] Chao-Tung Yang and Kuan-Chou Lai. A directivebased MPI code generator for Linux PC clusters. The Journal of Supercomputing, 50(2):177– 207, November 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
50&issue=2&spage=177.

Yang:2016:HTM

[YLC16] Fan Yang, Jinfeng Li, and James Cheng. Husky: towards a more efficient and expressive distributed computing framework. Proceedings of the VLDB Endowment, 9(5):420–431, January 2016. CODEN ???? ISSN 2150-8097.

Yan:2013:SFS

[YLZ13] Shengen Yan, Guoping Long, and Yunquan Zhang. StreamScan: fast scan algorithms for GPUs without global barrier synchronization. ACM SIGPLAN Notices, 48(8):229–238, August 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’13 Conference proceedings.

Yalamov:1997:BRT

[YM97] Plamen Y. Yalamov and Svetozar Margenov. Book reviews: Two books on MPI: Parallel Programming with MPI; MPI: The Complete Reference (2nd printing). IEEE Concurrency, 5(4): 81, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

                                        //dlib.computer.org/pd/
                                        books/pd1997/pdf/p4080.
                                        pdf.

Yilmaz:2011:RMS

[YMYI11] Erdal Yilmaz, Eray Molla, Cansin Yildiz, and Veysi Isler. Realistic modeling of spectator behavior for soccer videogames with CUDA. Computers and Graphics, 35(6):1063– 1069, December 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001476.

Yang:2021:SSG

[YNJS21] Lishan Yang, Bin Nie, Adwait Jog, and Evgenia Smirni. SUGAR: Speeding up GPGPU application resilience estimation with input sizing. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(1):01:1– 01:29, February 2021. CO-DEN ???? ISSN 2476-1249.

URL https://dl.acm.org/
doi/10.1145/3447375.

Yi:1994:PID

[YPA94] Sung Yi, K. H. Pierson, and M. F. Ahmad. Parallel implementation of dynamic simulation to filamentary composite structures with general rate dependent damping. Computing systems in engineering: an international journal, 5(4-6): 469–477, August–December 1994. CODEN COSEEO. ISSN 0956-0521.

Yilmaz:2009:HPC

[YPAE09] E. Yilmaz, R. U. Payli, H. U.

Akay, and A. Ecer. Hybrid parallelism for CFD simulations: Combining MPI with OpenMP. In Tuncer et al.

[TGEM09], pages 401–408. CODEN LNCSA6. ISBN 3-540-92743-3 (print), 3-540-92744-1 (e-book). ISSN 1439-7358. LCCN ???? URL

                                        http://link.springer.com/
                                        content/pdf/10.1007/978-
                                        3-540-92744-0_50. Parallel

CFD 2007 was held in Antalya, Turkey, from May 21 to 24, 2007.

You:1995:EIM

[YPZC95] J. You, E. Pissaloux, W. P. Zhu, and H. A. Cohen. Efficient image matching: a hierarchical Chamfer matching scheme via distributed system. Real-Time Imaging, 1 (4):245–259, October 1995. CODEN REIMFQ. ISSN 1077-2014.

Young:1993:PEN

[YS93] Y.-H. Young and K. Sikorski. Performance evaluation of network programming environments. In Mudge et al. [MMH93], pages 106–107 (vol. 2). ISBN 0-8186-3230-5. LCCN ???? Four volumes. IEEE catalog number 93TH0501-7.

Yuan:2012:PCS

[YSL^{+}12] Zhiyong Yuan, Weixin Si, Xiangyun Liao, Zhaoliang Duan, Yihua Ding, and Jianhui Zhao. Parallel computing of 3D smoking simulation based on OpenCL heterogeneous platform. The Journal of Supercomputing, 61(1):84–102, July 2012.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=
61&issue=1&spage=84.

Young-S:2017:OGI

[YSMA^{+}17] Luis E. Young-S., Paulsamy Muruganandam, Sadhan K. Adhikari, Vladimir Loncar, Dusan Vudragović, and Antun Balaz. OpenMP GNU and Intel Fortran programs for solving the timedependent Gross–Pitaevskii equation. Computer Physics Communications, 220(??): 503–506, November 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465517302321.

Yu:2005:HPB

[YSP^{+}05] Weikuan Yu, Sayantan Sur, Dhabaleswar K. Panda, Rob T. Aulwes, and Rich L. Graham. High performance broadcast support in LA-MPI over quadrics. The International Journal of High Performance Computing Applications, 19(4):453– 463, Winter 2005. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/19/
4/453.full.pdf+html.

Yeh:2017:PFG

[YSS^{+}17] Tsung Tai Yeh, Amit Sabne, Putt Sakdhnagool, Rudolf Eigenmann, and Timothy G. Rogers. Pagoda: Finegrained GPU resource virtualization for narrow tasks. ACM SIGPLAN Notices, 52 (8):221–234, August 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yeh:2019:PGR

[YSS^{+}19] Tsung Tai Yeh, Amit Sabne, Putt Sakdhnagool, Rudolf Eigenmann, and Timothy G. Rogers. Pagoda: a GPU runtime system for narrow tasks. ACM Transactions on Parallel Computing (TOPC), 6(4):21:1– 21:??, November 2019. CO-DEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Yang:2008:DPL

[YST08] Chao-Tung Yang, Wen-Chung Shih, and Shian-Shyong Tseng. Dynamic partitioning of loop iterations on heterogeneous PC clusters. The Journal of Supercomputing, 44 (1):1–23, April 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&

issn=0920-8542&volume=
44&issue=1&spage=1.

Young-S:2016:OFP

[YSVM^{+}16] Luis E. Young-S., Dusan Vudragović, Paulsamy Muruganandam, Sadhan K. Adhikari, and Antun Balaz. OpenMP Fortran and C programs for solving the timedependent Gross–Pitaevskii equation in an anisotropic trap. Computer Physics Communications, 204(??): 209–213, July 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S001046551630073X.

Yan:2014:OMB

[YSWY14] Xin Yan, Xiaohua Shi, Lina Wang, and Haiyan Yang. An OpenCL micro-benchmark suite for GPUs and CPUs. The Journal of Supercomputing, 69(2):693–713, August 2014. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11227-014-1112-2.

Yu:2020:EPW

[YT20] C. Yu and S. Tsao. Efficient and portable workgroup size tuning. IEEE Transactions on Parallel and Distributed Systems, 31(2): 455–469, February 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yoshinaga:2012:DBM

[YTH^{+}12] Kazumi Yoshinaga, Yuichi Tsujita, Atsushi Hori, Mikiko Sato, and Mitaro Namiki. Delegation-based MPI communications for a hybrid parallel computer with many-core architecture. Lecture Notes in Computer Science, 7490:47–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer.com/chapter/
                                        10.1007/978-3-642-33518-
                                        1_10/.

Yam-Uicab:2017:FHT

[YULMTS^{+}17] R. Yam-Uicab, J. L. Lopez-Martinez, J. A. Trejo-Sanchez, H. Hidalgo-Silva, and S. Gonzalez-Segura. A fast Hough transform algorithm for straight lines detection in an image using GPU parallel computing with CUDA-C. The Journal of Supercomputing, 73(11): 4823–4842, November 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yang:2021:HMC

[YW21] Sheng-Chun Yang and Yong-Lei Wang. A hybrid MPI-CUDA approach for nonequispaced discrete Fourier transformation. Computer Physics Communications, 258(??):Article 107513, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465520302393.

Yang:2011:PBP

[YWC11] Chao-Tung Yang, Chao-Chin Wu, and Jen-Hsiang Chang. Performance-based parallel loop self-scheduling using hybrid OpenMP and MPI programming on multicore SMP clusters. Concurrency and Computation: Practice and Experience, 23 (8):721–744, June 10, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Younge:2015:SHP

[YWCF15] Andrew J. Younge, John Paul Walters, Stephen P. Crago, and Geoffrey C. Fox. Supporting high performance molecular dynamics in virtualized clusters using IOMMU, SR-IOV, and GPUDirect. ACM SIGPLAN Notices, 50 (7):31–38, July 2015. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yonezawa:1995:IED

[YWO95] Naoki Yonezawa, Koichi Wada, and Motoko Obata. Implementation and evaluation of distributed shared data objects on a workstation cluster. In IEEE [IEE95e], pages 319–322. ISBN 0-7803-2553-2. LCCN TK 5101 A1 I34 1995. IEEE catalog number 95CH35765.

You:2015:VFO

[YWTC15] Yi-Ping You, Hen-Jung Wu, Yeh-Ning Tsai, and Yen-Ting Chao. VirtCL: a framework for OpenCL device abstraction and management. ACM SIGPLAN Notices, 50(8):161–172, August 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yong:1995:SOM

[YX95] Dou Yong and Zhou Xingming. Super-object model: implementing shared memory programming mode on distributed memory multicomputers. Chinese Journal of Computers, 18(7): 481–487, July 1995. CODEN JIXUDT. ISSN 0254-4164.

Yu:2012:SCC

[YYW^{+}12] Fang Yu, Shun-Ching Yang, Farn Wang, Guan-Cheng Chen, and Che-Chang Chan. Symbolic consistency checking of OpenMP parallel programs. ACM SIGPLAN Notices, 47(5):139–148, May 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCTES ’12 proceedings.

Yang:2014:CNR

[YZ14] Yi Yang and Huiyang Zhou.

CUDA-NP: realizing nested thread-level parallelism in GPGPU applications. ACM SIGPLAN Notices, 49(8): 93–106, August 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

You:1995:PIM

[YZPC95] J. You, W. P. Zhu, E. Pissaloux, and H. A. Cohen. Parallel image matching on a distributed system. In Narashimhan [Nar95], pages 870–873 (vol. 2). ISBN 0-7803-2018-2 (paperback), 0-7803-2019-0 (microfiche). LCCN QA76.6.I15 1995. Two volumes. IEEE catalog no. 95TH0682-5.

Zounmevo:2014:FRC

[ZA14] Judicael A. Zounmevo and Ahmad Afsahi. A fast and resource-conscious MPI message queue mechanism for large-scale jobs. Future Generation Computer Systems, 30(??):265–290, January 2014. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X13001489.

Zaza:2016:CBP

[ZAFAM16] Ayham Zaza, Abeeb A.

Awotunde, Faisal A. Fairag, and Mayez A. Al-Mouhamed. A CUDA based parallel multi-phase oil reservoir simulator. Computer Physics Communications, 206(??):2– 16, September 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465516300996.

Zahavi:2012:FTR

[Zah12] Eitan Zahavi. Fat-tree routing and node ordering providing contention free traffic for MPI global collectives. Journal of Parallel and Distributed Computing, 72(11): 1423–1432, November 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731512000305.

Zhong:2007:PPS

[ZAT^{+}07] Wei Zhong, Gulsah Altun, Xinmin Tian, Robert Harrison, Phang C. Tai, and Yi Pan. Parallel protein secondary structure prediction schemes using Pthread and OpenMP over hyperthreading technology. The Journal of Supercomputing, 41(1):1–16, July 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&

issn=0920-8542&volume=
41&issue=1&spage=1.

Zdetsis:1994:PMD

[ZB94] A. D. Zdetsis and R. Biswas.

A parallel molecular dynamics strategy for PVM. In Turchi and Gonis [TG94], pages 713–718. ISBN 0-306-44626-X. ISSN 0258-1221. LCCN TN690.S77 1994.

Zilli:1997:TBN

[ZB97] G. Zilli and L. Bergamaschi. Truncated block Newton and quasi-Newton methods for sparse systems of nonlinear equations. experiments on parallel platforms. Lecture Notes in Computer Science, 1332:390–400, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhu:2012:CDS

[ZBd12] Ke Zhu, Matthias Butenuth, and Pablo d’Angelo. Comparison of dense stereo using CUDA. Lecture Notes in Computer Science, 6554: 398–410, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/content/pdf/
10.1007/978-3-642-35740-
4_31.

Zhao:2010:GMP

[ZC10] Kaiyong Zhao and Xiaowen Chu. GPUMP: a multipleprecision integer library for GPUs. In IEEE, editor, IEEE 10th International Conference on Computer and Information Technology (CIT), 2010: June 29, 2010–July 1, 2010, Bradford, West Yorkshire, UK, pages 1164–1168. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 0-7695-4108-9 (print), 1-4244-7547-3. LCCN ???? IEEE Computer Society Order Number E4108. BMS Part Number: CFP10355-CDR.

Zhong:2022:ULV

[ZCBD22] Dong Zhong, Qinglei Cao, George Bosilca, and Jack Dongarra. Using long vector extensions for MPI reductions. Parallel Computing, 109(??):??, March 2022. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819121001137.

Zhang:1997:DED

[ZDD97] Xiaodong Zhang, Sandra G. Dykes, and Hong Deng. Distributed edge detection: Issues and implementations. IEEE Computational Science & Engineering, 4(1):72–82, January/ March 1997. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (elec-

tronic). URL http://dlib.

computer.org/cs/books/
cs1997/pdf/c1072.pdf;
http://www.computer.org/
cse/cs1998/c1072abs.htm.

Zhang:2001:PPV

[ZDR01] Xin Zhang, Lingli Ding, and Elke A. Rundensteiner. PVM: Parallel View Maintenance under concurrent data updates of distributed sources. Lecture Notes in Computer Science, 2114: 230–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/
link/service/series/0558/
bibs/2114/21140230.htm;
 http://link.springer-
ny.com/link/service/series/
0558/papers/2114/21140230.
pdf.

Zhang:2004:PMV

[ZDR04] Xin Zhang, Lingli Ding, and Elke A. Rundensteiner. Parallel multisource view maintenance. VLDB Journal: Very Large Data Bases, 13 (1):22–48, January 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zelek:1995:DPP

[Zel95]  J. S. Zelek.  Dynamic

path planning. In IEEE [IEE95a], pages 1285–1290 (vol. 2). ISBN 0-7803-2559-1. LCCN TA168.I19 1995.

Five volumes. IEEE catalog no. 95CH3576-7.

Zemla:1994:WTC

[Zem94] A. Zemla. Wavelet transforms computing on PVM. In Dongarra and Wasniewski [DW94], pages 534– 546. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1994. DM104.00.

Zhou:1995:FMP

[ZG95a] H. Zhou and A. Geist. Faster message passing in PVM. In Alnuweiri and Hamdi [AH95], pages 67–73. ISBN 0-8186-7124-6. LCCN TK5105.5 .H56 1995.

Zhou:1995:RMR

[ZG95b] Honbo Zhou and Al Geist. “receiver makes right” data conversion in PVM. In IEEE [IEE95b], pages 458– 464. ISBN 0-7803-2493-5, 0-7803-2492-7, 0-7803-2494-3. LCCN TK7885.A1 I567 1995. IEEE catalog no. 95CH35751.

Zhou:1996:FMP

[ZG96] Honbo Zhou and Al Geist. Faster message passing in PVM. Technical report, Mathematical Sciences Section, Oak Ridge National Laboratory, Knoxville, TN, USA, 1996. 7 pp. URL http:

                                        //www.epm.ornl.gov/~zhou/
                                        patm.ps.

Zhou:1998:LST

[ZG98] Honbo Zhou and Al Geist.

LPVM: a step towards multithread PVM. Concurrency: practice and experience, 10(5):407–416, April 25, 1998. CODEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract?ID=5385;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5385&PLACEBO=IE.pdf.

Zielinski:1994:PPS

[ZGC94] K. Zielinski, M. Gajecki, and G. Czajkowski. Parallel programming systems for LAN distributed computing. In IEEE [IEE94b], pages 600–607. ISBN 0-8186-6952-7 (casebound), 0-8186-6950-0 (paperback), 0-8186-6951-9 (microfiche). LCCN TA1637.I25 1994. Three volumes. IEEE catalog no. 94CH35708.

Zu:1994:OSM

[ZGN94] Hong Zu, Ya-Dong Gui, and L. M. Ni. Optimal software multicast in wormhole-routed multistage networks. In IEEE [IEE94h], pages 703–712. ISBN 0-8186-6607-2, 0-8186-6605-6, 0-8186-6606-4. ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Zhao:2022:SGM

[ZGNZ22] Chen Zhao, Wu Gao, Feiping Nie, and Huiyang Zhou. A survey of GPU multitasking methods supported by hardware architecture. IEEE Transactions on Parallel and Distributed Systems, 33 (6):1451–1463, June 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2020:CHM

[ZGZS20] Huan Zhou, José Gracia, Naweiluo Zhou, and Ralf Schneider. Collectives in hybrid MPI+MPI code: Design, practice and performance. Parallel Computing, 99(??):Article 102669, November 2020. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167819120300612.

Zheng:2006:PEA

[ZHK06] Gengbin Zheng, Chao Huang, and Laxmikant V. Kalé. Performance evaluation of automatic checkpoint-based fault tolerance for AMPI and Charm++. Operating Systems Review, 40(2):90– 99, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhou:2021:HPG

[Zho21] Chao Zhou. High performance graph data imputation on multiple GPUs. Future Internet, 13(2):36, January 31, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/2/36.

Zoraja:1999:SPD

[ZHS99] Ivan Zoraja, Hermann Hellwagner, and Vaidy Sunderam. SCIPVM: Parallel distributed computing on SCI workstation clusters. Concurrency: practice and experience, 11(3): 121–138, March 1999. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract?ID=61003667;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=61003667&PLACEBO=IE.
pdf.

Zhang:2018:IRP

[ZJDW18] Xuechen Zhang, Song Jiang, Alseny Diallo, and Lei Wang. IR+: Removing parallel I/O interference of MPI programs via data replication over heterogeneous storage devices. Parallel Computing, 76(??):91– 105, August 2018. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167819118300140.

Zarebavani:2020:CCB

[ZJHS20] B. Zarebavani, F. Jafarinejad, M. Hashemi, and S. Salehkaleybar. cuPC: CUDA-based parallel PC algorithm for causal structure learning on GPU. IEEE Transactions on Parallel and Distributed Systems, 31 (3):530–542, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zounmevo:2014:ESC

[ZKRA14] Judicael A. Zounmevo, Dries Kimpe, Robert Ross, and Ahmad Afsahi. Extremescale computing services over MPI: Experiences, observations and features proposal for next-generation message passing interface. The International Journal of High Performance Computing Applications, 28(4): 435–449, November 2014. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

http://hpc.sagepub.com/
content/28/4/435.

Zaky:1996:PDT

[ZL96] Amr Zaky and Ted Lewis, editors. Program development tools and environments for parallel and distributed systems: Session; 28th Hawaii international conference on system sciences — 1995, volume 2 of Kluwer International Series in Software Engineering. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-9675-8. LCCN QA76.58.T65 1996.

Zha:2017:IFM

[ZL17] Yue Zha and Jing Li.

IMEC: A fully morphable in-memory computing fabric enabled by resistive crossbar. IEEE Computer Architecture Letters, 16(2):123–126, July/December 2017. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zha:2018:LSM

[ZL18] Yue Zha and Jing Li. Liquid Silicon-Monona: a reconfigurable memory-oriented computing fabric with scalable multi-context support. ACM SIGPLAN Notices, 53 (2):214–228, February 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zaki:1999:TSP

[ZLGS99] Omer Zaki, Ewing Lusk, William Gropp, and Deborah Swider. Toward scalable performance visualization with Jumpshot. The International Journal of High Performance Computing Applications, 13(3):277– 288, Fall 1999. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Zhou:2012:DFD

[ZLL^{+}12] Xu Zhou, Kai Lu, Xicheng Lu, Xiaoping Wang, and Baohua Fan. dMPI: Facilitating debugging of MPI programs via deterministic message passing. Lecture Notes in Computer Science, 7513: 172–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/10.
                                        1007/978-3-642-35606-3_
                                        20/.

Zhang:2017:DLN

[ZLP17] Jie Zhang, Xiaoyi Lu, and Dhabaleswar K. (DK) Panda. Designing locality and NUMA aware MPI runtime for nested virtualization based HPC cloud with SR–IOV enabled InfiniBand. ACM SIGPLAN Notices, 52(7):187–200, July 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhu:2015:PIM

[ZLS^{+}15] Xiangyuan Zhu, Kenli Li, Ahmad Salah, Lin Shi, and Keqin Li. Parallel implementation of MAFFT on CUDAenabled graphics hardware. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 12(1):205– 218, January 2015. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Zhang:2020:CTE

[ZLWW20] T. Zhang, X. Liu, X. Wang, and A. Walid. cuTensor-Tubal: Efficient primitives for tubal-rank tensor learning operations on GPUs. IEEE Transactions on Parallel and Distributed Systems, 31(3):595–610, March 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhai:2011:CVH

[ZLZ^{+}11] Yan Zhai, Mingliang Liu, Jidong Zhai, Xiaosong Ma, and Wenguang Chen. Cloud versus in-house cluster: evaluating Amazon cluster compute instances for running MPI applications. In ACM [ACM11], pages 11:1–11:10. ISBN 1-4503-1139-3. LCCN ????

Zollweg:1993:OP

[Zol93] J. A. Zollweg. Overview of PVM. In Anonymous [Ano93f], pages 981–986. ISBN ???? ISSN 0254-6213. LCCN ????

Zarrelli:2006:EPE

[ZPI06] Roberto Zarrelli, Mario Petrone, and Angelo Iannaccio. Enabling PVM to exploit the SCTP protocol. Journal of Parallel and Distributed Computing, 66(11): 1472–1479, November 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zambonelli:1996:EPP

[ZPLS96] F. Zambonelli, M. Pugassi, L. Leonardi, and N. Scarabottolo. Experiences on porting a Parallel Objects environment from a transputer network to a PVM-based system. In IEEE [IEE96g]. ISBN 0-8186-7376-1. LCCN QA76.58 .E97 1996. IEEE order number PR07376.

Zheng:2011:GLO

[ZRQA11] Mai Zheng, Vignesh T. Ravi, Feng Qin, and Gagan Agrawal. GRace: a lowoverhead mechanism for detecting data races in GPU programs. ACM SIG-PLAN Notices, 46(8):135– 146, August 2011. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’11 Conference proceedings.

Zhao:2012:ASO

[ZSG12] Xin Zhao, Gopalakrishnan Santhanaraman, and William Gropp. Adaptive strategy for one-sided communication in MPICH2. Lecture Notes in Computer Science, 7490:16–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/
10.1007/978-3-642-33518-
1_7/.

Zarrabi:2015:GSA

[ZSK15] Amirreza Zarrabi, Khairulmizam Samsudin, and Ettikan K. Karuppiah. Gravitational search algorithm using CUDA: a case study in high-performance metaheuristics. The Journal of Supercomputing, 71(4):1277– 1296, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11227-014-1360-1.

Zoltani:2001:EPO

[ZSnH01] Csaba K. Zoltani, Punyam Satya-narayana, and Dixie Hisley. Evaluating performance of OpenMP and MPI on the SGI Origin 2000 with benchmarks of realistic problem sizes. Parallel and Distributed Computing Practices, 4(4):??, December 2001. CODEN ???? ISSN 1097-2803.

Zouaoui:2017:CNG

[ZT17] Chakib Mustapha Anouar Zouaoui and Nasreddine Taleb. CL ARRAY: a new generic library of multidimensional containers for C++ compilers with extension for OpenCL framework. Computer Languages, Systems and Structures, 50 (??):53–81, December 2017. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S147784241630135X.

Zhou:2020:EOP

[ZT20] Hongyang Zhou and Gábor Tóth. Efficient OpenMP parallelization to a complex MPI parallel magnetohydrodynamics code. Journal of Parallel and Distributed Computing, 139(??): 65–74, May 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0743731519304903.

Zaitsev:2019:SLD

[ZTD19] D. Zaitsev, S. Tomov, and J. Dongarra. Solving linear Diophantine systems on parallel architectures. IEEE Transactions on Parallel and Distributed Systems, 30 (5):1158–1169, May 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2021:POT

[ZWC21] Zijie Zhu, Yongxian Wang, and Xinghua Cheng. Parallel optimization of threedimensional wedge-shaped underwater acoustic propagation based on MPI + OpenMP hybrid programming model. The Journal of Supercomputing, 77 (5):4988–5018, May 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

https://link.springer.
com/article/10.1007/s11227-
020-03466-w.

Zareski:1995:EPG

[ZWHS95] D. Zareski, B. Wade, P. Hubbard, and P. Shirley. Efficient parallel global illumination using density estimation. In Uselton et al. [UCW95], pages 47– 54, 104–105. ISBN 0-89791-774-1 (softbound) [invalid checksum], 0-7803-3120-6 (microfiche). LCCN QA76.58.P3778 1995. ACM order number 428957. IEEE Computer Society Press order number 95TB8134.

Zheng:2005:SBP

[ZWJK05] Gengbin Zheng, Terry Wilmarth, Praveen Jagadishprasad, and Laxmikant V. Kalé. Simulation-based performance prediction for large parallel machines. International Journal of Parallel Programming, 33(2–3): 183–207, June 2005. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&

                                        issn=0885-7458&volume=
                                        33&issue=2&spage=183.

Zhang:2013:MPI

[ZWL13] Xiaohua Zhang, Sergio E. Wong, and Felice C. Lightstone. Message passing interface and multithreading hybrid for parallel molecular docking of large databases on petascale high performance computing machines. Journal of Computational Chemistry, 34(11):915–927, April 30, 2013. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Zhu:2017:OAP

[ZWL^{+}17] Huming Zhu, Yanfei Wu, Pei Li, Peng Zhang, Zhe Ji, and Maoguo Gong. An OpenCLaccelerated parallel immunodominance clone selection algorithm for feature selection. Concurrency and Computation: Practice and Experience, 29(9), May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:IRP

[ZWLZ21] Jingrong Zhang, Zihao Wang, Zhiyong Liu, and Fa Zhang. Improve the resolution and parallel performance of the threedimensional refine algorithm in RELION using CUDA and MPI. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 18(2):583–595, March 2021. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic). URL https://

dl.acm.org/doi/10.1109/
TCBB.2019.2929171.

Zhu:1995:RTC

[ZWZ^{+}95] Miaoliang Zhu, Chunming Wu, Youjun Zhang, Yi Jin, and Jie Li. A realtime and concurrent intelligent robotic system based on multi-agent architecture. High Technology Letters, 5 (10):20–24, October 1995. CODEN GTONE8. ISSN 1002-0470.

Zhang:2005:ULC

[ZWZ05] Youhui Zhang, Dongsheng Wong, and Weimin Zheng. User-level checkpoint and recovery for LAM/MPI. Operating Systems Review, 39(3): 72–81, July 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhuang:1995:PRS

[ZZ95] Xinglai Zhuang and Jianping Zhu. Parallelizing a reservoir simulator using MPI. In IEEE [IEE95j], pages 165– 174. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Zeyao:2004:AMI

[ZZ04] Mo Zeyao and Huang Zhengfeng. Application of MPI-IO in parallel particle transport Monte–Carlo simulation. Parallel Algorithms and Applications, 19 (4):227–236, ???? 2004. CO-DEN PAAPEC. ISSN 1063-

7192. URL http://www.
informaworld.com/smpp/
content~content=a714592658.

Zheng:2014:IMS

[ZZG^{+}14] Liang Zheng, Huai Zhang, Taras Gerya, Matthew Knepley, David A. Yuen, and Yaolin Shi. Implementation of a multigrid solver on a GPU for Stokes equations with strongly variable viscosity based on Matlab and CUDA. The International Journal of High Performance Computing Applications, 28(1):50–60, February 2014. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/28/
1/50.full.pdf+html.

Zhu:2015:PML

[ZZZ^{+}15] Leqing Zhu, Yadong Zhou, Daxing Zhang, Dadong Wang, Huiyan Wang, and Xun Wang. Parallel multilevel 2D-DWT on CUDA GPUs and its application in ring artifact removal. Concurrency and Computation: Practice and Experience, 27(17):5188–5202, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 REFERENCES 551

(print), 1532-0634 (electronic).